A Bibliography of Publications about Virtual Machines

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 October 2025 Version 1.452

Abstract This bibliography records books and other publications about virtual

machines.

Title word cross-reference

$32.95 [Ano97a]. 5 [ALW15, HH18]. ^{T M} [Cza00]. _{T P} [LTK17]. d [XDLS15]. HV^{2}M [CBZ^{+}16]. n [WZKP19]. ω [Arv02]. Π [Syr07]. V^{2} [DG05].

-dienste [WF03]. -Enabled [SB18]. -Tier [WZKP19].

.NET [Fra06, Fra09, Hee07, Hog06, Hog08, Men03].

/CLI [Fra06, Fra09, Hee07, Hog06, Hog08, Siv07, Wil06]. /dev/random [Fer11].

0 [Sim92, SCP93]. 0.9.0 [WR07]. 0.9.1 [WR08]. ’01 [Ano00, Ano01a, Ano01b, USE01c, USE01d]. ’02 [USE02]. ’03 [ACM03b, Ert03]. ’04 [Ano04a, Ano04b]. ’05 [ACM05d, Vra05].

1 [Fli77, Pul91, Sch94a, WDSW01]. 1-4 [Ano06a]. 1.x [KGG00]. ’10 [Ano10, Gal73, See10, VSC^{+}10, YCL^{+}18]. 10.0 [Bau06b]. 10GE [HB12]. 11 [Ham76, PK75a]. 11/40 [GBO87]. 11/60 [SP83]. 1100 [Kam75]. 11th [ACM04a]. 12th [IEE85]. 14-16 [ACM06f]. 148 [Ano15]. 15th [ACM06b]. 165 [IBM76a]. 168 [IBM76a]. 16th [BW03, Ano93]. 17th [Ano10]. 180x86 [Law00]. 19 [Cap21]. 1981 [Mar81]. 1985 [Jou85]. 1989 [Boa90, Shr89]. 1990 [ACM90]. 1991 [MR91]. 1992 [IEE92]. 1993 [IEE93a, IEE93b, JPTE94]. 19th [ACM03b, SS05]. 1st [ACM06c, Ano01a].

2 [Bri98, Com00, Com03, Kis08]. 2-Level [ZSR^{+}05]. 2.0 [Fra06, Ng01a, SUN97]. 2000 [ACM00]. 2001 [ACM01b]. 2003 [RM03, ACM03a, ACM03b, IEE03, Int05a]. 2004 [ACM04a, ACM04b]. 2005 [ACM05a, ACM05b, ACM05c, Wil06]. 2006 [ACM06c, ACM06b, ACM06d, IEE06b, IEE06a, Int06b, Int06c, Int06a]. 2008 [Lar09]. 2010 [Ano10]. 2011 [LCK11]. 2018 [Kol19]. 20th [IEE06a, Vra05]. 21st [IEE05]. 23272 [Int05b]. 26th [ACM99]. 29-state [Sig89]. 2nd [Ano02].

3 [McC08, PO09, vdK09]. 3.0 [MRGB91]. 3.1 [Bau06a, Skr01]. 3.5 [Fra09, Hog08]. 32 [Ano14b]. 32-bit [VED06]. 335 [ECM01, ECM02, ECM05, ECM06]. 360 [Kam75]. 360/40 [ABCC66]. 370 [Att79, Bar73, Bar78, Ber86, Cal75, Com82, GLC84, Gum83, IBM72, IBM73, IBM76a, IBM76b, McC74, Olb78, SM79]. 37th [ACM06d]. 390 [DBC^{+}00]. 3rd [ACM05b, ACM06e, Ano04a].

4 [Gal09b, G^{+}06, Lav10, Low09, NOK^{+}85]. 4-7 [M^{+}06]. 40 [Com82, GBO87]. 43rd [ACM06a]. 440 [R^{+}02]. 4th [USE00a].

5 [IEE02, War05]. 5.2 [McK04, P^{+}08]. 5.5 [Bau06c, LMG^{+}14]. 5G [CM18, HCB18, MVMHL24, RNA^{+}22, SP22, XWW^{+}21, ZLZ21a]. 5L [Mly09].

60 [SP83]. 6000 [ABDD^{+}91]. 64 [De 06, Don06]. 64-bit [VED06, VED07]. 67 [Bar73, Par72]. 6G [AAT^{+}22, PF23, RNS^{+}23, SCC^{+}23]. 6G-based [AAT^{+}22]. 6th [USE01b].

7 [HH08]. 7th [Tho93].

8 [LYBB14, She02]. 80 [BMWB86, BSUH87]. 84 [IT86]. 84/K [IT86]. ’89 [ACM89].

’90 [IEE90b]. 9000 [ADG^{+}92]. 91 [MR91]. ’92 [IEE92]. ’93 [GHH^{+}93, IEE93b, LFBB94]. ’96 [ACM96]. ’99 [ACM99, USE99].

A-DRM [WIS^{+}15]. A.NET [Men03]. Aachen [GHH^{+}93]. AADEBUG [ACM05a]. ABC [KSS^{+}20, KSS^{+}23, JFPL16, JFZL17]. ABC-BA [KSS^{+}20, KSS^{+}23]. Ability [RI00]. Abnormal [BSOK^{+}20, ZZW^{+}24]. above [LSPP^{+}23]. Absorbing [MTFK19, RHR20]. Abstract [Wel94, KMMV14, CK87]. Abstraction [YLH17, Sch09, YWTC15]. Abstractions [BJH^{+}16, MD12, Tsa14, UR15]. Abundant [ZCJ^{+}21]. Accelerated [MTFK19, NRdA^{+}20, SCSL12, SWF16, BHDS09, SPAK18]. Accelerates [Ano03a]. Accelerating [BSSM08, LSX^{+}24, MNS^{+}14, ZLBF14, KKC^{+}16]. Acceleration [DEK^{+}03, FLZ^{+}20, NTL^{+}24, WZL^{+}23, YML^{+}18, PRS16, PRS26, Wu13, XZZ^{+}16]. Accelerator [LWB13, RSV24, SXMX^{+}18, CLM24, GGK19, GKT17, JAC^{+}19, LKY^{+}17]. Accelerators [GLL^{+}21]. Access [DL19b, Fie68, KCWH14, KP15, LSX^{+}24, LZW^{+}17, MK13, Mar73, MMdMCPSR25, MSI18, SLZ^{+}25, WLCS17, WWL^{+}17b, YI24, Bor07, BTLNBF^{+}15a, CWC^{+}14, CFS^{+}12, HCB18, MN91, Oi08, PSC^{+}07, XJW^{+}18]. Accesses [WVT^{+}17]. account [Yel99]. accountability [HBP06]. Accounting [JSHM15, CMP^{+}13, HB08]. Acculock [XXZ13]. accuracy [CEG07, EG03]. Accurate [HLW^{+}23, RRB17, ZDLG17, SK13b, XXZ13]. ACDC [AHK^{+}15]. ACDC-JS [AHK^{+}15]. achieve [ZL13]. Achieving [HLPY16, KMK16, KLR^{+}20, MBK^{+}92, ZXL^{+}24, JYOB18]. Acid [LPSS19]. ACLE [PBR^{+}90]. ACM [ACM90, ACM01b, RM03, ACM03b, ACM04a, ACM04b, ACM05c, ACM05d, ACM06a, IEE04, Vra05, Cre10b]. ACM/IEEE [ACM05c]. ACM/USENIX [ACM05d]. ACO [KS20a]. ACO-based [KS20a]. ACOS [NOK^{+}85]. ACOS-4 [NOK^{+}85]. Acquiring [FG91]. Across [JWL^{+}18, TMMVL12, XLQL18]. ACStor [WWL^{+}17b]. action [Hol95, KB17, Siv07]. Activation [RSN^{+}18]. Actor [TCP^{+}17]. actors [GE85, Sub11]. Actra [TLD^{+}89]. Ada [Bak83, Dom80a, GBO87, GR80, Ibs84a, Kam83, Ker88, Ibs84b, SGS92, SM92, Vol90, WSX^{+}19, Wes98, ZEdlP13]. Ada-Things [WSX^{+}19]. Ada’97 [ACM97]. Adaptability [SV13]. Adaptable [GKSP99, CGM17]. adaptation [AJBJ23, ZBG^{+}05]. AdaptChain [ZHH^{+}24]. adapters [SAB^{+}07]. Adapting [BADM06, SJW^{+}13, WGLL13]. Adaption [BHI15]. Adaptive [AS14, Bir94, CYX^{+}17, DLH^{+}20, GA25, HHW10, HKKW13, JKB15, KHL17, KZ21, LMV12, Lee16, LCT^{+}15, LZL^{+}15, LYGG20, MCJ19, NZH20, OVI^{+}12, PSZ^{+}07, PYDG22, PMP23, SOAK23, SXCL14, XCSM18, XLL^{+}20, YLT^{+}23, ZDK^{+}22, ZLZ21a, ZFH^{+}22, ZHH^{+}24, dSOK17, BB12, BKT^{+}19, GKJ^{+}19, HPS23, JNR12, JC18, KKB14, KR16, RMP24, SENS16, SEPV19, SYMA17, WSX^{+}19, ZLCZ18]. adaptively [JDW^{+}14]. Adding [HHV^{+}02, Say67]. Address [HWHW18]. Addressing [XKKL23]. Administration [MJW^{+}06, NSHW10, Bor07, Dav04]. administrator [TC10]. Admitting [MLXG19]. Advanced [A^{+}04, Alf91, fLtNW14, AAB^{+}05c, B^{+}05, G^{+}05, IEE06a, MRM06, OH05, SXH^{+}19, SS05, Fit14]. Advances [Ert05]. Advancing [GEK25, ZKWH17]. Adversarial [HLP^{+}16, PST15b]. advisor [ZLV^{+}12, ZBS^{+}15]. affect [SMSH18]. Affinity [SK13c]. Affinity-aware [SK13c]. Against [HS19, CVWL13, CD12, GHD12, TMV25]. age [Bac11]. agenda [ZJRW19]. Agent [PXG^{+}17, RB01, ABV12, DHD20, PDM20]. agent-based [PDM20]. Aggregation [JZY^{+}22]. Agile [GHS16, GHS17, IMK^{+}13, ZLZ^{+}21b]. Agility [OSK15]. Aging [AJM^{+}25, AMA^{+}14]. Agnostic [IYAK23]. ahead [MA10]. AHP [AHRR22b]. Aid [KLF^{+}15]. aided [ME87, SI81, TB14]. AINA [SS05]. AIX [Gal11, Mly09]. ALEP [Sim92, SCP93]. ALEP-0 [Sim92, SCP93]. Alfa [WDSW01]. Alfa-1 [WDSW01]. Algorithm [AAR22, AAK18, BP99, LSC^{+}17, LLWM23, LWW16, LW12, LLS^{+}20, LW20, RK24, WCC20, YLQL25, ZHL16, AHRR22a, BB20, BRS^{+}22, EB20, FGZC23, FS19, FF22, GGQ^{+}13, GH20, GPR23, GA18, HAK22, Hog02, HS13, JGA^{+}88, LKR^{+}19, LZC^{+}16, LY23, MNB24, MM92, MS17, MMTM22, MHM19, NGN24, NAR19, PC21, PKS^{+}19, QBL^{+}23, RGAT18, RH17, RT18, SBI21, SEM^{+}20, SS19, SP23, TMLL14, Tho68, WBW^{+}19, XWW^{+}21, XXWG23, YLCH17, YYC^{+}19, YB24, ZYLY18]. Algorithm-Dependent [BP99]. Algorithms [ARAAA19, FGLI15, HHK94, KP99, LPSS19, Man15a, SHW^{+}15, AB16, BB12, CRB12, GLW23, HH19, Man18, ME87, MJ93, SGS92, WHW20, XTB17, YTS14]. aligned [AGIS94]. Alignment [EDS^{+}15]. allocate [LLF^{+}18]. allocating [MMTM22, XHW^{+}19]. Allocation [AIAR^{+}24, BFM^{+}21, CWL12, CPST14, Do11, GLBJ18, HKLM17, HLPY16, JFPL16, KRS^{+}17, KCY22, LLZ18, Man15a, MMdMCPSR25, NMG15, NHL22, PCC^{+}16, SJ21, SDM21, VTW16, WTJR22, XSC13, ZWFX17, ZLG^{+}20, CEPR22, CPST15, dCCDFdO15, DEG^{+}17, EdPG^{+}10, GLLJ16, GLW23, HMH17, HH19, JWH^{+}15, JC18, KDK20, KS20a, KS18a, LJYZ15, Mly09, RNA^{+}22, RCTY19, SGV13, SPSL25, THH^{+}14, WGY20, YGLY21, YYC^{+}19, YIR24, YI24, ZG13, ZLH^{+}15, ZWC^{+}19]. allocation-site-based [CPST15]. Allocations [YWH^{+}21]. allowing [Tho73]. already [RAT17]. Alternative [HBL^{+}10, MLG^{+}02, vMAT14, SPF^{+}07]. Alto [ACM01b]. Amazon [FFM^{+}23]. AMD [Str05]. AMD64 [Ano14a]. American [Boa90]. among [CDN02, LLF^{+}18, LTZ^{+}14, TtLcC13]. amplifying [DP11]. Analogy [Gai75]. analyses [BNS18, DAdBM^{+}24, HB13]. analysing [PV06]. Analysis [ACM05a, BE17, BFG^{+}14, BDG18, CC77, GMGF24, HT98, HKM^{+}18a, HB17, HWB03, JKK^{+}13, KNT02, LCK11, MM93, NMS^{+}14, Ost94, RI00, RRB19, SAR24, SM02, TKG89, VP16, WH99, WDL^{+}20, WLS^{+}18, ZTA^{+}21, ACM01a, AAH^{+}03, AMIA19, BDQR23, BBM09, BMER14, CBFH20, DD20, EBJ17, EMS15, FX06, GP13, GPW03, KSS^{+}20, KSS^{+}23, KS20a, LTZ^{+}14, MD73, MD74, MSG01, PFNC20, RRB17, RGS^{+}20, SMSB11, TLX17, Wün13, YJZY12, YSM^{+}21, YIR24, YI24, ZMD^{+}21, DHPW01]. Analysis-Driven [ACM05a]. Analytic [Bar73, Bar78]. Analytics [IGBKR19, KKE19, WTM18, KB17]. Analyzer [Ano03a, SHLJ13]. Analyzing [CVWL13, PV08, ZDK^{+}19]. ANCS [HLPY16]. Android [CXLX15, KLF^{+}15, MMP^{+}12, PDML25, STY^{+}14, THC^{+}14, ZYH^{+}19]. Anemia [RHV17]. Anemia-Like [RHV17]. Angeles [ACM06c, IEE84b]. Animated [PCR89]. annealing [RH17]. Annotated [MR04, RSF03]. annotation [ANH00]. annotation-aware [ANH00]. Announcement [Ano00]. Annual [ACM06a, Ano10, IEE85, IEE05, MS91b, Shr89, USE00a, USE01a, USE06, ACM06a]. anomalies [FRM^{+}15]. anomaly [Hui18, MW18, SIK^{+}16]. Anonymisation [VV18]. Ant [AAK18, PAC^{+}22, AP18, FS19, GGQ^{+}13, ZFL^{+}23]. Antfarm [JADAD06a]. Anti [SMA18, AKCP21, Sta07]. anti-debugging [AKCP21]. Anti-P2P [Sta07]. Anti-Virtual [SMA18]. anti-virtualization [AKCP21]. Antonio [ACM99, USE01b]. Anwendung [Bec09, Bor01, WF03, Zim06]. Any [WL96, FIF^{+}15]. AOT [WKJ17]. APA [JNR12]. Apache [FRM^{+}15]. apart [LBF12]. API [Ano14c, SM23b]. APL [Alf91]. App [ZYH^{+}19]. Apple [Sam22]. applets [Wes98]. Appliance [See10]. Appliances [BRX13, ZZW^{+}21, AEMWC^{+}12, BSM^{+}12]. Application [AJ18, AW17, BB17, BCZ19, CHW12, cCWS14, Cza00, HMH17, KNT02, KLF^{+}15, LWC^{+}17, LPZ^{+}22, MD73, MD74, PCW^{+}16, SFSN^{+}24, TB17, WGW^{+}18, ZYH^{+}19, AS14, BBS06, IBM88, Int88, IBM96, JSK^{+}13, JCZZ13, JDJ^{+}06, Kag09, Lia05, LBF12, LLS^{+}08, MRGB91, SE12, SWcCM12, SASG13, SL00, ZS01, ZBG^{+}05]. Application-Aware [AJ18]. application-specific [ZS01]. Application-transparent [AW17]. Applications [AHKC23, Ano99b, Ano03a, BAL15, Boa90, DSM^{+}18, DJS^{+}17, FBL18, HHV^{+}02, HSK17, HC17, HPB^{+}24, HCB18, IEE05, JW17, KKS^{+}19, LGJZ16, LSSC22, LH15, NKK^{+}06, Par71, PLMA18, Pfo13, PY93, RNS^{+}23, SN23, SS05, TR88, VP16, WLS^{+}18, WZKP19, XZL^{+}20, AS76, Alf91, AC16, AB16, ACT94, ABC^{+}07, BD11, BPM^{+}22, BSNB20, BTLNBF^{+}15b, BRS18, BOF17, BFS^{+}18, DMH18, DBC^{+}00, EBLM22, EF94, EMS15, GHD12, GTN^{+}06, GHH^{+}93, GLQ^{+}13, GKJ^{+}19, HKS19, HcC14, HKD^{+}13, HSC15, JPTE94, KRG^{+}12, LCL14, LF19, MCC18, MA19, NBS18, dOL12, PTM^{+}15, RNA^{+}22, R^{+}13, RSLAGCLB16, SZKY21, Sch13b, SGV12, SZ88, TDG^{+}18, TV18, WDCL08, WSX^{+}19, YYC^{+}19, YGN^{+}06, ZBP05, ZNSL14]. Applicative [AS85a, Abr82, AS85b]. applied [MM92]. Applying [FFM^{+}23, CSSE21]. Approach [ARAAA19, BC19, BFG^{+}14, BRX13, CFM17, CLW^{+}14, COV^{+}24, Cox09, DPCA11, DM75, EMAL17, Fie68, FPS^{+}02, FML^{+}22, Jen79, JQWG15, KC16, KAH83, LGS^{+}23, Mad69, MP16, MSC^{+}21, NLPV12, NSJ12, PF23, SDD^{+}16, SFSN^{+}24, VN06, WJ10, WVT^{+}17, WTJR22, XD17, YYY^{+}23, ZTWM17, AHRR22b, AH24, ADWM18, BML^{+}13, BHvR05, CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, CBZ^{+}16, DS22, GKP^{+}19, GHK24, GLLJ16, HLBZ20, KW13, KKB14, KF18, LH13, LU04, MD73, MD74, MK22, MAK18, MA21, MA24, NZH20, Par25, PSC^{+}07, PJZ^{+}19, Pon19, RWC21, SENS16, SHR19a, SHR19b, SEPV19, TZK17, TZY^{+}25, TSR19, XHCL15]. Approaches [BAL15, FMIF18, HM20, JK15, EYGS19, TIIN09]. Appropriate [ZRS^{+}16]. Approximation [GLW23, MNB24]. apps [MMP^{+}12]. April [Ano01b, IEE84a, USE01c]. Arbitration [SKJ^{+}17]. Architecting [SYC14, TZB19]. Architectural [DLLN18, DCP^{+}12, Gol73c, JR02, NMHS15, PJZ18, PEC^{+}14, SL12, CFS^{+}12, DLL^{+}16, PAKY16, RVJ^{+}01, WLL^{+}13]. Architecture [ASMA21, BBD^{+}91, BKMM87, Bos24, BDR^{+}12, BG73a, CAF^{+}91, DAH^{+}12, DSM^{+}18, DS09a, ECET18, EMW16, G^{+}05, Gol73a, Gol73b, Gum83, Han73, HW93, Hsu01, HWCH16, IEE85, KZB^{+}90, Kee77, KAA^{+}25, LW73, LMG00, LMG01, LGR14, MSS^{+}15, PCC^{+}16, PK75a, RC18, Rev11, SVM^{+}23, SJV^{+}05, SADP21, SSB03, SN05a, SJA^{+}17, SWF16, Sun99, TR88, TV12, Tur92, Uhl06, WIS^{+}15, You73, ZL18b, ZZW^{+}21, ZGW^{+}06, Ano94, DJP^{+}24, Ber86, BR01, BNS18, CCL^{+}17, CLDA07, DS09b, FS19, FMX^{+}25, FC98, GDSA^{+}17, GCARPC^{+}01, HIIG16, Hog02, HMS04, HPS23, IBM88, IIK^{+}06, Jou85, KW80, KNHH18, LLW^{+}12, LL14, MMTM22, MS01, MJ93, NOK^{+}85, OJG91, RFBLO01, Ros06, SIJPP11, SG09, SDN09, SOKE23, Wel02, YTS14, YYPA01, Yur02]. Architecture-aware [WIS^{+}15]. Architecture(R) [MBBS13]. Architectures [ACM06b, BN75, BDF19, EMAL17, ELC^{+}19, EG01, GG72, HW93, HHK94, Ian14, KKD24, PG74, PY93, QTR21, RD90, SXMX^{+}18, ZHH^{+}24, BGS13, DM93, EMI13, KMG^{+}18, NBS18, PNM^{+}20, PG73, Skr01, YZW^{+}13, ZP14, CEPR22]. Architektur [Dal97]. Archive [CLKEF21]. Area [BFG^{+}14, Fis01]. areas [BCZ19]. ARIMA [CSSE21]. ARINC [DPCL22]. Arizona [IEE05]. ARM [CJJ^{+}22, DN14, DLL^{+}16, DLLN18, GNDB16, MGL^{+}17, ZTWM17, PS19a]. Aroma [Sur01]. Arquillian [Ame13]. Array [MBK^{+}92, SV15]. arrays [dCJR16]. Arrivals [KMM13]. Arrives [Bai70]. Art [BGP00, SGB^{+}16, AEB19, BDF^{+}03, BDG18, MDD^{+}08]. Artificial [MR91, PS23, TVO92, AA24, BCM90, IM93, KCV11, NGN24, RK16]. arts [BB08]. ARVMEC [XLL^{+}20]. as-a-Service [ESY^{+}17, HPHV17]. ASM [SOAK23]. Aspect [YKP^{+}24, BADM06]. Aspect-oriented [YKP^{+}24]. Aspects [Hsu01, Kna93, PP73, EF94]. assembler [GBO87]. Assembly [BD01, SVB93, Ber86, Don88, Juo07]. Assembly-Language [SVB93]. Assessment [AJM^{+}25, FXL^{+}23, CDO24, PM19a, vCPWvT11]. assigned [AAT^{+}22]. assignment [AAM^{+}16, KMT14, WZV^{+}13]. Assist [Hor73, Olb78]. Assisted [AY24, CCML12, JSHM15, JAS^{+}15, PPG^{+}17, RTL^{+}18, AJH12, AEB19, GMK17, PDM20, TP25, ZYZ^{+}18]. Assists [OLZ16]. Association [Sof83]. Assurance [LJZ12, LLW^{+}12, MKM^{+}08]. Assuring [YDW18]. AST [ZLBF14]. asymmetric [CBGM12, KKJL14, KNHH18]. Asynchronous [Cav93, LJL^{+}11, MM93, RZPX19, SM01, WZKP19, WN17, vLSM01]. Atlanta [USE86, USE00a]. ATMS [CWG00]. atomicity [BHSB14]. attached [Mon97]. Attack [DL19b, WLCS17]. Attackers [CLS07]. Attacks [LSSC22, SL16, SYB12, TV12, WWL^{+}17a, GHD12, NS17, TMV25, VT14, WXW15]. attention [TP25]. Attestation [ZL16, KBC21, VT14]. attribute [FS89, SS19]. Auction [SZW^{+}16, TVKB16, ZG13, ZLH^{+}15]. auction-based [ZG13]. Auctions [ZHW^{+}17]. Auditing [SM90]. aufsetzen [RHM08]. augments [Bri98]. August [RM03, IEE96a, IEE96b, IEE97, IEE99, MR91, Ost94, USE93, USE00b, USE02]. Ausfalls [Mar08]. Austin [ACM75, IEE02, IEE03]. Australia [MR91]. Author [DM76]. Auto [TSCB19, TCTH23, KA25, TP25]. auto-encoder [KA25, TP25]. Auto-scaling [TSCB19, TCTH23]. AutoBoT [VS19]. Automata [JYOB18, NNK21, RGAT18, RT18, TLX17]. automata-based [RGAT18, RT18]. Automated [AD18a, ACM05a, Ano03b, BSSS14, HLP^{+}16, FGLI15]. Automatic [MS00, SMES01, SMA^{+}10, Sus76, WML02, ZLZ13, CL17b, MSZ09]. Automating [MJW^{+}06]. Automation [ACM06a]. automaton [Sig89]. Autonomic [LGJZ16, SKT^{+}19, SEK^{+}19, YWH^{+}21, SWC08, WDCL08]. Autonomous [SC17, NNK21]. autoscaler [MPM^{+}20]. Autoscaling [EGG^{+}24, Kov19]. Autoselection [KKE19]. Autotuning [KKE19]. Availability [LGXC23, RGS^{+}20, AAF^{+}09, Fu10, LDL^{+}08, MDZ^{+}21, MRC^{+}13, NMC18b, NMC18a, TUM18, YLH14]. Availability-aware [LGXC23]. Available [Ano03b, GI12, GVI13]. avatar [CKT08]. average [LDL14]. avionics [ABC^{+}07]. Avoidance [HS19, KAS13, LYS^{+}18, AHRR22a, OG16, PC21]. Avoiding [BLRC94]. Award [War11]. Aware [AJ18, AAK18, BMS16, BL17, CWH^{+}16, CGC16, CWL^{+}15, CTP^{+}17, CYX^{+}17, CHLY18, Do11, EGR15, EVCL21, GCL^{+}21, HC17, HTB22, HPP15, JJK^{+}11, JQWG15, KL14, LMM18, LXL^{+}22, LLY^{+}25, Man16, MA21, MA24, PYYG21, PHC20, RG17, SDD^{+}16, TB17, WZH^{+}16, XLL^{+}14, XLJ16, XLWX19, XZL^{+}20, YLH17, ZWFX17, ZCG^{+}17, ZLZ21a, ZWL^{+}18, dSdF16, AJBJ23, AT23, ADA^{+}19, AO16, AMAB17, ANH00, BSNB20, CD14, CCL^{+}20, CLL^{+}23, DXM^{+}17, DCMW17, EBJ17, FZS^{+}20, FA21, Fu10, GLK^{+}12, GA18, HKS19, HZL^{+}18, HH18, HH19, HLBZ20, HSC15, HC12, IRB19, IKU15, JNR12, JC18, JYOB18, KN18, KC16, KBDK22, KB21, KK21, KBB11, KCS14, KR16, KLF^{+}15, LYYY18, LYY^{+}20, LGXC23, LWL16, LWCZ22, LCL^{+}23a, LCL^{+}23b, LQD^{+}18, MMTM22, MHM19, MA19, PC21, PFPJ18, PF23, PS23, RNA^{+}22, RKT20, RH17, RHZ^{+}17, SSB^{+}14a, SHR19a, SHR19b, SSN12, SGV12]. aware [SS22, SZL^{+}14, SK13c, TDD20, WIS^{+}15, WCC^{+}16a, WDT18, XCJ^{+}14, XWW^{+}21, XLWZ18, XXWG23, YB24, YRJ18, YQZ19, ZHHC17, ZWC^{+}19, ZFL^{+}23, ZWH^{+}17, ZSRR22, JZY^{+}22]. Awareness [ZHL16, LCL14]. Azure [Fab13, RHV17]. Azure-Based [RHV17].

B [Req03]. B5500 [Ham76]. BA [KSS^{+}20, KSS^{+}23]. Back [KS08b]. backhaul [MCC18]. backpropagation [RWC21]. Backup [ACA16, KRS^{+}17, ZXW16]. Backup-Sharing [ACA16]. bad [RY10]. Bag [VS19]. Bahamas [Ano99b]. balance [YWH^{+}23]. Balanced [KAA^{+}25, LLW^{+}16, DS18, XZK^{+}20]. Balancing [ARAAA19, CJK^{+}12, CGC16, CL16a, DY17, Gua14, HPP15, KK19, LW12, LYS^{+}18, MKKE12, WWH^{+}16, WTL^{+}16, YWR^{+}14, Bir94, GH20, GPR23, KAZS14, RI25, TF16, Vac06, XH90, XTB17, ZWL09]. BalCon [GPR23].

Ballooning [LJL^{+}15, PGP19, LCLG26]. Baltimore [Ano93]. Band [ZSXZ07, PBYH^{+}08]. Bandwidth [ELC^{+}19, KDK20, LJFS17, LWZ^{+}18, YLH17, ZRS^{+}16, BAC15, GLLJ16, LZW^{+}15, THH^{+}14, WQG15, WXW15]. Bandwidths [LMM18]. bank [PAKY16]. Bare [AGH^{+}16, OSK15, RB24, ZXL^{+}24, GAH^{+}12]. Bare-Metal [ZXL^{+}24, AGH^{+}16, OSK15, RB24, GAH^{+}12]. barrier [Rix08]. barriers [LM99]. Base [UOKT84, WH08]. baseband [KWZ^{+}19]. Based [AHKC23, AAK18, Bad82, BAL15, BE17, CWL12, Cap21, CHW12, CLW^{+}14, CD12, CDD13, CDM^{+}25, DPCL22, DF96, ECET18, FD08, GGG03, HKM^{+}18b, HWHW18, JFPL16, JN15, KKD24, KP15, KLR^{+}20, KAZS14, LW11, LP14, LKL^{+}19, LCT^{+}15, LLWM23, LYY^{+}24, LGZ^{+}19, LW12, LZW^{+}17, LZM^{+}20, MJW^{+}14, MTFK19, MGL^{+}17, NASD21, Neu92, NL19, OVI^{+}12, PvDS08, Ran02, RZPX19, RHV17, RWX^{+}12, SAJ24, SJV^{+}05, SGK^{+}23, SXH^{+}19, SAR24, SDM21, SHZ^{+}14, SKJ^{+}17, TV12, WB81, WLS^{+}18, WTM18, YYY^{+}23, YWR^{+}14, YWW^{+}15, YLN^{+}17, ZSE^{+}25, ZDK^{+}22, ZQCZ16, ZLL^{+}20, ZXY^{+}15, ZB20, ZZW22, vLSM01, AD18a, ABB19a, AHRR22a, AAJD^{+}16, AAT^{+}22, Ano96, Ano06a, AB16, ALL06, AMA^{+}11, BD11, BLMP22, BL17, BSNB20, BY20, BNS18, CL17b, CBJ22, CPM^{+}18, CEPR22, CVWL13, CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, CWC^{+}14, CBZ^{+}16, CLcC13, CPST14, CPST15, CFRSSR19, CGV10]. based [CRG16, DD20, DP11, DS18, DC15, DLH^{+}20, DPCA11, EB20, ETAB22, EYG21, ESY^{+}17, FGZC23, FS89, FS19, FMJ15, FLCB10, FF96, FL13b, GTGB14, GDSA^{+}17, GH20, GR15, HM20, HKJ19, HBNK24a, HBNK24b, HL24, HO22, HOKO14, HWCH16, JWH^{+}15, JFZL17, Kag09, Kam13, KLY20, KS13, KS20a, KRCH14, KKB14, KDB16, KZ21, KA25, KK21, KM13a, KM13b, KJM^{+}07, KKJ^{+}13, gKEY13, KLK^{+}22, LMJ07, LBL16, LYYY17, LYYY18, LXRS19, LLZ^{+}19, LLX^{+}17, LLS^{+}08, LC13, LPZ^{+}22, LWCZ22, LCL^{+}23a, LCL^{+}23b, LMDP19, MCC18, MPA^{+}18, MW18, NZH20, NRdA^{+}20, NS17, Oi05, Oi06, Oi08, PFH^{+}16, Par25, PDM20, PGLG12, PMP23, uRQS20, QZDJ16, QBL^{+}23, QCM^{+}24, RGAT18, RH17, RHR20, RG19, RT18, RAP19, RCTY19, SJB14, SS13, SENS16, SG10a, SEM^{+}20, SGV13, SM23a, SS19, SM23b, SPF^{+}07, SYC14, SXMX^{+}18, SOKE23, SV17]. based [SCFP00, Sto07, SP23, TT96, TZY^{+}25, THB22, TXD^{+}24, TY14, TSCB19, TMV25, VT14, VGL23, Vog03, WKT08, WDCL08, WXZ^{+}17, WBW^{+}19, WGY20, WWL^{+}25, WW77, WYZAD20, XZ11, XZZ^{+}16, XJR^{+}17, XWX^{+}17, XYYY17, XA22, YC98a, YC98b, YZW^{+}13, YZLQ14, YLCH17, YBZ^{+}15, YWH^{+}23, YSM^{+}21, YLJ22, YC16, ZG13, ZMD^{+}21, ZLH^{+}15, ZWHC17, ZXB^{+}24, ZAI^{+}16, ZFL^{+}23, ZLL^{+}16, ZZW^{+}24, ZXR^{+}22, dSOK17, vKF13]. basic [A^{+}04]. basierende [Deu08]. Basis [Kar07]. Batch [KMM13, LD05, SS13]. Batched [GLL^{+}21]. bathymetry [MMG^{+}18]. Bay [Ano10]. Bayesian [LYYY17]. BCIM [ZSE^{+}25]. BCPL [Abr80, WW77]. BCPL-Slim [Abr80]. Be [Cox07]. beams [MC98]. Beautiful [SG09]. Bed [MMHP17]. Bedienung [KGG00]. Bee [PS23]. beetle [BRS^{+}22]. beginner [RR09, Wes98]. behave [Voe86]. Behavior [EG01, XWH^{+}16, ZDLG17, BSOK^{+}20, CL14, LWB^{+}15, Oi08, SEM^{+}20, Wol99, ZZW^{+}24]. behavior-based [SEM^{+}20]. behavioral [CL17b]. Behind [Cra98]. Belgium [ACM04a]. Benchmark [DHPW01, WZT19, GPW03, SMSB11]. Benchmarking [CGS06, MW25, RO16, AHK^{+}15, FLM^{+}08, KJ13, PRS26, ZS01, RB24]. benchmarks [LJN^{+}00]. Benefit [HB14]. Benefits [KWZ^{+}19, LS15, SIRP17, CM18]. Berkeley [USE01c]. Best [B^{+}07, BY20, GHS16, MS17, Sch13a]. best-fit-decreasing [BY20]. betreiben [RHM08]. Betriebssystem [CK06a, CK06b, CK06e, CK06c, CK06d, CK06g, CK06f, CK06i, CK06h, CK06j, CK06k, CK06m, CK06l, CK06n, CK06o, CK06p, CK06q, CK06t, CK06r, CK06s]. Betriebssysteme [WR07, WR08]. Better [MW05, Com00]. Between [Jen79, KLLT18, ZLHD15, BDJdS02, BRS18, CL17a, GLQ^{+}13, GSW^{+}17, KGS16, Mal73, EYGS19, SMR24]. Beyond [FPS^{+}02, ACM04a]. BGP [SAR24]. Bias [Lee16]. biased [ABDD^{+}91]. Big [ECET18, GTS^{+}15, MSG14, WXG^{+}25, WTM18, BOF17, DXM^{+}17, LMDP19]. Billing [RB17]. Bin [BB17, GR15, SXCL14, XDLS15]. Binaries [PA21]. Binary [BDG18, KLF^{+}15, WMUW19, ZSE^{+}25, ZFL15, dGG^{+}17, HLW^{+}13, JYW^{+}13, PGLG12, vKF13]. BIND [See10]. binding [KW13]. biodata [Wün13]. biogeography [ZLL^{+}16]. biogeography-based [ZLL^{+}16]. biology [Wün13]. Biopolis [Ano06a]. bird [Guy14]. birds [NGN24]. Birth [NOT^{+}17]. bison [Kag09]. bison/flex [Kag09]. bit [VED06, VED07]. Bitcoin [HB14]. BizOps [FBL18]. Black [NMMP15, VVB13, EB20, TZK17, WSVY09]. black-box [TZK17, WSVY09]. Blackboxes [KBK^{+}21]. blackhat [Ska07]. Blending [VSMC23]. Blessing [Kot10, Kot11]. BLFair [FMX^{+}25]. Block [LZY^{+}24, Sch94b, Sch94a, TLBW12, ZLL^{+}20, Zyt94a, Zyt94b, FFBG08, FLCB10, LLLE17, TKG89, WF07]. block-device [FFBG08]. Block-level [LZY^{+}24, FLCB10]. block-paging [TKG89]. Blockchain [CQLL18, DMH18, TDMP23, XKKL23, XJR^{+}17, GNK24, RI25]. Blockchain-based [XJR^{+}17]. blockchain-enabled [RI25]. Blocks [Lam75]. blowfish [GNK24]. blows [BBTK^{+}17]. Blue [SSU^{+}12]. BlueIO [JAD19]. Blueprinting [NLPV12]. board [CGV10]. Bochs [Ano14b]. bodies [AGIS94]. Bolton [ACM03b]. Book [Ano97a, Fro13, Lar09, Van98, B^{+}07, TC10, War02]. books [Van98]. boost [CBZ^{+}16]. boosting [AC16, LKY^{+}17, PGLG12]. Boot [NOT^{+}17, SB16, DBO^{+}18]. Bootstrapping [CBLFD12, Kam75]. Border [CDW^{+}24]. BOS} [RP07]. Boston [IEE85, USE01a, USE06]. Both [ZHL16]. Bottom [UOKT84]. Bottom-up [UOKT84]. bound [DJP^{+}24, JGA^{+}88, TSN^{+}23]. boundary [SBQZ14]. bounded [XHL^{+}13]. Bounds [SJMG24]. Box [NMMP15, TZK17, VVB13, WSVY09, XHCL15, MNS^{+}14, LGS^{+}23]. branch [CEG07, EG03, JGA^{+}88, JYW^{+}13, WHC16]. branch-and-bound [JGA^{+}88]. branch-and-price [WHC16]. branches [KJM^{+}07]. Breadth [MNS^{+}14]. Breaking [VMW^{+}19, GKBB15, Rix08]. breed [Arm98]. Brewing [WZL^{+}18]. Bridge [Men03]. BridgeGC [WXG^{+}25]. Bridging [ACM04b, FL13a, GSW^{+}17]. Brighton [Vra05]. bring [XKY^{+}11]. Bringing [BDR^{+}12, PPS^{+}18, STS^{+}13]. Broadcast [SXH^{+}19]. Brokered [BB17]. brokering [TMMVL12]. brokers [PAKY16]. browser [FIF^{+}15]. Browsers [YML^{+}18]. BSD [WF03]. Buch [KGG00, Tho08]. buddies [WTLS^{+}09]. Budget [BE17, RB17, SJ21]. Budget-Driven [RB17]. buffer [JADAD06b]. buffers [CFG^{+}13]. Bug [Ano97b, Ano15]. Build [Kol19]. Building [AAB^{+}05a, CGM17, DBC^{+}00, DF96, HWCH16, PEC^{+}14, SJV^{+}05, See10, TSP17, UIK19, Nie12, RG19, SG10b, WH08]. Burstable [WUNK17, DAdBM^{+}24]. Burstiness [RKT20]. Burstiness-aware [RKT20]. bursts [DP11]. bus [HHPV15]. Buying [YLN^{+}17, ZLH^{+}15]. buying-based [ZLH^{+}15]. BYOD [DMG^{+}15]. Bypass [LHAP06]. Bytecode [MO98, SEK^{+}19, HLW^{+}23]. bytecodes [SUH86].

C [Fra06, Fra09, Hee07, Hog06, Hog08, Wil06, ZB18, Blu02, CWG00, G^{+}01, Hee07, Hog06, Hog08, JM08, Men03, Siv07, Wil06, YIR24]. C# [G^{+}01]. c-mean [ZB18]. C-RAN [YIR24]. C/C [Blu02]. CA [ACM06a, ACM06c, Ano97a, IEE84b, IEE93a, USE01c]. Cache [HS21, JQWG15, KR18, KAA^{+}25, NsP16, RHR02, SDS^{+}21, TBS17, vSMK^{+}20, Boz89, JADAD06b, Oi05, RJK16, ZP14, AMA18]. CacheInspector [SDS^{+}21]. CacheOut [vSMK^{+}20]. caches [BLRC94, DSR23]. Caching [AMA18, ASMA21, KJL11, LGZ^{+}19, MM93, LM99, XWX^{+}17]. CAFE [ZFH^{+}22]. Calculations [Bad87, Hol95]. Calculus [ABV12, SJMG24, Wat86, Wat87, WK90]. Calif [ACM01b]. California [ACM05a, Ano01b, Ano04b, Ano10, IEE96a, IEE97, IEE99, USE91, USE99, USE01c, USE02, IEE84a, IEE90a, IEE91, Tho93]. Call [DEK^{+}03, Lee16, PULO16, PVRR14, SSB^{+}14a]. Call-site [SSB^{+}14a]. calling [HB13, SSB^{+}14a]. calls [VMBM12]. Cambridge [USE93]. CAMIG [HTB22]. Can [Cox07, GW07, THB06, Sig89]. Canada [ACM06f, Sof83]. CAOS [Sch86]. Cap [HC17]. Capabilities [TVO92, WZT19, Ame13, AAB^{+}05c, Fit14]. Capability [ECET18]. Capable [Ott18, PST^{+}15a]. Capacity [BB17, HMH17, LYGG20, SJ21, WUK^{+}18]. capo [SMSB11]. Capping [HSK17, JKK^{+}13]. Capture [SCFP00, Sur01]. Capture/Replay [SCFP00]. capturing [BKC^{+}13]. Carbon [SSS25]. Card [Siv04, SUN97, HM01, Req03, JCV99]. cards [GLV99, TLBW12]. carrier [FZS^{+}20]. carry [Ame13]. carrying [FCG^{+}05]. Cascade [YYL^{+}15]. cascading [HL13]. Case [FFM^{+}23, GGG03, HBL^{+}10, HWB03, Ian14, PK75a, PS19b, vCPWvT11, CDO24, GGK19, HIIG16, MN03, RK18, Sig89, SIRP17, Vit14]. Case-Based [GGG03]. Cases [FG91]. Cassandra [FRM^{+}15, SLC20]. Catalyst [Ano03a, GMK17]. Catching [SXH^{+}19]. Categories [Gai75]. catering [RNA^{+}22]. causes [FRM^{+}15]. CBase [ZLZ^{+}19b]. CCAP [JQWG15].

CCGrid [TLC06]. CCHybrid [Yu20]. CD [Joo06]. CDN [LYS^{+}18]. cell [MFT^{+}19]. Cells [DAH^{+}12]. Cellular [SLZ^{+}25, ALW15, Sig89]. Ceno [LZZ^{+}25]. Center [AT24, Ano93, Car14, CGC16, DY17, FML^{+}22, IEE90b, PCC^{+}16, RK24, WN17, XWJX15, ZDK^{+}22, CLL^{+}23, HKB19, HUWH14, IRB19, JYOB18, KBDK22, LZW^{+}15, Man15b, MRM06, MBM09, NGN24, NTH^{+}17, RGS^{+}20, RMP24, TDD20, VOS12, WDCL08, WZV^{+}13, YPLZ17, YGLY21, YLJ22, ZLZ^{+}19b, ZWH^{+}17, Car13]. Centers [AJ18, AAAF21, AGC18, BB13, CL17a, CTP^{+}17, EGR15, GA25, HTW^{+}19, HO22, JFPL16, KMM13, LMM18, LVM16, LLWM23, Man15a, Man16, SB16, SSS25, SJMG24, WZH^{+}16, YLH17, YWW^{+}15, ZHL16, dSdF16, ARA18, ARA20b, ARA20a, AD19, AJBJ23, AGH^{+}15b, AGH^{+}15a, AHRR22b, ATZP21, ATS16, AMAB17, ARMMA18, BB12, CFRSSR19, DLH^{+}20, FLL^{+}13, GH20, GHK24, GSKJ18, HM20, HTB19, HLBZ20, IKU15, IPRS21, JB24, KDK20, KZ21, KTB17, LCLG26, LZC^{+}16, MNB24, MAK18, MHM19, NZH20, PC21, Par25, PVRR14, Pon19, uRQS20, RK16, RH17, RT18, RK18, RJK^{+}17, RGS^{+}20, SBI21, SHR19b, SBNU18, SS19, SM23b, WCY^{+}17, WHW20, WTLS^{+}09, XLQL18, XXWG23, SHR19a]. centralized [Fis91]. centric [AAMBE21, PAKY16, SBBP20]. Certain [Han73, JHS12]. Certains [Han73]. Certified [Khn09, IIPB09]. Cetres [KAS13]. CéU [SIR^{+}17]. Chain [EMAL17, GR20, HYL^{+}25, HJG18, LW20, MTFK19, MSC^{+}21, RH17]. chain-based [RH17]. chained [TSN^{+}23]. Chaining [AP22, GGK18, KBK^{+}21, LLW^{+}16, LYL21, MP16, WCC20, XZL^{+}20, GHM^{+}18, LKR^{+}19, SHB19, TSCB19]. Chains [FBM^{+}21, JWL^{+}18, KLLT18, LGXC23, PHXL19, ZLZ^{+}21b, CGC^{+}24, NRS92, RHR20, XHW^{+}19, YXL^{+}20]. challenge [STMV18, Sam22]. Challenges [ARA20a, AFG^{+}17, AP22, Cap21, Car23, JW17, KK19, MZ20, Nie12, SCC^{+}23, SABL20, SG10b, AEB19, BCZ19, CM18, FJKK17, JAC^{+}19, LDDT12, MA10, MA17, PCB^{+}18, TIIN09, ARA20b]. change [ZL13]. Changing [Mac79]. Channel [LGR14, PF23, TTH^{+}19, LF19, MN03, RI25, WXW15]. Channel-aware [PF23]. Channels [Hu90]. characteristic [CJJ^{+}22]. Characteristics [SHW^{+}15, CWC^{+}14, CCW^{+}20]. Characterization [AMA^{+}14, CGS06, DS09a, IEE02, IEE03, NASD21, ACM06c, RVJ^{+}01, SSB^{+}16, TSN^{+}23]. characterize [LJN^{+}00]. Chatten [Joo06]. Cheat [Rul07]. checking [BHSB14]. checkpoint [BBHL08, EBLM22, GPS^{+}18]. checkpoint/restart [BBHL08, EBLM22]. Checkpointing [ECJ^{+}16, ZLW^{+}19b, PEL11, SGV12, TSLBYF08, dSOK17]. checkpointing-enabled [SGV12]. chemical [KK21]. Cherub [JCZZ13]. Chez [FDD^{+}19]. Chicago [ACM05d]. Chicken [PS23]. Chip [GGM^{+}16, Mon97]. Chips [FRD^{+}08, IEE97, IEE99, IEE96a]. Choices [NGRF19, XDLS15, Ano93]. chromatin [MFT^{+}19]. CICS [R^{+}13]. circuit [Bur02, KKC^{+}16]. Circuits [NLD^{+}23]. Cities [AEA^{+}23]. clairvoyant [KS18b, ZLW^{+}19a]. Class [LCWB^{+}11, LB98, Pat12, SS17, Won97]. classes [Bor07, OKAM17, Skr01]. classical [SGS92]. Classification [VLZL16, CWC^{+}14, YSM^{+}21]. classification-based [CWC^{+}14]. Cleancache [VTW16]. CLI [ECM01, ECM02, ECM05, ECM06, Int06b, Int06c, Int06a, Fra06, Fra09, Hee07, Hog06, Hog08, Siv07, SNS03, Vog03, Wil06]. CLI-based [Vog03]. Client [RSW^{+}06, DPW^{+}09, HIIG16]. CLIP7 [Lau87]. clock [DCA17]. Clones [ZCJ^{+}21]. Cloning [LCWB^{+}11]. Closing [ZLHD15]. Cloud [AJ18, AT24, AVNR19, AAAF21, AAR22, AGC18, AD18b, ASSB18, BB13, BLMP22, BCW20, BHEP14, CWL12, CPKL17, CFM17, Cap21, CPS17, CZX^{+}19, CTP^{+}17, DSM^{+}18, DKW15, ETAB22, ELC^{+}19, FBL18, FFM^{+}23, GB19, GA25, GLS15, GSW^{+}17, HMH17, HKLM17, HPB^{+}24, HW12, JE12, JQWG15, JW17, KC16, KKE19, KSVR23, KCY22, KMM13, KAS13, KAZS14, KK19, LCWB^{+}11, LKIL19, LGR14, LGJ^{+}18, LSSC22, LLWM23, LZY^{+}24, LW12, LH15, LWZ^{+}18, LS15, MSG14, MB15, MK13, Man15a, Man16, Man18, MJW^{+}14, MSLY24, MZ20, MPA^{+}18, MB20, MMHP17, NGN24, NLPV12, NSJ12, NHL22, PJZ18, PHXL19, PCW^{+}16, PXG^{+}17, PS16, PCC^{+}16, PG17, PG18, RK24, RSNK17, RSGG15, RWX^{+}12, SAJ24, SL14, SGK^{+}23, Sar16, SSS25, SJS^{+}17, SC18, SZW^{+}16, SV13, SLZ^{+}25, SB18, SXCL14, TB17, TVKB16, TMMVL12, TCTH23, VGL23, WDL^{+}20, WZH^{+}16, WVT^{+}17]. Cloud [WUNK17, WUK^{+}18, WLS^{+}18, WTM18, XSC13, XWJX15, XLL^{+}14, XLJ16, XLWX19, XLL^{+}20, XRL^{+}22, YLN^{+}17, YP15, ZDK^{+}22, ZDS^{+}22, ZQCZ16, ZL16, ZCG^{+}17, ZL18b, ZLZ^{+}21b, ZXL^{+}24, ZWL^{+}18, ZB20, ZHL16, ZLW18, ARA18, AD19, AJBJ23, AGH^{+}15b, AGH^{+}15a, AHRR22b, ATZP21, ADA^{+}19, AA24, AB16, AO16, AMA^{+}14, ATS16, AMAB17, ARMMA18, AP18, AEB19, AA18, AAC^{+}17, BB20, BD11, BTMS10, Beg12, BMF23, BCC^{+}15, BRS^{+}22, BSNB20, BFS^{+}18, CSMB15, CL14, CRX24, CEPR22, CSSS11, CCL^{+}20, CLL^{+}23, CBFH20, DL19a, DS18, DC15, DEG^{+}17, DQLW15, DLH^{+}20, DHD20, DCMW17, DS22, EYG21, FLL^{+}13, FAA17a, FAA17b, FS19, FF22, FA21, FPGK18, FMIF18, Fro13, GGQ^{+}13, GTGB14, GSG22, GDSA^{+}17, GH20, GHK24, GEK25, GLK^{+}12, GKJ^{+}19, GA18, GNK24, HKS19, HZL^{+}18, HBNK24a, HBNK24b, HTB19, HAK22, HLBZ20, HS13, Hui18, IRB19, IKU15]. cloud [JES^{+}15, JWH^{+}15, JC18, KN18, KDK20, KBDK22, KSS^{+}20, KB21, KSS^{+}23, KS20a, KSO^{+}15, KSRL10, KS18a, KMT14, KZ21, KTB17, KK21, KCS14, KJLY15, KCKC15, KU24, LLW^{+}12, LZWC13, LZWD15, LZC^{+}16, LLF^{+}18, LLWW18, LCL14, LLS14, LL14, LTZ^{+}14, LPZ^{+}22, LP11, LPBB^{+}18, LY23, MR23, MNB24, MK22, Man15b, MNA16, MK19, MK23, MMTM22, MW18, MA21, MA24, MA17, MHM19, MA19, MPM^{+}20, MMG^{+}18, NZH20, NNK21, NAR19, NS17, Nie12, NIA18, dOL12, OL13, PKS^{+}19, PAKY16, PM19a, PDM20, PFPJ18, Pon19, PS23, uRQS20, RK16, RKT20, RGAT18, RH17, RHR20, RG19, RWC21, RT18, RQD^{+}17, RK18, RJK^{+}17, RI25, Ros14, RGS^{+}20, RMP24, SZKY21, SSG^{+}20, SBI21, SHR19a, SHR19b, SG10a, SEM^{+}20, SGV13, SM23a, SASG13, SSEA18, SM23b, SBP^{+}17, Str13, SPSL25, TZK17]. cloud [TMLL14, TDD20, TXD^{+}24, TUM18, VT14, WCY^{+}17, WGY20, WWL^{+}25, WB16, WLL^{+}13, WRSvdM11, WRS^{+}15, WXW15, WCZ^{+}23, XHL^{+}13, XZZ^{+}16, XTB17, XLWZ18, XZK^{+}20, XA22, XXWG23, YLH14, YLHJ14, YLCH17, YW20, YJZ^{+}21, YBZ^{+}15, YWH^{+}23, YKP^{+}24, YB24, YRJ18, ZWKX17, ZYZ^{+}18, ZLZ13, ZWHC17, ZHHC17, ZWC^{+}19, ZFL^{+}23, ZWH^{+}17, ZYLY18, ZSRR22, BB12, CD14, CFVP12, CMG^{+}19, KKB14, KA25, KBB11, KMG^{+}18, XYYY17, YI24, SJL20]. cloud-assisted [ZYZ^{+}18]. Cloud-Based [Cap21, WLS^{+}18, BLMP22, MPA^{+}18, BSNB20, XYYY17]. cloud-computing [ZLZ13]. cloud-distributed [AB16]. Cloud-Edge [KSVR23]. cloud-Internet [KB21]. Cloud-of-Things [CMG^{+}19]. cloud-oriented [Beg12]. Cloud-Radio [SLZ^{+}25]. Cloud/Virtual [YP15]. ClouDiA [ZLV^{+}12, ZBS^{+}15]. cloudlet [YBZ^{+}15]. cloudlet-based [YBZ^{+}15]. Cloudlets [RSN^{+}18]. CloudMon [WLLZ16]. CloudNet [WRSvdM11, WRS^{+}15]. Clouds [AD11, CRZH15, ESY^{+}17, HTB22, HKM^{+}18b, HLPY16, HKKW13, HS21, KMK16, KDB16, KPHA20, Kov19, LWLL10, LLZ18, MLXG19, NMG15, OG16, OSK15, RG17, RB17, SBBP20, SDS^{+}21, SCL^{+}19, WZL15, WLLZ16, WHD^{+}16, WWL^{+}17b, XCSM18, YWY^{+}17, ZHW^{+}17, ZRZY15, ASB18, BB15, dCCDFdO15, DXM^{+}17, FBZS12, FGG14, HZZ^{+}14, KMK10, KR16, LMV12, LBZ^{+}11, LWLL16, LLZ^{+}19, PPO14, QXH18, RCTY19, STMV18, SYMA17, TSCB19, VSMC23, XJWW15, ZG13, ZLH^{+}15, ZLW^{+}19a, ZB18, ZLV^{+}12, ZBS^{+}15, EMS15]. Cloudscheduler [BCW20]. Cloudsim [OBSR16]. CloudSimSDN [SHB19]. CloudSimSDN-NFV [SHB19]. Cluster [BKM25, CL16b, GKSP99, LWZ^{+}18, SEF^{+}06, TLC06, ZCG^{+}17, FLCB10, KJLY15, LJL12, SBP^{+}17, SSN94, WDT18, WLG^{+}11, XLQL18, YLHJ14, YCL^{+}18, GWZ16]. Cluster-Aware [ZCG^{+}17]. cluster-based [FLCB10]. Clustered [DJS^{+}17]. Clustering [ARAAA19, SGK^{+}23, XZZ^{+}16, ZWHC17, GHK24, LQD^{+}18]. Clustering-based [XZZ^{+}16]. Clusters [CHPY17, EGG^{+}24, GSW^{+}17, LZ15, LWW16, PXG^{+}17, WIS^{+}15, YWCF15, ZLW^{+}14, AO16, CP17a, Fu10, HCJ07, KOY05, KS18b, PRS16, RHR20, SJJ^{+}12, ZWKX17]. CMD [CWC^{+}14]. CMI [MPM^{+}20]. CMS [SNC91, IBM96]. CNC [Lia05]. Co [DCG12, HS06, LGJZ16, LH16, SOAK23, WIDP12, CCW^{+}20, LF19, OG16, Wu13, YWGH13, Yu20, THB22]. Co-Design [LH16]. Co-Designed [HS06, DCG12, Wu13]. Co-evolution [WIDP12]. Co-Located [LGJZ16, SOAK23]. co-location [OG16]. Co-Location-Resistant [THB22]. co-resident [LF19]. co-scheduling [CCW^{+}20, YWGH13, Yu20]. COBOL [IBM88, Int88, TT96]. Cocoa [YLN^{+}17]. Code [AC98, CDN02, Dom80b, Fra83, GFH82, GHF83a, GHF83b, HWR^{+}24, MKZ^{+}25, RJK16, VGF16, WKJ20, WNL^{+}83, Ano15, DNR06, EL98, FC98, FCG^{+}05, HK07, HLW^{+}13, JM08, NG13, PV08, tTR82, UTO13, WKJ17, WGF11, Cox12]. code-copying [PV08]. Coded [ZLL^{+}20]. Codesign [KAJW93]. CodeWeavers [Ano03b]. Coding [CFLL19, Hsu01, IMBB20]. Cognitive [SN23, ZYZ^{+}18, AAJD^{+}16]. Coherence [YVCB17, YVCB18]. coherent [LKY^{+}17, ZP14]. COISA [AMB^{+}17]. Cold [BZD17, BBTK^{+}17, KSS24, WGF11]. collaboration [GLQ^{+}13].

Collaborative [Cap21, FXL^{+}23, IEE06a, SLW^{+}24, XWH^{+}16, ZCG^{+}17]. Collapsed [KBBM15]. Collecting [DS16]. Collection [ADM98, Ano03b, BS90, HPHV17, SHB^{+}03, URJ18, BOF17, DEE^{+}16, PBAM17]. Collection-Oriented [BS90]. collections [BDT13, SV15, SV17]. collective [SEM^{+}20]. Collector [GTS^{+}15, WXG^{+}25, WK08]. Collectoren [Sch13a]. collectors [Sch13a]. colocation [WTLS^{+}09]. Colony [AAK18, PAC^{+}22, AP18, FS19, GGQ^{+}13, ZFL^{+}23]. Colorado [USE00b]. Comandos [CTS^{+}93, MC93]. Combat [Cap21]. Combating [GG11]. Combinatorial [HMH17, ZG13]. Combining [BPP^{+}17, MK19, RSLAGCLB16, YJZY12]. COMMA [ZNSL14]. Commandos [MC93]. CommC [BKM25]. Commodity [KLK^{+}22, RTL^{+}18, Ros99, ZTWM17, BK14, CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, CLDA07, TLBW12, BKM25]. Common [CK87, Cro93, Int05a, Int05b, RB24, Int06a, ECM01, ECM02, ECM05, ECM06, Int06b, Int06c, Int06a, MR04, PW03, RSF03]. Communicating [LSPP^{+}23, KDK20, SK13c]. Communication [AAT^{+}22, CL17a, CZH^{+}25, CK06b, CK06e, DJ77, GGM^{+}16, HW15, Jen79, MTFK19, RLZ^{+}16, YC98a, YC98b, BML^{+}13, DSC^{+}08, DJ76, GI12, Kip21, Tho93]. Communications [NKK^{+}06, SGK^{+}23, CFVP12, HSC15, MN91]. communities [ACM04b]. Community [FMJ15, Var91, AAB^{+}05a]. Community-based [FMJ15]. compact [AMB^{+}17]. compaction [WK08]. companies [STMV18]. Comparative [LJL^{+}15, EYGS19, SAR24, SFSN^{+}24, Van98]. Comparing [Gal11]. Comparison [Do11, EDS^{+}15, Ng01a, Ng01b, QNC07, AA06]. Compatible [ZFL15]. compensation [XNH21]. Competition [CRZH15]. Competitive [BFG^{+}14]. Compilation [ACM06b, Cla97, FM90, JK13, KS13]. compile [RAT17]. compile-time [RAT17]. Compiler [GFH82, Har77, FS89, YC16, THL03]. Compiling [BS90, BSUH87, Ode87, Wak99]. Complete [Bod10, Fis09, LJN^{+}00, RRB17, War02]. completion [MNT14]. Complex [KAZS14, Sig89]. Complexity [SSH17, SP22, Bod88, FS08, GLK^{+}12, Sub08]. Compliance [HC18]. Compliant [CF00, HWCH16, LDRS18]. Component [Ano03b, BSNB20, WML02]. Component-aware [BSNB20]. Components [PM19b, HPHS04, IKU15, VWT13]. Composable [JHE14]. compose [RGS^{+}20]. Composed [Wel94]. Composite [DKW15]. composition [PFNC20]. compositional [Yel99]. compound [VMBM12]. Comprehensive [AP22, AEA^{+}23, HSN17b, LV99, PCW^{+}16, PS19a, TFtLcC15, GP13, GEK25, MFT^{+}19, MA17, NMC18b, NMC18a, RHR20, YWL^{+}18]. compressing [JDW^{+}14]. Compression [HKKW13, SHTE11]. compromise [CD01]. CompSC [PDC^{+}12]. Computatio [HW93]. Computation [MTFK19, RWC21, TCTH23, ZSE^{+}25, CMP^{+}13, CKP^{+}93, KJJ^{+}16]. Computational [MB20, THLK10, Wün13, YQZ14]. Computations [TDMP23, Kra90, NOR15]. Compute [GSW^{+}17, KL13]. Computer [ACM81, ACM06d, Ano93, Arm78, BGS89, BG74, CCO^{+}05, DM75, Gol73c, Hsu01, IEE85, IEE90a, IEE91, IEE05, Nel04, PBR^{+}90, SS75, SI81, Tho73, Tur92, WR07, WR08, ZR06, Ågr99, BR01, DG05, DTW07, DCA17, FFB^{+}00, GE85, GD08, Hog02, Jou85, Juo07, KW80, LBP^{+}07, ME87, MS01, Pou90, Ros06, Skr01, Spi06, SS72, Sus76, WO75, YYPA01, Yur02, Mon97, Osb01, War11]. Computers [BP99, BKMM87, BS90, KD78, MSS^{+}15, Say67, HP77, SGGB99, SGGB00]. Computing [ACM98, ACM04b, ACM05b, ACM06e, Abr80, AAMBE21, AGC18, AD18b, BMJ^{+}22, BCW20, Bos24, BHEP14, CWL12, CPKL17, CFM17, DDS^{+}94, DPCA11, Gei02, GB19, HCB18, HYL^{+}25, HW12, IEE96b, IEE04, IEE06a, IBBA20, KC16, KGZ^{+}04, KK19, LCK11, LGXC23, LW12, MSG14, MK13, MZ20, MO98, NLPV12, NSJ12, PCW^{+}16, PXG^{+}17, PLZ20, PS16, RCM^{+}12, RSNK17, RSN^{+}18, SGK^{+}23, SCSL12, SZW^{+}16, SEF^{+}06, SB18, SFSN^{+}24, TLC06, USE93, Vog03, WDL^{+}20, WB81, WCC20, WTM18, XSC13, YLN^{+}17, ZL18a, ZL16, ZZF06, ZAI^{+}16, ZD18, ZXG^{+}24, ZB20, ADA^{+}19, AA24, Ano96, AMA^{+}14, ARMMA18, AEB19, BB20, BMF23, BS96, CD14, CMGI^{+}23, CDM^{+}10, CCL^{+}20, CLM24, DQR^{+}13, DS18, DHD20, DCMW17, FGZC23, Fis91, FF96, Fro13, Fu10, GGQ^{+}13, GLA^{+}08, GEK25, GNK24, HKS19, HKJ19, HBNK24a, HBNK24b, HAK22, Hui18, JC18, JPTE94, dCJR16, KSS24]. computing [KHL17, KSO^{+}15, LBZ^{+}11, LLW^{+}12, LZC^{+}16, LCL14, LTZ^{+}14, LP11, LPBB^{+}18, LFHS23, LY23, MB21, MNA16, MK19, MK23, McG72, McK11, MFT^{+}19, MUKX06, M^{+}06, MA21, MA24, MA17, MA19, MMG^{+}18, NRdA^{+}20, NAR19, NIA18, PSZ^{+}07, PM19a, PDM20, QZDJ16, RNA^{+}22, RKT20, RGAT18, RHR20, RWC21, RHZ^{+}17, RQD^{+}17, Rob06, RI25, SSG^{+}20, SBI21, SEM^{+}20, SM23a, SJW^{+}13, SASG13, SSEA18, SB10, SHB19, Str05, TMLL14, TXD^{+}24, TMJ^{+}21, VGL23, WGY20, WWL^{+}25, WH08, XTB17, XLWZ18, XA22, YB24, YRJ18, ZLZ13, ZWHC17, ZLZ^{+}19a, ZYLY18, ZSRR22]. con [SMSB11]. Concept [AH68, Mad69, SIJPP11]. Concepts [PPTH72, Ågr99, Don88, MS01]. Concerning [Ker15]. Concerns [PM19b, VN08]. concolic [LLS^{+}12]. Concurrency [HTB22, MD12, CFS^{+}12, Sub11, UR15]. Concurrency-Aware [HTB22]. concurrency-safe [CFS^{+}12]. Concurrent [GMP89, Har77, KD78, IT86, WK08, YWGH13]. concurrently [SLC20]. Conditioned [WC01]. Conference [ACM81, ACM90, ACM96, ACM97, ACM00, ACM01b, ACM04b, ACM05d, ACM06a, ACM06b, ACM06f, Ano93, Ano99b, Ano01a, Ano02, Ano04a, Ano06a, BW03, DC15, IEE84b, IEE93a, IEE05, LCK11, Mar81, MS91b, MR91, Sof83, SS05, Shr89, USE99, USE00a, USE01a, USE01b, USE06, ACM05c, ACM06e, IEE06b, JPTE94, USE85, USE86, ACM00, IEE85]. Confidence [AHRR22a]. Configurable [WJGA12]. Configuration [BRX13, Lar09, A^{+}04, FL13b, SMA^{+}10]. Configurations [RB24, LDL^{+}08, RGS^{+}20]. configure [Car14]. Configuring [AL05, Rul07]. Confirmation [MTFK19, OG16]. conflict [BLRC94]. Conflicts [KPHA20]. conformation [MFT^{+}19]. CoNFV [ZSP^{+}21]. Congestion [CL16b, GR20, LYS^{+}18, PHC20, YLH17, ZWC^{+}14]. Congestion-Aware [PHC20, YLH17]. Congress [GHH^{+}93]. conjugate [MM92]. Connected [SMES01, MS00]. connection [MJ93, Tur84, XJW^{+}18, TR88]. connections [FBZS12, Ker15]. Connectivity [NVV^{+}24, TZB19, VOS12]. conservation [RK16, TDG^{+}18]. Conserving [DP11]. Consider [SC18]. consideration [YGLY21]. Considerations [GR20, G^{+}05, ZJRW19]. Considered [NMHS15, WC01]. Considering [XLWX19, JB24, LTZ^{+}14]. consistency [FRM^{+}15, LC14]. Consistent [DJS^{+}17]. Consolidated [HC17, HPP15, JJK^{+}11, KKJL14, OL13, SS13, ZLL^{+}16]. consolidating [BJ22, JB24]. Consolidation [AJ18, AAK18, BB13, BJC24, DCM22, GA25, HO22, HW12, LCLG26, LVM16, LLY^{+}25, MAK18, PZW^{+}07, SBK15, XCSM18, XLWX19, YWW^{+}15, ZDK^{+}22, ARA18, ARA20b, ARA20a, AGH^{+}15b, ATS16, AMAB17, AP18, BB12, BB15, BJG19, BJ20, BRS^{+}22, CD14, CSSE21, DLH^{+}20, DHD20, FAA17a, FAA17b, Fro13, GPR23, GKJ^{+}19, HM20, HMH17, HZZ^{+}14, HH19, HLBZ20, IRB19, JFZL17, JYOB18, KA25, gKEY13, KS20b, KCV11, KR16, KU24, LZC^{+}16, LBL16, LYYY17, LYYY18, LLWW18, LYY^{+}20, LL14, LWCZ22, LQD^{+}18, LY23, LDDT12, Man15b, MA19, NZH20, NGN24, NTH^{+}17, PC21, RT18, R^{+}02, SENS16, SHR19a, SHR19b, SBNU18, SSN12, TDD20, VGL23, WCC^{+}16a, WCZ^{+}23, YWH^{+}23, YRJ18, ZLCZ18, ZSRR22, RMP24]. consolidation-aware [WCC^{+}16a]. constituent [RHR02]. Constrained [EGR15, LTE12, LLS^{+}20, TV18]. Constraint [LFBB94, SAM14, DQLW15, HH18, LYYY18]. constraint-based [LYYY18]. Constraints [AD18b, BB13, FML^{+}22, KKS12, SJ21, LLZ^{+}19, NZH20, SZ13]. Constructing [CGC^{+}24, DM93]. Construction [BJPS73, SV24, XYYY17, YCL^{+}18]. consumer [PAKY16]. consumer-centric [PAKY16]. Consumption [DSM14, HKM^{+}18b, MV16, FAA17a, FAA17b, FFB^{+}00, KSS^{+}20, KSS^{+}23, LCLG26, DPBK16, RJK16, SMSH18, THG^{+}18, TUM18, VED07, VWT13]. Container [ECET18, EGG^{+}24, HO22, HWR^{+}24, QCM^{+}24, SPF^{+}07, YLN^{+}17, ZTA^{+}21, ZB20, ZLW18, CMG^{+}19, GKP^{+}19, HL24, Ker15, KA25, MG19, SG10a, Str13, TZY^{+}25]. Container-Based [ECET18, YLN^{+}17, HO22, QCM^{+}24, SPF^{+}07]. Containerization [HSL17, SFSN^{+}24]. Containerized [HSL17]. Containers [BJC24, EGG^{+}24, Kov19, MBK24, Ran20, SCA^{+}24, ZXR^{+}22, BJ22, DSS19, DL19a, MK19, MK23, MFT^{+}19, Ros14, SMSH18, SLC20, WGW^{+}18]. Containment [CLW^{+}14]. Content [CWH^{+}16, FLZ17, LYS^{+}18, MSC^{+}21, GVI13, HKN22, LLF^{+}18, LLWW18, XJR^{+}17]. Contention [JQWG15]. Contention-Aware [JQWG15]. contents [BTLNBF^{+}15b, BTLNBF^{+}15a]. Context [DMG^{+}15, LCMV17, TMV12, ZL18a, vLSM01, HB13, SSB^{+}14a, SM01]. Continuous [DL89, TSLBYF08]. Continuum [Bad87, MSLY24, SFSN^{+}24]. Contract [FZD^{+}24, SCF^{+}25]. Contraction [Par79]. contracts [ZBS^{+}22]. Contribution [Han73, ABB^{+}19b, KA25, Han73]. Control [Aln22, AGLM91, Att79, CFLL19, CL16b, Com65, Cre65, DL19b, GGK18, HS19, HHC^{+}16, LZ15, LGJZ16, LXL^{+}22, LXQ24, MMdMCPSR25, PSBG11a, RSNK17, RSN^{+}18, ST24, Sch94b, Sch94a, SDD^{+}16, Sur01, WJ10, WUK^{+}18, WN17, WSAJ13, WLCS17, Zyt94a, Zyt94b, AS76, AMIA19, BKH^{+}06, FP14, HZK^{+}25, HB08, Kee68, Kis08, KKS12, Lia05, Olb78, PSZ^{+}07, PSBG11b, PSC^{+}07, STS^{+}13, XHW^{+}19, ZBG^{+}05, ZSW^{+}06]. Control-Flow [WJ10]. controlled [KK79, Sto07]. controller [ZXB^{+}24]. controller-based [ZXB^{+}24]. Controllers [AMH^{+}16, SDM21, CWG00]. Controlling [HSK17, BKC^{+}13]. convection [BB95]. Convention [Ano93]. Conventional [Mad69]. converged [DPW^{+}09, SJL20]. Convergence [RM03, KKK^{+}18]. Conversion [GBO87, IBM94, YTY00]. convex [SJRS^{+}13]. convolution [QBL^{+}23]. Convolutional [EVCL21, TP25]. Cookbook [Car13, Car14, G^{+}06, P^{+}08, TH10]. cooling [ARMMA18]. Cooperative [KJL11, RIP18, GLLJ16]. Coordinated [BRX13, LZ15, CRB12, HH18, KKJ^{+}13, NS07, BBMA91, MSS91]. Coordinating [LH15, ZNSL14]. Coordination [ABV12, CRG16, Tho93]. COOTS [USE99]. Copley [USE01a]. Coprocessor [LRZ16]. Copy [AGJS16, LSC^{+}17, ZCJ^{+}21, HBNK24a, HBNK24b, HDG09, LXRS19]. Copy-on-Abundant-Write [ZCJ^{+}21]. copy-on-write [LXRS19]. copying [PV08]. CORBA [GCARPC^{+}01]. Core [JYM^{+}23, KR18, RTL^{+}18, CMP^{+}07, DQR^{+}13, JAD19, KW13, PNT12, SK13b, SWH^{+}13, YTS14]. Corel [Ano03b]. Corfu [DJS^{+}17]. Corner [Sch94b, Sch94a]. correct [DM93, IM75, Kou11]. Correction [ARA20b, HBNK24a, Lee16, LCL^{+}23a, NMC18b, SHR19a]. Correspondence [BDJdS02]. Corruption [CDW^{+}24]. CoShare [WTJR22]. Cosmology [Nel04]. Cost [AMA18, AMH^{+}16, CZX^{+}19, EGG^{+}24, EVCL21, HKS19, HKM^{+}18b, LGS^{+}23, SSS25, VS19, WDL^{+}20, XLWX19, ZB20, ADA^{+}19, Dre08, KJM^{+}07, LBZ^{+}11, MMTM22, NGN24, NMC18b, NMC18a, OMB^{+}15, SJRS^{+}13, WCY^{+}17, YRJ18, ZLZ15, ZLW^{+}19a]. Cost-Aware [EVCL21, YRJ18]. Cost-Effective [VS19, HKS19, MMTM22]. Cost-Efficient [AMA18, CZX^{+}19, ZB20, OMB^{+}15]. Cost-Performance [WDL^{+}20]. Costs [ZHW^{+}17, FLL^{+}13]. count [XWX^{+}17]. counter [NB11]. Counteracting [VT14]. Coupled [WN17]. Coupling [BJPS73]. course [AL05, Don88]. courses [BBS06, GD08]. Cover [Arm98, ZXB^{+}24]. Coverage [CSS^{+}16]. Coverage-directed [CSS^{+}16]. covert [WXW15]. COVID [Cap21]. COVID-19 [Cap21]. CP [Bar73, Com82, Par72]. CP-40 [Com82]. CP-67 [Bar73, Par72]. CPS [CCL^{+}20]. CPU [ASB18, BSSS14, GKJ^{+}19, HB08, JGW^{+}11, Kam13, LWC^{+}17, MB15, PDL^{+}23, Skr01, SK13c, TSN^{+}23, VWT13, WGLL13, Yu20]. CPU-bound [TSN^{+}23]. CPU-intensive [GKJ^{+}19]. CPUs [vSMK^{+}20]. crash [KY16]. create [Fit14]. creation [CK06b, CK06e]. Credit [KP15, KCS14]. Credit-Based [KP15]. Cricket [EBLM22]. crisis [AT16]. criteria [ATS16, KA25]. Critical [Ano15, CDO24]. Criticality [WLMD16, CCDR22, CDO24, LWM14]. Crop [UBF^{+}98, BDF^{+}98]. Cross [GAHL00, GSS^{+}18, JR02, JXL^{+}12, SWF16, SKT^{+}19, WLW^{+}15, WCC16b, WBHN18, WXG^{+}25, AWR05, BKC^{+}13, PKS^{+}19, CWH^{+}14]. Cross-Architectural [JR02]. Cross-Architecture [SWF16]. cross-cloud [PKS^{+}19]. Cross-ISA [WLW^{+}15, WCC16b, CWH^{+}14]. Cross-Language [GSS^{+}18]. Cross-Layer [SKT^{+}19]. Cross-Level [WXG^{+}25]. Cross-organizational [GAHL00]. Cross-Platform [JXL^{+}12, WBHN18]. cross-run [AWR05]. cross-thread [BKC^{+}13]. crossbars [JAC^{+}19]. Crosscut [CLG^{+}10]. CrossFlow [GAHL00]. CrossOver [Ano03b]. Crowd [SML18]. Crowd-Sensing [SML18]. cryptographic [QZDJ16]. cryptography [RY10, VDO14]. CSDA [War11]. CSDP [War11]. CTO [Cre08a, Cre08b, Cre09, Cre10b, Cre10a]. cuckoo [SBI21]. CUDA [EBLM22, MGL^{+}17, PRS16]. cultural [MHM19]. Current [AH12, BDG18, RG05]. Curse [HB14, Kot10, NGRF19, Kot11]. Customer [PPO14]. Customer-oriented [PPO14]. customisable [BNS18]. Customizable [LJFS17]. Customization [PCC^{+}16, CGV10]. customized [CSMB15, HB13, UIK19]. CVA6 [SVM^{+}23]. CVM [DSC^{+}08]. cyber [PTD^{+}18, XZK^{+}20]. cyber-physical [PTD^{+}18, XZK^{+}20]. CyberGuarder [LLW^{+}12]. Cybersecurity [Ott18, ADWM18].

D [WGXC25]. DADTA [ZLCZ18]. DAI [AKK^{+}07]. Dalvik [YC16]. damn [B^{+}07]. Dana [Ano10]. Dancing [DLX^{+}17]. Daonity [CCMY07]. Dark [Fer11]. Darling [MR91]. Dartmouth [Lee86]. Dartmouth-Smalltalk [Lee86]. Data [AJ18, AT24, AAAF21, Aln22, AGC18, Att73, BFHW75, BB13, BC19, CL17a, Cap21, CGC16, CTP^{+}17, DY17, EGR15, ECET18, FML^{+}22, FL13a, GTS^{+}15, GA25, HTW^{+}19, HO22, HWR^{+}24, IEE84b, JFPL16, KP15, KAS13, LMM18, LVM16, LLWM23, Man15a, Man16, MMdE19, Nel04, PCC^{+}16, RK24, SB16, SSS25, SJMG24, SMR24, UVL^{+}13, WKJ20, WZH^{+}16, WXG^{+}25, WN17, Wel94, WTM18, XWJX15, YLH17, YWW^{+}15, ZDK^{+}22, ZHH^{+}24, ZHL16, dSdF16, vSMK^{+}20, ARA18, ARA20b, ARA20a, AKK^{+}07, AD19, AJBJ23, AGH^{+}15b, AGH^{+}15a, AHRR22b, ATZP21, AH24, ATS16, AMAB17, ARMMA18, BK14, BB12, BDE^{+}03, BOF17, CKRJ17, CLL^{+}23, CFS^{+}12, Cla05, CFRSSR19, DLH^{+}20, DXM^{+}17, FLL^{+}13, GE85, GH91a, GH20, GHK24, GSKJ18, HM20, HN08, HKB19, HTB19, HLBZ20, HUWH14, IRB19, IKU15, IPRS21, JFZL17, JYOB18, JB24, KDK20, KBDK22, KZ21]. data [KTB17, KJJ^{+}16, KSLA08, KB17, LCLG26, LDL14, LZW^{+}15, LZC^{+}16, LRP^{+}19, LMDP19, MNB24, Man15b, MAK18, MRM06, MBM09, MHM19, NZH20, NGN24, NTH^{+}17, PC21, Par25, PVRR14, PRB07, Pon19, uRQS20, QXH18, RK16, RH17, RT18, RK18, RJK^{+}17, RGS^{+}20, RMP24, SZKY21, SBI21, SHR19a, SHR19b, SBNU18, She91, SS19, SM23b, TSLBYF08, TDD20, VOS12, WKJ17, WDCL08, WZV^{+}13, WCY^{+}17, WHW20, Wol99, WTLS^{+}09, WCG14, XXZ13, XHW^{+}19, XLQL18, XXWG23, YPLZ17, YGLY21, YLJ22, ZLZ^{+}19b, ZWH^{+}17]. data-control [XHW^{+}19]. data-driven [AH24]. data-flow [GE85]. Data-Intensive [WZH^{+}16, JFZL17, QXH18, SZKY21]. Data-Oriented [ECET18]. data-parallel [She91]. DataABC [JFZL17]. Database [WK90, BBS06, CSSS11, ECAE13, MN91, MRC^{+}13, PTM^{+}15, SI81, SMA^{+}10]. databases [GDSA^{+}17]. Datacenter [BBM^{+}15, CFLL19, KGGS17, BCP^{+}08, EYG21, GTGB14, MSG^{+}12, SG10b, ZLZ15, ZWC^{+}14]. datacenter-scale [MSG^{+}12]. Datacenters [FXL^{+}23, IBBA20, JWL^{+}18, KGGS18, KL14, LGJZ16, LGJ^{+}18, LCZ^{+}19, LW20, SC17, SC18, GLLJ16, KK21, LPBB^{+}18, WRS13]. datacentre [AH24]. Dataflow [HT98]. Datapath [TSP17]. Dataplane [BPP^{+}17]. datatypes [UIK19]. DAVmS [MA21, MA24]. DBMS [AY24]. DBMS-Assisted [AY24]. DBT [KS13]. DCN [CYX^{+}17]. DDG [PGLG12]. DDG-based [PGLG12]. DDGacc [PGLG12]. de-duplication [CLcC13]. de-facto [Rus08]. dead [SK13a]. Deadline [AD18b, DQLW15, HKS19]. deadline-aware [HKS19]. deadlocks [PRB07]. dealing [BG20]. deallocator [GPS^{+}18]. Death [NOT^{+}17]. Debian [CK06a, CK06b, CK06k, CK06q, Bau06a, CK06a, CK06b, ZMD^{+}21]. Debian-based [ZMD^{+}21]. Debues [Ano03b]. Debugger [MZG14, RB01, Sun99, But94, HH05]. Debugging [ACM05a, FS12, HHH04, AKCP21, Cia07, IMBB20, JHE14, KM13a, KM13b, KK79, PMC05, THL03]. decades [IMBB20]. December [ACM05b, HHK94, IEE05, M^{+}06]. decentralised [STMV18]. Decentralized [RI25, AJBJ23, HPS23, XJR^{+}17, ZBS^{+}22]. Deception [PC24]. Decision [CHW12, DJ77, SC17, CFRSSR19, DJ76, FA21, KA25, RK18, YIR24]. decision-making [FA21]. Decisions [HKKW13]. Declarative [CRG16, Dan86]. decoding [SPAK18]. Decomposition [JK15]. decreasing [BY20]. dedicated [GLV99, KOY05]. Deduplication [Li14, MJW^{+}14, PP16, CWC^{+}14, GMK17, HOKO14, LF19, SSG^{+}20, XZZ^{+}16]. Deduplication-Based [MJW^{+}14]. Deep [Che21, HTM^{+}24, KSVR23, MSC^{+}21, ZDS^{+}22, GKT17, HcC14, MK22, QBL^{+}23]. deeply [TMJ^{+}21]. defending [CVWL13]. defense [TMV25]. Defensive [BDJdS02, Coh97]. Defined [AFG^{+}17, CL17a, CPKL17, CMK^{+}16, FXHY21, FML^{+}22, JN15, LLW^{+}16, LXZ^{+}21, MP16, Ott18, Pap20, SN23, SLZ^{+}25, SB18, SKT^{+}19, TBS17, YWH^{+}21, ZKWH17, ALW15, DS19, HHSG18, LJR12, LWL16, RB24, TK20, XJW^{+}18]. Defining [DL89, Hir17, Lot91, BMWB86]. Definition [Dom80b, SSB14b, SMO84, EMS15, SSB01]. Definitive [Chi08, Oak14]. Defragmenting [SGV13]. degradation [BJ22]. Degree [GB19, KMM13]. DejaView [LBP^{+}07]. Delay [RSNK17, RKRK17, WCY^{+}17, ZRS^{+}16, HH18, KSS24, LCL14]. Delay-cost [WCY^{+}17]. delay-sensitive [LCL14]. delegated [CLM^{+}22]. Delivery [FLZ17, MSC^{+}21, TFtLcC15]. delta [SHTE11]. Demand [CWL12, KKJ^{+}13, MSS^{+}15, MK13, SC18, SEF^{+}06, ZZF06, DEG^{+}17, J^{+}05, JCZZ13, LZW^{+}15, SGV13]. Demand-based [KKJ^{+}13, SGV13]. demands [BG20]. Demilitarized [KBBM15]. Demon [XYD^{+}18]. Demystified [COV^{+}24]. Demystify [ZYH^{+}19]. Demystifying [PS19a]. Denelcor [Dun86]. denotational [Arv02]. Denver [USE00b]. Deoptimization [KRCH14]. Dependability [FP14, VW08, YI24]. Dependable [DPCA11, SJW^{+}13]. dependences [BKC^{+}13]. Dependent [BP99, BB17]. Deploy [Kol19, XHW^{+}19, CSMB15]. deployed [RY10]. Deploying [KLLT18, R^{+}13]. Deployment [FBM^{+}21, GMGF24, LXZ^{+}21, MSC^{+}21, SFSN^{+}24, ZLZ21a, AAB^{+}05b, Bor07, CGV10, LPZ^{+}22, QBL^{+}23, SASG13, WGW^{+}18, ZLZ13, ZLV^{+}12, ZBS^{+}15]. depth [CBFH20]. derivation [MSZ09]. Derivative [Pfo13]. derived [Int06c]. Deriving [HWB03]. Description [Cre65]. Design [ACM06a, AC16, Ano03a, Ano03b, fLtNW14, ACA16, BKR20, BGS89, CLKEF21, CPS17, Clo85, DAH^{+}12, Das91, Dom80a, DAdBM^{+}24, DLS^{+}01, ESY^{+}17, GFB^{+}92, JNR12, JJ02, KGGS17, KGGS18, KBBM15, Kut92, LH16, ML78, Mar73, Mar08, MKZ^{+}25, MNN05, NBS18, OH05, PP73, PDL^{+}23, PCW^{+}16, SVM^{+}23, SIR^{+}17, SGGB99, SGGB00, SJL20, SM02, Sur01, WC01, WFL^{+}25, WCSG05, WP97, XCJ^{+}14, ZSXZ07, ZL18b, ZAI^{+}16, AM16, Blu02, BT15, Bur02, CARB10, Car14, DN14, DCA04, DNR06, GR80, HH05, HH13, JAC^{+}19, Les74, Lia05, MSCK92, MRG18, Oi05, PMC05, PNM^{+}20, Pul91, RGS^{+}20, SI81, SNV10, SMSB11, SJW^{+}13, Tur84, ZJRW19, CMP^{+}07]. Designed [HS06, DCG12, Wu13]. Designing [Par79, PM19a, PDM20, TGCF08]. DesignJet [MSCK92]. Designs [DMS02, RGSJ17]. Desktop [Ano03b, BWD^{+}15, KGG00, CCWY05, EM06, LLX^{+}17, SWW^{+}18, WH05]. Desktopping [JKB15]. desktops [KKJL14]. destination [HM20]. Destruction [NOT^{+}17]. Detecting [CL14, CJJ^{+}22, JKDC05, TV12, CWdO^{+}06, LRC05]. Detection [AJM^{+}25, CWS12, CLW^{+}14, HTAY21, JHS12, LSSC22, SXH^{+}19, SCF^{+}25, AD18a, AMA^{+}11, BSOK^{+}20, FLM^{+}08, HL24, Hui18, LF19, LMDP19, MW18, MA17, NS17, PDM20, PBYH^{+}08, SIK^{+}16, TP25, WCG14, XXZ13, ZZW^{+}24]. detection/prevention [MA17]. detectors [LMJ07]. Determine [BP99]. Determining [ZRS^{+}16]. Deterministic [KD78, RTL^{+}18, BB12, KM13a, KM13b]. dev [Fer11]. Develop [DBMI92]. developers [SS17, Wil06]. Developing [HZZ^{+}14, PCR89, RHZ^{+}17, Win71, R^{+}13]. Development [IGBKR19, Kna93, LLWM23, Lia05, LXQ24, RT93, Wil01, Bor07, But94, CWG00, Her10, IBM88, Int88, STFH15, TT93]. Device [Ano03a, Bos24, JKJ^{+}10, KKTM17, Nou92, SGB^{+}16, XYD^{+}18, FFBG08, LU04, PRS26, SBQZ14, TtLcC13, WHSE15, YWTC15]. Deviceless [HYL^{+}25]. Devices [CXLX15, HYL^{+}25, KLK^{+}22, MV16, MTJ^{+}22, RC18, SSB03, SVL01, XD16, XD17, CMGI^{+}23, CT03, DPW^{+}09, PDC^{+}12, Rus08, Wal76]. Devirtualizable [LSS04]. devirtualization [KJM^{+}07]. Devirtualizing [HHS18]. DevOps [FBL18, SCL^{+}19]. Diagnosing [MST^{+}05]. diagnosis [PPO14]. dialect [BMWB86]. Diciclo [KA24]. Diego [Ano10, IEE93a, USE99]. dienste [WF03]. Difference [GLV^{+}09, GLV^{+}10, Wal10, WBW^{+}19]. Different [Nel04, PM19b, MFT^{+}19, WCG21]. Differential [FRM^{+}24, CSS^{+}16]. Differentiated [MSS^{+}15]. diffusion [DM93, MM92]. Digital [MBK^{+}92, TLBW12, vCPWvT11]. dimensional [BSSM08, HPcC04, ZMD^{+}21]. Dinamica [FSFP19]. DINO [RSW91]. Direct [MO98, TFtLcC15, BLRC94, LC09a]. direct-mapped [BLRC94]. Directed [AJM^{+}06, CSS^{+}16, NG13, RP07]. Directions [AEA^{+}23, FLZ^{+}20, GB19, WC01, ZXR^{+}22, NIA18]. directive [CPM^{+}18]. directive-based [CPM^{+}18]. director [KMK10]. Directory [BC19]. direkt [LC09a]. Dirty [LZM^{+}20, MZD^{+}18]. Disaggregated [ASPP22]. Disaggregation [CFC^{+}24, SM23b]. Disaster [HS19, KKLV16, MTFK19, AAF^{+}09, BGS13, RCOW12, Mar08]. disaster-recovery [BGS13, Mar08]. Disclosure [WWL^{+}17a, FSH^{+}13]. Discourse [MBWW86]. discovering [FBZS12]. Discovery [HPB^{+}24, PST15b]. Discrete [GLL^{+}21, YP15, YIR24, ZGL^{+}17]. Discrete-Time [GLL^{+}21, YIR24, ZGL^{+}17]. Discretization [Bos24]. Discussion [G^{+}01]. Disk [ECET18, WWL^{+}17b, AAH^{+}03, BC10, DSSP06, DP11, WTL^{+}16]. Disk-as-a-Resource [ECET18]. diskette [Ano97a]. disks [HJ10, hTMAC^{+}08]. Dispatch [DLS^{+}01, KKC^{+}16]. Distance [GPM21, KKLV16, MA24, AJD09, MA21]. Distributed [Ano10, BBD^{+}91, BDF^{+}99, Bos24, CEPR22, CLŁS12, Das91, FXL^{+}23, FKZ17, FD08, HKLM17, IEE93a, IEE96b, JWL^{+}18, JZY^{+}22, Kim84, KMG^{+}18, KAZS14, LLW98, LS15, MLXG19, PP16, PHXL19, SSS25, SC17, SCL^{+}19, SM02, TCP^{+}17, Vol90, WB81, WIS^{+}15, WVT^{+}17, WLS^{+}18, WN17, XWH^{+}16, ZZF06, AC95, Ano96, AB16, AFT01, Bir94, EBLM22, EMI13, FS19, Fis91, FF96, FX06, Fu10, GKP^{+}19, KTB17, KJJ^{+}16, KSLA08, LC14, NS17, SJB14, SSN12, SGGB99, SGGB00, SIK^{+}16, VOS12, WKC^{+}09, XLQL18, YYC^{+}19, ZWKX17, ZWHC17, ZB18]. Distributing [HHW10, PDML25]. Distribution [Deu08, Vol90, ZKWH17, BTLNBF^{+}15b, WRS13]. Distrusting [SOAK23]. Diversity [SJS^{+}17, WGLL13, WHD^{+}09]. divisible [HM18]. DJM [LLW98]. DMVL [THH^{+}14]. DMZ [Kar07]. DNA [Ano03a]. DNN [NTL^{+}24]. DNS [See10]. Do [AZEE17, AZEE18, GW07]. Do-It-Yourself [AZEE17, AZEE18]. Docker [Boe15, BSNB20, BNS18, BRS18, DSS19, DL19a, Isl19, Ker15, MFT^{+}19, SMSH18, Spa19, WGW^{+}18, XJR^{+}17, ZMD^{+}21, ZTA^{+}21]. Does [BC10, NKY^{+}18, SMSH18]. Dolly [CSSS11]. Domain [GGM^{+}16, HHV^{+}02, KLF^{+}15, WK90, BML^{+}13, SWL^{+}23, TK20]. Domain-aware [KLF^{+}15]. Domains [PNT12]. dominance [CPST14]. Dominant [ARAAA19]. done [Han16, HUL06]. Don’t [HHPV15]. Dortmund [Müh75]. DoubleChecker [BHSB14]. Down [COV^{+}24, JJ91, PBWH^{+}12]. Downing [Ano97a]. Downloading [BC19]. Dozens [War11]. DPMI [GMR93]. drafting [MSCK92]. Drive [SYC14]. Driven [ACM05a, NSJ12, NVV^{+}24, PY93, RB17, SV13, TVO92, XRL^{+}22, YLQL25, AH24, CSSS11, DLX^{+}17, EdPG^{+}10, HZK^{+}25, ZBS^{+}22]. Driver [JXL^{+}12]. DriverGuard [CDD13]. Drivers [Chu06, JKJ^{+}10, Nou92, LU04, MSZ09]. Drives [RB24]. DRL [ZDK^{+}22]. DRL-Based [ZDK^{+}22]. DRM [WIS^{+}15]. DRP [Mar08]. DSM [JZY^{+}22]. DSM-aware [JZY^{+}22]. Dual [KPHA20, FL13b, XHW^{+}19]. dual-VM [FL13b]. Duality [FS08]. dummies [Low08]. duplication [CLcC13]. Durham [Boa90]. during [JK13]. DuVisor [CLM^{+}22]. DVFS [Kam13]. DVM [MSG^{+}12, MSG14]. Dynabook [SV24]. Dynamic [Abr80, AT24, AMAB17, BB13, BHI15, BFS^{+}18, CFC^{+}24, CJK^{+}12, DHPW01, DMG^{+}15, DHD20, FBM^{+}21, GWZ16, GSN93, HTW^{+}19, HLPY16, JWH^{+}15, KKE19, Lee16, LLWM23, LB98, LJL^{+}15, MR23, MKZ^{+}25, MMdMCPSR25, MP16, MDGS98, NMG15, PTHH14, PHXL19, QLL^{+}21, RC18, RAP19, SAJ24, SZW^{+}16, SDM21, TMLL14, TB17, TV12, Vac06, WWH^{+}16, WCS09, XSC13, XCSM18, XML^{+}18, YLT^{+}23, YLN^{+}17, ZFL15, ZWL09, ABDD^{+}91, AT23, ARMMA18, AP18, BK14, BB12, BB15, BZA12, BOF17, CSV15, CPST15, DS18, FAA17a, FAA17b, GAHL00, GPW03, HM20, HTAY21, HLW^{+}13, HB13, IRB19, JK13, JYW^{+}13, JC18, JK17, KRCH14, KJM^{+}07, LMV12, LYYY18, LJL12, MRG18, Mly09, NZH20, NTH^{+}17, PGLG12, PMP23, PBAM17, RH17, RRB17, RI25, RMP24, SHR19a, SHR19b, SSEA18, THH^{+}14, TK20, Tho73, WWL^{+}25, WRSvdM11, WRS^{+}15, Wu13, WWH^{+}17, XWW^{+}21]. dynamic [XH90, YWF09, ZXB^{+}24, vKF13]. Dynamically [MZG14, SML18, BLRC94, BDT13, FC98, HH13]. dynamically-linked [FC98]. Dynamics [MB20, YWCF15, ACT94]. dynamo [Hol95].

E-Mail [Joo06]. e-Science [SGV12]. e-server [A^{+}04]. Eagle [KS18a]. early [HLW^{+}13]. early-exit [HLW^{+}13]. Earth [BC19]. Ease [Par79]. EASY [Fli77]. eBay [Joo06]. ECI [AMA18]. ECI-Cache [AMA18]. ECMA-335 [ECM01, ECM02, ECM05, ECM06]. ecological [KSSG16]. Economic [FBL18, CSV15]. Economy [YLT^{+}23]. Economy-Oriented [YLT^{+}23]. Ecosystem [VMW^{+}19, SCC^{+}23, DMH18]. Ecosystems [NVV^{+}24, PF23]. Edge [BMJ^{+}22, BBM^{+}15, Bos24, CPS17, Cre10b, HCB18, HS21, HYL^{+}25, KSVR23, LKIL19, LGXC23, LGZ^{+}19, MSLY24, MTJ^{+}22, MVMHL24, PF23, RSNK17, RSN^{+}18, Sar16, SFSN^{+}24, WCC20, XLL^{+}20, XZL^{+}20, ZLG^{+}20, ZXG^{+}24, CMGI^{+}23, Cre10a, FGZC23, LFHS23, MB21, MPA^{+}18, MA19, SHB19, TXD^{+}24, TMJ^{+}21, VSMC23, ZLZ^{+}19a, GEK25, KA25, ZXG^{+}24]. Edge-Cloud [MSLY24, XLL^{+}20, TXD^{+}24]. edge-intelligence [MPA^{+}18]. Edge-MPQ [ZXG^{+}24]. edge/cloud [MA19]. Edition [KGG00, LYBB14]. Editorial [Sed07, WYZAD20]. Editors [FDF05, KS08b]. EDSAC [CK96]. Education [ACM06d, GPM21, AJD09, DG05, GLA^{+}08, HMS04, DTW07]. educational [WDSW01, YMY17]. Effective [LW11, LWC^{+}17, VS19, WUK^{+}18, WYZY24, HKS19, MMTM22, Sto07, WKJ15]. Effectively [UR15]. Effectiveness [ELC^{+}19, Man15b]. Effects [JK17, PLMA18, KCV11]. Efficiency [AT24, BPP^{+}17, JFPL16, KDB16, AT23, CFRSSR19, DHD20, FGG14, GHK24, GKT17, GKJ^{+}19, IPRS21, KSSG16, MDZ^{+}21, Par25, PVRR14, PBAM17, QXH18, SEPV19, WTL^{+}16, XNH21, ZYLY18]. efficiency-aware [AT23]. Efficient [AMA18, ASMA21, BYZZ20, BWH^{+}19, BHDS09, BKH^{+}06, CWL12, CWH^{+}14, CZX^{+}19, CGV10, CHPY17, DSR23, DMR10, ECJ^{+}16, EG01, GHS17, GGK19, GA25, HTW^{+}19, HB13, HYL^{+}25, JYM^{+}23, JGSE13, KA24, KJL11, KAA^{+}25, KLR^{+}20, LM99, LFHQ19, LLY^{+}25, LSX^{+}24, LZM^{+}20, MZD^{+}18, MK13, MAK18, MBBS13, NSL^{+}06, ORPS09, PP16, PWJ16, PDY^{+}23, PCC^{+}16, RSF^{+}15, RK24, RMP24, SSG^{+}20, SGK^{+}23, SYZZ^{+}14, SMR24, SHZ^{+}14, SB73, TLX17, TCTH23, WLW^{+}15, WCC^{+}16a, WXZ^{+}17, WCC20, WXG^{+}25, WGXC25, WFL^{+}25, WHD^{+}16, WTJR22, XWX^{+}17, XYD^{+}18, YP15, ZSE^{+}25, ZDK^{+}22, ZZG^{+}23, ZLL^{+}20, ZLG^{+}20, ZB20, AJBJ23, AAM^{+}16, AMAB17, BHvR05, BB12, BB15, BRIdM10, BRS^{+}22, BHSB14, BDE^{+}03, CP17a, Car14, CGM17, CFS^{+}12, DQLW15, DCP^{+}12, DCA17, EGKP02, FM90, HM20, HM18, HMH17, HKJ19, HLBZ20, HZK^{+}25, IMK^{+}13, JFZL17, KMT14, Kha19, KK21, KMG^{+}18, KR16]. efficient [KU24, LLLE17, LZC^{+}16, LYY^{+}20, LFHS23, MR23, MHM19, NTH^{+}17, NBS18, OMB^{+}15, PEL11, PM19a, uRQS20, RHR20, RT18, RZ14, RCTY19, SBI21, SENS16, SJRS^{+}13, SSN12, SM23a, SGV12, SYMA17, SLA^{+}16, SP23, SHTE11, SPSL25, VSMC23, WKJ15, WHW20, XXZ13, XLQL18, YPLZ17, YYC^{+}19, YWH^{+}23, YLK^{+}10, ZXW16, ZDK^{+}19, ZL13, ZLCZ18]. Efficiently [CWL^{+}15, EGJS15, PHXL19, BKC^{+}13]. EGO [FSFP19]. Eighth [IEE01]. EIGRP [SAR24]. einem [See08a]. Einführung [CK06a, CK06b, CK06e, CK06c, CK06d, CK06g, CK06f, CK06i, CK06h, CK06j, CK06k, CK06m, CK06l, CK06n, CK06o, CK06p, CK06q, CK06t, CK06r, CK06s]. Einsatz [Zim05]. Einsatzmöglichkeiten [Zim06]. Einsatzszenarien [Sch13a]. eJSTK [UIK19]. Elaborate [WMUW19]. Elastic [AAMBE21, AD18b, KSO^{+}15, PLMA18, BKR20, LPBB^{+}18, NAR19, TSCB19]. Elasticity [GLS15, MMdE19, OSK15]. ElasticSearch [Ben21]. ElasticSFC [TSCB19]. electricity [LBZ^{+}11]. Electronic [MSCK92, ZR06]. Electronics [GPM21, BB08]. Elektronische [Mar08]. element [FF22]. Elémentaires [Han73]. Elementary [Han73]. EleVMate [AH24]. ELI [GAH^{+}12]. ELICA [KAA^{+}25]. elimination [VED07]. elliptic [AGIS94]. Elmau [IEE01]. em88110 [VdlFCC97]. embeddable [Web10]. Embedded [BHI15, CDM^{+}25, DPCL22, DEK^{+}03, DS09a, DK23, GGM^{+}16, GCL^{+}21, JAD19, JYM^{+}23, Kut92, MKZ^{+}25, Mon97, MW25, NKK^{+}06, PPG^{+}17, SMK02, SMP22, WLW^{+}15, AH12, Caa00, CT03, CGV10, HK07, Ive03, KKC^{+}16, LSS24, LTK17, MBBS13, RJK16, RMB02, TMJ^{+}21, UIK19]. Embedded-System [Kut92]. Embedding [AM16, BL17, Che21, EMW16, FXL^{+}23, LGS^{+}23, OMB^{+}15, PHXL19, PHC20, YLH17, AO16, BG20, BCC^{+}15, CRB12, EMI13, HKB19, HH18, JK15, KKM^{+}13, NTH^{+}17, OKAM17, SS19, SZL^{+}14, TK20, WHC16, WBW^{+}19, WZZ^{+}20]. Embeddings [RS20]. EMF [WIDP12]. Emission [SSS25]. Emphasis [Cre65]. EmuID [CJJ^{+}22]. emulate [tTR82]. Emulated [PC24, THC^{+}14]. emulating [VdlFCC97]. Emulation [Ano03a, BKMM87, CLKEF21, JN15, KKTM17, Mal72, BB08, CWH^{+}14, CJJ^{+}22, GD08, Kam13, YJZY12, Bro89]. emulations [Bod88]. Emulator [Ano14b, Bru07, CFH^{+}79, CFH^{+}80, CK87, FS11, MZG14, WCC16b, Bar06, KS13, Les74, She02]. Emulators [Ert03, HHC^{+}16, Mal73, Ert05, GFHML^{+}25]. Enable [XD17, HPS23, TMJ^{+}21]. Enabled [HTB22, LXZ^{+}21, MKZ^{+}25, SB18, DMH18, HTB19, KS20a, KZ21, RI25, SGV12, TUM18, VOS12, YKP^{+}24]. enabler [DPW^{+}09]. Enabling [AEA^{+}23, HD16, HS21, KMK10, NOT^{+}17, OVI^{+}12, SLZ^{+}25, Spa19, TY14, WHD^{+}16, ZZG^{+}23, FMX^{+}25, LSS04, ZBS^{+}22]. enCloud [YKP^{+}24]. encoder [KA25, TP25]. encoding [BDE^{+}03, SPAK18]. Encrypted [HB17]. Encrypting [Pro00]. Encryption [SXH^{+}19]. End [Ram93, SS17]. end-users [SS17]. Endurance [AMA18]. Energy [ADA^{+}19, AGC18, AAK18, BWD^{+}15, CWL12, CP17a, DMR10, DQLW15, Do11, DCMW17, EGR15, FML^{+}22, FLZ17, GA25, HTW^{+}19, HKM^{+}18b, IRB19, JJK^{+}11, JFPL16, JYM^{+}23, KC16, KSS^{+}20, KB21, KSS^{+}23, KDB16, KCS14, KL14, LMM18, LZC^{+}16, LYYY18, LGJ^{+}18, LYY^{+}20, LWCZ22, LFHS23, MDZ^{+}21, OBSR16, PHC20, RK16, RH17, RK24, SBNU18, SSS25, SYMA17, SZL^{+}14, SP23, TDD20, WZH^{+}16, WGXC25, XLWX19, YLK^{+}10, YRJ18, ZDK^{+}22, ZWC^{+}19, ZHL16, AJBJ23, AMAB17, ARMMA18, BAC15, BB12, BB15, BRIdM10, BJG19, BRS^{+}22, CD14, CFRSSR19, DP11, DHD20, DXM^{+}17, FAA17a, FAA17b, FFB^{+}00, GLK^{+}12, GTN^{+}06, GKJ^{+}19, HM20, HM18, HLBZ20, JWH^{+}15, JFZL17, JC18, KMT14, KTB17, KR16, LCLG26, LJYZ15, MR23, DPBK16, MHM19, NTH^{+}17, NBS18, dOL12, Par25, PVRR14, PTD^{+}18, QXH18, RHR20, RP07, RT18, RCTY19, SBI21, SENS16]. energy [SMSH18, SHR19a, SHR19b, SPSL25, THG^{+}18, VW08, WWL^{+}25, WDT18, WHW20, XNH21, XZK^{+}20, YPLZ17, YW20, YWH^{+}23, YLJ22, YB24, ZLCZ18, ZYLY18, ZSRR22, RNA^{+}22]. Energy- [WZH^{+}16]. Energy-Aware [AAK18, Do11, EGR15, LMM18, PHC20, XLWX19, ADA^{+}19, DCMW17, KC16, KB21, LYYY18, LWCZ22, RH17, SZL^{+}14, ZWC^{+}19, CD14, DXM^{+}17, GLK^{+}12, JC18, KCS14, MHM19, SHR19a, SHR19b, WDT18, YB24, ZSRR22]. Energy-Awareness [ZHL16]. Energy-credit [KCS14]. Energy-Efficiency [JFPL16, XNH21]. Energy-Efficient [DMR10, GA25, HTW^{+}19, JYM^{+}23, ZDK^{+}22, CP17a, LZC^{+}16, LYY^{+}20, LFHS23, SYMA17, SP23, YLK^{+}10, AJBJ23, BB15, BRIdM10, HM20, HM18, HLBZ20, JFZL17, MR23, NTH^{+}17, NBS18, RHR20, RCTY19, SBI21, SPSL25, WHW20, YPLZ17, YWH^{+}23]. Energy-Oriented [BWD^{+}15]. energy-performance [XZK^{+}20]. energy-saving [YLJ22]. Enforcement [LJFS17, NMMP15]. Enforcing [KC12, WZL15]. ’Engine [Wal10, GLV^{+}09, MO98, VG20, GLV^{+}10, J^{+}05, MIS^{+}05]. Engineering [GPM21, IEE84b, SDS^{+}21, ACM01a, MKM^{+}08, McG72, MPM^{+}20, WZV^{+}13]. Enhance [GLS15, CMGI^{+}23, MK19]. enhanced [MK23, SM23a, SDN09]. enhancement [DXM^{+}17, KS18a]. enhancements [AKK^{+}07]. Enhancing [CPKL17, GI12, GHK24, HWR^{+}24, PRS26, ZXB^{+}24]. ENIAC [ZR06]. Enlightened [AGJS16]. ensemble [RGAT18]. ensuring [Req03].

Enterprise [ADG^{+}92, CMK^{+}16, FPR^{+}06, G^{+}06, LVM16, BSNB20, EM06, Hal08, NS07, WH05, Ano03a, Gal11]. enterprises [GAHL00]. enthüllt [Joo06]. ENTICE [GKP^{+}19, HKM^{+}18a]. Entities [ZLG^{+}20]. Entity [LGZ^{+}19]. Entrepreneur [War11]. Entropia [CCWY05]. Entropy [TVO92]. Entropy-Driven [TVO92]. EnTruVe [RNA^{+}22]. enumeration [SSH17]. Environment [ACL72, BGM70, CL16b, GKSP99, Gen86, GGG03, HW93, IEE06a, J^{+}05, JADAD06a, LWC^{+}17, LW12, Mac79, MB15, RT93, SAJ24, TMV12, XSC13, XLL^{+}20, ZD18, AAB^{+}05b, BRS^{+}22, BH13, CLDA07, CWG00, DL19a, Don87, DS22, FCD09, FAA17a, GSG22, GD08, GMR93, GNK24, Hal09, HL13, JWH^{+}15, JXZ^{+}10, JADAD06b, KW13, KA25, KKK^{+}18, KMG^{+}18, KU24, LJYZ15, LPZ^{+}22, McG72, MST^{+}05, MW18, MPF^{+}06, NNK21, NS17, PM19a, RGAT18, RG19, RAP19, SPSL25, TMLL14, TP25, TT93, TV18, Van06, WLL^{+}13, XZZ^{+}16, YWH^{+}23, Yu20, ZBP05, ZFL^{+}23, ZZW^{+}24, ZLLL13, FAA17b]. Environments [ACM05d, ACM06f, AD18b, BB17, BMJ^{+}22, BE17, CWL12, CGMD19, GKXK13, HHW10, HKKW13, HYL^{+}25, KKH14, KGZ^{+}04, LH15, NKY^{+}18, PGP19, PWJ16, PLZ20, RIP18, RGSJ17, SV13, SKT^{+}19, TDMP23, XLWX19, ZWFX17, ZZF06, ARA18, ADA^{+}19, ATS14, BCC^{+}15, BRIdM10, BDK^{+}08, CFVP12, DP11, DS18, DEG^{+}17, FMIF18, GPS^{+}18, GMK17, GGK19, GEK25, HOKO14, HC12, dCJR16, KSO^{+}15, KKB14, LC14, PSZ^{+}07, PST15b, PRS26, SSG^{+}20, SJW^{+}13, SGV12, SHB19, TRG13, VDO14, WWWL13, WTL^{+}16, XHL^{+}13, YLK^{+}10]. Ephemeral [WHD^{+}16]. equilibrium [uRQS20]. equivalent [TLX17]. Erasure [ZLL^{+}20]. Erasure-Coded [ZLL^{+}20]. Erlang [TCP^{+}17]. Erratum [FAA17b]. Error [XH16, XHL^{+}13]. errors [AMIA19]. Ersatz [Hin08]. erstellen [Zim06]. Erstellung [See08a]. ESA [Fis91, GH91a, IBM94, MSS91, OJG91, SNC91]. ESA/390 [OJG91]. ESA/XC [GH91a]. Escape [WLCS17]. Escapers [SXH^{+}19]. eServer [R^{+}02, G^{+}05]. Espresso [WZL^{+}18]. ESPRIT [RD90]. Essentiality [FXL^{+}23]. Essentials [SNS03, MBM09, VSC^{+}10]. Estimation [DSM14, HSK17, KSSG16, NKY^{+}18, OBSR16, SJMG24, LBL16, MPA^{+}18, VWT13, WDT18]. ESX [AAH^{+}03, D^{+}04, MWHH05, OH05, Rul07, R^{+}02, Zim05, Hal08, MBM09, Wal02]. ESXi [GKBB15]. ET6 [Pul91]. ET6/1 [Pul91]. ETAS [IRB19]. Ethereum [FRM^{+}24, HLW^{+}23, Hir17]. Ethernet [YCL^{+}18]. ETICA [ASMA21]. Eucalyptus [AMA^{+}14]. European [ACM04a]. EUROTRA [Pul91]. Evaluating [Ben21, De 06, GLK^{+}12, GC00, HH19, HW93, MTJ^{+}22, RCM^{+}12, BJ22, SMSH18]. Evaluation [AD11, Bos24, CFH^{+}79, CFH^{+}80, DAH^{+}12, HB12, KD78, MG19, MMHP17, PZW^{+}07, RNS^{+}23, SJA^{+}17, SHB^{+}03, SP22, SHTE11, TMV25, TFtLcC15, VMBM12, ZZW22, ACM06c, ALW15, DSSP06, FSH^{+}13, GE85, HTB19, JFZL17, dCJR16, Kao17, Kee68, MCC18, Man18, NMC18b, NMC18a, SLC20, TUM18, VW08, WKT08, WWH^{+}17, YZW^{+}13, Hin08]. evaluations [SJW^{+}13]. Even [KBK^{+}21]. Event [DLX^{+}17, MV16, YP15, ZBS^{+}22]. Event-driven [DLX^{+}17, ZBS^{+}22]. events [LC13]. Everything [NBB^{+}19]. everywhere [Tre05]. Eviction [AGJS16]. Evictions [vSMK^{+}20]. Evil [HCJ07]. EVMFuzz [FRM^{+}24]. Evolution [BG73a, HH79, Kim84, SLM89, SL16, AGSS10, CD01, GBCW00, Kro09, WIDP12]. evolutionary [LKR^{+}19, LWCZ22]. Evolutions [BAL15]. evolving [Ano96, FF96]. Exact [WHW20, EYGS19]. examination [HN08]. Examining [NL00]. Examples [Gol71b]. Exascale [NLD^{+}23]. exceeding [GHS16]. Excelsior [MLG^{+}02]. exception [CGC^{+}24, Sal92]. Exceptionization [YKM17]. exceptions [Ven97b]. Excessive [IYAK23]. Exchange [SMR24]. exclusion [SGS92]. Executable [MP01, ST24]. Executables [LKL^{+}19, AD18a]. executing [ACT94, Lot91]. Execution [ACM05d, ACM06f, CGMD19, CDW^{+}24, FZD^{+}24, HWB03, KGZ^{+}04, LWC^{+}17, MM93, MKZ^{+}25, MO98, PY93, RT93, SV13, ZLSI17, vLSM01, AS76, AAB^{+}05b, BSD19, BFC02, BDK^{+}08, CLDA07, EBLM22, Fre05, GCARPC^{+}01, GK05, dCJR16, MMP^{+}12, OJG91, SM01, TT93, TV18, ZBS^{+}22, ZL13]. Execution-Driven [PY93]. executions [KM13a, KM13b]. Exercise [Lee86]. Exhaustive [PM19b]. existential [AT16]. Existing [JMSLM92, LTT92]. exit [HLW^{+}13]. exitless [AGH^{+}16]. exokernel [Cof99]. Expansion [Par79]. Expelliarmus [SBBP20]. Experience [San88, RM03, CARB10, CBLFD12, FDD^{+}19, PBAM17, RSC^{+}15, TGCF08]. Experiences [FBGS24, NV05, SCD90, Tsa14, CMP^{+}07]. experiment [HA79]. Experimental [Bro89, ACM06c, FSH^{+}13, HL13, SS72]. experimentation [ACG18]. Experimenting [Taf11]. experiments [Ker88]. Expert [Hee07]. ExpEther [NMS^{+}14]. Explaining [YYL^{+}15]. Explicit [WUK^{+}18]. Exploit [GLL^{+}21, RAP19]. Exploitation [SSMGD10]. Exploiting [CRZH15, EdPG^{+}10, GLS15, MPA^{+}18, MFT^{+}19, PA21, SJS^{+}17, YTS14, WTLS^{+}09]. Exploration [SVM^{+}23]. explorative [AHK^{+}15]. explore [Fit14]. Exploring [CLKEF21, CPM^{+}18, LSSC22, SE12, SIdLB15, WKJ20, YBZ^{+}15]. Expo [Ano06a]. Express [Ng01a, Ng01b]. Expression [Cox07, Cox09, Cox10, Cox12, Wat86, Wat87, Tho68]. Expressions [KP99]. extend [MK19]. Extended [DC15, FBGS24, Gum83, MT16, MT17, Olb78, You73, IBM88]. Extending [CT03, DLM^{+}06, PTHH14, YTY00]. Extensible [FLCB10, TSP17, DCA04, YJZY12]. extension [DCP^{+}12]. Extensions [Fis01, SCP93]. EXTERIOR [FL13b]. External [AA18, LHW^{+}20, FL13b]. extraction [KA25, WML02]. ExtraV [LKY^{+}17]. ExtraVirt [LRC05]. extreme [NOR15]. EXUS [SKC73]. eye [Guy14].

FaaS [CZH^{+}25, GEK25]. Fabric [ZL18a]. FACADE [GLV99]. faceted [BDQR23]. FACILE [GMP89]. Facilitating [cCWS14, SWcCM12]. Facilities [Gum83, GH91a, MN91]. Facility [McK04, MLA83, SM90, IBM72, IBM73, IBM76b, McC74, Olb78, SZ88]. Facility/370 [IBM72, IBM73, IBM76b, McC74, Olb78]. facto [Rus08]. Factor [SC18]. factors [BJ20]. Fad [Fra98]. Failover [SAR24]. Failure [Fu10, KLR^{+}20, MSI^{+}12, ZWH^{+}17]. Failure-aware [Fu10, ZWH^{+}17].

Failures [YYL^{+}15, PBYH^{+}08]. Fair [AIAR^{+}24, CL15, CFLL19, GLLJ16, HSN17a, KNHH18, TTH^{+}19, RZ14]. Fair-share [KNHH18]. FairGV [HSN17a]. Fairness [SWH^{+}13, SKJ^{+}17, HZK^{+}25]. Falle [Mar08]. familiarized [Ame13]. Farms [Do11]. Fast [CSS^{+}13, CLW^{+}14, Cox07, CHPY17, Hol95, HSN17a, JFPL16, Kou11, KLK^{+}22, NOT^{+}17, PEL11, ZLW^{+}14, ZFY18, ZLZ^{+}19b, ZLW^{+}19b, JFZL17, KMMV14, KJLY15, LKR^{+}19, MSZ09, SK13b, SV15]. Fast-Spreading [CLW^{+}14]. FastDesk [SWW^{+}18]. FASTSCALE [LKR^{+}19]. FAStT [D^{+}04]. Fault [FK03, JKJ^{+}10, Kim84, RZPX19, XXWG23, YWR^{+}14, YYL^{+}15, ZJXL11, SNV10, SM23a, YLH14, YI24]. Fault-Tolerant [FK03, Kim84, YWR^{+}14, SNV10]. faults [CP17b, LRC05]. FCP [SAB^{+}07]. Fe [ACM00]. feather [YGN^{+}06]. feather-weight [YGN^{+}06]. Feature [TP25, Bag76]. Features [Gal11, MB21, McC74, Bau06b, Bau06a, IT86, LPZ^{+}22]. features-based [LPZ^{+}22]. featuring [Wil06]. FEBI [HLW^{+}23]. February [Ano10, USE01b]. Federated [HTM^{+}24, MMdMCPSR25, RNS^{+}23, AO16, CFVP12, dCCDFdO15, KMG^{+}18]. federation [LWLL16]. Fedora [HH08]. feedback [NG13, ZBG^{+}05]. feedback-control [ZBG^{+}05]. feedback-directed [NG13]. feeding [NGN24]. FF [LSPP^{+}23]. FF-A [LSPP^{+}23]. FGP [FG91]. FHPCN [M^{+}06]. Fiber [GDSA^{+}17]. Fiber-based [GDSA^{+}17]. Fidelity [KKTM17]. Field [BBM^{+}15, KNT02]. Fifth [ACM75, IEE96b, USE99, IEE04]. File [AEMWC^{+}12, AvMT11, Li14, Neu92, SNC91, ZCJ^{+}21, ZZF06, FFBG08, HC12, Int06c, JXZ^{+}10, SBQZ14, TZY^{+}25, Vag10, WH08, WF07]. files [LLF^{+}18]. filesystem [ZYZ^{+}18]. filling [HUWH14]. film [SL00]. filtering [MG19]. FIMCE [ZD18]. final [Pul91]. find [Fab13]. finding [Bod88]. Fine [BSSS14, CHW12, CDD13, HSK17, JCZZ13, PG11, RB17, SCF^{+}25, YGLY21, YSS^{+}17, KWZ^{+}19, WJGA12, YTS14, YSM^{+}21]. fine-grain [WJGA12]. Fine-Grained [BSSS14, CHW12, CDD13, HSK17, RB17, SCF^{+}25, YSS^{+}17, JCZZ13, PG11, KWZ^{+}19, YTS14, YSM^{+}21]. Finite [SC17, GLW23]. Finite-Markov [SC17]. Firefly [KC16, SM23a]. Firefox [Joo06]. Firewall [TMV12, DS18, JES^{+}15]. firmware [ABB^{+}15, MSCK92]. First [ACM05d, IEE84b, LCWB^{+}11, MNS^{+}14, SMP22, ZR06, SS17, SHB^{+}03]. First-Class [LCWB^{+}11, SS17]. Fit [NKY^{+}18, BY20, LWB13]. Fixed [Lam75, Bod88]. FL [PF23]. Flash [SYC14, Pat12, ZXB^{+}24]. Flash-based [SYC14]. flaws [Ano07]. flex [Kag09]. FlexHM [PDY^{+}23]. Flexibilities [LS15]. Flexibility [BSI^{+}15, FPS^{+}02]. Flexibilizing [BG20]. Flexible [AvMT11, CGMD19, KA24, KWZ^{+}19, KS20b, KLK^{+}22, LZW^{+}17, LWB13, PDY^{+}23, vMAT14, AHRR22b, ACG18, CARB10, CCL^{+}17, TGCF08]. Flock [ZWC^{+}23]. Flow [FML^{+}22, MSLY24, WJ10, BSD19, BK14, BKH^{+}06, FLL^{+}13, GE85, RJK^{+}17, TK20, YKS16]. Flows [CDD13]. fluctuation [WWL^{+}25]. Fluid [MB20]. Flux [SML18]. Fly [MKZ^{+}25, AH24, URJ18]. Focused [SFSN^{+}24, BDG18]. Fog [NBS18, KZ21, MMTM22, RNA^{+}22, GEK25]. fog-cloud [MMTM22]. fog-enabled [KZ21]. folding [CPST14, Oi06]. Forecast [CWL12, TMLL14, VGL23]. Forecasting [PCW^{+}16, CBJ22, KSSG16]. Forensics [HN08, ZXY^{+}15]. Foreshadow [VMW^{+}19]. Formal [BDJdS02, BN75, CH78, Dom80b, JE12, Jen79, MP01, PG73, PG74, Qia99, YI24]. Formalism [UOKT84, Pul91]. Formalizing [HM01]. formation [HLW^{+}13]. FORSETI [CSV15]. FORTH [Mar81, Kna93, Ode87]. FORTRAN [IBM88, Int88]. Forum [CS76, DM76, Fra83, GHF83a, GHF83b, WNL^{+}83, DHPW01, GPW03]. Forward [Uhl06, YK13]. found [Ano97b]. foundation [OJG91]. Foundations [Hog08, HMS17]. Four [QNC07]. Fourth [Ano03b, MS91b]. Fourth-Generation [Ano03b]. FP [JFPL16]. FP-ABC [JFPL16]. FPGA [GP13, QTR21, ZDS^{+}22]. Fractional [PS23]. Fragmentation [GWZ16, HKM^{+}18a]. Frame [WH99]. Framework [DY17, GCL^{+}21, GH91b, JXL^{+}12, KCWH14, KAJW93, LGZ^{+}19, LWLL10, LWB13, MKZ^{+}25, MGL^{+}17, PC24, PXG^{+}17, PST^{+}15a, PLZ20, RNS^{+}23, SLW^{+}24, SZW^{+}16, SEK^{+}19, SCA^{+}24, TMV12, WGW^{+}18, XWH^{+}16, YWH^{+}21, ZFL15, ZWFX17, ZDS^{+}22, Ame13, AC16, BB15, BDE^{+}03, CD14, DS20, DLH^{+}20, FPGK18, FMJ15, Fre05, JSK^{+}13, Kag09, Kao17, KKM^{+}13, KJJ^{+}16, LLLE17, NB11, PM19a, PDM20, PV06, RH17, RSC^{+}15, RK18, RMP24, SJRS^{+}13, SSEA18, SL00, SIK^{+}16, STY^{+}14, WHC16, YWTC15, ZXW16, ZS01, ZSR^{+}05]. Frameworks [AP22, WXG^{+}25, ZLW18, AGH^{+}15b, HZZ^{+}14]. France [ACM90, ACM05b, Jou85, JPTE94]. Francisco [ACM06a, USE02]. Free [Ano03a, BRX13, SAM14, AJBJ23]. FreeBSD [McK04, MNN05, Sar01]. FreeDOS [WF03]. French [Apr09, AH68, Han73]. frequency [Kam13, SSEA18, AMAB17]. fresh [DSR23]. Friendly [ZBG^{+}05]. Front [Ram93]. Frontier [Sar16, Rob12]. Frontiers [ACM06e, M^{+}06]. Full [HHC^{+}16, HSL17, MZD^{+}18, MCE^{+}02, PGP19, Sch13b, SWF16, JK17, LLY^{+}18, YKS16]. Full-System [SWF16]. Fully [CGMD19, KBBM15, ZD18]. Function [AP22, Aln22, CEPR22, Che21, EMAL17, ELC^{+}19, FBM^{+}21, FR24, FLZ17, GGK18, HTAY21, HSL17, HYL^{+}25, JW17, KLR^{+}20, LGXC23, LLW^{+}16, LYL21, MLXG19, MDZ^{+}21, MSLY24, MSC^{+}21, Pap20, PHXL19, RKRK17, SN23, SLZ^{+}25, WCC20, XWW^{+}21, YWL^{+}18, ZZG^{+}23, ZSP^{+}21, ZLZ^{+}21b, ZWC^{+}23, ZKWH17, ALW15, BCC^{+}15, DS18, FZS^{+}20, HLW^{+}23, LMDP19, MCJ19, MHS21, QBL^{+}23, SHB19, SOKE23, TSR19, TP25, TSCB19, XHW^{+}19, YLTF20, YXL^{+}20, ZJRW19]. Function-as-a-Service [ZWC^{+}23]. function-virtualized [DS18]. Functional [ACM90, Dan86, DCG12, GMP89, Ame13, Wak99, Jou85]. functionality [MK19]. Functions [BYZZ20, BCZ19, CFC^{+}24, DL89, HTM^{+}24, KLLT18, MP16, NGRF19, NVV^{+}24, TF16, WZL^{+}23, DS19, FJKK17, HHSG18, HH19, KWZ^{+}19, LRP^{+}19, PMP23, PJZ^{+}19, PFNC20, QZDJ16, TSCB19, YCL^{+}19, ZGL^{+}17, CBJ22, GHM^{+}18, TSN^{+}23]. fundamental [BCZ19]. funfte [Müh75]. funnel [LMV12]. Fusion [Kis08]. Future [FLZ^{+}20, GB19, Her06, IBBA20, KS08b, LCMV17, RG05, Sup04, Var91, ZXR^{+}22, AH12, Bau05, NIA18, PTD^{+}18, Ros14, Str13, Yur02, SIJPP11]. fuzz [FRM^{+}24]. Fuzzing [KLF^{+}15, LSX^{+}24]. Fuzzy [AAR22, BY20, Hu90, LZ15, CFRSSR19, FA21, FLM^{+}08, SENS16, ZB18]. Fuzzy-logic-based [BY20]. FWNs [SIJPP11].

G [ALW15, HH18]. GA [HMH17]. Game [SAJ24, FK13, GLLJ16, LWCZ22, NS17]. Game-Based [SAJ24]. games [WKC^{+}09]. Gaming [CZX^{+}19, ZQCZ16, CZX^{+}19]. Gap [DGLZ^{+}11, FL13a, GSW^{+}17, ZLHD15]. gaps [HUWH14]. Garbage [ADM98, DS16, GTS^{+}15, HPHV17, PBAM17, Sch13a, SHB^{+}03, URJ18, WXG^{+}25, BOF17, DEE^{+}16]. Gast [WF03]. Gast-Systeme [WF03]. Gateway [CCO^{+}05]. Gateways [DW14]. gather [Wol99]. Gb [YCL^{+}18]. GC [HHPV15, SEPV19]. GC-Wise [SEPV19]. GCompris [CK06t, CK06r, CK06s, CK06q]. GCTrees [DS16]. GDB [MZG14]. gehärteten [See08a]. Geiger [JADAD06b]. Gelato [Ano06a]. Gene [SSU^{+}12]. Gene/P [SSU^{+}12]. General [Cre65, GFB^{+}92, XWH^{+}16, ZDS^{+}22, BJ22, BDE^{+}03, LSS04, SS72]. General-Purpose [GFB^{+}92, ZDS^{+}22]. Générateurs [Han73]. Generating [SCF^{+}25]. Generation [AEA^{+}23, Ano03b, AC98, BDF^{+}99, CF00, GFH82, MZG14, PG74, EL98, IIK^{+}06, LLS^{+}12, PG73, RGS^{+}20, Sus76, Web10]. generational [WK08]. generations [BOF17]. Generator [Han73, ABDD^{+}91, EGKP02]. Generators [Fra83, GHF83a, GHF83b, WNL^{+}83]. Generic [MM94, ZLZ^{+}21b, BKT^{+}19]. generics [Int06a]. Genetic [AAR22, MPM^{+}20, PC21]. Geo [JWL^{+}18, PHXL19, XLQL18]. Geo-Distributed [JWL^{+}18, PHXL19, XLQL18]. geographically [KTB17, ZB18]. geometry [Hol95]. George [ACM03b]. Georgia [USE86, USE00a]. German [Joo09, Bec09, Bod10, CK06a, CK06b, CK06e, Fis09, Lar09, Sch13a, Spr07, WR07]. Germany [RM03, GHH^{+}93, IEE01]. get [Ame13]. gets [Rou07]. Ghost [Arc07]. GI [Müh75]. Giants [FS12]. GiantVM [JZY^{+}22]. GINI [YMY17]. GKLEE [LLS^{+}12]. Glass [LHW^{+}20]. Global [LLW98, Neu92, Sta97]. GloudSim [DC15]. gMig [LZM^{+}20, MZD^{+}18]. GNAT [CDG97, MB98, Shi03]. go [BWH^{+}19, LWB^{+}15]. goes [RY10]. going [McK11]. good [RY10]. Goodness [BJC24]. Google [Cox12, Joo06, DC15]. Goto [Abr80]. GPGPU [CPM^{+}18, KLY20, MMG^{+}18, TY14]. GPOS [JK17]. GPU [CMGI^{+}23, DS09b, GMK17, HSN17a, HSN17b, IPRS21, KLY20, KSVR23, LYGG20, MZD^{+}18, MTFK19, MNS^{+}14, MGL^{+}17, NRdA^{+}20, NMS^{+}14, PGP19, PS19b, PRS26, RSC^{+}15, RS16, SCSL12, SM23b, SIRP17, SPAK18, SKYK16, TTH^{+}19, XML^{+}18, YLWH14, YCL^{+}18, YML^{+}18, YLT^{+}23, YSS^{+}17, ZZG^{+}23]. GPU-Accelerated [MTFK19, SCSL12, SPAK18]. GPU-assisted [GMK17]. GPU-Job [PS19b]. GPUDirect [YWCF15]. GPUs [LLS^{+}12]. GPUvm [SKYK16]. gQoS [LYGG20]. GRACE [M^{+}06]. gradient [MM92]. Gradual [RSF^{+}15, RAT17]. grain [WJGA12]. Grained [BSSS14, CHW12, CDD13, HSK17, RB17, SCF^{+}25, YSS^{+}17, JCZZ13, KWZ^{+}19, PG11, YTS14, YSM^{+}21]. grammar [FS89]. Grande [ACM01b, DHPW01, GPW03]. Grande/ISCOPE [ACM01b]. Granularity [PXG^{+}17, RRB19, LLS14, YGLY21]. Graph [CFM17, CRG16, LKY^{+}17, QBL^{+}23, Syr07, YTS14]. graph-based [CRG16]. graphic [Wal76]. graphic-simulator [Ber86]. graphical [Bur02]. Graphics [Ano03b, JXL^{+}12, PGP19, VLZL16, XML^{+}18, ME87, Sus76]. Graphs [Lee16, Bod88, PULO16]. gray [WSVY09]. gray-box [WSVY09]. Greedy [NMG15]. Green [KL14, MZ20, SLZ^{+}25, LLW^{+}12, LJL12, WZV^{+}13, XA22, YLHJ14]. Green-Energy-Aware [KL14]. Greene [War11]. Greener [BH13]. Grenoble [ACM05b, JPTE94]. Grid [ACM05b, CCMY07, IEE04, MFT^{+}19, SEF^{+}06, TLC06, ZZF06, vLSM01, Rob06, SJW^{+}13, SGV12, ZBP05, AKK^{+}07, CCO^{+}05, KGZ^{+}04, LP14, WKT08, ZBP07]. Grid-Based [vLSM01]. GridGIS [M^{+}06]. Grids [GPM21, CCWY05, MPA^{+}18, GTN^{+}06]. Group [Boa90, Sof83, YLN^{+}17, CKP78, KKK^{+}18, ZLH^{+}15]. Grouping [AAR22]. growth [LDL14]. GSX [Zim05]. GTP [M^{+}06]. Guarantee [LZ15, CMG^{+}19, MDZ^{+}21]. Guaranteed [LWZ^{+}18, ZWL^{+}18, KB21]. Guaranteeing [LZW^{+}15, YWR^{+}14, ZRS^{+}16]. guarantees [MSG01, ZHCB15]. Guest [CCML12, NOT^{+}17, ABG14, FL13b, JXZ^{+}10, LD11, MSZ09, XHCL15, FDF05, KS08b]. Guest-Assisted [CCML12]. guest-OS [FL13b]. guest-transparent [JXZ^{+}10]. guests [JK17]. GUI [PW03]. guidance [JSK^{+}13]. Guide [Ame13, BBD^{+}91, Bas04, Bas06, Chi08, Gal09a, IBM72, IBM73, IBM76a, Oak14, OH05, IBM88, Int88, IBM94, KSS09, KS10, MDD^{+}08, MIS^{+}05, RR09, TC10, War02, Wes98]. guided [HLW^{+}13, SSH17]. Guiltiness [PJZ^{+}19]. GVirtuS [MGL^{+}17]. gVMP [SM23b]. GWO [Par25].

H [JAS^{+}15, Wel02]. H-SVM [JAS^{+}15]. HA-VMSI [ZTWM17]. Hack [WMUW19]. hacking [Spi06]. Hadoop [GLBJ18, ZRD^{+}15]. Handbook [Bod10, Fis09, NSHW10, War05, Joo09]. Handbuch [Joo06, WF03, Bod10, Fis09, Joo09]. handler [Sal92]. handles [Ven97b, Ven97c]. Handling [AMB^{+}17, SB16, SMA18, CGC^{+}24]. Hands [Kol19, MDD^{+}08]. Hands-on [Kol19, MDD^{+}08]. Harbour [MR91]. Hard [JYM^{+}23, LTK17]. Hardness [RS20]. Hardware [AE01, BKM25, CWS12, Cla97, Gol71a, HHV^{+}02, HWF07, Hsu01, JAD19, JSHM15, JAS^{+}15, KAJW93, KLK^{+}22, LH16, LZW^{+}17, Mac79, NSL^{+}06, OT97, PvDS08, RTL^{+}18, SLW^{+}24, SYB12, SWF16, WCS06, YVCB17, YVCB18, ZTWM17, vD06, AA06, AJH12, AEB19, DJP^{+}24, BHDS09, CBGM12, CP17b, FP14, HH13, HP77, KW13, KJM^{+}07, Oi05, Oi06, Oi08, Olb78, PGLG12, PBB13, RPE12, SE12, TO96, WZW^{+}11, XZ11, YJZY12, ZDK^{+}19]. Hardware-Accelerated [SWF16]. Hardware-Assisted [JSHM15, JAS^{+}15, RTL^{+}18, AJH12]. Hardware-Based [PvDS08, KLK^{+}22, KJM^{+}07]. hardware-bound [DJP^{+}24]. hardware-translation [Oi06, Oi08]. Hardware/Software [KAJW93, LH16, HH13, HP77, WZW^{+}11]. Harmful [NMHS15, WC01]. Harmony [PPS^{+}18]. HARNESS [BDF^{+}99, GKSP99, MDGS98]. harnessing [GLV^{+}10]. hash [SV15]. hash-array [SV15]. Hawaii [MS91b, Shr89]. HBench [ZS01]. header [VED07]. Healing [BHI15, GK05]. Health [ZL16, ZL18b]. Healthcare [AAR22, KS20a]. Healthcare-Cloud [AAR22]. heap [CSV15, CH08, LDL14, LLS^{+}08, PNM^{+}20, TLX17, WSAJ13]. Heavy [HS19]. hedging [RY10]. Helix [Ano03a]. help [Car14, Men03]. HEP [Dun86]. herd [BB20, KS18a]. Hermes [ZLG^{+}20]. hesitant [FA21]. Heterogeneity [GLS15, KR16, XLJ16, AMB^{+}17, WCS09]. Heterogeneous [AJM^{+}25, GKSP99, HSK17, HHS18, HWCH16, HYL^{+}25, KGGS17, KGGS18, LMM18, LWW16, LLZ18, OVI^{+}12, PDY^{+}23, RG17, WFL^{+}25, YLH17, ZSP^{+}21, ZHH^{+}24, ZAI^{+}16, ZB20, Bac11, CDM^{+}10, CKRJ17, DCMW17, GTGB14, GEK25, GCARPC^{+}01, KHL17, KKB14, KSS^{+}18, LZW^{+}15, NRS92, PMC05, RAP19, SWH^{+}13, SWC08, ZLLL13]. HeteroOS [KGGS17, KGGS18]. HeteroVisor [GLS15]. Heuristic [BL17, LWW16, XH90, CD14, HAK22, KMT14, SM23b, TSR19]. Heuristics [ARMMA18, ATS16, BB12, KR16, Man15b, SBNU18]. HI [Shr89]. HICAMP [CFS^{+}12]. hidden [CWdO^{+}06, WQG15]. Hiding [CLS07]. Hierarchical [ABB19a, DM75, Kee68, SCA^{+}24, HPS23, SPAK18, YWF09]. Hierarchies [TBS17]. Hierarchy [SBK15]. High [ACM98, ACM04b, AMA18, Bad82, BPP^{+}17, BCW20, CW03, CZH^{+}25, DMS02, DYL^{+}12, Han16, Hog02, IEE96b, IEE06a, IBBA20, KCWH14, KBK^{+}21, KKTM17, KMM13, KKS^{+}19, LCK11, LMG01, LRP^{+}19, LJZ12, LHAP06, MLG^{+}02, RCM^{+}12, RB01, SD01, SCSL12, SV13, SYC14, URJ18, Vog03, WQG15, WCC16b, YWCF15, ZLSI17, dGG^{+}17, AAF^{+}09, Ano96, BML^{+}13, DQR^{+}13, EMS15, FF96, Fu10, G^{+}01, GTN^{+}06, GGJ^{+}92, GBCW00, HKJ19, LBZ^{+}11, LLLE17, LM99, LMG00, LDL^{+}08, ML78, MUKX06, M^{+}06, MRC^{+}13, MMG^{+}18, RQD^{+}17, SB10, SPF^{+}07, SPAK18, TZY^{+}25, WXW15, WWH^{+}17, XJW^{+}18, ZYZ^{+}18]. High-Assurance [LJZ12]. high-availability [Fu10, LDL^{+}08]. high-bandwidth [WXW15]. High-Endurance [AMA18]. High-Fidelity [KKTM17]. High-Level [DMS02, RB01]. High-Performance [ACM98, CZH^{+}25, IEE06a, IBBA20, KCWH14, LMG01, SD01, SCSL12, URJ18, WCC16b, dGG^{+}17, Han16, Hog02, KBK^{+}21, SYC14, HKJ19, LLLE17, LM99, LMG00, MUKX06, SPF^{+}07, SPAK18, TZY^{+}25, WWH^{+}17, ZYZ^{+}18]. high-performing [GBCW00]. High-Speed [KKS^{+}19, LRP^{+}19]. High-Throughput [BCW20]. Higher [BW03, MKM^{+}08]. higher-assurance [MKM^{+}08]. Highly [KD78, YYY^{+}23, ZFL15, CARB10, CGM17, GI12, GVI13, TGCF08]. Hilton [IEE90b]. HipHop [AEM^{+}14]. histograms [CL14]. History [Ran20, SKJ^{+}17]. History-Based [SKJ^{+}17]. HITAC [KAH83]. Hitless [ZWZ20]. HIVE [Tay76]. HLA [LCT^{+}15]. HLA-Based [LCT^{+}15]. hold [Yur02]. Holders [War11]. hole [EB20]. Holistic [LGJ^{+}18]. Home [DW14, See08b]. hones [Won97]. Honeypot [YYY^{+}23]. Honeypot-Based [YYY^{+}23]. honeypots [ALL06]. Hood [Ven96, Ven97b, Ven97c, Ven97d]. hooks [AKCP21]. Hop [WBHN18]. Hopping [DL19b]. horizontal [AH24]. Horus [LSX^{+}24]. Hose [YLH17]. Host [CLW^{+}14, NASD21, QNC07, HM20, LMJ07, TB14, TMV25, LSX^{+}24]. Host-Based [CLW^{+}14, NASD21, LMJ07, TMV25]. Hosted [SVL01, CBLFD12, CKT08, DS09b, SYZZ^{+}14]. hosting [RQD^{+}17, YMY17]. Hosts [BB13, Bau06c, CLL^{+}13, TtLcC13]. Hot [IEE96a, IEE97, IEE99, IEE01, BBTK^{+}17]. Hotel [USE01a]. HotOS [IEE01]. HotOS-VIII [IEE01]. Hotplug [LJL^{+}15]. HotSpot [Sch13a, IRB19, Arm98, BOF17, HHV^{+}02, WKJ20]. HotSpot^{T M} [RB01]. Houston [ACM06d]. HP [BKMM87, MSCK92]. HPC [M^{+}06, GPS^{+}18, HCJ07, JQWG15, PNT12, PCB^{+}18, Spa19]. HPC-GTP [M^{+}06]. HPC.NET [Vog03]. HPCC [DF96]. HPCS’06 [IEE06a]. HPVM [KSS^{+}18]. HSM [HYK^{+}23]. HSPT [WLW^{+}15]. HSSM [Wel02]. Huge [Got07, KYP^{+}17]. HVM [LTK17]. HVMs [CBZ^{+}16]. HW [DCG12, Wu13]. HW/SW [DCG12, Wu13]. Hybrid [GSW^{+}17, HD16, HAK22, KCWH14, LSC^{+}17, LGS^{+}23, PST^{+}15a, RSNK17, VVC^{+}17, WGLL13, FX06, GHK24, KN18, KSS^{+}20, KSS^{+}23, KS18a, LQW^{+}12, RJK^{+}17, STMV18, YB24, YWGH13, ZGW^{+}06, Gua14]. Hybrid-Copy [LSC^{+}17]. Hydrakon [PC24]. Hyper [Gal09b, Lar09, LC09a, TZB19, WXW15, Apr09, Car06, KVV09, KSS09, KS10, Lar09, LC09b, LC09a, MG08, MG09, SRS09]. hyper-space [WXW15]. Hyper-V [Gal09b, Lar09, LC09a, Apr09, Car06, KVV09, KSS09, KS10, Lar09, LC09b, LC09a, MG08, MG09, SRS09]. HyperBench [WZT19]. Hyperconverged [KAA^{+}25]. Hypercubes [HO92]. HyperMAMBO [dGG^{+}17]. HyperMAMBO-X64 [dGG^{+}17]. HyperMonitor [XZ11]. HyperSafe [WJ10]. hypertext [Alf91]. hyperthreading [BKR20]. Hypervisor [BAL15, CL16a, Chi08, HWCH16, JZY^{+}22, JSHM15, JAS^{+}15, KYP^{+}17, LKL^{+}19, NASD21, NOT^{+}17, PPG^{+}17, SJV^{+}05, SKYK16, WJ10, WHD^{+}16, XD16, XD17, YJZ^{+}21, You73, ABG14, ASB18, BBD^{+}10, CLM^{+}22, CDO24, DN14, MSZ09, NS17, RSLAGCLB16, Ste14, SL12, KSS09, KS10]. Hypervisor-as-a-Service [WHD^{+}16]. Hypervisor-Based [BAL15, LKL^{+}19]. hypervisor-secure [SL12]. Hypervisors [Rev11, SPF^{+}07]. hypervolume [EB20]. HYVI [Gua14].

I-Caching [MM93]. I-IoT [BSL^{+}18]. i.e [MC93, Müh75]. I/O [RM03, AJM^{+}06, AMA18, ASMA21, AD11, ABG14, ABB^{+}15, BMS16, BPM^{+}22, BHEP14, CWH^{+}16, CDD13, CRZH15, DCP^{+}12, DS09b, FMX^{+}25, GCL^{+}21, GAH^{+}12, HA79, HB12, HZK^{+}25, JAD19, KS08a, KBDK22, KAA^{+}25, KMN^{+}16, LLLE17, LMR18, LHAP06, LFHS23, NsP16, PST^{+}15a, RB24, Rus08, SBQZ14, SYC14, SVL01, THH^{+}14, TtLcC13, VW08, WR12, WYZY24, WTL^{+}16, XNH21, YJZ^{+}21, ZWFX17, ZSR^{+}05, ZXL^{+}24]. I/O-intensive [BPM^{+}22]. I/Os [OBSR16]. IA [Ano14b, De 06, Don06].

IA-32 [Ano14b]. IA-64 [De 06, Don06]. IaaS [FFM^{+}23, GLLJ16, GA18, HKM^{+}18b, KDB16, PPO14, RB17, SBBP20, XCSM18, ZLHD15, ZHW^{+}17]. IAS [FS11]. IAS/von [FS11]. IASSim [FS11]. IASTED [Ano99b]. iAware [XLL^{+}14]. IBBE [SXH^{+}19]. IBM [ADG^{+}92, A^{+}04, ABDD^{+}91, ABB^{+}15, Ber86, BBB^{+}25, B^{+}05, Bri98, D^{+}04, GBO87, G^{+}06, G^{+}05, Kam75, MIS^{+}05, Mly09, Olb78, P^{+}08, R^{+}06, R^{+}02, SZ88]. IBM/360 [Kam75]. ICE [Ano06a]. ICL [HP77, Kee77]. ICTree [FBZS12]. ID [SIJPP11]. ID/Locator [SIJPP11]. IDE [Ano03a]. idea [BBS06]. Ideal [Ran20]. identification [BZD17, HLW^{+}23, PMP23]. Identifying [CL17a, MD12]. Identity [SXH^{+}19]. Identity-based [SXH^{+}19]. Idiom [KKM^{+}13]. Idle [DEE^{+}16, SBK15, HKJ19]. IEC [Int05a, Int05b, Int06b, Int06c, Int06a]. IEEE [ACM04b, ACM05c, ACM06a, IEE90a, IEE91, IEE02, IEE03, IEE04]. IEEE/ACM [ACM04b, IEE04]. Igniting [ACM03a]. II [BBB^{+}25, Cre08a, TSR19]. IIoT [TDMP23]. IJCAI [MR91]. IJCAI-91 [MR91]. ILDJIT [CARB10]. Illinois [ACM05d]. illuminating [BK14]. im [KGG00, Mar08, Zim05]. IMA [XHCL15]. Image [AD11, CWH^{+}16, ETAB22, EF94, HKM^{+}18a, LZY^{+}24, NSJ12, ZCL^{+}21, IM93, KMG^{+}18, SBBP20, XZZ^{+}16, XWX^{+}17, ZXW16, ZFY18]. Image-Content-Aware [CWH^{+}16]. Images [Li14, WWL^{+}17b, ZTA^{+}21, BNS18, GKP^{+}19, Isl19, XJWW15, XJR^{+}17, ZMD^{+}21]. iMeter [YZLQ14]. iMIG [LZL^{+}15]. immutable [SV15]. Impact [KKS^{+}19, LSSC22, Ros06, WKJ20, WZKP19, BT15, WKJ17]. impacts [HH19, KWZ^{+}19]. Impasse [APST05]. Imperative [LFBB94]. implement [Sig89]. Implementation [fLtNW14, BBD^{+}91, DAH^{+}12, DJ77, DLS^{+}01, Hal79, JR02, JJ02, KR94, LXQ24, Mar73, MD12, MNN05, MN91, NsP16, Rev11, RNS^{+}23, SGS92, SIR^{+}17, SCD90, SB73, Sur01, TVO92, TO96, TFtLcC15, UOKT84, WLW^{+}15, War80, YLWH14, YCL^{+}19, YW20, ZSE^{+}25, ZSXZ07, ZL18b, AFT01, ANH00, BKR20, Blu02, BT15, CKP78, DN14, DJ76, DCA04, IT86, JNR12, Lav10, Man18, MJ93, PNM^{+}20, Sch09, SJW^{+}13, SGGB99, SGGB00, SJL20, Taf11, WW77, XCJ^{+}14, Lee86]. Implementations [AP22, HLP^{+}16, SVB93, VV18, AEMWC^{+}12, CSS^{+}16, CGC^{+}24]. Implementierung [Mar08]. Implementing [CTS^{+}93, D^{+}04, FR24, LFBB94, Tai98]. Implications [RM03, DLLN18, GTN^{+}06, MT16, MT17, ZTA^{+}21, DLL^{+}16, Pat12, RVJ^{+}01]. Important [SC18, CK06b]. Improve [GKXK13, GKBB15, KDB16, SAT09, YWGH13, YQZ14]. Improved [BRS^{+}22, War80, BTLNBF^{+}15b, KS20a, KU24, NGN24]. improvement [WWL^{+}25, YLH14]. Improving [AWR05, BHEP14, BCG73b, CFG^{+}13, CFRSSR19, HXZ^{+}16, HLW^{+}13, IPRS21, JKB15, KL13, LCT^{+}15, LBL16, LLWM23, LQD^{+}18, OSK15, QXH18, RSC^{+}15, RSLAGCLB16, SP83, TCP^{+}17, WKJ15, WGXC25, WHSE15, XNH21, AA24, BMF23, GVI13, HC12, JYW^{+}13, LC14, OL13, UTO13]. IMSA [Ano99b]. in-depth [CBFH20]. in-kernel [Uhl07]. In-Memory [TF16]. in-situ [CKRJ17]. In-VM [LWLL10]. in-VM-assisted [PDM20]. Inapproximability [RS20]. Incentive [XLWZ18]. Incentive-aware [XLWZ18]. included [Ano97a]. including [B^{+}07, CGW07, WG07]. Incorporating [GH91b]. Increasing [LWLL10]. Incremental [LXZ^{+}21]. Independent [DHPW01, DS09a, KAH83, USE93, GPW03, PW03, PFH^{+}16]. Index [Cox12]. indexed [JYW^{+}13]. Indicators [PC24]. Indirect [tTR82, CEG07, EG03, JYW^{+}13, KJM^{+}07]. individual [LWLL16]. induced [ZLZ^{+}19a]. Industrial [AAMBE21, LXQ24, QCM^{+}24, CCDR22, PTD^{+}18]. Industry [SXH^{+}19]. Inference [ZXG^{+}24]. Inferno [WP97]. Inferring [LHW^{+}20]. InfiniBand [PRS16, RS16, YCL^{+}18]. influence [Mly09]. influencing [BJ20]. Information [CAF^{+}91, IEE93a, Int05a, Int05b, Int06b, Int06c, Int06a, SS75, SS05, Ano93, BSD19, LC09a, MD73, MD74, RRB17]. Informed [HKKW13]. Infragistics [Ano03b]. Infrastructure [ECM01, ECM02, ECM05, ECM06, GMGF24, HW12, Int05a, Int05b, Int06b, Int06c, Int06a, LPSS19, MK13, McC08, MJW^{+}06, Nel04, NKK^{+}06, NSC^{+}22, OG16, Ott18, PP16, XH16, AO16, AMA^{+}14, AA18, BDS^{+}09, Car14, FF22, Hal09, HS13, HH13, Hui18, J^{+}05, KN18, KSRL10, KR16, LLY^{+}18, Low08, dOL12, YW20, MR04, PW03, RSF03, Fro13]. Infrastructures [KAA^{+}25, MVMHL24, WTM18, ZB20, ACG18, CSMB15, FPGK18, LPBB^{+}18, MPM^{+}20]. Ingens [KYP^{+}17]. inherently [TDG^{+}18]. injection [CP17b]. InkTag [HKD^{+}13]. Innovation [ACM03a]. innovations [ABB^{+}15]. Input [ACL72, Wal76]. Input-Output [ACL72]. insider [LC09a]. Insiderinformationen [LC09a]. insiders [KSS09, KS10]. Insights [Rev11]. Inspection [SKI^{+}17]. inspired [KHA22]. Installation [Bec09, Bor01, KGG00, Lar09, WF03, Zim05, Zim06, MIS^{+}05]. Instance [AMIA19, EMAL17, KCKC15]. Instances [WUNK17, ZG13]. Instant [HPP15, Joo06]. Instruction [Oi06, HW15]. instructional [DSSP06, DTW07, WO75]. Instructions [Qia99]. Instrumentation [ZFL15, BZA12]. Instrumenting [MZG14]. Instruments [BPB86]. Integer [Bos24, YTY00]. integer-reference [YTY00]. Integrated [BDF19, QLL^{+}21, SP22, ZXG^{+}24, vCPWvT11, CWG00, HKJ19, YZLQ14]. Integrating [JMSLM92, LTT92, LCL14, OBSR16, GEK25]. Integration [GMP89, VGF16, Ame13]. integrierten [Deu08]. Integrity [CW03, DL19a, DM75, (Fo71, (Fo78, QT06, WJ10, CS76, JXZ^{+}10, KBC21, LXRS19, XHCL15]. Intel [AJM^{+}06, COV^{+}24, CMP^{+}07, DLM^{+}06, Don06, KBC21, NSL^{+}06, NKK^{+}06, NBB^{+}19, RSW^{+}06, RI00, UNR^{+}05, Uhl06, vSMK^{+}20]. Intelligence [MR91, AA24, JNR12, MPA^{+}18]. Intelligent [GH91b, MVMHL24, PF23, SSS25, FGZC23, HTAY21, JYOB18, PTD^{+}18, SPSL25]. intelligente [PO09]. IntelliJ [Ano03a]. Intensive [WZH^{+}16, BPM^{+}22, GKJ^{+}19, IKU15, JFZL17, dCJR16, KBDK22, LFHQ19, QXH18, SZKY21, VVB13]. Inter [cCWS14, GGM^{+}16, RLZ^{+}16, BML^{+}13, CBZ^{+}16, SWcCM12, SBP^{+}17, SWL^{+}23, VOS12]. Inter-Application [cCWS14, SWcCM12]. inter-cloud [SBP^{+}17]. inter-connectivity [VOS12]. Inter-Domain [GGM^{+}16, BML^{+}13, SWL^{+}23]. Inter-Virtual-Machine [RLZ^{+}16]. inter-VM [CBZ^{+}16]. interact [EGD03]. Interacting [SK13a]. Interactions [cCWS14, SWcCM12]. Interactive [Hir17, LD05, MLA83, SSG90, WLS^{+}18, YYY^{+}23, Ber86, HMS04, KKJL14]. Interconnect [RCM^{+}12, SKJ^{+}17]. interdependencies [LBF12]. Interface [Cro93, SH04, Sun95a, Guz01, HP77, VL00]. Interfaces [CLKEF21, Mac79, PST^{+}15a, WML02]. Interfacing [MC93]. Interference [LXW^{+}23, NBH08, WYZY24, XLL^{+}14, XLJ16, ZRD^{+}15, ZLZ21a, HL13, gKEY13, LFHS23, SS13, VVB13]. Interference-Aware [XLL^{+}14, XLJ16, ZLZ21a]. Interferences [ZRZY15]. InterLISP [II79]. intermediate [GLV99]. internal [SI81]. internals [MKM^{+}08]. International [ACM00, ACM05a, ACM05b, ACM05d, ACM06b, ACM06f, Ano99b, BW03, IEE84b, IEE85, IEE93a, IEE96b, IEE02, IEE03, IEE04, IEE06b, IEE06a, LCK11, MS91b, MR91, Ost94, SS05, Shr89, Tho93, TLC06, ACM06c, JPTE94, M^{+}06, HHK94]. Internet [Ano99b, CK06b, KGG00, ASL^{+}20, AAMBE21, APST05, Ano03a, CHCC07, CK06b, CK06e, FR24, KB21, LLW98, Mon97, PTD^{+}18, SXH^{+}19, SDM21, WSX^{+}19, Wid01]. Internetkommunikation [CK06b, CK06e, CK06c, CK06d, CK06g, CK06f]. Internetprogramme [CK06b]. Internetprogrammen [CK06e, CK06c, CK06d, CK06g, CK06f]. Internship [HMS17]. Interoperability [GSS^{+}18, CPM^{+}18, Men03]. Interoperable [TDMP23, KKB14]. interposed [ZSR^{+}05]. Interpretation [FTNY69]. Interpreter [MSI18, SMK02, Ber86, KMMV14]. interpreter/graphic [Ber86]. interpreter/graphic-simulator [Ber86]. Interpreters [EG01, CEG07, EGKP02, EG03, Ert05, KKC^{+}16, SYZZ^{+}14, ZLBF14, Ert03]. Interpreting [Han05]. Interpretive [AS76, OJG91]. interpretive-execution [OJG91]. Interrupt [CL16a, TFtLcC15, AA18]. interrupts [AGH^{+}16]. interval [AHRR22a]. interval-based [AHRR22a]. Intranet [Ano03a]. Intrinsics [PSBG11a, PSBG11b]. introduce [MS01]. Introducing [BG74]. Introduction [A^{+}04, BG73b, CK06a, CK06b, CK06e, FDF05, IBM76b, KS08b, Sch94b, Sch94a, Boe15, Wün13]. introductory [BR01, Don88]. Introspection [CCML12, CLcC13, DGLZ^{+}11, FL13a, NBH08, Pfo13, SIdLB15, WWMG06, DSR23, FL13b, HN08, HL24, HcC14, TZY^{+}25]. Introspection-based [CLcC13]. Intrusion [HTAY21, AMA^{+}11, LMJ07, MA17, TP25]. intrusions [JKDC05]. intrusive [ZXY^{+}15]. Invariant [SCF^{+}25]. Invariants [PEC^{+}14]. Investigation [EVCL21]. invited [Piz17]. invocation [Ven97c]. Invocations [WZKP19]. IoE [SCC^{+}23]. IOMMU [YWCF15]. IoT [AHKC23, ABB^{+}19b, AMB^{+}17, BSL^{+}18, BLMP22, Bos24, GEK25, HCB18, KS20a, LXQ24, MKZ^{+}25, MPA^{+}18, MA19, NBS18, PFPJ18, RC18, RNS^{+}23, SJL20, TV18, XLL^{+}20, XZL^{+}20, ZYZ^{+}18]. IoT-Enabled [MKZ^{+}25, KS20a]. IOV [DYL^{+}12, DCP^{+}12, HB12, XD16, XD17, YWCF15, ZXL^{+}24]. IP [AM16, CF00, HWHW18, NTR18]. IPv6 [GLQ^{+}13, SAR24]. IPv6-Based [SAR24]. Iron [Ano05]. IronGrid [Ano03b]. irregular [AC16]. ISA [CWH^{+}14, DZ02, KNHH18, WLW^{+}15, WCC16b]. Ischia [ACM06e]. ISCOPE [ACM01b]. ISDF [M^{+}06]. ISDN [KGG00]. iShare [WTL^{+}16]. ISO [Int05a, Int05b, Int06b, Int06c, Int06a]. ISO/IEC [Int05a, Int05b, Int06b, Int06c, Int06a]. Isolated [Jen79, ZD18, KKK^{+}18]. Isolation [GGK18, HWR^{+}24, KA24, WZL15, ZZW^{+}21, ZTWM17, CDO24, Cza00, GNDB16, JK17, MD73, MK19, MK23, WTL^{+}16, YJZ^{+}21]. isolation-enhanced [MK23]. ISPA [M^{+}06]. ISPAN [HHK94]. ISPs [PST15b]. ISSTA [Ost94]. Issue [KM13b, TZB19, WYZAD20, Yur02]. Issues [AFG^{+}17, AD11, Car23, KS08a, KK19, MZ20, PZH13, SABL20, SEF^{+}06, Tur84, XKKL23, ARA20b, ARA20a, AGH^{+}15a, AEB19, BB08, CCDR22, CBFH20, PBB13]. Italy [BW03, M^{+}06, ACM06e]. Itanium [Ano06a]. Itanium-based [Ano06a]. Items [BB17]. iterators [ZLBF14]. IV [Int06c]. IVME [Ert03]. IX [BPP^{+}17, IEE97].

J [AC98]. J2EE [JDJ^{+}06]. J9 [WKJ15]. Jahrestagung [Müh75]. Jail [McK04, Sar01]. Jailed [Wid01]. Jalapeño [AAB^{+}00]. January [ACM99, IEE93a, Shr89, USE01b]. Japan [HHK94]. Java [ACM98, ACM01b, Ano00, Ano01a, Ano01b, Ano02, Ano03a, Sch13a, USE01c, USE01d, USE02, Wol99, ADM98, Ame13, AT16, Ano97b, Ano97c, Ano97d, Ano03b, AFT01, ABC^{+}07, AC98, ANH00, BDF^{+}98, BHDS09, BD01, BP01, BP03, Bri98, BZD17, Caa00, CW03, CT03, CGC^{+}24, CH08, Cla97, Coh97, CDG97, Cra98, Cza00, Dalxx, Dal97, DHPW01, DD20, DEK^{+}03, DS09a, DBC^{+}00, DCA04, DLS^{+}01, EGD03, Eng99, EL98, Eug06, FFB^{+}00, Fra98, FK03, G^{+}01, GGG03, GCARPC^{+}01, GC00, GPW03, GBCW00, HT98, Han05, HM01, HOKO14, HWB03, HB08, Ive03, JR02, JJ02, Juo07, Kal97, KS13, LM99, LMG00, LB98, LV99, LY97a, LY97b, LY99, LYxxa, LYxxb, LYBB13a, LYBB13b, LYBB14, LTK17, MSG01, MO98, Men03, MD97, MDxx, MLG^{+}02, MB98, Mon97, MP01, NG13, OT97]. Java [Oak14, Oi05, Oi06, Oi08, PTHH14, PNM^{+}20, PRB07, PV06, Qia99, RVJ^{+}01, RHR02, Ran02, R^{+}13, Req03, RRB19, SMK02, SSB^{+}14a, SD01, SE12, SH04, Sch13a, SSMGD10, Set13, SMSB11, SSB03, Shi03, SM01, SGV12, SEPV19, Siv04, Smi97, SSB01, SSB14b, SHB^{+}03, Sun95b, Sun95a, SUN97, JCV99, Sun99, STS^{+}13, SM02, Sur01, Tai98, Tol98, TO96, UBF^{+}98, UR15, Van98, Ven97a, Ven97b, Ven97c, Ven97d, Ven99a, Ven99b, VED06, VED07, VL00, WL96, WGF11, Wak99, WH99, Wes98, Wol99, Won97, WWMG06, WZL^{+}18, YC98a, YC98b, YME05, YKM17, Yel99, YTY00, ZP14, ZS01, vLSM01, Ano97a]. Java-based [Ano96, FF96, HOKO14, KS13, YC98b]. Java/CORBA [GCARPC^{+}01]. JavaCard [BDJdS02]. JavaScript [AHK^{+}15, CBLFD12, UIK19, VP16]. JavaScriptCore [Piz17]. Java^{T M} [LMG01, SMES01, CF00, RB01, vD00]. Javy [GGG03]. JCloudScale [ZLHD15]. JDMM [ZP14]. Jenga [TBS17]. JET [MLG^{+}02]. JetBrains [Ano03a]. jetzt [KGG00]. Jikes [AAB^{+}05a]. Jini [JJ02]. JiST [BHvR05]. JIT [JK13, PFH^{+}16, THL03, WKJ17]. JIT-based [PFH^{+}16]. JIT-compiler [THL03]. JITs [KRCH14]. JN [Mon97]. JnJVM [TGCF08].

Job [MNT14, PS19b, HKJ19, PC21, RAP19]. jobs [KS18b]. jockey [Hin97]. John [IEE06a]. Joint [CTP^{+}17, EGG^{+}24, FXHY21, KA25, MSLY24, MAK18, NTH^{+}17, RJK^{+}17, SJ21, WZV^{+}13, YXL^{+}20, ATZP21, LKR^{+}19, SBNU18, SM23b]. Jointly [LWL16, XHW^{+}19]. Jon [Ano97a]. Jose [Ano04b]. journaling [HC12]. JP2 [SSB^{+}14a]. JPDA [Sun99]. JPF [BA19, BA23, WKG17]. JPR [WKG17]. jRapture [SCFP00]. JS [AHK^{+}15]. judgment [CSV15]. July [IEE06b, Sof83]. Jump [WBHN18]. June [ACM90, ACM01a, ACM01b, ACM05d, ACM06f, IEE85, USE85, USE86, USE01a, USE06]. JVM [Ano00, Ano01a, Ano01b, USE01c, USE01d, USE02, AC16, CSS^{+}16, DBC^{+}00, Guy14, Kha19, R^{+}13, RRB17, SSB^{+}16, SYZZ^{+}14, SV15, Sub08, Sub11, Ven99b, WKJ20, WKG17]. JVMPI [Sun95a]. JVMs [BK14].

K. [Sch94a]. Kailua [Shr89]. Kailua-Kona [Shr89]. Kaleidoscope [LFBB94]. Kanazawa [HHK94]. Kanotix [CK06c, CK06h, CK06l, CK06r, CK06h]. Karlsruhe [RM03]. KDE [KGG00]. Keeping [NP13]. Kernel [FL13a, HD16, JJ91, KZB^{+}90, LSX^{+}24, SM90, SYB12, TY14, WLMD16, DD20, LWM14, LY23, Uhl07, VMBM12, KM13a, KM13b]. Kernel-based [TY14, KM13a, KM13b]. Kernelized [WCC16b]. kernels [HPHS04, RMB02]. Key [LCMV17, TF16, DPW^{+}09]. Key-Value [TF16]. Kinder [CK06q, CK06t, CK06r, CK06s]. Kingdom [Vra05]. kit [Car06, LC09b]. knapsack [EYGS19]. knew [RAT17]. Knob [WUK^{+}18, BR01]. Knoppix [CK06d, CK06i, CK06m, CK06s, Deu08, CK06i]. knot [LBF12]. Know [NBB^{+}19]. Knowledge [FG91, LZZ^{+}25, FS19, IT86, RAT17]. knowledge-based [FS19]. Kochbuch [PO09]. kompletten [Mar08]. Kona [Shr89]. Konfiguration [Bor01, Lar09, WF03, Zim06]. konfigurieren [RHM08]. Konsolidierung [See08a]. Konzept [Dal97]. Konzepte [Tho08]. Konzeption [Zim06]. krill [BB20, KS18a]. KScalar [MRL02]. Kubernetes [BSNB20, Car23, ZB20]. Kubernetes-Based [ZB20]. Kubuntu [CK06e, CK06j, CK06n, CK06t, CK06e, CK06j]. Kuck [War11]. Kundenserversystemen [See08a]. KVM [Deu08, Hin08, CJK^{+}12, DN14, GLC84, HWCH16, IYAK23, LZL^{+}15]. KVM-based [HWCH16]. KVM/370 [GLC84]. KVM/ARM [DN14]. KylinX [ZZW^{+}21].

L [Lot91]. lab [AL05, HMS04]. laboratories [DTW07]. Laboratory [GPM21, Kim84, SVN^{+}10]. Labs [See08b]. lag [ZMD^{+}21]. Lagrange [SS22]. Lagrangian [GR15]. Lagrangian-based [GR15]. Lake [ACM03b]. Lambda [Wat86, Wat87]. land [Tsa14]. Landing [ACM03b]. Language [CDM^{+}10, ECM01, ECM02, ECM05, ECM06, GSS^{+}18, Hog08, Int05a, Int05b, Int06b, Int06c, Int06a, Kam83, Luc97, MR04, PW03, PFH^{+}16, RSF03, SIR^{+}17, SVB93, SUN97, WIDP12, WBHN18, Arv02, Ber86, BD01, BMER14, DH01, Don88, GLV99, Hog06, IT86, Juo07, KRCH14, Les74, MD12, MC93, PRB07, RJK16, RSW91, SKC73, SMO84, Taf11, Tai98, WCG14, WWH^{+}17]. Language-independent [PFH^{+}16]. language-level [WCG14]. Language-Neutral [WBHN18]. Languages [BS90, Dan86, GC00, KP99, LFBB94, PTHH14, SSG90, Tol98, YKM17, ACM99, BDT13, Jou85, ML78, MRG18, PMC05, PULO16, SSB^{+}16, Sus76, TB14, Wel02, Wu13, YWF09]. LARD [WCG14]. Large [DK93, GKBB15, PHL^{+}12, RIP18, RGSJ17, SADP21, SLM89, XDLS15, ZSXZ07, ZLW^{+}14, ZTA^{+}21, BLRC94, DK75, FPGK18, LPD^{+}11, Nie12, Req03, STMV18, SZ13, SHTE11, WCG21, YZSC17]. Large-Scale [PHL^{+}12, SLM89, XDLS15, ZLW^{+}14, ZTA^{+}21, SZ13, WCG21, YZSC17]. last [Rob12]. Latency [ASSB18, BPP^{+}17, BL17, MV16, RZPX19, IMK^{+}13, MMTM22, ZSW^{+}06]. latency-aware [MMTM22]. Later [FS12]. launch [AMIA19]. launch-time [AMIA19]. Layer [SKT^{+}19, ZXG^{+}24, BTLNBF^{+}15b, BTLNBF^{+}15a, EBLM22, MA17, RSLAGCLB16, ZFY18]. Layer-Wise [ZXG^{+}24]. layered [PSC^{+}07]. layering [YWF09]. LayerMover [ZFY18]. lazy [Wak99]. LCVP [LLS^{+}20]. LDA* [YZSC17]. leadfoot [HHPV15]. Leaking [vSMK^{+}20]. Lean [WZL^{+}23, SV15, Ven96]. Learn [BWH^{+}19]. Learn-as-you-go [BWH^{+}19]. Learning [Bos24, BRX13, Che21, DS18, FFM^{+}23, GPM21, HTM^{+}24, HPS22, JYOB18, KKE19, KSVR23, MMdMCPSR25, MSC^{+}21, RNS^{+}23, XRL^{+}22, ZXR^{+}22, AD18a, AJBJ23, BMF23, GH20, GHK24, GKT17, HBNK24a, HBNK24b, KRG^{+}12, NNK21, Par25, QBL^{+}23, RGAT18, RT18, WBW^{+}19, WZZ^{+}20]. Learning-based [DS18, ZXR^{+}22, HBNK24a, HBNK24b, Par25]. Learning-Driven [XRL^{+}22]. legacy [LU04]. Legally [Sam22]. LegoSim [RMB02]. Length [GR20]. LEON [PDL^{+}23]. Lern [CK06q, CK06t, CK06r, CK06s]. Lern- [CK06q, CK06t, CK06r, CK06s]. Lernprogramme [CK06k, CK06m, CK06l, CK06n, CK06o]. Lernprogrammen [CK06k, CK06m, CK06l, CK06n, CK06o]. Lessons [RM03, LJZ12, Rob06, URJ18, HMS04]. Leuven [ACM04a]. Level [ASMA21, AC16, cCWS14, Chu06, DMS02, GCL^{+}21, KHW^{+}16, MMdE19, NTR18, RB01, SV13, WXG^{+}25, ZSR^{+}05, ZQCZ16, ZXL^{+}24, AD18a, AL05, BSM^{+}12, BSD19, BSOK^{+}20, But94, CLM^{+}22, Cia07, EGD03, FLCB10, IM75, JHE14, KA24, LZY^{+}24, LZW^{+}17, ML78, SVN^{+}10, SWcCM12, SSG90, WHSE15, WF07, WCG14, ZLZ13]. levels [CCMY07]. Leveraging [LLF^{+}18, LDL^{+}08, Pfo13, RTL^{+}18, SLZ^{+}25, WHD^{+}09, ZL13, AJD09, RAT17, ZBG^{+}05]. Libraries [DK93, Int05b, DSS19, Won97]. Library [Cro93, SJS^{+}17, KS20b, PBWH^{+}12]. libvfio [RB24]. libvfio-user [RB24]. libvirt [Ano14c]. License [HO22]. Life [ZR06]. Lifetime [BFM^{+}21, WJ10]. Light [WWL^{+}17a, HB08]. Light-Weight [WWL^{+}17a, HB08]. Lightweight [ABV12, CXLX15, CZH^{+}25, CDM^{+}25, PLZ20, Ran02, VN06, WJ10, YME05, ZLW^{+}19b, ZTWM17, vMAT14, AMA^{+}11, CCL^{+}17, DQR^{+}13, DL19a, KSS24, PDM20, RQD^{+}17, SSU^{+}12, TMJ^{+}21, TB14, XZ11]. Like [Abr80, RHV17, SSOT17, Voe86]. LILA [Dan86]. Limbo [Luc97]. LimeVI [WLG^{+}11]. limited [CH08]. Limits [WBB^{+}16, vKF13]. line [SV17]. linguistic [UR15]. Link [KLLT18, SAR24, CRB12, GGJ^{+}92, JK15]. Link-Failover [SAR24]. linked [FC98]. linking [FC98]. LINUX [KGG00, Ano06a, CK06a, CK06b, CK06g, CK06f, CK06i, CK06h, CK06j, CK06o, CK06p, G^{+}06, Mar08, USE00a, WF03, ABB19a, Bau05, Bau06c, BBHL08, Ble10, Bor01, CK06a, CK06b, Com00, Com03, DN14, Dav04, Fab13, G^{+}06, GNDB16, MZG14, NSHW10, NV05, P^{+}08, Ros14, Spr06, Spr07, VMBM12, Wün13]. Linux-based [ABB19a]. Linux-Server [Mar08]. Linux/OSS [Ble10]. Liquid [Li14, ZL18a]. LISP [ACM90, CK87]. List [TT96]. List-based [TT96]. Listing [LKL^{+}19]. Literature [BDF19, DCM22, MBK24, ARA18, ARA20b, ARA20a, ZJRW19]. LITL [Lam75]. little [Men03, YYPA01]. Live [AGC18, AY24, BWH^{+}19, CCZ^{+}06, Deu08, DK17, ECJ^{+}16, HKN22, HTB22, JFPL16, JDW^{+}14, KKLV16, LSC^{+}17, LZL^{+}15, LJL^{+}11, LH15, LZM^{+}20, MZD^{+}18, MB15, MK22, MSC^{+}21, RJS^{+}18, SHW^{+}15, SKI^{+}17, TUM18, XLL^{+}14, XD16, XD17, ZRS^{+}16, ZDLG17, ZXY^{+}15, ZZW22, AA24, AS14, BAC15, BB08, DS20, EYG21, FGLI15, GJK^{+}20, GNK24, HLW^{+}10, HBNK24a, HBNK24b, HTB19, HDG09, IMBB20, JKK^{+}13, JFZL17, JGW^{+}11, JGSE13, LFHQ19, NK22, NIA18, PKS^{+}19, PDC^{+}12, SS22, SSL^{+}13, SLA^{+}16, SHTE11, TDG^{+}06, WLG^{+}11, WRSvdM11, WRS^{+}15, YW20, ZLLL13, Isl19]. Live-Distribution [Deu08]. live-migration [JKK^{+}13]. Live-Streaming [MSC^{+}21]. lively [STFH15]. Liveness [ADM98, LDL14]. LLC [KKH14]. LLVM [LH13]. Load [ARAAA19, BFM^{+}21, CJK^{+}12, CL16a, DY17, KAA^{+}25, KAZS14, KK19, LW12, LYS^{+}18, SAR24, YWR^{+}14, Bir94, DS18, GH20, KNHH18, RI25, TF16, XH90, XTB17, YWH^{+}23]. load-balanced [DS18]. Load-balancing [KAZS14]. Load-Sharing [SAR24]. Loading [LB98, HSC15, WGF11]. Loads [LTE12, WCZ^{+}23]. Local [ADM98, Oi08, PCR89, HJ10, KMT14, Oi05]. Locality [HSC15, SZ88]. Localization [YYL^{+}15]. Located [LGJZ16, SOAK23]. Location [LLS^{+}20, THB22, USE93, OG16]. Location-Constrained [LLS^{+}20]. Location-Independent [USE93]. Locator [SIJPP11]. lock [YTS14, YQZ19]. lock-aware [YQZ19]. Logic [DMS02, FD08, GH91b, LSPP^{+}23, UOKT84, Alf91, BY20, Bur02]. Logic-Based [FD08]. Logical [Com65, RT93, Lia05, TT93]. Logically [Jen79, KKK^{+}18]. Logics [BW03]. Logisim [Bur02]. logistics [LZWC13]. LogP [CKP^{+}93]. Logs [HPB^{+}24]. Long [KKLV16, KGS16, SOKE23]. Long-Distance [KKLV16]. long-running [KGS16]. Longest [HWHW18]. Look [HMS17, SMP22]. lookaside [CFG^{+}13]. Looking [LHW^{+}20]. Lookup [HWHW18]. Loris [AvMT11]. Loss [XDLS15, CHCC07]. Lösungen [Tho08]. LOTOS [MS91a]. Louis [ACM97]. Low [BPP^{+}17, KKD24, RZPX19, WCG14, ZHCB15, GE85, IMK^{+}13, SJRS^{+}13]. low-cost [SJRS^{+}13]. Low-Latency [RZPX19, IMK^{+}13]. Low-level [WCG14]. Low-Overhead [KKD24, ZHCB15]. low-resolution [GE85].

LSTM [CEPR22, EVCL21, KU24]. LSTM-based [CEPR22]. LTTng [WKJ15]. Luminous [KNT02].

m [USE01c, Abr82, KAH83, AS85a, AS85b]. M-series [KAH83]. MA [USE06]. MAC [SJV^{+}05]. MAC-Based [SJV^{+}05]. Mach [USE91, MRGB91]. Machine [AGJS16, AS85a, ABCC66, AT24, AAAF21, AAR22, ABV12, Ano00, Ano01a, Ano01b, Ano02, Ano04a, Ano04b, fLtNW14, AE01, Apr09, Arc07, AAK18, AGIS94, BWP85, BFHW75, Bai70, Bak83, Bal91, BDF^{+}99, BH73, BN75, BWD^{+}15, Bos24, BFM^{+}21, BJH^{+}16, BG73a, BCG73b, BG74, CTS^{+}93, CW03, CFH^{+}79, CFH^{+}80, Car13, CF00, CGC16, CRZH15, Cox09, CWL^{+}15, CHPY17, CYX^{+}17, Dalxx, Dal97, DHPW01, Dan86, DCM22, DF96, DGLZ^{+}11, Dom80a, DL19b, DJ77, ETAB22, EG01, FZD^{+}24, FG91, Fie68, Fis01, FPS^{+}02, (Fo71, (Fo78, FFM^{+}23, FL13a, GKSP99, Gei02, Gen86, Gol69, Gol71a, GA25, GLBJ18, HHV^{+}02, HHW10, Hal79, HTW^{+}19, Han73, HH79, HKM^{+}18b, Hir17, Hor73, HKKW13, HPS22, IBM72, IBM73, IBM76b, Ibs84a, JHS12, JJK^{+}11, JMSLM92, JQWG15, JN15, JADAD06a, KC16, KS08a, KKE19, KSS^{+}20, KSS^{+}23]. Machine [KMK16, KNT02, KF91, Ken80, KDB16, Kim84, KAH83, KGZ^{+}04, KLF^{+}15, LCWB^{+}11, LMM18, Lau87, LW73, Law00, LW11, LSC^{+}17, LLW98, LTE12, Li14, LVM16, LGJ^{+}18, LLY^{+}25, LTT92, LY97b, LYxxa, LYxxb, LYBB14, LWLL10, LJL^{+}11, LLS^{+}20, LZZ^{+}25, LPB17, LFBB94, Loy92, LXM^{+}16, MSG14, Mac79, Mad69, MS91a, MK13, Man16, Mar73, MZ20, McC74, MS70, MD97, MDxx, MDGS98, MKKE12, MA21, MA24, II79, Mur69, NBH08, NBK16, NMG15, Nel04, NASD21, NLD^{+}23, NSJ12, NL19, Olb78, PPTH72, PP73, PAC^{+}22, PXG^{+}17, Pfo13, PF23, PCC^{+}16, PK75a, Pro00, Qia99, QT06, RNA^{+}22, RG17, RLZ^{+}16, Ren78, RK24, RI00, RSN^{+}18, RT93, Ros99, RG05, RMP24, Ibs84b, SL14, San88, SGK^{+}23, Sch94b, Sch94a, SSB03, SMA18, SCP93, SSG90, SHZ^{+}14, SB73, SHB^{+}03, SVL01, Sun95b, Sun95a, SUN97, JCV99]. Machine [TT96, TMV12, TY14, USE01c, USE01d, USE02, VTW16, Ven97a, VL00, WL96, WIDP12, Wak99, WH99, WDL^{+}20, WB81, WZH^{+}16, WWL^{+}17a, WGXC25, Wel94, WCSG05, WHD^{+}09, WP97, WLCS17, XWJX15, XLJ16, XLWX19, YWY^{+}17, YLQL25, YP15, ZDK^{+}22, ZLW^{+}14, ZRS^{+}16, ZL16, ZCG^{+}17, ZL18b, ZLZ^{+}19b, ZCL^{+}21, ZZF06, ZWL^{+}18, ZXR^{+}22, ZHL16, ZZW22, ZJXL11, ZTWM17, Zyt94a, Zyt94b, dSdF16, AD18a, Abr82, AS85b, AD19, AGSS10, AJBJ23, AGH^{+}15b, AGH^{+}15a, AHRR22a, AHRR22b, AT23, ATZP21, AAB^{+}00, AA24, AC95, Ame13, Ano94, Ano96, Ano99a, AO16, AFT01, ABC^{+}07, Arm98, AWR05, Arv02, AP18, ANH00, AMA^{+}11, BB20, BDF^{+}03, BBTK^{+}17, Beg12, BPC94, BMF23, BJ20, BCM90, BRS^{+}22, Bir94, Blu02, BADM06, BFC02, BY20, Bri98, CARB10, CL14, CD14, Car14, CEG07, Cav93, CFVP12, CS76, CHCC07]. machine [CCL^{+}20, CLL^{+}23, CGC^{+}24, CBLFD12, CK06a, CK06e, Clo85, Cof99, CGV10, dCCDFdO15, CWG00, CD01, DSR23, DH01, DSC^{+}08, DP11, DM93, DBC^{+}00, DLH^{+}20, Don87, DHD20, DJ76, DXM^{+}17, DS22, EYG21, EGKP02, EG03, FGZC23, FLL^{+}13, FS19, FF22, FM90, FA21, FSFP19, FMIF18, Fit14, FF96, FLM^{+}08, FCG^{+}05, Fre05, FRM^{+}24, GGQ^{+}13, GTGB14, GSG22, GH20, GHK24, GSKJ18, Gol74, GCARPC^{+}01, GPW03, GR80, GBCW00, GLW23, GA18, HZL^{+}18, HJ10, HKN22, HBNK24a, HBNK24b, HTB19, HL24, HUL06, HAK22, HK07, HcC14, HPHS04, HLBZ20, HSC15, Hui18, HPS23, IBM85, IBM88, Int88, IBM94, IBM96, IRB19, IKU15, JKK^{+}13, JNR12, JC18, JGW^{+}11, JADAD06b, Kal97, KOY05, KBDK22, KB21, KS13, KS20a, KSO^{+}15, KS18a, KZ21, KTB17, KK21, gKEY13, KCS14, KJLY15, KCKC15, KKC^{+}16, KMG^{+}18, KFF12]. machine [KHA22, KSS^{+}18, Kou11, KCV11, KRG^{+}12, Lam75, LSS24, LBZ^{+}11, Les74, LC02, LM99, LZWD15, LBL16, LWLL16, LYYY18, LLWW18, LFHQ19, LXRS19, LLZ^{+}19, LZLY20, Lia05, LL14, LPZ^{+}22, LWCZ22, LCL^{+}23a, LCL^{+}23b, LPBB^{+}18, Lot91, LG93, LFHS23, LQD^{+}18, LY23, MSG^{+}12, MR23, MD73, MD74, MSG01, MNB24, DPBK16, MS17, Man18, MNA16, MS00, McG72, MC93, McM11, MRG18, MN91, MST^{+}05, MW18, MHM19, EYGS19, MAK07, MJ93, NZH20, NGN24, NNK21, NK22, NOK^{+}85, NAR19, NIA18, OG16, Oi08, ORPS09, PEL11, Par25, PFPJ18, PCB^{+}18, Piz17, Pon19, Pul91, PS23, RKT20, RHR20, Raj79, RWC21, RZ14, Req03, RK18, RFBLO01, RY10, RJK^{+}17, RCTY19, SZKY21, SBI21, SBBP20, SHR19a, SHR19b, Sch13b, SSMGD10, SEM^{+}20, SHLJ13, SM23a, She91, SCEG08, SASG13, SSEA18, SL00, Sig89, SGGB99, SGGB00, SKC73]. machine [Smi97, SYMA17, SJL20, SMA^{+}10, SBP^{+}17, SSU^{+}12, Str05, SP23, SPSL25, TSLBYF08, TZY^{+}25, TMLL14, TDD20, Tay76, tTR82, THG^{+}18, TIIN09, TB14, TMV25, TT93, Tur84, Vag10, Van98, Ven96, Ven97b, Ven97c, Ven97d, Ven99b, VVB13, VGL23, WGF11, WKT08, WRX11, WZV^{+}13, WKJ15, WCY^{+}17, WSX^{+}19, WWL^{+}25, Web10, WHW20, WLL^{+}13, WW77, Won97, XHL^{+}13, XCJ^{+}14, XJWW15, XZZ^{+}16, XLWZ18, XZK^{+}20, XXWG23, YME05, YZW^{+}13, YLH14, YLHJ14, YPLZ17, YLCH17, YW20, YBZ^{+}15, YYC^{+}19, YWH^{+}23, YLK^{+}10, Yel99, YSM^{+}21, YC16, YIR24, YB24, YRJ18, YGN^{+}06, YQZ14, YQZ19, YTY00, ZG13, ZXW16, ZYZ^{+}18, ZLZ15, ZLH^{+}15, ZHHC17, ZFY18, ZWC^{+}19, ZLZ^{+}19a, ZBP07, ZLW^{+}19a, ZFL^{+}23, ZLL^{+}16, ZL13, ZLLL13, ZWH^{+}17, ZLCZ18, ZYLY18, ZWC^{+}14, dSOK17, AEM^{+}14, AAB^{+}05a, Ano97b, Ano97c, Ano97d, AC98, BD01, BP01, BP03, BZD17, Caa00, CCWY05, CK87, Cla97]. Machine [Coh97, CDG97, Cra98, Cza00, DD20, DCA04, DLS^{+}01, Eng99, FS11, FFB^{+}00, Fra98, FK03, Fuj91, GKP^{+}19, GGG03, GC00, HT98, HM01, HLW^{+}23, HWB03, HB08, Ive03, JR02, JDJ^{+}06, JJ02, Juo07, KM13a, KM13b, LMG00, LMG01, LB98, LV99, LY97a, LY99, LYBB13a, LYBB13b, LTK17, Men03, MB98, Mon97, MP01, OT97, Oi05, Oi06, PTHH14, PNM^{+}20, PRB07, Ran02, RRB19, RB01, SMK02, SSB^{+}14a, SH04, Sch13a, SMES01, Set13, SMSB11, Shi03, SGV12, Sim92, Siv04, SSB01, SSB14b, SM02, Sur01, Tai98, THB22, Tol98, TO96, TR88, UR15, Ven99a, Wel02, Wol99, WWMG06, YI24, vD00, Ano97a]. Machine-Based [LW11, WB81, CGV10, WKT08, YZW^{+}13]. Machines [Ano75, AY24, ASSB18, Att73, AH68, BMS16, BP99, BDJdS02, BSSS14, BWH^{+}19, Bee05, BB13, BJC24, BJPS73, BRX13, BG73b, BCG73a, CL17a, CWL12, CCML12, CWS12, CJK^{+}12, CGMD19, CSS^{+}13, CL16a, CCO^{+}05, CH78, CHLY18, CDN02, DSM14, DEK^{+}03, Den01, DK17, DMR10, DKW15, Do11, EGR15, EGJS15, ECJ^{+}16, EGG^{+}24, Ert03, EDS^{+}15, Gai75, Gal73, G^{+}01, GTS^{+}15, Gol71b, Gol73b, Gum83, Han73, HKLM17, HTB22, HB17, Hof20, HS06, HPP15, Ian14, JE12, Jen79, JXL^{+}12, JAS^{+}15, JKJ^{+}10, KCWH14, KJL11, KP15, KPHA20, KAH83, Kov19, KAS13, LMR18, LZL^{+}15, LYYY17, LLWM23, LD05, LXW^{+}23, LHAP06, LW12, LJL^{+}15, LLZ18, LSPP^{+}23, Mac79, MB15, Mal73, Man15a, MD12, MGL^{+}17, MM94, Par71, Par72, PSBG11a, PS16, Ran20, Rev11, Ros04, SD01, SCSL12, SV13, SN05a, SN05b, Sta97, SKI^{+}17, Sup04, TTH^{+}19]. Machines [TV12, UT87, Vog03, WLW^{+}15, WGLL13, WZL15, WLLZ16, Win71, XSC13, XLL^{+}14, XLL^{+}20, ZRD^{+}15, ZWC^{+}23, vLSM01, Ågr99, ABB19a, AAH^{+}03, ADA^{+}19, AGH^{+}16, ATS16, AAM^{+}16, AMAB17, AS14, BAC15, Bac11, Bag76, BML^{+}13, BDF^{+}98, BDQR23, BHvR05, Bel06, BB12, BB15, BJ22, BPM^{+}22, BBM09, BBS06, BB95, CL17b, CGM17, CSSE21, CCL^{+}17, CH08, Cra05, Cra06, CWdO^{+}06, CLL^{+}13, DDS^{+}94, DC15, DEG^{+}17, DQLW15, DAdBM^{+}24, DSZ11, DCMW17, EB20, EGD03, EM06, Ert05, EL98, EMS15, FBZS12, Fit14, FHL^{+}96, FGLI15, FX06, Fu10, GI12, GVI13, GJK^{+}20, Gol73a, GKJ^{+}19, GLV^{+}10, HKS19, HM18, HMH17, HZZ^{+}14, Hin97, HDG09, Hol95, IMBB20, JES^{+}15, JWH^{+}15, JDW^{+}14, JGSE13, JYOB18, JB24, KSS24, KDK20, KSSG16, KRCH14, KBB11, KBC21, KR16, LCLG26, LMJ07, LZC^{+}16, LLF^{+}18, LJL12, LQW^{+}12, LF19, LC13, LTZ^{+}14, LSS04]. machines [Man15b, Mat09, MK19, MK23, MG13, MRG17, MMTM22, hTMAC^{+}08, MPM^{+}20, NK10, NOR15, PKS^{+}19, PFH^{+}16, PSBG11b, PMC05, PDM20, PBYH^{+}08, PRS16, PV08, uRQS20, RK16, RH17, RHR02, RG19, RT18, RI25, SJB14, SS13, SENS16, SNV10, Sch09, SSN12, SJJ^{+}12, SJW^{+}13, SWH^{+}13, SLC20, SS22, SSL^{+}13, SPAK18, Ste14, Str13, SK13c, SLA^{+}16, SHTE11, Syr07, TZK17, TGCF08, TMMVL12, TDG^{+}06, TtLcC13, UIK19, VT14, VED07, VWT13, WQG15, WXZ^{+}17, WDT18, WCS06, WSVY09, WRSvdM11, WRS^{+}15, WCG21, WCZ^{+}23, XNH21, XHCL15, XWX^{+}17, XTB17, XA22, YC98b, YWF09, YLJ22, YWGH13, ZBG^{+}05, ZWHC17, ZWL09, ZSRR22, ADM98, BHDS09, CT03, Cla97, MLG^{+}02, PEC^{+}14, SM01, UBF^{+}98, VED06, YC98a, ZS01]. macro [Wel02]. macro-architecture [Wel02]. Made [Ste05, PDL^{+}23]. Mail [Joo06]. Main [AW17, AMH^{+}16]. mainframe [GBO87]. Mainstream [Uhl06, BBHL08]. maintaining [HBP06]. Maintenance [MBK24, LSS04]. Major [Cap21]. Make [THB06, BC10, DMH18]. makes [Wal10]. Making [HKKW13, NSC^{+}22, Voe86, XLL^{+}14, CFRSSR19, FA21, SJJ^{+}12]. Malicious [SMA18, Kip21]. Malware [CLS07, CD12, GG11, HL24, AD18a, CVWL13, CWdO^{+}06, PDM20, YJZY12]. MAN [TDG^{+}06, YYPA01]. MAN/WAN [TDG^{+}06]. manage [Car14, Fit14]. Manageability [Gua14, MW05]. managed [CBGM12, CFG^{+}13, GK05, RJK16]. Management [AW17, CMK^{+}16, CTP^{+}17, DMR10, HTW^{+}19, HC17, HTB22, KGGS17, KGGS18, KR18, KL14, Lar09, LJL^{+}15, LCMV17, LCFL12, LXM^{+}16, MBWW86, MDGS98, PLMA18, PYYG21, RC18, SMES01, SLW^{+}24, SC17, SDD^{+}16, SKT^{+}19, SP22, TB17, WIS^{+}15, WLW^{+}15, WYZY24, WGLL13, YYY^{+}23, ZCL^{+}21, AHK^{+}15, ATS16, ARMMA18, BAC15, Beg12, BBMA91, BHDS09, BN89, CH08, Cla05, EBJ17, Fit14, Fu10, GTGB14, GLK^{+}12, GAHL00, HKJ19, HB13, IMK^{+}13, IPRS21, KCKC15, KMG^{+}18, KF18, KB17, LLS^{+}08, MR23, MS00, MBA^{+}12, NGN24, NBS18, NS07, dOL12, PMP23, RH17, RHR20, RP07, RJK16, SBBP20, SG10b, SWC08, TRG13, Wal02, WDCL08, WWWL13, WB16, WCS06, WSVY09, YLCH17, YWTC15]. Management-Complexity [SP22]. Manager [Car13, Car14, KMT14, Apr09, MBA^{+}12]. Managing [BB13, KGZ^{+}04, LCZ^{+}19, BCP^{+}08, J^{+}05, YLHJ14]. Manipulating [GK05]. Manipulation [VGF16]. Mantle [BB95]. Manual [CRZ83]. Manufacturing [NHL22, LLS14]. Many [Bai70, JAD19, JYM^{+}23, LPB17, SXMX^{+}18, CLL^{+}13, DQR^{+}13, FGZC23, WR07]. Many-Accelerator [SXMX^{+}18]. Many-Core [JYM^{+}23, JAD19]. Many-Objective [LPB17, FGZC23]. Manycores [HPP15, KHW^{+}16]. Mapped [HW93, BLRC94, SV15]. Mapper [AJ18]. Mapping [Bak83, CFM17, LYY^{+}24, PS16, PCC^{+}16, BSOK^{+}20, CRB12, HS13, HSC15, JK15, UR15, YXL^{+}20, WK08]. MapReduce [HSC15, RAP19, XYYY17]. March [ACM06d, Ano10, SS05]. Marine [MMG^{+}18]. Market [LS15]. marketplace [KMK10]. Markets [TVKB16, XRL^{+}22]. Markov [BL17, MTFK19, NZH20, RH17, RHR20, SC17, WQG15, YIR24]. Markov-based [NZH20]. Marriott [USE01a]. Mars [BC19]. Maryland [Ano93]. Maschinen [Zim06]. Massachusetts [USE93, USE01a, IEE85]. Massively [BS90, Kra90, MM93]. Mastering [CBER09, Low09, Low11, LMG^{+}14, McC08, Sub11]. Matching [CFM17, Cox07, Cox09, Cox10, Cox12, YDW18]. Maté [LC02]. matrix [Kra90]. maximally [SS19]. Maximization [MLXG19, ZHW^{+}17, JWH^{+}15, KTB17, LWLL16]. Maximizing [BYBYT16, ZRD^{+}15]. May [ACM00, ACM06e, Ano04b, IEE84a, IEE90a, IEE91, IEE01, IEE06a, Mar81, TLC06, USE99, USE06, Yur02]. mayfly [DS22]. MBSA [CCL^{+}17]. MC [XJW^{+}18]. MC-VAP [XJW^{+}18]. MC68020 [MMM84]. MCG [ZGW^{+}06]. MCG-mesh [ZGW^{+}06]. MDev [PYDG22]. MDev-NVMe [PYDG22]. MDRUs [MTFK19]. Mean [ARAAA19, Ven96, ZB18]. Measurement [ACM81, Cal75, WLS^{+}18, LXRS19, XHCL15]. Measurements [WZL^{+}23, KBC21]. Measures [Att79, SM92]. Measuring [LWLL10, PC24, XHL^{+}13]. MEC [MMdMCPSR25, SDM21]. Mécanismes [Han73]. mechanics [MC98, Uhl07]. Mechanised [LSPP^{+}23]. Mechanism [LCT^{+}15, LLZ18, MD12, TVKB16, Mly09, SIRP17, SYMA17, TXD^{+}24, YLH14, YLWH14, YLJ22, ZLH^{+}15]. Mechanisms [Han73, NMG15, Nel04, XRL^{+}22, KHA22, LCLG26, MG13, RHZ^{+}17, TMMVL12]. MECOM [JDW^{+}14]. Media [JW17, ZCG^{+}17]. Mediated [PYDG22, XYD^{+}18, PRS26]. Meet [JW17, FHL^{+}96]. Meets [BBM^{+}15]. Megalos [SADP21]. Megh [BWH^{+}19]. mehr [Joo06]. Memento [CPST15]. MemFlex [ZLSI17]. memories [Pat12]. Memory [AW17, AZEE17, AZEE18, AMH^{+}16, Bad82, Bro89, VMW^{+}19, CLŁS12, CDW^{+}24, Cro93, GHS17, GGJ^{+}92, GKBB15, HHS18, HHC^{+}16, HPP15, JJK^{+}11, KLY20, KGGS17, KGGS18, LW11, LH16, LSSC22, LJL^{+}15, LSX^{+}24, LZW^{+}17, LXM^{+}16, MKKE12, NTL^{+}24, PGP19, PP73, PDY^{+}23, RC18, RLZ^{+}16, RWX^{+}12, RGSJ17, SMES01, SLM89, SLW^{+}24, VTW16, Wal02, WWH^{+}16, WWL^{+}17a, WK90, WTLS^{+}09, WZL^{+}18, XML^{+}18, ZSE^{+}25, ZL18a, ZLSI17, ZCL^{+}21, AHK^{+}15, ASPP22, ATS14, Ano15, BHDS09, BFS^{+}18, CWH^{+}14, CWC^{+}14, CLcC13, CH08, CMM^{+}06a, CMM^{+}06b, CMM^{+}06c, GPS^{+}18, GMK17, GVI13, GNDB16, GLV^{+}10, HKN22, HB13, HHPV15, HUWH14, JSK^{+}13, JDW^{+}14, KB17, LLWW18, LFHQ19, LJYZ15, LF19, LLS^{+}08, MS00, PNM^{+}20, PPO14, RO16, RJK16, SEPV19, SOKE23, VED07, WWS89, WZW^{+}11, WWWL13, WK08, ZP14, ZWKX17, ZHCB15, ZWL09, ZL13, TF16]. Memory-Aware [JJK^{+}11]. memory-based [SOKE23]. memory-limited [CH08]. Memory-Oriented [ZL18a]. memory-performance [SEPV19]. Memory-Resident [WK90]. merging [TLX17]. mesh [SJRS^{+}13, ZGW^{+}06]. Message [GGM^{+}16, DM93, TO91, UR15, XH90]. message-passing [TO91, UR15, XH90]. messaging [Joo06]. meta [BT15, HAK22, SBNU18, TSR19]. meta-heuristic [HAK22, TSR19]. meta-heuristics [SBNU18]. meta-tracing [BT15]. metacircular [PBAM17]. Metacomputing [MDGS98]. metaheuristic [ATZP21, EYGS19, XA22]. metaheuristic-based [XA22]. metaheuristics [ARMMA18, SEM^{+}20]. Metal [ZXL^{+}24, AGH^{+}16, GAH^{+}12, OSK15, RB24]. Method [AAT^{+}22, AAMBE21, AC16, BP99, BA19, DEK^{+}03, HT98, LZL^{+}15, LYY^{+}24, LLY^{+}25, LXQ24, Mar73, QLL^{+}21, RSNK17, SXH^{+}19, SSS25, TTH^{+}19, ZAI^{+}16, ATZP21, BJ22, CSSE21, DXM^{+}17, GSG22, JKK^{+}13, JXZ^{+}10, LYYY17, LYYY18, LXRS19, LYY^{+}20, MHM19, MA19, NS17, SEM^{+}20, THH^{+}14, Ven97c, XZK^{+}20, XA22, YLHJ14, ZFL^{+}23, ZZW^{+}24, ZSRR22]. Method-Level [AC16]. methodology [FS89]. Methods [BDG18, HSN17b, KKS^{+}19, Pfo13, Qia99, UT87, WH99, AAC^{+}17, BMWB86, MG19, NK22, XH90]. metric [SS17]. metrics [BSOK^{+}20, Sch13a]. Metriken [Sch13a]. Metron [KBK^{+}21]. Mexico [ACM00]. Meyer [Ano97a]. MGC’05 [ACM05b]. MI08 [Hin08]. Micon [BGS89]. Micro [LYY^{+}24, ZD18]. Micro-Computing [ZD18]. Micro-Segmentation [LYY^{+}24]. Microarchitectural [MSI18, CJJ^{+}22, EGD03, SK13b]. Microarchitecture [SVM^{+}23]. microcomputer [UBL^{+}82]. microcomputers [GBO87]. Microgrids [GPM21]. microkernel [GMR93, Sto07, Uhl07]. microkernel-based [Sto07]. Microkernels [FHL^{+}96, HUL06]. Micromachines [Sch73]. Microprocessor [Ran02, ACT94, WW77]. microprocessors [But94]. microprogrammable [Bag76]. microprogramming [ML78, SP83, Tho73]. Microservice [KKD24, BNS18, WGW^{+}18]. microservice-based [BNS18]. Microservices [Kol19]. Microsoft [Lar09, Zim05, Ano99a, B^{+}07, Car13, CBER09, Gal09b, Joo09, Kal97, KVV09, KSS09, KS10, Lar09, MRM06, Nou92, Ste05, Won97]. Middle [ZYH^{+}19]. Middleboxes [KRS^{+}17, YDW18]. Middleware [ACM05b, HOKO14]. Migratability [ZXL^{+}24]. Migrate [YBZ^{+}15, CLL^{+}13, KB21]. Migrating [JE12]. Migration [AGC18, ABV12, AY24, BWH^{+}19, BFG^{+}14, BWD^{+}15, CYX^{+}17, DK17, EMAL17, GWZ16, HTB22, HPS22, KC16, KGS16, KKLV16, LSC^{+}17, LZL^{+}15, LJL^{+}11, LH15, LZM^{+}20, MZD^{+}18, MB15, NBK16, PS19b, RSNK17, RSN^{+}18, RJS^{+}18, SL14, SHW^{+}15, TMV12, XWJX15, XLL^{+}14, XD16, XD17, XLWX19, YWR^{+}14, YWW^{+}15, ZRS^{+}16, ZCG^{+}17, ZDLG17, ZLZ21a, ZZW22, vLSM01, AGH^{+}15b, AGH^{+}15a, AA24, AS14, BAC15, BMF23, BB08, CLcC13, DS20, EYG21, FGZC23, FMIF18, FGLI15, GJK^{+}20, GNK24, HLW^{+}10, HKN22, HBNK24a, HBNK24b, HTB19, HH19, HAK22, HDG09, HPS23, JKK^{+}13, JGW^{+}11, JDW^{+}14, JGSE13, KN18, KLY20, KSS^{+}20, KSS^{+}23, KZ21, KTB17, KJLY15, LZWD15, LZC^{+}16, LFHQ19, LLZ^{+}19, DPBK16, MG13, NK22, NAR19, NIA18, PC21, PKS^{+}19, PDC^{+}12, Par25, PFPJ18, PCB^{+}18, RK16, RCTY19, SEM^{+}20, SM01, SS22, SYMA17, SSL^{+}13, SLA^{+}16, SHTE11, TK20, TMV25]. migration [TDG^{+}06, WCY^{+}17, WSX^{+}19, WDT18, WLG^{+}11, WRSvdM11, WRS^{+}15, XWW^{+}21, XA22, YW20, YBZ^{+}15, YKP^{+}24, ZLZ15, ZHHC17, ZFY18, ZLZ^{+}19b, ZLZ^{+}19a, ZFL^{+}23, ZNSL14, ZLLL13, ZYLY18, MK22, TUM18]. Migrations [WVT^{+}17, CBJ22, GSKJ18, JES^{+}15]. MigVisor [ZDLG17]. MIMO [LZ15]. Mini [ZXY^{+}15]. Mini-intrusive [ZXY^{+}15]. Miniboxing [UTO13]. minicomputer [KK79]. MiniComputers [Har77]. minidisk [Boz89]. Minimal [LPD^{+}11]. Minimal-overhead [LPD^{+}11]. Minimization [GR20, Kha19]. Minimizing [LGJ^{+}18, LGS^{+}23, RSNK17, RKRK17, SSS25, ZRD^{+}15, Par25, RK16, SZ13, THG^{+}18]. Minimum [BAC15, FSH^{+}13, ZLZ15]. Minimum-cost [ZLZ15]. Minimum-energy [BAC15]. Mining [NASD21]. MINIX [Kel06, vdK09]. Minneapolis [IEE92]. Minnesota [IEE92]. MIPS [MTJ^{+}22, RWX^{+}12]. MIRAGE [PC21]. mirror [Rob06]. Misalignment [SC18]. misses [BLRC94]. Misson [Ano10]. Mistakes [Ste05]. Misuse [Aln22]. Mitigate [WWL^{+}17a]. Mitigating [ASSB18, IYAK23, WZKP19, ASB18]. Mitigation [LGR14, IRB19, vCPWvT11]. Mixed [PA21, WLMD16, ZXG^{+}24, CCDR22, CDO24, LWM14]. Mixed-Criticality [WLMD16, CCDR22, CDO24]. Mixed-Precision [ZXG^{+}24]. Mixing [LD05]. MLN [Beg12]. MMU [XYD^{+}18]. MO [ACM97]. Mobile [CPKL17, CPS17, CWH^{+}16, LH16, LGXC23, LYS^{+}18, MV16, RSN^{+}18, SGB^{+}16, SML18, USE93, WVT^{+}17, WCC20, XZL^{+}20, ZLW^{+}19b, BD11, BBD^{+}10, CM18, FGZC23, FC98, HLW^{+}10, IIK^{+}06, ISE08, LLLE17, LFHS23, SASG13, WHSE15, ZLZ^{+}19a]. mobility [FX06, SBP^{+}17, ZLZ^{+}19a]. mobility-induced [ZLZ^{+}19a]. Mode [Dav04, CWH^{+}14, Cof99, YLJ22]. MODEF [SMO84]. Model [Bar73, BRX13, CHW12, DL19b, GGK18, HKM^{+}18b, IBM76a, KKTM17, KF91, KCY22, KAS13, KAZS14, LLWM23, MTFK19, MV16, MP01, Nel04, Neu92, NSJ12, SAJ24, ST24, SJMG24, WLCS17, XDLS15, YLH17, YLQL25, ZDLG17, AJBJ23, AA24, Bar78, BMF23, BCM90, Bir94, CKP^{+}93, EYG21, Fre05, HBNK24a, HBNK24b, JFZL17, NNK21, RHR20, Req03, SS13, SPSL25, TMJ^{+}21, TP25, WO75, YZLQ14, ZP14, ZBG^{+}05, ZGL^{+}17]. Model-Driven [NSJ12]. Model-Free [BRX13, AJBJ23]. Modeling [ACM81, CH78, GLL^{+}21, IN87, KRG^{+}12, LDL14, PFNC20, SHB19, TIIN09, WDL^{+}20, WLS^{+}18, WZZ^{+}20, XWH^{+}16, BPM^{+}22, BB95, FX06, gKEY13, SK13c, TLX17, YZSC17]. Modelling [DPBK16]. Models [DSM14, HBL^{+}10, HWB03, KKE19, Man15a, RSW^{+}06, SL16, TUM18, ADG^{+}92, BKR20, CPM^{+}18, CBFH20, HCJ07, Lia05, RO16, VVB13, WDT18, Ble89]. Modem [Ano03a]. Modern [BDG18, EG01, FKZ17, GG11, KKS^{+}19, FIF^{+}15, KB17, ZDK^{+}19]. Modification [Aln22]. modified [FS19]. Modular [AvMT11, ADWM18, DCA04, FC98, LH13, TO91]. Modularity [SVB93, DNR06]. Modulation [WUK^{+}18]. module [DJP^{+}24, ZXB^{+}24]. Modules [SMR24]. MOEA [WGXC25]. MOEA/D [WGXC25]. möglichen [Hin08]. moldable [HZZ^{+}14]. Molecular [YWCF15]. MOLP [ZB18]. monad [Dan12]. Monitor [LXM^{+}16, PDL^{+}23, QT06, Ren78, RI00, RT93, Ros99, SVL01, AGSS10, ALL06, AMA^{+}11, Cof99, KOY05, Kou11, SHLJ13, SSU^{+}12, TT93, XZ11, ZYZ^{+}18]. monitor-based [AMA^{+}11]. Monitoring [BAL15, CCML12, DLX^{+}17, LZW^{+}17, PC24, PLZ20, RHV17, WLLZ16, ZL16, ZL18b, ZXY^{+}15, ACT94, CL14, EYG21, JXZ^{+}10, JADAD06b, LMDP19, TZY^{+}25, WSX^{+}19, YCL^{+}19, YW20]. Monitors [JHS12, KS08a, KF91, RG05, WCSG05, BDF^{+}03, FLM^{+}08, HUL06, HPHS04, YME05]. Monona [ZL18a]. Monterey [ACM05a, Ano01b, USE91, USE01c]. mori [CPST15]. Mortar [HUWH14]. most [CK06b]. motion [Lia05]. Motorola [Ano03a, MMM84]. move [BGS13]. Moving [Cre10b, Cre10a, TMV25]. MPQ [ZXG^{+}24]. MPSoC [BHI15]. MPSoCs [OVI^{+}12]. MS [Tho08]. MU5 [MDFS72]. Multi [AVNR19, ABV12, AP18, BB17, BKM25, CLG^{+}10, CDM^{+}25, DY17, DLS^{+}01, Fie68, GA25, GSS^{+}18, GLBJ18, HMH17, HC17, HCB18, HPcC04, KSVR23, KR18, LZLY20, LLS14, LH15, LCZ^{+}19, MMdE19, MD12, MP16, MM94, PXG^{+}17, PNT12, RTL^{+}18, SL14, SCL^{+}19, TTH^{+}19, TSR19, TK20, WGXC25, WLL^{+}13, XCSM18, XZL^{+}20, ZL18a, ZRZY15, AD18a, AL05, ATS16, BB20, BDQR23, Bor07, BY20, DEG^{+}17, DHD20, DS22, FGG14, GGQ^{+}13, GKP^{+}19, GH20, GHK24, HZL^{+}18, JHE14, KMT14, LC14, LYYY18, LLZ^{+}19, LCL^{+}23a, LCL^{+}23b, MPM^{+}20, RK18, RPE12, STMV18, SE12, SWH^{+}13, SS19, SM23b, SIK^{+}16, SWW^{+}18, SOKE23, WDCL08, XZ11, XJW^{+}18, YKS16, YTS14, ZMD^{+}21, ZNSL14, ZLL^{+}16, JDJ^{+}06, NMS^{+}14]. Multi-Access [Fie68, HCB18]. Multi-Agent [PXG^{+}17, ABV12, DHD20]. multi-attribute [SS19]. Multi-Capacity [BB17, HMH17]. Multi-Channel [TTH^{+}19]. Multi-Cloud [AVNR19, DEG^{+}17]. multi-connection [XJW^{+}18]. Multi-Context [ZL18a]. Multi-Core [KR18, RTL^{+}18, PNT12, SWH^{+}13, YTS14]. multi-course [AL05]. multi-criteria [ATS16]. Multi-dimensional [HPcC04, ZMD^{+}21]. Multi-Dispatch [DLS^{+}01]. Multi-domain [TK20]. multi-faceted [BDQR23]. Multi-GPU [NMS^{+}14]. Multi-granularity [LLS14]. Multi-Language [GSS^{+}18, MD12]. Multi-Level [MMdE19, AD18a, JHE14]. Multi-Objective [GA25, GLBJ18, AP18, LZLY20, SL14, SCL^{+}19, TSR19, BB20, BY20, DS22, GGQ^{+}13, GKP^{+}19, GH20, GHK24, HZL^{+}18, LCL^{+}23a, LCL^{+}23b, MPM^{+}20, RK18, STMV18, SM23b, ZLL^{+}16]. multi-platform [XZ11]. Multi-processor [WLL^{+}13]. Multi-Provider [MP16]. Multi-Purpose [BKM25]. Multi-Resource [WGXC25, LYYY18, LLZ^{+}19, SOKE23]. multi-server [LC14, RPE12]. multi-source [SIK^{+}16]. Multi-stage [CLG^{+}10]. multi-start [KMT14]. Multi-tasking [JDJ^{+}06]. Multi-Tenancy [DY17]. Multi-Tenant [CDM^{+}25, LCZ^{+}19, ZRZY15, SWW^{+}18, YKS16]. Multi-threaded [HC17, FGG14, SE12]. Multi-Thresholds [XCSM18]. Multi-Tier [KSVR23, LH15, XZL^{+}20, WDCL08, ZNSL14]. multi-user [Bor07]. MultiCache [NsP16]. multicast [AAC^{+}17]. multicomponent [BRS18]. Multicore [FRD^{+}08, HHW10, Ian14, Man16, SOAK23, ZD18, CP17b, KNHH18, SE12, SSMGD10, SJJ^{+}12, Sub08, WCC^{+}16a, WCS09, WJGA12, YQZ14]. Multicore-Aware [Man16]. multicriteria [CFRSSR19]. multidimensional [dCJR16]. Multidomain [NVV^{+}24]. multigrid [AGIS94]. multihost [Bar06]. MultiLanes [KHW^{+}16]. Multilayer [VLZL16]. Multilayered [NsP16]. Multilevel [DD20]. Multimedia [Ano99b, CAF^{+}91, FLZ17, ZKWH17, BTLNBF^{+}15b]. Multimodal [ZZW^{+}24]. Multimodal-based [ZZW^{+}24]. Multipath [CFLL19]. Multiple [HTB22, LGS^{+}23, BG20, CSV15, Com00, GMR93, IKU15, OKAM17, SS22, SLA^{+}16, TMMVL12, TtLcC13]. multiplicity [SM79]. multiplier [SS22]. multiprocessing [DBO^{+}18, TLD^{+}89]. Multiprocessor [AGLM91, Dun86, NL19, KKJL14, WXZ^{+}17, Bro89]. Multiprocessors [Bad87, Cro93, SLM89, TO91, WWS89, WWT89, AGIS94]. multiprogramming [Abr82]. Multisensory [FBGS24]. multitarget [Bar06]. Multitasking [CD01, ZWC^{+}23, IBM96, TLD^{+}89]. multitasking/multiprocessing [TLD^{+}89]. Multitenant [KA24, LZW^{+}15, MK23]. Multithreading [LRZ16, ABB^{+}15, PV06]. Multitier [XRL^{+}22]. musical [BB08]. mutual [SGS92]. Mutually [SOAK23]. MVM [GMR93]. MVP [Lot91]. MySQL [Wün13].

Nam [Fro13]. Named [War11, XWJX15]. Nancy [Jou85]. Naplus [ZWKX17]. Narrow [YSS^{+}17]. Narrowing [DGLZ^{+}11]. Nash [uRQS20]. Nassau [Ano99b]. Nation [HS19]. Nation-Wide [HS19]. National [Ano93, SVN^{+}10, WZH^{+}16]. Native [AC98, UT87, EL98, RPE12, STS^{+}13]. NATUG [Boa90]. NATUG-2 [Boa90]. Nature [KHA22]. Nature-inspired [KHA22]. NC [Boa90]. NDSS [Ano10]. Near [ASPP22, LJFS17, UT87, XRL^{+}22, LKY^{+}17, RPE12, TDG^{+}18]. Near-memory [ASPP22]. Near-Native [UT87]. Near-Precise [LJFS17]. near-threshold [TDG^{+}18]. Neat [BB15]. need [BGS13, GLK^{+}12, WCS09]. needs [BKT^{+}19, STFH15]. Negotiation [ABV12]. Nested [HBL^{+}10, GHS16, KS20b, RQD^{+}17]. nested-virtualization [RQD^{+}17]. Nesting [BJC24]. Net [MBK^{+}92, Tur92]. NetAdvantage [Ano03b]. NetKernel [NSC^{+}22]. NetLCR [Joo06]. nets [NMC18a, NMC18b]. Netstumbler [Joo06]. NetWare [WF03]. Network [ACM98, RM03, AFG^{+}17, AP22, AVNR19, ASL^{+}20, Aln22, Ano10, AO16, ACA16, BYZZ20, BLMP22, BRIdM10, BL17, BHEP14, CFM17, CBJ22, CFC^{+}24, CPS17, CMK^{+}16, CFLL19, Che21, CKT08, Cre10b, CTP^{+}17, DW14, EMAL17, ELC^{+}19, EVCL21, EMW16, FXL^{+}23, FR24, Fis01, FML^{+}22, FLZ17, GHM^{+}18, HTAY21, HLPY16, HSL17, HB12, HTM^{+}24, HJG18, IKU15, JW17, KKE19, KKTM17, Ken80, KLR^{+}20, KAZS14, KLLT18, LXL^{+}22, LLW^{+}16, LHW^{+}20, LCZ^{+}19, LDRS18, LCFL12, MLXG19, MDZ^{+}21, MSLY24, MAK18, MMdMCPSR25, MP16, MCZ06, Mon97, MR06, NLD^{+}23, NSC^{+}22, NVV^{+}24, Nou92, PHL^{+}12, Pap20, PHXL19, PCR89, PST^{+}15a, PHC20, Rix08, RS20, RKRK17, SADP21, SAR24, SN23, SJMG24, SLZ^{+}25, SKT^{+}19, SSOT17, SP22, TSN^{+}23, UVL^{+}13, VV18, WB81, WZL^{+}23, XWH^{+}16, XWW^{+}21, XD16, XD17, YJZ^{+}21, YYY^{+}23, YWH^{+}21, ZSE^{+}25, ZWFX17, ZZG^{+}23, ZHHC17]. Network [ZSP^{+}21, ZLZ21a, ZWH^{+}17, ZKWH17, ACM06c, AM16, AMIA19, ALW15, BG20, BCC^{+}15, BCM90, BL90, BH13, BBS06, CBZ^{+}16, CB10, CRB12, Cre10a, DS19, DS18, DYL^{+}12, FCD09, FLL^{+}13, FZS^{+}20, FJKK17, FK13, FSH^{+}13, GLQ^{+}13, GLLJ16, HH18, HH19, HS13, HBP06, IM93, JAC^{+}19, JK15, KBDK22, KSO^{+}15, KK21, KKK^{+}18, KWZ^{+}19, KU24, LYYY17, LLZ^{+}19, LRP^{+}19, LMDP19, LQD^{+}18, DPBK16, MK22, MSZ09, MHS21, NTH^{+}17, OKAM17, OK90, PJZ^{+}19, PFNC20, PST15b, PBL^{+}16, QBL^{+}23, RK16, RWC21, SHB19, SOKE23, SZL^{+}14, SWL^{+}23, TSR19, TK20, TP25, TSCB19, Tur84, UBL^{+}82, VOS12, WWS89, WHC16, WCC16c, WBW^{+}19, WZZ^{+}20, WC91, WYZAD20, XHW^{+}19, YCL^{+}19, YLTF20, YXL^{+}20, ZLZ^{+}19a, ZJRW19, ZGL^{+}17, BCZ19, CEPR22, HTAY21, MCJ19, TF16, YWL^{+}18, YI24]. Network-Aware [CTP^{+}17, AO16, IKU15, ZHHC17, KK21, LQD^{+}18]. network-based [LYYY17]. Network-hosted [CKT08]. Network-I [RM03]. Network-I/O [RM03]. networked [CT03, HKN22, NBS18, SBNU18, SGGB99, SGGB00]. Networking [ACM04b, CPKL17, IEE06b, LCK11, MLA83, Pap20, SN23, SS05, SLZ^{+}25, SB18, XWJX15, ZKWH17, BTMS10, Bor07, BH13, GD08, Ker15, MCJ19, M^{+}06, Zho10]. Networks [AJM^{+}25, BSI^{+}15, CPKL17, CGC16, CFLL19, EVCL21, FXHY21, FML^{+}22, GMGF24, Hal79, HHK94, JN15, KKLV16, LLW^{+}16, LXZ^{+}21, LCMV17, MP16, MBWW86, MSC^{+}21, NGRF19, QLL^{+}21, SCC^{+}23, SLZ^{+}25, SIJPP11, TVO92, VVC^{+}17, XZL^{+}20, XRL^{+}22, ZDS^{+}22, ALW15, Alf91, AAC^{+}17, BTLNBF^{+}15a, CL15, CM18, DS19, FZS^{+}20, GCARPC^{+}01, GLQ^{+}13, GHM^{+}18, HHSG18, KCV11, LC02, LZW^{+}15, LWL16, MG19, Mon22, MAK07, NRS92, OMB^{+}15, RS16, THH^{+}14, TK20, TO91, WZV^{+}13, WT91, XWW^{+}21, XYYY17, XJW^{+}18, YKS16, YPLZ17, YLTF20, YMY17, AAJD^{+}16]. Netzwerk [KGG00]. Netzwerke [WF03]. Netzwerkkonfiguration [WF03]. Neumann [FS11, FS12, Sig89]. Neural [EVCL21, JAC^{+}19, MBK^{+}92, TVO92, Tur92, WWS89, ZSE^{+}25, ZDS^{+}22, Alf91, BCM90, BL90, IM93, KCV11, MK22, OK90, RK16, RWC21, TO91, WT91, WC91, HLW^{+}23]. Neural-FEBI [HLW^{+}23]. Neurocomputer [GFB^{+}92]. Neutral [WBHN18]. neutron [MM92]. Nevada [ACM81, ACM89]. newer [YK13]. Newfoundland [IEE06a]. News [Bri98, Kal97, Sta07]. Next [AEA^{+}23, BDF^{+}99, CF00, LPSS19, IIK^{+}06, RGS^{+}20]. Next-Generation [AEA^{+}23, IIK^{+}06, RGS^{+}20]. NFV [ALW15, Pap20, TF16, ASL^{+}20, BDF19, FS19, FLZ^{+}20, GDSA^{+}17, GLL^{+}21, JWL^{+}18, KBK^{+}21, LHW^{+}20, LXZ^{+}21, SDM21, SHB19, SP22, WTJR22, XRL^{+}22, ZHH^{+}24]. NFV-Based [SDM21]. NFV-Enabled [LXZ^{+}21]. NFV/SDN [BDF19]. NFVLearn [SOKE23]. NG2C [BOF17]. Nice [ACM90]. NICs [HB12]. Niklaus [BGP00]. Nimble [ZCJ^{+}21]. Ninth [USE00b]. NoC [FRD^{+}08]. NoCs [FD08]. Nodal [Che21]. Node [FXL^{+}23, NTR18, CRB12, JK15, KL13, LSS04, SS19]. Node.js [KKD24]. Node.js-Based [KKD24]. Nodes [SJ21, Vol90]. NoHype [KSRL10]. nom [BYBYT16]. Non [Aln22, AMH^{+}16, KS18b, LZZ^{+}25, PG17, PG18, WZL^{+}18, YKM17, KOY05, KM13a, KM13b, ZP14]. non-cache-coherent [ZP14]. Non-clairvoyant [KS18b]. Non-Control [Aln22]. non-dedicated [KOY05]. non-deterministic [KM13a, KM13b]. Non-Java [YKM17]. Non-Preemptive [PG17, PG18]. Non-uniform [LZZ^{+}25]. Non-Volatile [AMH^{+}16, WZL^{+}18]. Non-Volatility [WZL^{+}18]. nonaligned [AGIS94]. nonvolatile [PNM^{+}20]. normal [AM16]. Normalized [AAT^{+}22]. North [Boa90]. Nosv [RQD^{+}17]. Note [BCG73a, DMS02, KSS^{+}23, MA24]. notebook [IBM94]. Novel [AHKC23, ARAAA19, ATS16, BMJ^{+}22, JZY^{+}22, LSC^{+}17, LXQ24, NK10, PKS^{+}19, XCSM18, ZWFX17, DJP^{+}24, CBZ^{+}16, FF22, LXRS19, LJYZ15, SDN09, ZLCZ18]. Novell [WF03]. November [ACM75, ACM89, ACM96, ACM03a, ACM04b, ACM05b, ACM05c, IEE90b, IEE92, IEE93b, IEE02, IEE04, LCK11, USE91, ACM97]. NSGA [TSR19]. NSX [PPS^{+}18]. Nu [DNR06]. null [AT16]. NUMA [BMS16, GTS^{+}15, KP15, LL14, LXM^{+}16, SJA^{+}17, SKJ^{+}17]. NUMA-Aware [BMS16]. NumaGiC [GTS^{+}15]. Number [BP99, SZ13]. numbers [WCG21]. Numerical [Hol95]. nutzen [Zim06]. nützliche [LC09a]. NVMe [FMX^{+}25, HC18, HZK^{+}25, PYYG21, PYDG22]. NVRAM [ZLW^{+}19b].

O [RM03, AJM^{+}06, AMA18, ASMA21, AD11, ABG14, AIAR^{+}24, ABB^{+}15, BMS16, BHEP14, CWH^{+}16, CDD13, CRZH15, DCP^{+}12, DS09b, FMX^{+}25, GCL^{+}21, GAH^{+}12, HA79, HB12, HZK^{+}25, JAD19, KS08a, KBDK22, KAA^{+}25, KMN^{+}16, LLLE17, LMR18, LHAP06, LFHS23, NsP16, PST^{+}15a, RB24, Rus08, SBQZ14, SYC14, SVL01, THH^{+}14, TtLcC13, VW08, WR12, WYZY24, WTL^{+}16, XNH21, YJZ^{+}21, ZWFX17, ZSR^{+}05, ZXL^{+}24]. O-intensive [BPM^{+}22]. O-RAN [AIAR^{+}24]. Oak [SVN^{+}10]. Oakland [IEE84a, IEE90a, IEE91]. OAMulator [MS01]. OASIS [UBL^{+}82]. OB [XHCL15]. Oberon [WF03]. obfuscation [ZXB^{+}24]. Object [Bad82, BBD^{+}91, BP01, CAF^{+}91, Low88, PTHH14, PMC05, San88, STFH15, USE99, USE01b, BPB86, BP03, BZD17, DNR06, GSN93, IT86, LM99, VED07, WML02]. Object-Based [Bad82]. Object-Oriented [BBD^{+}91, USE99, USE01b, PTHH14, PMC05, San88, BPB86, GSN93, IT86, WML02]. Objective [GA25, GLBJ18, LPB17, AP18, BB20, BY20, DS22, FGZC23, GGQ^{+}13, GKP^{+}19, GH20, GHK24, HZL^{+}18, LZLY20, LCL^{+}23a, LCL^{+}23b, MPM^{+}20, RK18, STMV18, SL14, SM23b, SCL^{+}19, TSR19, ZLL^{+}16]. Objectives [AP22, ML78]. Objects [Qia99, ABB^{+}19b, SK13a]. Observation [NBH08, SCFP00]. observation-based [SCFP00]. Observations [LHW^{+}20]. occupied [SZ13]. OCTET [BKC^{+}13]. October [ACM03b, Ano99b, Ano06a, Boa90, IEE03, Tho93, USE00a, Vra05]. off [CGV10, MTJ^{+}22]. off-board [CGV10]. Offensive [BDJdS02]. Offers [Ano03a, Got07]. office [BRIdM10, Ano03b]. Offline [TRG13, SHLJ13]. Offloading [CL16a, GKXK13, TXD^{+}24]. offs [SIdLB15, XZK^{+}20]. OGSA [AKK^{+}07]. OGSA-DAI [AKK^{+}07]. Oktober [Müh75]. Old [Got07]. Older [SHB^{+}03]. Older-first [SHB^{+}03]. Oleco [Joo06]. On-Chip [GGM^{+}16]. On-Demand [MK13, SEF^{+}06, ZZF06, DEG^{+}17, JCZZ13]. On-Device [Bos24, XYD^{+}18]. On-Stack [WBHN18, LH13]. On-the-Fly [MKZ^{+}25, URJ18, AH24]. One [Bai70, Cre09, HPHV17, NKY^{+}18, JK15, Ste14]. one-shot [JK15]. Ongoing [Car23]. Online [BFM^{+}21, FL13a, GR15, HKLM17, HH18, HKKW13, JWL^{+}18, Joo06, KTB17, LGS^{+}23, LW20, MSC^{+}21, NG13, RG17, SZW^{+}16, SIK^{+}16, SXCL14, SCL^{+}19, XWW^{+}21, ZHW^{+}17, ZLZ21a, ZWC^{+}14, BB12, KS18b, LSS04, MPM^{+}20, NK10, THB22, ZXW16]. Online-Handbuch [Joo06]. Ontario [ACM06f, Sof83]. onto [AO16, Bak83, BS90, PS16]. Open [AFG^{+}17, AP22, GMGF24, LLWM23, LXQ24, MMdMCPSR25, SJV^{+}05, ARA20b, ARA20a, AGH^{+}15a, AAB^{+}05a, FP14, TSP17]. Open-Source [LLWM23, SJV^{+}05, AAB^{+}05a]. OpenCL [KJJ^{+}16, SXMX^{+}18, TY14, YWTC15]. OpenCL-based [SXMX^{+}18]. OpenFlow [YKS16]. OpenISA [AMB^{+}17]. OpenJDK [BFS^{+}18]. OpenNebula [KMT14]. OpenOffice [Joo06]. OpenQRM [Kar07]. OpenStack [AMIA19, BB15, BLMP22, HKJ19, YW20]. OpenSUSE [CK06g, CK06f, CK06o, CK06p, CK06p]. Operand [MSI18]. Operating [ACM75, ACM03b, BPP^{+}17, BH73, BYBYT16, CD12, Das91, HXZ^{+}16, IEE01, J^{+}05, Mar73, MNN05, MKKE12, MM94, RT93, SLM89, THB06, Vra05, ACT94, CCZ^{+}06, CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, CK06a, CK06b, CK06e, CKP78, Com00, CLDA07, Dav04, Don87, Fli77, HKD^{+}13, KSLA08, Kou11, KS20b, MW18, MDFS72, NV05, Ros06, SPF^{+}07, SS72, TT93, Vac06, Van06, WR07, WWT89, WHSE15, YK13, YLJ22, Mat10]. Operation [ZR06]. Operational [Dan12, LCMV17, Siv04, BG20, NMC18b, NMC18a]. Operations [NTL^{+}24, OLZ16, MPF^{+}06]. operator [GHM^{+}18]. Opportunistic [GJK^{+}20, KMK16, OMB^{+}15]. Opportunities [JAC^{+}19, CBFH20]. Optimal [BP99, BB12, DS19, DEG^{+}17, GSG22, HM18, HJG18, KAS13, LYL21, TCTH23, XYYY17, XRL^{+}22, ZB18, GSKJ18, HAK22, KB21, WHC16]. optimale [Sch13a]. Optimisation [SCL^{+}19, YWGH13, GKP^{+}19, PTD^{+}18]. optimise [DHD20]. Optimised [HKM^{+}18a]. Optimises [War80]. Optimistic [Pon19, WGF11]. Optimization [AGC18, CPS17, CWH^{+}16, DKW15, EGG^{+}24, GLBJ18, HO22, KC16, LW11, LKIL19, LGZ^{+}19, Man15a, MJW^{+}14, NIA18, PAC^{+}22, PS23, RK24, RRB19, SM06, SS22, SHZ^{+}14, SKT^{+}19, VG20, WDL^{+}20, WGXC25, WK90, YKM17, YLQL25, YWF09, AT23, BRS^{+}22, CLL^{+}23, EB20, FGZC23, GCARPC^{+}01, HLW^{+}13, JK13, KSS^{+}20, KSS^{+}23, KS13, KS18a, KK21, LLWW18, LZLY20, LCL^{+}23a, LCL^{+}23b, MS17, dOL12, RI25, SM23a, SPSL25, WGW^{+}18, WGY20, YXL^{+}20, ZFL^{+}23, ZLL^{+}16, ZYLY18]. Optimization-Based [SHZ^{+}14]. Optimizations [HB12, JZY^{+}22, NBK16, PDY^{+}23, RLZ^{+}16, CPST15, NG13, PGLG12]. Optimize [OLZ16, LDL^{+}08, RAT17]. Optimized [CGC16, MZD^{+}18, DS20, HZL^{+}18, KCV11, LWL16, RGS^{+}20, TMMVL12]. Optimizing [CEG07, dCCDFdO15, EG03, GKT17, HBNK24b, HHC^{+}16, JGW^{+}11, KRS^{+}17, LQW^{+}12, LL14, LXM^{+}16, MCZ06, SMK02, SAR24, SV15, WYZY24, WWL^{+}17b, ZLLL13, ZJXL11, FMIF18, HSC15, NNK21, ZLBF14, ZGL^{+}17, FLL^{+}13, HBNK24a]. Optimum [AT24]. Options [HDM08]. Oracle [SCF^{+}25, VSC^{+}10]. orbit [SSN94]. Orchestrated [MK23]. orchestrating [BRS18]. Orchestration [MVMHL24, NVV^{+}24, SCA^{+}24, ZB20, ZXR^{+}22, BSNB20]. Order [BW03, BFC02]. Ordering [HMH17, HTAY21]. ORE [OMB^{+}15]. Oregon [IEE93b, USE85]. O’Reilly [Ano97a]. Organization [BPC94, Kam83, RSGG15, Juo07, Skr01, Tho73]. Organizational [PXG^{+}17, GAHL00]. organizer [MS00, SMES01]. organizing [OK90]. Orient [IT86]. Oriented [BBD^{+}91, BWD^{+}15, BS90, CAF^{+}91, DY17, ECET18, HW12, LVM16, LYGG20, MP16, PYYG21, RSGG15, SYB12, USE99, USE01b, YLT^{+}23, ZL18a, Beg12, BPB86, Fro13, GSN93, IIK^{+}06, IT86, PTHH14, PMC05, PPO14, San88, WML02, YKP^{+}24]. Origin [Com82, Den01]. Original [BDR^{+}12]. Orthogonal [PNM^{+}20]. Orthogonally [LMG01, LMG00]. OS-Level [cCWS14, KHW^{+}16, SWcCM12]. OS/2 [Bri98]. OS/390 [DBC^{+}00]. OS6 [SS72]. OSCAR [VS06]. OSPF [SAR24]. OSS [Ble10]. Other [Den01, GC00, Mac79, KS13, Mat10]. OtOt [DKF94]. Ottawa [ACM06f]. Out-of-Band [ZSXZ07, PBYH^{+}08]. Out-of-order [BFC02].

Out-of-Process [RB01]. out-of-the-box [XHCL15]. Out-of-VM [ZFL15]. Outage [Che21]. Outline [Kee77]. Output [ACL72]. Outsourced [YDW18, CMP^{+}13, QZDJ16]. outsourcing [SASG13]. Over-Provisioning [SC18]. overbooked [LPBB^{+}18]. Overcoming [APST05]. Overcommitment [GKBB15]. Overcommitted [CWS12, WCS06, ZHHC17]. Overhead [KKD24, BJ20, BJ22, JB24, LPD^{+}11, LBL16, ZHCB15, ZLZ^{+}19a]. overheads [MST^{+}05]. Overlapped [LZM^{+}20]. Overlay [GMGF24]. Overload [KAS13, AHRR22a, LYYY18]. Overloaded [BB13]. Overshadow [CGL^{+}08a, CGL^{+}08b, CGL^{+}08c]. Oversubscription [YLT^{+}23]. Overview [Lau87, MLG^{+}02, ALW15, BB08, MNA16, NK22]. oVirt [Ano14d]. OVM [BFC02].

P [Dom80b, SSU^{+}12, Syr07]. P-Code [Dom80b]. P.R.O.S.E [Van06]. P2P [Sta07]. p5 [A^{+}04, B^{+}05, G^{+}05]. PA [ACM04b, ACM96, IEE04]. PaaS [ECET18, ZLHD15]. Pacifica [Str05]. Package [PBR^{+}90]. Packages [JMSLM92, LTT92]. Packet [VLZL16, LRP^{+}19, Ste14]. Packeteer [Ano03a]. Packing [BB17, GR15, RG17, SXCL14, XDLS15, LLZ^{+}19, SZ13]. PACO [PAC^{+}22]. PACT’06 [ACM06b]. Page [AW17, CWL^{+}15, CHLY18, KYP^{+}17, LH16, LLZ^{+}19, LZW^{+}17, LZM^{+}20, MZD^{+}18, MT16, MT17, WLW^{+}15, AJH12, BSSM08, CWC^{+}14, WTLS^{+}09]. Page-Aware [CWL^{+}15, CHLY18]. Page-level [LZW^{+}17]. Page-sharing-based [LLZ^{+}19]. Pages [GKBB15, Ano97a, JDW^{+}14]. Paging [BGM70, GHS17, HBL^{+}10, GHS16, TKG89]. Pagoda [YSS^{+}17]. Pallor [RHV17]. Palm [MS00, SMES01]. Palo [ACM01b]. Pandemic [Cap21]. Panel [G^{+}01, UBF^{+}98, BDF^{+}98]. Papers [DC15, KM13b, ACM90, G^{+}88]. PAPMSC [SDD^{+}16]. para [FMX^{+}25, LC13]. para-virtualization [FMX^{+}25]. para-virtualized [LC13]. paradigm [BD11]. PARALISP [CRZ83]. Parallax [hTMAC^{+}08]. Parallel [ACM06b, Arm78, BP99, BS90, CLM24, EGR15, FZD^{+}24, Fis01, HD16, HHK94, HTM^{+}24, IEE93a, IM93, JFPL16, JN15, KNT02, LZZ^{+}25, Loy92, LCFL12, MM92, MM93, MRG17, MM94, NOT^{+}17, PAC^{+}22, PY93, SSN94, TVO92, WCC16b, Wat86, Wat87, Wel94, YP15, ZRZY15, ZWZ20, AS14, AGIS94, BPC94, Bir94, BL90, BFC02, BB95, CARB10, Cav93, CDM^{+}10, dCCDFdO15, CRG16, CKP^{+}93, DKF94, DDS^{+}94, DM93, EF94, FM90, GSN93, HTAY21, Hol95, JGA^{+}88, KJLY15, KSS^{+}18, Kra90, Les74, LG93, McK11, MRG18, MN91, NOR15, NG13, Pou90, RH17, RSW91, She91, SL00, Taf11, WK08, YC98b, YYC^{+}19, Ble89, JPTE94, YC98a]. Parallelism [BYZZ20, HC18, YTS14]. Parallelization [LYL21, vKF13]. Parallelizing [SSL^{+}13]. Parallels [Tho08]. parameter [Kha19]. parameters [WWL^{+}25]. parametric [PULO16, UTO13]. Paranoid [Bau05, Bau06b, Bau06a]. ParaSail [Taf11]. Paravirtual [KMN^{+}16]. Paravirtualization [AD11, SBQZ14]. ParCo93 [JPTE94]. PARD [MSS^{+}15]. ParDMCom [M^{+}06]. Pareto [DS20, EB20]. PAROS [MM94]. PARS [CWL^{+}15]. Parser [UOKT84]. parsing [Kha19]. Part [Cre09, HO92, NSC^{+}22, RGSJ17, Sch94b, Sch94a, Cre08a, SS72, Zyt94a, Zyt94b]. Part-of-Memory [RGSJ17]. Partial [BWD^{+}15, FTNY69, KLY20, LYL21, ZLL^{+}20, WGF11, WWH^{+}17]. partiality [Dan12]. partially [HH13]. particle [AT23]. Partition [Int06c, LLS^{+}08]. Partition-based [LLS^{+}08]. partitioned [Van06]. Partitioning [Bad87, Ian14]. Partitions [Int06b, SJRS^{+}13]. Party [CRZH15]. Pascal [Har77, GBO87, SP83]. Pass [PYDG22, XYD^{+}18, PDC^{+}12, YLWH14]. Pass-Through [PYDG22, XYD^{+}18, PDC^{+}12, YLWH14, MLA83]. passé [BC10]. Passing [Fra98, GGM^{+}16, DM93, TO91, UR15, XH90]. Passthrough [XD16, XD17, PRS26]. Password [CD12]. Past [Sup04, Var91, BJG19, BS96, JKDC05]. PASTE’01 [ACM01a]. patches [Ano07, PDML25]. patching [PM19a]. Path [GR20, AM16]. PATHWORKS [Nou92]. Pattern [CFM17, HPP15, YDW18, ZDLG17, OK90]. Pattern-Aware [HPP15]. Patterns [Aln22, CL17a, ESY^{+}17, PMC05]. Paving [FLZ^{+}20]. PaVM [FZD^{+}24]. Paxos [HMS17]. PBS [ZLL^{+}20]. PC [ACM04a, GBO87, Mon97, Voe86]. PCI [YLWH14]. PCs [Ros99]. PCVM.ARIMA [CSSE21]. PDB [HHH04]. PDCE [M^{+}06]. PDP [Gal73, GBO87, Ham76, PK75a, SP83, She02]. PDP-10 [Gal73]. PDP-11 [GBO87, Ham76, PK75a, SP83]. PDP-11/40 [GBO87]. PDP-11/60 [SP83]. PDP-8 [She02]. PDS [AAB^{+}05b]. Peak [LTE12]. Pedagogy [CLKEF21]. PEMU [ZFL15]. penguin [Bau05, Bau06b, Bau06a, Fab13]. Pentium [RI00]. Perceiving [XWH^{+}16]. perception [MW18]. Perfctr [NB11]. Perfctr-Xen [NB11]. performability [EBJ17]. Performance [ACM98, ACM04b, Ano03b, AD11, Bad82, BPM^{+}22, BL90, Cal75, CFH^{+}79, CFH^{+}80, CGS06, CHW12, CZH^{+}25, DLLN18, De 06, DSZ11, EDS^{+}15, GE85, Gua14, GKBB15, HSK17, HTB19, Hor73, HB12, IEE96b, IEE06a, IN87, IBBA20, JR02, JK13, dCJR16, KCWH14, KS08a, KS20a, KMM13, KP15, KKS^{+}19, KD78, LZ15, LGJZ16, LCK11, LMR18, LMG01, LCT^{+}15, LXW^{+}23, LHAP06, LTZ^{+}14, MJW^{+}14, MT16, MT17, MLG^{+}02, MMHP17, MBK^{+}92, NBB^{+}19, NMS^{+}14, Oak14, OBSR16, PZW^{+}07, Pat12, PDY^{+}23, PNT12, Raj79, RCM^{+}12, RP07, RB24, SHW^{+}15, SD01, SAR24, SCSL12, SDD^{+}16, SLC20, SJA^{+}17, SM92, SP22, SM02, TSN^{+}23, THC^{+}14, URJ18, UT87, VP16, Vog03, WDL^{+}20, WKT08, WCC16b, WYZY24, WWL^{+}17b, WZL^{+}23, XLJ16, YC98a, You73, YWCF15, ZLSI17, ZXL^{+}24, ZRZY15, ZWL^{+}18, ZTA^{+}21, ZJXL11, dGG^{+}17, AKK^{+}07, AAH^{+}03, AA24, AGH^{+}16, Ano96]. performance [AWR05, ASB18, BML^{+}13, BDQR23, BB12, BJG19, BJ22, BBM09, BMER14, CBGM12, CBZ^{+}16, CCW^{+}20, CMP^{+}07, DQR^{+}13, DLL^{+}16, DSSP06, DLH^{+}20, DYL^{+}12, EMS15, Fit14, FF96, GP13, G^{+}01, GVI13, G^{+}05, GAH^{+}12, HKJ19, Han16, HHSG18, Hog02, HC12, HL13, KBK^{+}21, KZ21, KKJL14, KL13, Kou11, KCV11, LBZ^{+}11, LLLE17, LM99, LMG00, LL14, LQD^{+}18, MCC18, MA10, MST^{+}05, MUKX06, M^{+}06, MMG^{+}18, MW05, NB11, OL13, PJZ^{+}19, PRS26, PV08, QXH18, RHR02, RAP19, RQD^{+}17, Rix08, RGS^{+}20, RCTY19, SENS16, SE12, SBNU18, SP83, SEPV19, SB10, SPF^{+}07, SYC14, SPAK18, TZY^{+}25, TIIN09, VW08, WTL^{+}16, WWH^{+}17, XJW^{+}18, XZK^{+}20, YC98b, YZLQ14, YIR24, YQZ14, YQZ19, ZYZ^{+}18, ZSR^{+}05, ZSW^{+}06, ZFL^{+}23, ZLCZ18]. performance-aware [ZFL^{+}23]. Performance-Based [CHW12]. Performance-directed [RP07]. Performance-Guaranteed [ZWL^{+}18]. performance-optimized [RGS^{+}20]. performance-to-power [DLH^{+}20, KZ21, RCTY19]. Performances [GMGF24, KBBM15]. performing [BB08, GBCW00]. performs [Ven97d]. period [B^{+}07]. Periodic [LD05]. periodical [YQZ14]. Periods [RB17]. peripheral [VWT13]. Peripherals [BG74]. Persistence [SCD90, PNM^{+}20]. Persistent [GH91b, Low88, SMES01, SXH^{+}19, ZCL^{+}21, LM99, LMG00, MS00, PNM^{+}20, LMG01]. Personal [Hir92, LBP^{+}07]. Perspective [FLZ17, Han16, LCZ^{+}19, RSGG15, SMP22, FP14, LDDT12, PAKY16, Wal10]. perspectives [MA10]. Pervasive [HHH04, BTLNBF^{+}15b, HH05]. Petascale [Gei02]. Pete [Gal09a, Gal09b, Gal11]. PEVM [LMG00, LMG01]. Phantasy [RZPX19]. phase [JK13, SZKY21, TF16, ZL13]. phases [RHR02]. Phi [GGK19]. Philosophy [Com65]. Phoenix [ACM03a]. Phosphor [BK14]. Phylogeny [ASPP22]. Physical [BBM^{+}15, PS16, WLW^{+}17, AAM^{+}16, PTD^{+}18, XZK^{+}20, YLJ22, vCPWvT11]. physics [GTN^{+}06]. PI [RNS^{+}23, DCA17]. Piccolo [CHPY17]. PicoJava [MO98, TO96, OT97]. Picojava-I [OT97]. Pin [ZFL15]. Pioneer [War11]. Pipelines [RKRK17]. PIPPIN [DH01]. Pittsburgh [ACM96, ACM04b, IEE04]. PL [SKC73]. PL/EXUS [SKC73]. Place [USE01a, Fab13]. Placement [AAAF21, AAR22, BYZZ20, BMJ^{+}22, CGC16, GLBJ18, HTM^{+}24, HYL^{+}25, JQWG15, KP15, KAS13, LPSS19, LKIL19, LTE12, LLWM23, LYS^{+}18, LLS^{+}20, LPB17, MK13, Man16, MSLY24, MZ20, PAC^{+}22, PHXL19, PF23, RK24, SJ21, SSS25, SHZ^{+}14, WZH^{+}16, WCC20, WGXC25, XZL^{+}20, YWY^{+}17, YLQL25, ZWL^{+}18, ZHL16, dSdF16, AD19, AJBJ23, AHRR22a, ATZP21, BB20, CL17b, CCL^{+}20, DS19, EMS15, FLL^{+}13, FZS^{+}20, FS19, FF22, FA21, FMIF18, FMJ15, GGQ^{+}13, GH20, GHK24, GA18, HM18, HZL^{+}18, HAK22, IKU15, JC18, KN18, KBDK22, KHL17, KSO^{+}15, KK21, KHA22, LKR^{+}19, LBZ^{+}11, LZWD15, LLWW18, LZLY20, LCL^{+}23a, LCL^{+}23b, LPBB^{+}18, MNB24, MS17, Man18, MNA16, MHM19, EYGS19, Pon19, PS23, RKT20, RWC21, RK18, RJK^{+}17, SZKY21, SBI21, SM23a, SM23b, SP23, TMLL14, TSR19, THB22, TMMVL12, WWL^{+}25, WHW20, XTB17, YPLZ17]. placement [YB24, ZWHC17, ZLW^{+}19a, ZLL^{+}16, ZWH^{+}17]. placing [XWW^{+}21]. PLAN [CTP^{+}17]. Plane [GGK18, AMIA19, LRP^{+}19]. Planes [UVL^{+}13]. PlanetFlow [HBP06]. PlanetLab [MPF^{+}06]. Planning [IBM72, IBM73, Hal08, MIS^{+}05]. plans [Kal97, Lot91]. Plant [BYZZ20]. Planung [Zim05]. Platform [Bos24, CMK^{+}16, DHPW01, DMG^{+}15, Fra09, GWZ16, GPW03, HCB18, JXL^{+}12, JJ02, MCE^{+}02, PPS^{+}18, SML18, Sun99, TCP^{+}17, VGF16, WL96, Wal99, WBHN18, ZSP^{+}21, AMB^{+}17, BBD^{+}10, DJP^{+}24, CSMB15, CLM24, DCA17, Fra06, MW18, PW03, WQG15, WCC^{+}16a, WLG^{+}11, XZ11, YJZ^{+}21, YMY17, Ros99]. platform-independent [PW03]. Platforms [AMA18, ASMA21, Ano06a, AIAR^{+}24, BDG18, FR24, GLS15, RNS^{+}23, SN05b, Uhl06, YP15, BSL^{+}18, BSD19, DPW^{+}09, GLK^{+}12, MRM06, MBBS13, NV05, SWH^{+}13, SBP^{+}17]. Player [Joo06, Zim06]. Plex86 [Law00]. Pliant [KDB16]. Pliant-based [KDB16]. plotter [MSCK92]. plug [Kag09]. plug-in [Kag09]. Plural [UT87]. pocket [BBD^{+}10, FFB^{+}00]. POF [DS20]. POF-SVLM [DS20]. point [XJW^{+}18]. pointers [AT16]. points [TLX17]. points-to [TLX17]. Policies [CC77, KC12, NMMP15, KBDK22]. Policy [CTP^{+}17, EMW16, JFPL16, LDRS18, SL14, JFZL17, SZKY21]. Policy- [CTP^{+}17]. Policy-Compliant [LDRS18]. Polling [PYDG22]. polymer [NRS92]. polymorphism [PULO16, UTO13]. pooling [WRSvdM11, WRS^{+}15]. POPL [ACM99]. POPLOG [SSG90]. Port [DBMI92]. Portability [Hir92, JR02]. Portable [HWB03, Ibs84a, SMK02, Ibs84b, FCG^{+}05, HK07, LTK17, AEMWC^{+}12]. Porting [Caa00, JJ91, Kel06, MB98, Shi03, vdK09]. Portland [IEE93b, USE85]. position [Hin97]. posium [USE01c]. Possible [RB24]. Post [AGJS16, HDG09]. Post-Copy [AGJS16, HDG09]. Postroom [Osb01]. Potential [CLKEF21, FRD^{+}08, Got07, JK13]. Pour [Han73]. Power [AT24, AAM^{+}16, CLL^{+}23, DSM14, GPM21, HSK17, KBB11, KL14, LZ15, LGJZ16, LLLE17, MAK18, MV16, MJW^{+}06, PLZ20, RSNK17, RSN^{+}18, SSN12, SDD^{+}16, Sta07, VWT13, XDLS15, ZWL^{+}18, CBGM12, CMP^{+}07, DLH^{+}20, EBJ17, FLL^{+}13, HH18, HH19, IMK^{+}13, JKK^{+}13, JNR12, KZ21, KK21, NS07, RHZ^{+}17, RCTY19, SP23, TDG^{+}18, TUM18, THC^{+}14, WRS13, XHL^{+}13, YZLQ14, YLHJ14, YLCH17, YW20, A^{+}04, B^{+}05, G^{+}05, MBBS13]. Power-Aware [SDD^{+}16, ZWL^{+}18, KBB11, JNR12, RHZ^{+}17]. power-capping [JKK^{+}13]. Power-Efficiency [AT24]. Power-efficient [AAM^{+}16, LLLE17, SSN12, KK21]. POWER5 [AAB^{+}05c]. PowerPC [But94]. ppXen [ASB18]. Practical [BJH^{+}16, DLX^{+}17, HN10, Kna93, PDY^{+}23, WLW^{+}15, WBHN18, WWH^{+}17, FIF^{+}15, PJZ^{+}19, SNV10, TC10, Wün13]. Practice [Bec09, Cre08b, Lar09, SHB^{+}03]. Practices [MO98]. Praxis [Bec09]. Praxisbuch [Lar09]. Praxisführer [Bor01]. Pre [LUL^{+}05, HBNK24a, HBNK24b]. pre-copy [HBNK24a, HBNK24b]. Pre-virtualization [LUL^{+}05]. Precedence [EGR15]. Precedence-Constrained [EGR15]. Precise [LJFS17, BHSB14, CCW^{+}20, TLX17]. Precision [ADM98, BKMM87, KKS^{+}19, ZXG^{+}24]. preconditioned [MM92]. Predicate [UOKT84]. predicates [JKDC05]. Predictable [KR18, LTE12, XLJ16, LTK17, HK07]. predicting [WQG15]. Prediction [EVCL21, FFM^{+}23, HM20, LWC^{+}17, MB15, ZDLG17, ZFH^{+}22, ADA^{+}19, AA24, BKT^{+}19, CEG07, CCW^{+}20, EYG21, EG03, HBNK24a, HBNK24b, HLBZ20, KJM^{+}07, KCV11, PTD^{+}18, RGAT18, Raj79, SOKE23, SSN94, WCZ^{+}23]. Prediction-based [HM20, EYG21]. Predictions [BFM^{+}21]. predictive [CSSE21, XCJ^{+}14]. Predictor [BSMF08]. Preemptable [OL13]. Preempted [OLZ16]. preempting [SJB14]. preemption [YQZ14]. Preemptive [PG17, PG18, YXL^{+}20]. Preferences [AAAF21]. Preferred [Par72]. prefetch [KW13]. Prefetching [RZPX19]. Preliminary [HW93]. prep [IIPB09]. PreScheme [Ram93]. Presence [KBK^{+}21, CFG^{+}13, CJJ^{+}22]. Present [Var91, JKDC05, Yur02]. presented [ACM90]. Preservation [JE12, BB08]. preserve [STFH15]. Preserving [BS96, DNR06]. Presidio [Str05]. pretenuring [BOF17]. Prevent [KLY20, SYB12]. Preventing [DL19b, Kip21, WLCS17, PRB07]. prevention [MA17]. previous [STFH15]. Price [FFM^{+}23, WHC16]. Pricing [SAJ24, ADA^{+}19, DEG^{+}17, TXD^{+}24]. Primary [PP16]. Primitive [LCWB^{+}11, BMWB86, Pou90]. PRIMITIVES [Ble89]. Princeton [FS11]. principled [WSAJ13]. Principles [ACM75, ACM99, ACM03b, Gol73c, Juo07, PJZ18, SHW^{+}15, Vra05, SS72]. Prioritized [FBM^{+}21]. priority [OKAM17]. Privacy [IEE84a, IEE90a, IEE91, WLL^{+}13]. Private [GMGF24, HWR^{+}24, HW12, MMHP17, SCC^{+}23, Nie12, SYMA17, TUM18, WH08, ZLW^{+}19a, Fro13]. Privileged [MPF^{+}06]. Pro [SRS09, Fra06, Fra09, Wil06]. Proactive [MZ20, WB16, BKT^{+}19, CFRSSR19, IRB19, KU24]. Proactively [GKBB15]. probabilistic [PKS^{+}19]. probability [LYYY18]. Problem [AAR22, BL17, BFG^{+}14, GWZ16, Man15a, GLW23, LY23, MM92, EYGS19, SL00, XA22]. Problems [GR20]. Procedures [LSX^{+}24]. Proceedings [ACM96, ACM97, ACM99, ACM04b, ACM05b, ACM06a, ACM06b, Ano99b, Boa90, IEE96b, LCK11, USE99, USE00a, USE00b, USE01a, USE01b, ACM00, ACM03b, ACM05a, ACM06f, Ano93, GHH^{+}93, HHK94, IEE85, IEE04, JPTE94, Mat10, MR91, SS05, USE85, USE86, Vra05, ACM75, ACM81, ACM89, ACM90, ACM01b, RM03, ACM04a, ACM05c, ACM05d, ACM06e, ACM06c, ACM06d, Ano01b, Ano04b, Ano06a, BW03, IEE84b, IEE84a, IEE90a, IEE90b, IEE91, IEE92, IEE93a, IEE93b, IEE05, IEE06b, IEE06a, MS91b, Ost94, Sof83, Shr89, Tho93, USE91, USE93, USE01c, USE02, USE06, M^{+}06]. Process [AGLM91, Bal91, ETAB22, HPHV17, MZG14, RB01, SC17, Tho93, AC95, LZWD15, EYGS19, PAKY16, PTD^{+}18, XCJ^{+}14, YIR24]. process-aware [XCJ^{+}14]. Processes [JADAD06a, Kim84, SOAK23, SN05b, FA21, WT91]. Processing [ASPP22, DKW15, GLL^{+}21, Loy92, MMdE19, NTL^{+}24, VLZL16, DH01, EF94, GSN93, IM93, KHL17, KWZ^{+}19, LKY^{+}17, LRP^{+}19, LMDP19, LG93, MMG^{+}18, WWT89, Wün13, ZDK^{+}19, ZGL^{+}17]. Processing-in-Memory [NTL^{+}24]. Processor [BBB^{+}25, ISE08, NSL^{+}06, RWX^{+}12, SKJ^{+}17, BKR20, IIK^{+}06, LRC05, VdlFCC97, WDSW01, WLL^{+}13, WJGA12]. Processor-Interconnect [SKJ^{+}17]. Processors [DSM14, Gei02, MT16, MT17, MBK^{+}92, PDL^{+}23, PNT12, RTL^{+}18, KKC^{+}16, MN03]. product [IBM88, Int88, SV17]. production [SL00]. Products [Ano03a, Ano03b, Ano05]. Professional [vH08, IIPB09, Ham07, Khn09]. professionellen [Zim05]. Profile [WKJ20, AWR05, WKJ17]. Profiler [SH04, VL00]. Profiles [Int05b]. Profiling [LV99, Sun95a, YWW^{+}15, DSZ11, NK10, SSB^{+}14a, STY^{+}14, TZK17, TSN^{+}23, THC^{+}14, YZLQ14]. Profiling-Based [YWW^{+}15]. Profit [BYBYT16, MLXG19, ZHW^{+}17, LWLL16]. Profit-Maximizing [BYBYT16]. Profitability [WUK^{+}18]. Program [ACM01a, Com65, Cre65, FTNY69, Han05, HB08, MSG01, SZ88, ABDD^{+}91, BPB86, Olb78, She02, WGF11]. Programm [Mar08]. Programmability [EMW16]. Programmable [Bos24, DCG12, DMS02, FS11, Ken80, Kov19, MSS^{+}15]. Programmer [PSBG11a, PSBG11b]. programmers [Hee07]. Programming [ACM90, Arm78, DK75, Eng99, Gai75, GMP89, GH91b, LFBB94, Luc97, SYB12, Sub08, Sub11, Tho68, Tol98, ACM99, AS85b, Alf91, BCM90, CPM^{+}18, Ham76, Jou85, Kag09, ME87, MRG18, RSW91, SMO84, Tai98, AS85a]. Programming-in-the [DK75]. programming-in-the-small [DK75]. Programs [FS12, Kam83, NMMP15, Wel94, CK06b, CK06e, CRG16, DKF94, EGD03, GMR93, IM75, Kee68, Wak99, Wol99]. Progress [ZRD^{+}15, ZHCB15]. project [AAB^{+}05a, CKP78, Lot91, RD90]. projects [AL05]. PROLOG [Clo85, Ode87, War80]. Promenade [CFLL19]. Prometheus [ACG18]. Promoting [ACA16, WLW^{+}17]. Proof [FC98, LLZ18, Arv02, FP14, FCG^{+}05, ZLH^{+}15]. proof-carrying [FCG^{+}05]. Propagation [AD11]. Properties [BN75, BSD19]. property [VT14]. proportional [FMX^{+}25]. Proportionally [CFLL19]. Proposal [EVCL21, SLZ^{+}25]. proposed [GH91b]. Prospects [PCB^{+}18]. Prospero [Neu92]. protect [ZBP07]. Protected [BPP^{+}17, Cof99, GHD12]. Protecting [LMJ07, WTM18, WLL^{+}13]. Protection [VMW^{+}19, CD12, CDD13, SS75, CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, CBFH20, JCZZ13, PK75b, TSLBYF08, WJGA12]. Protectit [KSLA08]. Protocol [GKXK13, MN91]. protocols [DM93, RSLAGCLB16]. Prototype [Sim92]. Prototyping [SXMX^{+}18]. Provably [GNDB16, ZLG^{+}20]. provenance [RG19]. Proverb [Fer11]. Provers [Hir17]. Provide [ESY^{+}17, WJ10]. Provider [LPSS19, MP16, CFRSSR19]. providers [EdPG^{+}10, SG10a, TMMVL12]. Providerwahl [Joo06]. Providing [BDS^{+}09, Bos24, HC18, Isl19, KHW^{+}16, KKH14, KGZ^{+}04]. Proving [BW03, IM75]. Provision [LGXC23, WZH^{+}16, WN17]. provisioner [JNR12]. Provisioning [BSSS14, BCW20, ELC^{+}19, HJG18, LCT^{+}15, LWC^{+}17, LLZ18, NMG15, NSJ12, SC18, SZW^{+}16, SXCL14, XLJ16, ZLW^{+}14, ZRS^{+}16, CSSS11, CFVP12, FGG14, GSKJ18, KBB11, MHS21, PPO14, SJB14, VOS12]. proxies [Eug06, STFH15]. Prune [BYZZ20]. PS3 [Sta07]. pSeries [Mly09]. pseudo [ABDD^{+}91]. pseudo-random [ABDD^{+}91]. PSO [AT23, LW12]. PSO-Based [LW12]. Pthreads [RMB02]. Public [MB20, OG16, SDS^{+}21, WUNK17, WUK^{+}18, FBZS12, PKS^{+}19, ZLV^{+}12, ZBS^{+}15]. publications [Mat10]. Publisher [LCL^{+}23a]. Purpose [BKM25, GFB^{+}92, ZDS^{+}22]. Purposes [BHI15, BSL^{+}18, WDSW01, WO75]. Putting [LPSS19]. PV [Ano15]. pyramid [MJ93].

Q [AJBJ23, Che21]. Q-Learning [Che21, AJBJ23]. QEMU [WR07, WR08, CK06a, CK06b, CK06e, CK06c, CK06d, CK06g, CK06f, CK06i, CK06h, CK06j, CK06k, CK06m, CK06l, CK06n, CK06o, CK06p, CK06q, CK06t, CK06r, CK06s, Bar06, CJK^{+}12, MZG14, WR07, WR08, vdK09, CK06a, CK06b, CK06e, CK06c, CK06d, CK06g, CK06j, CK06k, CK06m, CK06l, CK06n, CK06q, CK06t, CK06r, CK06s, Deu08]. QEMU-KVM [CJK^{+}12]. QM [Fli77]. QM-1 [Fli77]. QoE [KS18a]. QoS [FAA17b, BAC15, DXM^{+}17, FAA17a, HLPY16, KN18, KCY22, KP15, LCL14, LWL16, LYGG20, PS23, WZH^{+}16, WWL^{+}25, XZL^{+}20]. QoS-Aware [WZH^{+}16, XZL^{+}20, KN18, LWL16]. QoS-Oriented [LYGG20]. QoS-Satisfied [KCY22]. qualitative [ALW15]. Quality [BB13, LLY^{+}25, MHS21, SV13, VOS12, WKJ20, CMG^{+}19, LYY^{+}20, NZH20, TDD20, WKJ17, XXWG23]. Quality-Aware [LLY^{+}25, LYY^{+}20]. quality-of-service [NZH20]. quantification [BKH^{+}06]. quantify [TZK17, TDG^{+}18]. Quantifying [FFB^{+}00, PJZ^{+}19]. Quantitative [YZW^{+}13]. Quantization [ZXG^{+}24]. Quantum [NLD^{+}23, CLM24]. Quelle [LC09a]. Quemu [CK06o]. Query [WK90, KHL17]. querying [CKRJ17]. queue [HZK^{+}25]. queuing [Pon19]. Quick [NOT^{+}17]. QuickDedup [SSG^{+}20]. QUICKTALK [BMWB86]. QUIS [CKRJ17].

R [Fro13, KMMV14, Vit14, Wün13]. R2 [Bod10, KS10, Apr09, Bod10, Car14, Gal09b]. Raccoon [ZWFX17]. race [HHPV15]. races [DKF94, PRB07, WCG14, XXZ13]. Racket [FDD^{+}19, KFF12]. Radio [AAJD^{+}16, MMdMCPSR25, SLZ^{+}25, SKT^{+}19, YI24, LJR12]. Radio/MEC [MMdMCPSR25]. RAID [SPAK18]. railway [FP14]. Rain [HS19]. RAM [GGJ^{+}92]. RAN [AIAR^{+}24, YIR24]. Random [CFLL19, ABDD^{+}91, Fer11]. randomized [JGA^{+}88, KFF12]. randomness [RY10]. range [HP77]. ranking [RAP19, SS19]. ranking-based [RAP19]. Rapid [But94, GMK17]. rapidly [BSM^{+}12]. Raspberry [RNS^{+}23, DCA17]. Rate [CFLL19]. Ratio [WDL^{+}20, DLH^{+}20, KZ21, RCTY19]. RBPSO [WGY20]. rCUDA [CPM^{+}18, PRS16, PS19b, RSC^{+}15, SIRP17]. RDMA [PST^{+}15a]. RDMA-capable [PST^{+}15a]. Re [MKM^{+}08]. Re-engineering [MKM^{+}08]. reachability [KY16]. reaction [KK21]. Reactive [DSM^{+}18, Kol19, NMMP15]. Read [MJW^{+}14]. Read-Performance [MJW^{+}14]. Real [AAR22, AE01, BE17, Ben21, CW03, CRX24, Cla97, FXHY21, GPM21, HcC14, JAD19, JYM^{+}23, KR18, LSSC22, LXL^{+}22, LD05, Mac79, Mat09, NL19, PPG^{+}17, QCM^{+}24, QT06, Ran20, Sta97, SCA^{+}24, Swa06, ABB19a, AS76, ABC^{+}07, BCC^{+}15, HK07, Ive03, KBB11, LTK17, NBS18, Nie12, PTD^{+}18, RK18, SBNU18, TZY^{+}25, WQG15, YCL^{+}19, ZEdlP13]. Real-Time [CW03, CRX24, FXHY21, GPM21, JAD19, JYM^{+}23, KR18, LSSC22, LXL^{+}22, NL19, PPG^{+}17, Sta97, HcC14, LD05, QCM^{+}24, QT06, SCA^{+}24, ABB19a, AS76, ABC^{+}07, HK07, Ive03, KBB11, LTK17, NBS18, PTD^{+}18, SBNU18, TZY^{+}25, WQG15, YCL^{+}19, ZEdlP13]. Real-World [AAR22, Ben21]. Realism [DSSP06]. realistic [CKP^{+}93]. Reality [BG20, FBGS24, CB07]. Realization [SJMG24]. Realizing [UT87, Syr07]. Reallocation [LWZ^{+}18, BY20]. RealNetworks [Ano03a]. Reap [HPHV17]. reasonable [KJ13]. reassignment [STMV18]. rebalancing [LZLY20]. Rebuilding [FDD^{+}19]. RECAP [Ben21]. Receives [War11]. Rechenzentrum [See08a]. Rechenzentrums [Mar08]. recipes [Car14]. Reclamation [Bad82]. recognition [KKM^{+}13, OK90]. Recommendation [XLL^{+}20, PAKY16]. Recompilation [THL03]. Reconciling [KPHA20, ABG14]. Reconfigurable [BHI15, IBBA20, KGS16, SML18, STY^{+}14, UVL^{+}13, ZL18a, FX06, HH13]. Reconfiguration [MKZ^{+}25, MDGS98, QLL^{+}21, ZWZ20, JES^{+}15, LJR12]. Reconsidered [Sta07]. reconstructed [AD18a]. Reconstruction [ASPP22, Sch13b]. Record [JKB15, IEE96a]. Record/Replay [JKB15]. recorder [LBP^{+}07]. recoverability [KY16]. recovering [LRC05]. Recovery [KKLV16, AAF^{+}09, BGS13, CHCC07, FL13b, Kou11, MSI^{+}12, STFH15, Tay76, ZXW16, BBMA91, Mar08, MSS91]. Recurrent [LCZ^{+}19]. Recursion [War80]. Recursive [BN75, LW73, FHL^{+}96]. Red [G^{+}06]. Redefining [CGS06]. RedHat [Gal11]. redirecting [JYW^{+}13]. Redirection [FL13a, LYS^{+}18]. redistribution [KNHH18]. reduce [FLL^{+}13, GPS^{+}18, LCLG26, LLZ^{+}19, RJK16]. reduced [VED07]. Reducing [ELC^{+}19, HPHS04, Hu90, HS06, KSS24, KY16, LBZ^{+}11, MV16, PLMA18, SC18, ZLZ^{+}19a, ZLW^{+}19a, KJM^{+}07, MA21, MA24]. Reduction [JJK^{+}11, Wat86, Wat87, ZHL16, HCJ07, LJYZ15, TDG^{+}18]. Redundancy [Tay76, WTJR22, GLV^{+}10]. redundant [KJJ^{+}16, ZWH^{+}17]. Reference [Ano03a, CRZ83, Hal79, HPP15, LC09a, XWX^{+}17, YTY00]. Referenz [LC09a]. Reflection [FPS^{+}02, ORPS09]. Reflections [MLA83]. Reflective [CGMD19]. region [HLW^{+}13, LXRS19, YC16, vKF13]. region-based [YC16, vKF13]. Register [CK87]. registers [SCEG08]. Regular [Cox07, Cox09, Cox10, Cox12, KP99, Tho68]. Regularity [DPCL22]. Regularity-Based [DPCL22]. reification [RRB17]. Reincarnation [Ros04]. REINFORCE [KLR^{+}20]. Reinforcement [HTM^{+}24, HPS22, MSC^{+}21, GHK24, Par25, QBL^{+}23, WZZ^{+}20]. Reinventing [Hof20]. Rejuvenation [AJM^{+}25, SAT09, AMA^{+}14, CRX24, MNT14, TUM18]. Relation [KLLT18]. Relational [WK90]. Relationship [Mal73]. Release [IBM73, IBM94, IBM96]. Releases [Ano03a, Ano03b]. relevant [NP13]. Reliability [BCG73a, BCG73b, ESY^{+}17, FZS^{+}20, HXZ^{+}16, XH16, MD74]. Reliability-aware [FZS^{+}20]. Reliable [PEC^{+}14, THB06, YWY^{+}17, Car14, SHR19a, SHR19b, SPSL25, Van06, WQG15, WXW15]. Reliably [TCP^{+}17]. relocation [KJLY15, MR23]. Remaining [XLWX19]. remapping [AS14, LJL12]. Remote [FLM^{+}08, JKB15, JHS12, KBC21, KMN^{+}16, Bor07, CPM^{+}18, CMGI^{+}23, GCARPC^{+}01, RSC^{+}15, RS16, SIRP17, SWW^{+}18]. Remoting [MGL^{+}17, SM23b]. remoting-based [SM23b]. removal [WGF11]. Remus} [dSOK17]. RemusDB [MRC^{+}13]. Renaissance [FDF05]. Rendezvous [SM92]. renewable [KTB17]. Renewal [WN17]. ReNIC [DCP^{+}12]. Reno [ACM89]. rental [FBZS12]. Repair [SEK^{+}19]. repeatability [Vit14]. Replacement [GHD12, WBHN18, LH13, uRQS20]. Replay [BJH^{+}16, JKB15, KM13a, KM13b, RTL^{+}18, SCFP00, CLG^{+}10, WXZ^{+}17]. Replaying [WKG17]. Replica [GLBJ18]. Replication [CWL^{+}15, LJL^{+}11, DCP^{+}12, KJJ^{+}16, LMV12, dSOK17]. replications [CBJ22]. reply [DM76]. Report [Ano01a, Ano02, Ano04a, CBLFD12, FDD^{+}19, Int06c, Int06a, PBAM17, Pul91]. repository [AWR05, GKP^{+}19]. representation [IT86]. representations [dCJR16]. reproducibility [Vit14]. Reproducible [MB20, Boe15]. reproducing [PTM^{+}15]. Request [LYS^{+}18]. Requests [MLXG19]. Requirement [YWR^{+}14]. Requirements [AP22, Gol71a, LCMV17, PG74, SCC^{+}23, BG20, PG73]. ReRanz [WWL^{+}17a]. Research [AP22, AEA^{+}23, AAB^{+}05a, Ano00, Ano01a, Ano01b, Ano02, Ano04a, Ano04b, Boa90, CLKEF21, Cre65, DMS02, IEE90a, IEE91, Kim84, Ten17, USE01c, USE01d, USE02, ARA20b, ARA20a, AGH^{+}15a, ADWM18, BJG19, Boe15, CBLFD12, Gol74, Her10, SVN^{+}10, Vit14, ZJRW19, HMS17]. ReSeer [WXZ^{+}17]. Reservation [HC18, ZWC^{+}19]. reservations [THG^{+}18]. reserved [DEG^{+}17]. reserving [YLJ22]. reset [RY10]. Reshaping [BHI15]. Resident [WK90, LF19]. Residual [AT24, AT23]. Resilience [NTR18, OMB^{+}15]. Resiliency [KLR^{+}20]. Resilient [CFB24, VS19, BGS13, OMB^{+}15, TDG^{+}18]. Resistant [THB22]. resistive [JAC^{+}19]. resolution [GE85]. resolving [ZWC^{+}14]. Resource [AJ18, AAMBE21, AIAR^{+}24, BKT^{+}19, BBMA91, BL17, CMK^{+}16, ECET18, EVCL21, FDF05, GWZ16, GLS15, GA18, HC17, HO22, JZY^{+}22, JSHM15, KCY22, LZWC13, LCT^{+}15, LLY^{+}25, LCFL12, MMdMCPSR25, MSS91, MBA^{+}12, MVMHL24, PFPJ18, RG17, SJB14, SC17, SC18, SZW^{+}16, SXCL14, SCA^{+}24, Sur01, WIS^{+}15, WGXC25, XSC13, YSS^{+}17, ZQCZ16, ZLG^{+}20, ATS16, AS14, BSOK^{+}20, Car06, CEPR22, CMP^{+}13, EdPG^{+}10, Fu10, GPR23, HZZ^{+}14, HH19, JWH^{+}15, JC18, KF18, LC09b, LYYY18, LLZ^{+}19, LLS14, MR23, MB21, MS01, Mly09, NBS18, PKS^{+}19, RGAT18, SBNU18, SGV13, SGV12, SOKE23, TV18, VSMC23, VVB13, Wal02, WDCL08, WGY20, WWL^{+}25, WB16, WSVY09, YGLY21, ZWC^{+}19, ZB18]. Resource-aware [GA18, PFPJ18, SGV12]. resource-constrained [TV18]. Resource-Efficient [LLY^{+}25, VSMC23]. Resource-Latency [BL17]. Resources [CRZH15, ELC^{+}19, HLPY16, KGS16, MK13, NHL22, PCC^{+}16, RK24, SDS^{+}21, ZB20, HMH17, HKJ19, KHL17, LTZ^{+}14, OKAM17, PSZ^{+}07, TZK17, WRSvdM11, WRS^{+}15, ZBP07]. Resourcing [MSS^{+}15]. Resourcing-on-Demand [MSS^{+}15]. Responding [BSM^{+}12]. Response [BE17, WZKP19, MA21, MA24]. Responsibility [GKXK13]. ReSQoV [KCY22]. Ressource [Mar08]. restart [BBHL08, EBLM22]. Restoration [AAC^{+}17, BS96, XWX^{+}17]. Restoring [EGJS15]. Results [HW93, Man15b]. Resurrecting [AKCP21]. Retargetable [GFH82, Fra83, GHF83a, GHF83b, WNL^{+}83]. Rethink [WRX11, XJWW15]. Rethinking [Ott18, PBWH^{+}12, RGSJ17, WCSG05]. Retraction [KSS^{+}23, MA24]. retrofitting [CGL^{+}08a, CGL^{+}08b, CGL^{+}08c]. Retrospect [GLC84]. Return [SYB12, Ven97c]. Return-Oriented [SYB12]. returned [BBS06]. Returning [PSBG11a, PSBG11b]. reuse [LU04]. Reverse [SDS^{+}21]. Review [AP22, AEA^{+}23, Ano97a, BDF19, BDG18, DCM22, Fro13, MBK24, Ng01a, Ng01b, QCM^{+}24, SLZ^{+}25, ARA18, ARA20a, AGH^{+}15a, BJG19, BJ20, KHA22, MB21, MA17, Van98, ZJRW19, Mat10, ARA20b]. Reviewer [Ano03b]. Reviewers [Ano06b]. Reviews [Ano03b]. Revised [Ram93]. Revisited [SCD90]. Revisiting [AJH12, CL16b, HMS17, IBBA20, IYAK23, Ran20, WWWL13]. revolution [McK11]. Reward [BL17, NMC18b, NMC18a]. Rewriting [WMUW19, XWX^{+}17]. RF [FF22]. RHEL [P^{+}08]. rich [RSLAGCLB16]. Ridge [SVN^{+}10]. Right [NBK16, HUL06]. rigidity [PRS26]. rigor [Vit14]. Rigorous [KJ13, Man15b]. RISC [ABDD^{+}91, BSUH87, DK23, MTJ^{+}22, SMP22, SVM^{+}23]. RISC-V [DK23, MTJ^{+}22, SMP22, SVM^{+}23]. Risk [HZL^{+}18, ZZW22, PMP23]. Risk-aware [HZL^{+}18]. risks [Bel06]. RNN [Par25]. roadside [YBZ^{+}15]. Rob [Bas04, Bas06]. Robinhood [PWJ16]. Robot [Arm78]. Robust [CCML12, LSPP^{+}23, PFNC20, SGV12, YZSC17]. robustness [SS19]. Rochester [Mar81]. Rockefeller [IEE90b]. role [GLA^{+}08]. Rollback [CHPY17]. Rollup [WFL^{+}25]. Rome [BW03]. Rose [Ano03b]. Rosenblum [War11]. Roundtable [Cre10b, Sta97, Cre08a, Cre08b, Cre09, Cre10a]. route [YPLZ17]. routed [AM16]. Router [GWZ16]. routers [GP13]. Routing [EMAL17, ELC^{+}19, FXHY21, FD08, GR20, HLP^{+}16, MSLY24, NGRF19, YWY^{+}17, FLL^{+}13, FS19, FSH^{+}13, LWL16, SJRS^{+}13, XWW^{+}21, YLTF20]. RPC [CSS^{+}13, KLY20]. RPC-based [KLY20]. RPython [MRG17]. RTLSim [YYPA01]. RTOS [JK17]. rule [HTAY21, Pul91]. Rules [Kov19, CFRSSR19]. rules-based [CFRSSR19]. Run [Bad87, ACT94, AWR05, CGM17, Com00]. Run-Time [Bad87, ACT94, CGM17]. Running [Bad87, MDD^{+}08, NL19, GMR93, KGS16, SLC20, SZ88]. runs [FIF^{+}15]. Runtime [CDM^{+}25, GSS^{+}18, Kam83, KP15, MB98, NMMP15, ST24, Shi03, XLWX19, HL24, KNHH18, ORPS09, RVJ^{+}01, STY^{+}14]. Runtimes [HD16, Han05, CSV15, GK05, PBAM17, WWH^{+}17]. Rust [Kol19].

S [M^{+}06, Ber86]. S-GRACE [M^{+}06]. S.u.S.E [KGG00]. S/370 [Ber86]. S2H [YJZ^{+}21]. SableSpMT [PV06]. Safe [BHI15, RSF^{+}15, SKI^{+}17, VVC^{+}17, CFS^{+}12, CLDA07, MSZ09, TV18]. Safety [BSI^{+}15, LSPP^{+}23, MTFK19, CDO24, HM01, MSG01, WWL^{+}25]. safety-critical [CDO24]. Sagamore [ACM03b]. Sampling [Lee16, THB22]. sampling-based [THB22]. San [ACM99, ACM06a, Ano04b, Ano10, IEE93a, USE99, USE01b, USE02]. sandboxed [MK23]. Sandboxing [GG11]. Sandpiper [WSVY09]. SANs [ZSXZ07]. Santa [ACM00]. Sapphire [URJ18]. Satellite [QLL^{+}21, CFVP12, SSN94]. Satellite-Terrestrial [QLL^{+}21]. Satisfaction [LVM16]. Satisfaction-Oriented [LVM16]. Satisfied [KCY22]. SAVE [GKJ^{+}19]. saving [YLCH17, YW20, YLJ22]. savings [WWL^{+}25]. SC’11 [LCK11]. SC2003 [ACM03a]. SCADA [ADWM18]. Scala [AT16, SMSB11, Sub08]. Scalability [KMK16, QNC07, TCP^{+}17, VP16, ZXL^{+}24, AH24, BFS^{+}18, JB24]. Scalable [ASPP22, CL17b, DSM^{+}18, FBL18, HYK^{+}23, HJ10, HPS22, JAD19, KCY22, Kol19, KLK^{+}22, Li14, RSN^{+}18, SD01, SADP21, SWL^{+}23, UVL^{+}13, XML^{+}18, ZL18a, ZSP^{+}21, DS18, HLW^{+}10, HTAY21, HPS23, LKR^{+}19, SJJ^{+}12, SPF^{+}07, SG10b, Uhl07]. Scalar [NTL^{+}24]. Scale [CZX^{+}19, HC17, PHL^{+}12, RIP18, RJS^{+}18, SLM89, XDLS15, ZLW^{+}14, ZTA^{+}21, FPGK18, LPD^{+}11, MSG^{+}12, SZ13, WWT89, WCG21, YZSC17]. scaled [KNHH18]. ScaleTrust [HYK^{+}23]. Scaling [CBJ22, CFC^{+}24, HC17, JWL^{+}18, JDJ^{+}06, LW20, PBL^{+}16, TCP^{+}17, AB16, SBNU18, SSEA18, TSCB19, TCTH23, XLQL18, AMAB17]. Scaling-Aware [HC17, AMAB17]. SCAN [Ble89]. Scenarios [MTFK19, SADP21, KCV11, Sch13a]. Scenes [Cra98]. Schedulability [NL19]. Scheduler [AGC18, IYAK23, ASB18, KCS14, RAP19, SWH^{+}13]. schedules [LC14]. Scheduling [AT24, ARAAA19, AD18b, BE17, Car23, EB20, EGR15, FML^{+}22, HSN17b, JJK^{+}11, KDB16, LMM18, LGJ^{+}18, LD05, LWW16, LC13, PG17, PG18, RB17, SGK^{+}23, TTH^{+}19, VS19, WDL^{+}20, WYZY24, WWT89, WCG21, ZWFX17, ZQCZ16, ZLW18, ABB19a, AT23, ATZP21, BC10, CRX24, CCL^{+}20, CLL^{+}23, CCW^{+}20, DEE^{+}16, DQLW15, DXM^{+}17, DCMW17, DS22, GSG22, HKS19, HZK^{+}25, JGW^{+}11, KS18b, KKJ^{+}13, KNHH18, KCV11, LFHS23, MMTM22, NAR19, PC21, RWC21, RZ14, RHZ^{+}17, SS13, SHLJ13, SSN12, Sto07, TMLL14, THG^{+}18, VVB13, WQG15, WCC^{+}16a, XCJ^{+}14, XLWZ18, XZK^{+}20, XXWG23, YPLZ17, YXL^{+}20, YWGH13, YQZ14, YQZ19, Yu20, ZSR^{+}05, ZXB^{+}24, ZB18, MA21, MA24]. schema [SI81]. Scheme [AJ18, AMA18, KAZS14, RSN^{+}18, SHZ^{+}14, YWR^{+}14, KK21, KJLY15, LJYZ15, SM23a, XCJ^{+}14, YPLZ17, YQZ14, YQZ19, FM90, FDD^{+}19, KR94]. Schemes [Do11, LSSC22, MNA16, YIR24, YWGH13]. Schloss [IEE01]. School [BGP00]. Science [ACM06d, BR01, DG05, SGV12]. Sciences [Shr89, MS91b]. Scientific [AD18b, Bad87, RB17, CSMB15, dCCDFdO15, EB20, MPM^{+}20, WCG21]. Scientists [THLK10]. scraping [DAdBM^{+}24]. Screening [LP14]. Scripting [MJW^{+}06]. SD [KKK^{+}18]. SDDSfL [CLŁS12]. SDN [Pap20, ASL^{+}20, BDF19, HTB19, HTB22, LLY^{+}18, SDM21, SB18, SP22, VVC^{+}17, WYZAD20]. SDN&NFV [ABB^{+}19b]. SDN-based [WYZAD20]. SDN-Enabled [HTB22, HTB19]. SDN-NFV [SP22]. SDNs [ALW15, BG20]. SDWN [AFG^{+}17]. SDWSN [AHKC23]. SDWSN-Based [AHKC23]. SE [LYBB14]. Seamless [Hir92, TDG^{+}06, XWJX15, BADM06, DS20]. Search [Cox12, MNS^{+}14, VG20, CWdO^{+}06, KMT14, LY23, SBI21, Tho68, WXZ^{+}17]. search-based [WXZ^{+}17]. Seattle [ACM05c, ACM06b, LCK11, Ost94]. Sebastopol [Ano97a]. sEc [SMK02]. SECD [Abr82, AS85a, AS85b]. SECD-M [Abr82, AS85a, AS85b]. Second [ACM06f, IEE93a, Shr89]. SecondSite [RCOW12]. Secure [AD19, AVNR19, AMH^{+}16, CCML12, CZH^{+}25, CLDA07, ETAB22, HYK^{+}23, JSHM15, JAS^{+}15, LJR12, LP11, PEC^{+}14, QZDJ16, RC18, RI00, RSGG15, SOAK23, THB06, TtLcC13, WF07, YML^{+}18, vD00, BDS^{+}09, GNDB16, GNK24, HKD^{+}13, ISE08, LLX^{+}17, Str05, SL12, TLBW12, ZXB^{+}24, ZBP05]. Secured [TMV12, WCC16c]. Securing [Sar01, Hal08, Hal09, PDM20]. Security [AKK^{+}07, Ano93, AEB19, Att79, Att73, BDG18, De 06, ESY^{+}17, FJKK17, GW07, GMGF24, HHSG18, HB17, IEE84a, IEE90a, IEE91, IEE05, JE12, KZB^{+}90, KS08a, KS08b, LWLL10, NMMP15, PM19b, PvDS08, Pfo13, Rob12, SJV^{+}05, SM90, SABL20, SEF^{+}06, Ste05, TMV12, TV12, USE00b, VN08, WHD^{+}09, WTM18, XKKL23, ZL16, ZL18b, ZYH^{+}19, Ano07, BTMS10, Bau05, Bau06b, Bau06a, Bel06, BCP^{+}08, Bor07, BBS06, CCMY07, CBFH20, EM06, FA21, Hal09, HMS04, IIK^{+}06, LLW^{+}12, MD73, MD74, Mat09, MKM^{+}08, MA17, PG11, PZH13, PBB13, PDML25, Sch13b, SDN09, VT14, WHSE15, YSM^{+}21, vCPWvT11, DTW07]. security-aware [FA21]. Security-focused [BDG18]. security-oriented [IIK^{+}06]. see [Yur02]. SEED [DTW07]. Segment [ELC^{+}19, LZZ^{+}25]. Segmentation [LYY^{+}24]. seinen [KGG00]. seL4 [DK23]. Selecting [GSKJ18, NBK16]. selection [AHRR22b, HM20, JK13, LZWC13, LLWW18, MCJ19, NNK21, ZB18]. Selective [WZW^{+}11]. Self [BHI15, BRX13, HHW10, JC18, dOL12, SEPV19, XCSM18, BKT^{+}19, CBLFD12, GK05, GKJ^{+}19, KKB14, NNK21, OK90]. Self-Adaption [BHI15]. Self-Adaptive [XCSM18, JC18, SEPV19, BKT^{+}19, GKJ^{+}19, KKB14]. Self-Configuration [BRX13]. Self-Healing [BHI15, GK05]. self-hosted [CBLFD12]. Self-management [dOL12]. self-optimizing [NNK21]. Semantic [Das91, DGLZ^{+}11, FL13a, GKP^{+}19, ST24, SBBP20, AD18a]. Semantic-centric [SBBP20]. Semantics [Gol71b, WIDP12, Dan12, EdPG^{+}10, Siv04, Wal76, ZHCB15]. Semi [SEK^{+}19, MSZ09]. semi-automatic [MSZ09]. Semi-Autonomic [SEK^{+}19]. Sensing [SML18]. sensitive [DK17, KSLA08, LCL14, MMTM22, ZBP07]. sensitivity [HB13, TZK17]. Sensor [BSI^{+}15, LC02, MAK07]. sensors [ALL06]. Separation [KF91, LSPP^{+}23, WLMD16, LWM14]. September [ACM81, ACM04a, ACM05a, ACM06c, ACM06b, Ano93, BW03, GHH^{+}93, Jou85, JPTE94]. Sequence [ARAAA19, EDS^{+}15]. sequential [Clo85]. Serialization [BP01, BP03]. Series [Kee77, KAH83]. Server [ARA18, Ano03a, Apr09, BE17, BJC24, Bod10, Car06, CGS06, Do11, HSK17, Joo09, KSS09, KS10, KLLT18, LZ15, Lar09, LC09b, LC09a, LXZ^{+}21, Mar08, MAK18, MG08, MG09, PZW^{+}07, RWX^{+}12, R^{+}02, SWC08, WN17, ZHW^{+}17, Zim05, Zim06, ARA20b, ARA20a, A^{+}04, AGH^{+}15b, AT23, BKR20, B^{+}07, DBC^{+}00, EBJ17, Hal08, IMK^{+}13, KF18, LC14, LLWW18, LLS^{+}08, LL14, LDDT12, MNT14, MRM06, NTH^{+}17, NMC18b, NMC18a, R^{+}13, RPE12, Wal02, WDT18, YZW^{+}13, AAH^{+}03, Ano03a, B^{+}07, D^{+}04, Ham07, Lar09, MWHH05, OH05, R^{+}06, Rul07, R^{+}02]. Serverless [FR24, KSS24, MK23, NRdA^{+}20]. Servern [Mar08]. Servers [DSM14, JJK^{+}11, KAZS14, SDD^{+}16, SKJ^{+}17, WLW^{+}17, A^{+}04, BJ22, BBHL08, G^{+}05, Hal08, JDJ^{+}06, JB24, Mly09, SZ13, YLJ22]. Service [AP22, AAMBE21, BB13, BCW20, BFG^{+}14, Bos24, DKW15, DPCA11, EMAL17, ESY^{+}17, FBM^{+}21, FR24, FFM^{+}23, GR20, GGK18, HS21, HYL^{+}25, HW12, HJG18, HPHV17, JWL^{+}18, KBK^{+}21, LP14, LGXC23, LZY^{+}24, LGZ^{+}19, LLW^{+}16, LYL21, LW20, MP16, MSC^{+}21, PHXL19, QLL^{+}21, RSNK17, RSGG15, SAM14, WVT^{+}17, WCC20, WHD^{+}16, XZL^{+}20, ZLG^{+}20, ZLZ^{+}21b, ZWC^{+}23, BSM^{+}12, BSOK^{+}20, CMG^{+}19, CHCC07, CFRSSR19, DS18, DXM^{+}17, EdPG^{+}10, ECAE13, EMI13, Fro13, GHM^{+}18, HKJ19, KKB14, KS20b, LZWC13, MCJ19, MDZ^{+}21, MHS21, NZH20, PAKY16, PFNC20, RCOW12, SHB19, SZ13, TDD20, TSCB19, VOS12, WB16, XXWG23, YCL^{+}19, YXL^{+}20, YKP^{+}24]. service-aware [TDD20]. Service-Based [LP14]. Service-centric [AAMBE21]. service-chaining [GHM^{+}18]. Service-Oriented [HW12, MP16, RSGG15, Fro13]. Serviceability [RB01]. Services [BFHW75, IEE06b, KA24, KCY22, KLR^{+}20, MSS^{+}15, MLXG19, WZH^{+}16, WC01, Wid01, ZLW18, BDS^{+}09, HBP06, KBB11, KSLA08, LKR^{+}19, LTZ^{+}14, ZEdlP13]. Set [AC98, EL98, NKY^{+}18, ZDLG17]. sets [HW15]. setter [YJZ^{+}21]. setups [RPE12]. several [FGG14]. SFC [FXHY21, SJMG24]. SFCs [LGS^{+}23]. SFT [LGS^{+}23]. SFT-Box [LGS^{+}23]. SGAM [ZLH^{+}15]. SGX [VMW^{+}19, KBC21, NBB^{+}19, YKP^{+}24]. SGX-enabled [YKP^{+}24]. Shadow [WLW^{+}15, GHS16]. ShadowReboot [YK13]. Shadows [TDMP23]. share [DSS19, KNHH18]. Shared [Bro89, CH08, Cro93, KR18, Low88, RLZ^{+}16, RKRK17, SLM89, SV13, SNC91, SNS03, ZLSI17, CFS^{+}12, JGSE13, PW03, TZK17, WWS89, WDCL08, ZWKX17]. Shared-Memory [Cro93, RLZ^{+}16, SLM89, WWS89]. shared-source [PW03]. Sharing [AGC18, ACA16, BFHW75, Cre65, CDN02, LYGG20, Mad69, MS70, PTM^{+}15, RG17, SAR24, SAB^{+}07, XML^{+}18, ZZG^{+}23, ZHH^{+}24, FMX^{+}25, GGK19, LCLG26, LLZ^{+}19, LLS14, LTZ^{+}14, OKAM17, TtLcC13, WTLS^{+}09]. Sharing-Aware [RG17]. shell [FL13b]. Shift [ARAAA19]. shock [BG20]. Shoot4U [OLZ16]. Short [HW15, KKC^{+}16, SOKE23]. Short-circuit [KKC^{+}16]. short-term [SOKE23]. shortest [AM16]. shot [JK15]. Should [NBB^{+}19]. Shoulders [FS12]. Showcase [USE00a]. showdown [SCEG08]. Shredder [AMH^{+}16]. Shredding [AMH^{+}16]. Shrink [LWB13]. Shrink-Fit [LWB13]. Shrinking [Ste14]. shuffling [ZWC^{+}14]. Shuttle [cCWS14]. Sibling [OG16]. side [LF19]. side-channel [LF19]. SIGACT [ACM99]. SIGCOMM [RM03]. SIGCSE [ACM06d]. SIGMETRICS [ACM81]. Signal [MBK^{+}92]. SIGOPS [ACM04a]. SIGPLAN [ACM01a, ACM99].

SIGPLAN-SIGACT [ACM99]. SIGSOFT [ACM01a]. SigVM [ZBS^{+}22]. Silent [AMH^{+}16]. Silicon [ZL18a]. Silicon-Monona [ZL18a]. SILLIAC [Gre10]. Sim [Skr01]. SIMD [PSBG11a, PSBG11b, PBR^{+}90, Sig89]. Simics [Ano14a, MCE^{+}02]. similarities [CL14, CL17b]. similarity [GVI13, LLF^{+}18, LLWW18]. Simple [Bak83, Cox07, NOR15, WDT18]. Simplicity [BGP00, DSSP06]. simplification [FS08]. Simplified [Beg12, ZZW^{+}21, PSC^{+}07]. simplifying [Cla05]. simulated [GE85, RH17, WDSW01]. Simulating [Ben21, HO92, NLD^{+}23, Pou90, RPE12, TO91, ZR06, FPGK18, Skr01, WC91]. Simulation [ADG^{+}92, AB16, DBMI92, JN15, KD78, Kut92, MCE^{+}02, MBK^{+}92, MJ93, PBR^{+}90, PY93, SXMX^{+}18, Tur92, WB81, WWMG06, YP15, Ano94, BHvR05, Bur02, BS96, Clo85, DSSP06, IMBB20, IM93, KK79, LJN^{+}00, NRS92, RMB02, SK13b, SHB19, UBL^{+}82, WWS89, YYC^{+}19, ZSRR22]. Simulations [LCT^{+}15, BL90, CLM24, DH01]. Simulator [Ben21, CK96, CRZ83, Dun86, FTNY69, PCR89, Ber86, BR01, CMP^{+}07, DC15, GBO87, Hog02, KW80, MRL02, YYPA01, Ano14a]. Simulators [NMHS15, Sup04, GFHML^{+}25, Man18, Yur02]. Simultaneous [LRZ16, ABB^{+}15, FS19]. Singapore [Ano06a, TLC06]. Single [CCO^{+}05, KP15, AGIS94, Fis91, KNHH18, LSS04, Mon97]. single-chip [Mon97]. Single-Computer [CCO^{+}05]. single-ISA [KNHH18]. single-node [LSS04]. single/multigrid [AGIS94]. site [CPST15, SSB^{+}14a]. situ [CKRJ17]. Sixth [ACM05a, TLC06]. Size [Lam75, NKY^{+}18, HPHS04, UTO13]. Sized [JJ02]. sizes [GPS^{+}18, HM18]. Sizing [LWB13, VTW16, CSV15, WSAJ13]. Skip [WBHN18]. Skype [Joo06]. SLA [AB16, EdPG^{+}10, GTGB14, KB21, KKB14, RT18, SS22, ZHL16]. SLA-based [AB16, GTGB14, KKB14]. SLA-driven [EdPG^{+}10]. SLA-guaranteed [KB21]. sledgehammer [LU04]. Slice [EMI13, KPHA20, ZLZ21a]. Slicing [AAT^{+}22, MMdMCPSR25]. Slicing-assigned [AAT^{+}22]. Slim [Abr80]. Slimming [WGF11]. SLO [GCL^{+}21, HC18, LJFS17]. SLO-Aware [GCL^{+}21]. Sloop [DZ02]. Small [JJ02, SSB03, AH24, DK75, HPHS04, SS72, WH08, WWT89]. small-scale [WWT89]. Small-Sized [JJ02]. smalltalk [FIF^{+}15, BMWB86, BSUH87, G^{+}88, Lee86, SUH86, TLD^{+}89]. Smalltalk-80 [BMWB86, BSUH87]. Smart [AHKC23, AEA^{+}23, Ano03b, CCW^{+}20, FZD^{+}24, GPM21, NAR19, NHL22, RHV17, SCF^{+}25, GEK25, GLV99, MPA^{+}18, Rou07, WTLS^{+}09, ZBS^{+}22]. SmartFVM [KLK^{+}22]. SmartOracle [SCF^{+}25]. Smartphone [DAH^{+}12]. SMIL [Bru07]. SMILemu [Bru07]. SMOK [DZ02]. Smoot [Fro13]. Smooth [DL89]. smoothed [CL14]. SMP [CL16a, KKJ^{+}13, RZ14]. SnapFiner [CHLY18]. SNAPS [RG19]. Snapshot [CHLY18, RG19]. Snapshots [CWL^{+}15, DS16]. Snowbird [ACM01a]. SnowFlock [LCWB^{+}11]. SOAR [SUH86]. SOC [LVM16]. social [BTLNBF^{+}15b, LWLL16]. Society [IEE90a, IEE91]. Soft [Ano03a, LXL^{+}22, XH16]. Software [AFG^{+}17, Ano94, Ano03a, Ano03b, AE01, AMA^{+}14, AJM^{+}25, BCG73a, BCG73b, CL17a, CPKL17, CLKEF21, CGMD19, CMK^{+}16, DBMI92, DL89, EDS^{+}15, FXHY21, FML^{+}22, HO22, Hsu01, IGBKR19, JMSLM92, JN15, KP99, Kna93, KAJW93, LH16, LTT92, LLW^{+}16, LXZ^{+}21, LZM^{+}20, MZD^{+}18, MBK24, MP16, NVV^{+}24, Ost94, Ott18, PJZ18, Pap20, Par79, PBR^{+}90, Sof83, SM06, SLW^{+}24, SN23, SMA18, Shr89, SAT09, SLZ^{+}25, SB18, SKT^{+}19, Sta07, SCL^{+}19, Tho93, TBS17, Win71, YWH^{+}21, YYL^{+}15, ZKWH17, vdK09, ACM01a, AA06, ALW15, AAB^{+}05b, AC95, BD11, CBGM12, CRX24, CFG^{+}13, DS19, FP14, Guz01, HHSG18, HH13, HP77, LJR12, LWL16, MNT14, PMP23, PV06, Sam22, SV17, TK20, WZW^{+}11, XJW^{+}18, YJZY12, ZWKX17, ZLZ13, ZHCB15, CK06q, CK06t, CK06r, CK06s]. Software-Based [LZM^{+}20]. Software-Defined [AFG^{+}17, CL17a, CMK^{+}16, FML^{+}22, JN15, LLW^{+}16, LXZ^{+}21, MP16, SLZ^{+}25, SB18, TBS17, YWH^{+}21, ZKWH17, ALW15, HHSG18, LJR12, TK20, XJW^{+}18]. Software-Driven [NVV^{+}24]. Softwarization [CM18, Mon22]. Softwarized [EVCL21]. Solaris [VSC^{+}10, WF03, Gal11, HDM08, See10]. Solid [SYC14]. Solid-State [SYC14]. Solution [CHW12, CXLX15, Coh10, DMG^{+}15, Gua14, KDB16, PYDG22, XYD^{+}18, BKT^{+}19, DSS19, MPA^{+}18]. Solutions [HN10, PM19b, SL16, ATS16, AGIS94, EMI13, PZH13]. Solve [Cap21, MTFK19]. solver [TB14]. solver-aided [TB14]. solvers [GCARPC^{+}01]. Solving [AAR22, XA22]. Some [Ker88, Par71, Man15b]. Sorrento [M^{+}06]. Sorting [BGM70]. SOSP [ACM03b, Vra05]. soul [McM11]. sound [BHSB14]. soundness [Req03]. SOUP [ZFH^{+}22]. Source [Ano03a, GMGF24, LLWM23, SJV^{+}05, SNS03, AAB^{+}05a, But94, CKRJ17, Cia07, JM08, LC09a, PW03, SIK^{+}16]. source-level [But94]. sous [Apr09]. Sova [YWH^{+}21]. SP [IBM94]. SP2 [Boz89]. Space [PGP19, SVM^{+}23, XML^{+}18, Kha19, PEL11, PG11, Web10, WXW15]. space-efficient [PEL11]. spaces [GH91a]. SPAN [RD90]. spares [CRX24]. Sparks [VN08]. sparse [Kra90, TP25]. sparse-matrix [Kra90]. Spatially [HW93]. Spatio [ZZG^{+}23]. Spatio-Temporal [ZZG^{+}23]. SPC [JYW^{+}13]. SPC-indexed [JYW^{+}13]. SPDK [FMX^{+}25]. speaking [Sam22]. Special [Bag76, Cre65, KM13b, TZB19, WYZAD20, Yur02]. Specialized [BDK^{+}08, ZZW^{+}21, PGLG12, Yur02]. Specific [HHV^{+}02, WIDP12, ZZW22, JKDC05, ZS01]. Specification [Coh97, DMS02, LY97b, LY99, LYBB13a, LYBB13b, LYBB14, LS15, II79, Qia99, Sun95b, SUN97, JCV99, Taf11]. SPECjvm98 [LJN^{+}00]. Speculation [AC16]. Speculative [ZLL^{+}20, GI12, PV06]. Speed [KKS^{+}19, GGJ^{+}92, LRP^{+}19, RPE12, UTO13]. SPEED08 [VW08]. speedup [JB24]. spherical [Hol95]. Spielesammlung [CK06q, CK06t, CK06r, CK06s]. Spin [CWS12, WCS06]. Spinlocks [KMK16, OL13]. Spinning [IYAK23]. SPIRE [JYW^{+}13]. Split [HWHW18, SIJPP11]. spoofing [SDN09]. Sporadic [BE17]. Spot [TVKB16, VS19]. Spotless [MS00, SMES01]. Spotlighting [Ano06a]. Spots [WBB^{+}16]. Sprache [Dalxx, Dal97]. Spreading [CLW^{+}14]. square [DG05]. squeak [Guz01]. SqueakJS [FIF^{+}15]. SR [AT23, DYL^{+}12, DCP^{+}12, HB12, XD16, XD17, YWCF15]. SR-IOV [DYL^{+}12, DCP^{+}12, HB12, XD16, XD17, YWCF15]. SR-PSO [AT23]. SRAM [NTL^{+}24]. SRVM [XD16]. SSD [FMX^{+}25, GFHML^{+}25]. SSDs [HC18]. St [IEE06a]. St. [ACM97]. Stable [XRL^{+}22]. Stack [AE01, Cia07, HB12, NSC^{+}22, Ran02, SSOT17, WH99, WBHN18, KRCH14, LH13, WW77, SCEG08]. Stack-Based [Ran02, KRCH14]. Stackdb [JHE14]. stacked [TP25]. stage [CLG^{+}10]. Standard [DPCL22, MR04, RSF03, WKG17, Ano94, Rus08]. Standards [Mar81, SG10a]. standards-based [SG10a]. Stanford [IEE96a, IEE97, IEE99]. start [KSS24, KMT14]. Startup [HS06]. Starvation [KLY20, AH24]. State [BDG18, CFC^{+}24, HZK^{+}25, LHW^{+}20, LJL^{+}11, SGB^{+}16, SYC14, Sur01, TV12, AEB19, MPA^{+}18, RI25, Sch13b, Sig89, Ven99b, Web10]. State-Based [TV12]. State-driven [HZK^{+}25]. State-of-the-art [BDG18]. stateful [XHW^{+}19]. Stateless [VDO14]. States [SBK15, IMK^{+}13, MC98, STFH15]. Static [JM08, YC16]. statistical [KF18]. Stealing [PWJ16]. STEP [BDE^{+}03]. Stephen [Fro13]. Sticky [KC12]. STM [Sub11]. Stochastic [CCL^{+}20, FX06, FK13, GR15, SDD^{+}16, HKS19, NMC18b, NMC18a, YLTF20]. stock [VGL23]. Stop [LWB^{+}15]. StopWatch [LGR14]. Storage [ACM04b, Att79, Bad82, BDT13, CC77, Cla05, ETAB22, FFBG08, FKZ17, GSW^{+}17, KCWH14, KHW^{+}16, KLK^{+}22, LCK11, LJFS17, MJW^{+}14, PPTH72, PP16, PYYG21, RB24, Rou07, SSOT17, VW08, ZSW^{+}06, ZLL^{+}20, ZTA^{+}21, BN89, CCL^{+}17, FLCB10, HJ10, HPcC04, JGSE13, LKY^{+}17, PFH^{+}16, Pat12, TLBW12, XJWW15, YLK^{+}10, ZLZ^{+}19b, ZLLL13]. Storages [TF16]. Store [Low88]. Storing [CWL^{+}15]. Storms [SB16]. Story [Arm98]. strange [Fab13]. Strategies [YLN^{+}17, BDT13, FGG14, GHK24, GHM^{+}18, LLS14, PFH^{+}16, TKG89, YI24]. Strategy [BMJ^{+}22, LLZ18, XCSM18, ZB20, DKF94, HKS19, KS18a, LPZ^{+}22, MW18, THB22, WSX^{+}19, Won97, YWH^{+}23, ZLZ15, ZLH^{+}15, ZLCZ18]. Strategy-Proof [LLZ18, ZLH^{+}15]. Stream [MMdE19, MV16, LMDP19, ZDK^{+}19]. Streaming [MSC^{+}21, BMER14, RSLAGCLB16, SIK^{+}16]. Streams [MM93]. Strengthen [GPM21]. stress [MC98]. String [HOKO14, YDW18]. Striping [DK93]. Stripped [JJ91]. Stripped-Down [JJ91]. Strong [ZZW^{+}21, ZHCB15]. structural [ORPS09]. Structure [Com65, LHW^{+}20, MDFS72, SS72, ZFY18, ZLZ^{+}19b]. Structured [Das91, Gai75, CFS^{+}12, IM75, Syr07]. Structures [AGLM91]. student [CKP78]. studies [vCPWvT11]. Studio [Ano03b]. Study [BBM^{+}15, LMR18, LSSC22, LJL^{+}15, PXG^{+}17, PK75a, ZAI^{+}16, CMG^{+}19, CDO24, HIIG16, HL13, KW13, MFT^{+}19, EYGS19, Pul91, RHR02, RK18, SM79, SASG13, Sig89]. Sub [GGM^{+}16]. Sub-System [GGM^{+}16]. Subroutines [HT98, Qia99]. Subset [SUN97, Req03]. substrate [OKAM17]. Subsystem [HH79, Ste14]. Suffix [HWHW18]. Sugar [YML^{+}18]. Suitable [Vog03, GSKJ18]. Suite [DHPW01, WZT19, DTW07, GPW03, SMSB11]. summaries [BA19]. Summary [CFH^{+}79]. Summer [HMS17, Sof83, USE85, USE86]. Sun [Gal09a, Gal09b, Gal11]. Superblock [KS13]. Supercloud [SJS^{+}17]. Supercomputer [MBK^{+}92, LPD^{+}11, XH90]. Supercomputing [ACM89, ACM96, ACM00, ACM04b, ACM05c, Hir92, IEE90b, IEE92, IEE93b]. Superconcurrent [NRS92]. superoptimization [HW15]. superscalar [VdlFCC97]. supertype [RRB17]. Supplement [McC74]. Support [BP01, DJ77, HHV^{+}02, HD16, HB12, KYP^{+}17, LV99, MSI18, NSL^{+}06, NLPV12, RI00, SCC^{+}23, SSG90, Tur92, XD16, ZL18a, dGG^{+}17, AC95, BADM06, BTLNBF^{+}15b, BP03, CHCC07, CFS^{+}12, DJ76, EBLM22, GK05, NBS18, Olb78, ORPS09, PGLG12, RK18, SJRS^{+}13, STFH15, SL12, TY14, THL03, WK08, WCS06, WLL^{+}13]. Supporting [BMS16, CWS12, Kim84, Kov19, MSS^{+}15, Mon97, RT93, XWJX15, YWCF15, ZZF06, BTLNBF^{+}15a, GD08, TT93]. Supports [Ano03a]. surgery [PBL^{+}16]. Survey [ASL^{+}20, AAAF21, BAL15, FLZ^{+}20, Gol74, HSN17b, KKLV16, KL14, KK19, LXW^{+}23, Mal72, Man15a, PM19b, PS16, PS19a, QTR21, SB16, SGB^{+}16, UOKT84, VV18, WMUW19, AGH^{+}15b, CB10, CCDR22, DS19, FMIF18, GEK25, GFHML^{+}25, HKB19, MG13, MHS21, NIA18, PBB13, XTB17, YWL^{+}18]. Surveyor [Fra83, GHF83a, GHF83b, WNL^{+}83]. Survivability [NHL22, YZW^{+}13]. Survivable [ACA16, AM16]. SUSE [Bau06b]. Sustainability [FBL18, SS17]. Sustainable [GB19]. SVGrid [ZBP05]. SVLM [DS20]. SVM [JAS^{+}15]. SVS [LJZ12]. SW [DCG12, Wu13]. swap [KB21]. Swapper [ZLSI17, ATS14]. Swapping [CC77, ABG14, LCLG26]. Swarm [PS23, AT23, BRS^{+}22, JNR12, KSS^{+}20, KSS^{+}23]. Sweet [WBB^{+}16]. Swift [NOT^{+}17]. Swiper [CRZH15]. switch [BR01, Ste14]. switches [YGLY21]. Switching [DMG^{+}15, LBL16, YLJ22]. Sy [USE01c]. Sydney [MR91, Gre10]. symbiotic [LD11]. symbolic [MMP^{+}12, TB14]. SymCall [LD11]. Symmetric [DBO^{+}18, GMP89]. symmetry [PBL^{+}16]. Symposium [ACM75, ACM03b, ACM05a, ACM06d, Ano00, Ano01a, Ano01b, Ano04a, Ano04b, Ano10, HHK94, IEE84a, IEE85, IEE90a, IEE91, IEE96b, IEE06a, Ost94, TLC06, USE91, USE93, USE00b, USE01d, USE02, Vra05, IEE96a, Ano02]. Synchronization [BC19, LJL^{+}11, ZHH^{+}24, ZJXL11, Sub11, Uhl07, Ven97d, YQZ19]. synchronized [KS18b]. Synchronous [SIR^{+}17]. synergy [BRS18]. syntax [KMMV14]. Synthesis [DMS02, BPB86]. Syracuse [IEE96b]. Sys [Mur69]. System [ACM75, Abr80, ABCC66, Ano10, AAK18, Bad82, BFHW75, BBD^{+}91, BPP^{+}17, BH73, BYBYT16, Ben21, BJPS73, BGS89, B^{+}05, Car13, CSS^{+}13, CZX^{+}19, Cre65, CWL^{+}15, CHPY17, CHLY18, DMR10, DM75, Fis01, GGM^{+}16, G^{+}06, GH91b, HXZ^{+}16, HW93, HHC^{+}16, HWCH16, IBM76a, IN87, JAD19, Kam83, Kee77, KP15, Kut92, LP14, Li14, LCZ^{+}19, LXQ24, LCFL12, LXM^{+}16, MCE^{+}02, Mar73, MMdMCPSR25, Mat10, MNN05, MS70, MDGS98, MB98, MS91b, MM94, NSHW10, Neu92, NMS^{+}14, P^{+}08, PHXL19, PDY^{+}23, QTR21, R^{+}06, RHV17, ST24, Sch86, SLM89, SVN^{+}10, Shi03, Shr89, SJA^{+}17, SWF16, Ste05, WLW^{+}15, WK90, ZCJ^{+}21, ZSXZ07, ZQCZ16, ZLL^{+}20, ZZF06, ZXY^{+}15, ZZW22, AD18a, AEMWC^{+}12, AL05, AH12, ACT94, AP18, Bar78, BSD19, Bor07, Bur02, Caa00, CWH^{+}14, CK06b, CK06e, CKP78, CBFH20, DHD20, DCA17]. system [FFBG08, Fis91, Fli77, GGQ^{+}13, HN08, HKD^{+}13, HC12, Hui18, IBM88, Int88, KB21, KCKC15, KK79, LJN^{+}00, Lia05, LLX^{+}17, LMDP19, LDL^{+}08, MR23, MD73, MD74, MDFS72, NMC18b, NMC18a, PRB07, PK75b, RG19, Rob06, RI25, SNV10, SPF^{+}07, SJL20, SWW^{+}18, SZ13, SS72, STY^{+}14, TC10, Vag10, Van06, VMBM12, VSC^{+}10, WKT08, WH08, WWT89, WHSE15, WF07, WC91, YLCH17, YZSC17, ADG^{+}92, ABDD^{+}91, Car14, Gum83, HTAY21, IBM76a, SNC91]. System-level [SVN^{+}10, AL05, BSD19, WHSE15]. System/370 [IBM76a, Gum83, IBM76a]. System/6000 [ABDD^{+}91]. System/9000 [ADG^{+}92]. Systemarchitecktur [See08a]. Systematic [BDF19, DCM22, MBK24, QCM^{+}24, ARA18, ARA20b, ARA20a, BJG19, BJ20, KHA22, NK22, ZJRW19]. Systeme [WF03]. Systèmes [Han73]. Systems [ACM81, ACM03b, AAT^{+}22, Ano99b, BBMA91, BHI15, BDG18, BG74, CD12, CC77, CAF^{+}91, CDM^{+}25, Das91, DJ77, Fie68, Gol69, Gol71a, Gol73c, Han73, HHS18, Her10, HBL^{+}10, IEE93a, IEE01, JAD19, JYM^{+}23, KSVR23, Lar09, LW11, LJZ12, Mad69, MM93, MJW^{+}14, MKZ^{+}25, MKKE12, MW25, NBB^{+}19, NL19, PC24, PPG^{+}17, QCM^{+}24, RT93, SMP22, SL14, SS75, Say66, SVB93, SL16, SAR24, SN05b, THB06, TCTH23, USE99, USE01b, Vra05, WN17, WLMD16, Win71, YVCB17, YVCB18, ZHH^{+}24, ZD18, ZTA^{+}21, AJH12, ALW15, AT16, Ano93, AAB^{+}05c, BKT^{+}19, BSOK^{+}20, BSSM08, CRX24, CCZ^{+}06, CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, CK06a, CP17b, CCDR22, CDO24, Com00, CGV10, CLDA07, Dav04, Don87, DJ76, DCMW17, EBJ17, FP14, FLCB10, GHH^{+}93, GK05, Ham76, HKN22, HH13, JSK^{+}13, Kee68, KCS14, Kou11, KS20b]. systems [LLLE17, LWM14, LZWD15, LCL14, LTK17, MRC^{+}13, MA17, NS07, NV05, PSC^{+}07, RVJ^{+}01, RKT20, RHZ^{+}17, RJK16, Ros06, RGS^{+}20, SJB14, SK13b, SSMGD10, SJJ^{+}12, Sto07, Syr07, TMJ^{+}21, TT93, THC^{+}14, UIK19, Vac06, Vit14, WR07, WKC^{+}09, XZK^{+}20, YK13, DPCL22]. Systemverwaltung [Lar09].

T [CZX^{+}19]. T-Gaming [CZX^{+}19]. Tables [MT16, MT17, WLW^{+}15]. tackle [Sub08]. tactics [OG16]. Tail [ASSB18, WZKP19, War80]. Tailor [PDL^{+}23]. Tailor-made [PDL^{+}23]. Taipei [SS05]. Taiwan [SS05]. Take [Kis08]. Taking [Uhl06]. talk [Piz17]. Taming [CZL08, HHPV15]. Tan [Fro13]. Tape [DK93]. Target [GC00, FCG^{+}05, TMV25]. Targeting [CDG97]. Targets [Sta07]. Task [ARAAA19, DS22, KMM13, LWW16, PCC^{+}16, RRB19, SGK^{+}23, YLQL25, ATZP21, LFHS23, MMTM22, ZB18].

Task-Based [SGK^{+}23]. Task-Driven [YLQL25]. Tasking [MB98, Shi03, JDJ^{+}06]. Tasks [KGS16, VS19, YSS^{+}17, ABB19a, FGG14, KLY20, YQZ14]. Taxonomy [Car23, FLZ^{+}20, GB19, SGB^{+}16, SB18, ZXR^{+}22, AGH^{+}15a, HKB19]. TCAM [HWHW18]. TCAM-Based [HWHW18]. TCB [HCJ07, HPHS04]. TCP [CL16b, GKXK13, GI12]. TD [WBW^{+}19]. TDX [COV^{+}24]. teach [Don88]. Teaching [Ågr99, Dav04, Don87, GGG03, ME87, Guz01, Ham76, KW80, MS01, NV05, WKC^{+}09, YYPA01]. teasing [LBF12]. Technical [ACM06d, Ano06b, Han16, OH05, USE01a, USE06, BB08, Int06c, Int06a, LC09a, Wal10, ZMD^{+}21]. Techniken [Tho08]. Technique [JHS12, JMSLM92, LTT92, SMK02, SGK^{+}23, WMUW19, ACT94, FAA17a, FAA17b, GNK24, KLY20, KU24, PS23, SLA^{+}16, XHL^{+}13, YKS16]. Techniques [ACM06b, ASL^{+}20, BDG18, BCG73b, BG74, KK19, LJL^{+}15, NKY^{+}18, OVI^{+}12, QTR21, SMA18, SIdLB15, Tho68, UOKT84, VV18, ZZF06, AD18a, ARA18, AA06, AH12, AA24, BADM06, CMGI^{+}23, HSC15, IM93, KS13, KRG^{+}12, SSN12, SHTE11, TSCB19, VGL23]. technische [LC09a]. technologie [Apr09]. Technologies [AEA^{+}23, DF96, FR24, LCMV17, PZW^{+}07, RC18, SABL20, USE99, USE01b, AMIA19, Cla05, Kao17, MPA^{+}18]. Technology [Ano00, Ano01a, Ano01b, Ano02, Ano04a, Ano04b, Cap21, DLM^{+}06, Don06, ELC^{+}19, Got07, Her06, LYY^{+}24, RG05, USE01c, USE01d, USE02, UNR^{+}05, VSMC23, WHD^{+}09, XKKL23, ZAI^{+}16, Apr09, BKR20, GNK24, Int05a, Int05b, Int06b, Int06c, Int06a, Str05, AJM^{+}06, NSL^{+}06, NKK^{+}06, RSW^{+}06, Uhl06]. TEE [WFL^{+}25]. TeeRollup [WFL^{+}25]. Tele [HMS04]. Tele-lab [HMS04]. telehealth [WQG15]. Telum [BBB^{+}25]. template [WRX11]. Temporal [CDO24, CWdO^{+}06, ZZG^{+}23, WBW^{+}19]. temporal-difference [WBW^{+}19]. Ten [Mur69]. Ten-Sys [Mur69]. Tenancy [DY17]. Tenant [CDM^{+}25, LCZ^{+}19, SWW^{+}18, YKS16, ZRZY15]. Tensor [NLD^{+}23]. term [SOKE23]. terminal [CKT08]. terminals [IIK^{+}06, ISE08]. Terra [BSI^{+}15]. Terrestrial [QLL^{+}21]. TerrierTail [ASSB18]. Tesseract [ABG14]. Test [Kol19, LPSS19, MMHP17, NL19, SM06, ABDD^{+}91, IIPB09, LLS^{+}12]. Testbed [AHKC23, HLW^{+}10, ZGW^{+}06]. testbeds [ACM06c, ADWM18]. Testboard [Kut92]. Testing [Ame13, CQLL18, DKF94, GFB^{+}92, HLP^{+}16, Kao17, KLF^{+}15, MMP^{+}12, Ost94, SAM14, VS06, BD11, CSS^{+}16, CGC^{+}24, FCD09, FRM^{+}24, KFF12, SCFP00]. Texas [ACM75, ACM06d, USE01b, IEE02, IEE03]. Theia [BA23]. their [EF94, KCV11, SS13]. Them [HPHV17]. Theorem [Hir17, SSH17, BW03]. theoretic [NS17]. theoretic-based [NS17]. Theoretical [Kna93]. Theory [ZZW22, LWCZ22, WSAJ13]. Thermal [LCL^{+}23b, AJBJ23, CLL^{+}23, IRB19, KR16, LCL^{+}23a]. Thermal-aware [LCL^{+}23b, AJBJ23, CLL^{+}23, IRB19, LCL^{+}23a]. Thermostat [AW17]. Things [Gal09a, Gal09b, Gal11, ASL^{+}20, AAMBE21, CMG^{+}19, FR24, KB21, PTD^{+}18, SXH^{+}19, WSX^{+}19]. Third [Ano04b, CRZH15, PG74, PG73]. Third-Party [CRZH15]. Thoroughly [PRS26]. Thoth [KB17]. thousand [SK13b]. thousand-core [SK13b]. Thread [GCL^{+}21, MP01, BKC^{+}13, Ven97d]. threaded [FGG14, HC17, SE12, tTR82]. threads [UR15]. Threat [Aln22, SL16, YYY^{+}23, LMDP19, PST15b]. Threats [ZYH^{+}19, PZH13, vCPWvT11]. Three [YYPA01, Vit14, YZW^{+}13, ZFY18]. three-layer [ZFY18]. threshold [SENS16, TDG^{+}18]. threshold-based [SENS16]. Thresholds [TCTH23, XCSM18]. Throughput [BPP^{+}17, BCW20, GKXK13, PYYG21, GI12, ZSW^{+}06]. Throughput-Oriented [PYYG21]. Thunderbird [Joo06]. ticket [OL13]. Tier [KSVR23, LH15, WZKP19, XZL^{+}20, WDCL08, ZNSL14]. Tiered [GGK18, SLW^{+}24, AW17]. Tiered-Memory [SLW^{+}24]. Tightly [ZXG^{+}24]. Time [Bad87, BE17, CW03, CRX24, Cre65, FXHY21, FML^{+}22, Fuj91, GLL^{+}21, GPM21, Hu90, HWB03, HS06, JAD19, JYM^{+}23, KR18, KPHA20, LTE12, LWC^{+}17, LSSC22, LXL^{+}22, Mad69, MS70, NL19, PPG^{+}17, Sta97, WZKP19, ABB19a, AS76, AMIA19, ACT94, ABC^{+}07, BBS06, CGM17, DEE^{+}16, HK07, HcC14, Ive03, KJ13, KBB11, LD05, LTK17, MNT14, MMTM22, MA21, MA24, NBS18, PTD^{+}18, QCM^{+}24, QT06, RAT17, SBNU18, She91, Ste14, SCA^{+}24, TSLBYF08, TZY^{+}25, TMV25, WQG15, YK13, YCL^{+}19, YIR24, ZEdlP13, ZGL^{+}17]. time-based [TMV25]. Time-Constrained [LTE12]. time-sensitive [MMTM22]. Time-Sharing [Cre65, Mad69, MS70]. timebombs [CWdO^{+}06]. Times [ELC^{+}19, PLMA18]. Timing [Hu90, HWB03, KKS^{+}19, LGR14]. tiny [LC02]. TLB [OLZ16, RGSJ17]. TM [Qia99]. Tolerance [JKJ^{+}10, RZPX19, ZJXL11, RCOW12, SM23a, XXWG23, YLH14]. Tolerant [FK03, Kim84, YWR^{+}14, SNV10, YI24]. Tool [Ano03b, Wil01, KK79, Lia05, Ska07, Skr01, SCFP00]. toolkit [ACG18, DZ02, PW03]. Tools [AC98, BDG18, Cal75, GG11, LC09a, MJW^{+}06, PY93, QNC07, ACM01a, EL98, YYPA01]. Toolset [Ott18, PTD^{+}18]. Top [COV^{+}24, KMT14, PBWH^{+}12, Won97]. Top-Down [COV^{+}24]. topic [YZSC17]. Topics [IEE01]. topological [KKM^{+}13]. Topology [CYX^{+}17, TB17, dSdF16, AM16, PST15b]. Topology-Adaptive [CYX^{+}17]. Topology-Aware [dSdF16]. TOPSIS [SS19]. Toronto [Sof83]. TOSCA [BSNB20, BRS18]. TosKer [BRS18]. Total [LGJ^{+}18, THG^{+}18]. TotalStorage [D^{+}04]. TPC [NP13]. TPHOLs [BW03]. TPM [KC12]. TR [Int05b, Int06c, Int06a]. Trace [BA23, MZG14, NASD21, BDE^{+}03, DC15]. Traces [BA23, WKG17, DD20]. Tracing [KKD24, BT15, PFH^{+}16, WKJ15, Wol99]. Track [Shr89]. Tracking [JADAD06a]. Tractable [KR94]. Trade [MTJ^{+}22, SIdLB15, XZK^{+}20]. Trade-off [MTJ^{+}22]. Trade-offs [SIdLB15, XZK^{+}20]. Tradeoff [MTFK19, UTO13, WCY^{+}17]. Tradeoffs [CMM^{+}06a, CMM^{+}06b, CMM^{+}06c]. trading [LWLL16, VGL23]. Traffic [BBM^{+}15, CGC16, CYX^{+}17, DK17, LXL^{+}22, PCW^{+}16, SJMG24, VV18, YLTF20, CBJ22, FLL^{+}13, HH19, IKU15, JYOB18, LLZ^{+}19, MG19, WZV^{+}13, XHW^{+}19, YCL^{+}19]. Traffic-Aware [CGC16, CYX^{+}17, JYOB18]. traffic-intensive [IKU15]. Traffic-sensitive [DK17]. Transactional [URJ18, CMM^{+}06a, CMM^{+}06b, CMM^{+}06c, ZHCB15]. Transcendent [VTW16]. Transfer [HHC^{+}16]. transfers [DPBK16]. Transformation [Bos24, WIDP12]. transformations [HB08]. transient [LRC05]. Transiently [LDRS18]. Transition [MBWW86, Syr07]. Translation [AZEE17, AZEE18, JXL^{+}12, LH16, YVCB17, YVCB18, dGG^{+}17, CFG^{+}13, JYW^{+}13, Oi05, Oi06, Oi08]. translation-based [Oi05]. Translational [WIDP12]. translations [UTO13]. Transmission [RSNK17, RSN^{+}18]. Transparency [ZXL^{+}24]. Transparent [BZA12, FK03, JKJ^{+}10, KKD24, KKH14, MSI^{+}12, dGG^{+}17, AW17, JXZ^{+}10, MRC^{+}13, YJZY12]. Transputer [Boa90, GHH^{+}93, Boa90, GHH^{+}93]. travel [TSLBYF08]. Traveling [YK13]. traversal [YTS14]. Treating [SSOT17]. Tree [Hal79, KMMV14]. Trenches [HN10]. Trends [RG05, AH12, CCDR22, CM18, JPTE94, vD06]. TRI [ACM97]. TRI-Ada’97 [ACM97]. trie [SV17]. trie-based [SV17]. tries [SV15]. Trigram [Cox12]. Troubleshooting [WF03]. Troy [Ano97a]. truly [ZBS^{+}22]. trust [XJR^{+}17, RNA^{+}22]. TRUst-aware [RNA^{+}22]. Trusted [CDW^{+}24, DPW^{+}09, ETAB22, SVB93, Str05, DJP^{+}24, BCP^{+}08, KSLA08, WH08, YKP^{+}24]. Trustworthy [CDM^{+}25]. TrustZone [PPG^{+}17, PS19a]. TrustZone-Assisted [PPG^{+}17]. Truthful [NMG15]. TSAC [WZL15]. Tucson [IEE05]. Tuning [EDS^{+}15, RS16, WZL^{+}23]. Tutoring [GH91b]. TVDc [BCP^{+}08]. Twelfth [MR91]. Twenty [MS91b, Shr89]. Twenty-Fourth [MS91b]. Twenty-Second [Shr89]. TwinDrivers [MSZ09]. twins [HCJ07]. Twitter [Guy14]. Two [AW17, ASMA21, IMBB20, SSG90, TF16, BSSM08, CCMY07, HCJ07, LUL^{+}05, SZKY21]. two-dimensional [BSSM08]. Two-Level [ASMA21, SSG90]. Two-phase [TF16, SZKY21]. Two-tiered [AW17]. TX [ACM99]. Type [ADM98, AT16, Arv02, KCV11, PRB07]. type- [Arv02]. Type-Precision [ADM98]. Typed [G^{+}88, BDT13, GLV99, KRCH14]. Types [Wel94, GLW23, MFT^{+}19]. TypeScript [RSF^{+}15]. Typing [RSF^{+}15, RAT17].

u.v.a [Tho08]. UCSD [SP83]. UKCF [JXL^{+}12]. UKFaaS [CZH^{+}25]. umfassende [Bod10, Fis09]. Umgebng [CK06p]. Umgebung [CK06a, CK06e, CK06c, CK06d, CK06g, CK06f, CK06i, CK06h, CK06j, CK06k, CK06m, CK06l, CK06n, CK06o, CK06q, CK06t, CK06r, CK06s]. UML [Fre05, RFBLO01]. UMLexe [Fre05]. Un-IOV [ZXL^{+}24]. Uncertain [ZZW22]. uncertainty [LPBB^{+}18]. underlying [FBZS12]. understand [DMH18]. Understanding [FRM^{+}15, Set13, ZRZY15, LWB^{+}15]. underutilized [HM20]. Undocumented [Sch94b, Sch94a]. Unexpected [Par71]. Unfairness [SJA^{+}17]. Unhooking [AKCP21]. Unified [ZDS^{+}22, MBA^{+}12]. Uniform [Eug06, Bod88, LZZ^{+}25]. Unifying [MD12]. Unikernel [CZH^{+}25]. unikernels [MK23]. unique [AM16]. Unit [DCG12, PXG^{+}17]. United [Vra05]. uniting [LUL^{+}05]. Units [VLZL16, Vol90, ZXG^{+}24]. UNIVAC [Kam75]. Universe [Nel04].

Universities [Sta07]. University [ACM75, ACM81, Gre10, IEE96a, IEE97, IEE99]. UNIX [JJ91, KAH83, NSHW10, Gen86, HO92, Kal97]. Unknown [CLW^{+}14]. unleashed [Ano97d, HH08, MG08, MG09]. Unmodified [HLP^{+}16, MKKE12]. Unpicking [LBF12]. unreliable [MPM^{+}20]. unsound [AT16]. Untrusted [CD12, HKD^{+}13, HPHS04, WLL^{+}13, ZBP05]. upcalls [LD11]. Update [FXHY21, LC14, SCL^{+}19, VVC^{+}17, J^{+}05]. Updates [LCZ^{+}19, LDRS18]. updating [CCZ^{+}06]. upfront [ZLW^{+}19a]. upgrade [CHCC07]. Upgrades [Ano03a]. uptrees [HB13]. UPWN [M^{+}06]. Urgent [AGJS16]. USA [ACM81, ACM01a, ACM03b, ACM05a, ACM06c, ACM06b, ACM06d, Boa90, IEE93a, Shr89, USE01c, ACM75, ACM05d, ACM06a, Ano01b, Ano04b, IEE84b, Ost94, USE85, USE86, USE91, USE93, USE99, USE00a, USE01a, USE01b, USE06]. Usage [KLLT18, MB15, RSW^{+}06, WH99, ZXL^{+}24, KTB17, RGAT18, SOKE23, SK13c, SP23, YW20]. USB [Ano03a]. Use [AAAF21, Bec09, CLŁS12, Guy14, GGK19, KK79, Sch13a, SJJ^{+}12]. use-case [GGK19]. used [tTR82]. useful [LC09a]. usefulness [SM79]. USENIX [ACM05d, Sof83, USE91, USE93, USE06]. User [Chu06, KA24, Mur69, RB24, ZQCZ16, Ano93, ACT94, Bor07, CLM^{+}22, Guz01, PG11, RSC^{+}15, Sto07, Tho73, ZLZ13, ZLZ^{+}19a, CKT08, Dav04, RB24]. user-controlled [Sto07]. User-defined [RB24]. User-Level [Chu06, ZQCZ16, KA24, CLM^{+}22, ZLZ13]. user-space [PG11]. User-terminal [CKT08]. Users [Boa90, IBM76a, SS17]. userspace [DD20, Ste14]. Using [AAF^{+}09, AT24, ARAAA19, ASL^{+}20, ABV12, ALL06, Bas04, Bas06, Ben21, BA23, BRX13, CMGI^{+}23, CQLL18, Che21, CCO^{+}05, DBMI92, Don88, ESY^{+}17, Guz01, HLW^{+}10, HWHW18, JMSLM92, LJN^{+}00, LTT92, LD05, Mar73, MV16, MZ20, NASD21, OLZ16, PEC^{+}14, RSW^{+}06, ST24, SGK^{+}23, Sar01, See10, SM06, SC17, SYB12, SAT09, SBK15, SXCL14, TDG^{+}18, WDSW01, WKG17, WUNK17, WFL^{+}25, Wil01, Wol99, XSC13, XCSM18, ZLG^{+}20, ZBP07, ZLW^{+}19b, dGG^{+}17, AD18a, Ågr99, AA24, ATS16, AWR05, AP18, AGIS94, BSM^{+}12, BHvR05, BSOK^{+}20, CL14, CPM^{+}18, CJK^{+}12, CCZ^{+}06, Dan12, DHD20, DS22, EB20, FFBG08, FF22, FA21, FL13b, GSG22, GHK24, GHM^{+}18, GNK24, HKJ19, HJ10, HBNK24a, HBNK24b, HTAY21, HN08, HPHS04, Hol95, HPS22, JNR12, JWH^{+}15, JGSE13, JYOB18, Juo07, KSS24, KSS^{+}20, KSS^{+}23]. using [KKM^{+}13, KS18a, KJJ^{+}16, Kip21, KGS16, KL13, Kou11, KAS13, KRG^{+}12, LCLG26, LDL14, LLWW18, LQW^{+}12, MB15, MHM19, NGN24, NMC18b, NMC18a, NV05, PBL^{+}16, Pon19, RP07, RWC21, SEM^{+}20, SGV13, SSN12, SS22, SIJPP11, SIK^{+}16, SSH17, STFH15, SSN94, Str05, SAM14, TSLBYF08, TSR19, TF16, VT14, WGW^{+}18, WZZ^{+}20, YK13, YLWH14, YWF09, YWCF15, ZLZ13, ZDLG17, ZXB^{+}24, ZB18]. usual [dCJR16]. UT [Ren78]. Utah [ACM01a, CK87]. Utility [LGZ^{+}19, CSV15, JWH^{+}15, PSZ^{+}07]. Utility-Based [LGZ^{+}19]. Utilization [HLBZ20, KBDK22, KCKC15, NL19, uRQS20, WWL^{+}25].

Utilization-Based [NL19]. Utilization-prediction-aware [HLBZ20]. Utilizing [GVI13, KOY05].

V [Gal09b, Lar09, LC09a, Apr09, Car06, DK23, KVV09, KSS09, KS10, Lar09, LC09b, LC09a, MTJ^{+}22, MG08, MG09, SMP22, SVM^{+}23, SRS09, AJ18]. v-Mapper [AJ18]. V2E [YJZY12]. Validation [BSL^{+}18, FZD^{+}24, SSB14b, SSB01]. Value [TF16]. VaNetLayer [BTLNBF^{+}15a]. VAP [PM19a, XJW^{+}18]. vApp [SG10a]. variability [ASB18, BDQR23]. Variable [ADM98, Lam75, Oi05, Oi08]. variation [CCL^{+}20]. variation-aware [CCL^{+}20]. VAX [KZB^{+}90, LJZ12]. VAX/SVS [LJZ12]. vCache [KKH14]. vCloud [KMK10]. VCP [Khn09]. VCPU [WCC^{+}16a, IYAK23, WYZY24]. vCPUs [OLZ16, WCG21]. vCUDA [SCSL12]. VDE [GD08]. VDI [ZFH^{+}22]. Vector [Abr80, LRZ16, NTL^{+}24, WWS89, Ble89, SZ88]. vectorized [SZ88]. vectorizing [LRP^{+}19]. VEE [ACM05d, ACM06f]. VEEs [LCT^{+}15]. Vegas [ACM81]. VEhicle [RNA^{+}22]. Vehicles [SDM21]. vehicular [BTLNBF^{+}15a, MCC18, YBZ^{+}15]. Velox [TV18]. Vergleich [Zim05]. verifiable [CMP^{+}13, PK75b]. Verification [ABDD^{+}91, JE12, JES^{+}15, LZM^{+}20, ST24, SSB14b, ZL18b, ZLZ^{+}21b, BSD19, DL19a, FC98, LLS^{+}12, PBL^{+}16, SSH17, SSB01, ZSRR22]. verified [SV24]. Versatile [EBJ17, SN05b, ZXG^{+}24]. Version [Bru07, Sim92, WR07, WR08, Ano94, Ano14a, IBM96, MIS^{+}05]. versioning [STFH15, WF07]. Versus [Ran20, DK75, HPHS04, SCEG08, VED06]. vertical [BFS^{+}18, STY^{+}14]. Verwaltung [Zim05]. Very [RGSJ17, SSB03]. VFe [Ano05]. vGPU [LZM^{+}20, SM23b]. vGreen [DMR10]. VHDL [FS89]. VI [Int06b]. via [CLM24, FL13a, GI12, GLLJ16, HSK17, HB13, HTM^{+}24, HWR^{+}24, KJM^{+}07, KNHH18, LF19, LJL^{+}11, MSS^{+}15, NGRF19, NTL^{+}24, QZDJ16, RZPX19, SP83, SDD^{+}16, SCF^{+}25, TDG^{+}18, WZL^{+}23, XWJX15, YTS14, ZSW^{+}06, vSMK^{+}20]. viable [HW15]. viele [WR07, WR08]. vieles [Joo06]. View [GB19, KKH14, AD18a, Guy14, LDDT12]. Viewer [BA23]. Viewpoint [LPSS19]. Views [PW03]. Vigilant [PBYH^{+}08]. VIII [IEE01, IEE96a]. VIKOR [GSG22]. VINEA [EMW16]. ViNEYard [CRB12]. Violation [ZHL16]. violations [BSM^{+}12]. VirtCL [YWTC15]. virtio [Rus08]. Virtual [ACM05d, ACM06f, AGJS16, AS85a, ABCC66, AEM^{+}14, ADM98, AGH^{+}15a, AT24, AZEE17, AZEE18, AAAF21, AAR22, AAB^{+}05a, ACL72, ABV12, Ano75, Ano97b, Ano97a, Ano97c, Ano97d, Ano00, Ano01a, Ano01b, Ano02, Ano04a, Ano04b, Ano05, fLtNW14, AE01, Apr09, Arc07, AD11, AAK18, AY24, ASSB18, Att79, Att73, AH68, AJM^{+}25, ACA16, AC98, AMA^{+}11, BWP85, BFHW75, Bai70, Bak83, Bal91, BMS16, BYZZ20, BP99, BDF^{+}03, BBTK^{+}17, BDJdS02, BSSS14, BWH^{+}19, BDF^{+}99, Bee05, BCC^{+}15, BH73, Bel06, BB13, BN75, BJ20, BJC24, BHDS09, BJPS73, BBHL08, BL17, BFG^{+}14, BWD^{+}15, BBM^{+}15, Blu02, Bos24, BBM09, BD01, BP01, BP03, BZD17, Bro89, BRX13, BFM^{+}21, VMW^{+}19, BBS06, BJH^{+}16, B^{+}07, BG73a, BG73b, BCG73a, BCG73b, BG74, Caa00, CTS^{+}93, CW03, CCWY05, CL17a, CFH^{+}79]. Virtual [CFH^{+}80, CWL12, CFM17, CCML12, Car13, CK87, CFVP12, CWS12, CJK^{+}12, CHCC07, CGMD19, CF00, CT03, CSS^{+}13, CGC16, CL16a, CL16b, Che21, CRZH15, CCO^{+}05, CC77, Cla97, Coh97, CDG97, Cox09, Cra05, Cra06, Cra98, CH78, CWG00, CWL^{+}15, CHPY17, CYX^{+}17, CHLY18, CDN02, Dalxx, DAH^{+}12, Dal97, DHPW01, Dan86, DD20, DSM14, DG05, DEK^{+}03, Den01, DK17, DMR10, DKW15, DCM22, DF96, Do11, DGLZ^{+}11, Dom80a, DL19b, DJ76, DJ77, DCA04, DLS^{+}01, EGR15, EGJS15, ECJ^{+}16, ETAB22, Eng99, EM06, EGG^{+}24, EMAL17, EG01, Ert03, EMW16, EDS^{+}15, FXL^{+}23, FZD^{+}24, FFB^{+}00, FG91, Fie68, Fis01, FPS^{+}02, (Fo71, (Fo78, Fra98, FK03, FL13a, Gai75, Gal73, G^{+}01, GWZ16, GKSP99, Gei02, Gen86, GMGF24, Gol69, Gol71a, Gol71b, Gol73c, Gol73b, GGG03, GC00, GA25, GLBJ18, Gum83, HHV^{+}02]. Virtual [HHW10, HT98, Hal79, HTW^{+}19, Han73, HKLM17, HM01, HA79, HTB22, HLW^{+}23, HH79, HB17, Hin97, HKM^{+}18b, Hir17, Hof20, Hor73, HKKW13, HS13, HWB03, HS06, HB08, HPS22, HPP15, IBM72, IBM73, IBM76b, IBM85, IBM88, Int88, Ian14, Ibs84a, Ive03, JR02, JHS12, JJK^{+}11, JE12, Jen79, JXL^{+}12, JMSLM92, JQWG15, JAS^{+}15, JN15, JKJ^{+}10, JADAD06a, JDJ^{+}06, JJ02, Juo07, KCWH14, KRS^{+}17, KC16, KS08a, KSS^{+}20, KSS^{+}23, KMK16, KNT02, KKTM17, KF91, Ken80, KDB16, Kim84, KJL11, gKEY13, KKJL14, KP15, KPHA20, KAH83, Kov19, KGZ^{+}04, KAS13, KLLT18, KLF^{+}15, LCWB^{+}11, LMM18, Lam75, Lau87, LW73, Law00, LW11, LP14, LSC^{+}17, LMR18, LLW98, LMG00, LMG01, LTE12, Li14, LZL^{+}15, LZWD15, LVM16, LWLL16, LYYY17, LGJ^{+}18, LLWM23, LLY^{+}25, LB98, LV99, LTT92, LD05, LWW16, LXW^{+}23, LY97a, LY97b, LY99, LYxxa, LYxxb, LYBB13a]. Virtual [LYBB13b, LYBB14, LHAP06, LWLL10, LJL^{+}11, LW12, LJL^{+}15, LLZ18, LWZ^{+}18, LCZ^{+}19, LLS^{+}20, LSPP^{+}23, LZZ^{+}25, LPB17, LPBB^{+}18, LFBB94, Loy92, LTK17, LXM^{+}16, MSG14, Mac79, Mad69, MB15, Mal73, MS91a, MK13, Man15a, Man16, MSLY24, Mar73, MD12, MP16, MZ20, McG72, MRG18, Men03, MS70, MD97, MDxx, MW18, MDGS98, MLG^{+}02, MB98, MKKE12, MA21, MA24, II79, MP01, MJW^{+}06, MM94, Mur69, NBH08, NBK16, NMG15, Nel04, NASD21, Neu92, NGRF19, NLD^{+}23, NSJ12, NL19, Nou92, OT97, OKAM17, Oi05, Oi06, Olb78, PTHH14, PAKY16, Par71, Par72, PPTH72, PP73, PSBG11a, PAC^{+}22, PHXL19, PXG^{+}17, PNM^{+}20, PRB07, Pfo13, PHC20, PF23, PS16, PCC^{+}16, PK75a, Pro00, Qia99, QBL^{+}23, QT06, RNA^{+}22, RG17, Ran20, Ran02, RLZ^{+}16, Ren78, Rev11, RIP18, RY10, RK24, RI00, RSN^{+}18, RRB19, Ros99]. Virtual [Ros04, RG05, RS20, RCTY19, RB01, SMK02, Ibs84b, SL14, San88, SGK^{+}23, SSB^{+}14a, SD01, Say66, SH04, Sch13a, SMES01, Sch09, Sch94b, Sch94a, Sch73, See10, Set13, SMSB11, SSB03, SC17, SCEG08, SCSL12, SMA18, Shi03, SM01, SGV12, SV13, Sim92, SCP93, Siv04, SSG90, SN05a, SN05b, SHZ^{+}14, SBP^{+}17, SXMX^{+}18, SB73, Sta97, SSB01, SSB14b, SHB^{+}03, SVL01, Sun95b, Sun95a, SUN97, JCV99, SKI^{+}17, Sup04, SM02, Sur01, TSLBYF08, Tai98, TT96, TTH^{+}19, TMV12, THB22, TY14, Tol98, TO96, TV12, USE01c, USE01d, USE02, UT87, UBF^{+}98, UR15, Vag10, VTW16, Ven97a, Ven99a, VGF16, VL00, Vog03, Vol90, WL96, WIDP12, Wak99, WH99, Wal99, WDL^{+}20, WB81, WLW^{+}15, WZH^{+}16, WWL^{+}17a, WGXC25, Wel94, WGLL13, WZL15, WLLZ16, WCSG05, WHD^{+}09, Win71, WP97]. Virtual [Wol99, Won97, WWMG06, WLCS17, WWL^{+}17b, XKY^{+}11, XSC13, XHL^{+}13, XWJX15, XLL^{+}14, XLJ16, XLWX19, XLL^{+}20, YC98a, YLH17, YWY^{+}17, YLCH17, YLQL25, YWH^{+}21, YP15, ZWFX17, ZDK^{+}22, ZS01, ZLW^{+}14, ZRD^{+}15, ZRS^{+}16, ZL16, ZCG^{+}17, ZL18b, ZLZ^{+}19b, ZZW^{+}21, ZCL^{+}21, ZZF06, ZWL^{+}18, ZWC^{+}23, ZLL^{+}16, Zho10, ZHL16, ZYLY18, ZZW22, ZJXL11, ZTWM17, Zim05, ZR06, Zyt94a, Zyt94b, dSdF16, vD00, vLSM01, Ågr99, AEMWC^{+}12, ABB19a, Abr82, AS85b, AD19, AGSS10, AJBJ23, AAH^{+}03, AGH^{+}15b, AHRR22a, AHRR22b, AT23, ATZP21, ADA^{+}19, AAB^{+}00, AAB^{+}05b, AA24, AC95, Ame13, AGH^{+}16, Ano94, Ano96, Ano99a, AO16, ATS16, AFT01, ABC^{+}07, Arm98, AWR05, AAM^{+}16, AMAB17, Arv02, AP18, AS14, AMB^{+}17, AAC^{+}17, ANH00, BB20, BAC15, Bag76, BML^{+}13, BSM^{+}12, BDF^{+}98, BDS^{+}09, BDQR23, BHvR05, BG20, Beg12, BPC94, BMF23, BB12]. virtual [BB15, BCP^{+}08, BJ22, BCM90, BRS^{+}22, BPM^{+}22, Bir94, BADM06, BFC02, BY20, Bri98, BB95, CSMB15, CARB10, CL14, CL17b, CD14, Car14, CEG07, Cav93, CS76, CGM17, CSSE21, CCL^{+}17, CCL^{+}20, CLL^{+}23, CGC^{+}24, CBLFD12, CH08, CRB12, CK06a, CK06e, CFRSSR19, Cof99, CGV10, dCCDFdO15, CWdO^{+}06, CLDA07, CLL^{+}13, CD01, DPW^{+}09, DSR23, DDS^{+}94, DS19, DSC^{+}08, DP11, DM93, DC15, DEG^{+}17, DBC^{+}00, DQLW15, DLH^{+}20, Don87, DHD20, DXM^{+}17, DAdBM^{+}24, DSZ11, DCMW17, DS22, DCA17, EB20, EGD03, EYG21, EGKP02, EG03, Ert05, EL98, EMS15, FCD09, FGZC23, FLL^{+}13, FZS^{+}20, FS19, FF22, FM90, FA21, FBZS12, FSFP19, FMIF18, Fit14, FHL^{+}96, FGLI15, FF96, FLM^{+}08, FCG^{+}05, Fre05, FX06, Fu10, FRM^{+}24, GP13, GGQ^{+}13, GTGB14, GSG22, GI12, GVI13, GH20, GHK24, GSKJ18]. virtual [GJK^{+}20, Gol73a, Gol74, GCARPC^{+}01, GAHL00, GPW03, GR80, GBCW00, GLQ^{+}13, GKJ^{+}19, GLW23, GLV^{+}10, GA18, HKS19, HM18, Hal09, HMH17, HZL^{+}18, HJ10, HKN22, HBNK24a, HBNK24b, HN08, HKB19, HZZ^{+}14, HTB19, HL24, HUL06, HH18, HH19, HAK22, HDG09, HcC14, HPHS04, Hol95, HLBZ20, HSC15, Hui18, HPS23, IBM94, IBM96, IRB19, IKU15, IMBB20, JSK^{+}13, JK15, JES^{+}15, JKK^{+}13, JNR12, JWH^{+}15, JC18, JGW^{+}11, JDW^{+}14, JGSE13, JYOB18, JADAD06b, JB24, dCJR16, Kal97, KOY05, KSS24, KDK20, KBDK22, KB21, KSSG16, KS20a, KSO^{+}15, KRCH14, KS18a, KZ21, KS18b, KTB17, KK21, KBB11, KCS14, KJLY15, KCKC15, KKC^{+}16, KNHH18, KKK^{+}18, KMG^{+}18, KFF12, KHA22, KF18, KSS^{+}18, Kou11, KCV11, KBC21, KR16, LBP^{+}07, LCLG26, LMJ07, LSS24, LBZ^{+}11, LC02, LM99, LC14, LZC^{+}16, LBL16, LYYY18, LLF^{+}18, LLWW18, LFHQ19, LXRS19]. virtual [LLZ^{+}19, LZLY20, Lia05, LJL12, LQW^{+}12, LF19, LC13, LL14, LTZ^{+}14, LPZ^{+}22, LWCZ22, LCL^{+}23a, LCL^{+}23b, LMDP19, Lot91, LSS04, LG93, LFHS23, LQD^{+}18, LY23, MSG^{+}12, MR23, MD73, MD74, MSG01, MNB24, DPBK16, Man15b, MS17, Man18, MRM06, MBM09, MNA16, MS00, Mat09, MK19, MK23, MN03, MC93, McM11, MG13, MRG17, MN91, MMTM22, MST^{+}05, hTMAC^{+}08, MHM19, MPM^{+}20, EYGS19, MAK07, NZH20, NGN24, NNK21, NK22, NK10, NOK^{+}85, NAR19, NOR15, NV05, NIA18, OG16, Oi08, OMB^{+}15, ORPS09, PKS^{+}19, PFH^{+}16, PEL11, PSBG11b, Par25, PMC05, PM19a, PDM20, PFPJ18, PBYH^{+}08, PJZ^{+}19, PCB^{+}18, Piz17, Pon19, PRS16, PRS26, PV08, Pul91, PS23, uRQS20, RK16, RKT20, RH17, RHR20, RHR02, Raj79, RG19, RWC21, RT18, RZ14, Req03, RK18, RFBLO01, RJK^{+}17, RI25, RGS^{+}20, Rus08]. virtual [SZKY21, SBI21, SJB14, SS13, SENS16, SBBP20, SHR19a, SHR19b, SNV10, Sch13b, SSMGD10, SEM^{+}20, SHLJ13, SSN12, SM23a, She91, SJJ^{+}12, SJW^{+}13, SWH^{+}13, SASG13, SLC20, SSEA18, SS19, SL00, SS22, SGGB99, SGGB00, SKC73, Smi97, SYMA17, SJL20, SSL^{+}13, SPAK18, SMA^{+}10, Spi06, SOKE23, Ste14, SSU^{+}12, Str13, Str05, SZL^{+}14, SK13c, SLA^{+}16, SP23, SHTE11, SPSL25, Syr07, TZK17, THH^{+}14, TZY^{+}25, TMLL14, TDD20, TSR19, Tay76, TK20, tTR82, TGCF08, THG^{+}18, TIIN09, TMMVL12, TB14, TMV25, TDG^{+}06, Tsa14, TtLcC13, Tur84, UIK19, Vac06, Van98, VT14, Ven96, Ven97b, Ven97c, Ven97d, Ven99b, VED07, VVB13, VWT13, VGL23, VDO14, WGF11, WKT08, WRX11, WZV^{+}13, WQG15, WKJ15, WHC16, WCY^{+}17, WXZ^{+}17, WSX^{+}19, WBW^{+}19, WZZ^{+}20, WGY20, WWL^{+}25, WR07, WDT18, Web10, WK08, WLG^{+}11, WHW20]. virtual [WH08, WCS06, WLL^{+}13, WW77, WSVY09, WRSvdM11, WRS^{+}15, WCG21, WCZ^{+}23, XNH21, XCJ^{+}14, XHW^{+}19, XHCL15, XJWW15, XZZ^{+}16, XWX^{+}17, XYYY17, XTB17, XLQL18, XLWZ18, XJW^{+}18, XZK^{+}20, XA22, XXWG23, YC98b, YME05, YZW^{+}13, YLH14, YLHJ14, YPLZ17, YCL^{+}18, YW20, YGLY21, YBZ^{+}15, YYC^{+}19, YWH^{+}23, YLK^{+}10, Yel99, YWF09, YSM^{+}21, YLJ22, YC16, YIR24, YB24, YRJ18, YMY17, YGN^{+}06, YWGH13, YQZ14, YQZ19, YTY00, ZG13, ZXW16, ZWKX17, ZYZ^{+}18, ZBG^{+}05, ZLZ15, ZLH^{+}15, ZWHC17, ZHHC17, ZFY18, ZWC^{+}19, ZLZ^{+}19a, ZJRW19, ZBP05, ZBP07, ZWL09, ZLW^{+}19a, ZFL^{+}23, ZL13, ZLLL13, ZWH^{+}17, ZLCZ18, ZSRR22, ZWC^{+}14, dSOK17, AGIS94, BPB86, CBJ22, Cza00, Fuj91, GKP^{+}19, GHM^{+}18, KM13a, KM13b, McC74, Mon97, PEC^{+}14, Ros99, RMP24, TSN^{+}23, VED06, Wel02, YI24]. virtual-machine [HUL06, HPHS04]. Virtual-Machine-Based [JN15]. virtual-time [She91]. Virtualbox [Deu08, Bec09]. Virtualisation [Ska07, Apr09, Rob12, SDN09]. virtualise [DJP^{+}24]. virtualised [MPF^{+}06]. virtualisierte [Mar08, Kar07]. Virtualisierung [Spr06, Spr07]. Virtualisierungs [Tho08]. Virtualisierungs-Buch [Tho08]. Virtualisierungslösung [See08a]. Virtualisierungslösungen [PO09]. Virtualisierungssoftware [Zim05]. Virtualisierungssystemen [Deu08]. Virtualities [Den01]. Virtualizable [GG72, HH13, PG74, PG73]. Virtualization [AFG^{+}17, AJM^{+}06, AP22, AAJD^{+}16, AVNR19, ASL^{+}20, AAT^{+}22, Aln22, AAMBE21, ADWM18, APST05, Ano03b, AvMT11, Bac11, BE17, BLMP22, BJG19, Ble10, BHEP14, BDR^{+}12, CZL08, CLS07, CGS06, CEPR22, CFB24, CHW12, CXLX15, CWH^{+}16, CMK^{+}16, CQLL18, CD12, CDD13, cCWS14, CLŁS12, Chu06, Coh10, Cre09, Cre10b, CGW07, DPCL22, DLLN18, DMS02, DW14, DPCA11, DLM^{+}06, Don06, DMG^{+}15, DK23, DY17, ECET18, EMAL17, ELC^{+}19, FPR^{+}06, FR24, Fer11, FDF05, FRD^{+}08, FLZ17, Gal09a, Gal11, GHS17, GW07, GCL^{+}21, Got07, GG11, HD16, HYK^{+}23, HWF07, HTAY21, Her06, HN10, HHC^{+}16, HSN17a, HSN17b, HDM08, HSL17, HB12, HW12, JAD19, JYM^{+}23, JW17, KHW^{+}16, KLY20, KS08a, KBBM15, KSVR23, KMM13, KR18, KS08b, KKS^{+}19, KGS16, Kot10, Kot11, KC12, KLR^{+}20, KLK^{+}22, LH16, LWC^{+}17, LXL^{+}22, LLW^{+}16, LRZ16, LZW^{+}17, LYGG20]. Virtualization [LCFL12, LDDT12, MZD^{+}18, MDZ^{+}21, MCC18, MA10, MCZ06, MUKX06, MA17, MMHP17, MGL^{+}17, MTJ^{+}22, MWHH05, NTR18, NSL^{+}06, NKK^{+}06, NsP16, NVV^{+}24, OVI^{+}12, PZW^{+}07, PHL^{+}12, Pap20, PGP19, PDL^{+}23, PM19b, PZH13, PYYG21, PYDG22, PvDS08, PNT12, PST^{+}15a, QCM^{+}24, QNC07, QTR21, RC18, RSW^{+}06, RCM^{+}12, R^{+}06, RTL^{+}18, RZPX19, RB24, RKRK17, RWX^{+}12, RR09, SMP22, SVM^{+}23, SADP21, Sed07, SM06, SLW^{+}24, SN23, SGB^{+}16, SYB12, SABL20, SAT09, SLZ^{+}25, SIJPP11, SYC14, SWF16, Spr07, Sta07, SFSN^{+}24, SAM14, SKYK16, Swa06, TDMP23, THLK10, TF16, Tre05, UNR^{+}05, Uhl06, UVL^{+}13, VN06, VN08, WBB^{+}16, WDCL08, WWH^{+}16, WZT19, WC01, WG07, WHD^{+}16, WH05, WLW^{+}17, WZL^{+}23, XH16, XYD^{+}18, XML^{+}18, YLT^{+}23, YSS^{+}17, ZDS^{+}22, ZZG^{+}23, ZSXZ07, ZQCZ16, ZYH^{+}19, ZSP^{+}21, ZZW^{+}21, ZXL^{+}24, ZZF06, ZAI^{+}16, ZXY^{+}15, ZLW^{+}19b]. Virtualization [ZKWH17, dGG^{+}17, vMAT14, vdK09, AA06, AKK^{+}07, AAF^{+}09, A^{+}04, AH12, AMIA19, ALW15, AJD09, Ano14c, Ano15, AKCP21, Apr09, AAB^{+}05c, AEB19, ABB^{+}19b, AA18, ABB^{+}15, BDF^{+}03, BBD^{+}10, BSL^{+}18, BRIdM10, BKR20, B^{+}05, BB08, Bor07, BH13, BC10, BTLNBF^{+}15b, BTLNBF^{+}15a, BSMF08, B^{+}07, CPM^{+}18, CSSS11, CMGI^{+}23, CMG^{+}19, CBER09, CDM^{+}10, CFG^{+}13, CWH^{+}14, CL15, CCZ^{+}06, CCMY07, CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, CLM^{+}22, CB10, CMM^{+}06a, CMM^{+}06b, CMM^{+}06c, Cia07, Cla05, CLM24, CBFH20, CM18, CKT08, Cre08a, Cre08b, Cre10a, CB07, DLL^{+}16, DBO^{+}18, DYL^{+}12, DCP^{+}12, DS09b, Dre08, EBLM22, EdPG^{+}10, ECAE13, FFBG08, FMX^{+}25, FP14, FJKK17, FLCB10, FS08, Fro13, FK13, FSH^{+}13, GMK17, GLA^{+}08, G^{+}06, GEK25, G^{+}05, GTN^{+}06, GAH^{+}12, GKT17, HLW^{+}10, Hal08, Han16, HIIG16, HHSG18, HPcC04, HC12, HZK^{+}25]. virtualization [IIK^{+}06, ISE08, IMK^{+}13, IPRS21, J^{+}05, JM08, JXZ^{+}10, JCZZ13, Kao17, KVV09, KSRL10, KKB14, Kip21, KWZ^{+}19, KL13, KS20b, Kro09, LPD^{+}11, LD11, LUL^{+}05, LLLE17, LLW^{+}12, LZWC13, LLY^{+}18, LLX^{+}17, LJYZ15, LQW^{+}12, LCL14, LWL16, LRP^{+}19, LLS14, LP11, LDL^{+}08, MG19, MB21, MRM06, MSI^{+}12, MDD^{+}08, MIS^{+}05, MBA^{+}12, MPA^{+}18, MBBS13, Mly09, Mon22, MMG^{+}18, MR06, MHS21, NTH^{+}17, NRdA^{+}20, NB11, P^{+}08, PG11, PBB13, PMP23, PFNC20, PST15b, PDML25, QZDJ16, RSC^{+}15, RS16, RSV24, RQD^{+}17, Rix08, RSLAGCLB16, Ros06, Rou07, Sam22, SVN^{+}10, SJRS^{+}13, SWcCM12, SM23b, SIRP17, SPF^{+}07, SHB19, SWW^{+}18, SAB^{+}07, SWC08, SWL^{+}23, SL12, TDG^{+}18, TZB19, TXD^{+}24, TMJ^{+}21, TP25, TSCB19, TLBW12, VW08, VSC^{+}10, VOS12, WR12, WZW^{+}11, WCC^{+}16a, WCC16c, WCS09, WJGA12, WHSE15, WYZAD20, XKY^{+}11, XZ11, YKS16, YJZY12]. virtualization [YTS14, YLH14, YLWH14, YCL^{+}18, YCL^{+}19, YLTF20, YXL^{+}20, Yu20, ZEdlP13, ZSR^{+}05, ZSW^{+}06, ZLZ13, ZXB^{+}24, ZZW^{+}24, vCPWvT11, vD06, vH08, Gua14, BCZ19, MCJ19, VSMC23, YWL^{+}18]. Virtualization-Aware [LXL^{+}22]. Virtualization-Based [CDD13, KLR^{+}20, RZPX19, AAJD^{+}16, DPCA11, MCC18, WDCL08, CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, LLX^{+}17, QZDJ16, TSCB19]. virtualization-driven [CSSS11]. Virtualized [AMA18, ASMA21, AIAR^{+}24, BB17, EGR15, GKXK13, GLBJ18, HO22, HBL^{+}10, HLPY16, HCB18, HTM^{+}24, KHW^{+}16, KKH14, LZ15, LGJZ16, MT16, MT17, MSC^{+}21, NBB^{+}19, NKY^{+}18, NSC^{+}22, PWJ16, PLZ20, RGSJ17, SB16, SL16, SDD^{+}16, WIS^{+}15, WKC^{+}09, WLMD16, WTM18, XWW^{+}21, YVCB17, YVCB18, YWW^{+}15, YWCF15, ARA18, ARA20b, ARA20a, AJH12, ATS14, ACG18, ASB18, BGS13, BSD19, BKT^{+}19, BSOK^{+}20, BSSM08, CP17a, CRX24, CP17b, CDO24, DS18, EBJ17, GPS^{+}18, GSG22, GGK19, HOKO14, HL13, JK17, KW13, KSRL10, KRG^{+}12, LKR^{+}19, LWM14, LC13, MNT14, MAK18, NBS18, NS07, NMC18b, NMC18a, PSZ^{+}07, PC21, PSC^{+}07, QXH18, RAP19, RHZ^{+}17, SBNU18, SG10b, TRG13, WWWL13, WB16, WTLS^{+}09, WTL^{+}16, YLJ22, ZGL^{+}17, ZWC^{+}14, YJZ^{+}21]. Virtualizing [BTMS10, CCDR22, Sar16, SB10, SVL01, WRS13]. VirtualKnotter [ZWC^{+}14]. Virtually [Say67, Spi06, WL96, Tre05]. VirtualPatch [PDML25]. VirtualPower [NS07]. virtuelle [WF03, WR07, WR08, Zim05, Zim06]. virtuellen [CK06a, CK06e, CK06c, CK06d, CK06g, CK06f, CK06i, CK06h, CK06j, CK06k, CK06m, CK06l, CK06n, CK06o, CK06p, CK06q, CK06t, CK06r, CK06s]. Virtuelles [AH68, Han73]. Virtuoso [DGLZ^{+}11]. VIRTUS [IIK^{+}06]. Vision [Arm78, SCC^{+}23]. Visual [Fra06, Fra09, MC98, Wil06, Hee07, Hog06, Hog08]. Visualization [Nel04]. Visualize [BA23]. Visualizing [WT91]. vKernel [HWR^{+}24]. VLAN [LYY^{+}24]. VLAN-VxLAN [LYY^{+}24]. VLISP [Ram93]. VLSI [IN87]. VM [Ano01a, Ano04a, Ano04b, FAA17b, AH24, Ano03a, AB16, ABG14, Att79, Bar73, Bar78, BCW20, BN89, BT15, Boz89, Cal75, CBZ^{+}16, CCW^{+}20, Com82, CTP^{+}17, DS20, ESY^{+}17, FAA17a, FMJ15, Fis91, FGG14, FL13b, GH91a, G^{+}06, GHD12, GPR23, GNK24, HM20, HKM^{+}18a, HKJ19, HXZ^{+}16, HC12, HW15, IBM94, IYAK23, JFPL16, JFZL17, KN18, KA25, KU24, LPSS19, LYY^{+}20, LBF12, LJZ12, LWLL10, LSX^{+}24, MK22, MSS91, MLA83, MA19, NOK^{+}85, NS17, Olb78, OJG91, P^{+}08, PDM20, PG17, PG18, RAT17, RSNK17, RJS^{+}18, STMV18, SSG^{+}20, SHW^{+}15, SM79, SSS25, SV24, SM23b, SBK15, SNC91, SIdLB15, TB17, TUM18, TV18, Var91, Wal10, WBHN18, XCSM18, YZLQ14, YKM17, YJZ^{+}21, YWR^{+}14, YKP^{+}24, ZFL15, ZWFX17, ZDLG17, ZLSI17, ZFL^{+}23]. VM-Agnostic [IYAK23]. VM-based [ESY^{+}17]. VM-protected [GHD12]. VM-scaling [AB16]. VM-to-hypervisor [NS17].

VM/370 [Att79, Bar73, Bar78, Cal75, Com82, Olb78, SM79]. VM/4 [NOK^{+}85]. VM/application [LBF12]. VM/ESA [Fis91, IBM94, MSS91, OJG91, SNC91]. VM/Pass [MLA83]. VM/Pass-Through [MLA83]. VM/XA [BN89, Boz89, IBM94]. VMBackup [ZXW16]. vmBBProfiler [TZK17]. VMbuddies [LH15]. VMDFS [SSEA18]. Vmgen [EGKP02]. VMI [LLF^{+}18]. VMIFresh [DSR23]. Vmknoppix [Deu08]. VMM [AD18a, ALL06, Car14, DQR^{+}13, DLX^{+}17, KZB^{+}90, LD11, LHAP06, OLZ16, RQD^{+}17, SM90, TUM18]. VMM-based [ALL06]. VMM-Bypass [LHAP06]. VMM-to-guest [LD11]. VMMB [MKKE12]. VMOR [MSI18]. vmOS [LLX^{+}17]. VMP [JNR12, PAC^{+}22, RK24]. VMP-ER [RK24]. VMPlanner [FLL^{+}13]. VMPlants [KGZ^{+}04]. VMPP [Loy92, LG93]. VMs [KMT14, KKJ^{+}13, PLMA18, RJK16, RB24, SEPV19, VS19, ZB18]. VMScatter [CLL^{+}13]. VMSI [ZTWM17]. VMSL [LSPP^{+}23]. VMThunder [ZLW^{+}14]. VMWare [Joo06, CK06f, Ham07, Khn09, KGG00, Tho08, Zim05, Zim06, Bas04, Bas06, War05, Wil01, AAH^{+}03, Ano03a, Ano03b, Ano07, BBD^{+}10, Bau06c, Bor01, BDR^{+}12, CK06f, Com00, Com03, DS09b, D^{+}04, Gal09b, GKBB15, Hal08, Hal09, Her10, HMS17, IIPB09, Kis08, KMK10, Lav10, Low08, Low09, Low11, LMG^{+}14, MRM06, MBM09, McC08, MWHH05, MJW^{+}06, Ng01a, Ng01b, NL00, OH05, PPS^{+}18, Ros99, Rul07, R^{+}02, See10, SIK^{+}16, SVL01, Ten17, TH10, Wal02, Wal99, War02, WF03, War11, Zim05, Zim06, B^{+}07]. VNC [RSLAGCLB16]. VNE [WBW^{+}19]. VNE-TD [WBW^{+}19]. VNET6 [GLQ^{+}13]. VNF [BMJ^{+}22, LKIL19, LW20, SJ21, XZL^{+}20, ZJRW19, ZLZ21a]. Vnode [KKD24]. VoIP [Mon22]. Vol.II [Shr89]. Volatile [AMH^{+}16, HN08, WZL^{+}18]. Volatility [WZL^{+}18]. voltage [TDG^{+}18, AMAB17]. Volume [AvMT11]. Vorstellung [CK06b, CK06e, CK06c, CK06d, CK06g, CK06f, CK06k, CK06m, CK06l, CK06n, CK06o, CK06q, CK06t, CK06r, CK06s]. VPC [KJM^{+}07]. VPFS [WH08]. VPN [MSI^{+}12]. VR [GWZ16]. VR-Cluster [GWZ16]. vs [Gal09b, Mad69, RB24, SFSN^{+}24, WKJ17]. VSA [SHLJ13]. vSAN [FKZ17]. VSched [LD05]. vSDN [ZWZ20]. vSFC [ZLZ^{+}21b]. Vshadow [WLW^{+}17]. VSim [RPE12]. vSphere [Gal09b, Lav10, Low09, LMG^{+}14, Fit14, Hal09]. vSphere5 [Low11]. VSPIM [NTL^{+}24]. VSwapper [ATS14]. vSwitch [TSP17]. vulnerabilities [RY10, YSM^{+}21]. Vulnerability [CRZH15, Ano99a, JKDC05, PM19a, PMP23]. vulnerability-based [PMP23]. vulnerability-specific [JKDC05]. Vulnerable [JSHM15, JAS^{+}15]. VxLAN [LYY^{+}24].

W [ALW15]. W-SDNs [ALW15]. WA [ACM05c, LCK11]. Wale [DSS19]. walks [AJH12, BSSM08]. WAN [KKK^{+}18, TDG^{+}06, WLG^{+}11, WRSvdM11, WRS^{+}15, ZFY18]. WAPPEN [Kag09]. WARDuino [LSS24]. warmup [BBTK^{+}17]. Washington [ACM06b, Ost94]. Wasm [CDM^{+}25]. Wasm-Based [CDM^{+}25]. WASMBOX [CDM^{+}25]. Watches [BDG18]. Watchmen [BDG18]. wavelength [AM16]. wavelength-routed [AM16]. Way [FLZ^{+}20, Ble10, Com00, WGF11]. WBATimeNet [MK22]. weak [RO16]. web [BTLNBF^{+}15a, DAdBM^{+}24, YSM^{+}21, Ano96, CVWL13, DF96, FF96, Kag09, LGJZ16, SJJ^{+}12, SDD^{+}16, VP16, WDCL08, YML^{+}18]. web-based [YSM^{+}21, CVWL13, Kag09]. Web/Java [FF96, Ano96]. Web/Java-based [FF96, Ano96]. WebAssembly [LSS24, MW25, SMR24]. Weight [WWL^{+}17a, HB08, YGN^{+}06]. Weir [BMER14]. Welfare [ZHW^{+}17, LWLL16]. Well [WC01]. Well-Conditioned [WC01]. Werkzeugen [KGG00]. Which [MS17, War80]. Whispers [WXW15]. White [LKL^{+}19]. Who [BDG18, LS15]. whole [BBM09]. whose [BBS06]. wichtigsten [CK06b]. Wide [BFG^{+}14, DF96, HS19]. Wide-Area [BFG^{+}14]. wie [Deu08]. WiFi [XKY^{+}11]. Wild [Cox10, STS^{+}13]. Win [War11]. Win4Lin [Ng01b, Ng01a]. WinCE [Kal97]. Windows [Bod10, Bor01, Joo09, Lar09, Sch94b, Sch94a, WF03, Apr09, Bod10, Car06, CK06a, CK06i, CK06h, CK06p, GMR93, KSS09, KS10, Lar09, LC09b, LC09a, MG08, MG09, Nou92, Sal92, YGN^{+}06, Zyt94a, Zyt94b]. WINRar [Joo06]. wired [XKY^{+}11]. Wireless [ACM06c, AFG^{+}17, ALW15, AAT^{+}22, BSI^{+}15, HLP^{+}16, KKTM17, SIJPP11, YMY17, DCA17, FK13, HLW^{+}10, WYZAD20, XKY^{+}11, XJW^{+}18]. Wirth [BGP00]. Wise [ZXG^{+}24, SEPV19]. Within [RD90, BTLNBF^{+}15a, YJZ^{+}21]. without [CD01, KSRL10, SUH86]. WLAN [KKTM17]. Wolves [DLX^{+}17]. WOMP [M^{+}06]. Work [HMS17, PWJ16, DMH18, KHL17]. Work-Stealing [PWJ16]. worked [Cox12]. Workers [VP16]. Workflow [MB20, GAHL00, HKS19, KCKC15, RWC21, WKT08, WCG21]. Workflows [AD18b, RB17, dCCDFdO15, EB20, FGG14, MPM^{+}20, QXH18, WB16, XYYY17]. Working [NKY^{+}18, ZDLG17, G^{+}88]. Working-Set [ZDLG17]. Workload [IEE02, IEE03, MA19, NASD21, PYYG21, SSB^{+}16, YWW^{+}15, ZWFX17, ZFH^{+}22, EBJ17, KCV11, SS13, SSN12, SLC20]. Workload-Aware [PYYG21, ZWFX17, EBJ17, SSN12]. Workloads [BB17, DS09a, GTGB14, IPRS21, LFHQ19, LL14, MK23, SMSH18, SMA^{+}10, SWC08, VVB13]. Workshop [ACM98, RM03, ACM05b, IEE01, IEE02, IEE03, IEE04, Mat10, Tho93, ACM01a, ACM04a, ACM06c]. workshops [M^{+}06]. Workstation [Bau06c, Bor01, BDR^{+}12, WF03, War05, SSN94, War02, SVL01]. World [AAR22, Ben21, DF96, GHH^{+}93, WLW^{+}17, BBM09, STS^{+}13]. World-Wide [DF96]. worlds [AJD09, LUL^{+}05]. Worm [CLW^{+}14]. Worst [HWB03]. Worst-Case [HWB03]. Write [ZCJ^{+}21, LFHQ19, LXRS19]. write-intensive [LFHQ19]. Writes [ZLL^{+}20]. Writing [Wes98]. written [MSG01]. WWC [IEE03, IEE02]. WWC-5 [IEE02]. WWC-6 [IEE03].

x3950 [R^{+}06]. X64 [dGG^{+}17]. x86 [AGSS10, BDR^{+}12, Cof99, MT16, MT17, MGL^{+}17, Rev11, AA06]. XA [BN89, Boz89, IBM94]. Xbox [Ste05]. XC [GH91a]. XEN [Hin08, PO09, Deu08, Kar07, Mar08, See08a, Tho08, RHM08, AJD09, Ano15, BDF^{+}03, B^{+}07, CBZ^{+}16, Chi08, CDO24, CGW07, De 06, DLM^{+}06, Don06, Fis09, Hab06, HWF07, HHH04, IGBKR19, Kar07, Kel06, LXL^{+}22, MDD^{+}08, MKM^{+}08, MST^{+}05, MCZ06, NB11, NOT^{+}17, PO09, PRS16, QT06, RHZ^{+}17, SJV^{+}05, SHLJ13, Spr06, Spr07, TC10, VS06, WG07, dSOK17, vH08]. Xen-based [CBZ^{+}16, dSOK17]. Xen-Basis [Kar07]. Xen-virtualisierte [Mar08]. XenEnterprise [CGW07, WG07]. XenExpress [CGW07, WG07]. XenServer [CGW07, WG07]. Xeon [GGK19]. XHive [KJL11]. XHPC [M^{+}06]. XINU [BWP85]. XIVE [AA18]. XML [Int06c, Kha19]. XPL [Kam75]. XSA [Ano15]. XScale [CMP^{+}07]. xSeries [R^{+}02]. XTREM [CMP^{+}07].

yang [CBGM12]. Years [FS12, BJG19]. yieldpoint [LWB^{+}15]. yin [CBGM12]. York [ACM03b, IEE90b, IEE96b, IEE90b]. Yountville [Tho93]. Yourself [AZEE17, AZEE18].

z [G^{+}06, P^{+}08]. z/VM [G^{+}06, P^{+}08]. z13 [ABB^{+}15]. Zero [AMH^{+}16, LZZ^{+}25, CHCC07]. Zero-Cost [AMH^{+}16]. Zero-Knowledge [LZZ^{+}25]. zeroloss [CHCC07]. Zeroing [GPS^{+}18]. ZNET [UBL^{+}82]. Zone [ETAB22, KBBM15]. Zone-based [ETAB22]. ZSim [SK13b]. ZSM [XKKL23]. zur [KGG00, See08a]. Zytaruk [Sch94b, Sch94a].

References

Adra:2004:APV

[A^{+}04] Bill Adra et al., editors. Advanced POWER virtualization on IBM e-server p5 servers. Introduction and basic configuration. IBM redbooks. IBM Corporation, San Jose, CA, USA, 2004. ISBN 0-7384-9081-4. xviii + 268 pp. LCCN

               QA76.9.V5 A378 2004. URL http://www.loc.gov/catdir/
               toc/fy0608/2005274479.html.

Adams:2006:CSH

[AA06] Keith Adams and Ole Agesen. A comparison of software and hardware techniques for x86 virtualization. Operating Systems Review, 40(5):2–13, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Auernhammer:2018:XEI

[AA18] F. Auernhammer and R. L. Arndt. XIVE: External interrupt virtualization for the cloud infrastructure. IBM Journal of Research and Development, 62(4–5):5:1–5:10, ???? 2018. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL https://ieeexplore.ieee.org/document/
8383690/.

Alubaidan:2024:PMI

[AA24] Haya A. Alubaidan and Sumayh S. Aljameel. A prediction model for improving virtual machine live migration performance in cloud computing using artificial intelligence techniques. International Journal of Computers and Applications, 46(12):1069–1087, 2024. CODEN IJ-CAFW. ISSN 1206-212X (print), 1925-7074 (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
1206212X.2024.2409988.

Alashaikh:2021:SUP

[AAAF21] Abdulaziz Alashaikh, Eisa Alanazi, and Ala Al-Fuqaha. A survey on the use of preferences for virtual machine placement in cloud data centers. ACM Computing Surveys, 54(5): 96:1–96:39, June 2021. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.acm.org/

doi/10.1145/3450517.

Alpern:2000:JAV

[AAB^{+}00] B. Alpern, C. R. Attanasio, J. J. Barton, M. G. Burke, P. Cheng, J.-D. Choi, A. Cocchi, S. J. Fink, D. Grove, M. Hind, S. F. Hummel, D. Lieber, V. Litvinov, M. F. Mergen, T. Ngo, J. R. Russell, V. Sarkar, M. J. Serrano, J. C. Shepherd, S. E. Smith, V. C. Sreedhar, H. Srinivasan, and J. Whaley. The Jalapeño virtual machine. IBM Systems Journal, 39(1):211–238, ???? 2000. CODEN IBMSA7. ISSN

0018-8670. URL http://www.research.ibm.com/journal/
sj/391/alpern.html.

Alpern:2005:JRV

[AAB^{+}05a] B. Alpern, S. Augart, S. M. Blackburn, M. Butrico, A. Cocchi, P. Cheng, J. Dolby, S. Fink, D. Grove, M. Hind, K. S. McKinley, M. Mergen, J. E. B. Moss, T. Ngo, V. Sarkar, and M. Trapp. The Jikes Research Virtual Machine project: Building an open-source research community. IBM Systems Journal, 44(2):399–417, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://

www.research.ibm.com/journal/sj/442/alpern.html;

http://www.research.ibm.com/journal/sj/442/alpern.
pdf;   http://www.research.ibm.com/journal/sj/442/
alpern.txt.

Alpern:2005:PVE

[AAB^{+}05b] Bowen Alpern, Joshua Auerbach, Vasanth Bala, Thomas Frauenhofer, Todd Mummert, and Michael Pigott. PDS: a virtual execution environment for software deployment. In ACM [ACM05d], pages 175–185. ISBN 1-59593-047-7. LCCN

QA76.9.V5 I575 2005. URL http://www.loc.gov/catdir/
toc/fy0611/2006530661.html. ACM order number 548059.

Armstrong:2005:AVC

[AAB^{+}05c] W. J. Armstrong, R. L. Arndt, D. C. Boutcher, R. G. Kovacs, D. Larson, K. A. Lucke, N. Nayar, and R. C. Swanberg. Advanced virtualization capabilities of POWER5 systems. IBM Journal of Research and Development, 49(4/5): 523–532, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.research.

ibm.com/journal/rd/494/armstrong.html.

Ayoubi:2017:RMC

[AAC^{+}17] Sara Ayoubi, Chadi Assi, Yiheng Chen, Tarek Khalifa, and Khaled Bashir Shaban. Restoration methods for cloud multicast virtual networks. Journal of Network and Computer Applications, 78(??):180–190, January 15, 2017. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302764.

Adeshiyan:2009:UVH

[AAF^{+}09] T. Adeshiyan, C. R. Attanasio, E. M. Farr, R. E. Harper, D. Pelleg, C. Schulz, L. F. Spainhower, P. Ta-Shma, and L. A. Tomek. Using virtualization for high availability and disaster recovery. IBM Journal of Research and Development, 53(4):??, ???? 2009. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.research.

ibm.com/journal/abstracts/rd/534/adeshiyan.html.

Ahmad:2003:ADP

[AAH^{+}03] I. Ahmad, J. M. Anderson, A. M. Holler, R. Kambo, and V. Makhija. An analysis of disk performance in VMware ESX Server virtual machines. In IEEE [IEE03], pages 65–76. ISBN 0-7803-8229-3. LCCN QA76.9.S88 W67 2003.

Al-Ayyoub:2016:VBC

[AAJD^{+}16] Mahmoud Al-Ayyoub, Yaser Jararweh, Ahmad Doulat, Haythem A. Bany Salameh, Ahmad Al Abed Al Aziz, Mohammad Alsmirat, and Abdallah A. Khreishah. Virtualizationbased Cognitive Radio Networks. The Journal of Systems and Software, 117(??):15–29, July 2016. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121216000479.

Aryania:2018:EAV

[AAK18] Azra Aryania, Hadi S. Aghdasi, and Leyli Mohammad Khanli. Energy-aware virtual machine consolidation algorithm based on ant colony system. Journal of Grid Computing, 16(3):477–491, September 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https://

link.springer.com/article/10.1007/s10723-018-9428-
4.

Aroca:2016:PEA

[AAM^{+}16] Jordi Arjona Aroca, Antonio Fernández Anta, Miguel A. Mosteiro, Christopher Thraves, and Lin Wang. Powerefficient assignment of virtual machines to physical machines. Future Generation Computer Systems, 54(??):82–94, January 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X15000072.

Alqahtani:2021:ECR

[AAMBE21] Fayez Alqahtani, Mohammed Al-Maitah, Khaldoun Besoul, and S. K. Elagan. Elastic computing resource virtualization method for a service-centric industrial Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print),

1872-7069 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1389128621000918.

Alharbe:2022:FGG

[AAR22] Nawaf Alharbe, Abeer Aljohani, and Mohamed Ali Rakrouki. A fuzzy grouping genetic algorithm for solving a real-world virtual machine placement problem in a healthcare-cloud. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/4/128.

Alharbi:2022:NSA

[AAT^{+}22] Abdullah Alharbi, Mohammed Aljebreen, Amr Tolba, Konstantinos A. Lizos, Saied Abd El-Atty, and Farid Shawki. A normalized slicing-assigned virtualization method for 6gbased wireless communication systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):134:1–134:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/10.1145/3546077.

Antonescu:2016:SSB

[AB16] Alexandru-Florian Antonescu and Torsten Braun. Simulation of SLA-based VM-scaling algorithms for clouddistributed applications. Future Generation Computer Systems, 54(??):260–273, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X15000321.

Axnix:2015:IZF

[ABB^{+}15] C. Axnix, G. Bayer, H. Bohm, J. von Buttlar, M. S. Farrell, L. C. Heller, J. P. Kubala, S. E. Lederer, R. Mansell, A. Nunez Mencias, and S. Usenbinz. IBM z13 firmware innovations for simultaneous multithreading and I/O virtualization. IBM Journal of Research and Development, 59(??): 11:1–11:11, ???? 2015. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Abeni:2019:HSR

[ABB19a] Luca Abeni, Alessandro Biondi, and Enrico Bini. Hierarchical scheduling of real-time tasks over Linux-based virtual machines. The Journal of Systems and Software, 149(??):234– 249, March 2019. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016412121830270X.

Atzori:2019:SCI

[ABB^{+}19b] L. Atzori, J. L. Bellido, R. Bolla, G. Genovese, A. Iera, A. Jara, C. Lombardo, and G. Morabito. SDN&NFV contribution to IoT objects virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 149(??):200–212, February 11, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-

7069 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1389128618312933.

Armbruster:2007:RTJ

[ABC^{+}07] Austin Armbruster, Jason Baker, Antonio Cunei, Chapman Flack, David Holmes, Filip Pizlo, Edward Pla, Marek Prochazka, and Jan Vitek. A real-time Java virtual machine with applications in avionics. ACM Transactions on Embedded Computing Systems, 7(1):5:1–5:49, December 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Adair:1966:VMS

[ABCC66] R. J. Adair, R. U. Bayles, L. W. Comeau, and R. J. Creasy. A virtual machine system for the 360/40. Technical Report 320-2007, International Business Machines (IBM), Data Processing Division, 1966.

Aharon:1991:VIR

[ABDD^{+}91] A. Aharon, A. Bar-David, B. Dorfman, E. Gofman, M. Leibowitz, and V. Schwartzburd. Verification of the IBM RISC System/6000 by a dynamic biased pseudo-random test program generator. IBM Systems Journal, 30(4):527–538, 1991. CODEN IBMSA7. ISSN 0018-8670.

Arya:2014:TRG

[ABG14] Kapil Arya, Yury Baskakov, and Alex Garthwaite. Tesseract: reconciling guest I/O and hypervisor swapping in a VM. ACM SIGPLAN Notices, 49(7):15–28, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abramson:1980:WGL

[Abr80] Harvey Abramson. Why is a goto like a dynamic vector in the BCPL-Slim computing system. Technical Report TR-80-09, Department of Computer Science, University of British Columbia, November 1980. Mon, 21 Jul 1997 19:29:13 GMT.

Abramsky:1982:SMV

[Abr82] S. Abramsky. SECD-M: a virtual machine for applicative multiprogramming. Technical Report QMW-DCS-1982-322; QMW-DCS-1982-339, Department of Computer Science, Queen Mary College, November 1982. (LA has).

Anderson:2012:MAN

[ABV12] Paul Anderson, Shahriar Bijani, and Alexandros Vichos. Multi-agent negotiation of virtual machine migration using the lightweight coordination calculus. Lecture Notes in Computer Science, 7327:124–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-30947-
2_16/.

Ambriola:1995:DVM

[AC95] Vincenzo Ambriola and Giovanni A. Cignoni. A distributed virtual machine to support software process. ACM SIGSOFT Software Engineering Notes, 20(1):85–89, January 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

AzanonEsteire:1998:JST

[AC98] Oscar Azañón Esteire and Juan Manual Cueva Lovelle. J — set of tools for native code generation for the Java Virtual Machine. ACM SIGPLAN Notices, 33(3):73–79, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anjo:2016:DML

[AC16] Ivo Anjo and João Cachopo. Design of a method-level speculation framework for boosting irregular JVM applications. Journal of Parallel and Distributed Computing, 87(??):13–25, January 2016. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731515001720.

Ayoubi:2016:TPB

[ACA16] Sara Ayoubi, Yiheng Chen, and Chadi Assi. Towards promoting backup-sharing in survivable virtual network design. IEEE/ACM Transactions on Networking, 24(5):3218–3231, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Anglano:2018:PFT

[ACG18] Cosimo Anglano, Massimo Canonico, and Marco Guazzone. Prometheus: a flexible toolkit for the experimentation with virtualized infrastructures. Concurrency and Computation: Practice and Experience, 30(11):??, June 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-

0634 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.4400.

Ancilotti:1972:VIO

[ACL72] R. Ancilotti, R. Cavina, and N. Lijtmaer. Virtual inputoutput in a virtual environment. In ????, editor, ACM AICA International Computer Symposium Proceedings, Venice, Italy, April 12–14, 1972, pages 302–312. ACM Press, New York, NY 10036, USA, 1972.

ACM:1975:PFS

[ACM75] ACM, editor. Proceedings of the Fifth Symposium on Operating System Principles, November 19–21, 1975, The University of Texas at Austin, Austin, Texas, USA, volume 9(5) of Operating Systems Review. ACM Press, New York, NY 10036, USA, 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

ACM:1981:ASC

[ACM81] ACM, editor. ACM SIGMETRICS Conference on Measurement and Modeling of Computer Systems, University of Nevada, Las Vegas, Nevada, USA, September 14–16, 1981, Proceedings, volume 10(3) of Performance Evaluation Review. ACM Press, New York, NY 10036, USA, Fall 1981. ISBN ???? LCCN ????

ACM:1989:PSN

[ACM89] ACM, editor. Proceedings, Supercomputing ’89: November 13–17, 1989, Reno, Nevada. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

ACM:1990:PAC

[ACM90] ACM, editor. Proceedings of the 1990 ACM Conference on LISP and Functional Programming: papers presented at the conference, Nice, France, June 27–29, 1990. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-368-X. LCCN QA 76.73 L23 A24 1990. ACM order no. 552900.

ACM:1996:SCP

[ACM96] ACM, editor. Supercomputing ’96 Conference Proceedings: November 17–22, Pittsburgh, PA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-89791-854-1. LCCN ???? URL http://www.

supercomp.org/sc96/proceedings/. ACM Order Num-

ber: 415962, IEEE Computer Society Press Order Number: RS00126.

ACM:1997:PTA

[ACM97] ACM, editor. Proceedings of the TRI-Ada’97 Conference, November 9–13, 1997, St. Louis, MO. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-981-5. LCCN QA 76.73 A35 T75 1997. Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

ACM:1998:AWJ

[ACM98] ACM, editor. ACM 1998 Workshop on Java for High-Performance Network Computing, Concurrency: Practice and Experience. ACM Press, New York, NY 10036, USA, 1998. CODEN CPEXEI. ISSN 1040-3108. LCCN ???? URL

http://www.cs.ucsb.edu/conferences/java98/program.

html. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

ACM:1999:PPA

[ACM99] ACM, editor. POPL ’99. Proceedings of the 26th ACM SIGPLAN-SIGACT on Principles of programming languages, January 20–22, 1999, San Antonio, TX. ACM Press, New York, NY 10036, USA, 1999. ISBN 1-58113-095-3.

LCCN ???? URL http://www.acm.org/pubs/contents/
proceedings/plan/292540/index.html.

ACM:2000:CPI

[ACM00] ACM, editor. Conference proceedings of the 2000 International Conference on Supercomputing: Santa Fe, New Mexico, May 8–11, 2000. ACM Press, New York, NY 10036, USA, 2000. ISBN 1-58113-270-0. LCCN QA76.88 .I573 2000.

URL http://www.acm.org/pubs/contents/proceedings/
supercomputing/335231.

ACM:2001:ASS

[ACM01a] ACM, editor. ACM SIGPLAN–SIGSOFT workshop on Program analysis for software tools and engineering: June 18–19, 2001, Snowbird, Utah, USA: PASTE’01. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-413-4. LCCN QA76.758. Supplement to ACM SIGPLAN Notices.

ACM:2001:PAJ

[ACM01b] ACM, editor. Proceedings of the ACM 2001 Java Grande/ ISCOPE Conference: Palo Alto, Calif., June 2–4, 2001. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-359-6. LCCN QA76.9.O35 A26 2001.

ACM:2003:SII

[ACM03a] ACM, editor. SC2003: Igniting Innovation. Phoenix, AZ, November 15–21, 2003. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 1-58113-695-1. LCCN ????

ACM:2003:SPA

[ACM03b] ACM, editor. SOSP ’03: proceedings of the 19th ACM Symposium on Operating Systems Principles: the Sagamore, Bolton Landing, Lake George, New York, USA, October 19– 22, 2003, volume 37(5) of Operating systems review. ACM Press, New York, NY 10036, USA, December 2003. ISBN 1-58113-757-5. ISSN 0163-5980 (print), 1943-586X (elec-

tronic). URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://uclibs.org/PID/34720. ACM or-

der number 534030.

ACM:2004:PWA

[ACM04a] ACM, editor. Proceedings of the 11th workshop on ACM SIGOPS European workshop: beyond the PC 2004, Leuven, Belgium, September 19–22, 2004. ACM Press, New York, NY 10036, USA, 2004. ISBN ???? LCCN ????

ACM:2004:SHP

[ACM04b] ACM, editor. SC 2004: High Performance Computing, Networking and Storage: Bridging communities: Proceedings of the IEEE/ACM Supercomputing 2004 Conference, Pittsburgh, PA, November 6–12, 2004. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2153-3. LCCN QA76.88. IEEE order number E2153. ACM order number 415043.

ACM:2005:APS

[ACM05a] ACM, editor. AADEBUG 2005: proceedings of the Sixth International Symposium on Automated and Analysis-Driven Debugging: Monterey, California, USA, September 19–21, 2005. ACM Press, New York, NY 10036, USA, 2005. ISBN 1-59593-050-7. LCCN QA76.9.D43 I58 2005.

ACM:2005:MPI

[ACM05b] ACM, editor. MGC’05: Proceedings of the 3rd International Workshop on Middleware for Grid Computing, Grenoble, France, November 28–December 02, 2005. ACM Press, New York, NY 10036, USA, 2005. ISBN 1-59593-269-0. LCCN ????

ACM:2005:PAI

[ACM05c] ACM, editor. Proceedings of the 2005 ACM/IEEE conference on Supercomputing 2005, Seattle, WA, November 12–18 2005. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 1-59593-061-2. LCCN ????

ACM:2005:PFA

[ACM05d] ACM, editor. Proceedings of the First ACM/USENIX International Conference on Virtual Execution Environments: VEE ’05: June 11–12, 2005, Chicago, Illinois, USA. ACM Press, New York, NY 10036, USA, 2005. ISBN 1-59593-047-7. LCCN QA76.9.V5 I575 2005. URL http://www.loc.gov/

catdir/toc/fy0611/2006530661.html. ACM order number

548059.

ACM:2006:AAI

[ACM06a] ACM, editor. Annual ACM IEEE Design Automation Conference, Proceedings of the 43rd Annual Conference on Design Automation, San Francisco, CA, USA. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-381-6. LCCN ????

ACM:2006:PPI

[ACM06b] ACM, editor. PACT’06: Proceedings of the 15th International Conference on Parallel Architectures and Compilation Techniques 2006, Seattle, Washington, USA, September 16– 20, 2006. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-264-X. LCCN ????

ACM:2006:PIW

[ACM06c] ACM, editor. Proceedings of the 1st international workshop on Wireless network testbeds, experimental evaluation and characterization 2006, Los Angeles, CA, USA, September 29, 2006. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-540-0. LCCN ????

ACM:2006:PST

[ACM06d] ACM, editor. Proceedings of the 37th SIGCSE Technical Symposium on Computer Science Education 2006, Houston, Texas, USA, March 03–05, 2006, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2006. CODEN SIGSD3. ISBN ???? ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA76.27. ACM order number 457060.

ACM:2006:PCC

[ACM06e] ACM, editor. Proceedings of the 3rd conference on Computing Frontiers, May 3–5, 2006, Ischia, Italy. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-302-6. LCCN ???? ACM order number 104060.

ACM:2006:VPS

[ACM06f] ACM, editor. VEE 2006: proceedings of the Second International Conference on Virtual Execution Environments, June 14-16, 2006, Ottawa, Ontario, Canada. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-332-6 (??invalid ISBN??). LCCN QA76.9.V4.

Argade:1994:TMR

[ACT94] Pramod V. Argade, David K. Charles, and Craig Taylor. A technique for monitoring run-time dynamics of an operating system and a microprocessor executing user applications. ACM SIGPLAN Notices, 29(11):122–131, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/proceedings/asplos/
195473/p122-argade/.

Armstrong:2011:PIC

[AD11] Django Armstrong and Karim Djemame. Performance issues in clouds: an evaluation of virtual image propagation and I/O paravirtualization. The Computer Journal, 54(6):836–849, June 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/54/6/836.full.pdf+html.

A:2018:AML

[AD18a] Ajay Kumara M. A. and Jaidhar C. D. Automated multi-level malware detection system based on reconstructed semantic view of executables using machine learning techniques at VMM. Future Generation Computer Systems, 79 (part 1)(??): 431–446, 2018. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167739X17311809.

Anwar:2018:ESS

[AD18b] Nazia Anwar and Huifang Deng. Elastic scheduling of scientific workflows under deadline constraints in cloud computing environments. Future Internet, 10(1):5, January 07, 2018. CODEN ???? ISSN 1999-5903. URL https://www.mdpi.

com/1999-5903/10/1/5.

Agarwal:2019:SVM

[AD19] Amit Agarwal and Ta Nguyen Binh Duong. Secure virtual machine placement in cloud data centers. Future Generation Computer Systems, 100(??):210–222, November 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X18326116.

Aldossary:2019:EAC

[ADA^{+}19] Mohammad Aldossary, Karim Djemame, Ibrahim Alzamil, Alexandros Kostopoulos, Antonis Dimakis, and Eleni Agiatzidou. Energy-aware cost prediction and pricing of virtual machines in cloud computing environments. Future Generation Computer Systems, 93(??):442–459, April 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167739X18310288.

Ackerman:1992:SIE

[ADG^{+}92] D. F. Ackerman, M. H. Decker, J. J. Gosselin, K. M. Lasko, M. P. Mullen, R. E. Rosa, E. V. Valera, and B. Wile. Simulation of IBM Enterprise System/9000 models 820 and 900. IBM Journal of Research and Development, 36(4):751–764, July 1992. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Agesen:1998:GCL

[ADM98] Ole Agesen, David Detlefs, and J. Eliot B. Moss. Garbage collection and local variable type-precision and liveness in Java Virtual Machines. ACM SIGPLAN Notices, 33(5):269– 279, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p269-agesen/; http://www.
cs.virginia.edu/pldi98/program.html.

Alves:2018:VST

[ADWM18] Thiago Alves, Rishabh Das, Aaron Werth, and Thomas Morris. Virtualization of SCADA testbeds for cybersecurity research: a modular approach. Computers & Security, 77(??):531–546, August 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818304905.

Aoki:2001:SVM

[AE01] Takashi Aoki and Takeshi Eto. On the software virtual machine for the real hardware stack machine. In USENIX [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.usenix.org/publications/
library/proceedings/jvm01/aoki.html.

Ali:2023:ETN

[AEA^{+}23] Shrouk A. Ali, Shaimaa Ahmed Elsaid, Abdelhamied A. Ateya, Mohammed ElAffendi, , and Ahmed A. Abd El-Latif. Enabling technologies for next-generation smart cities: a comprehensive review and research directions. Future Internet, 15(12):398, December 09, 2023. CODEN ???? ISSN

1999-5903. URL https://www.mdpi.com/1999-5903/15/
12/398.

Asvija:2019:SHA

[AEB19] B. Asvija, R. Eswari, and M. B. Bijoy. Security in hardware assisted virtualization for cloud computing — state of the art issues and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 151(??):68–92, March 14, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128618302998.

Adams:2014:HVM

[AEM^{+}14] Keith Adams, Jason Evans, Bertrand Maher, Guilherme Ottoni, Andrew Paroski, Brett Simmers, Edwin Smith, and Owen Yamauchi. The HipHop Virtual Machine. ACM SIG-PLAN Notices, 49(10):777–790, October 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abd-El-Malek:2012:FSV

[AEMWC^{+}12] Michael Abd-El-Malek, Matthew Wachs, James Cipar, Karan Sanghi, Gregory R. Ganger, Garth A. Gibson, and Michael K. Reiter. File system virtual appliances: Portable file system implementations. ACM Transactions on Storage, 8(3):9:1– 9:??, September 2012. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Abdelaziz:2017:SDW

[AFG^{+}17] Ahmed Abdelaziz, Ang Tan Fong, Abdullah Gani, Suleman Khan, Faiz Alotaibi, and Muhammad Khurram Khan. On software-defined wireless network (SDWN) network virtualization: Challenges and open issues. The Computer Journal, 60(10):1510–1519, October 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL https://academic.oup.com/comjnl/
article/60/10/1510/4321712.

Aridor:2001:DIV

[AFT01] Yariv Aridor, Michael Factor, and Avi Teperman. A distributed implementation of a virtual machine for Java. Concurrency and Computation: Practice and Experience, 13(3):

221–244, March 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/78003113/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=78003113&PLACEBO=IE.pdf.

Alshathri:2018:SLM

[AGC18] Samah Alshathri, Bogdan Ghita, and Nathan Clarke. Sharing with live migration energy optimization scheduler for cloud computing data centers. Future Internet, 10(9):86, September 06, 2018. CODEN ???? ISSN 1999-5903. URL https://www.

mdpi.com/1999-5903/10/9/86.

Ahmad:2015:VMM

[AGH^{+}15a] Raja Wasim Ahmad, Abdullah Gani, Siti Hafizah Ab. Hamid, Muhammad Shiraz, Feng Xia, and Sajjad A. Madani. Virtual machine migration in cloud data centers: a review, taxonomy, and open research issues. The Journal of Supercomputing, 71 (7):2473–2515, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://link.

springer.com/article/10.1007/s11227-015-1400-5.

Ahmad:2015:SVM

[AGH^{+}15b] Raja Wasim Ahmad, Abdullah Gani, Siti Hafizah Ab. Hamid, Muhammad Shiraz, Abdullah Yousafzai, and Feng Xia. A survey on virtual machine migration and server consolidation frameworks for cloud data centers. Journal of Network and Computer Applications, 52(??):11–25, June 2015. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000284.

Amit:2016:BMP

[AGH^{+}16] Nadav Amit, Abel Gordon, Nadav Har’El, Muli Ben-Yehuda, Alex Landau, Assaf Schuster, and Dan Tsafrir. Bare-metal performance for virtual machines with exitless interrupts. Communications of the ACM, 59(1):108–116, January 2016. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://cacm.acm.org/magazines/2015/1/
195734/fulltext.

Averbuch:1994:PES

[AGIS94] A. Averbuch, E. Gabber, S. Itzikowitz, and B. Shoham. On the parallel elliptic single/multigrid solutions about aligned and nonaligned bodies using the Virtual Machine for Multiprocessors. Scientific Programming, 3(1):13–32, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Abe:2016:UVM

[AGJS16] Yoshihisa Abe, Roxana Geambasu, Kaustubh Joshi, and Mahadev Satyanarayanan. Urgent virtual machine eviction with enlightened post-copy. ACM SIGPLAN Notices, 51(7):51– 64, July 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aral:1991:PCS

[AGLM91] Ziya Aral, Ilya Gertner, Alan Langerman, and Dave Mitchell. Process control structures for multiprocessor. In Milutinovic and Shriver [MS91b], pages 49–58. ISBN 0-8186-2119-2 (vol. 1), 0-8186-2120-6 (vol. 2), 0-8186-2010-2 (vol. 3), 0-8186-2122-2 (vol. 4). LCCN QA76.9.S88 H38 1991; QA76.9.S88; QA76.9.S88 H375 1991; Internet; TA168 .H37a 1991. Four volumes. IEEE catalog number 91TH0350-9.

Aagren:1999:TCC

[Ågr99] Ola Ågren. Teaching computer concepts using virtual machines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):84–85, June 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Agesen:2010:EXV

[AGSS10] Ole Agesen, Alex Garthwaite, Jeffrey Sheldon, and Pratap Subrahmanyam. The evolution of an x86 virtual machine monitor. Operating Systems Review, 44(4):3–18, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Auroux:1968:CMV

[AH68] A. Auroux and C. Han. Le concept de machines virtuelles. (French) [The concept of virtual machines]. Revue Française d’Informatique et de Recherche Opérationelle, 2e anné, 15: 45–51, 1968.

Aguiar:2012:CTF

[AH12] Alexandra Aguiar and Fabiano Hessel. Current techniques and future trends in embedded system’s virtualization. Software—Practice and Experience, 42(7):917–944, July 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alkhalaf:2024:EDD

[AH24] Asma Alkhalaf and Farookh Khadeer Hussain. EleVMate — a data-driven approach for ‘on-the-fly’ horizontal small datacentre scalability and VM starvation. Future Generation Computer Systems, 159(??):91–101, October 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X24001651.

Aigner:2015:AJE

[AHK^{+}15] Martin Aigner, Thomas Hütter, Christoph M. Kirsch, Alexander Miller, Hannes Payer, and Mario Preishuber. ACDC-JS: explorative benchmarking of JavaScript memory management. ACM SIGPLAN Notices, 50(2):67–78, February 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Al-Hamid:2023:NSB

[AHKC23] Duaa Zuhair Al-Hamid, Pejman A. Karegar, , and Peter Han Joo Chong. A novel SDWSN-based testbed for IoT smart applications. Future Internet, 15(9):291, August 28, 2023. CODEN ???? ISSN 1999-5903. URL https://www.

mdpi.com/1999-5903/15/9/291.

Ahmadi:2022:CIB

[AHRR22a] Javad Ahmadi, Abolfazl Toroghi Haghighat, Amir Masoud Rahmani, and Reza Ravanmehr. Confidence interval-based overload avoidance algorithm for virtual machine placement. Software—Practice and Experience, 52(10):2288–2311, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ahmadi:2022:FAV

[AHRR22b] Javad Ahmadi, Abolfazl Toroghi Haghighat, Amir Masoud Rahmani, and Reza Ravanmehr. A flexible approach for virtual machine selection in cloud data centers with AHP. Software—Practice and Experience, 52(5):1216–1241, May 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aslan:2024:FRA

[AIAR^{+}24] Fatih Aslan, George Iosifidis, Jose A. Ayala-Romero, Andres Garcia-Saavedra, and Xavier Costa-Perez. Fair resource allocation in virtualized O-RAN platforms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 8(1):17:1–17:??, March 2024. CODEN ????

ISSN 2476-1249. URL https://dl.acm.org/doi/10.1145/
3639043.

Abbasi:2018:VMA

[AJ18] Aaqif Afzaal Abbasi and Hai Jin. v-Mapper: An applicationaware resource consolidation scheme for cloud data centers. Future Internet, 10(9):90, September 15, 2018. CODEN ????

ISSN 1999-5903. URL https://www.mdpi.com/1999-5903/
10/9/90.

Aghasi:2023:DAM

[AJBJ23] Ali Aghasi, Kamal Jamshidi, Ali Bohlooli, and Bahman Javadi. A decentralized adaptation of model-free Q-learning for thermal-aware energy-efficient virtual machine placement in cloud data centers. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128623000695.

Anderson:2009:XWL

[AJD09] Benjamin R. Anderson, Amy K. Joines, and Thomas E. Daniels. Xen worlds: leveraging virtualization in distance education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):293–297, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Ahn:2012:RHA

[AJH12] Jeongseob Ahn, Seongwook Jin, and Jaehyuk Huh. Revisiting hardware-assisted page walks for virtualized systems. ACM SIGARCH Computer Architecture News, 40(3):476–487, June 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ISCA ’12 conference proceedings.

Abramson:2006:IVT

[AJM^{+}06] Darren Abramson, Jeff Jackson, Sridhar Muthrasanallur, Gil Neiger, Greg Regnier, Rajesh Sankaran, Ioannis Schoinas, Rich Uhlig, Balaji Vembu, and John Weigert. Intel Virtualization Technology for directed I/O. Intel Technology Journal, 10(3):179–192, August 10, 2006. ISSN 1535-

766X. URL http://developer.intel.com/technology/
itj/2006/v10i3/2-io/1-abstract.htm.

Avritzer:2025:SAD

[AJM^{+}25] Alberto Avritzer, Andrea Janes, Andrea Marin, Catia Trubiani, Andre van Hoorn, Matteo Camilli, Daniel S. Menasché, and André B. Bondi. Software aging detection and rejuvenation assessment in heterogeneous virtual networks. IEEE Transactions on Emerging Topics in Computing, 13(2):299– 313, 2025. ISSN 2168-6750 (print), 2376-4562 (electronic).

Apostolopoulos:2021:RAV

[AKCP21] Theodoros Apostolopoulos, Vasilios Katos, Kim-Kwang Raymond Choo, and Constantinos Patsakis. Resurrecting anti-virtualization and anti-debugging: Unhooking your hooks. Future Generation Computer Systems, 116(??):393– 405, March 2021. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167739X20330284.

Adamski:2007:SPE

[AKK^{+}07] Marcin Adamski, Michal Kulczewski, Krzysztof Kurowski, Jarek Nabrzyski, and Alastair Hume. Security and performance enhancements to OGSA-DAI for Grid data virtualization. Concurrency and Computation: Practice and Experience, 19(16):2171–2182, November 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Adams:2005:CMC

[AL05] Joel C. Adams and W. David Laverell. Configuring a multicourse lab for system-level projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (1):525–529, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alfonseca:1991:AAA

[Alf91] M. Alfonseca. Advanced applications of APL: logic programming, neural networks, and hypertext. IBM Systems Journal, 30(4):543–553, 1991. CODEN IBMSA7. ISSN 0018-8670.

Asrigo:2006:UVB

[ALL06] Kurniadi Asrigo, Lionel Litty, and David Lie. Using VMMbased sensors to monitor honeypots. In ACM [ACM06f], pages 13–23. ISBN 1-59593-332-6 (??invalid ISBN??). LCCN QA76.9.V4.

Alnaim:2022:MPT

[Aln22] Abdulrahman K. Alnaim. Misuse patterns from the threat of modification of non-control data in network function virtualization. Future Internet, 14(7):201, June 30, 2022. CODEN

???? ISSN 1999-5903. URL https://www.mdpi.com/1999-
5903/14/7/201.

Akyildiz:2015:WSD

[ALW15] Ian F. Akyildiz, Shih-Chun Lin, and Pu Wang. Wireless software-defined networks (W-SDNs) and network function virtualization (NFV) for 5G cellular systems: an overview and qualitative evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1)(??):66–79, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-

7069 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1389128615003862.

Agrawal:2016:EIU

[AM16] Gaurav Agrawal and Deep Medhi. Embedding IP unique shortest path topology on a wavelength-routed network: normal and survivable design. IEEE/ACM Transactions on Networking, 24(2):1109–1124, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Azmandian:2011:VMM

[AMA^{+}11] Fatemeh Azmandian, Micha Moffie, Malak Alshawabkeh, Jennifer Dy, Javed Aslam, and David Kaeli. Virtual machine monitor-based lightweight intrusion detection. Operating Systems Review, 45(2):38–53, July 2011. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Araujo:2014:SAE

[AMA^{+}14] Jean Araujo, Rubens Matos, Vandi Alves, Paulo Maciel, F. Vieira de Souza, Rivalino Matias, Jr., and Kishor S. Trivedi. Software aging in the Eucalyptus cloud computing infrastructure: Characterization and rejuvenation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(1):11:1–11:??, January 2014. CODEN ???? ISSN 1550-4832.

Ahmadian:2018:ECH

[AMA18] Saba Ahmadian, Onur Mutlu, and Hossein Asadi. ECI-Cache: a high-endurance and cost-efficient I/O caching scheme for virtualized platforms. ACM SIGMETRICS Performance Evaluation Review, 46(1):73, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arroba:2017:DVF

[AMAB17] Patricia Arroba, José M. Moya, José L. Ayala, and Rajkumar Buyya. Dynamic Voltage and Frequency Scaling-aware dynamic consolidation of virtual machines for energy efficient cloud data centers. Concurrency and Computation: Practice and Experience, 29(10), May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Auler:2017:HIP

[AMB^{+}17] Rafael Auler, Carlos Eduardo Millani, Alexandre Brisighello, Alisson Linhares, and Edson Borin. Handling IoT platform heterogeneity with COISA, a compact OpenISA virtual platform. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ament:2013:ATG

[Ame13] John Ament. Arquillian Testing Guide: get familiarized with the Arquillian framework and its capabilities to carry out integration and functional testing on a Java virtual machine.

Packt Publishing, Birmingham, UK, 2013. ISBN 1-78216-070-1. v + 224 pp. LCCN QA76.73.J3. URL http://

proquest.tech.safaribooksonline.de/9781782160700.

Awad:2016:SSZ

[AMH^{+}16] Amro Awad, Pratyusa Manadhata, Stuart Haber, Yan Solihin, and William Horne. Silent shredder: Zero-cost shredding for secure non-volatile main memory controllers. ACM SIGPLAN Notices, 51(4):263–276, April 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ahmed:2019:ILT

[AMIA19] Jawad Ahmed, Aqsa Malik, Muhammad U. Ilyas, and Jalal S. Alowibdi. Instance launch-time analysis of OpenStack virtualization technologies with control plane network errors. Computing: Archiv für Informatik und Numerik, 101(8):989– 1014, August 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Azevedo:2000:AAJ

[ANH00] Ana Azevedo, Alex Nicolau, and Joe Hummel. An annotation-aware Java virtual machine implementation. Concurrency: Practice and Experience, 12(6):423–444, May 2000. CODEN CPEXEI. ISSN 1040-3108. URL http://www3.

interscience.wiley.com/cgi-bin/abstract/72515731/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=72515731&PLACEBO=IE.pdf.

Anonymous:1975:VM

[Ano75] Anonymous. Virtual machines. In KB11-C Processor Manual (PDP-11/70), pages I.2.4–I.2.10 (24–30). Digital Equipment Corporation, Maynard, MA, USA, 1975. LCCN

???? URL http://bitsavers.trailing-edge.com/pdf/
dec/pdp11/1170/EK-KB11C-TM-001_1170procMan.pdf.

Anonymous:1993:NCS

[Ano93] Anonymous, editor. 16th National Computer Security Conference: September 20–23, 1993, Baltimore Convention Center, Baltimore, Maryland: proceedings: information systems security, user choices. National Institute of Standards and Technology/National Computer Security Center, Gaithersburg, MD, USA, 1993.

Anonymous:1994:SAS

[Ano94] Anonymous. Software architecture standard for simulation virtual machine: version 2.0. Technical Report NASA CR-188291, National Aeronautics and Space Administration, Washington, DC, USA, 1994. ???? pp.

Anonymous:1996:TWJb

[Ano96] Anonymous. Towards Web/Java-based high performance distributed computing — an evolving virtual machine. IEEE International Symposium on High Performance Distributed Computing, Proceedings, pages 308–317, 1996. CODEN PID-CFB. ISSN 1082-8907. IEEE catalog number 96TB100069.

Anonymous:1997:BRJe

[Ano97a] Anonymous. Book review: Java Virtual Machine: By Jon Meyer and Troy Downing. O’Reilly, Sebastopol, CA. (1997). 426 pages. $32.95 (diskette included). Computers and Mathematics with Applications, 34(10):135, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122197901899.

Anonymous:1997:BFJ

[Ano97b] Anonymous. Bug found in Java Virtual Machine. Network Security, 1997(3):2, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1353485897830334.

Anonymous:1997:IJV

[Ano97c] Anonymous. Inside the Java Virtual Machine. UNIX Review, 15(1):31, 32, 34–36, 38–39, January 1997. CODEN UNRED5. ISSN 0742-3136. Adapted, with permission, from [LY97b].

Anonymous:1997:JVM

[Ano97d] Anonymous. Java Virtual Machine unleashed. SAMS Publishing, Indianapolis, IN, USA, 1997. ISBN 1-57521-247-1 (paperback). ???? pp. LCCN 9710 BOOK NOT YET IN LC.

Anonymous:1999:MVM

[Ano99a] Anonymous. Microsoft virtual machine vulnerability. Network Security, 1999(12):2, December 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1353485899903240.

Anonymous:1999:PII

[Ano99b] Anonymous, editor. Proceedings of the IASTED International Conference on Internet and Multimedia Systems and Applications (IMSA). Nassau, Bahamas, October 1999. Acta Press, Anaheim, CA, USA, 1999. ISBN ???? LCCN ????

Anonymous:2000:AJV

[Ano00] Anonymous. Announcement: Java Virtual Machine Research and Technology Symposium (JVM ’01). ;login: the USENIX Association newsletter, 25(3):??, June 2000. CO-DEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/events/jvm01.

Anonymous:2001:CRJ

[Ano01a] Anonymous. Conference report: 1st Java VM Virtual Machine Research and Technology Symposium (JVM ’01). ;login: the USENIX Association newsletter, 26(5):??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:2001:PJV

[Ano01b] Anonymous, editor. Proceedings of the Java Virtual Machine Research and Technology Symposium (JVM ’01) April 23–24, 2001, Monterey, California, USA. USENIX, San Francisco, CA, USA, 2001. ISBN 1-880446-11-1. LCCN ????

Anonymous:2002:CRJ

[Ano02] Anonymous. Conference report: 2nd Java Virtual Machine Research and Technology Symposium. ;login: the USENIX Association newsletter, 27(5):??, October 2002. CODEN LOGNEM. ISSN 1044-6397. URL http:

//www.usenix.org/publications/library/proceedings/
javavm02/JavaVM2002.pdf.

Anonymous:2003:PJU

[Ano03a] Anonymous. Products: JetBrains upgrades IntelliJ Java IDE; Catalyst’s USB analyzer supports device emulation; VMware releases Enterprise Server VM software; Motorola offers free soft modem reference design; RealNetworks releases source for Helix DNA Server; Packeteer accelerates intranet and Internet applications. Computer, 36(3):80–81, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-

0814 (electronic). URL http://csdl.computer.org/dl/
mags/co/2003/03/r3080.htm;   http://csdl.computer.
org/dl/mags/co/2003/03/r3080.pdf.

Anonymous:2003:PVF

[Ano03b] Anonymous. Products: VMware’s fourth-generation desktop virtualization software; automated design reviews with Reviewer for Rose; CodeWeavers debues CrossOver Office; Corel Smart Graphics Studio now available; Iron-Grid’s Java performance tool; Infragistics releases NetAdvantage component collection. Computer, 36(6):98–99, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-

0814 (electronic). URL http://csdl.computer.org/dl/
mags/co/2003/06/r6098.htm;   http://csdl.computer.
org/dl/mags/co/2003/06/r6098.pdf.

Anonymous:2004:CRV

[Ano04a] Anonymous. Conference report: 3rd Virtual Machine Research and Technology Symposium (VM ’04). ;login: the USENIX Association newsletter, 29(4):??, August 2004. CODEN LOGNEM. ISSN 1044-6397. URL

http://www.usenix.org/publications/login/2004-08/
openpdfs/vm04reports.pdf.

Anonymous:2004:PTV

[Ano04b] Anonymous, editor. Proceedings of the Third Virtual Machine Research and Technology Symposium (VM ’04): May 6–7, 2004, San Jose, California, USA, Conference proceedings / USENIX Association. USENIX, San Francisco, CA, USA, 2004. ISBN 1-931971-20-X. LCCN QA76.9.V5 V565 2004.

Anonymous:2005:NPV

[Ano05] Anonymous. New products: Virtual iron (VFe). Linux Journal, 139:18, December 2005. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2006:PGI

[Ano06a] Anonymous, editor. Proceedings of Gelato ICE: Itanium Conference and Expo: Spotlighting Linux on Itanium-based Platforms, October 1-4, 2006, Biopolis, Singapore. ????, ????, 2006. ISBN ???? LCCN ???? URL http:/

/www.ice.gelato.org/;   http://www.ice.gelato.org/
about/oct06_presentations.php.

Anonymous:2006:TR

[Ano06b] Anonymous. Technical reviewers. Intel Technology Journal, 10(3):vii, August 10, 2006. ISSN 1535-766X.

Anonymous:2007:VPS

[Ano07] Anonymous. VMware patches security flaws. Network Security, 2007(10):2, October 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1353485807700892.

Anonymous:2010:NDS

[Ano10] Anonymous, editor. 17th Annual Network and Distributed System Symposium, NDSS ’10, The Dana on Misson Bay, San Diego, California. February 28–March 3, 2010. Internet Society, Reston, VA, USA, 2010. ISBN 1-891562-29-0, 1-891562-30-4. LCCN ???? URL http://www.isoc.org/

isoc/conferences/ndss/10/proceedings.shtml.

Anonymous:2014:ASS

[Ano14a] Anonymous. AMD64 Simics Simulator version 1. Web site.,

2014. URL http://www.windriver.com/simics/.

Anonymous:2014:BIE

[Ano14b] Anonymous. Bochs IA-32 emulator. Web site., 2014. URL

http://bochs.sourceforge.net/.

Anonymous:2014:LVA

[Ano14c] Anonymous. The libvirt virtualization API. Web site., 2014.

URL http://libvirt.org/.

Anonymous:2014:O

[Ano14d] Anonymous. oVirt. Web site., 2014. URL http://ovirt.

org/.

Anonymous:2015:CXB

[Ano15] Anonymous. Critical Xen bug in PV memory virtualization code (XSA 148). Web bug report, October 29,

2015. URL https://github.com/QubesOS/qubes-secpack/
blob/master/QSBs/qsb-022-2015.txt. The report notes

about this bug that allows memory pages to leak between Xen virtual machines on the same physical host: “… the bug is a very critical one. Probably the worst we have seen affecting the Xen hypervisor, ever. Sadly. … it is really shocking that such a bug has been lurking in the core of the hypervisor for so many years.”.

Aral:2016:NAE

[AO16] Atakan Aral and Tolga Ovatman. Network-aware embedding of virtual machine clusters onto federated cloud infrastructure. The Journal of Systems and Software, 120(??):89–104, October 2016. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0164121216301078.

Ashraf:2018:MOD

[AP18] Adnan Ashraf and Ivan Porres. Multi-objective dynamic virtual machine consolidation in the cloud using ant colony system. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 33(1):103–120, 2018. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Adoga:2022:NFV

[AP22] Haruna Umar Adoga and Dimitrios P. Pezaros. Network function virtualization and service function chaining frameworks: a comprehensive review of requirements, objectives, implementations, and open research challenges. Future Internet, 14(2):59, February 15, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.com/1999-5903/14/2/59.

Aprea:2009:HVS

[Apr09] Jean-François Apréa. Hyper-V et SC Virtual Machine Manager: technologie de virtualisation sous Windows Server 2008 R2. (French). [Hyper-V and SC Virtual Machine Manager: virtualization technology under Windows Server 2008 R2]. Expert IT. Éditions ENI, Saint Herblain, France, 2009. ISBN 2-7460-5061-7. 537 pp. LCCN ????

Anderson:2005:OII

[APST05] Thomas Anderson, Larry Peterson, Scott Shenker, and Jonathan Turner. Overcoming the Internet impasse through virtualization. Computer, 38(4):34–??, April 2005. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.org/dl/mags/co/
2005/04/r4034.htm;   http://csdl.computer.org/dl/
mags/co/2005/04/r4034.pdf.

Abadi:2018:SCT

[ARA18] Reza Mohamadi Bahram Abadi, Amir Masoud Rahmani, and Sasan H. Alizadeh. Server consolidation techniques in virtualized data centers of cloud environments: a systematic literature review. Software—Practice and Experience, 48(9): 1688–1726, September 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abadi:2020:CSC

[ARA20a] Reza Mohamadi Bahram Abadi, Amir Masoud Rahmani, and Sasan Hossein Alizadeh. Challenges of server consolidation in virtualized data centers and open research issues: a systematic literature review. The Journal of Supercomputing, 76(4): 2876–2927, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [ARA20b].

Abadi:2020:CCS

[ARA20b] Reza Mohamadi Bahram Abadi, Amir Masoud Rahmani, and Sasan Hossein Alizadeh. Correction to: Challenges of server consolidation in virtualized data centers and open research issues: a systematic literature review. The Journal of Supercomputing, 76(4):2928, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://link.springer.com/content/pdf/10.1007/
s11227-019-03105-z.pdf. See [ARA20a].

Al-Rahayfeh:2019:NAT

[ARAAA19] Amer Al-Rahayfeh, Saleh Atiewi, Abdullah Abuhussein, and Muder Almiani. Novel approach to task scheduling and load balancing using the dominant sequence clustering and mean shift clustering algorithms. Future Internet, 11(5):109, May 08, 2019. CODEN ???? ISSN 1999-5903. URL https://

www.mdpi.com/1999-5903/11/5/109.

Arce:2007:GVM

[Arc07] Iván Arce. Ghost in the virtual machine. IEEE Security & Privacy, 5(4):68–71, July/August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Armstrong:1978:PPC

[Arm78] J. L. Armstrong. Programming a parallel computer for robot vision. The Computer Journal, 21(3):215–218, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_21/Issue_03/tiff/215.
tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/216.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_21/Issue_03/tiff/
217.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/218.tif.

Armstrong:1998:CSH

[Arm98] Eric Armstrong. Cover story: HotSpot: a new breed of virtual machine. JavaWorld: IDG’s magazine for the Java community, 3(3):??, March 1998. CODEN ???? ISSN 1091-

8906. URL http://www.javaworld.com/javaworld/jw-03-
1998/jw-03-hotspot.htm.

Arroba:2018:HMD

[ARMMA18] Patricia Arroba, José L. Risco-Martín, José M. Moya, and José L. Ayala. Heuristics and metaheuristics for dynamic management of computing and cooling energy in cloud data centers. Software—Practice and Experience, 48(10):1775– 1804, October 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arvizo:2002:VMT

[Arv02] Teodoro Arvizo. A virtual machine for a type-ω denotational proof language. Thesis (M.Eng. and S.B.), Department of Electrical Engineering and Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, 2002. 123 pp. Supervised by Howard E. Shrobe.

Adix:1976:IER

[AS76] Mary S. Adix and Henrik A. Schutz. Interpretive execution of real-time control applications. ACM SIGPLAN Notices, 11 (4):78–87, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abramski:1985:SMV

[AS85a] S. Abramski and R. Sykes. Secd-m: A Virtual Machine for Applicative Programming. In Jean-Pierre Jouannaud, editor, FPCA ’85, Functional Programming Languages and Computer Architecture, volume 201 of Lecture Notes in Computer Science, pages 81–98. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985.

Abramsky:1985:SMV

[AS85b] Samson Abramsky and R. Sykes. SECD-M: a virtual machine for applicative programming. In Jouannaud [Jou85], pages 81–98. CODEN LNCSD9. ISBN 0-387-15975-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.201. URL http://link.springer-

ny.com/link/service/series/0558/tocs/t0201.htm;
http://www.springerlink.com/openurl.asp?genre=issue&
issn=0302-9743&volume=201.

Atif:2014:APA

[AS14] Muhammad Atif and Peter Strazdins. Adaptive parallel application resource remapping through the live migration of virtual machines. Future Generation Computer Systems, 37(??):148–161, July 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X13001428.

Asyabi:2018:PHC

[ASB18] Esmail Asyabi, Mohsen Sharifi, and Azer Bestavros. ppXen: A hypervisor CPU scheduler for mitigating performance variability in virtualized clouds. Future Generation Computer Systems, 83(??):75–84, June 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167739X17310555.

Alam:2020:SNV

[ASL^{+}20] Iqbal Alam, Kashif Sharif, Fan Li, Zohaib Latif, M. M. Karim, Sujit Biswas, Boubakr Nour, and Yu Wang. A survey of network virtualization techniques for Internet of Things using SDN and NFV. ACM Computing Surveys, 53 (2):35:1–35:40, July 2020. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3379444.

Ahmadian:2021:EET

[ASMA21] S. Ahmadian, R. Salkhordeh, O. Mutlu, and H. Asadi. ET-ICA: Efficient two-level I/O caching architecture for virtualized platforms. IEEE Transactions on Parallel and Distributed Systems, 32(10):2415–2433, October 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alachiotis:2022:SPR

[ASPP22] Nikolaos Alachiotis, Panagiotis Skrimponis, Manolis Pissadakis, and Dionisios Pnevmatikatos. Scalable phylogeny reconstruction with disaggregated near-memory processing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(3):25:1–25:32, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3484983.

Asyabi:2018:TMT

[ASSB18] Esmail Asyabi, SeyedAlireza SanaeeKohroudi, Mohsen Sharifi, and Azer Bestavros. TerrierTail: Mitigating tail latency of cloud virtual machines. IEEE Transactions on Parallel and Distributed Systems, 29(10):2346–2359, October 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2018/10/08338088-abs.html.

Amin:2016:JST

[AT16] Nada Amin and Ross Tate. Java and Scala’s type systems are unsound: the existential crisis of null pointers. ACM SIGPLAN Notices, 51(10):838–848, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ajmera:2023:SPS

[AT23] Kashav Ajmera and Tribhuwan Kumar Tewari. SR-PSO: server residual efficiency-aware particle swarm optimization for dynamic virtual machine scheduling. The Journal of Supercomputing, 79(14):15459–15495, September 2023. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-023-05270-8.

Ajmera:2024:DVM

[AT24] Kashav Ajmera and Tribhuwan Kumar Tewari. Dynamic virtual machine scheduling using residual optimum powerefficiency in the cloud data center. The Computer Journal, 67(3):1099–1110, March 2024. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://academic.

oup.com/comjnl/article/67/3/1099/7161852.

Amit:2014:VMS

[ATS14] Nadav Amit, Dan Tsafrir, and Assaf Schuster. VSwapper: a memory swapper for virtualized environments. ACM SIGARCH Computer Architecture News, 42(1):349– 366, March 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Arianyan:2016:NHC

[ATS16] Ehsan Arianyan, Hassan Taheri, and Saeed Sharifian. Novel heuristics for consolidation of virtual machines in cloud data centers using multi-criteria resource management solutions. The Journal of Supercomputing, 72(2):688–717, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL http://link.springer.com/article/
10.1007/s11227-015-1603-9.

Attansio:1973:VMD

[Att73] C. R. Attansio. Virtual machines and data security. In ????, editor, Proceedings ACM SIGARCH-SIGOPS Workshop on Virtual Computer Systems, Cambridge, MA, 1973, page ?? ACM Press, New York, NY 10036, USA, 1973.

Attanasio:1979:VCS

[Att79] C. R. Attanasio. Virtual control storage — security measures in VM/370. IBM Systems Journal, 18(1):93–110, 1979. CODEN IBMSA7. ISSN 0018-8670.

Alboaneen:2021:MMJ

[ATZP21] Dabiah Alboaneen, Hugo Tianfield, Yan Zhang, and Bernardi Pranggono. A metaheuristic method for joint task scheduling and virtual machine placement in cloud data centers. Future Generation Computer Systems, 115(??):201–212, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X20310967.

Appuswamy:2011:FMF

[AvMT11] Raja Appuswamy, David C. van Moolenbroek, and Andrew S. Tanenbaum. Flexible, modular file volume virtualization in Loris. In André Brinkmann and David Pease, editors, Proceedings of the 27th IEEE Conference on Mass Storage Systems and Technologies (MSST ’11), 23– 27 May 2011, Denver, CO, USA, pages 1–14. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4577-0428-5, 1-4577-0427-7. LCCN TK7895.M4. URL http://

ieeexplore.ieee.org/servlet/opac?punumber=5910685;
 http://www.minix3.org/docs/loris/MSST2011.pdf.

Alaluna:2019:SMC

[AVNR19] Max Alaluna, Eric Vial, Nuno Neves, and Fernando M. V. Ramos. Secure multi-cloud network virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 161(??):45–60, October 9, 2019. CODEN ???? ISSN 1389-1286 (print),

1872-7069 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1389128618312155.

Agarwal:2017:TAT

[AW17] Neha Agarwal and Thomas F. Wenisch. Thermostat: Application-transparent page management for two-tiered main memory. ACM SIGARCH Computer Architecture News, 45(1):631–644, March 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Arnold:2005:IVM

[AWR05] Matthew Arnold, Adam Welc, and V. T. Rajan. Improving virtual machine performance using a cross-run profile repository. ACM SIGPLAN Notices, 40(10):297–311, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Asanuma:2024:DAL

[AY24] Kota Asanuma and Hiroshi Yamada. DBMS-assisted live migration of virtual machines. IEEE Transactions on Computers, 73(2):380–393, February 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Alam:2017:DIY

[AZEE17] Hanna Alam, Tianhao Zhang, Mattan Erez, and Yoav Etsion. Do-it-yourself virtual memory translation. ACM SIGARCH Computer Architecture News, 45(2):457–468, May 2017. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Alam:2018:DIY

[AZEE18] Hanna Alam, Tianhao Zhang, Mattan Erez, and Yoav Etsion. Do-it-yourself virtual memory translation. Operating Systems Review, 52(1):1–12, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Blank:2005:APV

[B^{+}05] Annika Blank et al., editors. Advanced POWER virtualization on IBM System p5. IBM redbooks. IBM Corporation, San Jose, CA, USA, second edition, 2005. ISBN 0-7384-9404-6. xxii + 424 pp. LCCN QA76.9.V5 A38 2005. URL http://

www.loc.gov/catdir/toc/fy0608/2006273965.html.

Buytaert:2007:BDS

[B^{+}07] Kris Buytaert et al., editors. Best damn server virtualization book period: including Vmware, Xen, and Microsoft Virtual Server. Syngress Publishing, Inc., Rockland, MA, USA, 2007. ISBN 1-59749-217-5 (paperback). xxviii + 931 pp. LCCN QA76.9.C55 B475 2007.

Berglund:2019:MSJ

[BA19] Lasse Berglund and Cyrille Artho. Method summaries for JPF. ACM SIGSOFT Software Engineering Notes, 44(4):16, December 2019. CODEN SFENDP. ISSN 0163-5948 (print),

1943-5843 (electronic). URL https://dl.acm.org/doi/10.
1145/3364452.33644560.

Besseling:2023:UTT

[BA23] Johan Besseling and Cyrille Artho. Using Theia trace viewer to visualize JPF traces. ACM SIGSOFT Software Engineering Notes, 48(1):22–26, January 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/3573074.3573081.

Bacon:2011:VAH

[Bac11] David F. Bacon. Virtualization in the age of heterogeneous machines. ACM SIGPLAN Notices, 46(7):1–2, July 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baccarelli:2015:MEB

[BAC15] Enzo Baccarelli, Danilo Amendola, and Nicola Cordeschi. Minimum-energy bandwidth management for QoS live migration of virtual machines. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 1)(??):1–22, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-

7069 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1389128615003655.

Baden:1982:HPS

[Bad82] Scott B. Baden. High performance storage reclamation in an object-based memory system. Technical Report CSD-84-167, University of California, Berkeley, June 1982. 52 pp.

Baden:1987:RTP

[Bad87] Scott B. Baden. Run-time partitioning of scientific continuum calculations running on multiprocessors. Technical Report CSD-87-366, University of California, Berkeley, June 1987. 154 pp.

Bockisch:2006:AVMa

[BADM06] Christoph Bockisch, Matthew Arnold, Tom Dinkelaker, and Mira Mezini. Adapting virtual machine techniques for seamless aspect support. ACM SIGPLAN Notices, 41(10):109– 124, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bagley:1976:SFM

[Bag76] John D. Bagley. Special feature: microprogrammable virtual machines. Computer, 9(2):38–42, February 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bairstow:1970:MOV

[Bai70] J. N. Bairstow. Many from one: The virtual machine arrives. Computer Decisions, ??(??):29–31, January 1970.

Baker:1983:MAS

[Bak83] T. P. Baker. Mapping Ada onto a simple virtual machine. Technical Report ADA-83-02, Florida State University, 1983.

Balzer:1991:PVM

[Bal91] Robert Balzer. Process virtual machine. In Thomas [Tho93], pages 37–40. ISBN 0-8186-4050-2, 0-8186-4051-0. LCCN QA76.755 .I58 1991.

Bauman:2015:SHB

[BAL15] Erick Bauman, Gbadebo Ayoade, and Zhiqiang Lin. A survey on hypervisor-based monitoring: Approaches, applications, and evolutions. ACM Computing Surveys, 48(1):10:1–10:??, September 2015. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Bard:1973:AMC

[Bar73] Y. Bard. An analytic model of CP-67 — VM/370. In ????, editor, Proceedings ACM SIGARCH-SIGOPS Workshop on Virtual Computer Systems, Cambridge, MA, 1973, page ?? ACM Press, New York, NY 10036, USA, 1973.

Bard:1978:AMV

[Bar78] Y. Bard. An analytic model of the VM/370 system. IBM Journal of Research and Development, 22(5):498–508, September 1978. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Bartholomew:2006:QMM

[Bar06] Daniel Bartholomew. QEMU: a multihost, multitarget emulator. Linux Journal, 2006(145):??, May 2006. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Bastiaansen:2004:RGU

[Bas04] Rob Bastiaansen. Rob’s Guide to Using VMWARE. Books4Brains, ????, 2004. ISBN 90-808934-1-2. 268 (est.) pp. LCCN ????

Bastiaansen:2006:RGU

[Bas06] Rob Bastiaansen. Rob’s Guide to Using VMWARE. Books4Brains, ????, second edition, 2006. ISBN 90-808934-1-2. 350 (est.) pp. LCCN ????

Bauer:2005:PPF

[Bau05] Mick Bauer. Paranoid penguin: The future of Linux security. Linux Journal, 2005(136):10, August 2005. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Bauer:2006:PPSb

[Bau06a] Mick Bauer. Paranoid penguin: security features in Debian 3.1. Linux Journal, 2006(145):??, May 2006. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Bauer:2006:PPSa

[Bau06b] Mick Bauer. Paranoid penguin: security features in SUSE 10.0. Linux Journal, 2006(144):??, April 2006. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Bauer:2006:VWL

[Bau06c] Mick Bauer. VMware Workstation 5.5 for Linux hosts. Linux Journal, 2006(145):??, May 2006. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Bunge:1995:MCM

[BB95] Hans-Peter Bunge and John R. Baumgardner. Mantle convection modeling on parallel virtual machines. Computers in Physics, 9(2):207–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168525.

Bonardi:2008:PEM

[BB08] Alain Bonardi and Jérome Barthélemy. The preservation, emulation, migration, and virtualization of live electronics for performing arts: an overview of musical and technical issues. Journal on Computing and Cultural Heritage (JOCCH), 1(1): 6:1–6:??, June 2008. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Beloglazov:2012:OOD

[BB12] Anton Beloglazov and Rajkumar Buyya. Optimal online deterministic algorithms and adaptive heuristics for energy and performance efficient dynamic consolidation of virtual machines in Cloud data centers. Concurrency and Computation: Practice and Experience, 24(13):1397–1420, September 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Beloglazov:2013:MOH

[BB13] Anton Beloglazov and Rajkumar Buyya. Managing overloaded hosts for dynamic consolidation of virtual machines in cloud data centers under quality of service constraints. IEEE Transactions on Parallel and Distributed Systems, 24 (7):1366–1379, July 2013. CODEN ITDSEO. ISSN 1045-9219.

Beloglazov:2015:ONF

[BB15] Anton Beloglazov and Rajkumar Buyya. OpenStack Neat: a framework for dynamic and energy-efficient consolidation of virtual machines in OpenStack clouds. Concurrency and Computation: Practice and Experience, 27(5):1310–1333, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bassem:2017:MCP

[BB17] Christine Bassem and Azer Bestavros. Multi-capacity bin packing with dependent items and its application to the packing of brokered workloads in virtualized environments. Future Generation Computer Systems, 72(??):129–144, July 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X16302771.

Baalamurugan:2020:MOK

[BB20] K. M. Baalamurugan and S. Vijay Bhanu. A multi-objective krill herd algorithm for virtual machine placement in cloud computing. The Journal of Supercomputing, 76(6):4525–4542, June 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Berry:2025:ITI

[BBB^{+}25] Christopher Berry, Michael Becht, Tim Bubb, Howard Haynie, Robert Sonnelitter, Katie Seggerman, Jonathan Hsieh, Edward Malley, Mike Cadigan, Susan M. Eickhoff, Matthias Klein, Craig Walters, Christian G. Zoellin, and Cedric Lichtenau. The IBM Telum II processor. IEEE Micro, 45(3):66–75, May/June 2025. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Balter:1991:AIG

[BBD^{+}91] R. Balter, J. Bernadat, D. Decouchant, A. Duda, A. Freyssinet, S. Krakowiak, M. Meysembourg, P. Le Dot, H. Nguyen Van, E. Paire, M. Riveill, C. Roison, X. Rousset de Pina, R. Scioville, and G. Vandôme. Architecture and implementation of guide, an object-oriented distributed system. Computing Systems, 4(1):31–67, Winter 1991. CODEN CMSYE2. ISSN 0895-6340.

Barr:2010:VMV

[BBD^{+}10] Ken Barr, Prashanth Bungale, Stephen Deasy, Viktor Gyuris, Perry Hung, Craig Newell, Harvey Tuch, and Bruno Zoppis. The VMware mobile virtualization platform: is that a hypervisor in your pocket? Operating Systems Review, 44(4): 124–135, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bhattiprolu:2008:VSC

[BBHL08] Sukadev Bhattiprolu, Eric W. Biederman, Serge Hallyn, and Daniel Lezcano. Virtual servers and checkpoint/restart in mainstream Linux. Operating Systems Review, 42(5):104– 113, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bratanov:2009:VMW

[BBM09] Stanislav Bratanov, Roman Belenov, and Nikita Manovich. Virtual machines: a whole new world for performance analysis. Operating Systems Review, 43(2):46–55, April 2009. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birke:2015:WVM

[BBM^{+}15] Robert Birke, Mathias Björkqvist, Cyriel Minkenberg, Martin Schmatz, and Lydia Y. Chen. When virtual meets physical at the edge: a field study on datacenters’ virtual traffic. ACM SIGMETRICS Performance Evaluation Review, 43(1): 403–415, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bennett:1991:SMC

[BBMA91] R. B. Bennett, W. J. Bitner, M. A. Musa, and M. K. Ainsworth. Systems management for Coordinated Resource Recovery. IBM Systems Journal, 30(1):90–106, 1991. CO-DEN IBMSA7. ISSN 0018-8670.

Bullers:2006:VMI

[BBS06] William I. Bullers, Jr., Stephen Burd, and Alessandro F. Seazzu. Virtual machines — an idea whose time has returned: application to network, security, and database courses. In ACM [ACM06d], pages 102–106. CODEN SIGSD3. ISBN ???? ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA76.27. ACM order number 457060.

Barrett:2017:VMW

[BBTK^{+}17] Edd Barrett, Carl Friedrich Bolz-Tereick, Rebecca Killick, Sarah Mount, and Laurence Tratt. Virtual machine warmup blows hot and cold. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):52:1–52:??, October 2017. CODEN ???? ISSN 2475-1421.

Boutcher:2010:DVM

[BC10] David Boutcher and Abhishek Chandra. Does virtualization make disk scheduling passé? Operating Systems Review, 44 (1):20–24, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bertolazzi:2019:MED

[BC19] Marco Bertolazzi and Carlo Caini. Mars to Earth data downloading: a directory synchronization approach. Future Internet, 11(8):173, August 08, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.mdpi.com/1999-5903/11/8/173.

Bellavista:2015:VNF

[BCC^{+}15] Paolo Bellavista, Franco Callegati, Walter Cerroni, Chiara Contoli, Antonio Corradi, Luca Foschini, Alessandro Pernafini, and Giuliano Santandrea. Virtual network function embedding in real cloud environments. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3)(??):506–517, December 24, 2015. CODEN ???? ISSN 1389-1286 (print),

1872-7069 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1389128615003588.

Buzen:1973:NVM

[BCG73a] J. P. Buzen, P. P. Chen, and R. P. Goldberg. A note on virtual machines and software reliability. In ????, editor, Proceedings ACM SIGARCH-SIGOPS Workshop on Virtual Computer Systems, Cambridge, MA, 1973, page ?? ACM Press, New York, NY 10036, USA, 1973.

Buzen:1973:VMT

[BCG73b] J. P. Buzen, P. P. Chen, and R. P. Goldberg. Virtual machine techniques for improving software reliability. In ????, editor, Proceedings IEEE Symposium on Computer Software Reliability, New York, 1973, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1973.

Bessiere:1990:VMM

[BCM90] P. Bessiere, A. Chams, and T. Muntean. A virtual machine model for artificial neural network programming. In Proceedings of the International Neural Network Conference Paris, pages 689–692. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990.

Berger:2008:TMS

[BCP^{+}08] Stefan Berger, Ramón Cáceres, Dimitrios Pendarakis, Reiner Sailer, Enriquillo Valdez, Ronald Perez, Wayne Schildhauer, and Deepa Srinivasan. TVDc: managing security in the trusted virtual datacenter. Operating Systems Review, 42(1): 40–47, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Berghaus:2020:HTC

[BCW20] F. Berghaus, K. Casteels, and J. Weldon. High-throughput cloud computing with the cloudscheduler VM provisioning service. Computing and Software for Big Science, 4(1):??, December 2020. CODEN ???? ISSN 2510-2036 (print),

2510-2044 (electronic). URL https://link.springer.com/
article/10.1007/s41781-020-0036-1.

Bruschi:2019:AAF

[BCZ19] Roberto Bruschi, Florin Ciucu, and Thomas Zinner. Application areas and fundamental challenges in Network Functions Virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 154(??):26–27, May 8, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128619302348.

Bredlau:2001:ALT

[BD01] Carl Bredlau and Dorothy Deremer. Assembly language through the Java Virtual Machine. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (1):194–198, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baride:2011:CBS

[BD11] Srikanth Baride and Kamlesh Dutta. A cloud based software testing paradigm for mobile applications. ACM SIGSOFT Software Engineering Notes, 36(3):1–4, May 2011. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brown:2003:SFE

[BDE^{+}03] Rhodes Brown, Karel Driesen, David Eng, Laurie Hendren, John Jorgensen, Clark Verbrugge, and Qin Wang. STEP: a framework for the efficient encoding of general trace data. ACM SIGSOFT Software Engineering Notes, 28(1):27–34, January 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bak:1998:NCJ

[BDF^{+}98] Lars Bak, John Duimovich, Jesse Fang, Scott Meyer, and David Ungar. The new crop of Java virtual machines (panel). ACM SIGPLAN Notices, 33(10):179–182, October 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beck:1999:HNG

[BDF^{+}99] Micah Beck, Jack J. Dongarra, Graham E. Fagg, G. Al Geist, Paul Gray, James Kohl, Mauro Migliardi, Keith Moore, Terry Moore, Philip Papadopoulous, Stephen L. Scott, and Vaidy Sunderam. HARNESS: a next generation distributed virtual machine. Future Generation Computer Systems, 15(5–6):571–582, October 1, 1999. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.elsevier.com/gej-ng/10/19/19/30/21/20/
abstract.html;   http://www.netlib.org/utk/people/
JackDongarra/PAPERS/harness2.ps.

Barham:2003:VMM

[BDF^{+}03] Paul Barham, Boris Dragovic, Keir Fraser, Steven Hand, Tim Harris, Alex Ho, Rolf Neugebauer, Ian Pratt, and Andrew Warfield. Virtual machine monitors: Xen and the art of virtualization. In ACM [ACM03b], pages 164–177. ISBN 1-58113-757-5. ISSN 0163-5980 (print), 1943-586X (elec-

tronic). URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://uclibs.org/PID/34720. ACM or-

der number 534030.

Bonfim:2019:INS

[BDF19] Michel S. Bonfim, Kelvin L. Dias, and Stenio F. L. Fernandes. Integrated NFV/SDN architectures: a systematic literature review. ACM Computing Surveys, 51(6):114:1–114:??, February 2019. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-

7341 (electronic). URL https://dl.acm.org/ft_gateway.
cfm?id=3172866.

Botacin:2018:WWW

[BDG18] Marcus Botacin, Paulo Lício De Geus, and André Grégio. Who watches the watchmen: a security-focused review on current state-of-the-art techniques, tools, and methods for systems and binary analysis on modern platforms. ACM Computing Surveys, 51(4):69:1–69:??, September 2018. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Barthe:2002:FCB

[BDJdS02] Gilles Barthe, Guillaume Dufay, Line Jakubiec, and Simão Melo de Sousa. A formal correspondence between offensive and defensive JavaCard virtual machines. Lecture Notes in Computer Science, 2294:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2294/22940032.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2294/22940032.pdf.

Butrico:2008:SEE

[BDK^{+}08] Maria Butrico, Dilma Da Silva, Orran Krieger, Michal Ostrowski, Bryan Rosenburg, Dan Tsafrir, Eric Van Hensbergen, Robert W. Wisniewski, and Jimi Xenidis. Specialized execution environments. Operating Systems Review, 42(1):

106–107, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baresi:2023:MFA

[BDQR23] Luciano Baresi, Tommaso Dolci, Giovanni Quattrocchi, and Nicholas Rasi. A multi-faceted analysis of the performance variability of virtual machines. Software—Practice and Experience, 53(11):2067–2091, November 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bugnion:2012:BVX

[BDR^{+}12] Edouard Bugnion, Scott Devine, Mendel Rosenblum, Jeremy Sugerman, and Edward Y. Wang. Bringing virtualization to the x86 architecture with the original VMware workstation. ACM Transactions on Computer Systems, 30(4):12:1–12:51, November 2012. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Baldwin:2009:PSS

[BDS^{+}09] Adrian Baldwin, Chris Dalton, Simon Shiu, Krzysztof Kostienko, and Qasim Rajpoot. Providing secure services for a virtual infrastructure. Operating Systems Review, 43(1): 44–51, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bolz:2013:SSC

[BDT13] Carl Friedrich Bolz, Lukas Diekmann, and Laurence Tratt. Storage strategies for collections in dynamically typed languages. ACM SIGPLAN Notices, 48(10):167–182, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Beckert:2017:RTA

[BE17] Matthias Beckert and Rolf Ernst. Response time analysis for sporadic server based budget scheduling in real time virtualization environments. ACM Transactions on Embedded Computing Systems, 16(5s):161:1–161:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Becker:2009:VIA

[Bec09] Dirk Becker. VirtualBox Installation, Anwendung, Praxis. (German) [VirtualBox Installation, Use, and Practice].

Galileo computing. Galileo Press, Bonn, Germany, 2009. ISBN 3-8362-1374-5. 321 pp. LCCN ???? URL http://d-

nb.info/993523552/04;   http://www.galileocomputing.
de.

Beebe:2005:VM

[Bee05] Nelson H. F. Beebe. Virtual machines. World-Wide Web document., August 1, 2005. URL https://www.math.utah.

edu/~beebe/vm.html.

Begnum:2012:SCO

[Beg12] Kyrre Begnum. Simplified cloud-oriented virtual machine management with MLN. The Journal of Supercomputing, 61(2):251–266, August 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&
issn=0920-8542&volume=61&issue=2&spage=251.

Bellovin:2006:IRV

[Bel06] Steven M. Bellovin. Inside risks: Virtual machines, virtual security? Communications of the ACM, 49(10):104, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bendechache:2021:SER

[Ben21] Malika Bendechache. Simulating and evaluating a real-world ElasticSearch system using the RECAP DES simulator. Future Internet, 13(4):83, March 24, 2021. CODEN ???? ISSN

1999-5903. URL https://www.mdpi.com/1999-5903/13/4/
83.

Bernat:1986:IIG

[Ber86] Andrew P. Bernat. An interactive interpreter/graphicsimulator for IBM S/370 architecture assembly language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 18(2):13–16, June 1986. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bosilca:2002:OOE

[BFC02] George Bosilca, Gilles Fedak, and Franck Cappello. OVM: Out-of-order execution parallel virtual machine. Future Generation Computer Systems, 18(4):525–537, March 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.elsevier.com/gej-ng/10/19/
19/60/33/34/abstract.html.

Bienkowski:2014:WAV

[BFG^{+}14] Marcin Bienkowski, Anja Feldmann, Johannes Grassler, Gregor Schaffrath, and Stefan Schmid. The wide-area virtual service migration problem: a competitive analysis approach. IEEE/ACM Transactions on Networking, 22(1):165– 178, February 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bagley:1975:SDS

[BFHW75] John D. Bagley, E. R. Floto, S. C. Hsieh, and V. Watson. Sharing data and services in a virtual machine system. In ACM [ACM75], pages 82–88. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Buchbinder:2021:OVM

[BFM^{+}21] Niv Buchbinder, Yaron Fairstein, Konstantina Mellou, Ishai Menache, and Joseph (Seffi) Naor. Online virtual machine allocation with lifetime and load predictions. ACM SIG-METRICS Performance Evaluation Review, 49(1):9–10, June 2021. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3543516.
3456278.

Bruno:2018:DVM

[BFS^{+}18] Rodrigo Bruno, Paulo Ferreira, Ruslan Synytsky, Tetiana Fydorenchyk, Jia Rao, Hang Huang, and Song Wu. Dynamic vertical memory scalability for OpenJDK cloud applications. ACM SIGPLAN Notices, 53(5):59–70, May 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buzen:1973:EVM

[BG73a] J. P. Buzen and U. O. Gagliardi. The evolution of virtual machine architecture. In ????, editor, AFIPS Conference Proceedings, 1973 NCC, pages 290–299. AFIPS Press, Montvale,

NJ, USA, 1973. URL https://dl.acm.org/doi/pdf/10.
1145/1499586.1499667.

Buzen:1973:IVM

[BG73b] J. P. Buzen and U. O. Gagliardi. Introduction to virtual machines. Honeywell Computer Journal, 7(4):??, ???? 1973. CODEN HNCJA3. ISSN 0046-7847.

Buzen:1974:VMT

[BG74] J. P. Buzen and R. P. Goldberg. Virtual machine techniques for introducing peripherals into computer systems. In ????, editor, Computer Peripherals — Benefactor or Bottleneck? Digest of Papers COMPCON 74, San Francisco, February 1974, pages 157–160. ????, ????, 1974.

Bays:2020:RSV

[BG20] Leonardo Richter Bays and Luciano Paschoal Gaspary. Reality shock in virtual network embedding: Flexibilizing demands for dealing with multiple operational requirements in SDNs. Journal of Network and Computer Applications, 153(??):??, March 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804519303686.

Brawn:1970:SPE

[BGM70] B. S. Brawn, F. G. Gustavson, and E. S. Mankin. Sorting in a paging environment. Communications of the ACM, 13(8): 483–494, August 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boszormenyi:2000:SNW

[BGP00] László Böszörményi, Jurg Gutknecht, and Gustav Pomberger, editors. The School of Niklaus Wirth: The Art of Simplicity. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-723-4. viii + 260 pp. LCCN QA76.756

.S36 2000. US$39.95. URL http://www.mkp.com/books_
catalog/catalog.asp?ISBN=1-55860-723-4.

Birmingham:1989:MSC

[BGS89] William P. Birmingham, Anurag P. Gupta, and Daniel P. Siewiorek. The Micon system for computer design. IEEE Micro, 9(5):61–67, September/October 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bartholomy:2013:NMT

[BGS13] E. Bartholomy, G. Greenlee, and M. Sylvia. The need to move toward virtualized and more resilient disaster-recovery architectures. IBM Journal of Research and Development, 57(5):1:1–1:10, September–October 2013. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Bellino:1973:VMV

[BH73] J. Bellino and C. Hans. Virtual machine or virtual operating system. In ????, editor, Proceedings ACM SIGARCH-SIGOPS Workshop on Virtual Computer Systems, Cambridge, MA, 1973, page ?? ACM Press, New York, NY 10036, USA, 1973.

Botero:2013:GNN

[BH13] Juan Felipe Botero and Xavier Hesselbach. Greener networking in a network virtualization environment. Computer Networks (Amsterdam, Netherlands: 1999), 57(9):2021–2039, June 19, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-

7069 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1389128613001151.

Bertels:2009:EMM

[BHDS09] Peter Bertels, Wim Heirman, Erik D’Hollander, and Dirk Stroobandt. Efficient memory management for hardware accelerated Java Virtual Machines. ACM Transactions on Design Automation of Electronic Systems., 14(4):48:1–48:??, August 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bourguiba:2014:INV

[BHEP14] Manel Bourguiba, Kamel Haddadou, Ines El Korbi, and Guy Pujolle. Improving network I/O virtualization for cloud computing. IEEE Transactions on Parallel and Distributed Systems, 25(3):673–681, March 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Biedermann:2015:SDR

[BHI15] Alexander Biedermann, Sorin A. Huss, and Adeel Israr. Safe dynamic reshaping of reconfigurable MPSoC embedded systems for self-healing and self-adaption purposes. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 8(4):26:1–26:??, October 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Biswas:2014:DES

[BHSB14] Swarnendu Biswas, Jipeng Huang, Aritra Sengupta, and Michael D. Bond. DoubleChecker: efficient sound and precise atomicity checking. ACM SIGPLAN Notices, 49(6):28– 39, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barr:2005:JEA

[BHvR05] Rimon Barr, Zygmunt J. Haas, and Robbert van Renesse. JiST: an efficient approach to simulation using virtual machines. Software—Practice and Experience, 35(6):539–576, May 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Biradar:1994:ADL

[Bir94] Umesh V. Biradar. Adaptive distributed load balancing model for parallel virtual machine. Master of science in computer science, Department of Computer Science, College of Engineering, Lamar University, Beaumont, TX, USA, 1994. viii + 44 pp.

Bermejo:2020:VMC

[BJ20] Belen Bermejo and Carlos Juiz. Virtual machine consolidation: a systematic review of its overhead influencing factors. The Journal of Supercomputing, 76(1):324–361, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bermejo:2022:GME

[BJ22] Belen Bermejo and Carlos Juiz. A general method for evaluating the overhead when consolidating servers: performance degradation in virtual machines and containers. The Journal of Supercomputing, 78(9):11345–11372, June 2022. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-022-04318-5.

Bermejo:2024:GNC

[BJC24] Belen Bermejo, Carlos Juiz, and Maria Carla Calzarossa. The goodness of nesting containers in virtual machines for server consolidation. Journal of Grid Computing, 22(4):??, December 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL https://link.springer.com/article/
10.1007/s10723-024-09782-2.

Bermejo:2019:VCS

[BJG19] Belen Bermejo, Carlos Juiz, and Carlos Guerrero. Virtualization and consolidation: a systematic review of the past 10 years of research on energy and performance. The Journal of Supercomputing, 75(2):808–836, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Burtsev:2016:APV

[BJH^{+}16] Anton Burtsev, David Johnson, Mike Hibler, Eric Eide, and John Regehr. Abstractions for practical virtual machine replay. ACM SIGPLAN Notices, 51(7):93–106, July 2016. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berthaud:1973:CVM

[BJPS73] M. Berthaud, M. Jacolin, Ph. Potin, and H. Savary. Coupling virtual machines and system construction. In ????, editor, Proceedings ACM SIGARCH-SIGOPS Workshop on Virtual Computer Systems, Cambridge, MA, 1973, page ?? ACM Press, New York, NY 10036, USA, 1973.

Bell:2014:PID

[BK14] Jonathan Bell and Gail Kaiser. Phosphor: illuminating dynamic data flow in commodity JVMs. ACM SIGPLAN Notices, 49(10):83–101, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:2013:OCC

[BKC^{+}13] Michael D. Bond, Milind Kulkarni, Man Cao, Minjia Zhang, Meisam Fathi Salmi, Swarnendu Biswas, Aritra Sengupta, and Jipeng Huang. OCTET: capturing and controlling crossthread dependences efficiently. ACM SIGPLAN Notices, 48 (10):693–712, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOP-SLA ’13 conference proceedings.

Bockisch:2006:ECF

[BKH^{+}06] Christoph Bockisch, Sebastian Kanthak, Michael Haupt, Matthew Arnold, and Mira Mezini. Efficient control flow quantification. ACM SIGPLAN Notices, 41(12):125–138, December 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bompotas:2025:CMP

[BKM25] Agorakis Bompotas, Nikitas-Rigas Kalogeropoulos, and Christos Makris. CommC: a multi-purpose COMModity hardware cluster. Future Internet, 17(3):121, March 11, 2025. CODEN ???? ISSN 1999-5903. URL https://www.mdpi.

com/1999-5903/17/3/121.

Bergh:1987:HEH

[BKMM87] Arndt B. Bergh, Keith Keilman, Daniel J. Magenheimer, and James A. Miller. HP 3000 emulation on HP precision architecture computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38 (11):87–89, December 1987. CODEN HPJOAX. ISSN 0018-1153.

Bir:2020:DIE

[BKR20] Parth Bir, Shylaja Vinaykumar Karatangi, and Amrita Rai. Design and implementation of an elastic processor with hyperthreading technology and virtualization for elastic server models. The Journal of Supercomputing, 76(9):7394–7415, September 2020. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL https://link.springer.com/
article/10.1007/s11227-020-03174-5.

Benmakrelouf:2019:RNP

[BKT^{+}19] Souhila Benmakrelouf, Nadjia Kara, Hanine Tout, Rafi Rabipour, and Claes Edstrom. Resource needs prediction in virtualized systems: Generic proactive and selfadaptive solution. Journal of Network and Computer Applications, 148(??):??, December 15, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804519303030.

Board:1990:PPN

[BL90] J. A. Board, Jr. and J. Shue-Jen Lu. Performance of parallel neural network simulations. In Board [Boa90], pages 185–200.

Bianchi:2017:MRB

[BL17] Francesco Bianchi and Francesco Lo Presti. A Markov reward based resource-latency aware heuristic for the virtual network embedding problem. ACM SIGMETRICS Performance Evaluation Review, 44(4):57–68, March 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Blelloch:1989:SPP

[Ble89] G. E. Blelloch. SCAN PRIMITIVES AND PARALLEL VEC-TOR MODELS. Technical Report MIT/LCS/TR-463, Massachusetts Institute of Technology, Laboratory for Computer Science, July 1989. 256 pp. Cost is $26. Date: Oct. 1989 Keywords: parallel algorithms, parallel computing, P-RAM, parallel compilers, data-parallel, scan operation, vector models.

Bledsoe:2010:VLO

[Ble10] Greg Bledsoe. Virtualization the Linux/OSS way. Linux Journal, 2010(198):6:1–6:??, October 2010. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Benomar:2022:CBN

[BLMP22] Zakaria Benomar, Francesco Longo, Giovanni Merlino, and Antonio Puliafito. Cloud-based network virtualization in IoT with OpenStack. ACM Transactions on Internet Technology (TOIT), 22(1):19:1–19:26, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.1145/3460818.

Bershad:1994:ACM

[BLRC94] Brian N. Bershad, Dennis Lee, Theodore H. Romer, and J. Bradley Chen. Avoiding conflict misses dynamically in large direct-mapped caches. ACM SIGPLAN Notices, 29(11):158–170, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p158-bershad/.

Blunden:2002:VMD

[Blu02] Bill Blunden. Virtual machine design and implementation in C/C++. Wordware Publishing, Plano, TX, USA, 2002. ISBN 1-55622-903-8 (paperback). xvii + 668 pp. LCCN QA76.9.V5 B59 2002.

Burtsev:2014:WSL

[BMER14] Anton Burtsev, Nikhil Mishrikoti, Eric Eide, and Robert Ricci. Weir: a streaming language for performance analysis. Operating Systems Review, 48(1):65–70, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Belgacem:2023:MLM

[BMF23] Ali Belgacem, Saïd Mahmoudi, and Mohamed Amine Ferrag. A machine learning model for improving virtual machine migration in cloud computing. The Journal of Supercomputing, 79(9):9486–9508, June 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https://link.

springer.com/article/10.1007/s11227-022-05031-z.

Battisti:2022:NSV

[BMJ^{+}22] Anselmo Luiz Éden Battisti, Evandro Luiz Cardoso Macedo, Marina Ivanov Pereira Josué, Hugo Barbalho, Flávia C. Delicato, Débora Christina Muchaluat-Saade, Paulo F. Pires, Douglas Paulo de Mattos, and Ana Cristina Bernardo de Oliveira. A novel strategy for VNF placement in edge computing environments. Future Internet, 14(12):361, November 30, 2022. CODEN ???? ISSN 1999-5903. URL https://

www.mdpi.com/1999-5903/14/12/361.

Bai:2013:HPI

[BML^{+}13] Yuebin Bai, Yao Ma, Cheng Luo, Duo Lv, and Yuanfeng Peng. A high performance inter-domain communication approach for virtual machines. The Journal of Systems and Software, 86(2):367–376, February 2013. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121212002476.

Banerjee:2016:SNA

[BMS16] Amitabha Banerjee, Rishi Mehta, and Zach Shen. Supporting NUMA-aware I/O in virtual machines. IEEE Micro, 36 (4):28–36, July/August 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https://www.

computer.org/csdl/mags/mi/2016/04/mmi2016040028-abs.
html.

Ballard:1986:QSD

[BMWB86] M. B. Ballard, D. Maier, and A. Wirfs-Brock. QUICKTALK: a Smalltalk-80 dialect for defining primitive methods. ACM SIGPLAN Notices, 21(11):140, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Belpaire:1975:FPR

[BN75] Gerald Belpaire and Nai-Ting Nsu. Formal properties of recursive virtual machine architectures. In ACM [ACM75], pages 89–96. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Blandy:1989:VSM

[BN89] G. O. Blandy and S. R. Newson. VM/XA storage management. IBM Systems Journal, 28(1):175–191, 1989. CODEN IBMSA7. ISSN 0018-8670.

Brogi:2018:MBA

[BNS18] Antonio Brogi, Davide Neri, and Jacopo Soldani. A microservice-based architecture for (customisable) analyses of Docker images. Software—Practice and Experience, 48 (8):1461–1474, August 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/spe.2583.

Board:1990:TRA

[Boa90] J. A. Board, editor. Transputer Research and Applications 2. NATUG-2 Proceedings of the North American Transputer Users Group, Durham, NC, USA, October 18–19, 1989. IOS Press, Amsterdam, The Netherlands, 1990.

Bodlaender:1988:CFU

[Bod88] Hans L. Bodlaender. The complexity of finding uniform emulations on fixed graphs. Information Processing Letters, 29 (3):137–141, October 26, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boddenberg:2010:WSR

[Bod10] Ulrich B. Boddenberg. Windows Server 2008 R2 das umfassende Handbuch. (German) [Windows Server 2008 R2: the Complete Handbook]. Galileo computing. Galileo Press, Bonn, Germany, third edition, 2010. ISBN 3-8362-1528-4.

1410 pp. LCCN ???? URL http://www.galileocomputing.
de/.

Boettiger:2015:IDR

[Boe15] Carl Boettiger. An introduction to Docker for reproducible research. Operating Systems Review, 49(1):71–79, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bruno:2017:NPG

[BOF17] Rodrigo Bruno, Luís Picciochi Oliveira, and Paulo Ferreira. NG2C: pretenuring garbage collection with dynamic generations for HotSpot big data applications. ACM SIGPLAN Notices, 52(9):2–13, September 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Born:2001:VWP

[Bor01] Günter Born. VMware Workstation Praxisführer: Installation, Konfiguration, Anwendung unter Windows und Linux. SuSE-PRESS, Nürnberg, Germany, 2001. ISBN 3-934678-81-5. xxii + 372 pp. LCCN ???? DM 79.00, EUR 40.00.

Border:2007:DDM

[Bor07] Charles Border. The development and deployment of a multiuser, remote access virtualization system for networking, security, and system administration classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):576–580, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bosse:2024:VMP

[Bos24] Stefan Bosse. A virtual machine platform providing machine learning as a programmable and distributed service for IoT and edge on-device computing: Architecture, transformation, and evaluation of integer discretization. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/8/356.

Bozman:1989:VSM

[Boz89] G. P. Bozman. VM/XA SP2 minidisk cache. IBM Systems Journal, 28(1):165–174, 1989. CODEN IBMSA7. ISSN 0018-8670.

Barbosa:1999:ADM

[BP99] J. Barbosa and A. Padilha. Algorithm-dependent method to determine the optimal number of computers in parallel virtual machines. Lecture Notes in Computer Science, 1573: 508–521, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breg:2001:JVM

[BP01] Fabian Breg and Constantine Polychronopoulos. Java Virtual Machine support for object serialization. In ACM [ACM01b], pages 173–180. ISBN 1-58113-359-6. LCCN

QA76.9.O35 A26 2001. URL http://www.philippsen.
com/JGI2001/camerareadyabstracts/12.html;   http:
//www.philippsen.com/JGI2001/finalpapers/18500173.
ps.

Breg:2003:JVM

[BP03] Fabian Breg and Constantine D. Polychronopoulos. Java Virtual Machine support for object serialization. Concurrency and Computation: Practice and Experience, 15(3–5): 263–275, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhaskar:1986:VIO

[BPB86] K. S. Bhaskar, J. K. Peckol, and J. L. Beug. Virtual Instruments: object-oriented program synthesis. ACM SIG-PLAN Notices, 21(11):303, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beletsky:1994:OPV

[BPC94] V. Beletsky, T. Popova, and A. Chemeris. Organization of a parallel virtual machine. In Horiguchi et al. [HHK94], pages 421–426. ISBN 0-8186-6507-6. LCCN QA76.58 .I5673 1994. IEEE catalog no. 94TH0697-3.

Bhattacharya:2022:PMI

[BPM^{+}22] Tathagata Bhattacharya, Xiaopu Peng, Jianzhou Mao, Chaowei Zhang, Taha Takreeti, Ye Wang, Ting Cao, and Xiao Qin. Performance modeling for I/O-intensive applications on virtual machines. Concurrency and Computation: Practice and Experience, 34(10):e6823:1–e6823:??, May 1, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Belay:2017:IOS

[BPP^{+}17] Adam Belay, George Prekas, Mia Primorac, Ana Klimovic, Samuel Grossman, Christos Kozyrakis, and Edouard Bugnion. The IX operating system: Combining low latency, high throughput, and efficiency in a protected dataplane. ACM Transactions on Computer Systems, 34(4):11:1–11:??, January 2017. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Braught:2001:KSC

[BR01] Grant Braught and David Reed. The knob & switch computer: a computer architecture simulator for introductory computer science. ACM Journal on Educational Resources in Computing (JERIC), 1(4):31–45, December 2001. CO-DEN ???? ISSN 1531-4278.

Brier:1998:NIA

[Bri98] Steven E. Brier. News: IBM augments OS/2 with new Java virtual machine. JavaWorld: IDG’s magazine for the Java community, 3(3):??, March 1998. CODEN ???? ISSN 1091-

8906. URL http://www.javaworld.com/javaworld/jw-03-
1998/jw-03-infoworld.ibm.htm.

Berl:2010:NVE

[BRIdM10] Andreas Berl, Nicholas Race, Johnathan Ishmael, and Hermann de Meer. Network virtualization in energy-efficient office environments. Computer Networks (Amsterdam, Netherlands: 1999), 54(16):2856–2868, November 15, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Brorsson:1989:ESV

[Bro89] Mats Brorsson. Emulation of Shared Virtual Memory on an Experimental Multiprocessor. Technical report, Department of Computer Engineering, Lund University, P.O. Box 118, S-221 00 Lund, Sweden, October 1989.

Brogi:2018:TSB

[BRS18] Antonio Brogi, Luca Rinaldi, and Jacopo Soldani. TosKer: a synergy between TOSCA and Docker for orchestrating multicomponent applications. Software—Practice and Experience, 48(11):2061–2079, November 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bhagavathi:2022:IBS

[BRS^{+}22] Hariharan Bhagavathi, Siva Rathinavelayatham, Kaliraj Shanmugaiah, Kamaraj Kanagaraj, and Dinesh Elangovan. Improved beetle swarm optimization algorithm for energy efficient virtual machine consolidation on cloud environment. Concurrency and Computation: Practice and Experience, 34 (10):e6828:1–e6828:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brunschen:2007:SSE

[Bru07] Christian Brunschen. SMILemu: The SMIL emulator: Version 1.2. Web site, August 30, 2007. URL http:

//en.wikipedia.org/wiki/IAS_computer;   http://en.
wikipedia.org/wiki/SMIL_%28computer%29;   http://
video.ldc.lu.se/smil-50.htm; http://www.smilemu.org/

. SMIL (Siffermaskinen i Lund = Number Machine in Lund) was an early Swedish computer introduced in 1956, and in operation until 1970, that was based on John von Neumann’s 1952 IAS machine that was designed from 1945 to 1951, and was operational until 1958.

Bu:2013:CSC

[BRX13] Xiangping Bu, Jia Rao, and Cheng-Zhong Xu. Coordinated self-configuration of virtual machines and appliances using a model-free learning approach. IEEE Transactions on Parallel and Distributed Systems, 24(4):681–690, April 2013. CODEN ITDSEO. ISSN 1045-9219.

Blelloch:1990:CCO

[BS90] G. E. Blelloch and G. W. Sabot. Compiling collectionoriented languages onto massively parallel computers. Journal of Parallel and Distributed Computing, 8(2):119–134, February 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Burnet:1996:PCP

[BS96] Maxwell M. Burnet and Robert M. Supnik. Preserving computing’s past: restoration and simulation. Digital Technical Journal of Digital Equipment Corporation, 8(3):23–38, December 1996. CODEN DTJOEL.

ISSN 0898-901X. URL ftp://ftp.digital.com/pub/
Digital/info/DTJ/v8n3/Preserving_Computings_Past_
R_07jan1997DTJN02P8.ps;   http://artematrix.org/archive/
computer/restoration.simulation.htm;   http://simh.
trailing-edge.com/docs/dtjn02pf.pdf;   http://www.
digital.com:80/info/DTJN02/DTJN02AH.HTM;   http://
www.digital.com:80/info/DTJN02/DTJN02HM.HTM;   http:
//www.digital.com:80/info/DTJN02/DTJN02P8.PS;   http:
//www.digital.com:80/info/DTJN02/DTJN02PF.PDF;   http:
//www.digital.com:80/info/DTJN02/DTJN02SC.TXT.

Baumann:2019:VSL

[BSD19] Christoph Baumann, Oliver Schwarz, and Mads Dam. On the verification of system-level information flow properties for virtualized execution platforms. Journal of Cryptographic Engineering, 9(3):243–261, September 2019. CO-DEN ???? ISSN 2190-8508 (print), 2190-8516 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s13389-019-00216-4;   https://link.springer.
com/content/pdf/10.1007/s13389-019-00216-4.pdf.

Branco:2015:TFS

[BSI^{+}15] Adriano Branco, Francisco Sant’anna, Roberto Ierusalimschy, Noemi Rodriguez, and Silvana Rossetto. Terra: Flexibility and safety in wireless sensor networks. ACM Transactions on Sensor Networks, 11(4):59:1–59:??, December 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Batalla:2018:VVP

[BSL^{+}18] Jordi Mongay Batalla, Konrad Sienkiewicz, Waldemar Latoszek, Piotr Krawiec, Constandinos X. Mavromoustakis, and George Mastorakis. Validation of virtualization platforms for I-IoT purposes. The Journal of Supercomputing, 74(9):4227–4241, September 2018. CODEN JOSUED.

ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s11227-016-
1844-2.pdf.

Bairavasundaram:2012:RRS

[BSM^{+}12] Lakshmi N. Bairavasundaram, Gokul Soundararajan, Vipul Mathur, Kaladhar Voruganti, and Kiran Srinivasan. Responding rapidly to service level violations using virtual appliances. Operating Systems Review, 46(3):32–40, December 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Burcea:2008:PV

[BSMF08] Ioana Burcea, Stephen Somogyi, Andreas Moshovos, and Babak Falsafi. Predictor virtualization. Operating Systems Review, 42(2):157–167, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bogo:2020:CAO

[BSNB20] Matteo Bogo, Jacopo Soldani, Davide Neri, and Antonio Brogi. Component-aware orchestration of cloud-based enterprise applications, from TOSCA to Docker and Kubernetes. Software—Practice and Experience, 50(9):1793–1821, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Benmakrelouf:2020:ABD

[BSOK^{+}20] Souhila Benmakrelouf, Cédric St-Onge, Nadjia Kara, Hanine Tout, Claes Edstrom, and Yves Lemieux. Abnormal behavior detection using resource level to service level metrics mapping in virtualized systems. Future Generation Computer Systems, 102(??):680–700, January 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X19306600.

Bhargava:2008:ATD

[BSSM08] Ravi Bhargava, Benjamin Serebrin, Francesco Spadini, and Srilatha Manne. Accelerating two-dimensional page walks for virtualized systems. Operating Systems Review, 42(2):26– 35, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bartolini:2014:AFG

[BSSS14] Davide B. Bartolini, Filippo Sironi, Donatella Sciuto, and Marco D. Santambrogio. Automated fine-grained CPU provisioning for virtual machines. ACM Transactions on Architecture and Code Optimization, 11(3):27:1–27:??, October 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bush:1987:CSR

[BSUH87] W. R. Bush, A. D. Samples, D. Ungar, and P. N. Hilfinger. Compiling Smalltalk-80 to a RISC. In Proceedings of the Second International Conference on Architectural Support for Programming Languages and Operating Systems-ASPLOSII, pages 112–116. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 1987.

Bolz:2015:IMT

[BT15] Carl Friedrich Bolz and Laurence Tratt. The impact of meta-tracing on VM design and implementation. Science of Computer Programming, 98 (part 3)(??):408–421, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-

7964 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167642313000269.

Bravo-Torres:2015:VVL

[BTLNBF^{+}15a] Jack F. Bravo-Torres, Martín López-Nores, Yolanda Blanco-Fernández, José J. Pazos-Arias, and Esteban F. Ordóñez-Morales. VaNetLayer: a virtualization layer supporting access to web contents from within vehicular networks. Journal of Computational Science, 11:185–195, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1877750314001008.

Bravo-Torres:2015:IVL

[BTLNBF^{+}15b] Jack Fernando Bravo-Torres, Martín López-Nores, Yolanda Blanco-Fernández, José Juan Pazos-Arias, Manuel Ramos-Cabrer, and Alberto Gil-Solla. An improved virtualization layer to support distribution of multimedia contents in pervasive social applications. Journal of Network and Computer Applications, 51(??):1–17, May 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804515000120.

Basak:2010:VNS

[BTMS10] Debashis Basak, Rohit Toshniwal, Serge Maskalik, and Allwyn Sequeira. Virtualizing networking and security in the cloud. Operating Systems Review, 44(4):86–94, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Burch:2002:LGS

[Bur02] Carl Burch. Logisim: a graphical system for logic circuit design and simulation. ACM Journal on Educational Resources in Computing (JERIC), 2(1):5–16, March 2002. CODEN ???? ISSN 1531-4278.

Butt:1994:RDS

[But94] Farooq Butt. Rapid development of a source-level debugger for PowerPC microprocessors. ACM SIGPLAN Notices, 29 (12):73–77, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Basin:2003:TPH

[BW03] David Basin and Burkhart Wolff, editors. Theorem Proving in Higher Order Logics: 16th International Conference, TPHOLs 2003, Rome, Italy, September 8–12, 2003: Proceedings, volume 2758 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNCSD9. ISBN 3-540-40664-6. ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.9.A96. URL http://link.springer-ny.com/
link/service/series/0558/tocs/t2758.htm;   http://
www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=2758;   http://www.springerlink.
com/openurl.asp?genre=volume&id=doi:10.1007/b11828.

Bila:2015:EOP

[BWD^{+}15] Nilton Bila, Eric J. Wright, Eyal De Lara, Kaustubh Joshi, H. Andrés Lagar-Cavilla, Eunbyung Park, Ashvin Goel, Matti Hiltunen, and Mahadev Satyanarayanan. Energyoriented partial desktop virtual machine migration. ACM Transactions on Computer Systems, 33(1):2:1–2:??, March 2015. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Basu:2019:LYG

[BWH^{+}19] D. Basu, X. Wang, Y. Hong, H. Chen, and S. Bressan. Learn-as-you-go with Megh: Efficient live migration of virtual machines. IEEE Transactions on Parallel and Distributed Systems, 30(8):1786–1801, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bachrach:1985:XVM

[BWP85] Jonathan Bachrach, John Wallerius, and Jehan-François Paris. A XINU virtual machine. In USENIX Association [USE85], pages 348–355. LCCN QA76.8.U65 U8 1985.

Braiki:2020:FLB

[BY20] Khaoula Braiki and Habib Youssef. Fuzzy-logic-based multiobjective best-fit-decreasing virtual machine reallocation. The Journal of Supercomputing, 76(1):427–454, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ben-Yehuda:2016:NPM

[BYBYT16] Muli Ben-Yehuda, Orna Agmon Ben-Yehuda, and Dan Tsafrir. The nom profit-maximizing operating system. ACM SIGPLAN Notices, 51(7):145–160, July 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bao:2020:PPE

[BYZZ20] W. Bao, D. Yuan, B. B. Zhou, and A. Y. Zomaya. Prune and plant: Efficient placement and parallelism of virtual network functions. IEEE Transactions on Computers, 69(6):800–811, 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bruening:2012:TDI

[BZA12] Derek Bruening, Qin Zhao, and Saman Amarasinghe. Transparent dynamic instrumentation. ACM SIGPLAN Notices, 47(7):133–144, July 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Briggs:2017:COI

[BZD17] Kim T. Briggs, Baoguo Zhou, and Gerhard W. Dueck. Cold object identification in the Java Virtual Machine. Software— Practice and Experience, 47(1):79–95, January 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Caamano:2000:PJS

[Caa00] Paul Caamano. Porting a Java Virtual Machine to an embedded system. Thesis (m.s.), Department of Computer Science, University of California, Santa Cruz, Santa Cruz, CA, USA, 2000. viii + 56 pp.

Christodoulakis:1991:OOA

[CAF^{+}91] Stavros Christodoulakis, Natassa Ailamaki, Manolis Fragonikolakis, Yorgos Kapetanakis, and Leonidas Koveos. An object oriented architecture for multimedia information systems. Data Engineering, 14(3):4–15, September 1991.

Callaway:1975:PMT

[Cal75] Peter H. Callaway. Performance measurement tools for VM/ 370. IBM Systems Journal, 14(2):134–160, 1975. CODEN IBMSA7. ISSN 0018-8670.

Cappellari:2021:CBD

[Cap21] Max Cappellari. A cloud-based data collaborative to combat the COVID-19 pandemic and to solve major technology challenges. Future Internet, 13(3):61, February 27, 2021. CODEN

???? ISSN 1999-5903. URL https://www.mdpi.com/1999-
5903/13/3/61.

Carbone:2006:WSH

[Car06] Janique Carbone. Windows Server 2008 Hyper-V resource kit. Microsoft Press, Bellevue, WA, USA, 2006. ISBN 0-7356-2517-4. ???? pp. LCCN QA.

Cardoso:2013:MSC

[Car13] Edvaldo Alessandro Cardoso. Microsoft System Center Virtual Machine Manager 2012 Cookbook. Packt Publishing, Limited, Birmingham, UK, 2013. ISBN 1-84968-632-7 (paperback), 1-84968-633-5 (e-book). iii + 326 pp. LCCN QA76.774.M55. URL http://proquest.tech.

safaribooksonline.de/9781849686327.

Cardoso:2014:SCR

[Car14] Edvaldo Alessandro Cardoso. System Center 2012 R2 virtual machine manager cookbook: over 70 recipes to help you design, configure, and manage a reliable and efficient virtual infrastructure with VMM 2012 R2. Quick answers to common problems. Packt Publishing, Birmingham, UK, second edition, 2014. ISBN 1-78217-684-5, 1-78217-685-3 (ebook). 428 pp. LCCN QA76.76.C69. URL http://proquest.tech.

safaribooksonline.de/9781782176848.

Carrion:2023:KST

[Car23] Carmen Carrión. Kubernetes scheduling: Taxonomy, ongoing issues and challenges. ACM Computing Surveys, 55(7): 138:1–138:??, July 2023. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.acm.org/

doi/10.1145/3539606.

Campanoni:2010:HFP

[CARB10] Simone Campanoni, Giovanni Agosta, Stefano Crespi Reghizzi, and Andrea Di Biagio. A highly flexible, parallel virtual machine: design and experience of ILDJIT. Software—Practice and Experience, 40(2):177–207, February ??, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cavender:1993:APV

[Cav93] Mark Edward Cavender. Asynchronous parallel virtual machine. M.s. thesis, University of Texas at San Antonio. Division of Mathematics and Computer Science and Statistics, San Antonio, TX, USA, 1993. vi + 228 pp.

Crosby:2006:VR

[CB07] Simon Crosby and David Brown. The virtualization reality. ACM Queue: Tomorrow’s Computing Today, 4(10):34– 41, December/January 2006–2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Chowdhury:2010:SNV

[CB10] N. M. Mosharaf Kabir Chowdhury and Raouf Boutaba. A survey of network virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 54(5):862–876, April 8, 2010. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cerling:2009:MMV

[CBER09] Tim Cerling, Jeff Buller, Chuck Enstall, and Richard Ruiz. Mastering Microsoft virtualization. Wiley, New York, NY, USA, 2009. ISBN 0-470-44958-6. 576 (est.) pp. LCCN QA76.9.V5 M36 2009.

Compastie:2020:VSI

[CBFH20] Maxime Compastié, Rémi Badonnel, Olivier Festor, and Ruan He. From virtualization security issues to cloud protection opportunities: an in-depth analysis of system virtualization models. Computers & Security, 97(??):Article 101905, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820301814.

Cao:2012:YYP

[CBGM12] Ting Cao, Stephen M. Blackburn, Tiejun Gao, and Kathryn S. McKinley. The yin and yang of power and performance for asymmetric hardware and managed software. ACM SIGARCH Computer Architecture News, 40(3):225– 236, June 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ISCA ’12 conference proceedings.

Carpio:2022:SMR

[CBJ22] Francisco Carpio, Wolfgang Bziuk, and Admela Jukan. Scaling migrations and replications of Virtual Network Functions based on network traffic forecasting. Computer Networks (Amsterdam, Netherlands: 1999), 203(??):??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-

7069 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1389128621004898.

Chevalier-Boisvert:2012:BSH

[CBLFD12] Maxime Chevalier-Boisvert, Erick Lavoie, Marc Feeley, and Bruno Dufour. Bootstrapping a self-hosted research virtual machine for JavaScript: an experience report. ACM SIG-PLAN Notices, 47(2):61–72, February 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheng:2016:VMN

[CBZ^{+}16] Kun Cheng, Yuebin Bai, Yongwang Zhao, Yao Ma, Duo Lu, Yuanfeng Peng, and Minxuan Zhou. HV^{2}M: a novel approach to boost inter-VM network performance for Xen-based HVMs. The Journal of Systems and Software, 114(??):54– 68, April 2016. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0164121215002782.

Chow:1977:ASP

[CC77] We-Min Chow and W. W. Chiu. An analysis of swapping policies in virtual storage systems. IEEE Transactions on Software Engineering, SE-3(2):150–156, March/April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=1702417.

Cinque:2022:VMC

[CCDR22] Marcello Cinque, Domenico Cotroneo, Luigi De Simone, and Stefano Rosiello. Virtualizing mixed-criticality systems: a survey on industrial trends and issues. Future Generation Computer Systems, 129(??):315–330, April 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X21004787.

Chen:2017:MLF

[CCL^{+}17] Xian Chen, Wenzhi Chen, Zhongyong Lu, Yu Zhang, Rui Chang, Mohammad Mehedi Hassan, Abdulhameed Alelaiwi, and Yang Xiang. MBSA: a lightweight and flexible storage architecture for virtual machines. Concurrency and Computation: Practice and Experience, 29(16), August 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2020:SSV

[CCL^{+}20] Yunliang Chen, Xiaodao Chen, Wangyang Liu, Yuchen Zhou, Albert Y. Zomaya, Rajiv Ranjan, and Shiyan Hu. Stochastic scheduling for variation-aware virtual machine placement in a cloud computing CPS. Future Generation Computer Systems, 105(??):779–788, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X17320101.

Carbone:2012:SRM

[CCML12] Martim Carbone, Matthew Conover, Bruce Montague, and Wenke Lee. Secure and robust monitoring of virtual machines through guest-assisted introspection. Lecture Notes in Computer Science, 7462:22–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/chapter/10.1007/978-
3-642-33338-5_2/.

Chen:2007:DGS

[CCMY07] Haibo Chen, Jieyun Chen, Wenbo Mao, and Fei Yan. Daonity — grid security from two levels of virtualization. Information Security Technical Report, 12(3):123–138, ???? 2007. CODEN ISTRFR. ISSN 1363-4127 (print), 1873-605X (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1363412707000258.

Childs:2005:SCG

[CCO^{+}05] Stephen Childs, Brian Coghlan, David O’Callaghan, Geoff Quigley, and John Walsh. A single-computer Grid gateway using virtual machines. In Shih and Shibata [SS05], pages 310–315. ISBN 0-7695-2249-1 (paperback). ISSN 1550-445X. LCCN TK5105.5 .I5616 2005. URL http://ieeexplore.

ieee.org/servlet/opac?punumber=9746. IEEE Computer

Society Order Number P2249.

Cheng:2020:SVC

[CCW^{+}20] Yuxia Cheng, Wenzhi Chen, Zonghui Wang, Zhongxian Tang, and Yang Xiang. Smart VM co-scheduling with the precise prediction of performance characteristics. Future Generation Computer Systems, 105(??):1016–1027, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X16306616.

Chiueh:2014:SFI

[cCWS14] Tzi cker Chiueh, Xin Wang, and Zhiyong Shan. Shuttle: Facilitating inter-application interactions for OS-level virtualization. IEEE Transactions on Computers, 63(5):1220–1233, May 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Calder:2005:EVM

[CCWY05] Brad Calder, Andrew A. Chien, Ju Wang, and Don Yang. The Entropia Virtual Machine for desktop grids. In ACM [ACM05d], pages 186–185. ISBN 1-59593-047-7. LCCN

QA76.9.V5 I575 2005. URL http://www.loc.gov/catdir/
toc/fy0611/2006530661.html. ACM order number 548059.

Chen:2006:LUO

[CCZ^{+}06] Haibo Chen, Rong Chen, Fengzhe Zhang, Binyu Zang, and Pen-Chung Yew. Live updating operating systems using virtualization. In ACM [ACM06f], pages 35–44. ISBN 1-59593-332-6 (??invalid ISBN??). LCCN QA76.9.V4.

Czajkowski:2001:MCV

[CD01] Grzegorz Czajkowski and Laurent Daynés. Multitasking without compromise: a virtual machine evolution. ACM SIG-PLAN Notices, 36(11):125–138, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Cheng:2012:VBP

[CD12] Yueqiang Cheng and Xuhua Ding. Virtualization based password protection against malware in untrusted operating systems. Lecture Notes in Computer Science, 7344:201–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-30921-2_12/.

Cao:2014:EAH

[CD14] Zhibo Cao and Shoubin Dong. An energy-aware heuristic framework for virtual machine consolidation in Cloud computing. The Journal of Supercomputing, 69(1):429–451, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL http://link.springer.com/article/
10.1007/s11227-014-1172-3.

Cheng:2013:DVB

[CDD13] Yueqiang Cheng, Xuhua Ding, and Robert H. Deng. Driver-Guard: Virtualization-based fine-grained protection on I/O flows. ACM Transactions on Information and System Security, 16(2):6:1–6:??, September 2013. CODEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Comar:1997:TGJ

[CDG97] C. Comar, G. Dismukes, and F. Gasperoni. Targeting GNAT to the Java Virtual Machine. In ACM [ACM97], pages 149– 164. ISBN 0-89791-981-5. LCCN QA 76.73 A35 T75 1997. Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Chafi:2010:LVH

[CDM^{+}10] Hassan Chafi, Zach DeVito, Adriaan Moors, Tiark Rompf, Arvind K. Sujeeth, Pat Hanrahan, Martin Odersky, and Kunle Olukotun. Language virtualization for heterogeneous parallel computing. ACM SIGPLAN Notices, 45(10):835– 847, October 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coppolino:2025:WLW

[CDM^{+}25] Luigi Coppolino, Salvatore D’Antonio, Giovanni Mazzeo, Roberto Nardone, Luigi Romano, and Mathieu Schmitt. WASMBOX: a lightweight Wasm-based runtime for trustworthy multi-tenant embedded systems. IEEE Transactions on Emerging Topics in Computing, 13(2):467–480, 2025. ISSN 2168-6750 (print), 2376-4562 (electronic).

Czajkowski:2002:CSA

[CDN02] Grzegorz Czajkowski, Laurent Daynès, and Nathaniel Nystrom. Code sharing among virtual machines. Lecture Notes in Computer Science, 2374:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer-ny.com/link/service/series/0558/bibs/
2374/23740155.htm;   http://link.springer-ny.com/
link/service/series/0558/papers/2374/23740155.pdf.

Cinque:2024:TIA

[CDO24] Marcello Cinque, Luigi De Simone, and Daniele Ottaviano. Temporal isolation assessment in virtualized safety-critical mixed-criticality systems: a case study on Xen hypervisor. The Journal of Systems and Software, 216(??):??, October 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-

1228 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0164121224001924.

Cloosters:2024:MCB

[CDW^{+}24] Tobias Cloosters, Oussama Draissi, Johannes Willbold, Thorsten Holz, and Lucas Davi. Memory corruption at the border of trusted execution. IEEE Security & Privacy, 22(4): 87–96, July/August 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Casey:2007:OIB

[CEG07] Kevin Casey, M. Anton Ertl, and David Gregg. Optimizing indirect branch prediction accuracy in virtual machine interpreters. ACM Transactions on Programming Languages and Systems, 29(6):37:1–37:36, October 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Catena:2022:DLB

[CEPR22] Tiziana Catena, Vincenzo Eramo, Massimo Panella, and Antonello Rosato. Distributed LSTM-based cloud resource allocation in Network Function Virtualization Architectures. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1389128622002390.

Chelius:2000:ING

[CF00] Guillaume Chelius and Éric Fleury. An IP next generation compliant Java^{T M} virtual machine. Lecture Notes in Computer Science, 1800:528–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/series/0558/bibs/
1800/18000528.htm;   http://link.springer-ny.com/
link/service/series/0558/papers/1800/18000528.pdf.

Cerveira:2024:RV

[CFB24] Frederico Cerveira, António Howcroft Ferreira, and Raul Barbosa. Resilient virtualization. Computer, 57(2):70–78, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carvalho:2024:SDD

[CFC^{+}24] Fabrício B. Carvalho, Ronaldo A. Ferreira, Ítalo Cunha, Marcos A. M. Vieira, and Murali K. Ramanathan. State disaggregation for dynamic scaling of network functions. IEEE/ ACM Transactions on Networking, 32(1):81–95, 2024. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.acm.org/doi/10.1109/TNET.
2023.3282562.

Chang:2013:IVP

[CFG^{+}13] Xiaotao Chang, Hubertus Franke, Yi Ge, Tao Liu, Kun Wang, Jimi Xenidis, Fei Chen, and Yu Zhang. Improving virtualization in the presence of software managed translation lookaside buffers. ACM SIGARCH Computer Architecture News, 41(3):120–129, June 2013. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ICSA ’13 conference proceedings.

Canon:1979:VME

[CFH^{+}79] M. D. Canon, D. H. Fritz, John H. Howard, T. D. Howell, Michael F. Mitoma, and Juan Rodriguez-Rossel. A virtual machine emulator for performance evaluation (summary). In Proceedings of the 7th ACM Symposium on Operating Systems Principles (SOSP), Operating Systems Review, pages 1–?? ACM Press, New York, NY 10036, USA, 1979. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Canon:1980:VME

[CFH^{+}80] M. D. Canon, D. H. Fritz, John H. Howard, T. D. Howell, Michael F. Mitoma, and Juan Rodriguez-Rossel. A virtual machine emulator for performance evaluation. Communications of the ACM, 23(2):71–80, 1980. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chen:2019:PPF

[CFLL19] L. Chen, Y. Feng, B. Li, and B. Li. Promenade: Proportionally fair multipath rate control in datacenter networks with random network coding. IEEE Transactions on Parallel and Distributed Systems, 30(11):2536–2546, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cao:2017:VNM

[CFM17] Yang Cao, Wenfei Fan, and Shuai Ma. Virtual network mapping in cloud computing: a graph pattern matching approach. The Computer Journal, 60(3):60–??, March 2017. CODEN CMPJA6. ISSN ???? URL https:/

/academic.oup.com/comjnl/article/60/3/287/2608063/
Virtual-Network-Mapping-in-Cloud-Computing-A-Graph.

Cocana-Fernandez:2019:IEE

[CFRSSR19] Alberto Cocaña-Fernández, Julio Rodríguez-Soares, Luciano Sánchez, and José Ranilla. Improving the energy efficiency of virtual data centers in an IT service provider through proactive fuzzy rules-based multicriteria decision making. The Journal of Supercomputing, 75(3):1078–1093, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cheriton:2012:HAS

[CFS^{+}12] David Cheriton, Amin Firoozshahian, Alex Solomatnikov, John P. Stevenson, and Omid Azizi. HICAMP: architectural support for efficient concurrency-safe shared structured data access. ACM SIGARCH Computer Architecture News, 40(1): 287–300, March 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ASPLOS ’12 conference proceedings.

Celesti:2012:VMP

[CFVP12] Antonio Celesti, Maria Fazio, Massimo Villari, and Antonio Puliafito. Virtual machine provisioning through satellite communications in federated Cloud environments. Future Generation Computer Systems, 28(1):85–93, January 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X11001038.

Chen:2016:OVM

[CGC16] Tao Chen, Xiaofeng Gao, and Guihai Chen. Optimized virtual machine placement with traffic-aware balancing in data center networks. Scientific Programming, 2016(??): 3101658:1–3101658:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https://

www.hindawi.com/journals/sp/2016/3101658/.

Chen:2024:CEH

[CGC^{+}24] Bochuan Chen, Xiao Guo, Yuting Chen, Xiaofeng Yu, and Lei Bu. Constructing exception handling chains for testing Java virtual machine implementations. Journal of Software: Evolution and Process, 36(4):e2562:1–e2562:??, April 2024. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Chen:2008:OVBa

[CGL^{+}08a] Xiaoxin Chen, Tal Garfinkel, E. Christopher Lewis, Pratap Subrahmanyam, Carl A. Waldspurger, Dan Boneh, Jeffrey Dwoskin, and Dan R. K. Ports. Overshadow: a virtualizationbased approach to retrofitting protection in commodity operating systems. ACM SIGARCH Computer Architecture News, 36(1):2–13, March 2008. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Chen:2008:OVBb

[CGL^{+}08b] Xiaoxin Chen, Tal Garfinkel, E. Christopher Lewis, Pratap Subrahmanyam, Carl A. Waldspurger, Dan Boneh, Jeffrey Dwoskin, and Dan R. K. Ports. Overshadow: a virtualizationbased approach to retrofitting protection in commodity operating systems. Operating Systems Review, 42(2):2–13, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:2008:OVBc

[CGL^{+}08c] Xiaoxin Chen, Tal Garfinkel, E. Christopher Lewis, Pratap Subrahmanyam, Carl A. Waldspurger, Dan Boneh, Jeffrey Dwoskin, and Dan R. K. Ports. Overshadow: a virtualizationbased approach to retrofitting protection in commodity operating systems. ACM SIGPLAN Notices, 43(3):2–13, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chari:2017:BEH

[CGM17] Guido Chari, Diego Garbervetsky, and Stefan Marr. Building efficient and highly run-time adaptable virtual machines. ACM SIGPLAN Notices, 52(2):60–71, February 2017. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chari:2019:FRE

[CGMD19] G. Chari, D. Garbervetsky, S. Marr, and S. Ducasse. Fully reflective execution environments: Virtual machines for more flexible software. IEEE Transactions on Software Engineering, 45(9):858–876, September 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Casazza:2006:RSP

[CGS06] Jeffrey P. Casazza, Michael Greenfield, and Kan Shi. Redefining server performance characterization for virtualization benchmarking. Intel Technology Journal, 10(3):243– 251, August 10, 2006. ISSN 1535-766X. URL http:

//developer.intel.com/technology/itj/2006/v10i3/7-
benchmarking/1-abstract.htm.

Courbot:2010:EBD

[CGV10] Alexandre Courbot, Gilles Grimaud, and Jean-Jacques Vandewalle. Efficient off-board deployment and customization of virtual machine-based embedded systems. ACM Transactions on Embedded Computing Systems, 9(3):21:1–21:??, February 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Crosby:2007:VXI

[CGW07] Simon Crosby, Juan R. Garcia, and David E. Williams, editors. Virtualization with Xen: including XenEnterprise, XenServer, and XenExpress. Syngress Publishing, Inc., Rockland, MA, USA, 2007. ISBN 0-08-055393-1 (electronic), 1-59749-167-5. xx + 364 pp. LCCN QA76.9.V5 V57 2007eb.

Cremers:1978:FMV

[CH78] Armin B. Cremers and Thomas N. Hibbard. Formal modeling of virtual machines. IEEE Transactions on Software Engineering, SE-4(5):426–436, September 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Choi:2008:SHM

[CH08] Yoonseo Choi and Hwansoo Han. Shared heap management for memory-limited Java virtual machines. ACM Transactions on Embedded Computing Systems, 7(2):13:1–13:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chang:2007:VMS

[CHCC07] Da-Wei Chang, Cheng-En Hsieh, Yan-Pai Chen, and Kwo-Cheng Chiu. Virtual machine support for zero-loss Internet service recovery and upgrade. Software—Practice and Experience, 37(13):1349–1376, November 10, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chetty:2021:VNF

[Che21] Swarna Bindu Chetty. Virtual network function embedding under nodal outage using deep Q-learning. Future Internet, 13(3):82, March 23, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.com/1999-5903/13/3/82.

Chisnall:2008:DGX

[Chi08] David Chisnall. The Definitive Guide to the Xen Hypervisor. Prentice Hall open source software development series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2008. ISBN 0-13-234971-X. xx + 286 pp. LCCN QA76.9.V5 C427 2008.

Cui:2018:SPA

[CHLY18] Lei Cui, Zhiyu Hao, Lun Li, and Xiaochun Yun. SnapFiner: A page-aware snapshot system for virtual machines. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2613–2626, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:

//www.computer.org/csdl/trans/td/2018/11/08352722-
abs.html.

Cui:2017:PFE

[CHPY17] Lei Cui, Zhiyu Hao, Yaqiong Peng, and Xiaochun Yun. Piccolo: A fast and efficient rollback system for virtual machine clusters. IEEE Transactions on Parallel and Distributed Systems, 28(8):2328–2341, August 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2017/08/07852513-abs.html.

Chubb:2006:VUL

[Chu06] Peter Chubb. Virtualization and user-level drivers. In

Anonymous [Ano06a], page ??  ISBN ????  LCCN

???? URL http://www.ice.gelato.org/oct06/pres_pdf/
gelato_ICE06oct_vm_chubb_unsw.pdf.

Chen:2012:FGP

[CHW12] Jianhai Chen, Dawei Huang, and Bei Wang. A fine-grained performance-based decision model for virtualization application solution. Lecture Notes in Computer Science, 7144: 180–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-32627-1_13/.

Ciabrini:2007:SVS

[Cia07] Damien Ciabrini. Stack virtualization for source level debugging. Software—Practice and Experience, 37(7):693–725, June 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Choi:2022:EDP

[CJJ^{+}22] Yeseul Choi, Yunjong Jeong, Daehee Jang, Brent Byunghoon Kang, and Hojoon Lee. EmuID: Detecting presence of emulation through microarchitectural characteristic on ARM. Computers & Security, 113(??):Article 102569, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482100393X.

Chandak:2012:DLB

[CJK^{+}12] Akshay Chandak, Krishnakant Jaju, Akshay Kanfade, Amit Joshi, and Pushkar Lohiya. Dynamic load balancing of virtual machines using QEMU-KVM. International Journal of Computer Applications, 46(??):10–14, May 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume46/number6/6910-9263/.

Carr:1987:EUC

[CK87] Harold Carr and Robert R. Kessler. Emulator for Utah Common Lisp’s Abstract Virtual Register Machine. Journal of FORTH Application and Research, 5(1), 1987. CODEN JFAREL. ISSN 0738-2022.

Campbell-Kelly:1996:ES

[CK96] Martin Campbell-Kelly. The EDSAC simulator. Web

site, 1996. URL http://www.dcs.warwick.ac.uk/~edsac/

. Downloadable simulator software available for Microsoft Windows and GNU/Linux x86.

Chryselius:2006:DQE

[CK06a] Toralf Chryselius and Andrea Kuntz. Debian unter Qemu Einführung in das Betriebssystem Debian Linux in der virtuellen Umgebung Qemu unter Windows. (German) [Debian under Qemu: Introduction in the Debian Linux operating systems in the Qemu virtual machine under Windows], volume 17 of Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-116-3 (book), 3-86768-716-1 (DVD). 159 pp. LCCN ????

Chryselius:2006:IDQ

[CK06b] Toralf Chryselius and Andrea Kuntz. Internetkommunikation in Debian unter Qemu Einführung in das Betriebssystem Debian Linux in Qemu und Vorstellung der wichtigsten Internetprogramme. (German) [Internet Communication in Debian under Qemu: Introduction in the Debian Linux operating system in Qemu and creation of the most important Internet programs], volume 18 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-117-1 (book), 3-86768-717-X (DVD). 109 pp. LCCN ????

Chryselius:2006:IKQb

[CK06c] Toralf Chryselius and Andrea Kuntz. Internetkommunikation in Kanotix unter Qemu Einführung in das Betriebssystem Kanotix und Vorstellung von Internetprogrammen in der virtuellen Umgebung Qemu, volume 34 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-133-3 (book), 3-86768-733-1 (DVD). 114 pp. LCCN ????

Chryselius:2006:IKQc

[CK06d] Toralf Chryselius and Andrea Kuntz. Internetkommunikation in Knoppix unter Qemu Einführung in das Betriebssystem Knoppix und Vorstellung von Internetprogrammen in der virtuellen Umgebung Qemu, volume 50 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-149-X (book). 121 pp. LCCN ????

Chryselius:2006:IKQa

[CK06e] Toralf Chryselius and Andrea Kuntz. Internetkommunikation in Kubuntu unter Qemu Einführung in das Betriebssystem Kubuntu und Vorstellung von Internetprogrammen in der virtuellen Umgebung Qemu. (German) [Internet Communication in Kubuntu under Qemu: Introduction to the Kubuntu operating system and creation of Internet programs in the Qemu virtual machine], volume 6 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-105-8 (Buch), 3-86768-705-6 (DVD). 107 pp. LCCN ????

Chryselius:2006:IOV

[CK06f] Toralf Chryselius and Andrea Kuntz. Internetkommunikation in OpenSUSE unter VMware [Qemu] Einführung in das Betriebssystem OpenSUSE Linux und Vorstellung von Internetprogrammen in der virtuellen Umgebung VMware, volume 66 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-165-1 (book). 117 pp. LCCN ????

Chryselius:2006:IOQ

[CK06g] Toralf Chryselius and Andrea Kuntz. Internetkommunikation mit OpenSUSE unter Qemu: Einführung in das Betriebssystem OpenSUSE Linux und Vorstellung von Internetprogrammen in der virtuellen Umgebung Qemu, volume 66 of Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-165-1. 104 pp. LCCN ????

Chryselius:2006:KLQb

[CK06h] Toralf Chryselius and Andrea Kuntz. Kanotix Linux unter Qemu Einführung in das Betriebssystem Kanotix Linux in der virtuellen Umgebung Qemu unter Windows, volume 33 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-132-5 (book), 3-86768-732-3 (DVD). 156 pp. LCCN ????

Chryselius:2006:KLQa

[CK06i] Toralf Chryselius and Andrea Kuntz. Knoppix Linux unter Qemu Einführung in das Betriebssystem Knoppix Linux in der virtuellen Umgebung Qemu unter Windows, volume 49 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-148-1 (book), 3-86768-748-X (DVD). 142 pp. LCCN ????

Chryselius:2006:KQE

[CK06j] Toralf Chryselius and Andrea Kuntz. Kubuntu unter Qemu Einführung in das Betriebssystem Kubuntu Linux in der virtuellen Umgebung Qemu, volume 5 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-104-X (book), 3-86768-704-8 (DVD). 158 pp. LCCN ????

Chryselius:2006:LDQ

[CK06k] Toralf Chryselius and Andrea Kuntz. Lernprogramme in Debian unter Qemu Einführung in das Betriebssystem Debian und Vorstellung von Lernprogrammen in der virtuellen Umgebung Qemu, volume 19 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-118-X (book), 3-86768-718-8 (DVD). 141 pp. LCCN ????

Chryselius:2006:LKQb

[CK06l] Toralf Chryselius and Andrea Kuntz. Lernprogramme in Kanotix unter Qemu Einführung in das Betriebssystem Kanotix und Vorstellung von Lernprogrammen in der virtuellen Umgebung Qemu, volume 35 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-134-1 (book), 3-86768-734-X (DVD). 151 pp. LCCN ????

Chryselius:2006:LKQa

[CK06m] Toralf Chryselius and Andrea Kuntz. Lernprogramme in Knoppix unter Qemu Einführung in das Betriebssystem Knoppix und Vorstellung von Lernprogrammen in der virtuellen Umgebung Qemu, volume 51 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-150-3 (book), 3-86768-750-1 (DVD). 145 pp. LCCN ????

Chryselius:2006:LKQc

[CK06n] Toralf Chryselius and Andrea Kuntz. Lernprogramme in Kubuntu unter Qemu Einführung in das Betriebssystem Kubuntu und Vorstellung von Lernprogrammen in der virtuellen Umgebung Qemu, volume 7 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-106-6 (book), 3-86768-706-4 (DVD). 152 pp. LCCN ????

Chryselius:2006:LOL

[CK06o] Toralf Chryselius and Andrea Kuntz. Lernprogramme mit OpenSUSE Linux unter Qemu: Einführung in das Betriebssystem, OpenSUSE Linux und Vorstellung von Lernprogrammen in der virtuellen Umgebung Quemu, volume 63 of Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-166-X, 3-86768-766-8. 147 pp. LCCN ????

Chryselius:2006:OLQ

[CK06p] Toralf Chryselius and Andrea Kuntz. OpenSuSE Linux unter Qemu Einführung in das Betriebssystem OpenSUSE Linux in der virtuellen Umgebng Qemu unter Windows, volume 65 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-164-3 (book). 168 pp. LCCN ????

Chryselius:2006:SKD

[CK06q] Toralf Chryselius and Andrea Kuntz. Software für Kinder in Debian unter Qemu Einführung in das Betriebssystem Debian und Vorstellung der Lern- und Spielesammlung Gcompris in der virtuellen Umgebung Qemu, volume 20 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-119-8 (book), 3-86768-719-6 (DVD). 113 pp. LCCN ????

Chryselius:2006:SKKb

[CK06r] Toralf Chryselius and Andrea Kuntz. Software für Kinder in Kanotix unter Qemu Einführung in das Betriebssystem Kanotix und Vorstellung der Lern- und Spielesammlung GCompris in der virtuellen Umgebung Qemu, volume 36 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-135-X (book), 3-86768-735-8 (DVD). 113 pp. LCCN ????

Chryselius:2006:SKKc

[CK06s] Toralf Chryselius and Andrea Kuntz. Software für Kinder in Knoppix unter Qemu Einführung in das Betriebssystem Knoppix und Vorstellung der Lern- und Spielesammlung GCompris in der virtuellen Umgebung Qemu, volume 52 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-151-1 (book). 120 pp. LCCN ????

Chryselius:2006:SKKa

[CK06t] Toralf Chryselius and Andrea Kuntz. Software für Kinder in Kubuntu unter Qemu Einführung in das Betriebssystem Kubuntu und Vorstellung der Lern- und Spielesammlung GCompris in der virtuellen Umgebung Qemu, volume 8 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-107-4 (book), 3-86768-707-2 (DVD). 108 pp. LCCN ????

Comaa:1978:SGP

[CKP78] H. Comaa, J. Kramer, and B. K. Penney. A student group project in operating system implementation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):197–202, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Culler:1993:LTR

[CKP^{+}93] David E. Culler, Richard M. Karp, David A. Patterson, Abhijit Sahay, Klaus E. Schauser, Eunice Santos, Ramesh Subramonian, and Thorsten von Eicken. LogP: towards a realistic model of parallel computation. ACM SIGPLAN Notices, 28(7):1–12, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chamanara:2017:QSH

[CKRJ17] Javad Chamanara, Birgitta König-Ries, and H. V. Jagadish. QUIS: in-situ heterogeneous data source querying. Proceedings of the VLDB Endowment, 10(12):1877–1880, August 2017. CODEN ???? ISSN 2150-8097.

Conte:2008:NHA

[CKT08] Alberto Conte, Sylvaine Kerboeuf, and Laurent Thomas. Network-hosted avatar: User-terminal virtualization in the network. Bell Labs Technical Journal, 13(2):117–126, Summer 2008. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Canali:2014:DSV

[CL14] Claudia Canali and Riccardo Lancellotti. Detecting similarities in virtual machine behavior for cloud monitoring using smoothed histograms. Journal of Parallel and Distributed Computing, 74(8):2757–2769, August 2014. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514000343.

Checco:2015:FVN

[CL15] Alessandro Checco and Douglas J. Leith. Fair virtualization of 802.11 networks. IEEE/ACM Transactions on Networking, 23(1):148–160, February 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Cheng:2016:OIL

[CL16a] Luwei Cheng and Francis C. M. Lau. Offloading interrupt load balancing from SMP virtual machines to the hypervisor. IEEE Transactions on Parallel and Distributed Systems, 27(11):3298–3310, November 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2016/11/07425234-abs.html.

Cheng:2016:RTC

[CL16b] Luwei Cheng and Francis C. M. Lau. Revisiting TCP congestion control in a virtual cluster environment. IEEE/ACM Transactions on Networking, 24(4):2154–2167, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Canali:2017:ICP

[CL17a] Claudia Canali and Riccardo Lancellotti. Identifying communication patterns between virtual machines in softwaredefined data centers. ACM SIGMETRICS Performance Evaluation Review, 44(4):49–56, March 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Canali:2017:SAV

[CL17b] Claudia Canali and Riccardo Lancellotti. Scalable and automatic virtual machines placement based on behavioral similarities. Computing: Archiv für Informatik und Numerik, 99 (6):575–595, June 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cladingboel:1997:RJV

[Cla97] Christopher Cladingboel. Real Java Virtual Machines: Hardware compilation and the Java Virtual Machine. Thesis (M.Sc.), Board of the Faculty of Mathematical Sciences, Oxford University, Oxford, UK, 1997. 107 pp.

Clark:2005:SVT

[Cla05] Tom Clark. Storage virtualization: technologies for simplifying data storage and management. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-26251-4. xvii + 234 pp. LCCN QA76.9.V5 C55 2005.

Chiang:2013:IBM

[CLcC13] Jui-Hao Chiang, Han-Lin Li, and Tzi cker Chiueh. Introspectionbased memory de-duplication and migration. ACM SIG-PLAN Notices, 48(7):51–62, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Criswell:2007:SVA

[CLDA07] John Criswell, Andrew Lenharth, Dinakar Dhurjati, and Vikram Adve. Secure virtual architecture: a safe execution environment for commodity operating systems. Operating Systems Review, 41(6):351–366, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chow:2010:MSR

[CLG^{+}10] Jim Chow, Dominic Lucchetti, Tal Garfinkel, Geoffrey Lefebvre, Ryan Gardner, Joshua Mason, Sam Small, and Peter M. Chen. Multi-stage replay with Crosscut. ACM SIGPLAN Notices, 45(7):13–24, July 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cardoso-Llach:2021:AIE

[CLKEF21] Daniel Cardoso-Llach, Eric Kaltman, Emek Erdolu, and Zachary Furste. An archive of interfaces: Exploring the potential of emulation for software research, pedagogy, and design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW2):294:1–294:22, October 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https://

dl.acm.org/doi/10.1145/3476035.

Cui:2013:VMV

[CLL^{+}13] Lei Cui, Jianxin Li, Bo Li, Jinpeng Huai, Chunming Hu, Tianyu Wo, Hussain Al-Aqrabi, and Lu Liu. VMScatter: migrate virtual machines to many hosts. ACM SIGPLAN Notices, 48(7):63–72, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Chen:2023:PTA

[CLL^{+}23] Rui Chen, Bo Liu, WeiWei Lin, JianPeng Lin, HuiWen Cheng, and KeQin Li. Power and thermal-aware virtual machine scheduling optimization in cloud data center. Future Generation Computer Systems, 145(??):578–589, August 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X23001346.

Chrobot:2012:DMV

[CLŁS12] Arkadiusz Chrobot, Maciej Lasota, Grzegorz Łukawski, and Krzysztof Sapiecha. Distributed memory virtualization with the use of SDDSfL. Lecture Notes in Computer Science, 7204: 141–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-31500-8_15/.

Chen:2022:DUL

[CLM^{+}22] Jiahao Chen, Dingji Li, Zeyu Mi, Yuxuan Liu, Binyu Zang, Haibing Guan, and Haibo Chen. DuVisor: A user-level hypervisor through delegated virtualization,. arXiv.org, ??(??):1–

17, January 22, 2022. URL https://arxiv.org/abs/2201.
09652.

Claudino:2024:PQC

[CLM24] Daniel Claudino, Dmitry I. Lyakh, and Alexander J. Mc-Caskey. Parallel quantum computing simulations via quantum accelerator platform virtualization. Future Generation Computer Systems, 160(??):264–273, November 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X24003054.

Clocksin:1985:DSS

[Clo85] W. F. Clocksin. Design and simulation of a sequential PRO-LOG machine. New Generation Comput., 3(1):101–20, 1985.

Carpenter:2007:HVA

[CLS07] Matthew Carpenter, Tom Liston, and Ed Skoudis. Hiding virtualization from attackers and malware. IEEE Security & Privacy, 5(3):62–65, May/June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chen:2014:HBA

[CLW^{+}14] Songqing Chen, Lei Liu, Xinyuan Wang, Xinwen Zhang, and Zhao Zhang. A host-based approach for unknown fastspreading worm detection and containment. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(4): 21:1–21:??, January 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Condoluci:2018:SVM

[CM18] Massimo Condoluci and Toktam Mahmoodi. Softwarization and virtualization in 5G mobile networks: Benefits, trends and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 146(??):65–84, December 9, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128618302500.

Celesti:2019:SCV

[CMG^{+}19] Antonio Celesti, Davide Mulfari, Antonino Galletta, Maria Fazio, Lorenzo Carnevale, and Massimo Villari. A study on container virtualization for guarantee quality of service in Cloud-of-Things. Future Generation Computer Systems, 99(??):356–364, October 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X18325615.

Cecilia:2023:URG

[CMGI^{+}23] José M. Cecilia, Juan Morales-García, Baldomero Imbernón, Javier Prades, Juan-Carlos Cano, and Federico Silla. Using remote GPU virtualization techniques to enhance edge computing devices. Future Generation Computer Systems, 142(??):14–24, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X22004423.

Chen:2016:SDN

[CMK^{+}16] Jiann-Liang Chen, Yi-Wei Ma, Hung-Yi Kuo, Chu-Sing Yang, and Wen-Chien Hung. Software-defined network virtualization platform for enterprise network resource management. IEEE Transactions on Emerging Topics in Computing, 4(2):179–186, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chung:2006:TTMa

[CMM^{+}06a] JaeWoong Chung, Chi Cao Minh, Austen McDonald, Travis Skare, Hassan Chafi, Brian D. Carlstrom, Christos Kozyrakis, and Kunle Olukotun. Tradeoffs in transactional memory virtualization. ACM SIGARCH Computer Architecture News, 34(5):371–381, December 2006. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Chung:2006:TTMb

[CMM^{+}06b] JaeWoong Chung, Chi Cao Minh, Austen McDonald, Travis Skare, Hassan Chafi, Brian D. Carlstrom, Christos Kozyrakis, and Kunle Olukotun. Tradeoffs in transactional memory virtualization. Operating Systems Review, 40(5):371–381, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chung:2006:TTMc

[CMM^{+}06c] JaeWoong Chung, Chi Cao Minh, Austen McDonald, Travis Skare, Hassan Chafi, Brian D. Carlstrom, Christos Kozyrakis, and Kunle Olukotun. Tradeoffs in transactional memory virtualization. ACM SIGPLAN Notices, 41(11):371–381, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Contreras:2007:XPP

[CMP^{+}07] Gilberto Contreras, Margaret Martonosi, Jinzhang Peng, Guei-Yuan Lueh, and Roy Ju. The XTREM power and performance simulator for the Intel XScale core: Design and experiences. ACM Transactions on Embedded Computing Systems, 6(1):4:1–4:??, February 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2013:TVR

[CMP^{+}13] Chen Chen, Petros Maniatis, Adrian Perrig, Amit Vasudevan, and Vyas Sekar. Towards verifiable resource accounting for outsourced computation. ACM SIGPLAN Notices, 48 (7):167–178, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Coffing:1999:XPM

[Cof99] Charles L. Coffing. An x86 protected mode virtual machine monitor for the MIT exokernel. Thesis (S.B. and M.Eng.), Department of Electrical Engineering and Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, 1999. 109 pp. Supervised by M. Frans Kaashoek.

Cohen:1997:DJV

[Coh97] R. Cohen. The defensive Java Virtual Machine specification. Technical report, Computational Logic Inc., Austin, TX, USA, 1997.

Cohen:2010:VS

[Coh10] Fred Cohen. The virtualization solution. IEEE Security & Privacy, 8(3):60–63, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Comeau:1965:PLS

[Com65] Leslie W. Comeau. The philosophy and logical structure of the control program. Memorandum 2, IBM Cambridge SR&D Center Research Time-Sharing Computer, Cambridge, MA, USA, April 15, 1965. ???? pp.

Comeau:1982:COV

[Com82] Leslie W. Comeau. CP-40, the origin of VM/370. In ????, editor, Proceedings of SEAS AM82, September, 1982. ????, ????, 1982.

Compton:2000:VLB

[Com00] Jason Compton. VMware 2 for Linux: [a better way to run multiple operating systems on Linux]. Prima Tech, Rocklin, CA, USA, 2000. ISBN 0-7615-2764-8. xxii + 406 pp. LCCN QA76.76.O63 C656 2000.

Compton:2003:VL

[Com03] Jason Compton. VMware 2 for Linux. Premier Press, Rocklin, CA, USA, 2003. ISBN 0-7615-2764-8. xxii + 406 pp. LCCN QA76.76.O63 C656 2000. US$39.99. Includes CD-ROM.

Cheng:2024:ITD

[COV^{+}24] Pau-Chen Cheng, Wojciech Ozga, Enriquillo Valdez, Salman Ahmed, Zhongshu Gu, Hani Jamjoom, Hubertus Franke, and James Bottomley. Intel TDX demystified: a top-down approach. ACM Computing Surveys, 56(9):238:1–238:??, September 2024. CODEN CMSVAN. ISSN 0360-0300 (print),

1557-7341 (electronic). URL https://dl.acm.org/doi/10.
1145/3652597.

Cox:2007:REM

[Cox07] Russ Cox. Regular expression matching can be simple and fast. Report, swtch.com, Cambridge, MA, USA,

January 2007. URL http://swtch.com/~rsc/regexp/

regexp1.html. See also [Tho68, KP99, Cox09, Cox10, Cox12].

Cox:2009:REM

[Cox09] Russ Cox. Regular expression matching: the virtual machine approach. Report, swtch.com, Cambridge, MA, USA, Decem-

ber 2009. URL http://swtch.com/~rsc/regexp/regexp2.

html. See also [Tho68, KP99, Cox07, Cox10, Cox12].

Cox:2010:REM

[Cox10] Russ Cox. Regular expression matching in the wild. Report, swtch.com, Cambridge, MA, USA, March 2010. URL http:/

/swtch.com/~rsc/regexp/regexp3.html. See also [Tho68,

KP99, Cox07, Cox09, Cox12].

Cox:2012:REM

[Cox12] Russ Cox. Regular expression matching with a trigram index, or How Google Code search worked. Report, swtch.com, Cambridge, MA, USA, January 2012. URL http://swtch.

com/~rsc/regexp/regexp4.html. See also [Tho68, KP99,

Cox07, Cox09, Cox10].

Carabas:2017:EEV

[CP17a] Mihai Carabas and Pantelimon George Popescu. Energyefficient virtualized clusters. Future Generation Computer Systems, 74(??):151–157, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X15003313.

Cinque:2017:IHF

[CP17b] Marcello Cinque and Antonio Pecchia. On the injection of hardware faults in virtualized multicore systems. Journal of Parallel and Distributed Computing, 106(??):50–61, August 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731517300849.

Cao:2017:EMN

[CPKL17] Zizhong Cao, Shivendra S. Panwar, Murali Kodialam, and T. V. Lakshman. Enhancing mobile networks with software defined networking and cloud computing. IEEE/ACM Transactions on Networking, 25(3):1431–1444, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Castello:2018:EIR

[CPM^{+}18] Adrián Castelló, Antonio J. Peña, Rafael Mayo, Judit Planas, Enrique S. Quintana-Ortí, and Pavan Balaji. Exploring the interoperability of remote GPGPU virtualization using rCUDA and directive-based programming models. The Journal of Supercomputing, 74(11):5628–5642, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ceselli:2017:MEC

[CPS17] Alberto Ceselli, Marco Premoli, and Stefano Secci. Mobile edge cloud network design optimization. IEEE/ACM Transactions on Networking, 25(3):1818–1831, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Clifford:2014:AFB

[CPST14] Daniel Clifford, Hannes Payer, Michael Starzinger, and Ben L. Titzer. Allocation folding based on dominance. ACM SIGPLAN Notices, 49(11):15–24, November 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clifford:2015:MMD

[CPST15] Daniel Clifford, Hannes Payer, Michael Stanton, and Ben L. Titzer. Memento mori: dynamic allocation-site-based optimizations. ACM SIGPLAN Notices, 50(11):105–117, November 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2018:UVB

[CQLL18] Chen Chen, Zhuyun Qi, Yirui Liu, and Kai Lei. Using virtualization for blockchain testing. In Smart Computing and Communication, Lecture Notes in Computer Science, pages 289– 299. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2018. URL http://link.

springer.com/chapter/10.1007/978-3-319-73830-7_29.

Crawford:1998:BSJ

[Cra98] Ronald Crawford, II. Behind the scenes of the Java 1.1 Virtual Machine. Java Report: The Source for Java Development, 3(??):??, November 1998. CODEN JREPFI. ISSN

1086-4660. URL http://archive.javareport.com/9811/
html/from_pages/ftp_feature.shtml.

Craig:2005:VM

[Cra05] Iain D. Craig. Virtual machines. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-85233-969-1. xv + 269 pp. LCCN QA76.9 CRA.

Craig:2006:VM

[Cra06] Iain D. Craig. Virtual machines. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-85233-969-1. xv + 269 pp. LCCN QA76.9.V5 C73 2006.

Chowdhury:2012:VVN

[CRB12] Mosharaf Chowdhury, Muntasir Raihan Rahman, and Raouf Boutaba. ViNEYard: virtual network embedding algorithms with coordinated node and link mapping. IEEE/ACM Transactions on Networking, 20(1):206–219, February 2012. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Creasy:1965:GDR

[Cre65] Robert J. Creasy. General description of the research timesharing system with special emphasis on the control program. Memorandum 1, IBM Cambridge SR&D Center Research Time-Sharing Computer, Cambridge, MA, USA, January 29, 1965. ???? pp. This appears to be the earliest work on virtual machines that is cited in the IBM VM history [Var91]. That history reports on page 28: “Creasy and Comeau spent the last week of 1964 [36] joyfully brainstorming the design of CP-40, a new kind of operating system, a system that would provide not only virtual memory, but also virtual machines.

[37] They had seen that the cleanest way to protect users from one another (and to preserve compatibility as the new System/360 design evolved) was to use the System/360 Principles of Operations manual to describe the user’s interface to the Control Program. Each user would have a complete System/360 virtual machine (at first called a ‘pseudo-machine’).” Footnote 28 on page 28 says: “For the first few weeks, the CSC people referred to their concept as a ‘pseudo-machine’, but soon adopted the term ‘virtual machine’ after hearing Dave Sayre at IBM Research use it to describe a system he had built for a modified 7044.”.

Creeger:2008:CVR

[Cre08a] Mache Creeger. CTO virtualization roundtable, part II. Communications of the ACM, 51(12):43–49, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Creeger:2008:PCR

[Cre08b] Mache Creeger. Practice: CTO roundtable on virtualization. Communications of the ACM, 51(11):47–53, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Creeger:2009:CRV

[Cre09] Mache Creeger. CTO roundtable: Virtualization: Part one. ACM Queue: Tomorrow’s Computing Today, 7(1):26–35, January 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2010:MEC

[Cre10a] Mache Creeger. Moving to the edge: a CTO roundtable on network virtualization. Communications of the ACM, 53(8): 55–62, August 2010. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Creeger:2010:MEA

[Cre10b] Mache Creeger. Moving to the edge: An ACM CTO roundtable on network virtualization. ACM Queue: Tomorrow’s Computing Today, 8(7):20, July 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cruz:2016:DCG

[CRG16] Flavio Cruz, Ricardo Rocha, and Seth Copen Goldstein. Declarative coordination of graph-based parallel programs.

ACM SIGPLAN Notices, 51(8):4:1–4:??, August 2016. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crowl:1993:CLI

[Cro93] Lawrence A. Crowl. A common library interface to sharedmemory multiprocessors. Technical Report 93-80-08, Department of Computer Science, Oregon State University,

1993. URL http://www.cs.orst.edu/techpub/reports/
1993/1993R-ORSTCS-93-80-08/. Fri, 15 Dec 1995 03:17:47

GMT.

Carberry:2024:RTR

[CRX24] Joshua R. Carberry, John Rahme, and Haiping Xu. Realtime rejuvenation scheduling for cloud systems with virtualized software spares. The Journal of Systems and Software, 217(??):??, November 2024. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0164121224002139.

Cohen:1983:PSR

[CRZ83] Shismon Cohen, Roni Rosner, and Ari Zidon. PARALISP Simulator (Reference Manual). Research Report 83-2, Computer Science Department, Hebrew University, Jerusalem, Israel, January 1983.

Chiang:2015:SEV

[CRZH15] Ron C. Chiang, Sundaresan Rajasekaran, Nan Zhang, and H. Howie Huang. Swiper: Exploiting virtual machine vulnerability in third-party clouds with competition for I/O resources. IEEE Transactions on Parallel and Distributed Systems, 26(6):1732–1742, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.computer.org/csdl/trans/td/
2015/06/06824231-abs.html.

Chandersekaran:1976:FVM

[CS76] C. S. Chandersekaran and K. S. Shankar. Forum: On virtual machine integrity. IBM Systems Journal, 15(3): 264–269, 1976. CODEN IBMSA7. ISSN 0018-8670. See [DM75, DM76].

Caballer:2015:PDC

[CSMB15] Miguel Caballer, Damián Segrelles, Germán Moltó, and Ignacio Blanquer. A platform to deploy customized scientific virtual infrastructures on the cloud. Concurrency and Computation: Practice and Experience, 27(16):4318–4329, November 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2013:FRS

[CSS^{+}13] Hao Chen, Lin Shi, Jianhua Sun, Kenli Li, and Ligang He. A fast RPC system for virtual machines. IEEE Transactions on Parallel and Distributed Systems, 24(7):1267–1276, July 2013. CODEN ITDSEO. ISSN 1045-9219.

Chen:2016:CDD

[CSS^{+}16] Yuting Chen, Ting Su, Chengnian Sun, Zhendong Su, and Jianjun Zhao. Coverage-directed differential testing of JVM implementations. ACM SIGPLAN Notices, 51(6):85–99, June 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chehelgerdi-Samani:2021:PAP

[CSSE21] Maryam Chehelgerdi-Samani and Faramarz Safi-Esfahani. PCVM.ARIMA: predictive consolidation of virtual machines applying ARIMA method. The Journal of Supercomputing, 77(3):2172–2206, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https://link.

springer.com/article/10.1007/s11227-020-03354-3.

Cecchet:2011:DVD

[CSSS11] Emmanuel Cecchet, Rahul Singh, Upendra Sharma, and Prashant Shenoy. Dolly: virtualization-driven database provisioning for the cloud. ACM SIGPLAN Notices, 46(7):51– 62, July 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cameron:2015:JFE

[CSV15] Callum Cameron, Jeremy Singer, and David Vengerov. The judgment of FORSETI: economic utility for dynamic heap sizing of multiple runtimes. ACM SIGPLAN Notices, 50(11): 143–156, November 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2003:EJV

[CT03] D. Chen and V. Talwar. Extending Java Virtual Machines for networked embedded devices. Research Disclosure, 475: 475/076, 2003. CODEN ???? ISSN 0374-4353.

Cui:2017:PJP

[CTP^{+}17] Lin Cui, Fung Po Tso, Dimitrios P. Pezaros, Weijia Jia, and Wei Zhao. PLAN: Joint policy- and network-aware VM management for cloud data centers. IEEE Transactions on Parallel and Distributed Systems, 28(4):1163–1175, April 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2017/04/07556964-abs.html.

Cahill:1993:ICV

[CTS^{+}93] Vinny Cahill, Paul Taylor, Gradimir Starovic, Brendan Tangney, Darragh O’Grady, Rodger Lea, Christian Jacquemot, Peter Strarup Jensen, Paulo Amaral, Adam Mirowski, James Loveluck, Youcef Laribi, Xavier Rousset de Pina, and Pedro Sousa. Implementing the Comandos virtual machine. Technical Report TCD-CS-93-32, Trinity College; Chorus Systemes; OSF Grenoble Research Institute; Unite mixte BULL-IMAG; INESC, Dublin, Ireland; Grenoble, France, 1993. 50 pp. URL ftp://

ftp.cs.tcd.ie/pub/tcd/tech-reports/reports.93/TCD-
CS-93-32.ps.gz;   ftp://ftp.dsg.cs.tcd.ie/pub/doc/

TCD-CS-93-32.ps.gz. Also in Chapter 10 of The Comandos Distributed Application Platform.Cahill, V., Balter, R., Harris, N., and Rousset dePina, X. (Ed.s),Springer-Verlag, Berlin, 1993.

Chang:2013:ADA

[CVWL13] Jian Chang, Krishna K. Venkatasubramanian, Andrew G. West, and Insup Lee. Analyzing and defending against Webbased malware. ACM Computing Surveys, 45(4):49:1–49:??, August 2013. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Cai:2003:THI

[CW03] H. Cai and A. J. Wellings. Towards a high integrity realtime Java virtual machine. Lecture Notes in Computer Science, 2889:319–334, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:2014:CCB

[CWC^{+}14] Licheng Chen, Zhipeng Wei, Zehan Cui, Mingyu Chen, Haiyang Pan, and Yungang Bao. CMD: classification-based memory deduplication through page access characteristics. ACM SIGPLAN Notices, 49(7):65–76, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crandall:2006:TSD

[CWdO^{+}06] Jedidiah R. Crandall, Gary Wassermann, Daniela A. S. de Oliveira, Zhendong Su, S. Felix Wu, and Frederic T. Chong. Temporal search: detecting hidden malware timebombs with virtual machines. Operating Systems Review, 40 (5):25–36, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Crookston:2000:VCM

[CWG00] Newell W. Crookston, Brian Whipple, and William J. Grenney. Virtual C machine and integrated development environment for ATMS controllers. Technical Report 00-109, Utah Department of Transportation, Salt Lake City, UT, USA, April 2000. 23 pp.

Chang:2014:EMV

[CWH^{+}14] Chao-Jui Chang, Jan-Jan Wu, Wei-Chung Hsu, Pangfeng Liu, and Pen-Chung Yew. Efficient memory virtualization for Cross-ISA system mode emulation. ACM SIGPLAN Notices, 49(7):117–128, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2016:ICA

[CWH^{+}16] Renhai Chen, Yi Wang, Jingtong Hu, Duo Liu, Zili Shao, and Yong Guan. Image-content-aware I/O optimization for mobile virtualization. ACM Transactions on Embedded Computing Systems, 16(1):12:1–12:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cao:2012:EEA

[CWL12] Jian Cao, Yihua Wu, and Minglu Li. Energy efficient allocation of virtual machines in cloud computing environments based on demand forecast. Lecture Notes in Computer Science, 7296:137–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30767-6_12/

.

Cui:2015:PPA

[CWL^{+}15] Lei Cui, Tianyu Wo, Bo Li, Jianxin Li, Bin Shi, and Jinpeng Huai. PARS: a page-aware replication system for efficiently storing virtual machine snapshots. ACM SIGPLAN Notices, 50(7):215–228, July 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chakraborty:2012:SOV

[CWS12] Koushik Chakraborty, Philip M. Wells, and Gurindar S. Sohi. Supporting overcommitted virtual machines through hardware spin detection. IEEE Transactions on Parallel and Distributed Systems, 23(2):353–366, February 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2015:LVS

[CXLX15] Wenzhi Chen, Lei Xu, Guoxi Li, and Yang Xiang. A lightweight virtualization solution for Android devices. IEEE Transactions on Computers, 64(10):2741–2751, ???? 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cui:2017:TAV

[CYX^{+}17] Yong Cui, Zhenjie Yang, Shihan Xiao, Xin Wang, and Shenghui Yan. Traffic-aware virtual machine migration in topology-adaptive DCN. IEEE/ACM Transactions on Networking, 25(6):3427–3440, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Czajkowski:2000:AIJ

[Cza00] Grzegorz Czajkowski. Application isolation in the Java^{T M} Virtual Machine. ACM SIGPLAN Notices, 35(10):354–366, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.

acm.org/pubs/citations/proceedings/oops/353171/p354-
czajkowski/.

Chen:2025:ULH

[CZH^{+}25] Zhenqian Chen, Yuchun Zhan, Peng Hu, Xinkui Zhao, Muyu Yang, Siwei Tan, Lufei Zhang, Liqiang Lu, Jianwei Yin, and Zuoning Chen. UKFaaS: Lightweight, high-performance and secure FaaS communication with unikernel. IEEE Transactions on Computers, 74(10):3305–3318, October 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Carbone:2008:TV

[CZL08] Martim Carbone, Diego Zamboni, and Wenke Lee. Taming virtualization. IEEE Security & Privacy, 6(1):65–67, January/February 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chen:2019:GCE

[CZX^{+}19] H. Chen, X. Zhang, Y. Xu, J. Ren, J. Fan, Z. Ma, and W. Zhang. T-Gaming: A cost-efficient cloud gaming system at scale. IEEE Transactions on Parallel and Distributed Systems, 30(12):2849–2865, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dufrasne:2004:IVE

[D^{+}04] Bertrand Dufrasne et al., editors. Implementing VMware ESX Server with IBM TotalStorage FAStT. IBM redbooks. IBM, International Support Organization, ????, September 2004. ISBN 0-7384-9134-9. xvi + 246 pp. LCCN TK5105.86 I475 2004. Publication number SG24-6434-00.

Drummond:2024:DAW

[DAdBM^{+}24] Lúcia Maria A. Drummond, Luciano Andrade, Pedro de Brito Muniz, Matheus Marotti Pereira, Thiago do Prado Silva, and Luan Teylo. Design and analyses of web scraping on burstable virtual machines. Concurrency and Computation: Practice and Experience, 36(9):e7999:1–e7999:??, April 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dall:2012:DIE

[DAH^{+}12] Christoffer Dall, Jeremy Andrus, Alexander Van’t Hof, Oren Laadan, and Jason Nieh. The design, implementation, and evaluation of cells: a virtual Smartphone architecture. ACM Transactions on Computer Systems, 30(3):9:1–9:??, August 2012. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Dallmeier:1997:JVM

[Dal97] Matthias K. Dallmeier. Java Virtual Machine, Sprache, Konzept, Architektur. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 3-930673-73-8. 209 pp. LCCN ???? 25 DM. URL http://www.ora.de/. Includes CD-ROM.

Dalheimer:19xx:JVM

[Dalxx] Matthias K. Dalheimer. Java Virtual Machine, Sprache, Konzept. . .. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 19xx. ISBN 3-930673-73-8 (??invalid ISBN??).

LCCN ???? 29 DM. URL http://www.oreilly.de/german/
essential/javavm/index.html.

Danvy:1986:LVM

[Dan86] Olivier Danvy. LILA: a virtual machine for functional and declarative languages. Technical Report 86–38, LITP, France, May 1986. Workshop on Future Directions in Computer Science and Software.

Danielsson:2012:OSU

[Dan12] Nils Anders Danielsson. Operational semantics using the partiality monad. ACM SIGPLAN Notices, 47(9):127–138, September 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Daszczuk:1991:SSD

[Das91]  W. B. Daszczuk.  A structured semantic design of

distributed operating systems. The Computer Journal, 34(6):482–492, December 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www3.oup.co.uk/computer_journal/hdb/Volume_
34/Issue_06/tiff/482.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_34/Issue_06/tiff/483.
tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_34/Issue_06/tiff/484.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_34/Issue_06/tiff/
485.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/486.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_34/Issue_06/
tiff/487.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/488.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_34/Issue_06/
tiff/489.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/490.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_34/Issue_06/
tiff/491.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_06/tiff/492.tif.

Davoli:2004:TOS

[Dav04] Renzo Davoli. Teaching operating systems administration with User Mode Linux. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):112– 116, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dillenberger:2000:BJV

[DBC^{+}00] D. Dillenberger, R. Bordawekar, C. W. Clark, D. Durand, D. Emmes, O. Gohda, S. Howard, M. F. Oliver, F. Samuel, and R. W. St. John. Building a Java virtual machine for server applications: The JVM on OS/390. IBM Systems Journal, 39(1):194–210, ???? 2000. CODEN IBMSA7. ISSN

0018-8670. URL http://www.research.ibm.com/journal/
sj/391/dillenberger.html.

Darcy:1992:USD

[DBMI92] George A. Darcy III, Ronald F. Brender, Stephen J. Morris, and Michael V. Iles. Using simulation to develop and port software. Digital Technical Journal of Digital Equipment Corporation, 4(4):181–192, Fall 1992. CODEN DTJOEL.

ISSN 0898-901X. URL ftp://ftp.digital.com/pub/
Digital/info/DTJ/v4n4/Using_Simulation_to_Develop_
an_01apr1993DTJ812P8.ps;   http://www.digital.com:
80/info/DTJ812/DTJ812SC.TXT.

Denz:2018:SMB

[DBO^{+}18] Robert Denz, Scott Brookes, Martin Osterloh, Stephen Kuhn, and Stephen Taylor. Symmetric multiprocessing from boot to virtualization. Software—Practice and Experience, 48(3):681–718, March 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/spe.2505.

Di:2015:ECP

[DC15] Sheng Di and Franck Cappello. Extended conference papers: GloudSim: Google trace based cloud simulator with virtual machines. Software—Practice and Experience, 45(11):1571– 1590, November 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Doyle:2004:DIM

[DCA04] Patrick Doyle, Carlos Cavanna, and Tarek S. Abdelrahman. The design and implementation of a modular and extensible Java Virtual Machine. Software—Practice and Experience, 34(3):287–313, March 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dutra:2017:EVS

[DCA17] Diego L. C. Dutra, Edilson C. Corrêa, and Claudio L. Amorim. An efficient virtual system clock for the wireless Raspberry Pi computer platform. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coutinho:2015:OVM

[dCCDFdO15] Rafaelli de C. Coutinho, Lúcia M. A. Drummond, Yuri Frota, and Daniel de Oliveira. Optimizing virtual machine allocation for parallel scientific workflows in federated clouds. Future Generation Computer Systems, 46(??):51–68, May 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X14002027.

Deb:2012:HSC

[DCG12] Abhishek Deb, Josep Maria Codina, and Antonio Gonzalez. A HW/SW co-designed programmable functional unit. IEEE Computer Architecture Letters, 11(1):9–12, January/ June 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Junior:2016:PEV

[dCJR16] Francisco Heron de Carvalho Junior and Cenez Araújo Rezende. Performance evaluation of virtual execution environments for intensive computing on usual representations of multidimensional arrays. Science of Computer Programming, 132 (part 1)(??):29–49, December 15, 2016. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167642316300065.

Dias:2022:SLR

[DCM22] Alexandre H. T. Dias, Luiz. H. A. Correia, and Neumar Malheiros. A systematic literature review on virtual machine consolidation. ACM Computing Surveys, 54(8):176:1–176:38, November 2022. CODEN CMSVAN. ISSN 0360-0300 (print),

1557-7341 (electronic). URL https://dl.acm.org/doi/10.
1145/3470972.

Duan:2017:EAS

[DCMW17] Hancong Duan, Chao Chen, Geyong Min, and Yu Wu. Energy-aware scheduling of virtual machines in heterogeneous cloud computing systems. Future Generation Computer Systems, 74(??):142–150, September 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X16300292.

Dong:2012:RAE

[DCP^{+}12] Yaozu Dong, Yu Chen, Zhenhao Pan, Jinquan Dai, and Yunhong Jiang. ReNIC: Architectural extension to SR-IOV I/O virtualization for efficient replication. ACM Transactions on Architecture and Code Optimization, 8(4):40:1–40:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Daoud:2020:MAJ

[DD20] Houssem Daoud and Michel Dagenais. Multilevel analysis of the Java Virtual Machine based on kernel and userspace traces. The Journal of Systems and Software, 167(??):??, September 2020. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0164121220300698.

Dean:1994:CPV

[DDS^{+}94] C. E. Dean, R. C. Denny, P. C. Stephenson, G. J. Milne, and E. Pantos. Computing with parallel virtual machines. Journal de physique. IV, Colloque, 4(C9):C9/445–448, November 1994. CODEN JPICEI. ISSN 1155-4339.

DeRose:2006:EXI

[De 06] César De Rose. Evaluating Xen IA-64 security and performance. In Anonymous [Ano06a], page ?? ISBN ???? LCCN

???? URL http://www.ice.gelato.org/oct06/pres_pdf/
gelato_ICE06oct_xeneval_derose_pucrs.pdf.

Degenbaev:2016:ITG

[DEE^{+}16] Ulan Degenbaev, Jochen Eisinger, Manfred Ernst, Ross McIlroy, and Hannes Payer. Idle time garbage collection scheduling. ACM SIGPLAN Notices, 51(6):570–583, June 2016. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Diaz:2017:OAV

[DEG^{+}17] José Luis Díaz, Joaquín Entrialgo, Manuel García, Javier García, and Daniel Fernando García. Optimal allocation of virtual machines in multi-cloud environments with reserved and on-demand pricing. Future Generation Computer Systems, 71(??):129–144, June 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X17301954.

Debbabi:2003:MCA

[DEK^{+}03] M. Debbabi, M. Erhioui, L. Ketari, N. Tawbi, H. Yahyaoui, and S. Zhioua. Method call acceleration in embedded Java virtual machines. Lecture Notes in Computer Science, 2659: 750–759, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denning:2001:OVM

[Den01] Peter J. Denning. Origin of virtual machines and other virtualities. IEEE Annals of the History of Computing, 23(3):73, July 2001. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

computer.org/annals/an2001/a3073abs.htm;   http://
dlib.computer.org/an/books/an2001/pdf/a3073.pdf.

DELUG:2008:VKB

[Deu08] Deutsche Linux-User-Group, München, Germany. Vmknoppix 2008 05 19: auf Knoppix basierende Live-Distribution mit integrierten Virtualisierungssystemen wie Xen, KVM, Virtualbox oder Qemu, 2008. One CD-ROM.

Dincer:1996:BWW

[DF96] Kivanc Dincer and Geoffrey C. Fox. Building a worldwide virtual machine based on Web and HPCC technologies. In ACM [ACM96], page ?? ISBN 0-89791-854-

1. LCCN ???? URL http://www.supercomp.org/sc96/
proceedings/SC96PROC/DINCER/INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Davoli:2005:VSV

[DG05] Renzo Davoli and Michael Goldweber. Virtual square (V^{2}) in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):301– 305, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

dAntras:2017:HXU

[dGG^{+}17] Amanieu d’Antras, Cosmin Gorgovan, Jim Garside, John Goodacre, and Mikel Luján. HyperMAMBO-X64: Using virtualization to support high-performance transparent binary translation. ACM SIGPLAN Notices, 52(7):228–241, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dolan-Gavitt:2011:VNS

[DGLZ^{+}11] Brendan Dolan-Gavitt, Tim Leek, Michael Zhivich, Jonathon Giffin, and Wenke Lee. Virtuoso: Narrowing the semantic gap in virtual machine introspection. In 2011 IEEE Symposium on Security and Privacy. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 2011.

Decker:2001:PMS

[DH01] Rick Decker and Stuart Hirshfield. The PIPPIN machine: simulations of language processing. ACM Journal on Educational Resources in Computing (JERIC), 1(4):4–17, December 2001. CODEN ???? ISSN 1531-4278.

Donnell:2020:DVM

[DHD20] Nicola Mc Donnell, Enda Howley, and Jim Duggan. Dynamic virtual machine consolidation using a multi-agent system to optimise energy efficiency in cloud computing. Future Generation Computer Systems, 108(??):288–301, July 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X19314591.

Daly:2001:PID

[DHPW01] Charles Daly, Jane Horgan, James Power, and John Waldron. Platform independent dynamic Java Virtual Machine Analysis: the Java Grande Forum benchmark suite. In ACM [ACM01b], pages 106–115. ISBN 1-58113-359-6. LCCN

QA76.9.O35 A26 2001. URL http://www.philippsen.
com/JGI2001/camerareadyabstracts/16.html;   http:
//www.philippsen.com/JGI2001/finalpapers/18500106.
ps.

Donovan:1976:VMC

[DJ76] John J. Donovan and Henry D. Jacoby. Virtual machine communication for the implementation of decision support systems. Technical Report 884–76(28), MIT Center for Information Systems Research, Cambridge, MA, USA, 1976. 3 + 29 pp.

Donovan:1977:VMC

[DJ77] John J. Donovan and Henry D. Jacoby. Virtual machine communication for the implementation of decision support systems. IEEE Transactions on Software Engineering, SE-3(5): 333–342, September 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Benedictis:2024:NAV

[DJP^{+}24] Marco De Benedictis, Ludovic Jacquin, Ignazio Pedone, Andrea Atzeni, and Antonio Lioy. A novel architecture to virtualise a hardware-bound trusted platform module. Future Generation Computer Systems, 150(??):21–36, January 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X2300314X.

Dhawan:2017:CCA

[DJS^{+}17] Medhavi Dhawan, Gurprit Johal, Jim Stabile, Vjekoslav Brajkovic, James Chang, Kapil Goyal, Kevin James, Zeeshan Lokhandwala, Anny Martinez Manzanilla, Roger Michoud, Maithem Munshed, Srinivas Neginhal, Konstantin Spirov, Michael Wei, Scott Fritchie, Chris Rossbach, Ittai Abraham, and Dahlia Malkhi. Consistent clustered applications with Corfu. Operating Systems Review, 51(1):78–82, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

DeRemer:1975:PLV

[DK75] Frank DeRemer and Hans Kron. Programming-in-the large versus programming-in-the-small. ACM SIGPLAN Notices, 10(6):114–121, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Drapeau:1993:SLT

[DK93] Ann L. Drapeau and Randy H. Katz. Striping in large tape libraries. In IEEE [IEE93b], pages 378–387. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Deshpande:2017:TSL

[DK17] Umesh Deshpande and Kate Keahey. Traffic-sensitive live migration of virtual machines. Future Generation Computer Systems, 72(??):118–128, July 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X16301133.

Doran:2023:EVR

[DK23] Michael A. Doran and Nabeeh Kandalaft. Embedded virtualization on RISC-V with seL4. In IEEE, editor, 2023 IEEE 14th Annual Ubiquitous Computing, Electronics & Mobile Communication Conference (UEMCON), pages 0736–0740. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2023.

Damodaran-Kamal:1994:TRP

[DKF94] S. K. Damodaran-Kamal and J. M. Francioni. Testing races in parallel programs with an OtOt strategy. In Ostrand [Ost94].

CODEN SFENDP. ISBN 0-89791-683-2. ISSN 0163-5948. LCCN QA76.76.T48 I58 1994.

Di:2015:OCC

[DKW15] Sheng Di, D. Kondo, and Cho-Li Wang. Optimization of composite cloud service processing with virtual machines. IEEE Transactions on Computers, 64(6):1755–1768, ???? 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Demillo:1989:DSC

[DL89] Richard A. Demillo and Richard J. Lipton. Defining software by continuous, smooth functions. Technical Report SERC-TR-49-P, Software Engineering Research Centre, August 1989.

DeBenedictis:2019:IVD

[DL19a] Marco De Benedictis and Antonio Lioy. Integrity verification of Docker containers for a lightweight cloud environment. Future Generation Computer Systems, 97(??):236–246, August 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X18327201.

Dong:2019:ACM

[DL19b] Ying Dong and Zhou Lei. An access control model for preventing virtual machine hopping attack. Future Internet, 11 (3):82, March 26, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/1999-5903/11/3/82.

Ding:2020:AVM

[DLH^{+}20] Weichao Ding, Fei Luo, Liangxiu Han, Chunhua Gu, Haifeng Lu, and Joel Fuentes. Adaptive virtual machine consolidation framework based on performance-to-power ratio in cloud data centers. Future Generation Computer Systems, 111(??):254–270, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X19307769.

Dall:2016:AVP

[DLL^{+}16] Christoffer Dall, Shih-Wei Li, Jin Tack Lim, Jason Nieh, and Georgios Koloventzos. ARM virtualization: performance and architectural implications. ACM SIGARCH Computer Architecture News, 44(3):304–316, June 2016. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Dall:2018:AVP

[DLLN18] Christoffer Dall, Shih-Wei Li, Jin Tack Lim, and Jason Nieh. ARM virtualization: Performance and architectural implications. Operating Systems Review, 52(1):45–56, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dong:2006:EXI

[DLM^{+}06] Yaozu Dong, Shaofan Li, Asit Mallick, Jun Nakajim, Kun Tian, Xuefei Xu, Fred Yang, and Wilfred Yu. Extending Xen with Intel virtualization technology. Intel Technology Journal, 10(3):193–203, August 10, 2006. ISSN 1535-

766X. URL http://developer.intel.com/technology/
itj/2006/v10i3/3-xen/1-abstract.htm.

Dutchyn:2001:MDJ

[DLS^{+}01] Christopher Dutchyn, Paul Lu, Duane Szafron, Steven Bromling, and Wade Holst. Multi-dispatch in the Java Virtual Machine: Design and implementation. In USENIX [USE01b], page ?? ISBN 1-880446-12-X. LCCN ???? URL http:

//www.usenix.org/publications/library/proceedings/
coots01/dutchyn.html.

Deng:2017:DWT

[DLX^{+}17] Liang Deng, Peng Liu, Jun Xu, Ping Chen, and Qingkai Zeng. Dancing with wolves: Towards practical event-driven VMM monitoring. ACM SIGPLAN Notices, 52(7):83–96, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Donovan:1975:HAC

[DM75] John J. Donovan and S. E. Madnick. Hierarchical approach to computer system integrity. IBM Systems Journal, 14(2):188– 202, 1975. CODEN IBMSA7. ISSN 0018-8670. See letters [CS76, DM76].

Donovan:1976:FAR

[DM76] J. J. Donovan and S. E. Madnick. Forum: Authors’ reply. IBM Systems Journal, 15(3):270–278, 1976. CODEN IBMSA7. ISSN 0018-8670. See [DM75, CS76].

Despons:1993:CCP

[DM93] R. Despons and T. Muntean. Constructing correct protocols for a diffusion virtual machine in message passing parallel architectures. In Grebe et al. [GHH^{+}93], pages 465–480. ISBN 90-5199-140-1. LCCN ????

Dong:2015:VSB

[DMG^{+}15] YaoZu Dong, JunJie Mao, HaiBing Guan, Jian Li, and Yu Chen. A virtualization solution for BYOD with dynamic platform context switching. IEEE Micro, 35(1):34–43, January/February 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http://www.computer.

org/csdl/mags/mi/2015/01/mmi2015010034-abs.html.

Dhillon:2018:BEA

[DMH18] Vikram Dhillon, David Metcalf, and Max Hooper. Blockchain enabled applications: understand the blockchain ecosystem and how to make it work for you. Apress, Berkeley, CA, USA, 2018. ISBN 1-4842-3080-9 (print), 1-4842-3081-7 (e-book). xvii + 218 + 103 pp. LCCN QA76.9.D32. URL http://

link.springer.com/10.1007/978-1-4842-3081-7.

Dhiman:2010:VSE

[DMR10] Gaurav Dhiman, Giacomo Marchetti, and Tajana Rosing. vGreen: a system for energy-efficient management of virtual machines. ACM Transactions on Design Automation of Electronic Systems., 16(1):6:1–6:??, November 2010. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Diessel:2002:THL

[DMS02] O. Diessel, U. Malik, and K. So. Towards high-level specification, synthesis, and virtualization of programmable logic designs (research note). Lecture Notes in Computer Science, 2400:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2400/
24000314.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2400/24000314.pdf.

Dall:2014:KAD

[DN14] Christoffer Dall and Jason Nieh. KVM/ARM: the design and implementation of the Linux ARM hypervisor. ACM SIGARCH Computer Architecture News, 42(1):333– 348, March 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Dyer:2006:NPD

[DNR06] Robert Dyer, Harish Narayanappa, and Hridesh Rajan. Nu: preserving design modularity in object code. ACM SIGSOFT Software Engineering Notes, 31(6):1–2, November 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Do:2011:CAS

[Do11] Tien Van Do. Comparison of allocation schemes for virtual machines in energy-aware server farms. The Computer Journal, 54(11):1790–1797, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Oliveira:2012:SMC

[dOL12] Frederico Alvares de Oliveira, Jr. and Thomas Ledoux. Selfmanagement of cloud applications and infrastructure for energy optimization. Operating Systems Review, 46(2):10–18, July 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dommergaard:1980:DVM

[Dom80a] O. Dommergaard. The design of a virtual machine for Ada. In Abstract Software Specifications, pages 463–605. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980.

Dommergaard:1980:FDP

[Dom80b] S. Bodilsen O. Dommergaard. A formal definition of P-code. Technical report, Department of Comp. Sci., Techn. Univ. of Denmark, 1980.

Donaldson:1987:TOS

[Don87] John L. Donaldson. Teaching operating systems in a virtual machine environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 19(1):206–211, February 1987. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Donahue:1988:UAL

[Don88] Barry Donahue. Using assembly language to teach concepts in the introductory course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 20(1): 158–162, February 1988. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dong:2006:XIV

[Don06] Yaozu Dong. Xen and Intel virtualization technology for IA-64. In Anonymous [Ano06a], page ?? ISBN ???? LCCN

???? URL http://www.ice.gelato.org/oct06/pres_pdf/
gelato_ICE06oct_xenvt_dong_intel.pdf.

Deng:2011:CDE

[DP11] Yuhui Deng and Brandon Pung. Conserving disk energy in virtual machine based environments by amplifying bursts. Computing: Archiv für Informatik und Numerik, 91(1):3–21, January 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&
issn=0010-485X&volume=91&issue=1&spage=3.

Maio:2016:MEC

[DPBK16] Vincenzo De Maio, Radu Prodan, Shajulin Benedict, and Gabor Kecskemeti. Modelling energy consumption of network transfers and virtual machine migration. Future Generation Computer Systems, 56(??):388–406, March 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X15002307.

Dobre:2011:VBA

[DPCA11] Ciprian Dobre, Florin Pop, Valentin Cristea, and Ovidiu-Marian Achim. A virtualization-based approach to dependable service computing. Scalable Computing: Practice and Experience, 12(3):337–350, September 2011. CODEN ????

ISSN 1895-1767. URL http://www.scpe.org/index.php/
scpe/article/view/728.

Dai:2022:RBV

[DPCL22] Guangli Dai, Pavan Kumar Paluri, Albert Mo Kim Cheng, and Bozheng Liu. Regularity-based virtualization under the ARINC 653 Standard for Embedded Systems. IEEE Transactions on Computers, 71(10):2592–2605, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dalton:2009:TVP

[DPW^{+}09] Chris I. Dalton, David Plaquin, Wolfgang Weidner, Dirk Kuhlmann, Boris Balacheff, and Richard Brown. Trusted virtual platforms: a key enabler for converged client devices. Operating Systems Review, 43(1):36–43, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ding:2015:EES

[DQLW15] Youwei Ding, Xiaolin Qin, Liang Liu, and Taochun Wang. Energy efficient scheduling of virtual machines in cloud with deadline constraint. Future Generation Computer Systems, 50(??):62–74, September 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X15000369.

Dai:2013:LVM

[DQR^{+}13] Yuehua Dai, Yong Qi, Jianbao Ren, Yi Shi, Xiaoguang Wang, and Xuan Yu. A lightweight VMM on many core for high performance computing. ACM SIGPLAN Notices, 48(7):111– 120, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Drepper:2008:CV

[Dre08] Ulrich Drepper. The cost of virtualization. ACM Queue: Tomorrow’s Computing Today, 6(1):28–35, January 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Desai:2009:AIC

[DS09a] Aniruddha Desai and Jugdutt Singh. Architecture independent characterization of embedded Java workloads. IEEE Computer Architecture Letters, 8(1):29–32, January/June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Dowty:2009:GVV

[DS09b] Micah Dowty and Jeremy Sugerman. GPU virtualization on VMware’s hosted I/O architecture. Operating Systems Review, 43(3):73–82, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dragga:2016:GGC

[DS16] Chris Dragga and Douglas J. Santry. GCTrees: Garbage collecting snapshots. ACM Transactions on Storage, 12(1): 4:1–4:??, February 2016. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Dezhabad:2018:LBD

[DS18] Naghmeh Dezhabad and Saeed Sharifian. Learning-based dynamic scalable load-balanced firewall as a service in network function-virtualized cloud computing environments. The Journal of Supercomputing, 74(7):3329–3358, July 2018. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Demirci:2019:OPV

[DS19] Sedef Demirci and Seref Sagiroglu. Optimal placement of virtual network functions in software defined networks: a survey. Journal of Network and Computer Applications, 147(??):??, December 1, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804519302760.

Dhule:2020:PSP

[DS20] Chetan Dhule and Urmila Shrawankar. POF-SVLM: Pareto optimized framework for seamless VM live migration. Computing: Archiv für Informatik und Numerik, 102(10):2159– 2183, October 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Durairaj:2022:TSV

[DS22] Selvam Durairaj and Rajeswari Sridhar. Task scheduling to a virtual machine using a multi-objective mayfly approach for a cloud environment. Concurrency and Computation: Practice and Experience, 34(24):e7236:1–e7236:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deng:2008:CCV

[DSC^{+}08] Yi Deng, S. Masoud Sadjadi, Peter J. Clarke, Vagelis Hristidis, Raju Rangaswami, and Yingbo Wang. CVM — a communication virtual machine. The Journal of Systems and Software, 81(10):1640–1662, October 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

daSilva:2016:TAV

[dSdF16] Rodrigo A. C. da Silva and Nelson L. S. da Fonseca. Topology-aware virtual machine placement in data centers. Journal of Grid Computing, 14(1):75–90, March 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10723-015-9343-x.

Dargie:2014:PCE

[DSM14] Waltenegus Dargie, Alexander Schill, and Christoph Mobius. Power consumption estimation models for processors, virtual machines, and servers. IEEE Transactions on Parallel and Distributed Systems, 25(6):1600–1614, June 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Debski:2018:SRA

[DSM^{+}18] A. Debski, B. Szczepanik, M. Malawski, S. Spahr, and D. Muthig. A scalable, reactive architecture for cloud applications. IEEE Software, 35(2):62–71, March/April 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

daSilva:2017:ARA

[dSOK17] Marcelo Pereira da Silva, Rafael Rodrigues Obelheiro, and Guilherme Piegas Koslovski. Adaptive Remus: adaptive checkpointing for Xen-based virtual machine replication. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 32(4):348–367, 2017. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Dangl:2023:VEF

[DSR23] Thomas Dangl, Stewart Sentanoe, and Hans P. Reiser. VMIFresh: Efficient and fresh caches for virtual machine introspection. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823004376.

DUrso:2019:WSS

[DSS19] Fabio D’Urso, Corrado Santoro, and Federico Fausto Santoro. Wale: a solution to share libraries in Docker containers. Future Generation Computer Systems, 100(??):513–522, November 2019. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167739X18327511.

DeRosa:2006:RSD

[DSSP06] Peter DeRosa, Kai Shen, Christopher Stewart, and Jonathan Pearson. Realism and simplicity: disk simulation for instructional OS performance evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):308–312, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Du:2011:PPV

[DSZ11] Jiaqing Du, Nipun Sehrawat, and Willy Zwaenepoel. Performance profiling of virtual machines. ACM SIGPLAN Notices, 46(7):3–14, July 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Du:2007:SSI

[DTW07] Wenliang Du, Zhouxuan Teng, and Ronghua Wang. SEED: a suite of instructional laboratories for computer SEcurity EDucation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):486–490, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dunigan:1986:DHM

[Dun86] T. W. Dunigan. Denelcor HEP multiprocessor simulator. Report ORNL/TM-9971, Engineering Physics and Mathematics Division, Mathematical Sciences Section, Oak Ridge National Laboratory, Oak Ridge, TN 37831, USA, June 1986. iv

+ 22 pp. URL http://web.ornl.gov/info/reports/1986/
3445600639931.pdf.

Dillon:2014:VHN

[DW14] Marion Dillon and Timothy Winters. Virtualization of home network gateways. Computer, 47(11):62–65, November 2014. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.org/csdl/mags/co/
2014/11/mco2014110062-abs.html.

Dou:2017:EAV

[DXM^{+}17] Wanchun Dou, Xiaolong Xu, Shunmei Meng, Xuyun Zhang, Chunhua Hu, Shui Yu, and Jian Yang. An energy-aware virtual machine scheduling method for service QoS enhancement in clouds over big data. Concurrency and Computation: Practice and Experience, 29(14), July 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Duan:2017:LBM

[DY17] Jun Duan and Yuanyuan Yang. A load balancing and multi-tenancy oriented data center virtualization framework. IEEE Transactions on Parallel and Distributed Systems, 28(8):2131–2144, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:

//www.computer.org/csdl/trans/td/2017/08/07831441-
abs.html.

Dong:2012:HPN

[DYL^{+}12] Yaozu Dong, Xiaowei Yang, Jianhui Li, Guangdeng Liao, Kun Tian, and Haibing Guan. High performance network virtualization with SR-IOV. Journal of Parallel and Distributed Computing, 72(11):1471–1480, November 2012. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731512000329.

Dugan:2002:SIS

[DZ02] B. Dugan and J. Zahorjan. The Sloop ISA and the SMOK toolkit. ACM Journal on Educational Resources in Computing (JERIC), 2(1):49–71, March 2002. CODEN ???? ISSN 1531-4278.

Ebadifard:2020:SSW

[EB20] Fatemeh Ebadifard and Seyed Morteza Babamir. Scheduling scientific workflows on virtual machines using a Pareto and hypervolume based black hole optimization algorithm. The Journal of Supercomputing, 76(10):7635–7688, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL https://link.springer.com/article/
10.1007/s11227-020-03183-4.

Escheikh:2017:VWA

[EBJ17] Mohamed Escheikh, Kamel Barkaoui, and Hana Jouini. Versatile workload-aware power management performability analysis of server virtualized systems. The Journal of Systems and Software, 125(??):365–379, March 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228

(electronic). URL //www.sciencedirect.com/science/
article/pii/S0164121216302588.

Eiling:2022:CVL

[EBLM22] Niklas Eiling, Jonas Baude, Stefan Lankes, and Antonello Monti. Cricket: a virtualization layer for distributed execution of CUDA applications with checkpoint/restart support. Concurrency and Computation: Practice and Experience, 34 (14):e6474:1–e6474:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elmore:2013:TDV

[ECAE13] Aaron J. Elmore, Carlo Curino, Divyakant Agrawal, and Amr El Abbadi. Towards database virtualization for database as a service. Proceedings of the VLDB Endowment, 6(11):1194– 1195, August 2013. CODEN ???? ISSN 2150-8097.

Enes:2018:BDO

[ECET18] Jonatan Enes, Javier López Cacheiro, Roberto R. Expósito, and Juan Touriño. Big data-oriented PaaS architecture with disk-as-a-resource capability and container-based virtualization. Journal of Grid Computing, 16(4):587–605, December 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL https://link.springer.com/article/
10.1007/s10723-018-9460-4.

Egger:2016:ECL

[ECJ^{+}16] Bernhard Egger, Younghyun Cho, Changyeon Jo, Eunbyun Park, and Jaejin Lee. Efficient checkpointing of live virtual machines. IEEE Transactions on Computers, 65(10):3041– 3054, ???? 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

ECMA-335-1

[ECM01] ECMA. ECMA-335: Common Language Infrastructure (CLI). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 2001. xiii + 479 pp. URL http://www.

ecma.ch/ecma1/STAND/ecma-335.htm.

ECMA-335-2

[ECM02] ECMA. ECMA-335: Common Language Infrastructure (CLI). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, second edition, December 2002. iv + 99 (Part I), vii + 164 (Part II), iv + 124 (Part III), i + 16 (Part IV), ii + 78 (Part V) pp. URL http://www.ecma-

international.org/publications/files/ECMA-ST/Ecma-
335-xml.zip;   http://www.ecma-international.org/
publications/files/ECMA-ST/Ecma-335.pdf;   http://
www.ecma.ch/ecma1/STAND/ecma-335.htm.

ECMA-335-3

[ECM05] ECMA. ECMA-335: Common Language Infrastructure (CLI). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, third edition, June 2005. viii + 104 (Part I), viii + 192 (Part II), iv + 138 (Part III), ii + 20 (Part IV), i + 4 (Part V), ii + 60 (Part VI) pp. URL http://www.ecma-

international.org/publications/files/ECMA-ST/Ecma-
335-xml.zip;   http://www.ecma-international.org/
publications/files/ECMA-ST/Ecma-335.pdf;   http://
www.ecma.ch/ecma1/STAND/ecma-335.htm.

ECMA-335-4

[ECM06] ECMA. ECMA-335: Common Language Infrastructure (CLI). ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, fourth edition, June 2006. vii + 104 (Part I), viii + 191 (Part II), iv + 138 (Part III), ii + 20 (Part IV), i + 4 (Part V), ii + 57 (Part VI) pp. URL http://www.ecma-

international.org/publications/files/ECMA-ST/Ecma-
335.pdf; http://www.ecma-international.org/publications/
files/ECMA-ST/ECMA-335.zip;   http://www.ecma-international.
org/publications/standards/Ecma-335.htm.

Ejarque:2010:ESV

[EdPG^{+}10] Jorge Ejarque, Marc de Palol, Íñigo Goiri, Ferran Julià, Jordi Guitart, Rosa M. Badia, and Jordi Torres. Exploiting semantics and virtualization for SLA-driven resource allocation in service providers. Concurrency and Computation: Practice and Experience, 22(5):541–572, April 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Estrada:2015:PCT

[EDS^{+}15] Zachary John Estrada, Fei Deng, Zachary Stephens, Cuong Pham, Zbigniew Kalbarczyk, and Ravishankar Iyer. Performance comparison and tuning of virtual machines for sequence alignment software. Scalable Computing: Practice and Experience, 16(1):??, ???? 2015. CODEN ???? ISSN

1895-1767. URL http://www.scpe.org/index.php/scpe/
article/view/1061.

Erenyi:1994:IPA

[EF94] I. Erenyi and Z. Fazekas. Image processing applications and their parallel aspects. Computing and Control Engineering Journal, 5(2):71–4, 1994. CODEN CCEJEL. ISSN 0956-3385 (print), 1741-0460 (electronic).

Ertl:2001:BEV

[EG01] M. Anton Ertl and David Gregg. The behavior of efficient virtual machine interpreters on modern architectures. Lecture Notes in Computer Science, 2150:403–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2150/21500403.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2150/21500403.pdf.

Ertl:2003:OIB

[EG03] M. Anton Ertl and David Gregg. Optimizing indirect branch prediction accuracy in virtual machine interpreters. ACM SIGPLAN Notices, 38(5):278–288, May 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eeckhout:2003:HJP

[EGD03] Lieven Eeckhout, Andy Georges, and Koen De Bosschere. How Java programs interact with virtual machines at the microarchitectural level. ACM SIGPLAN Notices, 38(11): 169–186, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Entrialgo:2024:JAC

[EGG^{+}24] Joaquín Entrialgo, Manuel García, Javier García, José María López, and José Luis Díaz. Joint autoscaling of containers and virtual machines for cost optimization in container clusters. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10723-023-09732-4.

Egger:2015:ERV

[EGJS15] Bernhard Egger, Erik Gustafsson, Changyeon Jo, and Jeongseok Son. Efficiently restoring virtual machines. International Journal of Parallel Programming, 43(3):421–439, June 2015. CODEN IJPPE5. ISSN 0885-7458 (print),

1573-7640 (electronic). URL http://link.springer.com/
article/10.1007/s10766-013-0295-0.

Ertl:2002:VGE

[EGKP02] M. Anton Ertl, David Gregg, Andreas Krall, and Bernd Paysan. Vmgen — a generator of efficient virtual machine interpreters. Software—Practice and Experience, 32(3):265– 294, March 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://www.complang.

tuwien.ac.at/anton/vmgen/;   http://www.complang.
tuwien.ac.at/papers/ertl+02.ps.gz;   http://www3.
interscience.wiley.com/cgi-bin/abstract/90010508/
START;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=90010508&PLACEBO=IE.pdf.

Ebrahimirad:2015:EAS

[EGR15] Vahid Ebrahimirad, Maziar Goudarzi, and Aboozar Rajabi. Energy-aware scheduling for precedence-constrained parallel virtual machines in virtualized data centers. Journal of Grid Computing, 13(2):233–253, June 2015. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10723-015-9327-x.

Esteire:1998:STN

[EL98] Oscar Azañón Esteire and Juan Manuel Cueva Lovelle. Set of tools for native code generation for the Java virtual machines. ACM SIGPLAN Notices, 33(3):73–79, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eramo:2019:ESR

[ELC^{+}19] Vincenzo Eramo, Francesco G. Lavacca, Tiziana Catena, Marco Polverini, and Antonio Cianfrani. Effectiveness of segment routing technology in reducing the bandwidth and cloud resources provisioning times in network function virtualization architectures. Future Internet, 11(3):71, March 12, 2019. CODEN ???? ISSN 1999-5903. URL https://www.mdpi.

com/1999-5903/11/3/71.

England:2006:VME

[EM06] Paul England and John Manferdelli. Virtual machines for enterprise desktop security. Information Security Technical Report, 11(4):193–202, ???? 2006. CODEN ISTRFR. ISSN 1363-4127 (print), 1873-605X (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1363412706000501.

Eramo:2017:ASF

[EMAL17] Vincenzo Eramo, Emanuele Miucci, Mostafa Ammar, and Francesco Giacinto Lavacca. An approach for service function chain routing and virtual function network instance migration in network function virtualization architectures. IEEE/ ACM Transactions on Networking, 25(4):2008–2025, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Esposito:2013:SES

[EMI13] Flavio Esposito, Ibrahim Matta, and Vatche Ishakian. Slice embedding solutions for distributed service architectures.

ACM Computing Surveys, 46(1):6:1–6:??, October 2013. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Evoy:2015:ADP

[EMS15] Giacomo Mc Evoy, Antonio R. Mury, and Bruno Schulze. An analysis of definition and placement of virtual machines for high performance applications on Clouds. Concurrency and Computation: Practice and Experience, 27(7):1789–1814, May 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Esposito:2016:VAV

[EMW16] Flavio Esposito, Ibrahim Matta, and Yuefeng Wang. VINEA: An architecture for virtual network embedding policy programmability. IEEE Transactions on Parallel and Distributed Systems, 27(11):3381–3396, November 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2016/11/07401050-abs.html.

Engel:1999:PJV

[Eng99] Joshua Engel. Programming for the Java Virtual Machine. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-30972-6 , 0-201-61654-8 (CD-ROM). xix + 488 pp. LCCN QA76.73.J38E543 1999. US$39.95.

Ertl:2003:IVM

[Ert03] M. Anton Ertl, editor. Interpreters, Virtual Machines and Emulators (IVME ’03). ACM Press, New York, NY 10036, USA, 2003. ISBN 1-58311-655-2. LCCN QA76.7. URL http:

//www.complang.tuwien.ac.at/anton/ivme03/proceedings/
ivme.ps.gz.

Ertl:2005:AIV

[Ert05] M. Anton Ertl. Advances in interpreters, virtual machines and emulators. Science of Computer Programming, 57(3): 251–252, September 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Estrada:2017:UDP

[ESY^{+}17] Zachary J. Estrada, Read Sprabery, Lok Yan, Zhongzhi Yu, Roy Campbell, Zbigniew Kalbarczyk, and Ravishankar K.

Iyer. Using OS design patterns to provide reliability and security as-a-service for VM-based clouds. ACM SIGPLAN Notices, 52(7):157–170, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ElBalmany:2022:SVM

[ETAB22] Chawki El Balmany, Zakariae Tbatou, Ahmed Asimi, and Mohamed Bamarouf. Secure virtual machine image storage process into a trusted zone-based cloud storage. Computers & Security, 120(??):Article 102815, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002097.

Eugster:2006:UPJ

[Eug06] Patrick Eugster. Uniform proxies for Java. ACM SIGPLAN Notices, 41(12):139–152, December 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eramo:2021:PIC

[EVCL21] Vincenzo Eramo, Francesco Valente, Tiziana Catena, and Francesco Giacinto Lavacca. Proposal and investigation of a convolutional and LSTM neural network for the cost-aware resource prediction in softwarized networks. Future Internet, 13(12):316, December 16, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.mdpi.com/1999-5903/13/12/316.

ElMotaki:2021:PBM

[EYG21] Saloua El Motaki, Ali Yahyaouy, and Hamid Gualous. A prediction-based model for virtual machine live migration monitoring in a cloud datacenter. Computing: Archiv für Informatik und Numerik, 103(11):2711–2735, November 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00607-021-00981-3.

Motaki:2019:CSB

[EYGS19] Saloua El Motaki, Ali Yahyaouy, Hamid Gualous, and Jalal Sabor. Comparative study between exact and metaheuristic approaches for virtual machine placement process as knapsack problem. The Journal of Supercomputing, 75(10):6239– 6259, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Feizollahibarough:2021:SAV

[FA21] Sattar Feizollahibarough and Mehrdad Ashtiani. A securityaware virtual machine placement in the cloud using hesitant fuzzy decision-making processes. The Journal of Supercomputing, 77(6):5606–5636, June 2021. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.com/article/10.1007/s11227-
020-03496-4.

Fard:2017:DVC

[FAA17a] Seyed Yahya Zahedi Fard, Mohamad Reza Ahmadi, and Sahar Adabi. A dynamic VM consolidation technique for QoS and energy consumption in cloud environment. The Journal of Supercomputing, 73(10):4347–4368, October 2017. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See erratum [FAA17b].

Fard:2017:EDV

[FAA17b] Seyed Yahya Zahedi Fard, Mohamad Reza Ahmadi, and Sahar Adabi. Erratum to: A dynamic VM consolidation technique for QoS and energy consumption in cloud environment. The Journal of Supercomputing, 73(10):4369–4372, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11227-017-2079-6.pdf. See [FAA17a].

Fabbro:2013:LAS

[Fab13] Andrew Fabbro. Linux on Azure — a strange place to find a penguin. Linux Journal, 2013(226):2:1–2:??, February 2013. CODEN LIJOFX. ISSN 1075-3583.

Ferrise:2024:MEE

[FBGS24] Francesco Ferrise, Monica Bordegoni, Alberto Gallace, and Stefania Serafin. Multisensory experiences in extended reality. IEEE Computer Graphics and Applications, 44(4):11– 13, July/August 2024. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Fokaefs:2018:DBE

[FBL18] Marios Fokaefs, Cornel Barna, and Marin Litoiu. From DevOps to BizOps: Economic sustainability for scalable cloud applications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(4):25:1–25:??, January 2018. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Farkiani:2021:PDD

[FBM^{+}21] Behrooz Farkiani, Bahador Bakhshi, S. Ali MirHassani, Tim Wauters, Bruno Volckaert, and Filip De Turck. Prioritized deployment of dynamic service function chains. IEEE/ACM Transactions on Networking, 29(3):979–993, June 2021. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.acm.org/doi/10.1109/TNET.
2021.3055074.

Feng:2012:IDU

[FBZS12] Zhenqian Feng, Bing Bai, Baokang Zhao, and Jinshu Su. IC-Tree: discovering the underlying connections of your rental virtual machines in the public clouds. International Journal of Computer Systems Science and Engineering, 27(6):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Fong:1998:PLA

[FC98] Philip W. L. Fong and Robert D. Cameron. Proof linking: an architecture for modular verification of dynamically-linked mobile code. ACM SIGSOFT Software Engineering Notes, 23 (6):222–230, November 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Fagen:2009:VEN

[FCD09] Wade A. Fagen, João W. Cangussu, and Ram Dantu. A virtual environment for network testing. Journal of Network and Computer Applications, 32(1):184–214, January 2009. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S108480450800026X.

Franz:2005:PVM

[FCG^{+}05] Michael Franz, Deepak Chandra, Andreas Gal, Vivek Haldar, Christian W. Probst, Fermín Reig, and Ning Wang. A portable virtual machine target for proof-carrying code. Science of Computer Programming, 57(3):275–294, September 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Flich:2008:LBD

[FD08] J. Flich and J. Duato. Logic-based distributed routing for NoCs. IEEE Computer Architecture Letters, 7(1):13–16, January 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Flatt:2019:RRC

[FDD^{+}19] Matthew Flatt, Caner Derici, R. Kent Dybvig, Andrew W. Keep, Gustavo E. Massaccesi, Sarah Spall, Sam Tobin-Hochstadt, and Jon Zeppieri. Rebuilding Racket on Chez Scheme (experience report). Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):78:1–78:15,

July 2019. URL https://dl.acm.org/doi/abs/10.1145/
3341642.

Figueiredo:2005:GEI

[FDF05] Renato Figueiredo, Peter A. Dinda, and José Fortes. Guest Editors’ introduction: Resource virtualization renaissance. Computer, 38(5):28–??, May 2005. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.org/comp/mags/
co/2005/05/r5028.pdf;   http://csdl.computer.org/
comp/mags/co/2005/05/r5028abs.htm.

Ferrell:2011:DRV

[Fer11] Robert G. Ferrell. /dev/random: Virtualization: a dark proverb. ;login: the USENIX Association newsletter, 36(5): ??, October 2011. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/publications/login/october-
2011-volume-36-number-5/devrandom-virtualization-
dark-proverb.

Fox:1996:TWJ

[FF96] G. Fox and W. Furmanski. Towards Web/Java-based high performance distributed computing — an evolving virtual machine. In IEEE [IEE96b], page ?? ISBN 0-8186-7582-9. LCCN QA76.88.I52 1996. IEEE order plan catalog number 96TB100069. EEE Computer Society Press order number PR07582.

Farzaneh:2022:NVM

[FF22] Seyed Milad Farzaneh and Omid Fatemi. A novel virtual machine placement algorithm using RF element in cloud infrastructure. The Journal of Supercomputing, 78(1):1288–1329, January 2022. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL https://link.springer.com/
article/10.1007/s11227-021-03863-9.

Farkas:2000:QEC

[FFB^{+}00] Keith I. Farkas, Jason Flinn, Godmar Back, Dirk Grunwald, and Jennifer M. Anderson. Quantifying the energy consumption of a pocket computer and a Java Virtual Machine. ACM SIGMETRICS Performance Evaluation Review, 28(1):252– 263, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Faibish:2008:SVU

[FFBG08] Sorin Faibish, Stephen Fridella, Peter Bixby, and Uday Gupta. Storage virtualization using a block-device file system. Operating Systems Review, 42(1):119–126, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fragiadakis:2023:AML

[FFM^{+}23] George Fragiadakis, Evangelia Filiopoulou, Christos Michalakelis, Thomas Kamalakis, , and Mara Nikolaidou. Applying machine learning in cloud service price prediction: The case of Amazon IaaS. Future Internet, 15(8):277, August 19, 2023. CODEN ???? ISSN 1999-5903. URL https://www.

mdpi.com/1999-5903/15/8/277.

Fertig:1991:FVM

[FG91] Scott Fertig and David Gelernter. FGP: a virtual machine for acquiring knowledge from cases. In Mylopoulos and Reiter [MR91], pages 796–802. ISBN 1-55860-160-0. LCCN Q334 .I57 1991. Two volumes.

Frincu:2014:ESV

[FGG14] Marc E. Frincu, Stéphane Genaud, and Julien Gossa. On the efficiency of several VM provisioning strategies for workflows with multi-threaded tasks on clouds. Computing: Archiv für Informatik und Numerik, 96(11):1059–1086, November 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057

(electronic). URL http://link.springer.com/article/
10.1007/s00607-014-0410-0.

Forsman:2015:AAL

[FGLI15] Mattias Forsman, Andreas Glad, Lars Lundberg, and Dragos Ilie. Algorithms for automated live migration of virtual machines. The Journal of Systems and Software, 101(??):110– 126, March 2015. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0164121214002751.

Fan:2023:MOO

[FGZC23] Tian Fan, Wanwan Guo, Zhixia Zhang, and Zhihua Cui. A many-objective optimization based intelligent algorithm for virtual machine migration in mobile edge computing. Concurrency and Computation: Practice and Experience, 35(23): e7770:1–e7770:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ford:1996:MMR

[FHL^{+}96] Bryan Ford, Mike Hibler, Jay Lepreau, Patrick Tullmann, Godmar Back, and Stephen Clawson. Microkernels meet recursive virtual machines. Operating Systems Review, 30(SI): 137–151, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Field:1968:MAS

[Fie68] M. S. Field. Multi-access systems: The virtual machine approach. Report 320-2033, IBM Cambridge Scientific Center, Cambridge, MA, USA, September 1968.

Freudenberg:2015:SMP

[FIF^{+}15] Bert Freudenberg, Dan H. H. Ingalls, Tim Felgentreff, Tobias Pape, and Robert Hirschfeld. SqueakJS: a modern and practical smalltalk that runs in any browser. ACM SIGPLAN Notices, 50(2):57–66, February 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fischofer:1991:VSS

[Fis91] W. T. Fischofer. VM/ESA: a single system for centralized and distributed computing. IBM Systems Journal, 30(1):4– 13, 1991. CODEN IBMSA7. ISSN 0018-8670.

Fischer:2001:SAN

[Fis01] Markus Fischer. System area network extensions to the parallel virtual machine. Lecture Notes in Computer Science, 2131: 98–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-ny.

com/link/service/series/0558/bibs/2131/21310098.htm;
 http://link.springer-ny.com/link/service/series/
0558/papers/2131/21310098.pdf.

Fischer:2009:XUH

[Fis09] Marcus Fischer. Xen das umfassende Handbuch. (German) [Xen, the Complete Handbook]. Galileo computing. Galileo Press, Bonn, Germany, 2009. ISBN 3-8362-1118-1. 547 pp.

LCCN ???? URL http://d-nb.info/986059226/04;
http://deposit.d-nb.de/cgi-bin/dokserv?id=3015256;
 http://www.galileocomputing.de/.

Fitzhugh:2014:VVM

[Fit14] Rebecca Fitzhugh. vSphere virtual machine management: create vSphere virtual machines, manage performance, and explore advanced capabilities. Professional expertise distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78217-218-1, 1-78217-219-X. vi + 307 pp. LCCN QA76.9.V5 F58 2014.

Firoozjaei:2017:SCN

[FJKK17] Mahdi Daghmehchi Firoozjaei, Jaehoon (Paul) Jeong, Hoon Ko, and Hyoungshick Kim. Security challenges with network functions virtualization. Future Generation Computer Systems, 67(??):315–324, February 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X16302321.

Friedman:2003:TFT

[FK03] R. Friedman and A. Kama. Transparent fault-tolerant Java Virtual Machine. Proceedings of the Symposium on Reliable Distributed Systems, EDIT22:319–328, 2003. CODEN ???? ISSN 1060-9857.

Fu:2013:SGW

[FK13] Fangwen Fu and Ulas C. Kozat. Stochastic game for wireless network virtualization. IEEE/ACM Transactions on Networking, 21(1):84–97, February 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fink:2017:VMD

[FKZ17] Bryan Fink, Eric Knauft, and Gene Zhang. vSAN: Modern distributed storage. Operating Systems Review, 51(1):33–37, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fu:2013:BSG

[FL13a] Yangchun Fu and Zhiqiang Lin. Bridging the semantic gap in virtual machine introspection via online kernel data redirection. ACM Transactions on Information and System Security, 16(2):7:1–7:??, September 2013. CODEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Fu:2013:EUD

[FL13b] Yangchun Fu and Zhiqiang Lin. EXTERIOR: using a dual-VM based external shell for guest-OS introspection, configuration, and recovery. ACM SIGPLAN Notices, 48(7):97– 110, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Flouris:2010:EBL

[FLCB10] Michail D. Flouris, Renaud Lachaize, Konstantinos Chasapis, and Angelos Bilas. Extensible block-level storage virtualization in cluster-based systems. Journal of Parallel and Distributed Computing, 70(8):800–824, August 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Flink:1977:EOS

[Fli77] Charles W. Flink. EASY — an operating system for the QM-1. ACM SIGMICRO Newsletter, 8(3):7–14, September 1977. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL

https://dl.acm.org/doi/10.1145/1014196.803310.

Fang:2013:VOV

[FLL^{+}13] Weiwei Fang, Xiangmin Liang, Shengxin Li, Luca Chiaraviglio, and Naixue Xiong. VMPlanner: Optimizing virtual machine placement and traffic flow routing to reduce network power costs in cloud data centers. Computer Networks (Amsterdam, Netherlands: 1999), 57(1):179–196, January 16, 2013. CODEN ???? ISSN 1389-1286 (print), 1872-

7069 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1389128612003301.

Franklin:2008:RDV

[FLM^{+}08] Jason Franklin, Mark Luk, Jonathan M. McCune, Arvind Seshadri, Adrian Perrig, and Leendert van Doorn. Remote detection of virtual machine monitors with fuzzy benchmarking. Operating Systems Review, 42(3):83–92, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Anonymous:2014:AVM

[fLtNW14] Xiao feng Li, Jiu tao Nie, and Ligang Wang. Advanced Virtual Machine Design and Implementation. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4665-8260-X. 300 (est.) pp. LCCN ????

Fu:2017:MCD

[FLZ17] Silvery Fu, Jiangchuan Liu, and Wenwu Zhu. Multimedia content delivery with network function virtualization: The energy perspective. IEEE MultiMedia, 24(3):38–47, July/ September 2017. CODEN IEMUE4. ISSN 1070-986X (print),

1941-0166 (electronic). URL https://www.computer.org/
csdl/mags/mu/2017/03/mmu2017030038-abs.html.

Fei:2020:PWN

[FLZ^{+}20] Xincai Fei, Fangming Liu, Qixia Zhang, Hai Jin, and Hongxin Hu. Paving the way for NFV acceleration: a taxonomy, survey and future directions. ACM Computing Surveys, 53(4): 73:1–73:42, September 2020. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.acm.

org/doi/10.1145/3397022.

Feeley:1990:PVM

[FM90] Marc Feeley and James S. Miller. A parallel virtual machine for efficient Scheme compilation. In ACM [ACM90], pages 119–130. ISBN 0-89791-368-X. LCCN QA 76.73 L23

A24 1990. URL http://www.acm.org/pubs/citations/
proceedings/lfp/91556/p119-feeley/. ACM order no.

552900.

Filho:2018:AOV

[FMIF18] Manoel C. Silva Filho, Claudio C. Monteiro, Pedro R. M. Inácio, and Mário M. Freire. Approaches for optimizing virtual machine placement and migration in cloud environments: a survey. Journal of Parallel and Distributed Computing, 111(??):222–250, January 2018. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151730240X.

Filiposka:2015:CBV

[FMJ15] Sonja Filiposka, Anastas Mishev, and Carlos Juiz. Communitybased VM placement framework. The Journal of Supercomputing, 71(12):4504–4528, December 2015. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11227-015-1546-1.

Fraga:2022:FSD

[FML^{+}22] Martin Fraga, Matías Micheletto, Andrés Llinás, Rodrigo Santos, and Paula Zabala. Flow scheduling in data center networks with time and energy constraints: a software-defined network approach. Future Internet, 14(2):65, February 21, 2022. CODEN ???? ISSN 1999-5903. URL https://www.

mdpi.com/1999-5903/14/2/65.

Feng:2025:BEP

[FMX^{+}25] Yanchang Feng, Junchao Ma, Haojun Xia, Da Zhang, and Bibo Tu. BLFair: enabling proportional I/O sharing for NVMe SSD in SPDK para-virtualization architecture. The Computer Journal, 68(8):953–967, August 2025. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://academic.oup.com/comjnl/article/
68/8/953/8069015.

Forum:1971:VMI

[(Fo71] IBM (Forum). On Virtual Machine Integrity, 1971.

Forum:1978:VMI

[(Fo78] IBM (Forum). On Virtual Machine Integrity, 1978.

Feuser:2014:DOP

[FP14] Johannes Feuser and Jan Peleska. Dependability in open proof software with hardware virtualization — The railway control systems perspective. Science of Computer Programming, 91 (part B)(??):188–215, October 1, 2014. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167642313002001.

Filelis-Papadopoulos:2018:FSL

[FPGK18] Christos K. Filelis-Papadopoulos, George A. Gravvanis, and Panagiotis E. Kyziropoulos. A framework for simulating large scale cloud infrastructures. Future Generation Computer Systems, 79 (part 2)(??):703–714, 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167739X17303230.

Fabian:2006:VE

[FPR^{+}06] Patrick Fabian, Julia Palmer, Justin Richardson, Mic Bowman, Paul Brett, Rob Knauerhase, Jeff Sedayao, John Vicente, Cheng-Chee Koh, and Sanjay Rungta. Virtualization in the enterprise. Intel Technology Journal, 10(3): 227–242, August 10, 2006. ISSN 1535-766X. URL http:

//developer.intel.com/technology/itj/2006/v10i3/6-
enterprise/1-abstract.htm.

Folliot:2002:BFR

[FPS^{+}02] B. Folliot, I. Piumarta, L. Seinturier, C. Baillarguet, C. Khoury, A. Leger, and F. Ogel. Beyond flexibility and reflection: The virtual virtual machine approach. Lecture Notes in Computer Science, 2326:16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2326/23260016.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2326/23260016.pdf.

Femminella:2024:IIT

[FR24] Mauro Femminella and Gianluca Reali. Implementing Internet of Things service platforms with network function virtualization serverless technologies. Future Internet, 16(3):91, March 08, 2024. CODEN ???? ISSN 1999-5903. URL https:

//www.mdpi.com/1999-5903/16/3/91.

Fraser:1983:SFR

[Fra83] Christopher W. Fraser. Surveyor’s forum: Retargetable code generators. ACM Computing Surveys, 15(3):281–283, September 1983. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). See [GFH82, WNL^{+}83, GHF83a, GHF83b].

Franz:1998:JVM

[Fra98] Michael Franz. The Java Virtual Machine: a passing fad? IEEE Software, 15(6):26–29, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (elec-

tronic). URL http://computer.org/software/so1998/
s6026abs.htm;   http://dlib.computer.org/so/books/
so1998/pdf/s6026.pdf.

Fraser:2006:PVC

[Fra06] Stephen Fraser. Pro Visual C++/CLI and the .NET 2.0 platform. The expert’s voice in .NET. Apress, Berkeley, CA, USA, 2006. ISBN 1-4302-0109-6, 1-59059-640-4. xli + 917 pp. LCCN QA76.73.C153 F73 2006eb.

Fraser:2009:PVC

[Fra09] Stephen Fraser. Pro Visual C++/CLI and the .NET 3.5 Platform. Expert’s voice in .NET. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-1053-2 (paperback), 1-4302-1054-0 (electronic). xxx + 1048 pp. LCCN QA76.73.C153 F735 2009.

Flich:2008:PNV

[FRD^{+}08] J. Flich, S. Rodrigo, J. Duato, T. Sødring, Å. G. Solheim, T. Skeie, and O. Lysne. On the potential of NoC virtualization for multicore chips. Scalable Computing: Practice and Experience, 9(3):165–177, September 2008. CO-DEN ???? ISSN 1895-1767. URL http://www.scpe.org/

vols/vol09/no3/SCPE_9_3_02.pdf;   http://www.scpe.
org/vols/vol09/no3/SCPE_9_3_02.zip.

Fredriksen:2005:UUV

[Fre05] Kai Fredriksen. UMLexe — UML virtual machine: a framework for model execution. Masteroppgave i informatikk, Universitetet i Oslo, Oslo, Norway, 2005. 122 pp.

Fan:2015:UCC

[FRM^{+}15] Hua Fan, Aditya Ramaraju, Marlon McKenzie, Wojciech Golab, and Bernard Wong. Understanding the causes of consistency anomalies in Apache Cassandra. Proceedings of the VLDB Endowment, 8(7):810–813, February 2015. CODEN ???? ISSN 2150-8097.

Fu:2024:EDF

[FRM^{+}24] Ying Fu, Meng Ren, Fuchen Ma, Xin Yang, Heyuan Shi, Shanshan Li, and Xiangke Liao. EVMFuzz: Differential fuzz testing of Ethereum virtual machine. Journal of Software: Evolution and Process, 36(4):e2556:1–e2556:??, April 2024. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Froberg:2013:BRP

[Fro13] Scott Froberg. Book review: Private cloud computing: consolidation, virtualization, and service-oriented infrastructure by Stephen R. Smoot, Nam K. Tan. ACM SIGSOFT Software Engineering Notes, 38(2):35, March 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Farrow:1989:VCB

[FS89] Rodney Farrow and Alec G. Stanculescu. A VHDL compiler based on attribute grammar methodology. ACM SIG-PLAN Notices, 24(7):120–130, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org:80/pubs/citations/
proceedings/pldi/73141/p120-farrow/.

Fong:2008:DVS

[FS08] Liana Fong and Małgorzata Steinder. Duality of virtualization: simplification and complexity. Operating Systems Review, 42(1):96–97, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fagin:2011:IPE

[FS11] Barry S. Fagin and Dale J. Skrien. IASSim: a programmable emulator for the Princeton IAS/von Neumann Machine. In Thomas J. Cortina, editor, Proceedings of the 42nd ACM Technical Symposium on Computer Science Education (SIGCSE 11), pages 359–364. ACM Press, New York, NY 10036, USA, 2011. ISBN 1-4503-0500-8 (print). URL http:/

/dl.acm.org/citation.cfm?id=1953163;   http://www.
cs.colby.edu/djskrien/IASSim/. See [FS12] for an analy-

sis and debugging of von Neumann’s computer programs.

Fagin:2012:DSG

[FS12] Barry Fagin and Dale Skrien. Debugging on the shoulders of giants: von Neumann’s programs 65 years later. Computer, 45(11):59–68, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

cs.colby.edu/djskrien/IASSim/. See [FS11] for a descrip-

tion of the emulator on which von Neumann’s programs were run and debugged.

Farshin:2019:MKB

[FS19] Alireza Farshin and Saeed Sharifian. A modified knowledgebased ant colony algorithm for virtual machine placement and simultaneous routing of NFV in distributed cloud architecture. The Journal of Supercomputing, 75(8):5520–5550, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ferreira:2019:DEV

[FSFP19] Bruno Morais Ferreira, Britaldo Silveira Soares-Filho, and Fernando Magno Quintão Pereira. The Dinamica EGO virtual machine. Science of Computer Programming, 173(??):3–20, March 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0167642318300455.

Fukushima:2013:MDR

[FSH^{+}13] Masaki Fukushima, Kohei Sugiyama, Teruyuki Hasegawa, Toru Hasegawa, and Akihiro Nakao. Minimum disclosure routing for network virtualization and its experimental evaluation. IEEE/ACM Transactions on Networking, 21(6):1839– 1851, December 2013. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fuchi:1969:PSP

[FTNY69] K. Fuchi, H. Tanaka, Y. Namago, and T. Yuba. A program simulator by partial interpretation. In ????, editor, 2nd Symposium on Operating Systems Principles, Princeton, NJ, October 1969, pages 97–104. ????, ????, 1969.

Fu:2010:FAR

[Fu10] Song Fu. Failure-aware resource management for highavailability computing clusters with distributed virtual machines. Journal of Parallel and Distributed Computing, 70 (4):384–393, April 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fujimoto:1991:VTM

[Fuj91] Richard M. Fujimoto. The Virtual Time Machine. ACM SIGARCH Computer Architecture News, 19(1):35–44, March 1991. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Fu:2006:SMA

[FX06] Song Fu and Cheng-Zhong Xu. Stochastic modeling and analysis of hybrid mobility in reconfigurable distributed virtual machines. Journal of Parallel and Distributed Computing, 66 (11):1442–1454, November 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fan:2021:RTU

[FXHY21] Xingpeng Fan, Hongli Xu, He Huang, and Xuwei Yang. Realtime update of joint SFC and routing in software defined networks. IEEE/ACM Transactions on Networking, 29(6): 2664–2677, December 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://dl.acm.

org/doi/10.1109/TNET.2021.3095935.

Fan:2023:NEA

[FXL^{+}23] Weibei Fan, Fu Xiao, Mengjie Lv, Lei Han, Junchang Wang, and Xin He. Node essentiality assessment and distributed collaborative virtual network embedding in datacenters. IEEE Transactions on Parallel and Distributed Systems, 34(4): 1265–1280, April 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fang:2024:PPV

[FZD^{+}24] Yaozheng Fang, Zhiyuan Zhou, Surong Dai, Jinni Yang, Hui Zhang, and Ye Lu. PaVM: a parallel virtual machine for smart contract execution and validation. IEEE Transactions on Parallel and Distributed Systems, 35(1):186–202, 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fang:2020:RAV

[FZS^{+}20] Lang Fang, Xiaoning Zhang, Keshav Sood, Yunqing Wang, and Shui Yu. Reliability-aware virtual network function placement in carrier networks. Journal of Network and Computer Applications, 154(??):??, March 15, 2020. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520300102.

Gilbert:1988:TSW

[G^{+}88] P. (Paul) Gilbert et al. Typed Smalltalk working papers, 1988. Report, Department of Computer Science, University of Illinois at Urbana-Champaign, 1304 W. Springfield Ave., Urbana 61801, IL, USA, September 1988. 118 pp.

Gannon:2001:JCC

[G^{+}01] Dennis Gannon et al. Panel discussion: Java, C++, C#, and virtual machines for high performance. In ACM [ACM01b], page ?? ISBN 1-58113-359-6. LCCN QA76.9.O35 A26 2001.

Gibbs:2005:APV

[G^{+}05] G. Benton Gibbs et al., editors. Advanced POWER virtualization on IBM eserver p5 servers. Architecture and performance considerations. IBM redbooks. IBM Corporation, San Jose, CA, USA, second edition, 2005. ISBN 0-7384-9238-8 (paperback). xvi + 412 pp. LCCN QA76.9.V5 A377

2005. URL http://www.loc.gov/catdir/toc/fy0608/
2006275016.html.

Geiselhart:2006:IZV

[G^{+}06] Gregory Geiselhart et al., editors. IBM z/VM and Linux on IBM System z: virtualization cookbook for Red Hat Enterprise Linux 4. Number SG24-7272-00 in IBM redbooks. IBM Corporation, San Jose, CA, USA, 2006. ISBN 0-7384-9495-X (paperback). xiv + 218 pp. LCCN QA76.76.O63 I28 2006.

Gupta:2018:RAV

[GA18] Madnesh K. Gupta and Tarachand Amgoth. Resource-aware virtual machine placement algorithm for IaaS cloud. The Journal of Supercomputing, 74(1):122–140, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Goyal:2025:AMO

[GA25] Sahul Goyal and Lalit Kumar Awasthi. Adaptive multiobjective virtual machine consolidation for energy-efficient cloud data centers. Journal of Grid Computing, 23(2):??, June 2025. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL https://link.springer.com/article/
10.1007/s10723-025-09808-3.

Gordon:2012:EBM

[GAH^{+}12] Abel Gordon, Nadav Amit, Nadav Har’El, Muli Ben-Yehuda, Alex Landau, Assaf Schuster, and Dan Tsafrir. ELI: baremetal performance for I/O virtualization. ACM SIGARCH Computer Architecture News, 40(1):411–422, March 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ASPLOS ’12 conference proceedings.

Grefen:2000:CCO

[GAHL00] P. Grefen, K. Aberer, Y. Hoffner, and H. Ludwig. CrossFlow: Cross-organizational workflow management in dynamic virtual enterprises. International Journal of Computer Systems Science and Engineering, 15(5):??, September 2000. CODEN CSSEEI. ISSN 0267-6192.

Gaines:1975:ACV

[Gai75] Brian R. Gaines. Analogy categories, virtual machines, and structured programming. In Mühlbacher [Müh75], pages 691– 699. CODEN LNCSD9. ISBN 0-387-07410-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.34. English or German.

Galley:1973:PVM

[Gal73] S. W. Galley. PDP-10 virtual machines. In ????, editor, Proceedings ACM SIGARCH-SIGOPS Workshop on Virtual Computer Systems, Cambridge, MA, 1973, page ?? ACM Press, New York, NY 10036, USA, 1973.

Galvin:2009:PATb

[Gal09a] Peter Baer Galvin. Pete’s all things Sun: The Sun virtualization guide. ;login: the USENIX Association newsletter, 34(2):??, April 2009. CODEN LOGNEM. ISSN

1044-6397. URL https://www.usenix.org/publications/
login/april-2009-volume-34-number-2/petes-all-things-
sun-sun-virtualization-guide.

Galvin:2009:PATe

[Gal09b] Peter Baer Galvin. Pete’s all things Sun: VMware vSphere 4 vs. Microsoft Hyper-V R2. ;login: the USENIX Association newsletter, 34(5):58–67, October 2009. CO-DEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/october-2009-volume-
34-number-5/petes-all-things-sun-vmware-vsphere-4-
vs.

Galvin:2011:PATa

[Gal11] Peter Baer Galvin. Pete’s all things Sun: Comparing Solaris to RedHat Enterprise and AIX — virtualization features. ;login: the USENIX Association newsletter, 36(1):??, February 2011. CODEN LOGNEM. ISSN 1044-6397. URL https:

//www.usenix.org/publications/login/february-2011-
volume-36-number-1/petes-all-things-sun-comparing-
solaris-redhat.

Gill:2019:TFD

[GB19] Sukhpal Singh Gill and Rajkumar Buyya. A taxonomy and future directions for sustainable cloud computing: 360 degree view. ACM Computing Surveys, 51(5):104:1–104:??, January 2019. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341

(electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3241038.

Gu:2000:EHP

[GBCW00] W. Gu, N. A. Burns, M. T. Collins, and W. Y. P. Wong. The evolution of a high-performing Java virtual machine. IBM Systems Journal, 39(1):135–150, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.research.

ibm.com/journal/sj/391/gu.html.

Gayer:1987:CPA

[GBO87] Richard Gayer, Catherine M. Beise, and G. Scott Owen. Conversion of a PDP-11/40 assembler and simulator from mainframe Pascal to Ada on IBM PC microcomputers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 19(1):378–381, February 1987. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gough:2000:EJV

[GC00] K. John Gough and Diane Corney. Evaluating the Java Virtual Machine as a target for languages other than Java. In Modular Programming Languages, pages 278–290. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-44519-6. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/10722581_22.

Gonzalez-Castano:2001:JCV

[GCARPC^{+}01] F. J. González-Castaño, L. Anido-Rifón, J. M. Pousada-Carballo, P. S. Rodríguez-Hernández, and R. López-Gómez. A Java/CORBA virtual machine architecture for remote execution of optimization solvers in heterogeneous networks. Software—Practice and Experience, 31(1):1–16, January 2001. CODEN SPEXBL. ISSN 0038-0644 (print),

1097-024X (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/76502341/START;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=
76502341&PLACEBO=IE.pdf.

Gong:2021:TLS

[GCL^{+}21] X. Gong, D. Cao, Y. Li, X. Liu, Y. Li, J. Zhang, and T. Li. A thread level SLO-aware I/O framework for embedded virtualization. IEEE Transactions on Parallel and Distributed Systems, 32(3):500–513, March 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Goldweber:2008:VEE

[GD08] Michael Goldweber and Renzo Davoli. VDE: an emulation environment for supporting computer networking courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):138–142, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Gasiunas:2017:FBA

[GDSA^{+}17] Vaidas Gasiunas, David Dominguez-Sal, Ralph Acker, Aharon Avitzur, Ilan Bronshtein, Rushan Chen, Eli Ginot, Norbert Martinez-Bazan, Michael Müller, Alexander Nozdrin, Weijie Ou, Nir Pachter, Dima Sivov, and Eliezer Levy. Fiber-based architecture for NFV cloud databases. Proceedings of the VLDB Endowment, 10(12):1682–1693, August 2017. CODEN ???? ISSN 2150-8097.

Gaudiot:1985:PES

[GE85] J. L. Gaudiot and M. D. Ercegovac. Performance evaluation of a simulated data-flow computer with low-resolution actors. Journal of Parallel and Distributed Computing, 2(4):321–351, November 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Geist:2002:PVM

[Gei02] Al Geist. Petascale virtual machine: Computing on 100,000 processors. Lecture Notes in Computer Science, 2474:6–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.de/link/
service/series/0558/bibs/2474/24740006.htm; http:/
/link.springer.de/link/service/series/0558/papers/
2474/24740006.pdf.

Ghaseminya:2025:ACV

[GEK25] Mohammad Mahdi Ghaseminya, Elahe Eslami, and Seyed Mehdi Karbassi. Advancing cloud virtualization: a comprehensive survey on integrating IoT, Edge, and Fog computing with FaaS for heterogeneous smart environments. The Journal of Supercomputing, 81(14):??, September 2025. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-025-07799-2.

Genter:1986:UVM

[Gen86] Robert E. (Rick) Genter. Unix as a virtual machine environment. In USENIX Association [USE86], pages 475–485.

Garzon:1992:DTG

[GFB^{+}92] Max H. Garzon, Stanley P. Franklin, William Baggett, William S. Boyd, Jr., and Dinah Dickerson. Design and testing of a general-purpose neurocomputer. Journal of Parallel and Distributed Computing, 14(3):203–220, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ganapathi:1982:RCC

[GFH82] Mahadevan Ganapathi, Charles N. Fischer, and John L. Hennessy. Retargetable compiler code generation. ACM Computing Surveys, 14(4):573–592, December 1982. CODEN CMSVAN. ISSN 0010-4892. See also [WNL^{+}83, GHF83a, Fra83, GHF83b].

Gheibi-Fetrat:2025:SSS

[GFHML^{+}25] Atiyeh Gheibi-Fetrat, Fatemeh Serajeh Hassani, Masoud Mohammadi-Lak, Amir Mirzaei, Negar Akbarzadeh, Mahmoud Reza Kheyrati-Fard, Mohammad Hosseini, Ahmad Javadi Nezhad, Arash Tavakkol, Jeong A. Lee, and Hamid Sarbazi-Azad. A survey of SSD simulators and emulators. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL https://link.springer.com/article/
10.1007/s11227-025-07078-0.

Gagliardi:1972:VA

[GG72] U. O. Gagliardi and R. P. Goldberg. Virtualizable architectures. In ????, editor, Proceedings of 1972 ACMAICA International Comp. Symposium, Venice, Italy, April 1972, pages 527–538. ????, ????, 1972.

Greamo:2011:SVM

[GG11] Chris Greamo and Anup Ghosh. Sandboxing and virtualization: Modern tools for combating malware. IEEE Security & Privacy, 9(2):79–82, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

GomezMartin:2003:JVE

[GGG03] P. P. GomezMartin, M. A. GomezMartin, and P. A. GonzalezCalero. Javy: Virtual environment for case-based teaching of Java Virtual Machine. Lecture Notes in Computer Science, 2773:906–913, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gjessing:1992:MRL

[GGJ^{+}92] S. Gjessing, D. B. Gustavson, D. V. James, G. Stone, and H. Wiggers. Memory — a RAM link for high speed. IEEE Spectrum, 29(10):52–53, October 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gunleifsen:2018:TCP

[GGK18] Håkon Gunleifsen, Vasileios Gkioulos, and Thomas Kemmerich. A tiered control plane model for service function chaining isolation. Future Internet, 10(6):46, June 04, 2018. CODEN ???? ISSN 1999-5903. URL https://www.mdpi.

com/1999-5903/10/6/46.

Gerangelos:2019:EAS

[GGK19] Stefanos Gerangelos, Georgios Goumas, and Nectarios Koziris. Efficient accelerator sharing in virtualized environments: a Xeon Phi use-case. The Journal of Systems and Software, 150(??):37–50, April 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121218302863.

Garcia:2016:CMP

[GGM^{+}16] P. Garcia, T. Gomes, J. Monteiro, A. Tavares, and M. Ekpanyapong. On-chip message passing sub-system for embedded inter-domain communication. IEEE Computer Architecture Letters, 15(1):33–36, January/June 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Gao:2013:MOA

[GGQ^{+}13] Yongqiang Gao, Haibing Guan, Zhengwei Qi, Yang Hou, and Liang Liu. A multi-objective ant colony system algorithm for virtual machine placement in cloud computing. Journal of Computer and System Sciences, 79(8):1230–1242, December 2013. CODEN JCSSBM. ISSN 0022-0000 (print),

1090-2724 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0022000013000627.

Gdaniec:1991:VDS

[GH91a] J. M. Gdaniec and J. P. Hennessy. VM data spaces and ESA/XC facilities. IBM Systems Journal, 30(1):14–33, 1991. CODEN IBMSA7. ISSN 0018-8670.

Greenfield:1991:PIT

[GH91b] P. G. Greenfield and R. J. Hendley. A proposed intelligent tutoring system framework incorporating persistent logic programming. Technical Report CSR-91-1, University of Birmingham, School of Computer Science, February 1991.

Ghasemi:2020:MOL

[GH20] Arezoo Ghasemi and Abolfazl Toroghi Haghighat. A multiobjective load balancing algorithm for virtual machine placement in cloud data centers based on machine learning. Computing: Archiv für Informatik und Numerik, 102(9):2049– 2072, September 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ghosh:2012:RAA

[GHD12] Sudeep Ghosh, Jason Hiser, and Jack W. Davidson. Replacement attacks against VM-protected applications. ACM SIG-PLAN Notices, 47(7):203–214, July 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Ganapathi:1983:SFRa

[GHF83a] M. Ganapathi, J. L. Hennessy, and C. N. Fischer. Surveyor’s forum: Retargetable code generators. ACM Computing Surveys, 15(3):280–281, September 1983. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). See [GFH82, WNL^{+}83, Fra83, GHF83b].

Ganapathi:1983:SFRb

[GHF83b] M. Ganapathi, J. L. Hennessy, and C. N. Fischer. Surveyor’s forum: Retargetable code generators. ACM Computing Surveys, 15(3):283–284, September 1983. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). See [GFH82, WNL^{+}83, GHF83a, Fra83].

Grebe:1993:TAS

[GHH^{+}93] R. Grebe, J. Hektor, S. C. Hilton, M. R. Jane, and P. H. Welch, editors. Transputer applications and systems ’93: proceedings of the 1993 World Transputer Congress, 20–22 September 1993, Aachen, Germany. IOS Press, Amsterdam, The Netherlands, 1993. ISBN 90-5199-140-1. LCCN ????

Ghasemi:2024:EVM

[GHK24] Arezoo Ghasemi, Abolfazl Toroghi Haghighat, and Amin Keshavarzi. Enhancing virtual machine placement efficiency in cloud data centers: a hybrid approach using multiobjective reinforcement learning and clustering strategies. Computing: Archiv für Informatik und Numerik, 106(9): 2897–2922, September 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https://link.

springer.com/article/10.1007/s00607-024-01311-z.

Gupta:2018:SCS

[GHM^{+}18] Abhishek Gupta, M. Farhan Habib, Uttam Mandal, Pulak Chowdhury, Massimo Tornatore, and Biswanath Mukherjee. On service-chaining strategies using Virtual Network Functions in operator networks. Computer Networks (Amsterdam, Netherlands: 1999), 133(??):1–16, March 14, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128618300379.

Gandhi:2016:APE

[GHS16] Jayneel Gandhi, Mark D. Hill, and Michael M. Swift. Agile paging: exceeding the best of nested and shadow paging. ACM SIGARCH Computer Architecture News, 44(3): 707–718, June 2016. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Gandhi:2017:APE

[GHS17] Jayneel Gandhi, Mark D. Hill, and Michael M. Swift. Agile paging for efficient memory virtualization. IEEE Micro, 37 (3):80–86, May/June 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL https://www.

computer.org/csdl/mags/mi/2017/03/mmi2017030080-abs.
html.

Gerofi:2012:ETT

[GI12] Balazs Gerofi and Yutaka Ishikawa. Enhancing TCP throughput of highly available virtual machines via speculative communication. ACM SIGPLAN Notices, 47(7):87–96, July 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Gilesh:2020:OLM

[GJK^{+}20] Malayam Parambath Gilesh, Subham Jain, S. D. Madhu Kumar, Lillykutty Jacob, and Umesh Bellur. Opportunistic live migration of virtual machines. Concurrency and Computation: Practice and Experience, 32(5):e5477:1–e5477:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Griffith:2005:MME

[GK05] Rean Griffith and Gail Kaiser. Manipulating managed execution runtimes to support self-healing systems. ACM SIG-SOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Guo:2015:PBL

[GKBB15] Fei Guo, Seongbeom Kim, Yury Baskakov, and Ishan Banerjee. Proactively breaking large pages to improve memory overcommitment performance in VMware ESXi. ACM SIG-PLAN Notices, 50(7):39–51, July 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:2013:VMC

[gKEY13] Shin gyu Kim, Hyeonsang Eom, and Heon Y. Yeom. Virtual machine consolidation based on interference modeling. The Journal of Supercomputing, 66(3):1489–1506, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL http://link.springer.com/article/
10.1007/s11227-013-0939-2.

Guo:2019:SSA

[GKJ^{+}19] Wenxia Guo, Ping Kuang, Yaqiu Jiang, Xiang Xu, and Wenhong Tian. SAVE: self-adaptive consolidation of virtual machines for energy efficiency of CPU-intensive applications in the cloud. The Journal of Supercomputing, 75(11):7076–7100, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gec:2019:SAM

[GKP^{+}19] Sandi Gec, Dragi Kimovski, Uros Pascinski, Radu Prodan, and Vlado Stankovski. Semantic approach for multi-objective optimisation of the ENTICE distributed Virtual Machine and container images repository. Concurrency and Computation: Practice and Experience, 31(3):e4264:1–e4264:??, February 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Geist:1999:HAV

[GKSP99] G. A. Geist II, James Arthur Kohl, Stephen Scott, and Philip M. Papadopoulos. Harness: Adaptable virtual machine environment for heterogeneous cluster. Parallel Processing Letters, 9(2):253–??, June 1999. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Gschwind:2017:OED

[GKT17] M. Gschwind, T. Kaldewey, and D. K. Tam. Optimizing the efficiency of deep learning through accelerator virtualization. IBM Journal of Research and Development, 61(4): 12:1–12:11, ???? 2017. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://ieeexplore.

ieee.org/document/8030299/.

Gamage:2013:PRO

[GKXK13] Sahan Gamage, Ramana Rao Kompella, Dongyan Xu, and Ardalan Kangarlou. Protocol responsibility offloading to improve TCP throughput in virtualized environments. ACM Transactions on Computer Systems, 31(3):7:1–7:??, August 2013. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Gaspar:2008:RVC

[GLA^{+}08] Alessio Gaspar, Sarah Langevin, William Armitage, R. Sekar, and T. Daniels. The role of virtualization in computing education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):131–132, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Guerrero:2018:MOO

[GLBJ18] Carlos Guerrero, Isaac Lera, Belen Bermejo, and Carlos Juiz. Multi-objective optimization for virtual machine allocation and replica placement in virtualized Hadoop. IEEE Transactions on Parallel and Distributed Systems, 29(11):2568–2581, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print),

1558-2183 (electronic). URL https://www.computer.org/
csdl/trans/td/2018/11/08360490-abs.html.

Gold:1984:KR

[GLC84] B. D. Gold, R. R. Linde, and P. F. Cudney. KVM/370 in retrospect. In IEEE [IEE84a], pages 13–23. ISBN 0-8186-0532-4 (paperback), 0-8186-8532-8 (hard), 0-8186-4532-6 (microfiche). LCCN QA76.9.A25 S95 1984.

Ghumre:2012:ENC

[GLK^{+}12] Pooja Ghumre, Junwei Li, Mukil Kesavan, Ada Gavrilovska, and Karsten Schwan. Evaluating the need for complexity in energy-aware management for cloud platforms. ACM SIG-METRICS Performance Evaluation Review, 40(3):23–27, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Geissler:2021:DTM

[GLL^{+}21] Stefan Geissler, Stanislav Lange, Leonardo Linguaglossa, Dario Rossi, Thomas Zinner, and Tobias Hossfeld. Discretetime modeling of NFV accelerators that exploit batched processing. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 6(3):11:1– 11:27, September 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://dl.acm.org/

doi/10.1145/3488243.

Guo:2016:FNB

[GLLJ16] Jian Guo, Fangming Liu, John C. S. Lui, and Hai Jin. Fair network bandwidth allocation in IaaS datacenters via a cooperative game approach. IEEE/ACM Transactions on Networking, 24(2):873–886, April 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gu:2013:VIV

[GLQ^{+}13] Dujuan Gu, Xiaohan Liu, Gang Qin, Shuangjian Yan, Ze luo, and Baoping Yan. VNET6: IPv6 virtual network for the collaboration between applications and networks. Journal of Network and Computer Applications, 36(6):1579–1588, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print),

1095-8592 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1084804513000490.

Gupta:2015:HER

[GLS15] Vishal Gupta, Min Lee, and Karsten Schwan. HeteroVisor: Exploiting resource heterogeneity to enhance the elasticity of cloud platforms. ACM SIGPLAN Notices, 50(7):79–92, July 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grimaud:1999:FTI

[GLV99] Gilles Grimaud, Jean-Louis Lanet, and Jean-Jacques Vandewalle. FACADE: a typed intermediate language dedicated to smart cards. ACM SIGSOFT Software Engineering Notes, 24(6):476–493, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gupta:2009:DE

[GLV^{+}09] Diwaker Gupta, Sangmin Lee, Michael Vrable, Stefan Savage, Alex C. Snoeren, George Varghese, Geoffrey M. Voelker, and Amin Vahdat. Difference engine. ;login: the USENIX Association newsletter, 34(2):24–31, April 2009. CODEN LOGNEM. ISSN 1044-6397. URL https://www.usenix.

org/publications/login/april-2009-volume-34-number-
2/difference-engine.

Gupta:2010:DEH

[GLV^{+}10] Diwaker Gupta, Sangmin Lee, Michael Vrable, Stefan Savage, Alex C. Snoeren, George Varghese, Geoffrey M. Voelker, and Amin Vahdat. Difference engine: harnessing memory redundancy in virtual machines. Communications of the ACM, 53(10):85–93, October 2010. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Guo:2023:AAV

[GLW23] Lifeng Guo, Changhong Lu, and Guanlin Wu. Approximation algorithms for a virtual machine allocation problem with finite types. Information Processing Letters, 180(??):Article 106339, February 2023. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0020019022000965.

Gentile:2024:OVP

[GMGF24] Antonio Francesco Gentile, Davide Macrì, Emilio Greco, and Peppino Fazio. Overlay and virtual private networks security performances analysis with open source infrastructure deployment. Future Internet, 16(8):??, August 7, 2024. CO-DEN ???? ISSN 1999-5903. URL https://www.mdpi.com/

1999-5903/16/8/283.

Garg:2017:CGA

[GMK17] Anshuj Garg, Debadatta Mishra, and Purushottam Kulkarni. Catalyst: GPU-assisted rapid memory deduplication in virtualization environments. ACM SIGPLAN Notices, 52(7): 44–59, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Giacalone:1989:FSI

[GMP89] Alessandro Giacalone, Prateek Mishra, and Sanjiva Prasad. FACILE: a symmetric integration of concurrent and functional programming. International Journal of Parallel Programming, 18(2):121–160, April 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&
issn=0885-7458&volume=18&issue=2&spage=121. Also in

TAPSOFT ’89, ed. J. Diaz and F. Orejas, pp. 184-209, Springer-Verlag, Lecture Notes in Computer Science 352 (1989).

Golub:1993:MER

[GMR93] D. B. Golub, R. Manikundalam, and F. L. Rawson. MVM— an environment for running multiple DOS, Windows and DPMI programs on the microkernel. In USENIX [USE93], pages 173–190. ISBN 1-880446-51-0. LCCN QA 76.76 O63

U86 1993. URL http://www.usenix.org/publications/
library/proceedings/mobile93/.

Guanciale:2016:PSM

[GNDB16] Roberto Guanciale, Hamed Nemati, Mads Dam, and Christoph Baumann. Provably secure memory isolation for Linux on ARM. Journal of Computer Security, 24(6):793– 837, ???? 2016. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gupta:2024:SVL

[GNK24] Ambika Gupta, Suyel Namasudra, and Prabhat Kumar. A secure VM live migration technique in a cloud computing environment using blowfish and blockchain technology. The Journal of Supercomputing, 80(19):27370–27393, December 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL https://link.springer.com/article/
10.1007/s11227-024-06461-7.

Goldberg:1969:VMS

[Gol69] R. P. Goldberg. Virtual machine systems. Report MS-2687 (also 28L-0036), MIT Lincoln Laboratory, Lexington, MA, USA, September 1969.

Goldberg:1971:HRV

[Gol71a] R. P. Goldberg. Hardware requirements for virtual machine systems. In ????, editor, HICSS-4, Hawaii International Conference on System Sciences, Honolulu, January 1971, page ?? ????, ????, 1971.

Goldberg:1971:VMS

[Gol71b] R. P. Goldberg. Virtual machines: Semantics and examples. In ????, editor, Proceedings IEEE Computer Society Conference, Boston, MA, September 1971, pages 141–142. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1971.

Goldberg:1973:AVM

[Gol73a] R. P. Goldberg. Architecture of virtual machines. In Proceedings of the ACM Workshop on Virtual Computer Systems, pages 74–112. ACM Press, New York, NY 10036,

USA, 1973. URL http://portal.acm.org/citation.cfm?
id=803950&dl=ACM&coll=portal#.

Goldberg:1973:VMA

[Gol73b] R. P. Goldberg. Virtual machines architecture. Honeywell Computer Journal, 7(4):??, ???? 1973. CODEN HNCJA3. ISSN 0046-7847.

Goldberg:1973:APV

[Gol73c] Robert Philip Goldberg. Architectural Principles for Virtual Computer Systems. Ph.D. thesis, Division of Engineering and Applied Physics Harvard University, Cambridge, MA,

USA, 1973. URL https://www.proquest.com/pqdtglobal/
docview/302694242.

Goldberg:1974:SVM

[Gol74] Robert P. Goldberg. Survey of virtual machine research. Computer, 7(6):34–45, June 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goth:2007:VOT

[Got07] Greg Goth. Virtualization: Old technology offers huge new potential. IEEE Distributed Systems Online, 8(2):??, February 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683

(electronic). URL http://csdl.computer.org/comp/mags/
ds/2007/02/o2003.pdf.

Ganegedara:2013:CPA

[GP13] Thilan Ganegedara and Viktor Prasanna. A comprehensive performance analysis of virtual routers on FPGA. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(2):9:1–9:??, July 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Gutierrez:2021:RTP

[GPM21] Juan Roberto López Gutiérrez, Pedro Ponce, and Arturo Molina. Real-time power electronics laboratory to strengthen distance learning engineering education on smart grids and microgrids. Future Internet, 13(9):237, September 17, 2021. CODEN ???? ISSN 1999-5903. URL https://www.mdpi.

com/1999-5903/13/9/237.

Gudkov:2023:BRB

[GPR23] Andrei Gudkov, Pavel Popov, and Stepan Romanov. Bal-Con — resource balancing algorithm for VM consolidation. Future Generation Computer Systems, 147(??):265–274, October 2023. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167739X23001760.

Gad:2018:ZMD

[GPS^{+}18] Ramy Gad, Simon Pickartz, Tim Süß, Lars Nagel, Stefan Lankes, Antonello Monti, and André Brinkmann. Zeroing memory deallocator to reduce checkpoint sizes in virtualized HPC environments. The Journal of Supercomputing, 74(11):

6236–6257, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gregg:2003:PID

[GPW03] David Gregg, James Power, and John Waldron. Platform independent dynamic Java virtual machine analysis: the Java Grande Forum benchmark suite. Concurrency and Computation: Practice and Experience, 15(3–5):459–484, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Groves:1980:DVM

[GR80] L. J. Groves and W. J. Rogers. The design of a virtual machine for Ada. ACM SIGPLAN Notices, 15(11):223–234, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gupta:2015:LBO

[GR15] Varun Gupta and Ana Radovanovic. Lagrangian-based online stochastic bin packing. ACM SIGMETRICS Performance Evaluation Review, 43(1):467–468, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Gao:2020:CMS

[GR20] Lingnan Gao and George N. Rouskas. Congestion minimization for service chain routing problems with path length considerations. IEEE/ACM Transactions on Networking, 28(6): 2643–2656, December 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://dl.acm.

org/doi/10.1109/TNET.2020.3017792.

Green:2010:SUS

[Gre10] David Green. The Sydney University SILLIAC. Web site,

August 14, 2010. URL http://members.iinet.net.au/~

dgreen/silliac.html. The SILLIAC was the first computer installed at Sydney University, and was operational from 1956 to 1968. The Web site links to the SILLIAC Emulator, a C program for Microsoft Windows.

Garg:2022:OVM

[GSG22] Neha Garg, Damanpreet Singh, and Major Singh Goraya. Optimal virtual machine scheduling in virtualized cloud environment using VIKOR method. The Journal of Supercomputing, 78(4):6006–6034, March 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.com/article/10.1007/s11227-
021-04081-z.

Gilesh:2018:SSV

[GSKJ18] M. P. Gilesh, Sanjay Satheesh, S. D. Madhu Kumar, and Lillykutty Jacob. Selecting suitable virtual machine migrations for optimal provisioning of virtual data centers. ACM SIGAPP Applied Computing Review, 18(2):22–32, July 2018. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3243064.3243066.

Grimshaw:1993:DOP

[GSN93] Andrew S. Grimshaw, W. Timothy Strayer, and P. Narayan. Dynamic, object-oriented parallel processing. IEEE parallel and distributed technology: systems and applications, 1(2):33– 47, May 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Grimmer:2018:CLI

[GSS^{+}18] Matthias Grimmer, Roland Schatz, Chris Seaton, Thomas Würthinger, and Mikel Luján. Cross-language interoperability in a multi-language runtime. ACM Transactions on Programming Languages and Systems, 40(2):8:1–8:??, June 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gupta:2017:HCS

[GSW^{+}17] Abhishek Gupta, Rick Spillane, Wenguang Wang, Maxime Austruy, Vahid Fereydouny, and Christos Karamanolis. Hybrid cloud storage: Bridging the gap between compute clusters and cloud storage. Operating Systems Review, 51(1):48– 53, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Garg:2014:SBV

[GTGB14] Saurabh Kumar Garg, Adel Nadjaran Toosi, Srinivasa K. Gopalaiyengar, and Rajkumar Buyya. SLA-based virtual machine management for heterogeneous workloads in a cloud datacenter. Journal of Network and Computer Applications, 45(??):108–120, October 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804514001787.

Gilbert:2006:IVG

[GTN^{+}06] Laura Gilbert, Jeff Tseng, Rhys Newman, Saeed Iqbal, Ronald Pepper, Onur Celebioglu, Jenwei Hsieh, Victor Mashayekhi, and Mark Cobban. Implications of virtualization on Grids for high energy physics applications. Journal of Parallel and Distributed Computing, 66(7):922–930, July 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gidra:2015:NGC

[GTS^{+}15] Lokesh Gidra, Gaël Thomas, Julien Sopena, Marc Shapiro, and Nhan Nguyen. NumaGiC: a garbage collector for big data on big NUMA machines. ACM SIGPLAN Notices, 50 (4):661–673, April 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guan:2014:HHV

[Gua14] Haibing Guan. HYVI: A HYbrid VIrtualization solution balancing performance and manageability. IEEE Transactions on Parallel and Distributed Systems, 25(9):2332–2341, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.computer.org/

csdl/trans/td/2014/09/06605688-abs.html.

Gum:1983:SEA

[Gum83] Peter H. Gum. System/370 extended architecture: Facilities for virtual machines. IBM Journal of Research and Development, 27(6):530–544, November 1983. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Guyer:2014:UJT

[Guy14] Samuel Z. Guyer. Use of the JVM at Twitter: a bird’s eye view. ACM SIGPLAN Notices, 49(11):1, November 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guzdial:2001:UST

[Guz01] Mark Guzdial. Using squeak for teaching user interface software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):219–223, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gerofi:2013:UMC

[GVI13] Balazs Gerofi, Zoltan Vass, and Yutaka Ishikawa. Utilizing memory content similarity for improving the performance of highly available virtual machines. Future Generation Computer Systems, 29(4):1085–1095, June 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X12001446.

Garfinkel:2007:WVC

[GW07] Tal Garfinkel and Andrew Warfield. What virtualization can do for security. ;login: the USENIX Association newsletter, 32(6):??, December 2007. CODEN LOGNEM. ISSN

1044-6397. URL https://www.usenix.org/publications/
login/december-2007-volume-32-number-6/what-virtualization-
can-do-security.

Gao:2016:VCD

[GWZ16] Xianming Gao, Baosheng Wang, and Xiaozhe Zhang. VR-Cluster: Dynamic migration for resource fragmentation problem in virtual router platform. Scientific Programming, 2016 (??):3976965:1–3976965:14, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https:

//www.hindawi.com/journals/sp/2016/3976965/.

Hartung:1979:VE

[HA79] R. L. Hartung and A. Ammerman. Virtual I/O, an experiment. ACM SIGMICRO Newsletter, 10(4):109–113, December 1979. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X.

URL https://dl.acm.org/doi/10.1145/1014188.803019.

Habib:2006:X

[Hab06] Irfan Habib. Xen. Linux Journal, 2006(145):??, May 2006. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic). URL http://www.cl.cam.ac.uk/Research/SRG/

netos/xen/documentation.html;   http://www.cl.cam.
ac.uk/xeno; http://www.xensource.com/.

Henry:2022:HMH

[HAK22] Niroshini Infantia Henry, C Anbuananth, and S Kalarani. Hybrid meta-heuristic algorithm for optimal virtual machine placement and migration in cloud computing. Concurrency and Computation: Practice and Experience, 34(28):e7353:1– e7353:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Halstead:1979:RTN

[Hal79] R. H. Halstead. Reference Tree Networks: Virtual Machine and Implementation. Thesis (Ph.D.), Department of Electrical Engineering and Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, 1979. 252 pp. Also issued as Technical Report MIT/LCS/TR-222.

Haletky:2008:VES

[Hal08] Edward Haletky. VMware ESX server in the enterprise: planning and securing virtualization servers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2008. ISBN 0-13-230207-1 (paperback). xxiv + 551 pp. LCCN QA76.9.V5 H35

2008. URL http://www.loc.gov/catdir/toc/ecip083/
2007044443.html.

Haletky:2009:VVV

[Hal09] Edward Haletky. VMware vSphere and virtual infrastructure security: securing the virtual environment. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2009. ISBN 0-13-715800-9 (paperback). xxx + 521 pp. LCCN QA76.9.V5 H36 2009.

Hamlet:1976:PBT

[Ham76] Richard Hamlet. The PDP-11 as B5500 in teaching systems programming. ACM SIGPLAN Notices, 11(5):47–52, May 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hammersley:2007:PVS

[Ham07] Eric Hammersley. Professional VMware Server. Wiley, New York, NY, USA, 2007. ISBN 0-470-07988-6. xviii + 437 pp. LCCN QA76.76.O63 H344637 2007. URL http://

www.loc.gov/catdir/enhancements/fy0741/2006033472-
b.html;   http://www.loc.gov/catdir/enhancements/
fy0741/2006033472-d.html;   http://www.loc.gov/catdir/
toc/ecip072/2006033472.html.

Hans:1973:CAM

[Han73] C. Hans. Contribution a l’Architecture de Mécanismes Elémentaires Pour Certains Systèmes Générateurs de Machines Virtuelles. (French) [Contribution to the Architecture of Elementary Mechanisms for Certain Virtual Machine Generator Systems]. Thèse, L’Université Scientifique et Medicale de Grenoble, Grenoble, France, November 24, 1973.

Hansen:2005:IJP

[Han05] Stuart A. Hansen. Interpreting Java program runtimes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):36–40, March 2005. CO-DEN SIGSD3. ISBN 1-58113-997-7. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2005.
bib.

Hand:2016:TPH

[Han16] Steve Hand. Technical perspective: High-performance virtualization: are we done? Communications of the ACM, 59 (1):107, January 2016. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://cacm.acm.org/

magazines/2015/1/195736/fulltext.

Hartmann:1977:CPC

[Har77] A. C. Hartmann. A Concurrent Pascal Compiler for Mini-Computers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977.

Hulaas:2008:PTL

[HB08] Jarle Hulaas and Walter Binder. Program transformations for light-weight CPU accounting and control in the Java Virtual Machine. Higher-Order and Symbolic Computation, 21(1–2):119–146, June 2008. CODEN LSCOEX.

ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&
issn=1388-3690&volume=21&issue=1&spage=119.

Huang:2012:PEN

[HB12] Shu Huang and Ilia Baldine. Performance evaluation of 10GE NICs with SR-IOV support: I/O virtualization and network stack optimizations. Lecture Notes in Computer Science, 7201:197–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28540-0_14/

.

Huang:2013:ECS

[HB13] Jipeng Huang and Michael D. Bond. Efficient context sensitivity for dynamic analyses via calling context uptrees and customized memory management. ACM SIGPLAN Notices, 48(10):53–72, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOP-SLA ’13 conference proceedings.

Hurlburt:2014:BBC

[HB14] G. F. Hurlburt and I. Bojanova. Bitcoin: Benefit or curse? IT Professional, 16(3):10–15, May 2014. CODEN IPMAFM. ISSN 1520-9202 (print), 1941-045x (electronic).

Hetzelt:2017:SAE

[HB17] Felicitas Hetzelt and Robert Buhren. Security analysis of encrypted virtual machines. ACM SIGPLAN Notices, 52(7): 129–142, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hoang:2010:CAN

[HBL^{+}10] Giang Hoang, Chang Bae, John Lange, Lide Zhang, Peter Dinda, and Russ Joseph. A case for alternative nested paging models for virtualized systems. IEEE Computer Architecture Letters, 9(1):17–20, January/June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Haris:2024:COP

[HBNK24a] Raseena M. Haris, Mahmoud Barhamgi, Armstrong Nhlabatsi, and Khaled M. Khan. Correction to: Optimizing precopy live virtual machine migration in cloud computing using machine learning-based prediction model. Computing: Archiv für Informatik und Numerik, 106(12):4249, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057

(electronic). URL https://link.springer.com/article/
10.1007/s00607-024-01350-6. See [HBNK24b].

Haris:2024:OPC

[HBNK24b] Raseena M. Haris, Mahmoud Barhamgi, Armstrong Nhlabatsi, and Khaled M. Khan. Optimizing pre-copy live virtual machine migration in cloud computing using machine learning-based prediction model. Computing: Archiv für Informatik und Numerik, 106(9):3031–3062, September 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL https://link.springer.com/article/10.

1007/s00607-024-01318-6. See correction [HBNK24a].

Huang:2006:PMA

[HBP06] Mark Huang, Andy Bavier, and Larry Peterson. PlanetFlow: maintaining accountability for network services. Operating Systems Review, 40(1):89–94, January 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Huang:2012:VAJ

[HC12] Ting-Chang Huang and Da-Wei Chang. VM aware journaling: improving journaling file system performance in virtualization environments. Software—Practice and Experience, 42 (3):303–330, March 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hankendi:2017:SCS

[HC17] Can Hankendi and Ayse Kivilcim Coskun. Scale & cap: Scaling-aware resource management for consolidated multithreaded applications. ACM Transactions on Design Automation of Electronic Systems., 22(2):30:1–30:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2018:PSC

[HC18] Sheng-Min Huang and Li-Pin Chang. Providing SLO compliance on NVMe SSDs through parallelism reservation. ACM Transactions on Design Automation of Electronic Systems., 23(3):28:1–28:??, April 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hsieh:2018:VMA

[HCB18] Han-Chuan Hsieh, Jiann-Liang Chen, and Abderrahim Benslimane. 5G virtualized multi-access edge computing platform for IoT applications. Journal of Network and Computer Applications, 115(??):94–102, August 1, 2018. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301577.

Hizver:2014:RTD

[HcC14] Jennia Hizver and Tzi cker Chiueh. Real-time deep virtual machine introspection and its applications. ACM SIGPLAN Notices, 49(7):3–14, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:2007:ETT

[HCJ07] Jacob Gorm Hansen, Eske Christiansen, and Eric Jul. Evil twins: two models for TCB reduction in HPC clusters. Operating Systems Review, 41(4):20–29, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hale:2016:EHP

[HD16] Kyle C. Hale and Peter A. Dinda. Enabling hybrid parallel runtimes through kernel and virtualization support. ACM SIGPLAN Notices, 51(7):161–175, July 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hines:2009:PCL

[HDG09] Michael R. Hines, Umesh Deshpande, and Kartik Gopalan. Post-copy live migration of virtual machines. Operating Systems Review, 43(3):14–26, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hu:2008:SVO

[HDM08] Wenjin Hu, Todd Deshane, and Jeanna Matthews. Solaris virtualization options. ;login: the USENIX Association newsletter, 33(5):??, October 2008. CODEN LOGNEM. ISSN

1044-6397. URL https://www.usenix.org/publications/
login/october-2008-volume-33-number-5/solaris-virtualization-
options.

Heege:2007:ECC

[Hee07] Marcus Heege. Expert C++/CLI: .NET for Visual C++ programmers. The expert’s voice in .NET; Books for professionals by professionals. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-756-7. xvii + 330 pp. LCCN QA76.73.C153 H44 2007.

Herrod:2006:FVT

[Her06] Steve Herrod. The future of virtualization technology. ACM SIGARCH Computer Architecture News, 34(2):352, 2006. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Herrod:2010:SRD

[Her10] Stephen Alan Herrod. Systems research and development at VMware. Operating Systems Review, 44(4):1–2, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hendricks:1979:EVM

[HH79] Edson C. Hendricks and Timothy C. Hartmann. Evolution of a virtual machine subsystem. IBM Systems Journal, 18(1): 111–142, 1979. CODEN IBMSA7. ISSN 0018-8670.

Ho:2005:DPD

[HH05] Alex Ho and Steven Hand. On the design of a pervasive debugger. In ACM [ACM05a], pages 117–122. ISBN 1-59593-050-7. LCCN QA76.9.D43 I58 2005.

Hudson:2008:FU

[HH08] Andrew Hudson and Paul Hudson. Fedora 7 unleashed. SAMS Publishing, Indianapolis, IN, USA, 2008. ISBN 0-672-32942-5. xxix + 903 pp. LCCN QA76.76.O63 H794

2008. URL http://www.loc.gov/catdir/toc/ecip0723/
2007030725.html.

Huang:2013:VHS

[HH13] Chun-Hsian Huang and Pao-Ann Hsiung. Virtualizable hardware/software design infrastructure for dynamically partially reconfigurable systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(2):11:1–11:??, July 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Hejja:2018:OPA

[HH18] Khaled Hejja and Xavier Hesselbach. Online power aware coordinated virtual network embedding with 5G delay constraint. Journal of Network and Computer Applications, 124(??):121–136, December 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804518303151.

Hejja:2019:EIT

[HH19] Khaled Hejja and Xavier Hesselbach. Evaluating impacts of traffic migration and virtual network functions consolidation on power aware resource allocation algorithms. Future Generation Computer Systems, 101(??):83–98, December 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X19304108.

Hong:2016:OCT

[HHC^{+}16] Ding-Yong Hong, Chun-Chen Hsu, Cheng-Yi Chou, Wei-Chung Hsu, Pangfeng Liu, and Jan-Jan Wu. Optimizing control transfer and memory virtualization in full system emulators. ACM Transactions on Architecture and Code Optimization, 12(4):47:1–47:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ho:2004:PPD

[HHH04] Alex Ho, Steven Hand, and Tim Harris. PDB: Pervasive debugging with Xen. In IEEE [IEE04], pages 260– 265. ISBN 0-7695-2256-4. LCCN QA76.9.C58; QA76.9.C58

I585 2004eb; Internet. URL http://ieeexplore.ieee.org/
servlet/opac?punumber=9495. IEEE Computer Society Or-

der Number P2256.

Horiguchi:1994:ISP

[HHK94] S. Horiguchi, D. F. Hsu, and M. Kimura, editors. International Symposium on Parallel Architectures, Algorithms, and Networks (ISPAN): proceedings of the 1994, December 14– 16, 1994, Kanazawa, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6507-6. LCCN QA76.58 .I5673 1994. IEEE catalog no. 94TH0697-3.

Hussein:2015:DRM

[HHPV15] Ahmed Hussein, Antony L. Hosking, Mathias Payer, and Christopher A. Vick. Don’t race the memory bus: taming the GC leadfoot. ACM SIGPLAN Notices, 50(11):15–27, November 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haria:2018:DMH

[HHS18] Swapnil Haria, Mark D. Hill, and Michael M. Swift. Devirtualizing memory in heterogeneous systems. ACM SIGPLAN Notices, 53(2):637–650, February 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hausheer:2018:SPS

[HHSG18] David Hausheer, Oliver Hohlfeld, Stefan Schmid, and Guofei Gu. Security and performance of software-defined networks and functions virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 138(??):15–17, June 19, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128618301452.

Ha:2002:AHS

[HHV^{+}02] Yajun Ha, Radovan Hipik, Serge Vernalde, Diederik Verkest, Marc Engels, Rudy Lauwereins, and Hugo De Man. Adding hardware support to the HotSpot virtual machine for domain specific applications. Lecture Notes in Computer Science, 2438:1135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2438/24381135.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2438/24381135.pdf.

Haase:2010:SDV

[HHW10] Jan Haase, Andreas Hofmann, and Klaus Waldschmidt. A self distributing virtual machine for adaptive multicore environments. International Journal of Parallel Programming, 38(1):19–37, February 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&
issn=0885-7458&volume=38&issue=1&spage=19.

Haque:2016:ACV

[HIIG16] Syed Arefinul Haque, Salekul Islam, Md. Jahidul Islam, and Jean-Charles Grégoire. An architecture for client virtualization: a case study. Computer Networks (Amsterdam, Netherlands: 1999), 100(??):75–89, May 8, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128616300421.

Hines:1997:VMJ

[Hin97] J. R. Hines. Virtual machines jockey for position. IEEE Spectrum, 34(7):16, July 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hinkelmann:2008:EKM

[Hin08] Matthias Hinkelmann. Evaluation von KVM als möglichen Ersatz für XEN (MI08). Bachelor thesis, Hochschule, Stuttgart, Germany, 2008. 85 pp.

Hirschsohn:1992:PSS

[Hir92] Ian Hirschsohn. Personal supercomputing: Seamless portability. Dr. Dobb’s Journal of Software Tools, 17(7):40, 42–44, 46, 48, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Hirai:2017:DEV

[Hir17] Yoichi Hirai. Defining the Ethereum virtual machine for interactive theorem provers. In Financial Cryptography and Data Security, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017.

URL http://link.springer.com/chapter/10.1007/978-
3-319-70278-0_33.

Hansen:2010:SVM

[HJ10] Jacob Gorm Hansen and Eric Jul. Scalable virtual machine storage using local disks. Operating Systems Review, 44(4): 71–79, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Huin:2018:ONS

[HJG18] Nicolas Huin, Brigitte Jaumard, and Frederic Giroire. Optimal network service chain provisioning. IEEE/ACM Transactions on Networking, 26(3):1320–1333, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Henzinger:2007:EMP

[HK07] Thomas A. Henzinger and Christoph M. Kirsch. The embedded machine: Predictable, portable real-time code. ACM Transactions on Programming Languages and Systems, 29(6): 33:1–33:29, October 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hbaieb:2019:STV

[HKB19] Ameni Hbaieb, Mahdi Khemakhem, and Maher Ben Jemaa. A survey and taxonomy on virtual data center embedding. The Journal of Supercomputing, 75(10):6324–6360, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hofmann:2013:ISA

[HKD^{+}13] Owen S. Hofmann, Sangman Kim, Alan M. Dunn, Michael Z. Lee, and Emmett Witchel. InkTag: secure applications on an untrusted operating system. ACM SIGARCH Computer Architecture News, 41(1):265–278, March 2013. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Han:2019:EJM

[HKJ19] Seok-Hyeon Han, Hyun-Woo Kim, and Young-Sik Jeong. An efficient job management of computing service using integrated idle VM resources for high-performance computing based on OpenStack. The Journal of Supercomputing, 75(8): 4388–4407, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hovestadt:2013:AOC

[HKKW13] Matthias Hovestadt, Odej Kao, Andreas Kliem, and Daniel Warneke. Adaptive online compression in clouds — making informed decisions in virtual machine environments. Journal of Grid Computing, 11(2):167–186, June 2013. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10723-013-9249-4.

Hao:2017:OAV

[HKLM17] Fang Hao, Murali Kodialam, T. V. Lakshman, and Sarit Mukherjee. Online allocation of virtual machines in a distributed cloud. IEEE/ACM Transactions on Networking, 25 (1):238–249, February 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hajnal:2018:EVI

[HKM^{+}18a] Akos Hajnal, Gabor Kecskemeti, Attila Csaba Marosi, Jozsef Kovacs, Peter Kacsuk, and Robert Lovas. ENTICE VM image analysis and optimised fragmentation. Journal of Grid Computing, 16(2):247–263, June 2018. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10723-018-9430-x.

Hinz:2018:CMI

[HKM^{+}18b] Mauro Hinz, Guilherme Piegas Koslovski, Charles C. Miers, Laércio L. Pilla, and Maurício A. Pillon. A cost model for IaaS clouds based on virtual machine energy consumption. Journal of Grid Computing, 16(3):493–512, September 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL https://link.springer.com/article/
10.1007/s10723-018-9440-8.

Haris:2022:LMV

[HKN22] Raseena M. Haris, Khaled M. Khan, and Armstrong Nhlabatsi. Live migration of virtual machine memory content in networked systems. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128622000962.

Haidri:2019:CED

[HKS19] R. A. Haidri, C. P. Katti, and P. C. Saxena. Cost-effective deadline-aware stochastic scheduling strategy for workflow applications on virtual machines in cloud computing. Concurrency and Computation: Practice and Experience, 31(7): e5006:1–e5006:??, April 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2013:ESC

[HL13] Qun Huang and Patrick P. C. Lee. An experimental study of cascading performance interference in a virtualized environment. ACM SIGMETRICS Performance Evaluation Review, 40(4):43–52, March 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

He:2024:MDC

[HL24] Xinfeng He and Riyang Li. Malware detection for container runtime based on virtual machine introspection. The Journal of Supercomputing, 80(6):7245–7268, April 2024. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-023-05727-w.

Hsieh:2020:UPA

[HLBZ20] Sun-Yuan Hsieh, Cheng-Sheng Liu, Rajkumar Buyya, and Albert Y. Zomaya. Utilization-prediction-aware virtual machine consolidation approach for energy-efficient cloud data centers. Journal of Parallel and Distributed Computing, 139(??):99–109, May 2020. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S074373151930190X.

Hoque:2016:AAT

[HLP^{+}16] Endadul Hoque, Hyojeong Lee, Rahul Potharaju, Charles Killian, and Cristina Nita-Rotaru. Automated adversarial testing of unmodified wireless routing implementations. IEEE/ACM Transactions on Networking, 24(6):3369–3382, December 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hong:2016:AAQ

[HLPY16] Cheol-Ho Hong, Kyungwoon Lee, Hyunchan Park, and Chuck Yoo. ANCS: Achieving QoS through dynamic allocation of network resources in virtualized clouds. Scientific Programming, 2016(??):4708195:1–4708195:10, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.hindawi.com/journals/sp/
2016/4708195/.

Hahn:2010:UVL

[HLW^{+}10] Dongwoon Hahn, Ginnah Lee, Brenton Walker, Matt Beecher, and Padma Mundur. Using virtualization and live migration in a scalable mobile wireless testbed. ACM SIG-METRICS Performance Evaluation Review, 38(3):21–25, December 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hsu:2013:IDB

[HLW^{+}13] Chun-Chen Hsu, Pangfeng Liu, Jan-Jan Wu, Pen-Chung Yew, Ding-Yong Hong, Wei-Chung Hsu, and Chien-Min Wang. Improving dynamic binary optimization through early-exit guided code region formation. ACM SIGPLAN Notices, 48(7):23–32, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

He:2023:NFA

[HLW^{+}23] Jiahao He, Shuangyin Li, Xinming Wang, Shing-Chi Cheung, Gansen Zhao, and Jinji Yang. Neural-FEBI: Accurate function identification in Ethereum Virtual Machine bytecode. The Journal of Systems and Software, 199(??):??, May 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-

1228 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0164121223000225.

Hartel:2001:FSJ

[HM01] Pieter H. Hartel and Luc Moreau. Formalizing the safety of Java, the Java Virtual Machine, and Java card. ACM Computing Surveys, 33(4):517–558, December 2001. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Halacsy:2018:OEE

[HM18] Gergely Halácsy and Zoltán Ádám Mann. Optimal energyefficient placement of virtual machines with divisible sizes. Information Processing Letters, 138(??):44–50, October 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019018301285.

Haghshenas:2020:PBU

[HM20] Kawsar Haghshenas and Siamak Mohammadi. Predictionbased underutilized and destination host selection approaches for energy-efficient dynamic VM consolidation in data centers. The Journal of Supercomputing, 76(12):10240–10257, December 2020. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL https://link.springer.com/
article/10.1007/s11227-020-03248-4.

Hallawi:2017:MCC

[HMH17] Huda Hallawi, Jörn Mehnen, and Hongmei He. Multicapacity combinatorial ordering GA in application to cloud resources allocation and efficient virtual machines consolidation. Future Generation Computer Systems, 69(??):1–10, April 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X16304630.

Hu:2004:TLI

[HMS04] Ji Hu, Christoph Meinel, and Michael Schmitt. Tele-lab IT security: an architecture for interactive lessons for security education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):412–416, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Howard:2017:RPF

[HMS17] Heidi Howard, Dahlia Malkhi, and Sasha Spiegelman. Revisiting the Paxos Foundations: a look at summer internship work at VMware Research. Operating Systems Review, 51 (1):67–71, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hay:2008:FEV

[HN08] Brian Hay and Kara Nance. Forensics examination of volatile system data using virtual introspection. Operating Systems Review, 42(3):74–82, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hess:2010:PVS

[HN10] Kenneth Hess and Amy Newman. Practical Virtualization Solutions: Virtualization from the Trenches. Negus software solutions series. Prentice Hall/Pearson Education, Upper Saddle River, NJ, USA, 2010. ISBN 0-13-714297-8 (paperback). xxiii + 304 pp. LCCN QA76.9.V5 H47 2010.

Hamilton:1992:SHU

[HO92] Jeffrey W. Hamilton and Eileen M. Ornsby. Simulating hypercubes in Unix. part I. Dr. Dobb’s Journal of Software Tools, 17(12):72, 74, 76, 108–110, 112–113, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Helali:2022:SLC

[HO22] Leila Helali and Mohamed Nazih Omri. Software license consolidation and resource optimization in container-based virtualized data centers. Journal of Grid Computing, 20(2):??, June 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL https://link.springer.com/article/
10.1007/s10723-022-09602-5.

Hoffmann:2020:RVM

[Hof20] Leah Hoffmann. Reinventing virtual machines. Communications of the ACM, 63(4):128ff, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoganson:2002:HPC

[Hog02] Kenneth E. Hoganson. High-performance computer architecture and algorithm simulator. ACM Journal on Educational Resources in Computing (JERIC), 2(1):131–148, March 2002. CODEN ???? ISSN 1531-4278.

Hogenson:2006:CCV

[Hog06] Gordon Hogenson. C++/CLI: the Visual C++ language for .NET. The Expert’s voice in .NET. Apress, Berkeley, CA, USA, 2006. ISBN 1-4302-0282-3, 1-59059-705-2. xxxiv + 412 pp. LCCN QA76.73.C153 H59 2006eb.

Hogenson:2008:FCC

[Hog08] Gordon Hogenson. Foundations of C++/CLI: the Visual C++ Language for .NET 3.5. Apress, Berkeley, CA, USA, 2008. ISBN 1-4302-1024-9, 1-4302-1023-0. LCCN QA76.73.C153 H54 2008eb.

Horie:2014:SDJ

[HOKO14] Michihiro Horie, Kazunori Ogata, Kiyokuni Kawachiya, and Tamiya Onodera. String deduplication for Java-based middleware in virtualized environments. ACM SIGPLAN Notices, 49(7):177–188, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hollerbach:1995:FDA

[Hol95] Rainer Hollerbach. Fast dynamo action in spherical geometry: Numerical calculations using parallel virtual machines. Computers in Physics, 9(4):460–??, July 1995. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168547.

Horton:1973:VMA

[Hor73] F. R. Horton. Virtual machine assist: Performance. Guide 37, Boston, MA, 1973.

Huxtable:1977:HSI

[HP77] D. H. R. Huxtable and J. M. M. Pinkerton. The hardware/software interface of the ICL 2900 range of computers. The Computer Journal, 20(4):290–295, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_20/Issue_04/tiff/290.
tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/291.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_20/Issue_04/tiff/
292.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/293.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/294.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/295.tif.

Harutyunyan:2024:DCA

[HPB^{+}24] Ashot Harutyunyan, Arnak Poghosyan, Tigran Bunarjyan, Andranik Haroyan, Marine Harutyunyan, Lilit Harutyunyan, and Nelson Baloian. Discovery of cloud applications from logs. Future Internet, 16(6):216, June 18, 2024. CODEN

???? ISSN 1999-5903. URL https://www.mdpi.com/1999-
5903/16/6/216.

Huang:2004:MDS

[HPcC04] Lan Huang, Gang Peng, and Tzi cker Chiueh. Multidimensional storage virtualization. ACM SIGMETRICS Performance Evaluation Review, 32(1):14–24, June 2004. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hohmuth:2004:RTS

[HPHS04] Michael Hohmuth, Michael Peter, Hermann Härtig, and Jonathan S. Shapiro. Reducing TCB size by using untrusted components: small kernels versus virtual-machine monitors. In ACM [ACM04a], page ?? ISBN ???? LCCN ????

Hussein:2017:OPR

[HPHV17] Ahmed Hussein, Mathias Payer, Antony L. Hosking, and Chris Vick. One process to reap them all: Garbage collection as-a-service. ACM SIGPLAN Notices, 52(7):171–186, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hwang:2015:RPA

[HPP15] Woomin Hwang, Ki-Woong Park, and Kyu Ho Park. Reference pattern-aware instant memory balancing for consolidated virtual machines on manycores. IEEE Transactions on Parallel and Distributed Systems, 26(7):2036–2050, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183

(electronic). URL http://www.computer.org/csdl/trans/
td/2015/07/06860294-abs.html.

Hummaida:2022:SVM

[HPS22] Abdul Rahman Hummaida, Norman W. Paton, and Rizos Sakellariou. Scalable virtual machine migration using reinforcement learning. Journal of Grid Computing, 20(2):??, June 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL https://link.springer.com/article/
10.1007/s10723-022-09603-4.

Hummaida:2023:HDA

[HPS23] Abdul R. Hummaida, Norman W. Paton, and Rizos Sakellariou. A hierarchical decentralized architecture to enable adaptive scalable virtual machine migration. Concurrency and Computation: Practice and Experience, 35(2):e7487:1– e7487:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2006:RST

[HS06] Shiliang Hu and James E. Smith. Reducing startup time in co-designed virtual machines. ACM SIGARCH Computer Architecture News, 34(2):277–288, 2006. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Hsu:2013:VNM

[HS13] Wu-Hsiao Hsu and Yuh-Pyng Shieh. Virtual network mapping algorithm in the cloud infrastructure. Journal of Network and Computer Applications, 36(6):1724–1734, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1084804513000660.

Honda:2019:NWD

[HS19] Hirotada Honda and Hiroshi Saito. Nation-wide disaster avoidance control against heavy rain. IEEE/ACM Transactions on Networking, 27(3):1084–1097, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2021:ESC

[HS21] Chih-Kai Huang and Shan-Hsiang Shen. Enabling service cache in edge clouds. ACM Transactions on Internet of Things (TIOT), 2(3):18:1–18:24, July 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL https:

//dl.acm.org/doi/10.1145/3456564.

Hsu:2015:LLA

[HSC15] Ching-Hsien Hsu, Kenn D. Slagter, and Yeh-Ching Chung. Locality and loading aware virtual machine mapping techniques for optimizing communications in MapReduce applications. Future Generation Computer Systems, 53(??):43–54, December 2015. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167739X15000989.

Ha:2017:PPE

[HSK17] Tuan Minh Ha, Masaki Samejima, and Norihisa Komoda. Power and performance estimation for fine-grained server power capping via controlling heterogeneous applications. ACM Transactions on Management Information Systems (TMIS), 8(4):11:1–11:??, September 2017. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Hu:2017:TFC

[HSL17] Yang Hu, Mingcong Song, and Tao Li. Towards “Full Containerization” in containerized network function virtualization. ACM SIGARCH Computer Architecture News, 45(1): 467–481, March 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Hong:2017:FFF

[HSN17a] Cheol-Ho Hong, Ivor Spence, and Dimitrios S. Nikolopoulos. FairGV: Fair and fast GPU virtualization. IEEE Transactions on Parallel and Distributed Systems, 28(12):3472–3485, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print),

1558-2183 (electronic). URL https://www.computer.org/
csdl/trans/td/2017/12/07954729-abs.html.

Hong:2017:GVS

[HSN17b] Cheol-Ho Hong, Ivor Spence, and Dimitrios S. Nikolopoulos. GPU virtualization and scheduling methods: a comprehensive survey. ACM Computing Surveys, 50(3):35:1–35:??, October 2017. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Hsu:2001:CAS

[Hsu01] John Y. Hsu. Computer Architecture: Software Aspects, Coding, Hardware. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. ISBN 0-8493-1026-1, 1-351-83604-8, 1-4200-4110-X (e-book). 427 pp. LCCN A76.9.A73 H758 2001. US$89.95, UK£59.99.

Hagiya:1998:NMD

[HT98] M. Hagiya and A. Tozawa. On a new method for dataflow analysis of Java Virtual Machine subroutines. Lecture Notes in Computer Science, 1503:17–32, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haugerud:2021:DSP

[HTAY21] Hårek Haugerud, Huy Nhut Tran, Nadjib Aitsaadi, and Anis Yazidi. A dynamic and scalable parallel Network Intrusion Detection System using intelligent rule ordering and network function virtualization. Future Generation Computer Systems, 124(??):254–267, November 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X21001874.

He:2019:PEL

[HTB19] TianZhang He, Adel N. Toosi, and Rajkumar Buyya. Performance evaluation of live virtual machine migration in SDNenabled cloud data centers. Journal of Parallel and Distributed Computing, 131(??):55–68, September 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151830474X.

He:2022:CCA

[HTB22] TianZhang He, Adel N. Toosi, and Rajkumar Buyya. CAMIG: Concurrency-aware live migration management of multiple virtual machines in SDN-enabled clouds. IEEE Transactions on Parallel and Distributed Systems, 33(10):2318–2331, October 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2024:PPV

[HTM^{+}24] Haojun Huang, Jialin Tian, Geyong Min, Hao Yin, Cheng Zeng, Yangming Zhao, and Dapeng Oliver Wu. Parallel placement of virtualized network functions via federated deep reinforcement learning. IEEE/ACM Transactions on Networking, 32(4):2936–2949, June 2024. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://dl.

acm.org/doi/10.1109/TNET.2024.3366950.

Meyer:2008:PVD

[hTMAC^{+}08] Dut h T. Meyer, Gitika Aggarwal, Brendan Cully, Geoffrey Lefebvre, Michael J. Feeley, Norman C. Hutchinson, and Andrew Warfield. Parallax: virtual disks for virtual machines. Operating Systems Review, 42(4):41–54, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Han:2019:EED

[HTW^{+}19] Zhenhua Han, Haisheng Tan, Rui Wang, Guihai Chen, Yupeng Li, and Francis Chi Moon Lau. Energy-efficient dynamic virtual machine management in data centers. IEEE/ ACM Transactions on Networking, 27(1):344–360, February 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Hu:1990:RTC

[Hu90] W. Hu. Reducing timing channels with fuzzy time. In IEEE [IEE90a], pages 8–20. ISBN 0-8186-2060-9, 0-8186-6060-0, 0-8186-9060-7. LCCN QA76.9.A25; QA76.9.A25 S95 1990eb; QA76.9.A25 I34 1990. IEEE Computer Society Order Number 2060. IEEE Catalog Number 90CH2884-5.

Hui:2018:VMA

[Hui18] Yang Hui. A virtual machine anomaly detection system for cloud computing infrastructure. The Journal of Supercomputing, 74(11):6126–6134, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Heiser:2006:VMM

[HUL06] Gernot Heiser, Volkmar Uhlig, and Joshua LeVasseur. Are virtual-machine monitors microkernels done right? Operating Systems Review, 40(1):95–99, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hwang:2014:MFG

[HUWH14] Jinho Hwang, Ahsen Uppal, Timothy Wood, and Howie Huang. Mortar: filling the gaps in data center memory. ACM SIGPLAN Notices, 49(7):53–64, July 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herbordt:1993:EEA

[HW93] Martin C. Herbordt and Charles C. Weems. An environment for evaluating architectures for spatially mapped computatio: System architecture and preliminary results. Technical Report UM-CS-1993-070, University of Massachusetts, Amherst, Computer Science, September 1993. URL ftp://

ftp.cs.umass.edu/pub/techrept/1993/UM-CS-1993-070.
ps.

Hudic:2012:PCC

[HW12] Aleksandar Hudic and Edgar Weippl. Private cloud computing: Consolidation, virtualization, and service-oriented infrastructure. Computers & Security, 31(4):629, June 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404812000612.

Hume:2015:SCS

[HW15] Tom Hume and Des Watson. Short communication: Is superoptimization viable for VM instruction sets? Software— Practice and Experience, 45(4):571–579, April 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hu:2003:DJV

[HWB03] E. Y. S. Hu, A. J. Wellings, and G. Bernat. Deriving Java Virtual Machine timing models for portable worst-case execution time analysis. Lecture Notes in Computer Science, 2889: 411–424, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:2016:BKB

[HWCH16] Yu-Ju Huang, Hsuan-Heng Wu, Yeh-Ching Chung, and Wei-Chung Hsu. Building a KVM-based hypervisor for a heterogeneous system architecture compliant system. ACM SIG-PLAN Notices, 51(7):3–15, July 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hand:2007:HVX

[HWF07] Steven Hand, Andrew Warfield, and Keir Fraser. Hardware virtualization with Xen. ;login: the USENIX Association newsletter, 32(1):??, February 2007. CODEN LOGNEM. ISSN 1044-6397. URL https://www.usenix.

org/publications/login/february-2007-volume-32-number-
1/hardware-virtualization-xen.

Huang:2018:TBI

[HWHW18] Jhih-Yu Huang, Pi-Chung Wang, Jhih-Yu Huang, and Pi-Chung Wang. TCAM-based IP address lookup using longest suffix split. IEEE/ACM Transactions on Networking, 26(2): 976–989, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Huang:2024:VEC

[HWR^{+}24] Hang Huang, Honglei Wang, Jia Rao, Song Wu, Hao Fan, Chen Yu, Hai Jin, Kun Suo, and Lisong Pan. vKernel: Enhancing container isolation via private code and data. IEEE Transactions on Computers, 73(7):1711–1723, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hao:2016:IRO

[HXZ^{+}16] Zheng Hao, Dong Xiaoshe, Zhu Zhengdong, Chen Baoke, Bai Xiuxiu, Zhang Xingjun, and Wang Endong. Improving the reliability of the operating system inside a VM. The Computer Journal, 59(5):715–740, May 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/59/5/715.

Han:2023:SSV

[HYK^{+}23] Juhyeng Han, Insu Yun, Seongmin Kim, Taesoo Kim, Sooel Son, and Dongsu Han. Scalable and secure virtualization of HSM with ScaleTrust. IEEE/ACM Transactions on Networking, 31(4):1595–1610, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://dl.

acm.org/doi/10.1109/TNET.2022.3220427.

Huang:2025:ESF

[HYL^{+}25] Yaodong Huang, Tingting Yao, Zelin Lin, Xiaojun Shang, Yukun Yuan, Laizhong Cui, and Yuanyuan Yang. Efficient service function chain placement over heterogeneous devices in deviceless edge computing environments. IEEE Transactions on Computers, 74(1):222–236, January 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2025:SDF

[HZK^{+}25] Zhaoyang Huang, Yifu Zhu, Xin Kuang, Yanjie Tan, Huailiang Tan, and Keqin Li. State-driven fairness control for efficient I/O queue scheduling in NVMe virtualization. Future Generation Computer Systems, 173(??):??, December 2025. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X25001979.

Han:2018:RAM

[HZL^{+}18] Jin Han, Wangyu Zang, Li Liu, Songqing Chen, and Meng Yu. Risk-aware multi-objective optimized virtual machine placement in the cloud. Journal of Computer Security, 26 (5):707–730, ???? 2018. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

He:2014:DRC

[HZZ^{+}14] Ligang He, Deqing Zou, Zhang Zhang, Chao Chen, Hai Jin, and Stephen A. Jarvis. Developing resource consolidation frameworks for moldable virtual machines in clouds. Future Generation Computer Systems, 32(??):69–81, March 2014.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X12001112.

Iancu:2014:CPV

[Ian14] Costin Iancu. The case for partitioning virtual machines on multicore architectures. IEEE Transactions on Parallel and Distributed Systems, 25(10):2683–2696, October 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://www.computer.org/csdl/trans/td/
2014/10/06613989-abs.html.

Ijaz:2020:RHP

[IBBA20] Qaiser Ijaz, El-Bay Bourennane, Ali Kashif Bashir, and Hira Asghar. Revisiting the high-performance reconfigurable computing for future datacenters. Future Internet, 12(4):64, April 06, 2020. CODEN ???? ISSN 1999-5903. URL https://www.

mdpi.com/1999-5903/12/4/64.

IBM:1972:IVM

[IBM72] IBM. IBM Virtual Machine Facility/370 Planning Guide. IBM Corporation, San Jose, CA, USA, 1972. ???? pp. Publication No. GC20-1801-0.

IBM:1973:IVM

[IBM73] IBM. IBM Virtual Machine Facility/370: Release 2 Planning Guide. IBM Corporation, San Jose, CA, USA, 1973. ???? pp. Publication No. GC20-1814-0.

IBM:1976:GIS

[IBM76a] IBM. A Guide to the IBM System/370 Model 168 for System/370 Model 165 Users. IBM Corporation, San Jose, CA, USA, fourth edition, September 1976. URL http://www.

bitsavers.org/pdf/ibm/370/model168/GC20-1755-3_A_
Guide_to_the_IBM_System_370_Model_168_for_Model_165_

Users_Sep76.pdf. Publication GC20-1755-3. See Section 40, Virtual Machines, pages 86–95.

IBM:1976:IVM

[IBM76b] IBM. IBM Virtual Machine Facility/370: Introduction. IBM Corporation, Poughkeepsie, NY, USA, seventh edition, Oc-

tober 1976. 54 pp. URL http://bitsavers.trailing-
edge.com/pdf/ibm/370/VM_370/Release_3/GC20-1800-6_

VM370_Introduction_Rel_3_Oct76.pdf. Order Number

GC20-1800-6.

IBM:1985:VM

[IBM85] IBM. Virtual machine. IBM Corporation, Endicott, NY, USA, fourth edition, 1985. 282 pp.

IBM:1988:VMSa

[IBM88] IBM Corporation. Virtual machine / extended architecture system product: application development guide for FOR-TRAN and COBOL. IBM Corporation, San Jose, CA, USA, 1988. iii–x + 263 pp.

IBM:1994:CGN

[IBM94] IBM. Conversion guide and notebook for VM/XA SP and VM/ESA, release 2.2: virtual machine. International Business Machines Corporation, 1701 North St., Endicott 13760-5553 NY, USA, fifth edition, 1994. 1061 pp. Publication number SC24-5525-04.

IBM:1996:CAM

[IBM96] IBM. CMS application multitasking, version 2, release 2.0: virtual machine. International Business Machines Corporation, 1701 North St., Enicott 13760-5553, NY, USA, second edition, 1996. xiv + 380 pp. Publication number SC24-5766-01.

Ibsen:1984:PVM

[Ibs84a] Leif Ibsen. A portable virtual machine for Ada. Software— Practice and Experience, 14(1):17–29, January 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

SPE::Ibsen1984

[Ibs84b] Leif Ibsen. A portable virtual machine for Ada. Software— Practice and Experience, 14(1):17–29, January 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

IEEE:1984:PSS

[IEE84a] IEEE, editor. Proceedings of the 1984 Symposium on Security and Privacy, April 29–May 2, 1984, Oakland, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. ISBN 0-8186-0532-4 (paperback), 0-8186-8532-8 (hard), 0-8186-4532-6 (microfiche). LCCN QA76.9.A25 S95 1984.

IEEE:1984:DE

[IEE84b] IEEE, editor. Proceedings of the First International Conference on Data Engineering, Los Angeles, CA, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. ISBN 0-8186-0533-2 (paperback), 0-8186-8533-6 (hard), 0-8186-4533-4 (microfiche). LCCN QA76.9.D3 I5582 1984.

IEEE:1985:CPA

[IEE85] IEEE, editor. Conference proceedings: the 12th Annual International Symposium on Computer Architecture, June 17– 19, 1985, Boston, Massachusetts, volume 13(3) of SIGARCH newsletter. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 1985. ISBN 0-8186-0634-7, 0-8186-4634-9, 0-8186-8634-0. LCCN QA76.9.A73; QA76.9.A73 C65; QA76.9.A73 S97 1985; QA76.9.A73 I56 1985.

IEEE:1990:PIC

[IEE90a] IEEE, editor. Proceedings, 1990 IEEE Computer Society Symposium on Research in Security and Privacy: May 7–9, 1990, Oakland, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2060-9, 0-8186-6060-0, 0-8186-9060-7. LCCN QA76.9.A25; QA76.9.A25 S95 1990eb; QA76.9.A25 I34 1990. IEEE Computer Society Order Number 2060. IEEE Catalog Number 90CH2884-5.

IEEE:1990:PSN

[IEE90b] IEEE, editor. Proceedings, Supercomputing ’90: November 12–16, 1990, New York Hilton at Rockefeller Center, New York, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

IEEE:1991:PIC

[IEE91] IEEE, editor. Proceedings, 1991 IEEE Computer Society Symposium on Research in Security and Privacy: May 20–22, 1991, Oakland, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-9168-9, 0-8186-2168-0, 0-8186-6168-2. LCCN QA76.9.A25; QA76.9.A25 I34 1991; QA76.9.A25 S95 1991eb; QA76.9.A25 S95 1991. IEEE Catalog Number 91CH2986-8. IEEE Computer Society Order Number 2168.

IEEE:1992:PSM

[IEE92] IEEE Computer Society. Technical Committee on Computer Architecture, editor. Proceedings, Supercomputing ’92: Minneapolis, Minnesota, November 16–20, 1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

IEEE:1993:PSI

[IEE93a] IEEE, editor. Proceedings of the Second International Conference on Parallel and Distributed Information Systems, San Diego, CA, USA, January 20–22, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3330-1. LCCN QA76.58 .I54 1993. IEEE catalog number 93TH0493-7.

IEEE:1993:PSP

[IEE93b] IEEE, editor. Proceedings, Supercomputing ’93: Portland, Oregon, November 15–19, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

IEEE:1996:HCV

[IEE96a] IEEE, editor. Hot chips VIII: symposium record: Stanford University, Stanford, California, August 18–20, 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN ???? LCCN ????

IEEE:1996:PFIa

[IEE96b] IEEE, editor. Proceedings of the Fifth IEEE International Symposium on High Performance Distributed Computing, August 6–9, 1996, Syracuse, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7582-9. LCCN QA76.88.I52 1996. IEEE order plan catalog number 96TB100069. EEE Computer Society Press order number PR07582.

IEEE:1997:HCI

[IEE97] IEEE, editor. Hot Chips IX: Stanford University, Stanford, California, August 24–26, 1997. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN ???? LCCN ????

IEEE:1999:HCS

[IEE99] IEEE, editor. Hot Chips 11: Stanford University, Stanford, California, August 15–17, 1999. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN ???? LCCN ???? URL http://

www.hotchips.org/hotc11_index.html.

IEEE:2001:EIW

[IEE01] IEEE, editor. Eighth IEEE Workshop on Hot Topics in Operating Systems (HotOS-VIII). May 20–23, 2001, Schloss Elmau, Germany. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7695-1040-X. US$135.00. URL http://computer.

org/CSPRESS/CATALOG/pr01040.htm. IEEE catalog number

PR01040.

IEEE:2002:WII

[IEE02] IEEE, editor. WWC-5: 2002 IEEE International Workshop on Workload Characterization: November 25, 2002, Austin, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7803-7681-1. LCCN QA76.9.S88; QA76.9.S88 W67 2002eb; Inter-

net. URL http://ieeexplore.ieee.org/servlet/opac?

punumber=8689. IEEE Catalog Number 02EX633.

IEEE:2003:IIW

[IEE03] IEEE, editor. 2003 IEEE International Workshop on Workload Characterization: WWC-6: October 27, 2003, Austin, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7803-8229-3. LCCN QA76.9.S88 W67 2003.

IEEE:2004:FIA

[IEE04] IEEE, editor. Fifth IEEE/ACM International Workshop on Grid Computing: proceedings; 8 November 2004, Pittsburgh, PA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2256-4. LCCN QA76.9.C58; QA76.9.C58 I585 2004eb; Inter-

net. URL http://ieeexplore.ieee.org/servlet/opac?

punumber=9495. IEEE Computer Society Order Number P2256.

IEEE:2005:PAC

[IEE05] IEEE, editor. Proceedings of the 21st Annual Computer Security Applications Conference 2005, December 05–09, 2005, Tucson, Arizona. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2461-3. ISSN 1063-9527. LCCN L787.5. IEEE Computer Society Order Number P2461.

IEEE:2006:PIS

[IEE06a] IEEE, editor. Proceedings of the 20th International Symposium on High-Performance Computing in an Advanced Collaborative Environment (HPCS’06), St. John’s, Newfoundland, May 14–17, 2006. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 0-7695-2582-2. ISSN 1550-5243. LCCN QA76.88. ACM product number E2582.

IEEE:2006:PIC

[IEE06b] IEEE, editor. Proceedings of the International conference on Networking and Services 2006, July 16–18, 2006. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 0-7695-2622-5. LCCN ????

Izquierdo:2019:SDA

[IGBKR19] D. Izquierdo, J. M. Gonzalez-Barahona, L. Kurth, and G. Robles. Software development analytics for Xen: Why and how. IEEE Software, 36(3):28–32, May/June 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Moore:1979:IVM

[II79] J. S. Moore II. The InterLISP virtual machine specification. Report CSL 76-5, Xerox, Palo Alto Research Center (PARC), Palo Alto, CA, USA, 1979. Sep. 1976, Rvsd. March 1979.

Inoue:2006:VNP

[IIK^{+}06] Hiroaki Inoue, Akihisa Ikeno, Masaki Kondo, Junji Sakai, and Masato Edahiro. VIRTUS: a new processor virtualization architecture for security-oriented next-generation mobile terminals. In ACM [ACM06a], pages 484–489. ISBN 1-59593-381-6. LCCN ????

Ilgenfritz:2009:VCP

[IIPB09] Merle Ilgenfritz, John Ilgenfritz, John Powell, and Steven Baca. VMware certified professional test prep. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-6599-8. xxi + 867 pp. LCCN

QA76.3 .I56 2009eb. URL http://www.engnetbase.com/
ejournals/books/book\%5Fsummary/summary.asp?id=7409.

Ilkhechi:2015:NAV

[IKU15] Amir Rahimzadeh Ilkhechi, Ibrahim Korpeoglu, and Özgür Ulusoy. Network-aware virtual machine placement in cloud data centers with multiple traffic-intensive components. Computer Networks (Amsterdam, Netherlands: 1999), 91(??):508–527, November 14, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1389128615003023.

Infante:1975:PSP

[IM75] R. Infante and U. Montanari. Proving structured programs correct, level by level. ACM SIGPLAN Notices, 10(6):427– 436, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Inouchi:1993:PTI

[IM93] H. Inouchi and N. McLoughlin. Parallel techniques for image processing and artificial neural network simulation. Ai and Cognitive Science ’91, pages 177–89, 1993.

Ingalls:2020:TDL

[IMBB20] Daniel Ingalls, Eliot Miranda, Clément Béra, and Elisa Gonzalez Boix. Two decades of live coding and debugging of virtual machines through simulation. Software—Practice and Experience, 50(9):1629–1650, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Isci:2013:AEV

[IMK^{+}13] Canturk Isci, Suzanne McIntosh, Jeffrey Kephart, Rajarshi Das, James Hanson, Scott Piper, Robert Wolford, Thomas Brey, Robert Kantner, Allen Ng, James Norris, Abdoulaye Traore, and Michael Frissora. Agile, efficient virtualization power management with low-latency server power states. ACM SIGARCH Computer Architecture News, 41(3):96–107, June 2013. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ICSA ’13 conference proceedings.

Iacobovici:1987:VSP

[IN87] S. Iacobovici and C. C. Ng. VLSI and system performance modeling. IEEE Micro, 7(4):59–72, July/August 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

IBM:1988:VMSb

[Int88] International Business Machines Corporation. Virtual machine / system product: application development guide for FORTRAN and COBOL. IBM Corporation, San Jose, CA, USA, fourth edition, 1988. x + 280 pp.

ISO:2005:IIIa

[Int05a] International Organization for Standardization. ISO/IEC 23271:2003: Information technology — Common Language Infrastructure. International Organization for Standardization, Geneva, Switzerland, 2005. ISBN ???? xi + 99 (Part. I), ix + 164 (Part. II), vi + 125 (Part. III), iii + 16 (Part. IV), iv + 79 (Part. V) pp. LCCN ????

328 Sfr (print). URL http://standards.iso.org/ittf/
PubliclyAvailableStandards/c036769_ISO_IEC_23271_
2003(E).zip;   http://www.iso.org/iso/en/CatalogueDetailPage.
CatalogueDetail?CSNUMBER=36769.

ISO:2005:IIIb

[Int05b] International Organization for Standardization. ISO/IEC TR 23272: Information technology — Common Language Infrastructure — Profiles and Libraries. International Organization for Standardization, Geneva, Switzerland, 2005. ISBN ???? 6 pp. LCCN ???? 328 Sfr (print). URL http://

standards.iso.org/ittf/PubliclyAvailableStandards/
c036770_ISO_IEC_TR_23272_2003(E).zip;   http://www.
iso.org/iso/en/CatalogueDetailPage.CatalogueDetail?
CSNUMBER=36770.

ISO:2006:ITCb

[Int06a] International Organization for Standardization . ISO/IEC TR 25438:2006: Information technology: Common Language Infrastructure (CLI): technical report: common generics. Technical report. International Organization for Standardization, Geneva, Switzerland, 2006. LCCN ????

ISO:2006:III

[Int06b] International Organization for Standardization. ISO/IEC 23271:2006: Information technology: Common Language Infrastructure (CLI) Partitions I to VI. International standard. International Organization for Standardization, Geneva, Switzerland, second edition, 2006. LCCN ????

ISO:2006:ITCa

[Int06c] International Organization for Standardization. ISO/IEC TR 23272:2006: Information technology: Common Language Infrastructure (CLI): technical report on information derived from Partition IV XML file. Technical report. International Organization for Standardization, Geneva, Switzerland, second edition, 2006. ???? pp. LCCN ????

Iserte:2021:IME

[IPRS21] Sergio Iserte, Javier Prades, Carlos Reaño, and Federico Silla. Improving the management efficiency of GPU workloads in data centers through GPU virtualization. Concurrency and Computation: Practice and Experience, 33(2): e5275:1–e5275:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ilager:2019:EET

[IRB19] Shashikant Ilager, Kotagiri Ramamohanarao, and Rajkumar Buyya. ETAS: Energy and thermal-aware dynamic virtual machine consolidation in cloud data center with proactive hotspot mitigation. Concurrency and Computation: Practice and Experience, 31(17):e5221:1–e5221:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Inoue:2008:PVS

[ISE08] Hiroaki Inoue, Junji Sakai, and Masato Edahiro. Processor virtualization for secure mobile terminals. ACM Transactions on Design Automation of Electronic Systems., 13(3): 48:1–48:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

IT:2019:PDI

[Isl19] Island of TEX. Providing Docker images for TEX Live and ConTEXt. TUGboat, 40(3):231, 2019. CODEN ????

ISSN 0896-3207. URL https://tug.org/TUGboat/tb40-3/
tb126island-docker.pdf.

Ishikawa:1986:COO

[IT86] Y. Ishikawa and M. Tokoro. A concurrent object-oriented knowledge representation language Orient 84/K: its features and implementation. ACM SIGPLAN Notices, 21(11):232, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ive:2003:TER

[Ive03] Anders Ive. Towards an embedded real-time Java Virtual Machine. Thesis (fil.dr.), Lund University, Department of Computer Science, Lund, Sweden, 2003. viii + 139 pp.

Ishiguro:2023:RVA

[IYAK23] Kenta Ishiguro, Naoki Yasuno, Pierre-Louis Aublin, and Kenji Kono. Revisiting VM-agnostic KVM vCPU scheduler for mitigating excessive vCPU spinning. IEEE Transactions on Parallel and Distributed Systems, 34(10):2615–2628, October 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jacob:2005:DOE

[J^{+}05] Bart Jacob et al., editors. On demand Operating Environment: managing the infrastructure (virtualization engine update). IBM redbooks. IBM Corporation, San Jose, CA, USA, second edition, 2005. ISBN 0-7384-9124-1 (paperback). xxiv + 364 pp. LCCN QA76.754 .O5 2005.

Jain:2019:NNA

[JAC^{+}19] S. Jain, A. Ankit, I. Chakraborty, T. Gokmen, M. Rasch, W. Haensch, K. Roy, and A. Raghunathan. Neural network accelerator design with resistive crossbars: Opportunities and challenges. IBM Journal of Research and Development, 63(6): 10:1–10:13, November 2019. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Jiang:2019:BSR

[JAD19] Zhe Jiang, Neil Audsley, and Pan Dong. BlueIO: a scalable real-time hardware I/O virtualization system for many-core embedded systems. ACM Transactions on Embedded Computing Systems, 18(3):19:1–19:??, June 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:

//dl.acm.org/ft_gateway.cfm?id=3309765.

Jones:2006:ATP

[JADAD06a] Stephen T. Jones, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Antfarm: Tracking processes in a virtual machine environment. In USENIX [USE06], page ?? ISBN 1-931971-44-7. LCCN ???? URL http://www.usenix.org/

events/usenix06/tech/jones.html.

Jones:2006:GMB

[JADAD06b] Stephen T. Jones, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Geiger: monitoring the buffer cache in a virtual machine environment. Operating Systems Review, 40 (5):14–24, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jin:2015:HSH

[JAS^{+}15] Seongwook Jin, Jeongseob Ahn, Jinho Seol, Sanghoon Cha, Jaehyuk Huh, and Seungryoul Maeng. H-SVM: Hardwareassisted secure virtual machines under a vulnerable hypervisor. IEEE Transactions on Computers, 64(10):2833–2846, ???? 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Juiz:2024:SSC

[JB24] Carlos Juiz and Belen Bermejo. On the scalability of the speedup considering the overhead of consolidating virtual machines in servers for data centers. The Journal of Supercomputing, 80(9):12463–12511, June 2024. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-024-05943-y.

Jiang:2018:SAR

[JC18] Han-Peng Jiang and Wei-Mei Chen. Self-adaptive resource allocation for energy-aware virtual machine placement in dynamic computing cloud. Journal of Network and Computer Applications, 120(??):119–129, October 15, 2018. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302352.

Sun:1999:JCV

[JCV99] SUN Microsystems, Inc. Java Card 2.1 Virtual Machine Specification, March 3, 1999. Final Revision 1.0.

Jin:2013:CFG

[JCZZ13] Hai Jin, Ge Cheng, Deqing Zou, and Xinwen Zhang. Cherub: Fine-grained application protection with on-demand virtualization. Computers and Mathematics with Applications, 65(9):1326–1338, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0898122112001149.

Jordan:2006:SJT

[JDJ^{+}06] Mick Jordan, Laurent Daynès, Marcin Jarzab, Ciarán Bryce, and Grzegorz Czajkowski. Scaling J2EE ^{T M} application servers with the Multi-tasking Virtual Machine. Software— Practice and Experience, 36(6):557–580, May 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jin:2014:MLM

[JDW^{+}14] Hai Jin, Li Deng, Song Wu, Xuanhua Shi, Hanhua Chen, and Xiaodong Pan. MECOM: Live migration of virtual machines by adaptively compressing memory pages. Future Generation Computer Systems, 38(??):23–35, September 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X13002100.

Jarraya:2012:FVS

[JE12] Yosr Jarraya and Arash Eghtesadi. Formal verification of security preservation for migrating virtual machines in the cloud. Lecture Notes in Computer Science, 7596:111–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-33536-5_12/.

Jensen:1979:FAC

[Jen79] Richard M. Jensen. A formal approach for communication between logically isolated virtual machines. IBM Systems Journal, 18(1):71–92, 1979. CODEN IBMSA7. ISSN 0018-8670.

Jarraya:2015:VFR

[JES^{+}15] Yosr Jarraya, Arash Eghtesadi, Sahba Sadri, Mourad Debbabi, and Makan Pourzandi. Verification of firewall reconfiguration for virtual machines migrations in the cloud. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3)(??):480–491, December 24, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S138912861500376X.

Jiang:2016:FAF

[JFPL16] Jianhua Jiang, Yunzhao Feng, Milan Parmar, and Keqin Li. FP-ABC: Fast and parallel ABC based energy-efficiency live VM allocation policy in data centers. Scientific Programming, 2016(??):9524379:1–9524379:9, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https:

//www.hindawi.com/journals/sp/2016/9524379/.

Jiang:2017:DFA

[JFZL17] Jianhua Jiang, Yunzhao Feng, Jia Zhao, and Keqin Li. DataABC: a fast ABC based energy-efficient live VM consolidation policy with data-intensive energy evaluation model. Future Generation Computer Systems, 74(??):132–141, September 2017. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167739X16301236.

Janakiram:1988:RPB

[JGA^{+}88] V. K. Janakiram, E. F. Gehringer, D. P. Agrawal, Mehrotra, and R. A randomized parallel branch-and-bound algorithm. International Journal of Parallel Programming, 17 (3):277–301, June 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Jo:2013:ELM

[JGSE13] Changyeon Jo, Erik Gustafsson, Jeongseok Son, and Bernhard Egger. Efficient live migration of virtual machines using shared storage. ACM SIGPLAN Notices, 48(7):41–50, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Jin:2011:OLM

[JGW^{+}11] Hai Jin, Wei Gao, Song Wu, Xuanhua Shi, Xiaoxin Wu, and Fan Zhou. Optimizing the live migration of virtual machine by CPU scheduling. Journal of Network and Computer Applications, 34(4):1088–1096, July 2011. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001116.

Johnson:2014:CML

[JHE14] David Johnson, Mike Hibler, and Eric Eric. Composable multi-level debugging with Stackdb. ACM SIGPLAN Notices, 49(7):213–226, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jamthagen:2012:TRD

[JHS12] Christopher Jämthagen, Martin Hell, and Ben Smeets. A technique for remote detection of certain virtual machine monitors. Lecture Notes in Computer Science, 7222:129– 137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-32298-3_9/.

Jolitz:1991:PUS

[JJ91] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a stripped-down kernel. Dr. Dobb’s Journal of Software Tools, 16(7):32, 34–36, 38, 40, 84–88, 88, 90, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Jung:2002:DIS

[JJ02] Jun-Young Jung and Min-Soo Jung. Design and implementation of small-sized Java Virtual Machine on Java platform Jini. Lecture Notes in Computer Science, 2343:571–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.de/link/
service/series/0558/bibs/2343/23430571.htm; http:/
/link.springer.de/link/service/series/0558/papers/
2343/23430571.pdf.

Jang:2011:ERC

[JJK^{+}11] Jae-Wan Jang, Myeongjae Jeon, Hyo-Sil Kim, Heeseung Jo, Jin-Soo Kim, and Seungryoul Maeng. Energy reduction in consolidated servers through memory-aware virtual machine scheduling. IEEE Transactions on Computers, 60(4):552– 564, April 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jantz:2013:PPO

[JK13] Michael R. Jantz and Prasad A. Kulkarni. Performance potential of optimization phase selection during dynamic JIT compilation. ACM SIGPLAN Notices, 48(7):131–142, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Jarray:2015:DAV

[JK15] Abdallah Jarray and Ahmed Karmouch. Decomposition approaches for virtual network embedding with one-shot node and link mapping. IEEE/ACM Transactions on Networking, 23(3):1012–1025, June 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Joe:2017:EDI

[JK17] Hyunwoo Joe and Hyungshin Kim. Effects of dynamic isolation for full virtualized RTOS and GPOS guests. Future Generation Computer Systems, 70(??):26–41, May 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X16308020.

Jaffer:2015:IRD

[JKB15] Shehbaz Jaffer, Piyus Kedia, and Sorav Bansal. Improving remote desktopping through adaptive record/replay. ACM SIGPLAN Notices, 50(7):161–172, July 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joshi:2005:DPP

[JKDC05] Ashlesha Joshi, Samuel T. King, George W. Dunlap, and Peter M. Chen. Detecting past and present intrusions through vulnerability-specific predicates. Operating Systems Review, 39(5):91–104, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jo:2010:TFT

[JKJ^{+}10] Heeseung Jo, Hwanju Kim, Jae-Wan Jang, Joonwon Lee, and Seungryoul Maeng. Transparent fault tolerance of device drivers for virtual machines. IEEE Transactions on Computers, 59(11):1466–1479, November 2010. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5432158.

Jeong:2013:AVM

[JKK^{+}13] Jinkyu Jeong, Sung-Hun Kim, Hwanju Kim, Joonwon Lee, and Euiseong Seo. Analysis of virtual machine livemigration as a method for power-capping. The Journal of Supercomputing, 66(3):1629–1655, December 2013. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11227-013-0956-1.

Jansen:2008:SVC

[JM08] Sam Jansen and Anthony McGregor. Static virtualization of C source code. Software—Practice and Experience, 38(4):

397–416, April 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jim-Min:1992:IES

[JMSLM92] Lin Jim-Min, Rong Tsai Shang, and Tseng Li-Ming. Integrating existing software packages using the virtual machine technique. The Journal of Systems and Software, 18(3):207– ??, July 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jin:2015:PSV

[JN15] Dong Jin and David M. Nicol. Parallel simulation and virtualmachine-based emulation of software-defined networks. ACM Transactions on Modeling and Computer Simulation, 26(1): 8:1–8:??, December 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Jeyarani:2012:DIA

[JNR12] R. Jeyarani, N. Nagaveni, and R. Vasanth Ram. Design and implementation of adaptive power-aware virtual machine provisioner (APA–VMP) using swarm intelligence. Future Generation Computer Systems, 28(5):811–821, May 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X11001130.

Joos:2006:OHE

[Joo06] Thomas Joos. Das Online-Handbuch: enthüllt: Providerwahl, E-Mail, eBay, Google, Skype, Chatten, Instant messaging und vieles mehr; [auf der CD: Firefox 1.5, Thunderbird 1.5, OpenOffice 2.01, Oleco NetLCR, Netstumbler, VMWare Player, WINRar und vieles mehr]. Markt-+-Technik-Verl., München, Germany, 2006. ISBN 3-8272-4076-X. 905 pp. LCCN ???? EUR 39.90 (DE).

Joos:2009:MWS

[Joo09] Thomas Joos. Microsoft Windows Server 2008 — das Handbuch. (German) Microsoft Windows Server 2008 — the Handbook. Microsoft Press, Unterschleißheim, second edition, 2009. ISBN 3-86645-130-X. 1356 pp. LCCN ???? URL http:

//deposit.d-nb.de/cgi-bin/dokserv?id=3161715.

Jouannaud:1985:FPL

[Jou85] Jean-Pierre Jouannaud, editor. Functional programming languages and computer architecture: Nancy, France, September 1985, volume 201 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNCSD9. ISBN 0-387-15975-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1

L43 no.201. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0201.htm;   http://www.
springerlink.com/openurl.asp?genre=issue&issn=0302-
9743&volume=201.

Joubert:1994:PCT

[JPTE94] G. R. Joubert, F. J. Peters, D. Trystram, and D. J. Evans, editors. Parallel computing: trends and applications: proceedings of the international conference ParCo93, Grenoble, France, 7–10 September 1993, volume 9 of Advances in parallel computing. North-Holland, Amsterdam, The Netherlands, 1994. ISBN 0-444-81841-3. LCCN QA76.58 .P3794 1993.

Jin:2015:CCC

[JQWG15] Hai Jin, Hanfeng Qin, Song Wu, and Xuerong Guo. CCAP: A cache contention-aware virtual machine placement approach for HPC cloud. International Journal of Parallel Programming, 43(3):403–420, June 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10766-013-0286-1.

Jacob:2002:CAP

[JR02] Matthias Jacob and Keith Randall. Cross-architectural performance portability of a Java Virtual Machine implementation. In USENIX [USE02], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38 J42 2002. URL http:

//www.usenix.org/publications/library/proceedings/
javavm02/jacob.html.

Jin:2015:HAS

[JSHM15] Seongwook Jin, Jinho Seol, Jaehyuk Huh, and Seungryoul Maeng. Hardware-assisted secure resource accounting under a vulnerable hypervisor. ACM SIGPLAN Notices, 50(7):201– 213, July 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jantz:2013:FAG

[JSK^{+}13] Michael R. Jantz, Carl Strickland, Karthik Kumar, Martin Dimitrov, and Kshitij A. Doshi. A framework for application guidance in virtual memory systems. ACM SIGPLAN Notices, 48(7):155–166, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Juola:2007:PCO

[Juo07] Patrick Juola. Principles of computer organization and assembly language: using the Java Virtual Machine. Pearson/Prentice Hall, Upper Saddle River, NJ, USA, 2007. ISBN 0-13-148683-7 (paperback). xv + 317 pp. LCCN

QA76.9.C643 J96 2007. URL http://www.loc.gov/catdir/
toc/ecip073/2006034154.html.

Jin:2017:WCM

[JW17] Yichao Jin and Yonggang Wen. When cloud media meet network function virtualization: Challenges and applications. IEEE MultiMedia, 24(3):72–82, July/September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (elec-

tronic). URL https://www.computer.org/csdl/mags/mu/
2017/03/mmu2017030072-abs.html.

Jia:2015:DRA

[JWH^{+}15] Xiaohua Jia, Jinhai Wang, Chuanhe Huang, Qin Liu, Kai He, Jing Wang, and Peng Li. Dynamic resource allocation based on energy utility maximization using virtual machines in cloud environment. International Journal of Computer Systems Science and Engineering, 30(6):??, November 2015. CODEN CSSEEI. ISSN 0267-6192.

Jia:2018:OSN

[JWL^{+}18] Yongzheng Jia, Chuan Wu, Zongpeng Li, Franck Le, Alex Liu, Zongpeng Li, Yongzheng Jia, Chuan Wu, Franck Le, and Alex Liu. Online scaling of NFV service chains across geodistributed datacenters. IEEE/ACM Transactions on Networking, 26(2):699–710, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jiang:2012:UNG

[JXL^{+}12] Haitao Jiang, Yun Xu, Yin Liao, Guojie Jin, and Guoliang Chen. UKCF: a new graphics driver cross-platform translation framework for virtual machines. Lecture Notes in Computer Science, 7513:156–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35606-
3_18/.

Jin:2010:GTF

[JXZ^{+}10] Hai Jin, Guofu Xiang, Deqing Zou, Feng Zhao, Min Li, and Chen Yu. A guest-transparent file integrity monitoring method in virtualization environment. Computers and Mathematics with Applications, 60(2):256–266, July 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122110000180.

Jiang:2023:THR

[JYM^{+}23] Zhe Jiang, Kecheng Yang, Yunfeng Ma, Nathan Fisher, Neil Audsley, and Zheng Dong. Towards hard real-time and energy-efficient virtualization for many-core embedded systems. IEEE Transactions on Computers, 72(1):111–126, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jobava:2018:AIT

[JYOB18] Akaki Jobava, Anis Yazidi, B. John Oommen, and Kyrre Begnum. On achieving intelligent traffic-aware consolidation of virtual machines in a data center using learning automata. Journal of Computational Science, 24:290–312, January 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1877750317300054.

Jia:2013:SID

[JYW^{+}13] Ning Jia, Chun Yang, Jing Wang, Dong Tong, and Keyi Wang. SPIRE: improving dynamic binary translation through SPC-indexed indirect branch redirecting. ACM SIG-PLAN Notices, 48(7):1–12, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Jia:2022:GND

[JZY^{+}22] Xingguo Jia, Jin Zhang, Boshi Yu, Xingyue Qian, Zhengwei Qi, and Haibing Guan. GiantVM: a novel distributed hypervisor for resource aggregation with DSM-aware optimizations. ACM Transactions on Architecture and Code Optimization, 19(2):20:1–20:27, June 2022. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https://dl.acm.org/

doi/10.1145/3505251.

Kappes:2024:DFU

[KA24] Giorgos Kappes and Stergios V. Anastasiadis. Diciclo: Flexible user-level services for efficient multitenant isolation. ACM Transactions on Computer Systems, 42(1–2):3:1–3:??, May 2024. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-

7333 (electronic). URL https://dl.acm.org/doi/10.1145/
3639404.

Kiaee:2025:JVC

[KA25] Farkhondeh Kiaee and Ehsan Arianyan. Joint VM and container consolidation with auto-encoder based contribution extraction of decision criteria in Edge–Cloud environment. Journal of Network and Computer Applications, 233(??):??, January 2025. CODEN JNCAF3. ISSN 1084-8045 (print),

1095-8592 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1084804524002261.

Kishani:2025:EEL

[KAA^{+}25] Mostafa Kishani, Sina Ahmadi, Saba Ahmadian, Reza Salkhordeh, Zdenek Becvar, Onur Mutlu, André Brinkmann, and Hossein Asadi. ELICA: Efficient and load balanced I/O cache architecture for hyperconverged infrastructures. IEEE Transactions on Parallel and Distributed Systems, 36(10): 2152–2168, October 2025. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kagawa:2009:WWB

[Kag09] Koji Kagawa. WAPPEN: a Web-based application framework for programming and its bison/flex plug-in. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):383, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Kojima:1983:AMI

[KAH83] Tomihiko Kojima, Hidehiko Akita, and Hisashi Hashimoto. An approach to a machine independent UNIX — UNIX on HITAC M-series virtual machines. In Software Tools Users Group [Sof83], pages 363–369. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Kumar:1993:FHS

[KAJW93] Sanjaya Kumar, James H. Aylor, Barry W. Johnson, and Wm. A. Wulf. A framework for hardware/software codesign. Computer, 26(12):39–45, December 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kalin:1997:NMP

[Kal97] Sari Kalin. News: Microsoft plans Java virtual machine for Unix, WinCE by 1998. JavaWorld: IDG’s magazine for the Java community, 2(5):??, May 1997. CODEN ???? ISSN

1091-8906. URL http://www.javaworld.com/javaworld/
jw-05-1997/jw-05-infoworld.microsoftvm.htm.

Kamnitzer:1975:BXI

[Kam75] Stephen H. Kamnitzer. Bootstrapping XPL from IBM/360 to UNIVAC 1100. ACM SIGPLAN Notices, 10(5):14–20, May 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kamrad:1983:ROA

[Kam83] J. M. Kamrad III. Runtime organization for the Ada language system programs. ACM SIGADA Ada Letters, III(3):58–68, November 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kamga:2013:CFE

[Kam13] Christine Mayap Kamga. CPU frequency emulation based on DVFS. Operating Systems Review, 47(3):34–41, December 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kao:2017:TEF

[Kao17] Chia Hung Kao. Testing and evaluation framework for virtualization technologies. Computing: Archiv für Informatik und Numerik, 99(7):657–677, July 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Karcher:2007:VDX

[Kar07] René Karcher. Virtualisierte DMZ auf Xen-Basis: OpenQRM und Xen. Diplomarbeit, Fachhochschule für Oekonomie und Management, Essen, Germany, 2007.

Kumar:2013:OAM

[KAS13] Narander Kumar, Shalini Agarwal, and Vipin Saxena. Overload avoidance model using optimal placement of virtual machines in cloud data cetres. International Journal of Computer Applications, 73(??):18–25, July 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume73/number11/12786-9963/.

Kumar:2014:DLB

[KAZS14] Narander Kumar, Shalini Agarwal, Taskeen Zaidi, and Vipin Saxena. A distributed load-balancing scheme based on a complex network model of cloud servers. ACM SIGSOFT Software Engineering Notes, 39(6):1–6, November 2014. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kunjir:2017:TAM

[KB17] Mayuresh Kunjir and Shivnath Babu. Thoth in action: memory management in modern data analytics. Proceedings of the VLDB Endowment, 10(12):1917–1920, August 2017. CODEN ???? ISSN 2150-8097.

Karthikeyan:2021:EAS

[KB21] Ramamoorthy Karthikeyan and Venkatachalam Balamurugan. Energy-aware and SLA-guaranteed optimal virtual machine swap and migrate system in cloud-Internet of Things. Concurrency and Computation: Practice and Experience, 33 (10):e6171:1–e6171:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2011:PAP

[KBB11] Kyong Hoon Kim, Anton Beloglazov, and Rajkumar Buyya. Power-aware provisioning of virtual machines for real-time Cloud services. Concurrency and Computation: Practice and Experience, 23(13):1491–1505, September 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kasapovic:2015:PVD

[KBBM15] Suad Kasapovic, Esmir Biberkic, and Lejla Banjanovic-Mehmedovic. Performances of virtualization in design of fully collapsed demilitarized zone. International Journal of Computer Applications, 131(??):1–6, December 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume131/number13/23506-2015907475/.

Kucab:2021:RAI

[KBC21] Michał Kucab, Piotr Boryło, and Piotr Chołda. Remote attestation and integrity measurements with Intel SGX for virtual machines. Computers & Security, 106(??):Article 102300, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001243.

Karmakar:2022:UAN

[KBDK22] Kamalesh Karmakar, Somrita Banerjee, Rajib K. Das, and Sunirmal Khatua. Utilization aware and network I/O intensive virtual machine placement policies for cloud data center. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print),

1095-8592 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1084804522000959.

Katsikas:2021:MHP

[KBK^{+}21] Georgios P. Katsikas, Tom Barbette, Dejan Kostić, Gerald Q. Maguire, Jr., and Rebecca Steinert. Metron: Highperformance NFV service chaining even in the presence of blackboxes. ACM Transactions on Computer Systems, 38(1– 2):3:1–3:45, July 2021. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic). URL https://dl.acm.org/

doi/10.1145/3465628.

Kounga:2012:ESP

[KC12] Gina Kounga and Liqun Chen. Enforcing sticky policies with TPM and virtualization. Lecture Notes in Computer Science, 7222:32–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-32298-3_3/.

Kansal:2016:EAV

[KC16] Nidhi Jain Kansal and Inderveer Chana. Energy-aware virtual machine migration for cloud computing — a firefly optimization approach. Journal of Grid Computing, 14(2):327– 345, June 2016. CODEN ???? ISSN 1570-7873 (print),

1572-9184 (electronic). URL http://link.springer.com/
article/10.1007/s10723-016-9364-0.

Kim:2015:UWM

[KCKC15] Huioon Kim, Kyungwon Chun, Hyounggyu Kim, and Youngjoo Chung. Utilization of workflow management system for virtual machine instance management on cloud. Concurrency and Computation: Practice and Experience, 27(17): 5350–5373, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2014:ECS

[KCS14] Nakku Kim, Jungwook Cho, and Euiseong Seo. Energycredit scheduler: an energy-aware virtual machine scheduler for cloud systems. Future Generation Computer Systems, 32(??):128–137, March 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0167739X1200115X.

Kousiouris:2011:ESW

[KCV11] George Kousiouris, Tommaso Cucinotta, and Theodora Varvarigou. The effects of scheduling, workload type and consolidation scenarios on virtual machine performance and their prediction through optimized artificial neural networks. The Journal of Systems and Software, 84(8):1270–1291, August 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kang:2014:HSA

[KCWH14] Chih-Kai Kang, Yu-Jhang Cai, Chin-Hsien Wu, and Pi-Cheng Hsiu. A hybrid storage access framework for highperformance virtual machines. ACM Transactions on Embedded Computing Systems, 13(5s):157:1–157:??, September 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Khan:2022:RSR

[KCY22] Hassan Mahmood Khan, Fang-Fang Chua, and Timothy Tzen Vun Yap. ReSQoV: a scalable resource allocation model for QoS-satisfied cloud services. Future Internet, 14(5):131, April 26, 2022. CODEN ???? ISSN 1999-5903. URL https:

//www.mdpi.com/1999-5903/14/5/131.

Kumar:1978:PEH

[KD78] B. Kumar and Edward S. Davidson. Performance evaluation of highly concurrent computers by deterministic simulation. Communications of the ACM, 21(11):904–913, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kertesz:2016:PBV

[KDB16] A. Kertesz, J. D. Dombi, and A. Benyi. A pliant-based virtual machine scheduling solution to improve the energy efficiency of IaaS clouds. Journal of Grid Computing, 14(1): 41–53, March 2016. CODEN ???? ISSN 1570-7873 (print),

1572-9184 (electronic). URL http://link.springer.com/
article/10.1007/s10723-015-9336-9.

Karmakar:2020:BAC

[KDK20] Kamalesh Karmakar, Rajib K. Das, and Sunirmal Khatua. Bandwidth allocation for communicating virtual machines in cloud data centers. The Journal of Supercomputing, 76(9): 7268–7289, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https://link.

springer.com/article/10.1007/s11227-019-03128-6.

Keefe:1968:HCP

[Kee68] D. D. Keefe. Hierarchical control programs for systems evaluation. IBM Systems Journal, 7(2):123–133, 1968. CODEN IBMSA7. ISSN 0018-8670.

Keedy:1977:OIS

[Kee77] J. L. Keedy. An outline of the ICL 2900 series system architecture. Australian Computer Journal, 9(2):53–62, 1977. CODEN ACMJB2. ISSN 0004-8917.

Kelly:2006:PMX

[Kel06] Ivan Kelly. Porting MINIX to Xen. Final year project, Department of Computer Science, University of Limerick, Limerick, Ireland, May 8, 2006. vii + 45 pp. URL http:/

/minixonxen.skynet.ie/cgi-bin/trac.cgi/attachment/
wiki/Report/Report.pdf?format=raw; http://www.minix3.
org/theses/kelly-xen.pdf.

Kent:1980:PNV

[Ken80] S. A. Kent. A programmable network virtual machine. Computer Networks: The International Journal of Distributed Informatique, 4(3):125–137, June 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Kermarrec:1988:SEA

[Ker88] Y. Kermarrec. Some experiments with Ada. Ada User, 9(2): 79–82, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Kereki:2015:CCC

[Ker15] Federico Kereki. Concerning containers’ connections: on Docker networking. Linux Journal, 2015(254):2:1–2:??, June 2015. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-

3827 (electronic). URL http://dl.acm.org/ft_gateway.
cfm?id=2807680.

Kelem:1991:SMV

[KF91] N. L. Kelem and R. J. Feiertag. A separation model for virtual machine monitors. In IEEE [IEE91], pages 78–87 (or 78–86??). ISBN 0-8186-9168-9, 0-8186-2168-0, 0-8186-6168-2. LCCN QA76.9.A25; QA76.9.A25 I34 1991; QA76.9.A25 S95 1991eb; QA76.9.A25 S95 1991. IEEE Catalog Number 91CH2986-8. IEEE Computer Society Order Number 2168.

Kontoudis:2018:SAV

[KF18] Dimitris Kontoudis and Panayotis Fouliras. A statistical approach to virtual server resource management. Concurrency and Computation: Practice and Experience, 30 (4):??, February 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4335.

Klein:2012:RVM

[KFF12] Casey Klein, Matthew Flatt, and Robert Bruce Findler. The Racket virtual machine and randomized testing. Higher-Order and Symbolic Computation, 25(2–4):209–253, December 2012. CODEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10990-013-9091-1;
 http://link.springer.com/article/10.1007/s10990-
013-9091-1.

Klappheck:2000:BLE

[KGG00] Günther Klappheck, Peter Glinsky, and Frank Gehrke. Das Buch — LINUX Edition 2000: [jetzt zur S.u.S.E. 6.4, Installation und Bedienung von LINUX und seinen Werkzeugen, LINUX im Netzwerk und Internet, LINUX und ISDN, der Desktop KDE 1.x, VMware unter LINUX]. Sybex, Düsseldorf, Germany, 2000. ISBN 3-8155-0175-X. xxviii + 906 pp. LCCN ???? DM 69.95.

Kannan:2017:HDH

[KGGS17] Sudarsun Kannan, Ada Gavrilovska, Vishal Gupta, and Karsten Schwan. HeteroOS: OS design for heterogeneous memory management in datacenter. ACM SIGARCH Computer Architecture News, 45(2):521–534, May 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Kannan:2018:HDH

[KGGS18] Sudarsun Kannan, Ada Gavrilovska, Vishal Gupta, and Karsten Schwan. HeteroOS: OS design for heterogeneous memory management in datacenters. Operating Systems Review, 52(1):13–26, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Knodel:2016:MLR

[KGS16] Oliver Knodel, Paul R. Genssler, and Rainer G. Spallek. Migration of long-running tasks between reconfigurable resources using virtualization. ACM SIGARCH Computer Architecture News, 44(4):56–61, September 2016. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Krsul:2004:VPM

[KGZ^{+}04] Ivan Krsul, Arijit Ganguly, Jian Zhang, Jose A. B. Fortes, and Renato J. Figueiredo. VMPlants: Providing and managing virtual machine execution environments for Grid computing. In ACM [ACM04b], pages 7–?? ISBN 0-7695-2153-3. LCCN QA76.88. IEEE order number E2153. ACM order number 415043.

Khan:2019:TEX

[Kha19] Minhaj Ahmad Khan. Towards efficient XML parsing through minimization of JVM parameter space. The Journal of Supercomputing, 75(7):3693–3711, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kong:2022:NIV

[KHA22] Yuqiang Kong, Yaoping He, and Karlo Abnoosian. Natureinspired virtual machine placement mechanisms: a systematic review. Concurrency and Computation: Practice and Experience, 34(11):e6900:1–e6900:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karnagel:2017:AWP

[KHL17] Tomas Karnagel, Dirk Habich, and Wolfgang Lehner. Adaptive work placement for query processing on heterogeneous computing resources. Proceedings of the VLDB Endowment, 10(7):733–744, March 2017. CODEN ???? ISSN 2150-8097.

Khnaser:2009:VVC

[Khn09] Elias N. Khnaser. VCP: VMware Certified Professional. Exam cram. Que Corporation, Indianapolis, IN, USA, 2009. ISBN 0-7897-3805-8 (paperback). xv + 316 pp. LCCN QA76.3 .K495 2009. Includes CD-ROM.

Kang:2016:MPV

[KHW^{+}16] Junbin Kang, Chunming Hu, Tianyu Wo, Ye Zhai, Benlong Zhang, and Jinpeng Huai. MultiLanes: Providing virtualized storage for OS-level virtualization on manycores. ACM Transactions on Storage, 12(3):12:1–12:??, June 2016. CO-DEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Kim:1984:EVM

[Kim84] K. H. Kim. Evolution of a virtual machine supporting faulttolerant distributed processes at a research laboratory. In IEEE [IEE84b], pages 620–628. ISBN 0-8186-0533-2 (paperback), 0-8186-8533-6 (hard), 0-8186-4533-4 (microfiche). LCCN QA76.9.D3 I5582 1984. URL http://dblp.uni-

trier.de/db/conf/icde/Kim84.html.

Kiperberg:2021:PMC

[Kip21] Michael Kiperberg. Preventing malicious communication using virtualization. Journal of Information Security and Applications (JISA), 61(??):??, September 2021. CODEN ????

ISSN 2214-2126. URL http://www.sciencedirect.com/
science/article/pii/S2214212621001009.

Kissell:2008:TCV

[Kis08] Joe Kissell. Take control of VMware Fusion 2. Take control. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 1-933671-49-1. ???? pp. LCCN QA76.76.O63 K57 2008.

Kalibera:2013:RBR

[KJ13] Tomas Kalibera and Richard Jones. Rigorous benchmarking in reasonable time. ACM SIGPLAN Notices, 48(11):63–74, November 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ISMM ’13 conference proceedings.

Kim:2016:DOF

[KJJ^{+}16] Junghyun Kim, Gangwon Jo, Jaehoon Jung, Jungwon Kim, and Jaejin Lee. A distributed OpenCL framework using redundant computation and data replication. ACM SIGPLAN Notices, 51(6):553–569, June 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:2011:XEC

[KJL11] Hwanju Kim, Heeseung Jo, and Joonwon Lee. XHive: Efficient cooperative caching for virtual machines. IEEE Transactions on Computers, 60(1):106–119, January 2011. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2015:PMS

[KJLY15] Changhyeon Kim, Changho Jeon, Wonjoo Lee, and Sungil Yang. A parallel migration scheme for fast virtual machine relocation on a cloud cluster. The Journal of Supercomputing, 71(12):4623–4645, December 2015. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11227-015-1563-0.

Kim:2007:VPR

[KJM^{+}07] Hyesoon Kim, José A. Joao, Onur Mutlu, Chang Joo Lee, Yale N. Patt, and Robert Cohn. VPC prediction: reducing the cost of indirect branches via hardware-based dynamic devirtualization. ACM SIGARCH Computer Architecture News, 35(2):424–435, May 2007. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Kobayashi:1979:SMC

[KK79] Y. Kobayashi and T. L. Kunii. A simulation of a minicomputer controlled system and its use as a debugging tool. The Computer Journal, 22(1):41–49, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/41.tif; http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/42.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/43.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/44.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/45.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/46.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/47.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/48.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/49.tif.

Kumar:2019:ICL

[KK19] Pawan Kumar and Rakesh Kumar. Issues and challenges of load balancing techniques in cloud computing: a survey. ACM Computing Surveys, 51(6):120:1–120:??, February 2019. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341

(electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3281010.

Kiani:2021:NAP

[KK21] Mohsen Kiani and Mohammad Reza Khayyambashi. A network-aware and power-efficient virtual machine placement scheme in cloud datacenters based on chemical reaction optimization. Computer Networks (Amsterdam, Netherlands: 1999), 196(??):??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1389128621002954.

Kertesz:2014:ISA

[KKB14] A. Kertesz, G. Kecskemeti, and I. Brandic. An interoperable and self-adaptive approach for SLA-based service virtualization in heterogeneous Cloud environments. Future Generation Computer Systems, 32(??):54–68, March 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X12001124.

Kim:2016:SCD

[KKC^{+}16] Channoh Kim, Sungmin Kim, Hyeon Gyu Cho, Dooyoung Kim, Jaehyeok Kim, Young H. Oh, Hakbeom Jang, and Jae W. Lee. Short-circuit dispatch: accelerating virtual machine interpreters on embedded processors. ACM SIGARCH Computer Architecture News, 44(3):291–303, June 2016. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Kabamba:2024:VLO

[KKD24] Herve M. Kabamba, Matthew Khouzam, and Michel R. Dagenais. Vnode: Low-overhead transparent tracing of Node.jsbased microservice architectures. Future Internet, 16(1):13, December 29, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/1999-5903/16/1/13.

Karn:2019:DAA

[KKE19] R. R. Karn, P. Kudva, and I. A. M. Elfadel. Dynamic autoselection and autotuning of machine learning models for cloud network analytics. IEEE Transactions on Parallel and Distributed Systems, 30(5):1052–1064, May 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2014:VPT

[KKH14] Daehoon Kim, Hwanju Kim, and Jaehyuk Huh. vCache: Providing a transparent view of the LLC in virtualized environments. IEEE Computer Architecture Letters, 13(2):109–112, July/December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Kim:2013:DBC

[KKJ^{+}13] Hwanju Kim, Sangwook Kim, Jinkyu Jeong, Joonwon Lee, and Seungryoul Maeng. Demand-based coordinated scheduling for SMP VMs. ACM SIGARCH Computer Architecture News, 41(1):369–380, March 2013. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Kim:2014:VAM

[KKJL14] Hwanju Kim, Sangwook Kim, Jinkyu Jeong, and Joonwon Lee. Virtual asymmetric multiprocessor for interactive performance of consolidated desktops. ACM SIGPLAN Notices, 49(7):29–40, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:2018:LIG

[KKK^{+}18] Dongkyun Kim, Yong-Hwan Kim, Ki-Hyun Kim, Joo-Beom Kim, Gi-Seong You, and Joon-Min Gil. Logically isolated group network for virtual convergence environment over SD– WAN. The Journal of Supercomputing, 74(12):6742–6752, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kokkinos:2016:SLM

[KKLV16] Panagiotis Kokkinos, Dimitris Kalogeras, Anna Levin, and Emmanouel Varvarigos. Survey: Live migration and disaster recovery over long-distance networks. ACM Computing Surveys, 49(2):26:1–26:??, September 2016. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Kawahito:2013:IRF

[KKM^{+}13] Motohiro Kawahito, Hideaki Komatsu, Takao Moriyama, Hiroshi Inoue, and Toshio Nakatani. Idiom recognition framework using topological embedding. ACM Transactions on Architecture and Code Optimization, 10(3):13:1–13:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Koksal:2012:CC

[KKS12] Ali Sinan Köksal, Viktor Kuncak, and Philippe Suter. Constraints as control. ACM SIGPLAN Notices, 47(1):151–164, January 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kirova:2019:IMV

[KKS^{+}19] Veronika Kirova, Kirill Karpov, Eduard Siemens, Irina Zander, Oksana Vasylenko, Dmitry Kachan, and Sergii Maksymov. Impact of modern virtualization methods on timing precision and performance of high-speed applications. Future Internet, 11(8):179, August 16, 2019. CODEN ???? ISSN

1999-5903. URL https://www.mdpi.com/1999-5903/11/8/
179.

Kawai:2017:VWD

[KKTM17] Takaaki Kawai, Shigeru Kaneda, Mineo Takai, and Hiroshi Mineno. A virtual WLAN device model for high-fidelity wireless network emulation. ACM Transactions on Modeling and Computer Simulation, 27(3):17:1–17:??, September 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kocoloski:2013:ICN

[KL13] Brian Kocoloski and John Lange. Improving compute node performance using virtualization. The International Journal of High Performance Computing Applications, 27(2):124–135, May 2013. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-

2846 (electronic). URL http://hpc.sagepub.com/content/
27/2/124.full.pdf+html.

Kong:2014:SGE

[KL14] Fanxin Kong and Xue Liu. A survey on green-energy-aware power management for datacenters. ACM Computing Surveys, 47(2):30:1–30:??, November 2014. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Kyle:2015:ADA

[KLF^{+}15] Stephen Kyle, Hugh Leather, Björn Franke, Dave Butcher, and Stuart Monteith. Application of domain-aware binary fuzzing to aid Android virtual machine testing. ACM SIG-PLAN Notices, 50(7):121–132, July 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kwon:2022:SFF

[KLK^{+}22] Dongup Kwon, Wonsik Lee, Dongryeong Kim, Junehyuk Boo, and Jangwoo Kim. SmartFVM: a fast, flexible, and scalable hardware-based virtualization for commodity storage devices. ACM Transactions on Storage, 18(2):12:1–12:27, May 2022. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3511213.

Kuo:2018:DCV

[KLLT18] Tung-Wei Kuo, Bang-Heng Liou, Kate Ching-Ju Lin, and Ming-Jer Tsai. Deploying chains of virtual network functions: On the relation between link and server usage. IEEE/ ACM Transactions on Networking, 26(4):1562–1576, August 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kulkarni:2020:RAE

[KLR^{+}20] Sameer G. Kulkarni, Guyue Liu, K. K. Ramakrishnan, Mayutan Arumaithurai, Timothy Wood, and Xiaoming Fu. RE-INFORCE: Achieving efficient failure resiliency for network function virtualization-based services. IEEE/ACM Transactions on Networking, 28(2):695–708, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

URL https://dl.acm.org/doi/abs/10.1109/TNET.2020.
2969961.

Kang:2020:PMT

[KLY20] JiHun Kang, JongBeom Lim, and HeonChang Yu. Partial migration technique for GPGPU tasks to prevent GPU memory starvation in RPC-based GPU virtualization. Software— Practice and Experience, 50(6):948–972, June 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kiefer:2013:RDN

[KM13a] Kurt E. Kiefer and Louise E. Moser. Replay debugging of non-deterministic executions in the Kernel-based Virtual Machine. Software—Practice and Experience, 43(11):1261–1281, November 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kiefer:2013:SIP

[KM13b] Kurt E. Kiefer and Louise E. Moser. Special issue papers: Replay debugging of non-deterministic executions in the Kernelbased Virtual Machine. Software—Practice and Experience, 43(11):1261–1281, November 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kimovski:2018:DEE

[KMG^{+}18] Dragi Kimovski, Attila Marosi, Sandi Gec, Nishant Saurabh, Attila Kertesz, Gabor Kecskemeti, Vlado Stankovski, and Radu Prodan. Distributed environment for efficient virtual machine image management in federated Cloud architectures. Concurrency and Computation: Practice and Experience, 30 (20):e4220:1–e4220:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krieger:2010:EMC

[KMK10] Orran Krieger, Phil McGachey, and Arkady Kanevsky. Enabling a marketplace of clouds: VMware’s vCloud director. Operating Systems Review, 44(4):103–114, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kashyap:2016:OSA

[KMK16] Sanidhya Kashyap, Changwoo Min, and Taesoo Kim. Opportunistic spinlocks: Achieving virtual machine scalability in the clouds. Operating Systems Review, 50(1):9–16, January 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Khazaei:2013:PCC

[KMM13] Hamzeh Khazaei, Jelena Misic, and Vojislav B. Misic. Performance of cloud centers with high degree of virtualization under batch task arrivals. IEEE Transactions on Parallel and Distributed Systems, 24(12):2429–2438, December 2013. CODEN ITDSEO. ISSN 1045-9219.

Kalibera:2014:FAS

[KMMV14] Tomas Kalibera, Petr Maj, Floreal Morandat, and Jan Vitek. A fast abstract syntax tree interpreter for R. ACM SIGPLAN Notices, 49(7):89–102, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kuperman:2016:PR

[KMN^{+}16] Yossi Kuperman, Eyal Moscovici, Joel Nider, Razya Ladelsky, Abel Gordon, and Dan Tsafrir. Paravirtual remote I/O. ACM SIGPLAN Notices, 51(4):49–65, April 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kessaci:2014:MSL

[KMT14] Yacine Kessaci, Nouredine Melab, and El-Ghazali Talbi. A multi-start local search heuristic for an energy efficient VMs assignment on top of the OpenNebula cloud manager. Future Generation Computer Systems, 36(??):237–256, July 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X13001520.

Kamran:2018:QAV

[KN18] Kamran and Babar Nazir. QoS-aware VM placement and migration for hybrid cloud infrastructure. The Journal of Supercomputing, 74(9):4623–4646, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Knaggs:1993:PTA

[Kna93] Peter J. Knaggs. Practical and Theoretical Aspects of Forth Software Development. PhD thesis, School of Computing and Mathematics, University of Teesside, Middlesbrough, Cleve-

land. UK, March 1993. URL ftp://cis.paisley.ac.uk/
forth/thesis.

Kim:2018:FSS

[KNHH18] Myungsun Kim, Soonhyun Noh, Jinhwa Hyeon, and Seongsoo Hong. Fair-share scheduling in single-ISA asymmetric multicore architecture via scaled virtual runtime and load redistribution. Journal of Parallel and Distributed Computing, 111(??):174–186, January 2018. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731517302423.

Kasprzyk:2002:APV

[KNT02] Leszek Kasprzyk, Ryszard Nawrowski, and Andrzej Tomczewski. Application of a parallel virtual machine for the analysis of a luminous field. Lecture Notes in Computer Science, 2474:122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/series/0558/bibs/
2474/24740122.htm; http://link.springer.de/link/service/
series/0558/papers/2474/24740122.pdf.

Kolodin:2019:HMR

[Kol19] Denis Kolodin. Hands-on Microservices with Rust: Build, Test, and Deploy Scalable and Reactive Microservices with Rust 2018. Packt Publishing, Birmingham, UK, 2019. ISBN 1-78934-198-1, 1-78934-275-9. 511 (est.) pp. LCCN QA76.73.R87. URL http://proquest.

safaribooksonline.com/?fpi=9781789342758.

Kotsovinos:2010:VBC

[Kot10] Evangelos Kotsovinos. Virtualization: Blessing or curse? ACM Queue: Tomorrow’s Computing Today, 8(11):40, November 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kotsovinos:2011:VBC

[Kot11] Evangelos Kotsovinos. Virtualization: blessing or curse? Communications of the ACM, 54(1):61–65, January 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kourai:2011:FCP

[Kou11] Kenichi Kourai. Fast and correct performance recovery of operating systems using a virtual machine monitor. ACM SIGPLAN Notices, 46(7):99–110, July 2011. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kovacs:2019:SPA

[Kov19] József Kovács. Supporting programmable autoscaling rules for containers and virtual machines on clouds. Journal of Grid Computing, 17(4):813–829, December 2019. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10723-019-09488-w;   https://link.springer.
com/content/pdf/10.1007/s10723-019-09488-w.pdf.

Kaneda:2005:VMM

[KOY05] Kenji Kaneda, Yoshihiro Oyama, and Akinori Yonezawa. A virtual machine monitor for utilizing non-dedicated clusters. In Vrable [Vra05], pages 1–11. ISBN 1-59593-079-5. LCCN QA76.6 .S9196 2005; QA76.6; QA76.6 .S9196 2005eb; Internet.

Kernighan:1999:REL

[KP99] Brian W. Kernighan and Rob Pike. Regular expressions: Languages, algorithms, software. Dr. Dobb’s Journal of Software Tools, 24(4):19–22, April 1999. CODEN DDJOEB. ISSN

1044-789X. URL http://www.ddj.com/ftp/1999/1999_04/
regexp.txt;   http://www.ddj.com/ftp/1999/1999_04/

regexp.zip. See also [Tho68, Cox07, Cox09, Cox10, Cox12].

Kim:2015:CBR

[KP15] Chulmin Kim and Kyu Ho Park. Credit-based runtime placement of virtual machines on a single NUMA system for QoS of data access performance. IEEE Transactions on Computers, 64(6):1633–1646, ???? 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2020:RTS

[KPHA20] T. Kim, C. H. Park, J. Huh, and J. Ahn. Reconciling time slice conflicts of virtual machines with dual time slice for clouds. IEEE Transactions on Parallel and Distributed Systems, 31(10):2453–2465, 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kelsey:1994:TSI

[KR94] Richard A. Kelsey and Jonathan A. Rees. A tractable Scheme implementation. Lisp and Symbolic Computation, 7(4):315– 335, ???? 1994. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Kumar:2016:HTA

[KR16] Mohan Raj Velayudhan Kumar and Shriram Raghunathan. Heterogeneity and thermal aware adaptive heuristics for energy efficient consolidation of virtual machines in infrastructure clouds. Journal of Computer and System Sciences, 82(2):191–212, March 2016. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002200001500080X.

Kim:2018:PSC

[KR18] Hyoseung Kim and Ragunathan (Raj) Rajkumar. Predictable shared cache management for multi-core real-time virtualization. ACM Transactions on Embedded Computing Systems, 17(1):22:1–22:??, January 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kratzer:1990:MPS

[Kra90] Steven G. Kratzer. Massively parallel sparse-matrix computations. Technical report SRC-TR-90-008, Supercomputing Research Center: IDA, Lanham, MD, USA, February 1, 1990. 14 pp.

Kedlaya:2014:DDL

[KRCH14] Madhukar N. Kedlaya, Behnam Robatmili, Cglin Cascaval, and Ben Hardekopf. Deoptimization for dynamic language JITs on typed, stack-based virtual machines. ACM SIG-PLAN Notices, 49(7):103–114, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kundu:2012:MVA

[KRG^{+}12] Sajib Kundu, Raju Rangaswami, Ajay Gulati, Ming Zhao, and Kaushik Dutta. Modeling virtualized applications using machine learning techniques. ACM SIGPLAN Notices, 47(7): 3–14, July 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Kroeker:2009:EV

[Kro09] Kirk L. Kroeker. The evolution of virtualization. Communications of the ACM, 52(3):18–20, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kanizo:2017:OVB

[KRS^{+}17] Yossi Kanizo, Ori Rottenstreich, Itai Segall, Jose Yallouz, Yossi Kanizo, Ori Rottenstreich, Itai Segall, and Jose Yallouz. Optimizing virtual backup allocation for middleboxes. IEEE/ ACM Transactions on Networking, 25(5):2759–2772, October 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Karger:2008:VVM

[KS08a] Paul A. Karger and David R. Safford. Virtualization: I/O for virtual machine monitors: Security and performance issues. IEEE Security & Privacy, 6(5):16–23, September/October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

King:2008:GEI

[KS08b] Samuel T. King and Sean W. Smith. Guest Editors’ introduction: Virtualization and security: Back to the future. IEEE Security & Privacy, 6(5):15, September/October 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kelbley:2010:WSR

[KS10] John Kelbley and Mike Sterling. Windows Server 2008 R2 Hyper-V: insiders guide to Microsoft’s Hypervisor. Wiley, New York, NY, USA, 2010. ISBN 0-470-62700-X (paperback). ???? pp. LCCN QA76.9.V5 K453 2010.

Kaufmann:2013:SCO

[KS13] Marco Kaufmann and Rainer G. Spallek. Superblock compilation and other optimization techniques for a Java-based DBT machine emulator. ACM SIGPLAN Notices, 48(7):33– 40, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Kesavaraja:2018:QEC

[KS18a] D. Kesavaraja and A. Shenbagavalli. QoE enhancement in cloud virtual machine allocation using Eagle strategy of hybrid krill herd optimization. Journal of Parallel and Distributed Computing, 118 (Part 2)(??):267–279, ???? 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731517302459.

Khorandi:2018:NCO

[KS18b] Sina Mahmoodi Khorandi and Mohsen Sharifi. Nonclairvoyant online scheduling of synchronized jobs on virtual clusters. The Journal of Supercomputing, 74(6):2353–2384, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kavitha:2020:PAA

[KS20a] Kadarla Kavitha and S. C. Sharma. Performance analysis of ACO-based improved virtual machine allocation in cloud for IoT-enabled healthcare. Concurrency and Computation: Practice and Experience, 32(21):e5613:1–e5613:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kourai:2020:FSC

[KS20b] Kenichi Kourai and Kouta Sannomiya. Flexible service consolidation with nested virtualization and library operating systems. Software—Practice and Experience, 50(1):3–21, January 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kong:2008:PTD

[KSLA08] Jiantao Kong, Karsten Schwan, Min Lee, and Mustaque Ahamad. Protectit: trusted distributed services operating on sensitive data. Operating Systems Review, 42(4):137–147, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kavvadia:2015:EVM

[KSO^{+}15] Eleni Kavvadia, Spyros Sagiadinos, Konstantinos Oikonomou, Giorgos Tsioutsiouliklis, and Sonia Aïssa. Elastic virtual machine placement in cloud computing network environments. Computer Networks (Amsterdam, Netherlands: 1999), 93 (part 3)(??):435–447, December 24, 2015. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1389128615003631.

Keller:2010:NVC

[KSRL10] Eric Keller, Jakub Szefer, Jennifer Rexford, and Ruby B. Lee. NoHype: virtualized cloud infrastructure without the virtualization. ACM SIGARCH Computer Architecture News, 38 (3):350–361, June 2010. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Kelbley:2009:WSH

[KSS09] John Kelbley, Mike Sterling, and Allen Stewart. Windows Server 2008 Hyper-V: insiders guide to Microsoft’s Hypervisor. Wiley, New York, NY, USA, 2009. ISBN 0-470-50398-X (electronic), 0-470-44096-1 (paperback). xxii + 361 pp. LCCN QA76.9.V5 K45 2009eb. URL http://www.

books24x7.com/marc.asp?bookid=31993.

Kotsifakou:2018:HHP

[KSS^{+}18] Maria Kotsifakou, Prakalp Srivastava, Matthew D. Sinclair, Rakesh Komuravelli, Vikram Adve, and Sarita Adve. HPVM: heterogeneous parallel virtual machine. ACM SIGPLAN Notices, 53(1):68–80, January 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karthikeyan:2020:ECA

[KSS^{+}20] K. Karthikeyan, R. Sunder, K. Shankar, S. K. Lakshmanaprabu, V. Vijayakumar, Mohamed Elhoseny, and Gunasekaran Manogaran. Energy consumption analysis of virtual machine migration in cloud using hybrid swarm optimization (ABC-BA). The Journal of Supercomputing, 76(5): 3374–3390, May 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Karthikeyan:2023:RNE

[KSS^{+}23] K. Karthikeyan, R. Sunder, K. Shankar, S. K. Lakshmanaprabu, V. Vijayakumar, Mohamed Elhoseny, and Gunasekaran Manogaran. Retraction note: Energy consumption analysis of virtual machine migration in cloud using hybrid swarm optimization (ABC-BA). The Journal of Supercomputing, 79(5):5832–5833, March 2023. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-022-04860-2. See [KSS^{+}20].

Karamzadeh:2024:RCS

[KSS24] Amirmohammad Karamzadeh and Alireza Shameli-Sendi. Reducing cold start delay in serverless computing using lightweight virtual machines. Journal of Network and Computer Applications, 232(??):??, December 2024. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1084804524002078.

Katsaros:2016:EFE

[KSSG16] Gregory Katsaros, Pascal Stichler, Josep Subirats, and Jordi Guitart. Estimation and forecasting of ecological efficiency of virtual machines. Future Generation Computer Systems, 55(??):480–494, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X15000035.

Kennedy:2023:MTG

[KSVR23] Jason Kennedy, Vishal Sharma, Blesson Varghese, and Carlos Reaño. Multi-tier GPU virtualization for deep learning in cloud-edge systems. IEEE Transactions on Parallel and Distributed Systems, 34(7):2107–2123, July 2023. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Khosravi:2017:OVM

[KTB17] Atefeh Khosravi, Adel Nadjaran Toosi, and Rajkumar Buyya. Online virtual machine migration for renewable energy usage maximization in geographically distributed cloud data centers. Concurrency and Computation: Practice and Experience, 29(18), September 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2024:EPV

[KU24] K. Dinesh Kumar and E. Umamaheswari. An efficient proactive VM consolidation technique with improved LSTM network in a cloud environment. Computing: Archiv für Informatik und Numerik, 106(1):1–28, January 2024. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00607-023-01214-5.

Kutter:1992:STE

[Kut92] Michael Kutter. Simulation and testboard for embeddedsystem design. Dr. Dobb’s Journal of Software Tools, 17(12): 62, 66–68, December 1992. CODEN DDJOEB. ISSN 1044-789X.

Kappel:2009:MVH

[KVV09] Jason Kappel, Toby J. Velte, and Anthony T. Velte. Microsoft virtualization with Hyper-V. McGraw-Hill, New York, NY, USA, 2009. ISBN 0-07-161403-6. xviii + 430 pp. LCCN QA76.9.V5 K37 2009.

Kerridge:1980:STC

[KW80] J. M. Kerridge and N. Willis. A simulator for teaching computer architecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 12(2):65–71, July 1980. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kang:2013:HPP

[KW13] Hui Kang and Jennifer L. Wong. To hardware prefetch or not to prefetch?: a virtualized environment study and core binding approach. ACM SIGPLAN Notices, 48(4):357–368, April 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kist:2019:FFG

[KWZ^{+}19] Maicon Kist, Juliano Araújo Wickboldt, Lisandro Zambenedetti Granville, Juergen Rochol, Luiz A. DaSilva, and Cristiano Bonato Both. Flexible fine-grained baseband processing with network functions virtualization: Benefits and impacts. Computer Networks (Amsterdam, Netherlands: 1999), 151(??):158–165, March 14, 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128619301033.

Koskinen:2016:RCR

[KY16] Eric Koskinen and Junfeng Yang. Reducing crash recoverability to reachability. ACM SIGPLAN Notices, 51(1):97–108, January 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kwon:2017:IHP

[KYP^{+}17] Youngjin Kwon, Hangchen Yu, Simon Peter, Christopher J. Rossbach, and Emmett Witchel. Ingens: Huge page support for the OS and hypervisor. Operating Systems Review, 51 (1):83–93, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Khaleel:2021:AVM

[KZ21] Mustafa I. Khaleel and Michelle M. Zhu. Adaptive virtual machine migration based on performance-to-power ratio in fog-enabled cloud data centers. The Journal of Supercomputing, 77(10):11986–12025, October 2021. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-021-03753-0.

Karger:1990:VSK

[KZB^{+}90] P. A. Karger, M. E. Zurko, D. W. Bonin, A. H. Mason, and C. E. Kahn. A VMM security kernel for the VAX architecture.

In IEEE [IEE90a], pages 2–19. ISBN 0-8186-2060-9, 0-8186-6060-0, 0-8186-9060-7. LCCN QA76.9.A25; QA76.9.A25 S95 1990eb; QA76.9.A25 I34 1990. IEEE Computer Society Order Number 2060. IEEE Catalog Number 90CH2884-5.

Lamming:1975:LVM

[Lam75] M. Lamming. LITL virtual machine. fixed or variable size blocks. Technical Report QMW-DCS-1975-085; QMW-DCS-1975-091, Queen Mary College, Department of Computer Science, June 1975.

Larisch:2009:PMH

[Lar09] Dirk Larisch. Praxisbuch Microsoft Hyper-V: Installation, Konfiguration und Systemverwaltung von Hyper-V für Windows Server 2008 und Hyper-V Server. (German) [Microsoft Hyper-V Practice Book: Installation, Configuration, and Systems Management of Hyper-V for Windows Server 2008 and Hyper-V Server]. Carl Hanser, München, Germany, 2009. ISBN 3-446-41687-0. 500 pp. LCCN ???? URL http://

deposit.d-nb.de/cgi-bin/dokserv?id=3200334.

Lau:1987:OCV

[Lau87] A. S. K. Lau. Overview of the CLIP7 virtual machine. Report 87/6, Department of Physics and Astronomy, University College London, London, UK, 1987.

Laverick:2010:VVI

[Lav10] Mike Laverick. VMware vSphere 4 implementation. McGraw-Hill, New York, NY, USA, 2010. ISBN 0-07-166452-1. xxviii + 659 pp. LCCN QA76.9.V5 L39 2010.

Lawton:2000:PVM

[Law00] Kevin P. Lawton. Plex86: An 180x86 virtual machine. In USENIX [USE00a], page ?? ISBN 1-880446-17-0.

LCCN ???? URL http://www.usenix.org/publications/
library/proceedings/als2000/lawton.html.

Liang:1998:DCL

[LB98] Sheng Liang and Gilad Bracha. Dynamic class loading in the Java Virtual Machine. ACM SIGPLAN Notices, 33(10): 36–44, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lin:2012:UKT

[LBF12] Yi Lin, Stephen M. Blackburn, and Daniel Frampton. Unpicking the knot: teasing apart VM/application interdependencies. ACM SIGPLAN Notices, 47(7):181–190, July 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Li:2016:ICV

[LBL16] Mingfu Li, Jingping Bi, and Zhongcheng Li. Improving consolidation of virtual machine based on virtual switching overhead estimation. Journal of Network and Computer Applications, 59(??):158–167, January 2016. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451500171X.

Laadan:2007:DPV

[LBP^{+}07] Oren Laadan, Ricardo A. Baratto, Dan B. Phung, Shaya Potter, and Jason Nieh. DejaView: a personal virtual computer recorder. Operating Systems Review, 41(6):279–292, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Le:2011:REC

[LBZ^{+}11] Kien Le, Ricardo Bianchini, Jingru Zhang, Yogesh Jaluria, Jiandong Meng, and Thu D. Nguyen. Reducing electricity cost through virtual machine placement in high performance computing clouds. In Lathrop et al. [LCK11], pages 22:1– 22:12. ISBN 1-4503-0771-X. LCCN ????

Levis:2002:MTV

[LC02] Philip Levis and David Culler. Maté: a tiny virtual machine for sensor networks. ACM SIGARCH Computer Architecture News, 30(5):85–95, December 2002. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Larson:2009:WSHb

[LC09a] Robert Larson and Janique Carbone. Windows Server 2008 Hyper-V — die technische Referenz: [Insiderinformationen und nützliche Tools direkt von der Quelle]: [Windows Server 2008 Hyper-V: the technical reference — insider information and useful tools — direct from the source]. Microsoft Press, Unterschleißheim, Germany, 2009. ISBN 3-86645-926-2. xxviii + 739 pp. LCCN ????

Larson:2009:WSHa

[LC09b] Robert Larson and Janique Carbone. Windows Server 2008 Hyper-V resource kit. Microsoft Press, Bellevue, WA, USA, 2009. ISBN 0-7356-2517-4. xxxv + 744 pp. LCCN QA76.9.V5

L37 2009. URL http://proquest.safaribooksonline.
com/9780735625174.

Liu:2013:SPV

[LC13] Denghui Liu and Jinli Cao. Scheduling para-virtualized virtual machines based on events. Future Generation Computer Systems, 29(5):1130–1139, July 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X12002348.

Li:2014:USI

[LC14] Yusen Li and Wentong Cai. Update schedules for improving consistency in multi-server distributed virtual environments. Journal of Network and Computer Applications, 41(??):263– 273, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print),

1095-8592 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1084804514000083.

Luo:2012:PNV

[LCFL12] Juan Luo, Lei Chen, Shan Fu, and Renfa Li. Parallel network virtualization resource management system. Lecture Notes in Computer Science, 7419:69–77, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33050-6_8/.

Lathrop:2011:SPI

[LCK11] Scott Lathrop, Jim Costa, and William Kramer, editors. SC’11: Proceedings of 2011 International Conference for High Performance Computing, Networking, Storage and Analysis, Seattle, WA, November 12–18 2011. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4503-0771-X. LCCN ????

Lin:2014:IQA

[LCL14] Jenn-Wei Lin, Chien-Hung Chen, and Chi-Yi Lin. Integrating QoS awareness with virtualization in cloud computing systems for delay-sensitive applications. Future Generation Computer Systems, 37(??):478–487, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X13002987.

Liu:2023:PCT

[LCL^{+}23a] Bo Liu, Rui Chen, Weiwei Lin, Wentai Wu, Jianpeng Lin, and Keqin Li. Publisher correction to: Thermal-aware virtual machine placement based on multi-objective optimization. The Journal of Supercomputing, 79(15):17756–17757, October 2023. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL https://link.springer.com/
article/10.1007/s11227-023-05345-6. See [LCL^{+}23b].

Liu:2023:TAV

[LCL^{+}23b] Bo Liu, Rui Chen, Weiwei Lin, Wentai Wu, Jianpeng Lin, and Keqin Li. Thermal-aware virtual machine placement based on multi-objective optimization. The Journal of Supercomputing, 79(11):12563–12590, July 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL https://link.springer.com/article/
10.1007/s11227-023-05136-z. See publisher correction

[LCL^{+}23a].

Lambert:2026:CVM

[LCLG26] Simon Lambert, Eddy Caron, Laurent Lefevre, and Rémi Grivel. Consolidation of virtual machines to reduce energy consumption of data centers by using ballooning, sharing and swapping mechanisms. Future Generation Computer Systems, 174(??):??, January 2026. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X25002638.

Lopez:2017:KTC

[LCMV17] Lorena Isabel Barona López, Ángel Leonardo Valdivieso Caraguay, Marco Antonio Sotelo Monge, and Luis Javier García Villalba. Key technologies in the context of future networks: Operational and management requirements. Future Internet, 9(1):1, December 22, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.com/1999-5903/9/1/1.

Li:2015:ARP

[LCT^{+}15] Zengxiang Li, Wentong Cai, Stephen John Turner, Xiaorong Li, Ta Nguyen Binh Duong, and Rick Siow Mong Goh. Adaptive resource provisioning mechanism in VEEs for improving performance of HLA-based simulations. ACM Transactions on Modeling and Computer Simulation, 26(1):1:1–1:??, December 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lagar-Cavilla:2011:SVM

[LCWB^{+}11] H. Andrés Lagar-Cavilla, Joseph A. Whitney, Roy Bryant, Philip Patchin, Michael Brudno, Eyal de Lara, Stephen M. Rumble, M. Satyanarayanan, and Adin Scannell. SnowFlock: Virtual machine cloning as a first-class cloud primitive. ACM Transactions on Computer Systems, 29(1):2:1–2:45, February 2011. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Liu:2019:MRV

[LCZ^{+}19] Z. Liu, Y. Cao, X. Zhang, C. Zhu, and F. Zhang. Managing recurrent virtual network updates in multi-tenant datacenters: A system perspective. IEEE Transactions on Parallel and Distributed Systems, 30(8):1816–1825, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2005:VMB

[LD05] Bin Lin and Peter A. Dinda. VSched: Mixing batch and interactive virtual machines using periodic real-time scheduling. In ACM [ACM05c], pages 8–?? ISBN 1-59593-061-2. LCCN ????

Lange:2011:SSV

[LD11] John R. Lange and Peter Dinda. SymCall: symbiotic virtualization through VMM-to-guest upcalls. ACM SIGPLAN Notices, 46(7):193–204, July 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lv:2012:VCV

[LDDT12] Hui Lv, Yaozu Dong, Jiangang Duan, and Kevin Tian. Virtualization challenges: a view from server consolidation perspective. ACM SIGPLAN Notices, 47(7):15–26, July 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Loveland:2008:LVO

[LDL^{+}08] Scott Loveland, Eli M. Dow, Frank LeFevre, Duane Beyer, and Philip F. Chan. Leveraging virtualization to optimize high-availability system configurations. IBM Systems Journal, 47(4):591–604, ???? 2008. CODEN IBMSA7. ISSN

0018-8670. URL http://www.research.ibm.com/journal/
sj/474/loveland.pdf.

Li:2014:MHD

[LDL14] Pengcheng Li, Chen Ding, and Hao Luo. Modeling heap data growth using average liveness. ACM SIGPLAN Notices, 49 (11):71–82, November 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ludwig:2018:TPC

[LDRS18] Arne Ludwig, Szymon Dudycz, Matthias Rost, and Stefan Schmid. Transiently policy-compliant network updates. IEEE/ACM Transactions on Networking, 26(6):2569–2582, December 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lee:1986:DSE

[Lee86] Joon Sup Lee. Dartmouth-Smalltalk: An Exercise in Implementation. Technical Report PCS-TR86-108, Dartmouth College, Computer Science, Hanover, NH, USA, January

Lee:2016:ACS

[Lee16] Byeongcheol Lee. Adaptive correction of sampling bias in dynamic call graphs. ACM Transactions on Architecture and Code Optimization, 12(4):45:1–45:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lesser:1974:DEP

[Les74] Victor R. Lesser. The design of an emulator for a parallel machine language. ACM SIGPLAN Notices, 9(8):23–36, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lindemann:2019:DAC

[LF19] Jens Lindemann and Mathias Fischer. On the detection of applications in co-resident virtual machines via a memory deduplication side-channel. ACM SIGAPP Applied Computing Review, 18(4):31–46, January 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3307624.3307628.

Lopez:1994:ICI

[LFBB94] Gus Lopez, Bjorn Freeman-Benson, and Alan Borning. Implementing constraint imperative programming languages: The Kaleidoscope ’93 virtual machine. ACM SIGPLAN Notices, 29(10):259–271, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2019:ELV

[LFHQ19] Chunguang Li, Dan Feng, Yu Hua, and Leihua Qin. Efficient live virtual machine migration for memory writeintensive workloads. Future Generation Computer Systems, 95(??):126–139, June 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0167739X18320491.

Lu:2023:EET

[LFHS23] Baoshan Lu, Junli Fang, Xuemin Hong, and Jianghong Shi. Energy-efficient task scheduling for mobile edge computing with virtual machine I/O interference. Future Generation Computer Systems, 148(??):538–549, November 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X23002431.

Loyot:1993:VVM

[LG93] E. C. Loyot, Jr. and A. S. Grimshaw. VMPP: a virtual machine for parallel processing. In IEEE [IEE93a], pages 735– 740. ISBN 0-8186-3330-1. LCCN QA76.58 .I54 1993. IEEE catalog number 93TH0493-7.

Li:2018:HVM

[LGJ^{+}18] Xiang Li, Peter Garraghan, Xiaohong Jiang, Zhaohui Wu, and Jie Xu. Holistic virtual machine scheduling in cloud datacenters towards minimizing total energy. IEEE Transactions on Parallel and Distributed Systems, 29(6):1317–1331, June 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-

2183 (electronic). URL https://www.computer.org/csdl/
trans/td/2018/06/07888576-abs.html.

Lama:2016:APP

[LGJZ16] Palden Lama, Yanfei Guo, Changjun Jiang, and Xiaobo Zhou. Autonomic performance and power control for colocated Web applications in virtualized datacenters. IEEE Transactions on Parallel and Distributed Systems, 27(5): 1289–1302, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.computer.

org/csdl/trans/td/2016/05/07152934-abs.html.

Li:2014:SCA

[LGR14] Peng Li, Debin Gao, and Michael K. Reiter. StopWatch: a cloud architecture for timing channel mitigation. ACM Transactions on Information and System Security, 17(2):8:1– 8:??, November 2014. CODEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Lin:2023:SBO

[LGS^{+}23] Xu Lin, Deke Guo, Yulong Shen, Guoming Tang, Bangbang Ren, and Ming Xu. SFT-Box: an online approach for minimizing the embedding cost of multiple hybrid SFCs. IEEE/ACM Transactions on Networking, 31(4):1463–1477, 2023. CODEN IEANEP. ISSN 1063-6692 (print), 1558-

2566 (electronic). URL https://dl.acm.org/doi/10.1109/
TNET.2022.3221868.

Li:2023:AAP

[LGXC23] Jinxi Li, Deke Guo, Junjie Xie, and Sheng Chen. Availabilityaware provision of service function chains in mobile edge computing. ACM Transactions on Sensor Networks, 19(3):57:1– 57:??, August 2023. CODEN ???? ISSN 1550-4859 (print),

1550-4867 (electronic). URL https://dl.acm.org/doi/10.
1145/3565483.

Liang:2019:UBO

[LGZ^{+}19] Y. Liang, J. Ge, S. Zhang, J. Wu, Z. Tang, and B. Luo. A utility-based optimization framework for edge service entity caching. IEEE Transactions on Parallel and Distributed Systems, 30(11):2384–2395, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lameed:2013:MAS

[LH13] Nurudeen A. Lameed and Laurie J. Hendren. A modular approach to on-stack replacement in LLVM. ACM SIGPLAN Notices, 48(7):143–154, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Liu:2015:VCL

[LH15] Haikun Liu and Bingsheng He. VMbuddies: Coordinating live migration of multi-tier applications in cloud environments. IEEE Transactions on Parallel and Distributed Systems, 26(4):1192–1205, April 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/csdl.computer.org/csdl/trans/td/2015/04/06784491-
abs.html.

Lee:2016:HSC

[LH16] Yuan-Cheng Lee and Chih-Wen Hsueh. Hardware/software co-design of memory page translation for mobile virtualization. IEEE Transactions on Computers, 65(10):3070–3082, ???? 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2006:HPV

[LHAP06] Jiuxing Liu, Wei Huang, Bulent Abali, and Dhabaleswar K. Panda. High performance VMM-bypass I/O in virtual machines. In USENIX [USE06], page ?? ISBN 1-931971-44-7. LCCN ????

Lin:2020:LGN

[LHW^{+}20] Yilei Lin, Ting He, Shiqiang Wang, Kevin Chan, and Stephen Pasteris. Looking glass of NFV: Inferring the structure and state of NFV network from external observations. IEEE/ ACM Transactions on Networking, 28(4):1477–1490, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-

2566 (electronic). URL https://dl.acm.org/doi/10.1109/
TNET.2020.2985908.

Li:2014:LSD

[Li14] Keqin Li. Liquid: A scalable deduplication file system for virtual machine images. IEEE Transactions on Parallel and Distributed Systems, 25(5):1257–1266, May 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liang:2005:DLM

[Lia05] Jiangang Liang. Development of logical models for CNC machine tool motion control system with application to virtual machine tool design. Dissertations (Ph.D.), University of California, Davis, Davis, CA, USA, March 2005. URL http://

uclibs.org/PID/11984.

Li:2017:CSN

[LJFS17] Ning Li, Hong Jiang, Dan Feng, and Zhan Shi. Customizable SLO and its near-precise enforcement for storage bandwidth. ACM Transactions on Storage, 13(1):6:1–6:??, March 2017. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Liu:2011:LVM

[LJL^{+}11] Haikun Liu, Hai Jin, Xiaofei Liao, Chen Yu, and Cheng-Zhong Xu. Live virtual machine migration via asynchronous replication and state synchronization. IEEE Transactions on Parallel and Distributed Systems, 22(12):1986–1999, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liao:2012:TGC

[LJL12] Xiaofei Liao, Hai Jin, and Haikun Liu. Towards a green cluster through dynamic remapping of virtual machines. Future Generation Computer Systems, 28(2):469–477, February 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X11000641.

Liu:2015:HBC

[LJL^{+}15] Haikun Liu, Hai Jin, Xiaofei Liao, Wei Deng, Bingsheng He, and Cheng zhong Xu. Hotplug or ballooning: A comparative study on dynamic memory management techniques for virtual machines. IEEE Transactions on Parallel and Distributed Systems, 26(5):1350–1363, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.computer.org/csdl/trans/td/
2015/05/06807799-abs.html.

Li:2000:UCS

[LJN^{+}00] Tao Li, Lizy Kurian John, Vijaykrishnan Narayanan, Anand Sivasubramaniam, Jyotsna Sabarinathan, and Anupama Murthy. Using complete system simulation to characterize SPECjvm98 benchmarks. In ACM [ACM00], pages 22–33. ISBN 1-58113-270-0. LCCN QA76.88 .I573 2000.

URL http://www.acm.org/pubs/contents/proceedings/
supercomputing/335231.

Li:2012:SRS

[LJR12] Chunxiao Li, Niraj K. Jha, and Anand Raghunathan. Secure reconfiguration of software-defined radio. ACM Transactions on Embedded Computing Systems, 11(1):10:1–10:??, March 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liao:2015:NMA

[LJYZ15] Xiaofei Liao, Hai Jin, Shizhan Yu, and Yu Zhang. A novel memory allocation scheme for memory energy reduction in virtualization environment. Journal of Computer and System Sciences, 81(1):3–15, February 2015. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022000014001056.

Lipner:2012:LVS

[LJZ12] Steve Lipner, Trent Jaeger, and Mary Ellen Zurko. Lessons from VAX/SVS for high-assurance VM systems. IEEE Security & Privacy, 10(6):26–35, November/December 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Leivadeas:2019:VPO

[LKIL19] Aris Leivadeas, George Kesidis, Mohamed Ibnkahla, and Ioannis Lambadaris. VNF placement optimization at the edge and cloud. Future Internet, 11(3):69, March 09, 2019. CO-DEN ???? ISSN 1999-5903. URL https://www.mdpi.com/

1999-5903/11/3/69.

Leon:2019:HBW

[LKL^{+}19] R. S. Leon, M. Kiperberg, A. A. Leon Zabag, A. Resh, A. Algawi, and N. J. Zaidenberg. Hypervisor-based white listing of executables. IEEE Security & Privacy, 17(5):58–67, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Laaziz:2019:FFS

[LKR^{+}19] Lahlou Laaziz, Nadjia Kara, Rafi Rabipour, Claes Edstrom, and Yves Lemieux. FASTSCALE: a fast and scalable evolutionary algorithm for the joint placement and chaining of virtualized services. Journal of Network and Computer Applications, 148(??):??, December 15, 2019. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519302875.

Lee:2017:EBG

[LKY^{+}17] Jinho Lee, Heesu Kim, Sungjoo Yoo, Kiyoung Choi, H. Peter Hofstee, Gi-Joon Nam, Mark R. Nutter, and Damir Jamsek. ExtraV: boosting graph processing near storage with a coherent accelerator. Proceedings of the VLDB Endowment, 10 (12):1706–1717, August 2017. CODEN ???? ISSN 2150-8097.

Liu:2014:OVM

[LL14] Ming Liu and Tao Li. Optimizing virtual machine consolidation performance on NUMA server architecture for cloud workloads. ACM SIGARCH Computer Architecture News, 42 (3):325–336, June 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Li:2018:LCS

[LLF^{+}18] Huixi Li, Wenjun Li, Qilong Feng, Shigeng Zhang, Haodong Wang, and Jianxin Wang. Leveraging content similarity among VMI files to allocate virtual machines in cloud. Future Generation Computer Systems, 79 (part 2)(??):528–542, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167739X17300146.

Lee:2017:PEH

[LLLE17] Kihong Lee, DongWoo Lee, Sungkil Lee, and Young Ik Eom. Power-efficient and high-performance block I/O framework for mobile virtualization systems. The Journal of Supercomputing, 73(4):1307–1321, April 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Liu:2008:PBH

[LLS^{+}08] Tiancheng Liu, Ying Li, Andrew Schofield, Matt Hogstrom, Kewei Sun, and Ying Chen. Partition-based heap memory management in an application server. Operating Systems Review, 42(1):98, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2012:GCV

[LLS^{+}12] Guodong Li, Peng Li, Geof Sawaya, Ganesh Gopalakrishnan, Indradeep Ghosh, and Sreeranga P. Rajan. GKLEE: concolic verification and test generation for GPUs. ACM SIG-PLAN Notices, 47(8):215–224, August 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Liu:2014:MGR

[LLS14] Ning Liu, Xiaoping Li, and Weiming Shen. Multi-granularity resource virtualization and sharing strategies in cloud manufacturing. Journal of Network and Computer Applications, 46(??):72–82, November 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804514001933.

Liu:2020:LCV

[LLS^{+}20] Zhenpeng Liu, Jiahuan Lu, Nan Su, Bin Zhang, and Xiaofei Li. Location-constrained virtual machine placement (LCVP) algorithm. Scientific Programming, 2020(1):8846087:1– 8846087:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1155/2020/8846087.

Leung:1998:DGD

[LLW98] Kwong-Sak Leung, Kin-Hong Lee, and Yuk-Yin Wong. DJM: a global distributed virtual machine on the Internet. Software —Practice and Experience, 28(12):1269–1297, October 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract?ID=1763;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=1763&PLACEBO=IE.pdf.

Li:2012:CVS

[LLW^{+}12] Jianxin Li, Bo Li, Tianyu Wo, Chunming Hu, Jinpeng Huai, Lu Liu, and K. P. Lam. CyberGuarder: a virtualization security assurance architecture for green cloud computing. Future Generation Computer Systems, 28(2):379–390, February 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X1100063X.

Lin:2016:BSC

[LLW^{+}16] Po-Ching Lin, Ying-Dar Lin, Cheng-Ying Wu, Yuan-Cheng Lai, and Yi-Chih Kao. Balanced service chaining in software-defined networks with network function virtualization. Computer, 49(11):68–76, November 2016. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://www.computer.org/csdl/mags/co/
2016/11/mco2016110068-abs.html.

Li:2023:IDP

[LLWM23] Na Li, XiaoLing Liu, Yu Wang, and Musa Mojarad. Improving dynamic placement of virtual machines in cloud data centers based on open-source development model algorithm. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10723-023-09651-4.

Li:2018:OVM

[LLWW18] Huixi Li, Wenjun Li, Haodong Wang, and Jianxin Wang. An optimization of virtual machine selection and placement by using memory content similarity for server consolidation in cloud. Future Generation Computer Systems, 84(??):98– 107, July 2018. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167739X17323063.

Liang:2017:VVB

[LLX^{+}17] Hongliang Liang, Mingyu Li, Jian Xu, Wenying Hu, Xiaoxiao Pei, Xiaodong Jia, and Yan Song. vmOS: a virtualization-based, secure desktop system. Computers & Security, 65(??):329–343, March 2017. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404816301390.

Li:2018:TFV

[LLY^{+}18] Junfeng Li, Dan Li, Yirong Yu, Yukai Huang, Jing Zhu, and Jinkun Geng. Towards full virtualization of SDN infrastructure. Computer Networks (Amsterdam, Netherlands: 1999), 143(??):ii, October 9, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1389128618303852.

Li:2025:REQ

[LLY^{+}25] Zhihua Li, Zhaonan Li, Ran Yang, Junhao Qian, and Ning Yu. Resource-efficient and quality-aware virtual machine consolidation method. Journal of Grid Computing, 23(1): ??, March 2025. CODEN ???? ISSN 1570-7873 (print),

1572-9184 (electronic). URL https://link.springer.com/
article/10.1007/s10723-024-09793-z.

Liu:2018:SPM

[LLZ18] Xi Liu, Weidong Li, and Xuejie Zhang. Strategy-proof mechanism for provisioning and allocation virtual machines in heterogeneous clouds. IEEE Transactions on Parallel and Distributed Systems, 29(7):1650–1663, July 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2018/07/08241849-abs.html.

Li:2019:PSB

[LLZ^{+}19] Huixi Li, Wenjun Li, Shigeng Zhang, Haodong Wang, Yi Pan, and Jianxin Wang. Page-sharing-based virtual machine packing with multi-resource constraints to reduce network traffic in migration for clouds. Future Generation Computer Systems, 96(??):462–471, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X18302619.

Lewis:1999:EBP

[LM99] Brian T. Lewis and Bernd Mathiske. Efficient barriers for persistent object caching in a high-performance Java virtual machine. Technical report, Sun Microsystems, Palo Alto, CA, USA, December 1999. 9 pp.

Lopez:2019:TMT

[LMDP19] Martin Andreoni Lopez, Diogo M. F. Mattos, Otto Carlos M. B. Duarte, and Guy Pujolle. Toward a monitoring and threat detection system based on stream processing as a virtual network function for big data. Concurrency and Computation: Practice and Experience, 31(20):e5344:1–e5344:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lewis:2000:APH

[LMG00] Brian T. Lewis, Bernd Mathiske, and Neal M. Gafter. Architecture of the PEVM: a high-performance orthogonally persistent Java Virtual Machine. Technical Report 2000-93, Sun Microsystems, Palo Alto, CA, USA, October 2000. 19 pp.

Lewis:2001:APH

[LMG01] Brian Lewis, Bernd Mathiske, and Neal Gafter. Architecture of the PEVM: a high-performance orthogonally Persistent Java^{T M} Virtual Machine. Lecture Notes in Computer Science, 2135:18–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2135/
21350018.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2135/21350018.pdf.

Lowe:2014:MVV

[LMG^{+}14] Scott Lowe, Nick Marshall, Forbes Guthrie, Matt Liebowitz, and Josh Atwell. Mastering VMware vSphere 5.5. Wiley, New York, NY, USA, 2014. ISBN 1-118-66114-1 (paperback). xxix + 808 pp. LCCN ????

Laureano:2007:PHB

[LMJ07] M. Laureano, C. Maziero, and E. Jamhour. Protecting hostbased intrusion detectors through virtual machines. Computer Networks (Amsterdam, Netherlands: 1999), 51(5): 1275–1283, April 11, 2007. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lago:2018:EAV

[LMM18] Daniel Guimaraes Lago, Edmundo R. M. Madeira, and Deep Medhi. Energy-aware virtual machine scheduling on data centers with heterogeneous bandwidths. IEEE Transactions on Parallel and Distributed Systems, 29(1):83–98, January 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2018/01/08039210-abs.html.

Lettieri:2018:SPV

[LMR18] Giuseppe Lettieri, Vincenzo Maffione, and Luigi Rizzo. A study of I/O performance of virtual machines. The Computer Journal, 61(6):808–831, June 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/academic.oup.com/comjnl/article/61/6/808/4259797.

Laden:2012:ADF

[LMV12] Guy Laden, Roie Melamed, and Ymir Vigfusson. Adaptive and dynamic funnel replication in clouds. Operating Systems Review, 46(1):40–46, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lott:1991:DVM

[Lot91] Christopher M. Lott. Defining a virtual machine for executing MVP–L project plans. Master’s thesis, Department of Computer Science, The University of Maryland, College Park, MD 20742, August 1991.

Low:1988:SPO

[Low88] C. Low. A shared, persistent object store. In S. Gjessing and K. Nygaard, editors, ECOOP ’88 European Conference on Object-Oriented Programming, pages 390–410. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., August 1988.

Lowe:2008:VID

[Low08] William J. Lowe. VMware infrastructure 3 for dummies. Wiley, New York, NY, USA, 2008. ISBN 0-470-27793-9. xviii + 335 pp. LCCN ????

Lowe:2009:MVV

[Low09] Scott Lowe. Mastering VMware vSphere 4. Sybex serious skills. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2009. ISBN 0-470-48138-2. xxi + 673 pp. LCCN QA76.9.V5 L67 2009.

Lowe:2011:MVV

[Low11] Scott Lowe. Mastering VMware vSphere5. Serious skills. Wiley, New York, NY, USA, 2011. ISBN 0-470-89080-0 (paperback), 1-118-18010-0 (ebk.), 1-118-18011-9 (ebk.), 1-118-18012-7 (ebk.). xxv + 742 pp. LCCN QA76.9.V5 L685 2011.

Loyot:1992:VVM

[Loy92] Edmond C. Loyot, Jr. VMPP: a virtual machine for parallel processing. Technical Report CS-92-30, University of Virginia, September 29, 1992. URL ftp://ftp.cs.virginia.

edu/pub/techreports/CS-92-30.ps.Z.

Lombardi:2011:SVC

[LP11] Flavio Lombardi and Roberto Di Pietro. Secure virtualization for cloud computing. Journal of Network and Computer Applications, 34(4):1113–1122, July 2011. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001062.

Lee:2014:GSB

[LP14] Hwamin Lee and Doosoon Park. A Grid service-based virtual screening system. The Computer Journal, 57(2):

302–307, February 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/57/2/302.full.pdf+html.

Lopez-Pires:2017:MOV

[LPB17] Fabio López-Pires and Benjamín Barán. Many-objective virtual machine placement. Journal of Grid Computing, 15(2): 161–176, June 2017. CODEN ???? ISSN 1570-7873 (print),

1572-9184 (electronic). URL https://link.springer.com/
article/10.1007/s10723-017-9399-x.

Lopez-Pires:2018:VMP

[LPBB^{+}18] Fabio López-Pires, Benjamín Barán, Leonardo Benítez, Saúl Zalimben, and Augusto Amarilla. Virtual machine placement for elastic infrastructures in overbooked cloud computing datacenters under uncertainty. Future Generation Computer Systems, 79 (part 3)(??):830–848, February 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X17303126.

Lange:2011:MOV

[LPD^{+}11] John R. Lange, Kevin Pedretti, Peter Dinda, Patrick G. Bridges, Chang Bae, Philip Soltero, and Alexander Merritt. Minimal-overhead virtualization of a large scale supercomputer. ACM SIGPLAN Notices, 46(7):169–180, July 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lebre:2019:PNV

[LPSS19] Adrien Lebre, Jonathan Pastor, Anthony Simonet, and Mario Sudholt. Putting the next 500 VM placement algorithms to the acid test: The infrastructure provider viewpoint. IEEE Transactions on Parallel and Distributed Systems, 30(1):204–217, January 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:

//www.computer.org/csdl/trans/td/2019/01/08409978-
abs.html.

Liu:2022:AFB

[LPZ^{+}22] Danxu Liu, Junjie Peng, Xin Zhang, Yongkang You, and Bo Ning. Application features-based virtual machine deployment strategy in cloud environment. Concurrency and Computation: Practice and Experience, 34(5):e6691:1–e6691:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2018:IPN

[LQD^{+}18] Gangyi Luo, Zhuzhong Qian, Mianxiong Dong, Kaoru Ota, and Sanglu Lu. Improving performance by network-aware virtual machine clustering and consolidation. The Journal of Supercomputing, 74(11):5846–5864, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lin:2012:OVM

[LQW^{+}12] Qian Lin, Zhengwei Qi, Jiewei Wu, Yaozu Dong, and Haibing Guan. Optimizing virtual machines using hybrid virtualization. The Journal of Systems and Software, 85(11):2593–2603, November 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121212001677.

Lucchetti:2005:EDR

[LRC05] Dominic Lucchetti, Steven K. Reinhardt, and Peter M. Chen. ExtraVirt: detecting and recovering from transient processor faults. In Vrable [Vra05], pages 1–8. ISBN 1-59593-079-5. LCCN QA76.6 .S9196 2005; QA76.6; QA76.6 .S9196 2005eb; Internet.

Linguaglossa:2019:HSD

[LRP^{+}19] Leonardo Linguaglossa, Dario Rossi, Salvatore Pontarelli, Dave Barach, Damjan Marjon, and Pierre Pfister. Highspeed data plane and network functions virtualization by vectorizing packet processing. Computer Networks (Amsterdam, Netherlands: 1999), 149(??):187–199, February 11, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128618312957.

Lu:2016:VCV

[LRZ16] Yaojie Lu, Seyedamin Rooholamin, and Sotirios G. Ziavras. Vector coprocessor virtualization for simultaneous multithreading. ACM Transactions on Embedded Computing Systems, 15(3):57:1–57:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ludwig:2015:DCM

[LS15] Arne Ludwig and Stefan Schmid. Distributed cloud market: Who benefits from specification flexibilities? ACM SIGMET-RICS Performance Evaluation Review, 43(3):38–41, December 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lei:2017:NHC

[LSC^{+}17] Zhou Lei, Exiong Sun, Shengbo Chen, Jiang Wu, and Wenfeng Shen. A novel hybrid-copy algorithm for live migration of virtual machine. Future Internet, 9(3):37, July 18, 2017. CODEN ???? ISSN 1999-5903. URL https://www.mdpi.

com/1999-5903/9/3/37.

Liu:2023:VSL

[LSPP^{+}23] Zongyuan Liu, Sergei Stepanenko, Jean Pichon-Pharabod, Amin Timany, Aslan Askarov, and Lars Birkedal. VMSL: a separation logic for mechanised robust safety of virtual machines communicating above FF-A. Proceedings of the ACM on Programming Languages (PACMPL), 7(PLDI): 165:1–165:??, June 2023. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3591279.

Lowell:2004:DVM

[LSS04] David E. Lowell, Yasushi Saito, and Eileen J. Samberg. Devirtualizable virtual machines enabling general, single-node, online maintenance. ACM SIGARCH Computer Architecture News, 32(5):211–223, December 2004. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Lauwaerts:2024:WEW

[LSS24] Tom Lauwaerts, Robbert Gurdeep Singh, and Christophe Scholliers. WARDuino: an embedded WebAssembly virtual machine. Journal of Computer Languages, 79(??):1–35, June 2024. CODEN ???? ISSN 2665-9182 (print), 2590-

1184 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S259011842400011X.

Li:2022:SIM

[LSSC22] Zhuozhao Li, Tanmoy Sen, Haiying Shen, and Mooi Choo Chuah. A study on the impact of memory DoS attacks on cloud applications and exploring real-time detection schemes.

IEEE/ACM Transactions on Networking, 30(4):1644–1658, 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-

2566 (electronic). URL https://dl.acm.org/doi/10.1109/
TNET.2022.3144895.

Liu:2024:HAK

[LSX^{+}24] Jianzhong Liu, Yuheng Shen, Yiru Xu, Hao Sun, and Yu Jiang. Horus: Accelerating kernel fuzzing through efficient Host–VM memory access procedures. ACM Transactions on Software Engineering and Methodology, 33(1):11:1– 11:??, January 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https://dl.acm.org/

doi/10.1145/3611665.

Li:2012:VMP

[LTE12] Wubin Li, Johan Tordsson, and Erik Elmroth. Virtual machine placement for predictable and time-constrained peak loads. Lecture Notes in Computer Science, 7150:120–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-28675-9_9/.

Luckow:2017:HTP

[LTK17] Kasper Søe Luckow, Bent Thomsen, and Stephan Erbs Korsholm. HVM_{T P}: a time predictable and portable Java Virtual Machine for hard real-time embedded systems. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:1992:IES

[LTT92] Jim-Min Lin, Shang Rong Tsai, and Li-Ming Tseng. Integrating existing software packages using the virtual machine technique. The Journal of Systems and Software, 18(3):207– 218, July 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2014:PAC

[LTZ^{+}14] Xiaodong Liu, Weiqin Tong, Xiaoli Zhi, Fu ZhiRen, and Liao WenZhao. Performance analysis of cloud computing services considering resources sharing among virtual machines. The Journal of Supercomputing, 69(1):357–374, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL http://link.springer.com/article/
10.1007/s11227-014-1156-3.

LeVasseur:2004:SAR

[LU04] Joshua LeVasseur and Volkmar Uhlig. A sledgehammer approach to reuse of legacy device drivers. In ACM [ACM04a], page ?? ISBN ???? LCCN ????

Lucent:1997:LPL

[Luc97] Lucent Technologies Inc. The Limbo Programming Language,

1997. URL http://inferno.lucent.com/inferno/.

LeVasseur:2005:PVU

[LUL^{+}05] Joshua LeVasseur, Volkmar Uhlig, Ben Leslie, Matthew Chapman, and Gernot Heiser. Pre-virtualization: uniting two worlds. In Vrable [Vra05], pages 1–2. ISBN 1-59593-079-5. LCCN QA76.6 .S9196 2005; QA76.6; QA76.6 .S9196 2005eb; Internet.

Liang:1999:CPS

[LV99] Sheng Liang and Deepa Viswanathan. Comprehensive profiling support in the Java Virtual Machine. In USENIX

[USE99], page ??  ISBN ????  LCCN QA76.64

.U84 1999. URL http://www.usenix.org/publications/
library/proceedings/coots99/liang.html.

Li:2016:SSO

[LVM16] Xi Li, Anthony Ventresque, and John Murphy. SOC: Satisfaction-oriented virtual machine consolidation in enterprise data centers. International Journal of Parallel Programming, 44(1):130–150, February 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10766-014-0333-6.

Lauer:1973:RVM

[LW73] H. C. Lauer and D. Wyeth. A recursive virtual machine architecture. In ????, editor, Proceedings ACM SIGARCH-SIGOPS Workshop on Virtual Computer Systems, Cambridge, MA, 1973, page ?? ACM Press, New York, NY 10036, USA, 1973.

Le:2011:EMO

[LW11] Duy Le and Haining Wang. An effective memory optimization for virtual machine-based systems. IEEE Transactions on Parallel and Distributed Systems, 22(10):1705–1713, November 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2012:PBA

[LW12] Zhanghui Liu and Xiaoli Wang. A PSO-based algorithm for load balancing in virtual machines of cloud computing environment. Lecture Notes in Computer Science, 7331:142– 147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-30976-2_17/.

Luo:2020:OAV

[LW20] Ziyue Luo and Chuan Wu. An online algorithm for VNF service chain scaling in datacenters. IEEE/ACM Transactions on Networking, 28(3):1061–1073, June 2020. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.acm.org/doi/10.1109/TNET.
2020.2979263.

Lyons:2013:SFF

[LWB13] Michael Lyons, Gu-Yeon Wei, and David Brooks. Shrink-Fit: A framework for flexible accelerator sizing. IEEE Computer Architecture Letters, 12(1):17–20, January/June 2013. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lin:2015:SGU

[LWB^{+}15] Yi Lin, Kunshan Wang, Stephen M. Blackburn, Antony L. Hosking, and Michael Norrish. Stop and go: understanding yieldpoint behavior. ACM SIGPLAN Notices, 50(11):70–80, November 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2017:AET

[LWC^{+}17] Hong-Wei Li, Yu-Sung Wu, Yi-Yung Chen, Chieh-Min Wang, and Yen-Nun Huang. Application execution time prediction for effective CPU provisioning in virtualization environment. IEEE Transactions on Parallel and Distributed Systems, 28(11):3074–3088, November 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2017/11/07933268-abs.html.

Liu:2022:EAV

[LWCZ22] Xialin Liu, Junsheng Wu, Lijun Chen, and Lili Zhang. Energy-aware virtual machine consolidation based on evolutionary game theory. Concurrency and Computation: Practice and Experience, 34(10):e6830:1–e6830:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2016:JOQ

[LWL16] Shih-Chun Lin, Pu Wang, and Min Luo. Jointly optimized QoS-aware virtualization and routing in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 96(??):69–78, February 26, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S138912861500256X.

Liu:2010:VMF

[LWLL10] Qian Liu, Chuliang Weng, Minglu Li, and Yuan Luo. An in-VM measuring framework for increasing virtual machine security in clouds. IEEE Security & Privacy, 8(6):56–62, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Li:2016:VMT

[LWLL16] Hongxing Li, Chuan Wu, Zongpeng Li, and Francis C. M. Lau. Virtual machine trading in a federation of clouds: individual profit and social welfare maximization. IEEE/ACM Transactions on Networking, 24(3):1827–1840, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2014:VSK

[LWM14] Ye Li, Richard West, and Eric Missimer. A virtualized separation kernel for mixed criticality systems. ACM SIGPLAN Notices, 49(7):201–212, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lin:2016:HTS

[LWW16] Weiwei Lin, Wentai Wu, and James Z. Wang. A heuristic task scheduling algorithm for heterogeneous virtual clusters. Scientific Programming, 2016(??):7040276:1–7040276:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X

(electronic). URL https://www.hindawi.com/journals/
sp/2016/7040276/.

Liu:2018:TBG

[LWZ^{+}18] Jialei Liu, Shangguang Wang, Ao Zhou, Xiang Xu, Sathish A. P. Kumar, and Fangchun Yang. Towards bandwidth guaranteed virtual cluster reallocation in the cloud. The Computer Journal, 61(9):1284–1295, September 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://academic.oup.com/comjnl/article/
61/9/1284/4779880.

Li:2022:VAT

[LXL^{+}22] Chong Li, Sisu Xi, Chenyang Lu, Roch Guérin, and Christopher D. Gill. Virtualization-aware traffic control for soft real-time network traffic on Xen. IEEE/ACM Transactions on Networking, 30(1):257–270, February 2022. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.acm.org/doi/10.1109/TNET.
2021.3114055.

Luo:2016:OMM

[LXM^{+}16] Qiuming Luo, Feng Xiao, Zhong Ming, Hao Li, Jianyong Chen, and Jianhua Zhang. Optimizing the memory management of a virtual machine monitor on a NUMA system. Computer, 49(6):66–74, June 2016. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.org/csdl/mags/co/
2016/06/mco2016060066-abs.html.

Liu:2024:DNO

[LXQ24] Lisi Liu, Zijie Xu, and Xiaobin Qu. Development of a novel open control system implementation method under industrial IoT. Future Internet, 16(8):??, August 14, 2024. CODEN

???? ISSN 1999-5903. URL https://www.mdpi.com/1999-
5903/16/8/293.

Li:2019:NIM

[LXRS19] Shupan Li, Limin Xiao, Li Ruan, and Shubin Su. A novel integrity measurement method based on copy-on-write for region in virtual machine. Future Generation Computer Systems, 97(??):714–726, August 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X18317618.

Lin:2023:PIV

[LXW^{+}23] Weiwei Lin, Chennian Xiong, Wentai Wu, Fang Shi, Keqin Li, and Minxian Xu. Performance interference of virtual machines: a survey. ACM Computing Surveys, 55(12):254:1– 254:??, December 2023. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.acm.org/

doi/10.1145/3573009.

Liu:2021:ISD

[LXZ^{+}21] Jianchun Liu, Hongli Xu, Gongming Zhao, Chen Qian, Xingpeng Fan, Xuwei Yang, and He Huang. Incremental server deployment for software-defined NFV-enabled networks. IEEE/ ACM Transactions on Networking, 29(1):248–261, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-

2566 (electronic). URL https://dl.acm.org/doi/10.1109/
TNET.2020.3030298.

Lindholm:1997:IJV

[LY97a] Tim Lindholm and Frank Yellin. Inside the Java Virtual Machine. UNIX Review, 15(1):31, 32, 34–36, 38, 39, January 1997. CODEN UNRED5. ISSN 0742-3136.

Lindholm:1997:JVM

[LY97b] Tim Lindholm and Frank Yellin. The Java Virtual Machine Specification. The Java Series. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-63452-X. xvi + 475 pp. LCCN QA76.73.J38L56 1997. US$36.53. URL http://www.aw.com/

cp/javaseries.html.

Lindholm:1999:JVM

[LY99] Tim Lindholm and Frank Yellin. The Java Virtual Machine Specification. Addison-Wesley, Reading, MA, USA, second edition, 1999. ISBN 0-201-43294-3. xv + 473 pp. LCCN QA76.73.J38L56 1999. US$42.95.

Lindholm:19xx:JVMa

[LYxxa] Tim Lindholm and Frank Yellin. The Java Virtual Machine. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN ???? URL http://www.gotop.com. tw. Chinese translation by Thi Shiang Workshop.

Lindholm:19xx:JVMb

[LYxxb] Tim Lindholm and Frank Yellin. The Java Virtual Machine. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? LCCN ???? ???? Chinese translation by Thi Shiang Workshop.

Luo:2023:KSA

[LY23] Jiang-Yao Luo and Jian-Hua Yuan. A kernel search algorithm for virtual machine consolidation problem in cloud computing. The Journal of Supercomputing, 79(17):19277–19296, November 2023. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL https://link.springer.com/
article/10.1007/s11227-023-05406-w.

Lindholm:2013:JVMa

[LYBB13a] Tim Lindholm, Frank Yellin, Gilad Bracha, and Alex Buckley. The Java Virtual Machine Specification. The Java series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2013. ISBN 0-13-326044-5 (paperback). ???? pp. LCCN ????

Lindholm:2013:JVMb

[LYBB13b] Tim Lindholm, Frank Yellin, Gilad Bracha, and Alex Buckley. The Java Virtual Machine Specification. Addison-Wes-ley, Addison-Wesley, Java SE 7 edition, 2013. ISBN 0-13-326049-6, 0-13-326044-5. xvii + 587 (est.) pp. LCCN QA76.73.J38 L56 1999. URL http://proquest.tech.

safaribooksonline.de/9780133260496.

Lindholm:2014:JVM

[LYBB14] Tim Lindholm, Frank Yellin, Gilad Bracha, and Alex Buckley. The Java Virtual Machine Specification: Java SE 8 edition. Addison-Wesley, Addison-Wesley, 2014. ISBN 0-13-390590-X (paperback), 0-13-392274-X (e-book). xvi + 584 pp. LCCN QA76.73.J38 L56 2014.

Lu:2020:GQO

[LYGG20] Q. Lu, J. Yao, H. Guan, and P. Gao. gQoS: a QoS-oriented GPU virtualization with adaptive capacity sharing. IEEE Transactions on Parallel and Distributed Systems, 31(4):843– 855, April 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2021:TOP

[LYL21] I-Chieh Lin, Yu-Hsuan Yeh, and Kate Ching-Ju Lin. Toward optimal partial parallelization for service function chaining. IEEE/ACM Transactions on Networking, 29(5):2033–2044, October 2021. CODEN IEANEP. ISSN 1063-6692 (print),

1558-2566 (electronic). URL https://dl.acm.org/doi/10.
1109/TNET.2021.3075709.

Liu:2018:CAL

[LYS^{+}18] Jiayi Liu, Qinghai Yang, Gwendal Simon, Jiayi Liu, Gwendal Simon, and Qinghai Yang. Congestion avoidance and load balancing in content placement and request redirection for mobile CDN. IEEE/ACM Transactions on Networking, 26 (2):851–863, April 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Li:2020:EEQ

[LYY^{+}20] Zhihua Li, Xinrong Yu, Lei Yu, Shujie Guo, and Victor Chang. Energy-efficient and quality-aware VM consolidation method. Future Generation Computer Systems, 102(??):789–809, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0167739X18324713.

Li:2024:MSM

[LYY^{+}24] Di Li, Zhibang Yang, Siyang Yu, Mingxing Duan, and Shenghong Yang. A micro-segmentation method based on VLAN-VxLAN mapping technology. Future Internet, 16(9): 320, September 4, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/1999-5903/16/9/320.

Li:2017:BNB

[LYYY17] Zhihua Li, Chengyu Yan, Xinrong Yu, and Ning Yu. Bayesian network-based virtual machines consolidation method. Future Generation Computer Systems, 69(??):75–87, April 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X16307415.

Li:2018:EAM

[LYYY18] Zhihua Li, Chengyu Yan, Lei Yu, and Xinrong Yu. Energyaware and multi-resource overload probability constraintbased virtual machine dynamic consolidation method. Future Generation Computer Systems, 80(??):139–156, March 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X16307476.

Lama:2015:CPP

[LZ15] P. Lama and Xiaobo Zhou. Coordinated power and performance guarantee with fuzzy MIMO control in virtualized server clusters. IEEE Transactions on Computers, 64(1): 97–111, January 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2016:EEM

[LZC^{+}16] Hongjian Li, Guofeng Zhu, Chengyuan Cui, Hong Tang, Yusheng Dou, and Chen He. Energy-efficient migration and consolidation algorithm of virtual machines in data centers for cloud computing. Computing: Archiv für Informatik und Numerik, 98(3):303–317, March 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Li:2015:ITA

[LZL^{+}15] Jianxin Li, Jieyu Zhao, Yi Li, Lei Cui, Bo Li, Lu Liu, and John Panneerselvam. iMIG: Toward an adaptive live migration method for KVM virtual machines. The Computer Journal, 58(6):1227–1242, June 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/58/6/1227.

Li:2020:MOO

[LZLY20] Rui Li, Qinghua Zheng, Xiuqi Li, and Zheng Yan. Multiobjective optimization for rebalancing virtual machine placement. Future Generation Computer Systems, 105(??):824– 842, April 2020. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167739X1731840X.

Lu:2020:GEV

[LZM^{+}20] Q. Lu, X. Zheng, J. Ma, Y. Dong, Z. Qi, J. Yao, B. He, and H. Guan. gMig: Efficient vGPU live migration with overlapped software-based dirty page verification. IEEE Transactions on Parallel and Distributed Systems, 31(5):1209–1222, May 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2015:GHB

[LZW^{+}15] Dan Li, Jing Zhu, Jianping Wu, Junjie Guan, and Ying Zhang. Guaranteeing heterogeneous bandwidth demand in multitenant data center networks. IEEE/ACM Transactions on Networking, 23(5):1648–1660, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lu:2017:FPL

[LZW^{+}17] Kai Lu, Wenzhe Zhang, Xiaoping Wang, Mikel Luján, and Andy Nisbet. Flexible page-level memory access monitoring based on virtualization hardware. ACM SIGPLAN Notices, 52(7):201–213, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2013:RVS

[LZWC13] Wenfeng Li, Ye Zhong, Xun Wang, and Yulian Cao. Resource virtualization and service selection in cloud logistics. Journal of Network and Computer Applications, 36(6):1696–1704, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print),

1095-8592 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S108480451300057X.

Li:2015:VMP

[LZWD15] Kangkang Li, Huanyang Zheng, Jie Wu, and Xiaojiang Du. Virtual machine placement in cloud systems through migration process. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(5):393–410, 2015. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Li:2024:BLI

[LZY^{+}24] Huiba Li, Zhihao Zhang, Yifan Yuan, Rui Du, Kai Ma, Lanzheng Liu, Yiming Zhang, and Windsor Hsu. Block-level image service for the cloud. ACM Transactions on Storage, 20(1):1:1–1:??, February 2024. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic). URL https://dl.acm.org/

doi/10.1145/3620672.

Liu:2025:ECN

[LZZ^{+}25] Tianyi Liu, Zhenfei Zhang, Yuncong Zhang, Wenqing Hu, and Ye Zhang. Ceno: Non-uniform, segment and parallel zero-knowledge virtual machine. Journal of Cryptology: the journal of the International Association for Cryptologic Research, 38(2):??, April 2025. CODEN JOCREQ. ISSN 0933-2790 (print), 1432-1378 (electronic). URL https://link.

springer.com/article/10.1007/s00145-024-09533-2.

Min:2006:FHP

[M^{+}06] Geyong Min et al., editors. Frontiers of high performance computing and networking: ISPA 2006 workshops: ISPA 2006 international workshops, FHPCN, XHPC, S-GRACE, GridGIS, HPC-GTP, PDCE, ParDMCom, WOMP, ISDF, and UPWN, Sorrento, Italy, December 4-7, 2006: proceedings, volume 4331 of Lecture notes in computer science, 0302-9743. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-49860-5 (paperback). LCCN QA76.58 .I88 2006a; QA76.88 .F76

2006. URL http://www.loc.gov/catdir/enhancements/
fy0825/2006937143-d.html.

McDougall:2010:VPP

[MA10] Richard McDougall and Jennifer Anderson. Virtualization performance: perspectives and challenges ahead. Operating Systems Review, 44(4):40–56, December 2010. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Modi:2017:VLS

[MA17] Chirag N. Modi and Kamatchi Acha. Virtualization layer security challenges and intrusion detection/prevention systems in cloud computing: a comprehensive review. The Journal of Supercomputing, 73(3):1192–1234, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mohiuddin:2019:WAV

[MA19] Irfan Mohiuddin and Ahmad Almogren. Workload aware VM consolidation method in edge/cloud computing for IoT applications. Journal of Parallel and Distributed Computing, 123(??):204–214, January 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518306762.

Mirobi:2021:DDA

[MA21] G. Justy Mirobi and L. Arockiam. DAVmS: Distance Aware Virtual Machine Scheduling approach for reducing the response time in cloud computing. The Journal of Supercomputing, 77(7):6664–6675, July 2021. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-020-03563-w.

Mirobi:2024:RND

[MA24] G. Justy Mirobi and L. Arockiam. Retraction note: DAVmS: Distance Aware Virtual Machine Scheduling approach for reducing the response time in cloud computing. The Journal of Supercomputing, 80(13):19945, September 2024. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-024-05970-9. See [MA21].

MacKinnon:1979:CVM

[Mac79] Richard A. MacKinnon. The changing virtual machine environment: Interfaces to real hardware, virtual hardware, and other virtual machines. IBM Systems Journal, 18(1):18–46, 1979. CODEN IBMSA7. ISSN 0018-8670.

Madnick:1969:TSS

[Mad69] S. E. Madnick. Time-sharing systems: Virtual machine concept vs. conventional approach. Modern Data, 2(3):34–36, March 1969.

Muller:2007:VMS

[MAK07] René Müller, Gustavo Alonso, and Donald Kossmann. A virtual machine for sensor networks. Operating Systems Review, 41(3):145–158, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Marotta:2018:JPE

[MAK18] Antonio Marotta, Stefano Avallone, and Andreas Kassler. A joint power efficient server and network consolidation approach for virtualized data centers. Computer Networks (Amsterdam, Netherlands: 1999), 130(??):65–80, January 15, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128617304012.

Mallach:1972:ES

[Mal72] E. G. Mallach. Emulation — a survey. Honeywell Computer Journal, 6(4):??, ???? 1972. CODEN HNCJA3. ISSN 0046-7847.

Mallach:1973:RBE

[Mal73] E. G. Mallach. On the relationship between emulators and virtual machines. In ????, editor, Proceedings ACM SIGOPS-SIGARCH Workshop on Virtual Computer Systems, Cambridge, MA, 1973, page ?? ACM Press, New York, NY 10036, USA, 1973.

Mann:2015:AVM

[Man15a] Zoltán Ádám Mann. Allocation of virtual machines in cloud data centers — a survey of problem models and optimization algorithms. ACM Computing Surveys, 48(1):11:1–11:??, September 2015. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Mann:2015:RRE

[Man15b] Zoltán Ádám Mann. Rigorous results on the effectiveness of some heuristics for the consolidation of virtual machines in a cloud data center. Future Generation Computer Systems, 51(??):1–6, October 2015. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X15000965.

Mann:2016:MAV

[Man16] Zoltán Ádám Mann. Multicore-aware virtual machine placement in cloud data centers. IEEE Transactions on Computers, 65(11):3357–3369, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mann:2018:CSI

[Man18] Zoltán Ádám Mann. Cloud simulators in the implementation and evaluation of virtual machine placement algorithms. Software—Practice and Experience, 48(7):1368–1389, July 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-

024X (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/spe.2579.

March:1973:DIV

[Mar73] J. H. March. The design and implementation of a virtual machine operating system using a virtual access method. In ????, editor, Proceedings ACM SIGARCH-SIGOPS Workshop on Virtual Computer Systems, Cambridge, MA, 1973, page ?? ACM Press, New York, NY 10036, USA, 1973.

Martin:1981:RFS

[Mar81] Thea Martin, editor. 1981 Rochester FORTH Standards Conference, May 12–15, 1981. The Institute for Applied Forth Research, Inc., 70 Elmwood Avenue, Rochester, New York 14611, USA, 1981. ISBN ???? LCCN QA76.73.F24 R63 1981.

Marcy:2008:DRP

[Mar08] Hervé Marcy. Disaster-Recovery Programm mit Xen auf Linux Servern [Elektronische Ressource]: Design und Implementierung eines DRP für Xen-virtualisierte Linux-Server im Falle eines kompletten Ausfalls eines Rechenzentrums. Verlag Dr. Müller, Saarbrücken, Germany, 2008. ISBN 3-639-09099-3. ???? pp. LCCN ????

Mattsson:2009:RSV

[Mat09] Ulf Mattsson. Real security for virtual machines. Network Security, 2009(4):15–17, April 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1353485809700418.

Matthews:2010:WPO

[Mat10] Jeanna Neefe Matthews. Workshop proceedings and other publications in operating system review. Operating Systems Review, 44(1):1, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Millet:1998:PGT

[MB98] Laurent Millet and Ted Baker. Porting the GNAT tasking runtime system to the Java Virtual Machine. Lecture Notes in Computer Science, 1411:19–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/1411/14110019.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
1411/14110019.pdf.

Malik:2015:LMV

[MB15] Vikas Malik and C. R. Barde. Live migration of virtual machines in cloud environment using prediction of CPU usage. International Journal of Computer Applications, 117(??):1–5, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume117/number23/20691-3604/.

Mesnard:2020:RWP

[MB20] O. Mesnard and L. A. Barba. Reproducible workflow on a public cloud for computational fluid dynamics. Computing in Science and Engineering, 22(1):102–116, January/February 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mansouri:2021:REC

[MB21] Yaser Mansouri and M. Ali Babar. A review of edge computing: Features and resource virtualization. Journal of Parallel and Distributed Computing, 150(??):155–183, April 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304317.

Mayer:2012:URM

[MBA^{+}12] C. Mayer, F. Baitinger, S. Amann, G. McAfee, and A. Nunez Mencias. Unified Resource Manager virtualization management. IBM Journal of Research and Development, 56(1–2): 18:1–18:9, ???? 2012. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Mittal:2013:EVE

[MBBS13] Aashish Mittal, Dushyant Bansal, Sorav Bansal, and Varun Sethi. Efficient virtualization on embedded Power Architecture(R) platforms. ACM SIGPLAN Notices, 48(4):445–458, April 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muller:1992:ASP

[MBK^{+}92] Urs A. Muller, Bernhard Baumle, Peter Kohler, Anton Gunzinger, and Walter Guggenbuhl. Achieving supercomputer performance for neural net simulation with an array of digital signal processors. IEEE Micro, 12(5):55–65, September/ October 1, 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Malhotra:2024:SLR

[MBK24] Ruchika Malhotra, Anjali Bansal, and Marouane Kessentini. A systematic literature review on maintenance of software containers. ACM Computing Surveys, 56(8):193:1–193:??, August 2024. CODEN CMSVAN. ISSN 0360-0300 (print),

1557-7341 (electronic). URL https://dl.acm.org/doi/10.
1145/3645092.

Marshall:2009:VEE

[MBM09] David (David W.) Marshall, Stephen S. Beaver, and Jason W. McCarty. VMware ESX essentials in the virtual data center. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-7027-4. xiv + 237 pp. LCCN QA76.76.O63 M3647 2009.

McDonald:1986:TND

[MBWW86] D. McDonald, J. Brooks, B. Woolf, and P. Werner. Transition networks for discourse management. Technical Report UM-CS-1986-034, University of Massachusetts, Amherst, Computer Science, December 31, 1986.

McHugh:1993:ILC

[MC93] C. (Colm) McHugh and V. (Vinny) Cahill. Interfacing a language to the Commandos [i.e., Comandos] virtual machine. External technical reports, Department of Computer Science, Trinity College, Dublin, Ireland, 1993. 23 pp.

Miller:1998:VMB

[MC98] Gregory R. Miller and Stephen C. Cooper. Visual mechanics: beams and stress states. PWS Publishing Company, Boston, MA, USA, 1998. ISBN 0-534-95587-8. x + 149 pp. LCCN TA660.B4 M55 1998.

McCullough:1974:VMF

[McC74] Barbara McCullough. Virtual Machine Facility/370 Features Supplement. IBM Corporation, ????, January 1974. ???? pp. Publication GC20-1757.

McCain:2008:MVI

[McC08]  Chris McCain.  Mastering VMware Infrastructure 3.

Wiley Technology Publishing, Indianapolis, IN, USA, 2008. ISBN 0-470-18313-6 (paperback). xxi + 554 pp. LCCN QA76.76.O63 M37483 2008. URL http://

www.loc.gov/catdir/enhancements/fy0805/2007045713-
d.html;   http://www.loc.gov/catdir/enhancements/
fy0805/2007045713-t.html;   http://www.loc.gov/catdir/
enhancements/fy0829/2007045713-b.html.

Malandrino:2018:VBE

[MCC18] Francesco Malandrino, Carla-Fabiana Chiasserini, and Claudio Casetti. Virtualization-based evaluation of backhaul performance in vehicular applications. Computer Networks (Amsterdam, Netherlands: 1999), 134(??):93–104, April 7, 2018. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128618300276.

Magnusson:2002:SFS

[MCE^{+}02] Peter S. Magnusson, Magnus Christensson, Jesper Eskilson, Daniel Forsgren, Gustav Hållberg, Johan Högberg, Fredrik Larsson, Andreas Moestedt, and Bengt Werner. Simics: a full system simulation platform. Computer, 35(2):50–58, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-

0814 (electronic). URL http://dlib.computer.org/co/
books/co2002/pdf/r2050.pdf;   http://www.computer.
org/computer/co2002/r2050abs.htm.

McGrath:1972:VMC

[McG72] Mark McGrath. Virtual machine computing in an engineering environment. IBM Systems Journal, 11(2):131–149, 1972. CODEN IBMSA7. ISSN 0018-8670.

Ma:2019:ASF

[MCJ19] Yi-Wei Ma, Jiann-Liang Chen, and Jia-Yi Jhou. Adaptive service function selection for Network Function Virtualization networking. Future Generation Computer Systems, 91(??):108–123, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167739X17328406.

McKusick:2004:JFF

[McK04] Kirk McKusick. The jail facility in FreeBSD 5.2. ;login: the USENIX Association newsletter, 29(4):??, August 2004. CODEN LOGNEM. ISSN 1044-6397. URL http:

//www.usenix.org/publications/login/2004-08/pdfs/
mckusick.pdf.

McKinley:2011:HPC

[McK11] Kathryn S. McKinley. How’s the parallel computing revolution going? ACM SIGPLAN Notices, 46(8):123–124, August 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’11 Conference proceedings.

McMillan:2011:SVM

[McM11] William W. McMillan. The soul of the virtual machine. IEEE Spectrum, 48(7):44–59, July 2011. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Menon:2006:ONV

[MCZ06] Aravind Menon, Alan L. Cox, and Willy Zwaenepoel. Optimizing network virtualization in Xen. In USENIX [USE06], page ?? ISBN 1-931971-44-7. LCCN ???? URL http://

www.usenix.org/events/usenix06/tech/menon.html.

Madnick:1973:AAV

[MD73] Stuart E. Madnick and John J. Donovan. Application and analysis of the virtual machine approach to information system security and isolation. In Proceedings of the ACM Workshop on Virtual Computer Systems, pages 210–224. ACM Press, New York, NY 10036, USA, 1973.

URL http://portal.acm.org/citation.cfm?id=803961&
coll=portal&dl=ACM.

Madnick:1974:AAV

[MD74] Stuart E. Madnick and John J. Donovan. Application and analysis of the virtual machine approach to information system security and reliability. Technical Report 722–74(2), MIT Center for Information Systems Research, Cambridge, MA, USA, 1974. 25 pp.

Meyer:1997:JVM

[MD97] Jon Meyer and Troy Downing. Java Virtual Machine. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 1997. ISBN 1-56592-194-1. xxiv + 426 pp. LCCN QA76.73.J38 M49 1997.

US$32.95. URL http://www.ora.com/www/item/javavm.
html;   http://www.oreilly.com/catalog/javavm.

Meyer:19xx:JVMb

[MDxx] Jon Meyer and Troy Downing. The Java Virtual Machine. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 19xx. ISBN 4-900900-63-X. LCCN ???? 4,500 yen. Japanese translation.

Marr:2012:IUM

[MD12] Stefan Marr and Theo D’Hondt. Identifying a unifying mechanism for the implementation of concurrency abstractions on multi-language virtual machines. Lecture Notes in Computer Science, 7304:171–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/chapter/10.1007/978-
3-642-30561-0_13/.

Matthews:2008:RXH

[MDD^{+}08] Jeanna Matthews, Eli M. Dow, Todd Deshane, Wenjin Hu, Jeremy Bongio, Patrick F. Wilbur, and Brendan Johnson, editors. Running Xen: a hands-on guide to the art of virtualization. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2008. ISBN 0-13-207467-2, 0-13-234966-3 (paperback). xxxiii + 586 pp. LCCN QA76.9.V5 R86 2008. URL http://www.

loc.gov/catdir/toc/ecip087/2007052439.html.

Morris:1972:SMO

[MDFS72] D. Morris, G. D. Detlefsen, G. R. Frank, and T. J. Sweeney. The structure of the MU5 operating system. The Computer Journal, 15(2):113–116, May 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/150113.sgm.abs.html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/
113.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/114.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/
tiff/115.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/116.tif.

Migliardi:1998:DRV

[MDGS98] M. Migliardi, J. Dongarra, A. Geist, and V. Sunderam. Dynamic reconfiguration and virtual machine management in the Harness metacomputing system. Lecture Notes in Computer Science, 1505:127–134, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

netlib.org/utk/people/JackDongarra/PAPERS/harness1.
ps.

Mai:2021:EES

[MDZ^{+}21] Liuyang Mai, Yi Ding, Xiaoning Zhang, Lang Fan, Shui Yu, and Zhichao Xu. Energy efficiency with service availability guarantee for network function virtualization. Future Generation Computer Systems, 119(??):140–153, June 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X21000479.

Maxim:1987:TPA

[ME87] Bruce R. Maxim and Bruce S. Elenbogen. Teaching programming algorithms aided by computer graphics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 19(1):297–301, February 1987. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mengant:2003:NBJ

[Men03] J. Y. Mengant. A.NET bridge to a Java Virtual Machine: Java and .NET interoperability, with a little help from C++. C/C++ Users Journal, 21(10):28, 2003. CODEN CCUJEX. ISSN 1075-2838.

Merelli:2019:EDC

[MFT^{+}19] Ivan Merelli, Federico Fornari, Fabio Tordini, Daniele D’Agostino, Marco Aldinucci, and Daniele Cesini. Exploiting Docker containers over grid computing for a comprehensive study of chromatin conformation in different cell types. Journal of Parallel and Distributed Computing, 134(??):116–127, December 2019. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519305593.

Morimoto:2008:WSH

[MG08] Rand Morimoto and Jeff Guillet. Windows Server 2008 Hyper-V unleashed. SAMS Publishing, Indianapolis, IN, USA, 2008. ISBN 0-672-33028-8. xv + 459 pp. LCCN QA76.9.V5 M656 2009.

Morimoto:2009:WSH

[MG09] Rand Morimoto and Jeff Guillet. Windows Server 2008 Hyper-V unleashed. SAMS Publishing, Indianapolis, IN, USA, 2009. ISBN 0-7686-8567-2, 0-672-33028-8. xv + 459 pp. LCCN QA76.9.V5 M656 2009. URL http://proquest.

safaribooksonline.com/9780768685671.

Medina:2013:SMM

[MG13] Violeta Medina and Juan Manuel García. A survey of migration mechanisms of virtual machines. ACM Computing Surveys, 46(3):30:1–30:??, January 2013. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Makowski:2019:EVT

[MG19] Lukasz Makowski and Paola Grosso. Evaluation of virtualization and traffic filtering methods for container networks. Future Generation Computer Systems, 93(??):345–357, April 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167739X18302371.

Montella:2017:VCB

[MGL^{+}17] Raffaele Montella, Giulio Giunta, Giuliano Laccetti, Marco Lapegna, Carlo Palmieri, Carmine Ferraro, Valentina Pelliccia, Cheol-Ho Hong, Ivor Spence, and Dimitrios S. Nikolopoulos. On the virtualization of CUDA based GPU remoting on ARM and x86 machines in the GVirtuS framework. International Journal of Parallel Programming, 45(5):1142–1163, October 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Mohammadhosseini:2019:EEA

[MHM19] Mahdieh Mohammadhosseini, Abolfazl Toroghi Haghighat, and Ebrahim Mahdipour. An efficient energy-aware method for virtual machine placement in cloud data centers using the cultural algorithm. The Journal of Supercomputing, 75(10): 6904–6933, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mostafavi:2021:QSP

[MHS21] Seyedakbar Mostafavi, Vesal Hakami, and Maryam Sanaei. Quality of service provisioning in network function virtualization: a survey. Computing: Archiv für Informatik und Numerik, 103(5):917–991, May 2021. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00607-021-00925-x.

Matthys:2005:IVE

[MIS^{+}05] Christian Matthys, Hiroshi Ito, Bill Souhrada, Lynn Winkelbauer, Holger M. Wittmann, et al. IBM virtualization engine, version 1: planning and installation guide. IBM redbooks. IBM Corporation, San Jose, CA, USA, 2005. ISBN 0-7384-9277-9. xvi + 398 pp. LCCN MLCM 2006/42092 (Q).

Mzaik:1993:SPA

[MJ93] T. Mzaik and J. Jagadeesh. Simulation of pyramid architecture on the connection machine and implementation of algorithms. In National Aerospace and Electronics Conference. NAECON 1993, pages 838–44. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993.

Muller:2006:SVP

[MJW^{+}06] Al Muller, Andy Jones, David E. Williams, Stephen Beaver, David A. Payne, Jeremy Pries, and David E. Hart. Scripting VMware Power Tools: Automating Virtual Infrastructure Administration. Syngress Publishing, Inc., Rockland, MA, USA, 2006. ISBN 1-59749-059-8. xviii + 398 pp. LCCN QA76.76.O63 S385 2006eb.

Mao:2014:RPO

[MJW^{+}14] Bo Mao, Hong Jiang, Suzhen Wu, Yinjin Fu, and Lei Tian. Read-performance optimization for deduplication-based storage systems in the cloud. ACM Transactions on Storage, 10 (2):6:1–6:??, March 2014. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Mandal:2013:EVM

[MK13] Sameer Kumar Mandal and Pabitra Mohan Khilar. Efficient virtual machine placement for on-demand access to infrastructure resources in cloud computing. International Journal of Computer Applications, 68(??):6–11, April 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume68/number12/11629-7101/.

Mavridis:2019:CCV

[MK19] Ilias Mavridis and Helen Karatza. Combining containers and virtual machines to enhance isolation and extend functionality on cloud computing. Future Generation Computer Systems, 94(??):674–696, May 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X18305764.

Mangalampalli:2022:WDN

[MK22] Ashish Mangalampalli and Avinash Kumar. WBATimeNet: a deep neural network approach for VM Live Migration in the cloud. Future Generation Computer Systems, 135(??):438–449, October 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0167739X2200187X.

Mavridis:2023:OSC

[MK23] Ilias Mavridis and Helen Karatza. Orchestrated sandboxed containers, unikernels, and virtual machines for isolationenhanced multitenant workloads and serverless computing in cloud. Concurrency and Computation: Practice and Experience, 35(11):e6365:1–e6365:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Min:2012:VVM

[MKKE12] Changwoo Min, Inhyeok Kim, Taehyoung Kim, and Young Ik Eom. VMMB: Virtual machine memory balancing for unmodified operating systems. Journal of Grid Computing, 10(1):69–84, March 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&
issn=1570-7873&volume=10&issue=1&spage=69.

McDermott:2008:REX

[MKM^{+}08] J. McDermott, J. Kirby, B. Montrose, T. Johnson, and M. Kang. Re-engineering Xen internals for higher-assurance security. Information Security Technical Report, 13(1):17– 24, ???? 2008. CODEN ISTRFR. ISSN 1363-4127 (print),

1873-605X (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1363412708000022.

Marevac:2025:FDD

[MKZ^{+}25] Elmin Marevac, Esad Kadusić, Natasa Zivić, Nevzudin Buzadija, and Samir Lemes. Framework design for the dynamic reconfiguration of IoT-enabled embedded systems and “on-the-fly” code execution. Future Internet, 17(1):23, January 7, 2025. CODEN ???? ISSN 1999-5903. URL https:/

/www.mdpi.com/1999-5903/17/1/23.

Malik:1978:DOH

[ML78] Kamran Malik and Ted Lewis. Design objectives for high level microprogramming languages. ACM SIGMICRO Newsletter, 9(4):154–160, December 1978. CODEN SIGMDJ. ISSN 0163-

5751, 1050-916X. URL https://dl.acm.org/doi/10.1145/
1014198.804325.

Mendelsohn:1983:RVF

[MLA83] Noah Mendelsohn, Mark H. Linehan, and William J. Anzick. Reflections on VM/Pass-Through: a facility for interactive networking. IBM Systems Journal, 23(1-2):63–79, 1983. CO-DEN IBMSA7. ISSN 0018-8670.

Mikheev:2002:OEJ

[MLG^{+}02] V. Mikheev, N. Lipsky, D. Gurchenkov, P. Pavlov, V. Sukharev, A. MarkovS. Kuksenko, S. Fedoseev, D. Leskov, and A. Yeryomin. Overview of Excelsior JET, a high performance alternative to Java Virtual Machines. In Proceedings of the third international workshop on Software and performance: WOSP ’02, July 24–26, 2002 Rome, Italy, pages 104–113. ACM Press, New York, NY 10036, USA, 2002.

Ma:2019:PMA

[MLXG19] Y. Ma, W. Liang, Z. Xu, and S. Guo. Profit maximization for admitting requests with network function services in distributed clouds. IEEE Transactions on Parallel and Distributed Systems, 30(5):1143–1157, May 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mlynski:2009:IIP

[Mly09] Maciej Mlynski. The influence of the IBM pSeries servers virtualization mechanism on dynamic resource allocation in AIX 5L. Scalable Computing: Practice and Experience, 10 (2):189–199, June 2009. CODEN ???? ISSN 1895-1767.

URL http://www.scpe.org/vols/vol10/no2/SCPE_10_2_
05.pdf;   http://www.scpe.org/vols/vol10/no2/SCPE_
10_2_05.zip.

Majumdar:1992:PPC

[MM92] A. Majumdar and W. R. Martin. Parallel preconditioned conjugate gradient algorithm applied to neutron diffusion problem. Transactions of the American Nuclear Society, 65:209– 210, 1992. CODEN TANSAO. ISSN 0003-018X.

Manning:1993:AAE

[MM93] Serge M. Manning and David G. Meyer. Analysis of asynchronous execution streams with I-caching in massively parallel systems. Journal of Parallel and Distributed Computing, 19(3):279–291, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.1993.1112/
production;   http://www.idealibrary.com/links/doi/
10.1006/jpdc.1993.1112/production/pdf.

Muntean:1994:PGM

[MM94] T. Muntean and A. Melo. PAROS: a generic multi virtual machines parallel operating system. In Joubert et al. [JPTE94], pages 319–328. ISBN 0-444-81841-3. LCCN QA76.58 .P3794 1993.

Marangozova-Martin:2019:MLE

[MMdE19] V. Marangozova-Martin, N. de Palma, and A. El Rheddane. Multi-level elasticity for data stream processing. IEEE Transactions on Parallel and Distributed Systems, 30(10):2326– 2337, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Martinez-Morfa:2025:FLS

[MMdMCPSR25] Mario Martínez-Morfa, Carlos Ruiz de Mendoza, Cristina Cervelló-Pastor, and Sebastia Sallent-Ribes. Federated learning system for dynamic radio/MEC resource allocation and slicing control in open radio access network. Future Internet, 17(3):106, February 26, 2025. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.com/1999-5903/17/3/106.

Montella:2018:MBP

[MMG^{+}18] Raffaele Montella, Livia Marcellino, Ardelio Galletti, Diana Di Luccio, Sokol Kosta, Giuliano Laccetti, and Giulio Giunta. Marine bathymetry processing through GPGPU virtualization in high performance cloud computing. Concurrency and Computation: Practice and Experience, 30(24):e4895:1– e4895:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mohapatra:2017:PEV

[MMHP17] Subasish Mohapatra, Subhadarshini Mohanty, Arunima Hota, and Shradha Pattanayak. Performance evaluation of virtualization in a private cloud test bed. International Journal of Computer Applications, 162(??): 20–27, March 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume162/number7/27256-2017913386/.

MacGregor:1984:MM

[MMM84] Doug MacGregor, Dave Mothersole, and Bill Moyer. The Motorola MC68020. IEEE Micro, 4(4):101–118, July/August 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mirzaei:2012:TAA

[MMP^{+}12] Nariman Mirzaei, Sam Malek, Corina S. Păsăreanu, Naeem Esfahani, and Riyadh Mahmood. Testing Android apps through symbolic execution. ACM SIGSOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Memari:2022:LAT

[MMTM22] Pedram Memari, Seyedeh Samira Mohammadi, and Reza Tavakkoli-Moghaddam. A latency-aware task scheduling algorithm for allocating virtual machines in a cost-effective and time-sensitive fog-cloud architecture. The Journal of Supercomputing, 78(1):93–122, January 2022. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-021-03868-4.

Meleshchuk:1991:IPP

[MN91] S. B. Meleshchuk and A. N. Nedumov. Implementation of a protocol for parallel database access with virtual machine communications facilities. Programmirovanie, 17(1):35–42, January/February 1991. CODEN PCSODA. ISSN 0132-3474, 0361-7688. English translation in Programming and Computer Software, vol. 17, no. 1, pp. 27–32, November 1991.

McAuley:2003:CVC

[MN03] Derek McAuley and Rolf Neugebauer. A case for virtual channel processors. In Romanow and Mogul [RM03], pages 237– 242. ISBN ???? LCCN TK5105.5. ACM order number 534032.

Masdari:2016:OVM

[MNA16] Mohammad Masdari, Sayyid Shahab Nabavi, and Vafa Ahmadi. An overview of virtual machine placement schemes in cloud computing. Journal of Network and Computer Applications, 66(??):106–127, May 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804516000291.

Mahmoodabadi:2024:AAV

[MNB24] Zahra Mahmoodabadi and Mostafa Nouri-Baygi. An approximation algorithm for virtual machine placement in cloud data centers. The Journal of Supercomputing, 80(1):915–941, January 2024. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL https://link.springer.com/
article/10.1007/s11227-023-05505-8.

McKusick:2005:DIF

[MNN05] Marshall Kirk McKusick and George V. Neville-Neil. The Design and Implementation of the FreeBSD Operating System. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-201-70245-2 (hardcover). xxviii + 683 pp. LCCN QA76.76.O63 M398745 2005. URL http://www.mckusick.

com/FreeBSDbook.html.

Mitsuishi:2014:ABF

[MNS^{+}14] Takuji Mitsuishi, Shimpei Nomura, Jun Suzuki, Yuki Hayashi, Masaki Kan, and Hideharu Amano. Accelerating breadth first search on GPU–BOX. ACM SIGARCH Computer Architecture News, 42(4):81–86, 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Machida:2014:JCT

[MNT14] Fumio Machida, Victor F. Nicola, and Kishor S. Trivedi. Job completion time on a virtualized server with software rejuvenation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(1):10:1–10:??, January 2014. CO-DEN ???? ISSN 1550-4832.

McGhan:1998:CPP

[MO98] Harlan McGhan and Mike O’Connor. Computing practices: PicoJava: a direct execution engine for Java bytecode. Computer, 31(10):22–30, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/books/co1998/pdf/rx022.pdf;
 http://www.computer.org/computer/co1998/rx022abs.
htm.

Montague:1997:JEJ

[Mon97] Bruce R. Montague. JN: OS for an embedded Java Network Computer — supporting Java’s Virtual Machine on a singlechip embedded PC attached to the Internet. IEEE Micro, 17 (3):54–60, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http://pascal.

computer.org/mi/books/mi1997/pdf/m3054.pdf.

Montazerolghaem:2022:SVV

[Mon22] Ahmadreza Montazerolghaem. Softwarization and virtualization of VoIP networks. The Journal of Supercomputing, 78(12):14471–14503, August 2022. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.com/article/10.1007/s11227-
022-04448-w.

Moore:2001:EFJ

[MP01] J. Strother Moore and George M. Porter. An executable formal Java Virtual Machine thread model. In USENIX [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.usenix.org/publications/
library/proceedings/jvm01/moore.html.

Martini:2016:SOA

[MP16] Barbara Martini and Federica Paganelli. A service-oriented approach for dynamic chaining of virtual network functions over multi-provider software-defined networks. Future Internet, 8(2):24, June 01, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.com/1999-5903/8/2/24.

Meloni:2018:CBI

[MPA^{+}18] A. Meloni, P. A. Pegoraro, L. Atzori, A. Benigni, and S. Sulis. Cloud-based IoT solution for state estimation in smart grids: Exploiting virtualization and edge-intelligence technologies. Computer Networks (Amsterdam, Netherlands: 1999), 130(??):156–165, January 15, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128617303912.

Muir:2006:POP

[MPF^{+}06] Steve Muir, Larry Peterson, Marc Fiuczynski, Justin Cappos, and John Hartman. Privileged operations in the PlanetLab virtualised environment. Operating Systems Review, 40(1): 75–88, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Monge:2020:COM

[MPM^{+}20] David A. Monge, Elina Pacini, Cristian Mateos, Enrique Alba, and Carlos García Garino. CMI: an online multiobjective genetic autoscaler for scientific and engineering workflows in cloud infrastructures with unreliable virtual machines. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804519303248.

Mylopoulos:1991:IPT

[MR91] John Mylopoulos and Ray Reiter, editors. IJCAI-91: proceedings of the Twelfth International Conference on Artificial Intelligence, Darling Harbour, Sydney, Australia, 24–30 August 1991. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1991. ISBN 1-55860-160-0. LCCN Q334 .I57 1991. Two volumes.

Miller:2004:CLI

[MR04] James S. Miller and Susann Ragsdale. The Common Language Infrastructure Annotated Standard. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-15493-2. xxxii + 891 pp. LCCN QA76.7 .M52 2003.

Moreno:2006:NV

[MR06] Victor Moreno and Kumar Reddy. Network virtualization. Cisco Press, Indianapolis, IN, USA, 2006. ISBN 1-58705-248-2. xx + 365 pp. LCCN TK5105.875.E87

M65 2006. URL http://www.loc.gov/catdir/toc/fy0703/
2005926713.html.

Madireddy:2023:DVM

[MR23] Amarendhar Reddy Madireddy and Kongara Ravindranath. Dynamic virtual machine relocation system for energyefficient resource management in the cloud. Concurrency and Computation: Practice and Experience, 35(3):e7520:1– e7520:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Minhas:2013:RTH

[MRC^{+}13] Umar Farooq Minhas, Shriram Rajagopalan, Brendan Cully, Ashraf Aboulnaga, Kenneth Salem, and Andrew Warfield.

RemusDB: transparent high availability for database systems. VLDB Journal: Very Large Data Bases, 22(1):29–45, February 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Meier:2017:PVM

[MRG17] Remigius Meier, Armin Rigo, and Thomas R. Gross. Parallel virtual machines with RPython. ACM SIGPLAN Notices, 52(2):48–59, February 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meier:2018:VMD

[MRG18] Remigius Meier, Armin Rigo, and Thomas R. Gross. Virtual machine design for parallel dynamic programming languages. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):109:1–109:25, October 2018. URL

https://dl.acm.org/doi/abs/10.1145/3276479.

Malan:1991:MA

[MRGB91] G. Malan, R. Rashid, D. Golub, and R. Baron. DOS as a Mach 3.0 application. In USENIX [USE91], pages 27–40. LCCN QAX 27.

Moure:2002:KS

[MRL02] J. C. Moure, Dolores I. Rexachs, and Emilio Luque. The KScalar simulator. ACM Journal on Educational Resources in Computing (JERIC), 2(1):73–116, March 2002. CODEN ???? ISSN 1531-4278.

Marshall:2006:ASV

[MRM06] David (David W.) Marshall, Wade A. Reynolds, and Dave McCrory. Advanced server virtualization: VMware and Microsoft platforms in the virtual data center. Auerbach, Princeton, NJ, USA, 2006. ISBN 0-8493-3931-6. xvi + 742 pp. LCCN QA76.76.O63 M3646 2006. URL http://

www.loc.gov/catdir/enhancements/fy0702/2006008008-
d.html;   http://www.loc.gov/catdir/toc/ecip0610/
2006008008.html.

Meyer:1970:VMT

[MS70] Richard A. Meyer and Love H. Seawright. A virtual machine time-sharing system. IBM Systems Journal, 9(3):199–218, 1970. CODEN IBMSA7. ISSN 0018-8670.

Manas:1991:VLM

[MS91a] J. A. Manas and J. Salvachua. A virtual LOTOS machine. In Kenneth R. Parker and Gordon A. Rose, editors, Proceedings of Formal Description Techniques IV, pages 441–456. North-Holland, Amsterdam, The Netherlands, 1991.

Milutinovic:1991:PTA

[MS91b] V. Milutinovic and B. D. Shriver, editors. Proceedings of the Twenty-Fourth Annual Hawaii International Conference on System Sciences. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-2119-2 (vol. 1), 0-8186-2120-6 (vol. 2), 0-8186-2010-2 (vol. 3), 0-8186-2122-2 (vol. 4). LCCN QA76.9.S88 H38 1991; QA76.9.S88; QA76.9.S88 H375 1991; Internet; TA168 .H37a 1991. Four volumes. IEEE catalog number 91TH0350-9.

Mathiske:2000:APM

[MS00] Bernd Mathiske and Daniel Schneider. Automatic persistent memory management for the Spotless virtual machine on the Palm connected organizer. SMLI TR 2000-89, Sun Microsystems, Palo Alto, CA, USA, June 2000. 18 pp.

Menczer:2001:OTR

[MS01] Filippo Menczer and Alberto Maria Segre. OAMulator: a teaching resource to introduce computer architecture concepts. ACM Journal on Educational Resources in Computing (JERIC), 1(4):18–30, December 2001. CODEN ???? ISSN 1531-4278.

Mann:2017:WBA

[MS17] Zoltán Ádám Mann and Máté Szabó. Which is the best algorithm for virtual machine placement optimization? Concurrency and Computation: Practice and Experience, 29(10), May 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moreno:2021:OSF

[MSC^{+}21] Jesús Fernando Cevallos Moreno, Rebecca Sattler, Raúl P. Caulier Cisterna, Lorenzo Ricciardi Celsi, Aminael Sánchez Rodríguez, and Massimo Mecella. Online service function chain deployment for live-streaming in virtualized content delivery networks: a deep reinforcement learning approach.

Future Internet, 13(11):278, October 29, 2021. CODEN ????

ISSN 1999-5903. URL https://www.mdpi.com/1999-5903/
13/11/278.

Mebane:1992:EFD

[MSCK92] Alfred Holt Mebane, IV, James R. Schmedake, Iue-Shuenn Chen, and Anne P. Kadonaga. Electronic and firmware design of the HP DesignJet drafting plotter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(6):16–23, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Maessen:2001:PAS

[MSG01] Jan-Willem Maessen, Vivek Sarkar, and David Grove. Program analysis for safety guarantees in a Java virtual machine written in Java. In ACM [ACM01a], pages 62–65. ISBN 1-58113-413-4. LCCN QA76.758. URL http://

www.acm.org/pubs/articles/proceedings/soft/379605/
p62-maessen/p62-maessen.pdf;   http://www.acm.org/
pubs/citations/proceedings/soft/379605/p62-maessen/

. Supplement to ACM SIGPLAN Notices.

Ma:2012:DTD

[MSG^{+}12] Zhiqiang Ma, Zhonghua Sheng, Lin Gu, Liufei Wen, and Gong Zhang. DVM: towards a datacenter-scale virtual machine. ACM SIGPLAN Notices, 47(7):39–50, July 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Ma:2014:DBV

[MSG14] Zhiqiang Ma, Zhonghua Sheng, and Lin Gu. DVM: A big virtual machine for cloud computing. IEEE Transactions on Computers, 63(9):2245–2258, September 2014. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Matsuhashi:2012:TVF

[MSI^{+}12] Yohei Matsuhashi, Takahiro Shinagawa, Yoshiaki Ishii, Nobuyuki Hirooka, and Kazuhiko Kato. Transparent VPN failure recovery with virtualization. Future Generation Computer Systems, 28(1):78–84, January 2012. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X11001026.

Mashimo:2018:VMS

[MSI18] Susumu Mashimo, Ryota Shioya, and Koji Inoue. VMOR: Microarchitectural support for operand access in an interpreter. IEEE Computer Architecture Letters, 17(2):217–220, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mao:2024:JVN

[MSLY24] Yingling Mao, Xiaojun Shang, Yu Liu, and Yuanyuan Yang. Joint virtual network function placement and flow routing in edge-cloud continuum. IEEE Transactions on Computers, 73 (3):872–886, March 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Maslak:1991:CRR

[MSS91] B. A. Maslak, J. M. Showalter, and T. J. Szczygielski. Coordinated Resource Recovery in VM/ESA. IBM Systems Journal, 30(1):72–89, 1991. CODEN IBMSA7. ISSN 0018-8670.

Ma:2015:SDS

[MSS^{+}15] Jiuyue Ma, Xiufeng Sui, Ninghui Sun, Yupeng Li, Zihao Yu, Bowen Huang, Tianni Xu, Zhicheng Yao, Yun Chen, Haibin Wang, Lixin Zhang, and Yungang Bao. Supporting differentiated services in computers via programmable architecture for resourcing-on-demand (PARD). ACM SIGPLAN Notices, 50 (4):131–143, April 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Menon:2005:DPO

[MST^{+}05] Aravind Menon, Jose Renato Santos, Yoshio Turner, G. (John) Janakiraman, and Willy Zwaenepoel. Diagnosing performance overheads in the Xen virtual machine environment. In ACM [ACM05d], pages 13–23. ISBN 1-59593-047-7. LCCN QA76.9.V5 I575 2005. URL http://www.loc.gov/

catdir/toc/fy0611/2006530661.html. ACM order number

548059.

Menon:2009:TSA

[MSZ09] Aravind Menon, Simon Schubert, and Willy Zwaenepoel. TwinDrivers: semi-automatic derivation of fast and safe hypervisor network drivers from guest OS drivers. ACM SIGARCH Computer Architecture News, 37(1):301–312, March 2009. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Merrifield:2016:PIE

[MT16] Timothy Merrifield and H. Reza Taheri. Performance implications of extended page tables on virtualized x86 processors. ACM SIGPLAN Notices, 51(7):25–35, July 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Merrifield:2017:PIE

[MT17] Timothy Merrifield and H. Reza Taheri. Performance implications of extended page tables on virtualized x86 processors. Operating Systems Review, 51(1):38–47, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mao:2019:AMC

[MTFK19] B. Mao, F. Tang, Z. M. Fadlullah, and N. Kato. An absorbing Markov chain based model to solve computation and communication tradeoff in GPU-accelerated MDRUs for safety confirmation in disaster scenarios. IEEE Transactions on Computers, 68(9):1256–1268, September 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Moratelli:2022:MRV

[MTJ^{+}22] Carlos Moratelli, Ramão Tiburski, Sérgio F. Johann, Emanuel Moura, Everton De Matos, and Fabiano Hessel. MIPS and RISC-V: Evaluating virtualization trade-off for edge devices. In IEEE, editor, 2022 IEEE 8th World Forum on Internet of Things (WF-IoT), pages 1–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2022.

Muhlbacher:1975:GIF

[Müh75] Jörg Mühlbacher, editor. GI — 5. [i.e. funfte] Jahrestagung: Dortmund, 8.–10. Oktober 1975, volume 34 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNCSD9. ISBN 0-387-07410-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.34. English or German.

Mergen:2006:VHP

[MUKX06] Mark F. Mergen, Volkmar Uhlig, Orran Krieger, and Jimi Xenidis. Virtualization for high-performance computing. Operating Systems Review, 40(2):8–11, April 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Murphy:1969:TSU

[Mur69] Daniel Murphy. The Ten-Sys user virtual machine. TEN-SYS Memorandum 8, Digital Equipment Corporation, Marlboro, MA, USA, April 30, 1969. 6 pp. URL https://walden-

family.com/bbn/10-SYS/TEN-SYS-8.pdf.

Marz:2016:RPC

[MV16] Stephen Marz and Brad Vander Zanden. Reducing power consumption and latency in mobile devices using an event stream model. ACM Transactions on Embedded Computing Systems, 16(1):11:1–11:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Moreno-Vozmediano:2024:IRO

[MVMHL24] Rafael Moreno-Vozmediano, Rubén S. Montero, Eduardo Huedo, and Ignacio M. Llorente. Intelligent resource orchestration for 5G edge infrastructures. Future Internet, 16(3): 103, March 19, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/1999-5903/16/3/103.

Munawar:2005:BPB

[MW05] Mohammad A. Munawar and Paul A. S. Ward. Better performance or better manageability? ACM SIGSOFT Software Engineering Notes, 30(4):1–4, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Miao:2018:VMA

[MW18] Xuna Miao and Xiaobo Wu. Virtual machine anomaly detection strategy based on cloud platform operating environment perception. Concurrency and Computation: Practice and Experience, 30(22):e4656:1–e4656:??, November 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moron:2025:BWE

[MW25] Konrad Moron and Stefan Wallentowitz. Benchmarking WebAssembly for embedded systems. ACM Transactions on Architecture and Code Optimization, 22(3):83:1–83:21, September 2025. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Muller:2005:VVE

[MWHH05] Al Muller, Seburn Wilson, Don Happe, and Gary J. Humphrey, editors. Virtualization with VMware ESX Server. Syngress Publishing, Inc., Rockland, MA, USA, 2005. ISBN 1-59749-019-9. xxiii + 463 pp. LCCN QA76.76.O63 V465 2005; QA76.9.V5 V57 2005eb. URL http://site.ebrary.

com/lib/ucsc/Doc?id=10110269.

Masdari:2020:GCC

[MZ20] Mohammad Masdari and Mehran Zangakani. Green cloud computing using proactive virtual machine placement: Challenges and issues. Journal of Grid Computing, 18(4):727– 759, December 2020. CODEN ???? ISSN 1570-7873 (print),

1572-9184 (electronic). URL https://link.springer.com/
article/10.1007/s10723-019-09489-9.

Ma:2018:GEG

[MZD^{+}18] Jiacheng Ma, Xiao Zheng, Yaozu Dong, Wentai Li, Zhengwei Qi, Bingsheng He, and Haibing Guan. gMig: Efficient GPU live migration optimized by software dirty page for full virtualization. ACM SIGPLAN Notices, 53(3):31–44, March 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mihajlovic:2014:DIQ

[MZG14] Bojan Mihajlović, Zeljko Zilić, and Warren J. Gross. Dynamically instrumenting the QEMU emulator for Linux process trace generation with the GDB debugger. ACM Transactions on Embedded Computing Systems, 13(5s):167:1–167:??, November 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nashaat:2019:SES

[NAR19] Heba Nashaat, Nesma Ashry, and Rawya Rizk. Smart elastic scheduling algorithm for virtual machine migration in cloud computing. The Journal of Supercomputing, 75(7):3842–3865, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nemati:2021:HBV

[NASD21] Hani Nemati, Seyed Vahid Azhari, Mahsa Shakeri, and Michel Dagenais. Host-based virtual machine workload characterization using hypervisor trace mining. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 6(1):4:1–4:25, June 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/10.1145/3460197.

Nikolaev:2011:PXF

[NB11] Ruslan Nikolaev and Godmar Back. Perfctr-Xen: a framework for performance counter virtualization. ACM SIGPLAN Notices, 46(7):15–26, July 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ngoc:2019:EYS

[NBB^{+}19] Tu Dinh Ngoc, Bao Bui, Stella Bitchebe, Alain Tchana, Valerio Schiavoni, Pascal Felber, and Daniel Hagimont. Everything you should know about Intel SGX performance on virtualized systems. ACM SIGMETRICS Performance Evaluation Review, 47(1):77–78, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:

//dl.acm.org/doi/abs/10.1145/3376930.3376979.

Nance:2008:VMI

[NBH08] Kara Nance, Matt Bishop, and Brian Hay. Virtual machine introspection: Observation or interference? IEEE Security & Privacy, 6(5):32–37, September/October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Nathan:2016:SRO

[NBK16] Senthil Nathan, Umesh Bellur, and Purushottam Kulkarni. On selecting the right optimizations for virtual machine migration. ACM SIGPLAN Notices, 51(7):37–49, July 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Naranjo:2018:DEE

[NBS18] Paola G. Vinueza Naranjo, Enzo Baccarelli, and Michele Scarpiniti. Design and energy-efficient resource management of virtualized networked fog architectures for the real-time support of IoT applications. The Journal of Supercomputing, 74(6):2470–2507, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nelson:2004:CDC

[Nel04] T. Nelson. A cosmology for a different computer universe: Data model, mechanisms, virtual machine and visualization infrastructure. Journal of Digital Information, 5(1):??, July 2004. CODEN ???? ISSN 1368-7506. URL http://jodi.

tamu.edu/Articles/v05/i01/Nelson.

Neuman:COMPSYS-5-4-407

[Neu92] B. Clifford Neuman. The Prospero file system: a global file system based on the virtual system model. Computing Systems, 5(4):407–432, Fall 1992. CODEN CMSYE2. ISSN 0895-6340.

Ng:2001:VEWa

[Ng01a] Choong Ng. VMware Express 2.0 and Win4Lin 2.0: a comparison review. Linux Journal, 85:??, May 2001. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic). URL http://www.linuxjournal.com/articles/

linux_review/0036.html. Not in journal issue: Web only.

Ng:2001:VEWb

[Ng01b] Choong Ng. VMware express 2.0 and win4lin 2.0: a comparison review. Linux Journal, 86:??, June 2001. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic). URL http://www.linuxjournal.com/articles/

linux_review/0036.html. Not in journal issue: Web only.

Noll:2013:OFD

[NG13] Albert Noll and Thomas Gross. Online feedback-directed optimizations for parallel Java code. ACM SIGPLAN Notices, 48(10):713–728, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Naeen:2024:CDC

[NGN24] Mohammad Ali Monshizadeh Naeen, Hamid Reza Ghaffari, and Hossein Monshizadeh Naeen. Cloud data center cost management using virtual machine consolidation with an improved artificial feeding birds algorithm. Computing: Archiv für Informatik und Numerik, 106(6):1795–1823, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00607-024-01267-0.

Nguyen:2019:RFV

[NGRF19] Thi-Minh Nguyen, Andre Girard, Catherine Rosenberg, and Serge Fdida. Routing via functions in virtual networks: The curse of choices. IEEE/ACM Transactions on Networking, 27(3):1192–1205, June 2019. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nong:2022:ARC

[NHL22] Mengxin Nong, Lingfeng Huang, and Mingtao Liu. Allocation of resources for cloud survivability in smart manufacturing. ACM Transactions on Management Information Systems (TMIS), 13(4):45:1–45:??, December 2022. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic). URL

https://dl.acm.org/doi/10.1145/3533701.

Noshy:2018:OLV

[NIA18] Mostafa Noshy, Abdelhameed Ibrahim, and Hesham Arafat Ali. Optimization of live virtual machine migration in cloud computing: A survey and future directions. Journal of Network and Computer Applications, 110(??):1–10, May 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1084804518300833.

Nieh:2012:CBR

[Nie12] Jason Nieh. Challenges in building a real, large private cloud. ACM SIGPLAN Notices, 47(7):1–2, July 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Namjoshi:2010:NOP

[NK10] Manjiri A. Namjoshi and Prasad A. Kulkarni. Novel online profiling for virtual machines. ACM SIGPLAN Notices, 45 (7):133–144, July 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Najjari:2022:SOL

[NK22] Ahmadreza Hassannezhad Najjari and Ali Asghar Pourhaji Kazem. A systematic overview of live virtual machine migration methods. Concurrency and Computation: Practice and Experience, 34(17):e6915:1–e6915:??, August 1, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Neumann:2006:IVT

[NKK^{+}06] Dean Neumann, Dileep Kulkarni, Aaron Kunze, Gerald Rogers, and Edwin Verplanke. Intel Virtualization Technology in embedded and communications infrastructure applications. Intel Technology Journal, 10(3), August 10,

2006. ISSN 1535-766X. URL http://developer.intel.
com/technology/itj/2006/v10i3/5-communications/1-
abstract.htm.

Nitu:2018:WSS

[NKY^{+}18] Vlad Nitu, Aram Kocharyan, Hannas Yaya, Alain Tchana, Daniel Hagimont, and Hrachya Astsatryan. Working set size estimation techniques in virtualized environments: One size does not fit all. ACM SIGMETRICS Performance Evaluation Review, 46(1):62–63, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nieh:2000:EV

[NL00] Jason Nieh and Ozgur Can Leonard. Examining VMware. Dr. Dobb’s Journal of Software Tools, 25(8):70, 72–74, 76, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Niyizamwiyitira:2019:UBS

[NL19] Christine Niyizamwiyitira and Lars Lundberg. A utilizationbased schedulability test of real-time systems running on a multiprocessor virtual machine. The Computer Journal, 62 (6):884–904, June 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://academic.oup.

com/comjnl/article/62/6/884/5319152.

Nguyen:2023:TNQ

[NLD^{+}23] Thien Nguyen, Dmitry Lyakh, Eugene Dumitrescu, David Clark, Jeff Larkin, and Alexander McCaskey. Tensor network quantum virtual machine for simulating quantum circuits at exascale. ACM Transactions on Quantum Computing (TQC), 4(1):6:1–6:??, March 2023. CODEN ???? ISSN 2643-6809 (print), 2643-6817 (electronic). URL https://dl.

acm.org/doi/10.1145/3547334.

Nguyen:2012:BAS

[NLPV12] Dinh Khoa Nguyen, Francesco Lelli, Mike P. Papazoglou, and Willem-Jan Van den Heuvel. Blueprinting approach in support of cloud computing. Future Internet, 4(1):322–346, March 21, 2012. CODEN ???? ISSN 1999-5903. URL https:

//www.mdpi.com/1999-5903/4/1/322.

Nguyen:2018:CEA

[NMC18a] Tuan Anh Nguyen, Dugki Min, and Eunmi Choi. A comprehensive evaluation of availability and operational cost for a virtualized server system using stochastic reward nets. The Journal of Supercomputing, 74(1):222–276, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [NMC18b].

Nguyen:2018:CCE

[NMC18b] Tuan Anh Nguyen, Dugki Min, and Eunmi Choi. Correction to: A comprehensive evaluation of availability and operational cost for a virtualized server system using stochastic reward nets. The Journal of Supercomputing, 74(6):2903, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11227-018-2428-0.pdf. See [NMC18a].

Nejad:2015:TGM

[NMG15] Mahyar Movahed Nejad, Lena Mashayekhy, and Daniel Grosu. Truthful greedy mechanisms for dynamic virtual machine provisioning and allocation in clouds. IEEE Transactions on Parallel and Distributed Systems, 26(2):594–603, February 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.computer.org/

csdl/trans/td/2015/02/06748071-abs.html.

Nowatzki:2015:ASC

[NMHS15] Tony Nowatzki, Jaikrishnan Menon, Chen-Han Ho, and Karthikeyan Sankaralingam. Architectural simulators considered harmful. IEEE Micro, 35(6):4–12, November/December 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143

(electronic). URL http://csdl.computer.org/csdl/mags/
mi/2015/06/mmi2015060004-abs.html.

Ngo:2015:RES

[NMMP15] Minh Ngo, Fabio Massacci, Dimiter Milushev, and Frank Piessens. Runtime enforcement of security policies on black box reactive programs. ACM SIGPLAN Notices, 50(1):43–54, January 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nomura:2014:PAM

[NMS^{+}14] Shimpei Nomura, Takuji Mitsuishi, Jun Suzuki, Yuki Hayashi, Masaki Kan, and Hideharu Amano. Performance analysis of the Multi-GPU system with ExpEther. ACM SIGARCH Computer Architecture News, 42(4):9–14, 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Najafizadegan:2021:AMS

[NNK21] Negin Najafizadegan, Eslam Nazemi, and Vahid Khajehvand. An autonomous model for self-optimizing virtual machine selection by learning automata in cloud environment. Software —Practice and Experience, 51(6):1352–1386, June 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nanba:1985:VAV

[NOK^{+}85] S. Nanba, N. Ohno, H. Kubo, H. Morisue, T. Ohshima, and H. Yamagishi. VM/4: ACOS-4 virtual machine architecture. ACM SIGARCH Computer Architecture News, 13(3): 171–178, June 1985. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Nejad:2015:SPV

[NOR15] Bijan Chokoufe Nejad, Thorsten Ohl, and Jürgen Reuter. Simple, parallel virtual machines for extreme computations.

Computer Physics Communications, 196(??):58–69, November 2015. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465515001952.

Nitu:2017:SBQ

[NOT^{+}17] Vlad Nitu, Pierre Olivier, Alain Tchana, Daniel Chiba, Antonio Barbalace, Daniel Hagimont, and Binoy Ravindran. Swift birth and quick death: Enabling fast parallel guest boot and destruction in the Xen hypervisor. ACM SIGPLAN Notices, 52(7):1–14, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nourse:1992:MWN

[Nou92] Andrew W. Nourse. Microsoft Windows network virtual device drivers in PATHWORKS for DOS. Digital Technical Journal of Digital Equipment Corporation, 4(1):47–55 (or 47–56??), Winter 1992. CODEN DTJOEL. ISSN 0898-

901X. URL ftp://ftp.digital.com/pub/Digital/info/
DTJ/v4n1/Microsoft_Windows_Network_Virt_01oct1992DTJ506P8.
ps;   http://www.digital.com:80/info/DTJ506/DTJ506SC.
TXT.

Nambiar:2013:KTR

[NP13] Raghunath Nambiar and Meikel Poess. Keeping the TPC relevant! Proceedings of the VLDB Endowment, 6(11):1186– 1187, August 2013. CODEN ???? ISSN 2150-8097.

Naranjo:2020:ASC

[NRdA^{+}20] Diana M. Naranjo, Sebastián Risco, Carlos de Alfonso, Alfonso Pérez, Ignacio Blanquer, and Germán Moltó. Accelerated serverless computing based on GPU virtualization. Journal of Parallel and Distributed Computing, 139(??):32– 42, May 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519303533.

Nakanishi:1992:SSP

[NRS92] H. Nakanishi, V. Rego, and V. Sunderam. Superconcurrent simulation of polymer chains on heterogeneous networks. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 561–569. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Nathuji:2007:VCP

[NS07] Ripal Nathuji and Karsten Schwan. VirtualPower: coordinated power management in virtualized enterprise systems. Operating Systems Review, 41(6):265–278, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nezarat:2017:GTB

[NS17] Amin Nezarat and Yaser Shams. A game theoretic-based distributed detection method for VM-to-hypervisor attacks in cloud environment. The Journal of Supercomputing, 73(10): 4407–4427, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Niu:2022:NMN

[NSC^{+}22] Zhixiong Niu, Qiang Su, Peng Cheng, Yongqiang Xiong, Dongsu Han, Keith Winstein, Chun Jason Xue, and Hong Xu. NetKernel: Making network stack part of the virtualized infrastructure. IEEE/ACM Transactions on Networking, 30 (3):999–1013, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://dl.acm.org/

doi/10.1109/TNET.2021.3129806.

Nemeth:2010:ULS

[NSHW10] Evi Nemeth, Garth Snyder, Trent R. Hein, and Ben Whaley. UNIX and Linux System Administration Handbook. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2010. ISBN 0-13-148005-7. xlvii + 1279 pp. LCCN QA76.76.O63 N45 2010.

Nhan:2012:MDA

[NSJ12] Tam Le Nhan, Gerson Sunyé, and Jean-Marc Jézéquel. A model-driven approach for virtual machine image provisioning in cloud computing. Lecture Notes in Computer Science, 7592:107–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33427-6_8/.

Neiger:2006:IVT

[NSL^{+}06] Gil Neiger, Amy Santoni, Felix Leung, Dion Rodgers, and Rich Uhlig. Intel Virtualization Technology: Hardware support for efficient processor virtualization. Intel Technology Journal, 10(3):167–177, August 10, 2006. ISSN 1535-

766X. URL http://developer.intel.com/technology/
itj/2006/v10i3/1-hardware/1-abstract.htm.

No:2016:MMC

[NsP16] Jaechun No and Sung soon Park. MultiCache: Multilayered cache implementation for I/O virtualization. Scientific Programming, 2016(??):3780163:1–3780163:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X

(electronic). URL https://www.hindawi.com/journals/
sp/2016/3780163/.

Nam:2017:JNE

[NTH^{+}17] Tran Manh Nam, Nguyen Huu Thanh, Hoang Trung Hieu, Nguyen Tien Manh, Nguyen Van Huynh, and Hoang Duong Tuan. Joint network embedding and server consolidation for energy-efficient dynamic data center virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 125(??):76–89, October 9, 2017. CODEN ???? ISSN 1389-1286 (print),

1872-7069 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1389128617302517.

Nie:2024:VSP

[NTL^{+}24] Chen Nie, Chenyu Tang, Jie Lin, Huan Hu, Chenyang Lv, Ting Cao, Weifeng Zhang, Li Jiang, Xiaoyao Liang, Weikang Qian, Yanan Sun, and Zhezhi He. VSPIM: SRAM processingin-memory DNN acceleration via vector–scalar operations. IEEE Transactions on Computers, 73(10):2378–2390, October 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nagy:2018:NVI

[NTR18] Mate Nagy, Janos Tapolcai, and Gabor Retvari. Node virtualization for IP level resilience. IEEE/ACM Transactions on Networking, 26(3):1250–1263, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Nieh:2005:ETO

[NV05] Jason Nieh and Chris Vaill. Experiences teaching operating systems using virtual platforms and Linux. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):520–524, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nogales:2024:SDC

[NVV^{+}24] Borja Nogales, Ivan Vidal, Francisco Valera, Victor Sanchez-Aguero, and Diego R. Lopez. Software-driven connectivity orchestration for multidomain network functions virtualization ecosystems. IEEE Software, 41(2):88–97, March/April 2024. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Naeen:2020:AMB

[NZH20] Hossein Monshizadeh Naeen, Esmaeil Zeinali, and Abolfazl Toroghi Haghighat. Adaptive Markov-based approach for dynamic virtual machine consolidation in cloud data centers with quality-of-service constraints. Software—Practice and Experience, 50(2):161–183, February 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Oaks:2014:JPD

[Oak14] Scott Oaks. Java Performance: The Definitive Guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2014. ISBN 1-4493-5845-4 (paperback), 1-4493-6354-7 (e-book). xiv + 408 pp. LCCN

QA76.73.J38. URL http://proquest.safaribooksonline.
com/9781449363512.

Ouarnoughi:2016:ICP

[OBSR16] Hamza Ouarnoughi, Jalil Boukhobza, Frank Singhoff, and Stéphane Rubini. Integrating I/Os in Cloudsim for performance and energy estimation. Operating Systems Review, 50 (3):27–36, December 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Odette:1987:CPF

[Ode87] L. L. Odette. Compiling Prolog to Forth. Journal of Forth Applications and Research, 4(4):487–534, 1987.

OLoughlin:2016:SVM

[OG16] John O’Loughlin and Lee Gillam. Sibling virtual machine colocation confirmation and avoidance tactics for public infrastructure clouds. The Journal of Supercomputing, 72(3):961– 984, March 2016. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11227-016-1627-9.pdf.

Oglesby:2005:VES

[OH05] Ron Oglesby and Scott Herold. VMware ESX Server: Advanced Technical Design Guide. Brian Madden Publishing, Silver Spring, MD, USA, 2005. ISBN 0-9711510-6-7. 444 pp. LCCN QA76.9.V5 O35 2005eb; **See. URL http://

library.books24x7.com/library.asp?B&bookid=12048.

Oi:2005:DLV

[Oi05] Hitoshi Oi. On the design of the local variable cache in a hardware translation-based Java Virtual Machine. ACM SIG-PLAN Notices, 40(7):87–94, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oi:2006:IFH

[Oi06] Hitoshi Oi. Instruction folding in a hardware-translation based Java Virtual Machine. In ACM [ACM06e], pages 139– 146. ISBN 1-59593-302-6. LCCN ???? ACM order number 104060.

Oi:2008:LVA

[Oi08] Hitoshi Oi. Local variable access behavior of a hardwaretranslation based Java virtual machine. The Journal of Systems and Software, 81(11):2059–2068, November 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Osisek:1991:EIA

[OJG91] D. L. Osisek, K. M. Jackson, and P. H. Gum. ESA/390 interpretive-execution architecture, foundation for VM/ESA. IBM Systems Journal, 30(1):34–51, 1991. CODEN IBMSA7. ISSN 0018-8670.

Ozgur:1990:SON

[OK90] A. Ozgur and O. Kaynak. A self organizing neural network for pattern recognition. In INNC 90 Paris. International Neural Network Conference, volume 1, page 107. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0831-X.

Ogino:2017:VNE

[OKAM17] Nagao Ogino, Takeshi Kitahara, Shin’ichi Arakawa, and Masayuki Murata. Virtual network embedding with multiple priority classes sharing substrate resources. Computer Networks (Amsterdam, Netherlands: 1999), 112(??):52–66, January 15, 2017. CODEN ???? ISSN 1389-1286 (print),

1872-7069 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1389128616303395.

Ouyang:2013:PTS

[OL13] Jiannan Ouyang and John R. Lange. Preemptable ticket spinlocks: improving consolidated performance in the cloud. ACM SIGPLAN Notices, 48(7):191–200, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Olbert:1978:ECP

[Olb78] Arthur G. Olbert. Extended control program support: VM/ 370: a hardware assist for the IBM Virtual Machine Facility/370. ACM SIGMICRO Newsletter, 9(3):8–25, September 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL

https://dl.acm.org/doi/10.1145/1096532.1096534.

Ouyang:2016:SUV

[OLZ16] Jiannan Ouyang, John R. Lange, and Haoqiang Zheng. Shoot4U: Using VMM assists to optimize TLB operations on preempted vCPUs. ACM SIGPLAN Notices, 51(7):17– 23, July 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oliveira:2015:ORE

[OMB^{+}15] Rodrigo R. Oliveira, Daniel S. Marcon, Leonardo R. Bays, Miguel C. Neves, Luciano P. Gaspary, Deep Medhi, and Marinho P. Barcellos. Opportunistic Resilience Embedding (ORE): Toward cost-efficient resilient virtual networks. Computer Networks (Amsterdam, Netherlands: 1999), 89(??):59– 77, 2015. CODEN ???? ISSN 1389-1286 (print), 1872-

7069 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1389128615002339.

Ortin:2009:EVM

[ORPS09] Francisco Ortin, Jose Manuel Redondo, and J. Baltasar García Perez-Schofield. Efficient virtual machine support of runtime structural reflection. Science of Computer Programming, 74(10):836–860, August 15, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Osborne:2001:PC

[Osb01] Hugh Osborne. The Postroom Computer. ACM Journal on Educational Resources in Computing (JERIC), 1(4):81–110, December 2001. CODEN ???? ISSN 1531-4278.

Omote:2015:IAE

[OSK15] Yushi Omote, Takahiro Shinagawa, and Kazuhiko Kato. Improving agility and elasticity in bare-metal clouds. ACM SIGPLAN Notices, 50(4):145–159, April 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ostrand:1994:PIS

[Ost94] Thomas Ostrand, editor. Proceedings of the 1994 International Symposium on Software Testing and Analysis (ISSTA): August 17–19, 1994, Seattle, Washington, USA, ACM SIG-SOFT Software Engineering Notes. ACM Press, New York, NY 10036, USA, 1994. CODEN SFENDP. ISBN 0-89791-683-2. ISSN 0163-5948. LCCN QA76.76.T48 I58 1994.

OConnor:1997:PJV

[OT97] J. Michael O’Connor and Marc Tremblay. Picojava-I — the Java Virtual Machine in hardware. IEEE Micro, 17(2):45– 53, March/April 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ott:2018:SDI

[Ott18] David E. Ott. Software defined infrastructure: Rethinking cybersecurity with a more capable toolset. Operating Systems Review, 52(1):129–133, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ost:2012:EAT

[OVI^{+}12] Luciano Ost, Sameer Varyani, Leandro Soares Indrusiak, Marcelo Mandelli, Gabriel Marchesan Almeida, Eduardo Wachter, Fernando Moraes, and Gilles Sassatelli. Enabling adaptive techniques in heterogeneous MPSoCs based on virtualization. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 5(3):17:1–17:??, October 2012. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Parziale:2008:ZVL

[P^{+}08] Lydia Parziale et al., editors. z/VM and Linux on IBM System z: the virtualization cookbook for RHEL 5.2. Number SG24-7492-00 in IBM redbooks. IBM Corporation, San Jose, CA, USA, 2008. ISBN 0-7384-3181-8. xvi + 250 pp. LCCN QA76.76.O63 Z867 2008. URL http://proquest.

safaribooksonline.com/?fpi=0738431818.

Papaevripides:2021:EMB

[PA21] Michalis Papaevripides and Elias Athanasopoulos. Exploiting mixed binaries. ACM Transactions on Privacy and Security (TOPS), 24(2):7:1–7:29, February 2021. CODEN ???? ISSN 2471-2566 (print), 2471-2574 (electronic). URL https://dl.

acm.org/doi/10.1145/3418898.

Peake:2022:PVP

[PAC^{+}22] Joshua Peake, Martyn Amos, Nicholas Costen, Giovanni Masala, and Huw Lloyd. PACO–VMP: Parallel ant colony optimization for virtual machine placement. Future Generation Computer Systems, 129(??):174–186, April 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X21004568.

Park:2016:VCB

[PAKY16] Joonseok Park, Youngmin An, Taejun Kang, and Keunhyuk Yeom. Virtual cloud bank: consumer-centric service recommendation process and architectural perspective for cloud service brokers. Computing: Archiv für Informatik und Numerik, 98(11):1153–1184, November 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Papavassiliou:2020:SDN

[Pap20] Symeon Papavassiliou. Software defined networking (SDN) and network function virtualization (NFV). Future Internet, 12(1):7, January 02, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.com/1999-5903/12/1/7.

Parmelee:1971:VMS

[Par71] R. P. Parmelee. Virtual machines — some unexpected applications. In ????, editor, Proceedings IEEE Computer Society Conference, Boston, MA, September 1971, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1971.

Parmelee:1972:PVM

[Par72] R. P. Parmelee. Preferred virtual machines for CP-67. Report G320-2068, IBM Corporation, Cambridge Scientific Center, Cambridge, MA, USA, ???? 1972.

Parnas:1979:DSE

[Par79] D. L. Parnas. Designing software for ease of expansion and contraction. IEEE Transactions on Software Engineering, SE-5(2):128–138, March 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Parsafar:2025:RLB

[Par25] Parsa Parsafar. A reinforcement learning-based GWO–RNN approach for energy efficiency in data centers by minimizing virtual machine migration. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https://link.

springer.com/article/10.1007/s11227-024-06510-1.

Patel:2012:PIF

[Pat12] Naresh M. Patel. Performance implications of flash and storage class memories. ACM SIGMETRICS Performance Evaluation Review, 40(1):1–2, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pimas:2017:GCE

[PBAM17] Javier Pimás, Javier Burroni, Jean Baptiste Arnaud, and Stefan Marr. Garbage collection and efficiency in dynamic metacircular runtimes: an experience report. ACM SIG-PLAN Notices, 52(11):39–50, November 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pek:2013:SSI

[PBB13] Gábor Pék, Levente Butty’an, and Boldizsár Bencsáth. A survey of security issues in hardware virtualization. ACM Computing Surveys, 45(3):40:1–40:??, June 2013. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Plotkin:2016:SNV

[PBL^{+}16] Gordon D. Plotkin, Nikolaj Bjørner, Nuno P. Lopes, Andrey Rybalchenko, and George Varghese. Scaling network verification using symmetry and surgery. ACM SIGPLAN Notices, 51(1):69–83, January 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Plata:1990:ASP

[PBR^{+}90] O. G. Plata, J. D. Bruguera, F. F. Rivera, R. Doallo, and E. L. Zapata. ACLE: a software package for SIMD computer simulation. The Computer Journal, 33(3):194–203, June 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_33/Issue_03/tiff/194.
tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_33/Issue_03/tiff/195.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_03/tiff/
196.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/197.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_03/
tiff/198.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/199.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_03/
tiff/200.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/201.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_03/
tiff/202.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_03/tiff/203.tif.

Porter:2012:RLT

[PBWH^{+}12] Donald E. Porter, Silas Boyd-Wickizer, Jon Howell, Reuben Olinsky, and Galen C. Hunt. Rethinking the library OS from the top down. ACM SIGPLAN Notices, 47(4):291–304, April 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pelleg:2008:VBD

[PBYH^{+}08] Dan Pelleg, Muli Ben-Yehuda, Rick Harper, Lisa Spainhower, and Tokunbo Adeshiyan. Vigilant: out-of-band detection of failures in virtual machines. Operating Systems Review, 42 (1):26–31, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Padhy:2021:MCA

[PC21] Satyajit Padhy and Jerry Chou. MIRAGE: a consolidation aware migration avoidance genetic job scheduling algorithm for virtualized data centers. Journal of Parallel and Distributed Computing, 154(??):106–118, August 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000575.

Papazis:2024:HFM

[PC24] Kon Papazis and Naveen Chilamkurti. Hydrakon, a framework for measuring indicators of deception in emulated monitoring systems. Future Internet, 16(12):455, December 4, 2024. CODEN ???? ISSN 1999-5903. URL https://www.

mdpi.com/1999-5903/16/12/455.

Pickartz:2018:PCV

[PCB^{+}18] Simon Pickartz, Carsten Clauss, Jens Breitbart, Stefan Lankes, and Antonello Monti. Prospects and challenges of virtual machine migration in HPC. Concurrency and Computation: Practice and Experience, 30(9), May 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1002/cpe.4412.

Piraghaj:2016:VMC

[PCC^{+}16] Sareh Fotuhi Piraghaj, Rodrigo N. Calheiros, Jeffrey Chan, Amir Vahid Dastjerdi, and Rajkumar Buyya. Virtual machine customization and task mapping architecture for efficient allocation of cloud data center resources. The Computer Journal, 59(2):208–224, February 2016. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/59/2/208.

Perez-Cazares:1989:DAL

[PCR89] J.-R. Perez-Cazares and R. Rannou. Developing an animated local network simulator. Computer Graphics Forum, 8(2): 139–142, June 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peng:2016:TCT

[PCW^{+}16] Yang Peng, Kai Chen, Guohui Wang, Wei Bai, Yangming Zhao, Hao Wang, Yanhui Geng, Zhiqiang Ma, and Lin Gu. Towards comprehensive traffic forecasting in cloud computing: Design and application. IEEE/ACM Transactions on Networking, 24(4):2210–2222, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Pan:2012:CLM

[PDC^{+}12] Zhenhao Pan, Yaozu Dong, Yu Chen, Lei Zhang, and Zhijiao Zhang. CompSC: live migration with pass-through devices. ACM SIGPLAN Notices, 47(7):109–120, July 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Parra:2023:TMV

[PDL^{+}23] Pablo Parra, Antonio Da Silva, Borja Losa, J. Ignacio García, Óscar R. Polo, Agustín Martínez, and Sebastián Sánchez. Tailor-made virtualization monitor design for CPU virtualization on LEON processors. ACM Transactions on Embedded Computing Systems, 22(4):60:1–60:??, July 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

URL https://dl.acm.org/doi/10.1145/3584702.

Patil:2020:DVA

[PDM20] Rajendra Patil, Harsha Dudeja, and Chirag Modi. Designing in-VM-assisted lightweight agent-based malware detection framework for securing virtual machines in cloud computing. International Journal of Information Security, 19(2): 147–162, April 2020. CODEN ???? ISSN 1615-5262 (print),

1615-5270 (electronic). URL http://link.springer.com/
article/10.1007/s10207-019-00447-w.

Pizzi:2025:VDA

[PDML25] Simeone Pizzi, Samuele Doria, Nicholas Miazzo, and Eleonora Losiouk. VirtualPatch: Distributing Android security patches through Android virtualization. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404825003049.

Peng:2023:FPS

[PDY^{+}23] Bo Peng, Yaozu Dong, Jianguo Yao, Fengguang Wu, and Haibing Guan. FlexHM: a practical system for heterogeneous memory with flexible and efficient performance optimizations. ACM Transactions on Architecture and Code Optimization, 20(1):13:1–13:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https://dl.acm.org/

doi/10.1145/3565885.

Pham:2014:BRS

[PEC^{+}14] Cuong Pham, Zachary J. Estrada, Phuong Cao, Zbigniew Kalbarczyk, and Ravishankar K. Iyer. Building reliable and secure Virtual Machines using architectural invariants. IEEE Security & Privacy, 12(5):82–85, September/October 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (elec-

tronic). URL http://www.computer.org/csdl/mags/sp/
2014/05/msp2014050082-abs.html.

Park:2011:FSE

[PEL11] Eunbyung Park, Bernhard Egger, and Jaejin Lee. Fast and space-efficient virtual machine checkpointing. ACM SIG-PLAN Notices, 46(7):75–86, July 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Picano:2023:CAF

[PF23] Benedetta Picano and Romano Fantacci. A channel-aware FL approach for virtual machine placement in 6G edge intelligent ecosystems. ACM Transactions on Internet of Things (TIOT), 4(2):12:1–12:??, May 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL https://dl.

acm.org/doi/10.1145/3584705.

Pape:2016:LIS

[PFH^{+}16] Tobias Pape, Tim Felgentreff, Robert Hirschfeld, Anton Gulenko, and Carl Friedrich Bolz. Language-independent storage strategies for tracing JIT-based virtual machines. ACM SIGPLAN Notices, 51(2):104–113, February 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pham:2020:MAR

[PFNC20] Tuan-Minh Pham, Serge Fdida, Thi-Thuy-Lien Nguyen, and Hoai-Nam Chu. Modeling and analysis of robust service composition for network functions virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):Article 106989, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1389128619305080.

Pfoh:2013:LDV

[Pfo13] Jonas Pfoh. Leveraging Derivative Virtual Machine Introspection Methods for Security Applications. Doktors der Naturwissenschaften (Dr. rer. nat.), Fakultät für Informatik der Technischen Universität München, Lehrstuhl für Sicherheit in der Informatik, Munich, Germany, 2013.

xviii + 117 pp. URL http://mediatum.ub.tum.de/
?id=1115811;   http://mediatum.ub.tum.de/download/
1115811/1115811.pdf;   http://nbn-resolving.de/urn/
resolver.pl?urn:nbn:de:bvb:91-diss-20130318-1115811-
0-5.

Paulraj:2018:RAV

[PFPJ18] Getzi Jeba Leelipushpam Paulraj, Sharmila Anand John Francis, J. Dinesh Peter, and Immanuel Johnraja Jebadurai. Resource-aware virtual machine migration in IoT cloud. Future Generation Computer Systems, 85(??):173–183, August 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167739X17322471.

Popek:1973:FRV

[PG73] Gerald J. Popek and Robert P. Goldberg. Formal requirements for virtualizable third generation architectures. Operating Systems Review, 7(4):121, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Popek:1974:FRV

[PG74] Gerald J. Popek and Robert P. Goldberg. Formal requirements for virtualizable third generation architectures. Communications of the ACM, 17(7):412–421, July 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Payer:2011:FGU

[PG11] Mathias Payer and Thomas R. Gross. Fine-grained user-space security through virtualization. ACM SIGPLAN Notices, 46 (7):157–168, July 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Psychas:2017:NPV

[PG17] Konstantinos Psychas and Javad Ghaderi. On nonpreemptive VM scheduling in the cloud. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):35:1–35:29, December 2017. CODEN ????

ISSN 2476-1249. URL https://dl.acm.org/doi/10.1145/
3154493.

Psychas:2018:NPV

[PG18] Konstantinos Psychas and Javad Ghaderi. On nonpreemptive VM scheduling in the cloud. ACM SIGMET-RICS Performance Evaluation Review, 46(1):67–69, June 2018. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Pavlou:2012:DBD

[PGLG12] Demos Pavlou, Enric Gibert, Fernando Latorre, and Antonio Gonzalez. DDGacc: boosting dynamic DDG-based binary optimizations through specialized hardware support. ACM SIGPLAN Notices, 47(7):159–168, July 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Park:2019:BGM

[PGP19] Younghun Park, Minwoo Gu, and Sungyong Park. Ballooning graphics memory space in full GPU virtualization environments. Scientific Programming, 2019(1):5240956:1– 5240956:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1155/2019/5240956.

Pham:2020:CAE

[PHC20] Minh Pham, Doan B. Hoang, and Zenon Chaczko. Congestionaware and energy-aware virtual network embedding. IEEE/ ACM Transactions on Networking, 28(1):210–223, February 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566

(electronic). URL https://dl.acm.org/doi/abs/10.1109/
TNET.2019.2958367.

Papadimitriou:2012:TLS

[PHL^{+}12] Panagiotis Papadimitriou, Ines Houidi, Wajdi Louati, Djamal Zeghlache, and Christoph Werle. Towards large-scale network virtualization. Lecture Notes in Computer Science, 7277:13–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-30630-3_2/.

Pei:2019:EES

[PHXL19] J. Pei, P. Hong, K. Xue, and D. Li. Efficiently embedding service function chains with dynamic virtual network function placement in geo-distributed cloud system. IEEE Transactions on Parallel and Distributed Systems, 30(10):2179–2192, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pizlo:2017:JVM

[Piz17] Filip Pizlo. The JavaScriptCore virtual machine (invited talk). ACM SIGPLAN Notices, 52(11):1, November 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pahl:2018:APC

[PJZ18] Claus Pahl, Pooyan Jamshidi, and Olaf Zimmermann. Architectural principles for cloud software. ACM Transactions on Internet Technology (TOIT), 18(2):17:1–17:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Pfitscher:2019:GPA

[PJZ^{+}19] Ricardo José Pfitscher, Arthur Selle Jacobs, Luciano Zembruzki, Ricardo Luis dos Santos, Eder John Scheid, Muriel Figueredo Franco, Alberto Schaeffer-Filho, and Lisandro Zambenedetti Granville. Guiltiness: a practical approach for quantifying virtual network functions performance. Computer Networks (Amsterdam, Netherlands: 1999), 161(??):14–31, October 9, 2019. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1389128618310508.

Popek:1975:PVM

[PK75a] Gerald J. Popek and Charles S. Kline. The PDP-11 virtual machine architecture: a case study. In ACM [ACM75], pages 97–105. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Popek:1975:VPS

[PK75b] Gerald J. Popek and Charles S. Kline. A verifiable protection system. ACM SIGPLAN Notices, 10(6):294–304, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pal:2019:NPR

[PKS^{+}19] Souvik Pal, Raghvendra Kumar, Le Hoang Son, Krishnan Saravanan, Mohamed Abdel-Basset, Gunasekaran Manogaran, and Pham Huy Thong. Novel probabilistic resource migration algorithm for cross-cloud live migration of virtual machines in public cloud. The Journal of Supercomputing, 75(9): 5848–5865, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Pascual:2018:ERV

[PLMA18] Jose A. Pascual, Jose A. Lozano, and Jose Miguel-Alonso. Effects of reducing VMs management times on elastic applications. Journal of Grid Computing, 16(3):513–530, September 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL https://link.springer.com/article/
10.1007/s10723-018-9441-7.

Phung:2020:LPM

[PLZ20] J. Phung, Y. C. Lee, and A. Y. Zomaya. Lightweight power monitoring framework for virtualized computing environments. IEEE Transactions on Computers, 69(1):14–25, January 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Patil:2019:DEF

[PM19a] Rajendra Patil and Chirag Modi. Designing an efficient framework for vulnerability assessment and patching (VAP) in virtual environment of cloud computing. The Journal of Supercomputing, 75(5):2862–2889, May 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Patil:2019:ESS

[PM19b] Rajendra Patil and Chirag Modi. An exhaustive survey on security concerns and solutions at different components of virtualization. ACM Computing Surveys, 52(1):12:1– 12:??, February 2019. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3287306.

Parson:2005:OOD

[PMC05] Dale Parson, David J. Murray, and Yu Chen. Object-oriented design patterns for debugging heterogeneous languages and virtual machines. Software—Practice and Experience, 35(3): 255–279, March 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Perez:2023:AVB

[PMP23] Alberto García Pérez, Antonio López Martínez, and Manuel Gil Pérez. Adaptive vulnerability-based risk identification software with virtualization functions for dynamic management. Journal of Network and Computer Applications, 219(??):??, October 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804523001479.

Perez:2020:OPN

[PNM^{+}20] Taciano D. Perez, Marcelo V. Neves, Diego Medaglia, Pedro H. G. Monteiro, and César A. F. De Rose. Orthogonal persistence in nonvolatile memory architectures: a persistent heap design and its implementation for a Java Virtual Machine. Software—Practice and Experience, 50(4):368–387, April 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Petrides:2012:HPD

[PNT12] Panayiotis Petrides, George Nicolaides, and Pedro Trancoso. HPC performance domains on multi-core processors with virtualization. Lecture Notes in Computer Science, 7179: 123–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-28293-5_11/.

Picht:2009:XKI

[PO09] Hans-Joachim Picht and Sebastian Ott. Xen Kochbuch: intelligente Virtualisierungslösungen mit XEN 3. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 3-89721-729-5. xix + 461 pp. LCCN ???? EUR 44.90. URL

http://www.gbv.de/dms/ilmenau/toc/572426836.PDF.

Ponraj:2019:OVM

[Pon19] Anitha Ponraj. Optimistic virtual machine placement in cloud data centers using queuing approach. Future Generation Computer Systems, 93(??):338–344, April 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167739X18304692.

Pountain:1990:SPP

[Pou90] D. Pountain. Simulating a primitive parallel computer. Byte Magazine, 15(13):72IS/25–26, 28, 30, 32, 34, 36–37, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Parnas:1973:DVM

[PP73] D. L. Parnas and W. R. Price. he design of the virtual memory aspects of a virtual machine. In ????, editor, Proceedings ACM SIGARCH-SIGOPS Workshop on Virtual Computer Systems, Cambridge, MA, 1973, page ?? ACM Press, New York, NY 10036, USA, 1973.

Paulo:2016:EDD

[PP16] João Paulo and José Pereira. Efficient deduplication in a distributed primary storage infrastructure. ACM Transactions on Storage, 12(4):20:1–20:??, August 2016. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Pinto:2017:TTA

[PPG^{+}17] Sandro Pinto, Jorge Pereira, Tiago Gomes, Mongkol Ekpanyapong, and Adriano Tavares. Towards a TrustZone-assisted hypervisor for real-time embedded systems. IEEE Computer Architecture Letters, 16(2):158–161, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Pfitscher:2014:COD

[PPO14] Ricardo J. Pfitscher, Mauricio A. Pillon, and Rafael R. Obelheiro. Customer-oriented diagnosis of memory provisioning for IaaS clouds. Operating Systems Review, 48(1):2–10, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pettit:2018:BPH

[PPS^{+}18] Justin Pettit, Ben Pfaff, Joe Stringer, Cheng-Chun Tu, Brenden Blanco, and Alex Tessmer. Bringing platform harmony to VMware NSX. Operating Systems Review, 52(1):123–128, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Parmelee:1972:VSV

[PPTH72] Richard P. Parmelee, Theodore I. Peterson, Coyt C. Tillman, Jr., and Donald J. Hatfield. Virtual storage and virtual machine concepts. IBM Systems Journal, 11(2):99–130, 1972. CODEN IBMSA7. ISSN 0018-8670. URL https://dlnext.

acm.org/doi/10.1147/sj.112.0099.

Permandla:2007:TSP

[PRB07] Pratibha Permandla, Michael Roberson, and Chandrasekhar Boyapati. A type system for preventing data races and deadlocks in the Java Virtual Machine language: 1. ACM SIG-PLAN Notices, 42(7):10, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Provos:2000:EVM

[Pro00] Niels Provos. Encrypting virtual machine. In USENIX [USE00b], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.usenix.org/publications/
library/proceedings/sec2000/provos.html.

Prades:2016:CAX

[PRS16] Javier Prades, Carlos Reaño, and Federico Silla. CUDA acceleration for Xen virtual machines in InfiniBand clusters with rCUDA. ACM SIGPLAN Notices, 51(8):35:1–35:??, August 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prades:2026:EPG

[PRS26] Javier Prades, Carlos Reaño, and Federico Silla. Enhancing the performance of GPU acceleration in virtual environments: Thoroughly benchmarking the rigidity of mediated device passthrough. Future Generation Computer Systems, 174(??): ??, January 2026. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167739X25002390.

Pietri:2016:MVM

[PS16] Ilia Pietri and Rizos Sakellariou. Mapping virtual machines onto physical machines in cloud computing: a survey. ACM Computing Surveys, 49(3):49:1–49:??, November 2016. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Pinto:2019:DAT

[PS19a] Sandro Pinto and Nuno Santos. Demystifying Arm Trust-Zone: a comprehensive survey. ACM Computing Surveys, 51 (6):130:1–130:??, February 2019. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=3291047.

Prades:2019:GJM

[PS19b] J. Prades and F. Silla. GPU-job migration: The rCUDA case. IEEE Transactions on Parallel and Distributed Systems, 30 (12):2718–2729, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pushpa:2023:FAB

[PS23] Ramaiah Pushpa and Maadappa Siddappa. Fractional artificial bee chicken swarm optimization technique for QoS aware virtual machine placement in cloud. Concurrency and Computation: Practice and Experience, 35(4):e7532:1–e7532:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Parri:2011:RCPa

[PSBG11a] Jonathan Parri, Daniel Shapiro, Miodrag Bolic, and Voicu Groza. Returning control to the programmer: SIMD intrinsics for virtual machines. ACM Queue: Tomorrow’s Computing Today, 9(2):30, February 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Parri:2011:RCPb

[PSBG11b] Jonathan Parri, Daniel Shapiro, Miodrag Bolic, and Voicu Groza. Returning control to the programmer: SIMD intrinsics for virtual machines. Communications of the ACM, 54 (4):38–43, April 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Payne:2007:LAS

[PSC^{+}07] Bryan D. Payne, Reiner Sailer, Ramón Cáceres, Ron Perez, and Wenke Lee. A layered approach to simplified access control in virtualized systems. Operating Systems Review, 41(4): 12–19, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pfefferle:2015:HVF

[PST^{+}15a] Jonas Pfefferle, Patrick Stuedi, Animesh Trivedi, Bernard Metzler, Ionnis Koltsidas, and Thomas R. Gross. A hybrid I/O virtualization framework for RDMA-capable network interfaces. ACM SIGPLAN Notices, 50(7):17–30, July 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pignolet:2015:ATD

[PST15b] Yvonne Anne Pignolet, Stefan Schmid, and Gilles Tredan. Adversarial topology discovery in network virtualization environments: a threat for ISPs? Distributed Computing, 28(2):91–109, April 2015. CODEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (electronic). URL https://link.

springer.com/article/10.1007/s00446-014-0217-4.

Padala:2007:ACV

[PSZ^{+}07] Pradeep Padala, Kang G. Shin, Xiaoyun Zhu, Mustafa Uysal, Zhikui Wang, Sharad Singhal, Arif Merchant, and Kenneth Salem. Adaptive control of virtualized resources in utility computing environments. Operating Systems Review, 41(3): 289–302, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pease:2018:IRT

[PTD^{+}18] Sarogini Grace Pease, Russell Trueman, Callum Davies, Jude Grosberg, Kai Hin Yau, Navjot Kaur, Paul Conway, and Andrew West. An intelligent real-time cyber-physical toolset for energy and process prediction and optimisation in the future industrial Internet of Things. Future Generation Computer Systems, 79 (part 3)(??):815–829, February 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X1630382X.

Pape:2014:EJV

[PTHH14] Tobias Pape, Arian Treffer, Robert Hirschfeld, and Michael Haupt. Extending a Java Virtual Machine to Dynamic Object-oriented Languages, volume 82 of Technische Berichte des Hasso-Plattner-Instituts für Softwaresystemtechnik an der Universität Potsdam. Universitätsverlag Potsdam, Potsdam, Germany, 2014. ISBN 3-86956-266-8. 163 pp. LCCN

???? URL http://d-nb.info/1046379119/04;   http://
opus.kobv.de/ubp/volltexte/2013/6743/.

Pham:2015:SRD

[PTM^{+}15] Quan Pham, Severin Thaler, Tanu Malik, Ian Foster, and Boris Glavic. Sharing and reproducing database applications. Proceedings of the VLDB Endowment, 8(12):1988–1991, August 2015. CODEN VLDBFR. ISSN 2150-8097.

Pulman:1991:EER

[Pul91] Stephen G. Pulman, editor. EUROTRA ET6/1: rule formalism and virtual machine design study – final report. Commission of the European Communities, Luxembourg, Luxembourg, 1991. Contributors: H. Alshawi, D.J. Arnold, R. Backofen, D.M. Carter, J. Lindop, K. Netter, S.G. Pulman, J. Tsujii and H. Uszkoreit.

Petrashko:2016:CGL

[PULO16] Dmitry Petrashko, Vlad Ureche, Ondrej Lhoták, and Martin Odersky. Call graphs for languages with parametric polymorphism. ACM SIGPLAN Notices, 51(10):394–409, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pickett:2006:SSF

[PV06] Christopher J. F. Pickett and Clark Verbrugge. SableSpMT: a software framework for analysing speculative multithreading in Java. ACM SIGSOFT Software Engineering Notes, 31(1): 59–66, January 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Prokopski:2008:APC

[PV08] Gregory B. Prokopski and Clark Verbrugge. Analyzing the performance of code-copying virtual machines. ACM SIG-PLAN Notices, 43(10):403–422, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perez:2008:VHB

[PvDS08] Ronald Perez, Leendert van Doorn, and Reiner Sailer. Virtualization and hardware-based security. IEEE Security & Privacy, 6(5):24–31, September/October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pawlish:2014:CEE

[PVRR14] Michael Pawlish, Aparna S. Varde, Stefan A. Robila, and Anand Ranganathan. A call for energy efficiency in data centers. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(1):45–51, March 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Panesar-Walawege:2003:VHM

[PW03] Rajwinder Kaur Panesar-Walawege. Views: a platformindependent GUI toolkit for the shared-source Common Language Infrastructure. Thesis (M. Sc.), University of Victoria, Victoria, BC, Canada, 2003.

Peng:2016:RTE

[PWJ16] Yaqiong Peng, Song Wu, and Hai Jin. Robinhood: Towards efficient work-stealing in virtualized environments. IEEE Transactions on Parallel and Distributed Systems, 27(8):2363–2376, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/csdl.computer.org/csdl/trans/td/2016/08/07300465-
abs.html.

Peng:2017:SMA

[PXG^{+}17] Zhiping Peng, Bo Xu, Antonio Marcel Gates, Delong Cui, and Weiwei Lin. A study of a multi-agent organizational framework with virtual machine clusters as the unit of granularity in cloud computing. The Computer Journal, 60(7): 1032–1043, July 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://academic.

oup.com/comjnl/article/60/7/1032/2608048.

Poulsen:1993:ETP

[PY93] David K. Poulsen and Pen-Chung Yew. Execution-driven tools for parallel simulation of parallel architectures and applications. In IEEE [IEE93b], pages 860–869. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Peng:2022:MNM

[PYDG22] Bo Peng, Jianguo Yao, Yaozu Dong, and Haibing Guan. MDev-NVMe: Mediated pass-through NVMe virtualization solution with adaptive polling. IEEE Transactions on Computers, 71(2):251–265, February 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Peng:2021:TON

[PYYG21] Bo Peng, Ming Yang, Jianguo Yao, and Haibing Guan. A throughput-oriented NVMe storage virtualization with workload-aware management. IEEE Transactions on Computers, 70(12):2112–2124, December 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pearce:2013:VIS

[PZH13] Michael Pearce, Sherali Zeadally, and Ray Hunt. Virtualization: Issues, security threats, and solutions. ACM Computing Surveys, 45(2):17:1–17:??, February 2013. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Padala:2007:PEV

[PZW^{+}07] Pradeep Padala, Xiaoyun Zhu, Zhikui Wang, Sharad Singhal, and Kang G. Shin. Performance evaluation of virtualization technologies for server consolidation. Technical Report HPL-2007-59, Enterprise Systems and Software Laboratory, HP Laboratories, Palo Alto, CA, USA, April 11,

2007. URL http://www.hpl.hp.com/techreports/2007/
HPL-2007-59.pdf.

Qiu:2023:VNF

[QBL^{+}23] Rixuan Qiu, Jiawen Bao, Yuancheng Li, Xin Zhou, Liang Liang, Hui Tian, Yanting Zeng, and Jie Shi. Virtual network function deployment algorithm based on graph convolution deep reinforcement learning. The Journal of Supercomputing, 79(6):6849–6870, April 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https://link.

springer.com/article/10.1007/s11227-022-04947-w.

Queiroz:2024:CBV

[QCM^{+}24] Rui Queiroz, Tiago Cruz, Jérôme Mendes, Pedro Sousa, and Paulo Simões. Container-based virtualization for real-time industrial systems — a systematic review. ACM Computing Surveys, 56(3):59:1–59:??, March 2024. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https:

//dl.acm.org/doi/10.1145/3617591.

Qian:1999:FSJ

[Qia99] Z. Qian. A formal specification of Java[TM] virtual machine instructions for objects, methods and subroutines. Lecture Notes in Computer Science, 1523:271–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qiao:2021:DSR

[QLL^{+}21] Wenxin Qiao, Hao Lu, Yu Lu, Lijie Meng, and Yicen Liu. A dynamic service reconfiguration method for satelliteterrestrial integrated networks. Future Internet, 13(10):260, October 09, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/1999-5903/13/10/260.

Quetier:2007:SCF

[QNC07] Benjamin Quétier, Vincent Neri, and Franck Cappello. Scalability comparison of four host virtualization tools. Journal of Grid Computing, 5(1):83–98, March 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&volume=5&issue=1&
spage=83.

Quynh:2006:RTI

[QT06] Nguyen Anh Quynh and Yoshiyasu Takefuji. A real-time integrity monitor for Xen virtual machine. In IEEE [IEE06b], pages 90–?? ISBN 0-7695-2622-5. LCCN ????

Quraishi:2021:SSA

[QTR21] M. H. Quraishi, E. B. Tavakoli, and F. Ren. A survey of system architectures and techniques for FPGA virtualization. IEEE Transactions on Parallel and Distributed Systems, 32 (9):2216–2230, September 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qu:2018:IEE

[QXH18] Xilong Qu, Peng Xiao, and Lirong Huang. Improving the energy efficiency and performance of data-intensive workflows in virtualized clouds. The Journal of Supercomputing, 74(7): 2935–2955, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Qiang:2016:SCF

[QZDJ16] Weizhong Qiang, Kang Zhang, Weiqi Dai, and Hai Jin. Secure cryptographic functions via virtualization-based outsourced computing. Concurrency and Computation: Practice and Experience, 28(11):3149–3163, August 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Russell:2002:SCI

[R^{+}02] Steve Russell et al., editors. Server consolidation with the IBM eServer xSeries 440 and VMware ESX Server. IBM redbooks. IBM, International Technical Support Organization, ????, November 2002. ISBN 0-7384-2684-9. xiv + 222 pp. LCCN QA76.9.C55 S378 2002. Publication number SG24-6852-00.

ReFerre:2006:VIS

[R^{+}06] Massimo Re Ferre’ et al., editors. Virtualization on the IBM System x3950 Server. IBM redbooks; IBM eserver. IBM Corporation, San Jose, CA, USA, 2006. ISBN 0-7384-9709-6 (paperback). xii + 272 pp. LCCN

QA76.9.V5 V578 2006. URL http://www.loc.gov/catdir/
toc/fy0706/2006284881.html.

Rayns:2013:CJS

[R^{+}13] Chris Rayns et al. CICS and the JVM server developing and deploying Java applications. IBM redbooks. IBM Corporation, International Technical Support Organization, Poughkeepsie, NY, USA, 2013. ISBN 0-7384-3833-

2. ???? pp. LCCN ???? URL http://proquest.tech.
safaribooksonline.de/0738438332.

Rajaraman:1979:PPV

[Raj79] M. K. Rajaraman. Performance prediction of a virtual machine. ACM SIGMETRICS Performance Evaluation Review, 8(1–2):57–62, Spring-Summer 1979. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ramsdell:1993:RVP

[Ram93] John D. Ramsdell. The revised VLISP prescheme front end. Technical report, MITRE, 1993. 91 pp. URL ftp:

//cs.indiana.edu/pub/scheme-repository/txt/vlisp/
preschemerevised.dvi.Z.

Raner:2002:LJV

[Ran02] Mirko Raner. A lightweight Java Virtual Machine for a stackbased microprocessor. In USENIX [USE02], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38 J42 2002. URL http:

//www.usenix.org/publications/library/proceedings/
javavm02/raner.html.

Randal:2020:IVR

[Ran20] Allison Randal. The ideal versus the real: Revisiting the history of virtual machines and containers. ACM Computing Surveys, 53(1):5:1–5:31, May 2020. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3365199.

Rathinaraja:2019:DRB

[RAP19] J. Rathinaraja, V. S. Ananthanarayana, and Anand Paul. Dynamic ranking-based MapReduce job scheduler to exploit heterogeneous performance in a virtualized environment. The Journal of Supercomputing, 75(11):7520–7549, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Richards:2017:VAK

[RAT17] Gregor Richards, Ellen Arteca, and Alexi Turcotte. The VM already knew that: leveraging compile-time knowledge to optimize gradual typing. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):55:1–55:??, October 2017. CODEN ???? ISSN 2475-1421.

Russell:2001:HSA

[RB01] Kenneth Russell and Lars Bak. The HotSpot^{T M} serviceability agent: An out-of-process high-level debugger for a Java^{T M} Virtual Machine. In USENIX [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.usenix.org/publications/
library/proceedings/jvm01/russell.html.

Rodriguez:2017:BDS

[RB17] Maria A. Rodriguez and Rajkumar Buyya. Budget-driven scheduling of scientific workflows in IaaS clouds with finegrained billing periods. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(2):5:1–5:??, May 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Rolon:2024:BMP

[RB24] Sebastián Rolón and Oana Balmau. Is bare-metal I/O performance with user-defined storage drives inside VMs possible? Benchmarking libvfio-user vs. common storage virtualization configurations. Operating Systems Review, 58(1):45– 52, June 2024. CODEN OSRED8. ISSN 0163-5980 (print),

1943-586X (electronic). URL https://dl.acm.org/doi/10.
1145/3689051.3689059.

R:2018:SDM

[RC18] Jithin R and Priya Chandran. Secure and dynamic memory management architecture for virtualization technologies in IoT devices. Future Internet, 10(12):119, November 30, 2018. CODEN ???? ISSN 1999-5903. URL https://www.

mdpi.com/1999-5903/10/12/119.

Ramakrishnan:2012:EIV

[RCM^{+}12] Lavanya Ramakrishnan, R. Shane Canon, Krishna Muriki, Iwona Sakrejda, and Nicholas J. Wright. Evaluating interconnect and virtualization performance for high performance computing. ACM SIGMETRICS Performance Evaluation Review, 40(2):55–60, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Rajagopalan:2012:SDT

[RCOW12] Shriram Rajagopalan, Brendan Cully, Ryan O’Connor, and Andrew Warfield. SecondSite: disaster tolerance as a service. ACM SIGPLAN Notices, 47(7):97–108, July 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Ruan:2019:VMA

[RCTY19] Xiaojun Ruan, Haiquan Chen, Yun Tian, and Shu Yin. Virtual machine allocation and migration based on performanceto-power ratio in energy-efficient clouds. Future Generation Computer Systems, 100(??):380–394, November 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X18321629.

Rounce:1990:AWE

[RD90] Peter A. Rounce and Jose Delgado. Architectures within the ESPRIT SPAN Project. IEEE Micro, 10(6):24–27, 88–97, November/December 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Renaud:1978:UVM

[Ren78] D. J. Renaud. The UT virtual machine monitor. Technical Report CS-TR-78-78, University of Texas at Austin, Department of Computer Sciences, Austin, TX, USA, May 1, 1978.

URL ftp://ftp.cs.utexas.edu/pub/techreports/tr78-
78a.pdf;   ftp://ftp.cs.utexas.edu/pub/techreports/

tr78-78b.pdf. Wed, 23 Jun 104 21:38:33 GMT.

Requet:2003:BME

[Req03] Antoine Requet. A B model for ensuring soundness of a large subset of the Java Card virtual machine. Science of Computer Programming, 46(3):283–306, March 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Revelle:2011:HVM

[Rev11] Don Revelle. Hypervisors and virtual machines: Implementation insights on the x86 architecture. ;login: the USENIX Association newsletter, 36(5):17–22, October 2011. CO-DEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/october-2011-volume-
36-number-5/hypervisors-and-virtual-machines-implementation.

Riehle:2001:AUV

[RFBLO01] Dirk Riehle, Steven Fraleigh, Dirk Bucka-Lassen, and Nosa Omorogbe. The architecture of a UML virtual machine. ACM SIGPLAN Notices, 36(11):327–341, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Rosenblum:2005:VMM

[RG05] Mendel Rosenblum and Tal Garfinkel. Virtual machine monitors: Current technology and future trends. Computer, 38(5): 39–??, May 2005. CODEN CPTRB4. ISSN 0018-9162 (print),

1558-0814 (electronic). URL http://csdl.computer.org/
comp/mags/co/2005/05/r5039abs.htm;   http://csdl.
computer.org/dl/mags/co/2005/05/r5039.pdf.

Rampersaud:2017:SAO

[RG17] Safraz Rampersaud and Daniel Grosu. Sharing-aware online virtual machine packing in heterogeneous resource clouds. IEEE Transactions on Parallel and Distributed Systems, 28(7):2046–2059, July 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:

//www.computer.org/csdl/trans/td/2017/07/07792170-
abs.html.

Raju:2019:STB

[RG19] BKSP Kumar Raju and G. Geethakumari. SNAPS: Towards building snapshot based provenance system for virtual machines in the cloud environment. Computers & Security, 86(??):92–111, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404818301895.

Rahmanian:2018:LAB

[RGAT18] Ali Asghar Rahmanian, Mostafa Ghobaei-Arani, and Sajjad Tofighy. A learning automata-based ensemble resource usage prediction algorithm for cloud computing environment. Future Generation Computer Systems, 79 (part 1)(??):54–71, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167739X17309378.

Rosendo:2020:AAD

[RGS^{+}20] Daniel Rosendo, Demis Gomes, Guto Leoni Santos, Leylane Silva, Andre Moreira, Judith Kelner, Djamel Sadok, Glauco Gonçalves, Amardeep Mehta, Mattias Wildeman, and Patricia Takako Endo. Availability analysis of design configurations to compose virtual performance-optimized data center systems in next-generation cloud data centers. Software— Practice and Experience, 50(6):805–826, June 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ryoo:2017:RTD

[RGSJ17] Jee Ho Ryoo, Nagendra Gulur, Shuang Song, and Lizy K. John. Rethinking TLB designs in virtualized environments: a very large part-of-memory TLB. ACM SIGARCH Computer Architecture News, 45(2):469–480, May 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Rajabzadeh:2017:EAF

[RH17] Mehdi Rajabzadeh and Abolfazl Toroghi Haghighat. Energyaware framework with Markov chain-based parallel simulated annealing algorithm for dynamic management of virtual machines in cloud data centers. The Journal of Supercomputing, 73(5):2001–2017, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Radonic:2008:XAK

[RHM08] Andrej Radonic, Thomas Halinka, and Frank Meyer. XEN 3.2: aufsetzen, konfigurieren, betreiben. Franzis Professional Series. Franzis-Verlag, Poing, Germany, second edition, 2008. ISBN 3-7723-7247-3. 499 pp. LCCN ????

Rajan:2002:CPJ

[RHR02] A. S. Rajan, Shiwen Hu, and J. Rubio. Cache performance in Java virtual machines: a study of constituent phases. In IEEE [IEE02], pages 81–90. ISBN 0-7803-7681-1. LCCN QA76.9.S88; QA76.9.S88 W67 2002eb; Inter-

net. URL http://ieeexplore.ieee.org/servlet/opac?

punumber=8689. IEEE Catalog Number 02EX633.

Rajabzadeh:2020:NCM

[RHR20] Mehdi Rajabzadeh, Abolfazl Toroghi Haghighat, and Amir Masoud Rahmani. New comprehensive model based on virtual clusters and absorbing Markov chains for energyefficient virtual machine management in cloud computing. The Journal of Supercomputing, 76(9):7438–7457, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL https://link.springer.com/article/
10.1007/s11227-020-03169-2.

Roychowdhury:2017:ABS

[RHV17] Sohini Roychowdhury, Paul Hage, and Joseph Vasquez. Azure-based smart monitoring system for anemia-like pallor. Future Internet, 9(3):39, July 26, 2017. CODEN ????

ISSN 1999-5903. URL https://www.mdpi.com/1999-5903/
9/3/39.

Ren:2017:DPA

[RHZ^{+}17] Shenyuan Ren, Ligang He, Huanzhou Zhu, Zhuoer Gu, Wei Song, and Jiandong Shang. Developing power-aware scheduling mechanisms for computing systems virtualized by Xen. Concurrency and Computation: Practice and Experience, 29 (3):??, February 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Robin:2000:AIP

[RI00] John Scott Robin and Cynthia E. Irvine. Analysis of the Intel Pentium’s ability to support a secure virtual machine monitor. In USENIX [USE00b], page ?? ISBN 1-880446-18-9.

LCCN ???? URL http://www.usenix.org/publications/
library/proceedings/sec2000/robin.html.

Roselin:2025:DDL

[RI25] J. Roselin and Israelin J. Insulata. Decentralized dynamic load balancing for virtual machines in cloud computing: a blockchain-enabled system with state channel optimization. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-025-06922-7.

Ricci:2018:LSC

[RIP18] Laura Ricci, Alexandru Iosup, and Radu Prodan. Large scale cooperative virtual environments. Concurrency and Computation: Practice and Experience, 30(20):e4878:1–e4878:??, October 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rixner:2008:NVB

[Rix08] Scot Rixner. Network virtualization: breaking the performance barrier. ACM Queue: Tomorrow’s Computing Today, 6(1):36–ff, January 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Robinson:2016:CCM

[RJK16] Forrest J. Robinson, Michael R. Jantz, and Prasad A. Kulkarni. Code cache management in managed language VMs to reduce memory consumption for embedded systems. ACM SIGPLAN Notices, 51(5):11–20, May 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roh:2017:JFV

[RJK^{+}17] Heejun Roh, Cheoulhoon Jung, Kyunghwi Kim, Sangheon Pack, and Wonjun Lee. Joint flow and virtual machine placement in hybrid cloud data centers. Journal of Network and Computer Applications, 85(??):4–13, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516303101.

Ruprecht:2018:VLM

[RJS^{+}18] Adam Ruprecht, Danny Jones, Dmitry Shiraev, Greg Harmon, Maya Spivak, Michael Krebs, Miche Baker-Harvey, and Tyler Sanderson. VM live migration at scale. ACM SIG-PLAN Notices, 53(3):45–56, March 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Radhakrishnan:2016:ECC

[RK16] A. Radhakrishnan and V. Kavitha. Energy conservation in cloud data centers by minimizing virtual machines migration through artificial neural network. Computing: Archiv für Informatik und Numerik, 98(11):1185–1202, November 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Riahi:2018:MOD

[RK18] Montassar Riahi and Saoussen Krichen. A multi-objective decision support framework for virtual machine placement in cloud data centers: a real case study. The Journal of Supercomputing, 74(7):2984–3015, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rjeib:2024:VEV

[RK24] Hasanein D. Rjeib and Gabor Kecskemeti. VMP-ER: an efficient virtual machine placement algorithm for energy and resources optimization in cloud data center. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/
7/295.

Rottenstreich:2017:MDN

[RKRK17] Ori Rottenstreich, Isaac Keslassy, Yoram Revah, and Aviran Kadosh. Minimizing delay in network function virtualization with shared pipelines. IEEE Transactions on Parallel and Distributed Systems, 28(1):156–169, January 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2017/01/07457328-abs.html.

Rahmani:2020:BAV

[RKT20] Somayeh Rahmani, Vahid Khajehvand, and Mohsen Torabian. Burstiness-aware virtual machine placement in cloud computing systems. The Journal of Supercomputing, 76(1): 362–387, January 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ren:2016:SMO

[RLZ^{+}16] Yi Ren, Ling Liu, Qi Zhang, Qingbo Wu, Jianbo Guan, Jinzhu Kong, Huadong Dai, and Lisong Shao. Sharedmemory optimizations for inter-virtual-machine communication. ACM Computing Surveys, 48(4):49:1–49:??, May 2016. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

ACM:2003:ATA

[RM03] Allyn Romanow and Jeff Mogul, editors. Proceedings of the ACM SIGCOMM Workshop on Network-I/O Convergence: experience, Lessons, Implications 2003, Karlsruhe, Germany, August 25–27, 2003. ACM Press, New York, NY 10036, USA, 2003. ISBN ???? LCCN TK5105.5. ACM order number 534032.

Roblitz:2002:LSE

[RMB02] Thomas Röblitz, Frank Mueller, and Oliver Bühn. LegoSim: simulation of embedded kernels over Pthreads. ACM Journal on Educational Resources in Computing (JERIC), 2(1):117– 130, March 2002. CODEN ???? ISSN 1531-4278.

Rozehkhani:2024:ECD

[RMP24] Seyyed Meysam Rozehkhani, Farnaz Mahan, and Witold Pedrycz. Efficient cloud data center: an adaptive framework for dynamic Virtual Machine Consolidation. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000626.

Rahman:2022:EET

[RNA^{+}22] Fatin Hamadah Rahman, S. H. Shah Newaz, Thien-Wan Au, Wida Susanty Suhaili, M. A. Parvez Mahmud, and Gyu Myoung Lee. EnTruVe: ENergy and TRUst-aware virtual machine allocation in VEhicle fog computing for catering applications in 5G. Future Generation Computer Systems, 126(??):196–210, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X21002983.

Ridolfi:2023:IEF

[RNS^{+}23] Lorenzo Ridolfi, David Naseh, Swapnil Sadashiv Shinde, , and Daniele Tarchi. Implementation and evaluation of a federated learning framework on raspberry PI platforms for IoT 6G applications. Future Internet, 15(11):358, October 31, 2023. CODEN ???? ISSN 1999-5903. URL https://www.mdpi.

com/1999-5903/15/11/358.

Ritson:2016:BWM

[RO16] Carl G. Ritson and Scott Owens. Benchmarking weak memory models. ACM SIGPLAN Notices, 51(8):24:1–24:??, August 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Robbins:2006:LGC

[Rob06] Stuart Robbins. Lessons in grid computing: the system is a mirror. Wiley, New York, NY, USA, 2006. ISBN 0-471-79010-9 (cloth). xviii + 363 pp. LCCN HD30.2 .R627

2006. URL http://www.loc.gov/catdir/enhancements/
fy0740/2006002910-b.html;   http://www.loc.gov/catdir/
enhancements/fy0740/2006002910-d.html;   http://www.
loc.gov/catdir/toc/ecip067/2006002910.html.

Robertson:2012:SVL

[Rob12] Brian Robertson. Security: virtualisation’s last frontier. Network Security, 2012(4):12–15, April 2012. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1353485812700260.

Rosenblum:1999:VVP

[Ros99] Mendel Rosenblum. VMware’s Virtual Platform: a virtual machine monitor for commodity PCs. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http://www.

hotchips.org/hotc11_index.html.

Rosenblum:2004:RVM

[Ros04] Mendel Rosenblum. The reincarnation of virtual machines. ACM Queue: Tomorrow’s Computing Today, 2(5):34–40, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Rosenblum:2006:IVC

[Ros06] Mendel Rosenblum. Impact of virtualization on computer architecture and operating systems. Operating Systems Review, 40(5):1, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rosen:2014:LCF

[Ros14] Rami Rosen. Linux containers and the future cloud. Linux Journal, 2014(240):3:1–3:??, April 2014. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Roussos:2007:SVG

[Rou07] Kostadis Roussos. Storage virtualization gets smart. ACM Queue: Tomorrow’s Computing Today, 5(6):38–44, September 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ramamurthy:2007:PDE

[RP07] Pratap Ramamurthy and Ramanathan Palaniappan. Performancedirected energy management using BOS. Operating Systems Review, 41(1):66–77, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ryckbosch:2012:VSM

[RPE12] Frederick Ryckbosch, Stijn Polfliet, and Lieven Eeckhout. VSim: Simulating multi-server setups at near native hardware speed. ACM Transactions on Architecture and Code Optimization, 8(4):52:1–52:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ren:2017:NLN

[RQD^{+}17] Jianbao Ren, Yong Qi, Yuehua Dai, Yu Xuan, and Yi Shi. Nosv: a lightweight nested-virtualization VMM for hosting high performance computing on cloud. The Journal of Systems and Software, 124(??):137–152, February 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121216302151.

Ruest:2009:VBG

[RR09] Danielle Ruest and Nelson Ruest. Virtualization: a beginner’s guide. Network professional’s library. McGraw-Hill, New York, NY, USA, 2009. ISBN 0-07-161401-X. xx + 442 pp. LCCN QA76.9.V5 R88 2009.

Rosa:2017:ARC

[RRB17] Andrea Rosà, Eduardo Rosales, and Walter Binder. Accurate reification of complete supertype information for dynamic analysis on the JVM. ACM SIGPLAN Notices, 52(12):104– 116, December 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosa:2019:AOT

[RRB19] Andrea Rosà, Eduardo Rosales, and Walter Binder. Analysis and optimization of task granularity on the Java Virtual Machine. ACM Transactions on Programming Languages and Systems, 41(3):19:1–19:??, July 2019. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL https:

//dl.acm.org/ft_gateway.cfm?id=3338497.

Reano:2016:TRG

[RS16] Carlos Reaño and Federico Silla. Tuning remote GPU virtualization for InfiniBand networks. The Journal of Supercomputing, 72(12):4520–4545, December 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rost:2020:HIV

[RS20] Matthias Rost and Stefan Schmid. On the hardness and inapproximability of virtual network embeddings. IEEE/ ACM Transactions on Networking, 28(2):791–803, April 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566

(electronic). URL https://dl.acm.org/doi/abs/10.1109/
TNET.2020.2975646.

Reano:2015:IUE

[RSC^{+}15] Carlos Reaño, Federico Silla, Adrián Castelló, Antonio J. Peña, Rafael Mayo, Enrique S. Quintana-Ortí, and José Duato. Improving the user experience of the rCUDA remote GPU virtualization framework. Concurrency and Computation: Practice and Experience, 27(14):3746–3770, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ragsdale:2003:CLI

[RSF03] Susann Ragsdale, David Sussman, and Mark Fussell. The Common Language Infrastructure Annotated Standard. Microsoft.NET development series. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-321-15493-2. 826 pp. LCCN ????

Rastogi:2015:SEG

[RSF^{+}15] Aseem Rastogi, Nikhil Swamy, Cédric Fournet, Gavin Bierman, and Panagiotis Vekris. Safe & efficient gradual typing for TypeScript. ACM SIGPLAN Notices, 50(1):167–180, January 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roy:2015:SCP

[RSGG15] Arpan Roy, Santonu Sarkar, Rajeshwari Ganesan, and Geetika Goel. Secure the cloud: From the perspective of a service-oriented organization. ACM Computing Surveys, 47 (3):41:1–41:??, April 2015. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Rodriguez-Silva:2016:IVR

[RSLAGCLB16] Daniel A. Rodríguez-Silva, Jaime Loureiro-Acuña, Francisco J. González-Castaño, and Cristina López-Bravo. Improving the virtualization of rich applications by combining VNC and streaming protocols at the hypervisor layer. Software—Practice and Experience, 46(5):625–644, May 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rodrigues:2018:CAS

[RSN^{+}18] Tiago Gama Rodrigues, Katsuya Suto, Hiroki Nishiyama, Nei Kato, and Katsuhiro Temma. Cloudlets activation scheme for scalable mobile edge computing with transmission power control and virtual machine migration. IEEE Transactions on Computers, 67(9):1287–1300, ???? 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/document/8322166/.

Rodrigues:2017:HMM

[RSNK17] T. G. Rodrigues, K. Suto, H. Nishiyama, and N. Kato. Hybrid method for minimizing service delay in edge cloud computing through VM migration and transmission power control. IEEE Transactions on Computers, 66(5):810–819, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Reano:2024:AV

[RSV24] Carlos Reaño, Federico Silla, and Blesson Varghese. Accelerator virtualization. Concurrency and Computation: Practice and Experience, 36(10):e6254:1–e6254:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rosing:1991:DPP

[RSW91] Matthew Rosing, Robert B. Schnabel, and Robert P. Weaver. The DINO parallel programming language. Journal of Parallel and Distributed Computing, 13(1):30–42, September 1, 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ramachandran:2006:NCV

[RSW^{+}06] Mahendra Ramachandran, Ned Smith, Matthew Wood, Sharad Garg, Jim Stanley, Eswar Eduri, Rinat Rappoport, Arie Chobotaro, Carl Klotz, and Lori Janz. New client virtualization usage models using Intel Virtualization Technology. Intel Technology Journal, 10(3):205–216, August 10, 2006.

ISSN 1535-766X. URL http://developer.intel.com/
technology/itj/2006/v10i3/4-models/1-abstract.htm.

Rong:1993:LMM

[RT93] Tsai Shang Rong and Lian-Jou Tsai. A logical machine monitor supporting an environment for development and execution of operating systems. The Journal of Systems and Software, 21(1):27–39, April 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ranjbari:2018:LAB

[RT18] Milad Ranjbari and Javad Akbari Torkestani. A learning automata-based algorithm for energy and SLA efficient consolidation of virtual machines in cloud data centers. Journal of Parallel and Distributed Computing, 113(??):55–62, March 2018. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151730285X.

Ren:2018:LHA

[RTL^{+}18] Shiru Ren, Le Tan, Chunqi Li, Zhen Xiao, and Weijia Song. Leveraging hardware-assisted virtualization for deterministic replay on commodity multi-core processors. IEEE Transactions on Computers, 67(1):45–58, January 2018. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.org/document/7982675/.

Rule:2007:HCC

[Rul07] David Rule. How to Cheat at Configuring VMware ESX Server. Syngress Publishing, Inc., Rockland, MA, USA, 2007. ISBN 1-59749-194-2. xvii + 372 pp. LCCN QA76.9.V5

H69 2007. URL http://www.sciencedirect.com/science/
book/9781597491945.

Russell:2008:VTF

[Rus08] Rusty Russell. virtio: towards a de-facto standard for virtual I/O devices. Operating Systems Review, 42(5):95–103, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Radhakrishnan:2001:JRS

[RVJ^{+}01] R. Radhakrishnan, N. Vijaykrishnan, L. K. John, A. Sivasubramaniam, J. Rubio, and J. Sabarinathan. Java runtime systems: characterization and architectural implications. IEEE Transactions on Computers, 50(2):131–146, February 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Raman:2021:CWS

[RWC21] Narayani Raman, Aisha Banu Wahab, and Sutherson Chandrasekaran. Computation of workflow scheduling using backpropagation neural network in cloud computing: a virtual machine placement approach. The Journal of Supercomputing, 77(9):9454–9473, September 2021. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-021-03648-0.

Ruan:2012:MVM

[RWX^{+}12] Li Ruan, Huixiang Wang, Limin Xiao, Mingfa Zhu, and Feibo Li. Memory virtualization for MIPS processor based cloud server. Lecture Notes in Computer Science, 7296: 54–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-30767-6_5/.

Ristenpart:2010:WGR

[RY10] Thomas Ristenpart and Scott Yilek. When good randomness goes bad: Virtual machine reset vulnerabilities and hedging deployed cryptography. In Anonymous [Ano10], page ?? ISBN 1-891562-29-0, 1-891562-30-

4. LCCN ????  URL http://www.isoc.org/isoc/
conferences/ndss/10/pdf/15.pdf;   http://www.isoc.
org/isoc/conferences/ndss/10/proceedings.shtml.

Rao:2014:TFE

[RZ14] Jia Rao and Xiaobo Zhou. Towards fair and efficient SMP virtual machine scheduling. ACM SIGPLAN Notices, 49(8): 273–286, August 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ren:2019:PLL

[RZPX19] Shiru Ren, Yunqi Zhang, Lichen Pan, and Zhen Xiao. Phantasy: Low-latency virtualization-based fault tolerance via asynchronous prefetching. IEEE Transactions on Computers, 68(2):225–238, ???? 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://ieeexplore.

ieee.org/document/8438984/.

Srikrishnan:2007:SFA

[SAB^{+}07] J. Srikrishnan, S. Amann, G. Banzhaf, F. W. Brice, R. Dugan, G. R. Frazier, G. P. Kuch, and J. Leopold. Sharing FCP adapters through virtualization. IBM Journal of Research and Development, 51(1/2):103–118, January/March 2007. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-

8556 (electronic). URL http://www.research.ibm.com/
journal/rd/511/srikrishnan.html.

Sierra-Arriaga:2020:SIC

[SABL20] Federico Sierra-Arriaga, Rodrigo Branco, and Ben Lee. Security issues and challenges for virtualization technologies. ACM Computing Surveys, 53(2):45:1–45:37, July 2020. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3382190.

Scazzariello:2021:MSA

[SADP21] Mariano Scazzariello, Lorenzo Ariemma, Giuseppe Di Battista, and Maurizio Patrignani. Megalos: a scalable architecture for the virtualization of large network scenarios. Future Internet, 13(9):227, August 30, 2021. CODEN ???? ISSN

1999-5903. URL https://www.mdpi.com/1999-5903/13/9/
227.

Saadatfar:2024:NDG

[SAJ24] Hamid Saadatfar, Hamid Gholampour Ahangar, and Javad Hassannataj Joloudari. A new dynamic game-based pricing model for cloud environment. Future Internet, 16(2):49, January 31, 2024. CODEN ???? ISSN 1999-5903. URL https:

//www.mdpi.com/1999-5903/16/2/49.

Salter:1992:EHW

[Sal92] Brett Salter. An exception handler for Windows 3. Dr. Dobb’s Journal of Software Tools, 17(9):32, 34, 36, 102, 104–107, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Subbiah:2014:CFT

[SAM14] Dharmalingam Subbiah, Balaji Arulmozhi, and Hariharasudhan Maruthamuthu. Constraint free testing using service virtualization. International Journal of Computer Applications, 105(??):23–26, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume105/number17/18470-9882/.

Samuelson:2022:LSA

[Sam22] Pamela Samuelson. Legally speaking: Apple’s challenge to virtualization software. Communications of the ACM, 65 (7):24–26, July 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.acm.org/

doi/10.1145/3538636.

Sandberg:1988:EOO

[San88] David W. Sandberg. Experience with an object-oriented virtual machine. Software—Practice and Experience, 18(5):415– 426 (or 415–425??), May 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sarmiento:2001:SFU

[Sar01] Evan Sarmiento. Securing FreeBSD using Jail. Sys Admin: The Journal for UNIX Systems Administrators, 10(5):31, 32, 34, 36–37, May 2001. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Sarkar:2016:VEC

[Sar16] Vivek Sarkar. Virtualizing the edge of the cloud: the new frontier. ACM SIGPLAN Notices, 51(7):1, July 2016. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shahid:2024:ONP

[SAR24] Kamal Shahid, Saleem Naseer Ahmad, and Syed Tahir Hussain Rizvi. Optimizing network performance: a comparative analysis of EIGRP, OSPF, and BGP in IPv6-based loadsharing and link-failover systems. Future Internet, 16(9):339, September 20, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/1999-5903/16/9/339.

Shiraz:2013:SVM

[SASG13] Muhammad Shiraz, Saeid Abolfazli, Zohreh Sanaei, and Abdullah Gani. A study on virtual machine deployment for application outsourcing in mobile cloud computing. The Journal of Supercomputing, 63(3):946–964, March 2013. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11227-012-0846-y.

Silva:2009:UVI

[SAT09] L. M. Silva, J. Alonso, and J. Torres. Using virtualization to improve software rejuvenation. IEEE Transactions on Computers, 58(11):1525–1538, November 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5184821.

Sayre:1966:VS

[Say66] D. Sayre. On virtual systems. Report, IBM Corporation T. J. Watson Research Laboratory, Yorktown Heights, NY, USA, April 15, 1966. ???? pp.

Sayre:1967:ACV

[Say67] D. Sayre. Adding computers virtually. IBM Corporation Computer Report, 3(2):12–15, March 1967.

Srodawa:1973:EVM

[SB73] R. J. Srodawa and L. A. Bates. An efficient virtual machine implementation. In ????, editor, Proceedings AFIPS National Computer Conference 1973. ????, ????, 1973.

Simons:2010:VHP

[SB10] Joshua E. Simons and Jeffrey Buell. Virtualizing high performance computing. Operating Systems Review, 44(4):136–145, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Samant:2016:HBS

[SB16] Durgesh Samant and Umesh Bellur. Handling boot storms in virtualized data centers — a survey. ACM Computing Surveys, 49(1):16:1–16:??, July 2016. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Son:2018:TSD

[SB18] Jungmin Son and Rajkumar Buyya. A taxonomy of softwaredefined networking (SDN)-enabled cloud computing. ACM Computing Surveys, 51(3):59:1–59:??, July 2018. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Saurabh:2020:ESC

[SBBP20] Nishant Saurabh, Shajulin Benedict, Jorge G. Barbosa, and Radu Prodan. Expelliarmus: Semantic-centric virtual machine image management in IaaS clouds. Journal of Parallel and Distributed Computing, 146(??):107–121, December 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731520303415.

Salami:2021:EEC

[SBI21] Hamza Onoruoiza Salami, Abubakar Bala, and Idris Ismail. An energy-efficient cuckoo search algorithm for virtual machine placement in cloud computing data centers. The Journal of Supercomputing, 77(11):13330–13357, November 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-021-03807-3.

Singh:2015:TVC

[SBK15] Rayman Preet Singh, Tim Brecht, and S. Keshav. Towards VM consolidation using a hierarchy of idle states. ACM SIG-PLAN Notices, 50(7):107–119, July 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scarpiniti:2018:EPH

[SBNU18] Michele Scarpiniti, Enzo Baccarelli, Paola G. Vinueza Naranjo, and Aurelio Uncini. Energy performance of heuristics and meta-heuristics for real-time joint resource scaling and consolidation in virtualized networked data centers. The Journal of Supercomputing, 74(5):2161–2198, May 2018. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sotiriadis:2017:VMC

[SBP^{+}17] Stelios Sotiriadis, Nik Bessis, Euripides G. M. Petrakis, Cristiana Amza, Catalin Negru, and Mariana Mocanu. Virtual machine cluster mobility in inter-cloud platforms. Future Generation Computer Systems, 74(??):179–189, September 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X16300206.

Sani:2014:PDF

[SBQZ14] Ardalan Amiri Sani, Kevin Boos, Shaopu Qin, and Lin Zhong. I/O paravirtualization at the device file boundary. ACM SIGARCH Computer Architecture News, 42(1):319– 332, March 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Shen:2017:DAV

[SC17] Haiying Shen and Liuhua Chen. Distributed autonomous virtual resource management in datacenters using finite-Markov decision process. IEEE/ACM Transactions on Networking, 25(6):3836–3849, December 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shen:2018:RDM

[SC18] Haiying Shen and Liuhua Chen. Resource demand misalignment: an important factor to consider for reducing resource over-provisioning in cloud datacenters. IEEE/ACM Transactions on Networking, 26(3):1207–1221, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Struhar:2024:HRO

[SCA^{+}24] Václav Struhár, Silviu S. Craciunas, Mohammad Ashjaei, Moris Behnam, and Alessandro V. Papadopoulos. Hierarchical resource orchestration framework for real-time containers. ACM Transactions on Embedded Computing Systems, 23 (1):4:1–4:??, January 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://dl.acm.org/

doi/10.1145/3592856.

Serodio:2023:ESI

[SCC^{+}23] Carlos Serôdio, José Cunha, Guillermo Candela, Santiago Rodriguez, Xosé Ramón Sousa, , and Frederico Branco. The 6G ecosystem as support for IoE and private networks: Vision, requirements, and challenges. Future Internet, 15(11):348, October 25, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/1999-5903/15/11/348.

Schuh:1990:PRI

[SCD90] Daniel T. Schuh, Michael J. Carey, and David J. DeWitt. Persistence in E revisited—implementation experiences. Technical Report CS-TR-1990-957, University of Wisconsin, Madison, August 1990.

Shi:2008:VMS

[SCEG08] Yunhe Shi, Kevin Casey, M. Anton Ertl, and David Gregg. Virtual machine showdown: Stack versus registers. ACM Transactions on Architecture and Code Optimization, 4(4): 2:1–2:??, January 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Su:2025:SGS

[SCF^{+}25] Jianzhong Su, Jiachi Chen, Zhiyuan Fang, Xingwei Lin, Yutian Tang, and Zibin Zheng. SmartOracle: Generating smart contract oracle via fine-grained invariant detection. IEEE Transactions on Software Engineering, 51(4):947–959, April 2025. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Steven:2000:JCR

[SCFP00] John Steven, Pravir Chandra, Bob Fleck, and Andy Podgurski. jRapture: a capture/replay tool for observationbased testing. ACM SIGSOFT Software Engineering Notes, 25(5):158–167, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schwenk:1973:VM

[Sch73] H. Schwenk. Virtual micromachines. In ????, editor, Proceedings ACM SIGARCH-SIGOPS Workshop on Virtual Computer Systems, Cambridge, MA, 1973, page ?? ACM Press, New York, NY 10036, USA, 1973.

Schoen:1986:CS

[Sch86] Eric Schoen. The CAOS system. Technical Report CS-TR-86-1125, Stanford University, Department of Computer Science, March 1986. 76 pp.

Schulman:1994:UCI

[Sch94a] Andrew Schulman. Undocumented corner: Introduction to ‘The Windows 3.1 Virtual Machine Control Block Part 1’ (K. Zytaruk). Dr. Dobb’s Journal of Software Tools, 19(1): 115–??, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Schulman:1994:IWV

[Sch94b] Andrew Schulman. Undocumented corner: Introduction to ‘The Windows 3.1 Virtual Machine Control Block Part 2’ (K. Zytaruk). Dr. Dobb’s Journal of Software Tools, 19(2): 107–??, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Schocken:2009:VMA

[Sch09] Shimon Schocken. Virtual machines: abstraction and implementation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):203–207, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Schmeisser:2013:MOE

[Sch13a] Michael Schmeißer. Metriken und optimale Einsatzszenarien für Garbage Collectoren der Java HotSpot Virtual Machine. (German) [Metrics and best use scenarios for garbage collectors of the Java HotSpot Virtual Machine]. Masterarbeit, Hochschule für Technik, Wirtschaft und Kultur, Leipzig, Germany, 2013. iii + 103 pp.

Schneider:2013:FVM

[Sch13b] Christian A. Schneider. Full virtual machine state reconstruction for security applications. Doktors der Naturwissenschaften (Dr. rer. nat.), Fakultät für Informatik der Technischen Universität München, Lehrstuhl für Sicherheit in der Informatik, Munich, Germany, April 23, 2013. xvi + 153

pp. URL http://mediatum.ub.tum.de/node?id=1142206;
 http://nbn-resolving.de/urn/resolver.pl?urn:nbn:
de:bvb:91-diss-20131029-1142206-0-0.

Sun:2019:MOO

[SCL^{+}19] Daniel Sun, Shiping Chen, Guoqiang Li, Yuanyuan Zhang, and Muhammad Atif. Multi-objective optimisation of online distributed software update for DevOps in clouds. ACM Transactions on Internet Technology (TOIT), 19(3):43:1– 43:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Simpkins:1993:AVM

[SCP93] N. K. Simpkins, G. Cruickshank, and P.E.International. ALEP-0 Virtual Machine extensions. Technical report, CEC, 1993.

Shi:2012:VGA

[SCSL12] Lin Shi, Hao Chen, Jianhua Sun, and Kenli Li. vCUDA: GPU-accelerated high-performance computing in virtual machines. IEEE Transactions on Computers, 61(6):804–816, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sarkar:2001:HPS

[SD01] Vivek Sarkar and Julian Dolby. High-performance scalable Java virtual machines. Lecture Notes in Computer Science, 2228:151–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2228/
22280151.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2228/22280151.pdf.

Shi:2016:PPA

[SDD^{+}16] Xiaoyu Shi, Jin Dong, Seddik M. Djouadi, Yong Feng, Xiao Ma, and Yefu Wang. PAPMSC: Power-aware performance management approach for virtualized Web servers via stochastic control. Journal of Grid Computing, 14(1):171– 191, March 2016. CODEN ???? ISSN 1570-7873 (print),

1572-9184 (electronic). URL http://link.springer.com/
article/10.1007/s10723-015-9341-z.

Simoes:2021:DAS

[SDM21] Rhodney Simões, Kelvin Dias, and Ricardo Martins. Dynamic allocation of SDN controllers in NFV-based MEC for the Internet of Vehicles. Future Internet, 13(11):270, October 26, 2021. CODEN ???? ISSN 1999-5903. URL https://www.

mdpi.com/1999-5903/13/11/270.

Sridevi:2009:NAE

[SDN09] R. Sridevi, A. Damodaram, and S. V. L. Narasimham. A novel architecture for enhanced security through virtualisation and spoofing. Network Security, 2009(9):12–16, September 2009. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-

9371 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135348580970100X.

Song:2021:CRE

[SDS^{+}21] Weijia Song, Christina Delimitrou, Zhiming Shen, Robbert Van Renesse, Hakim Weatherspoon, Lotfi Benmohamed, Frederic De Vaulx, and Charif Mahmoudi. CacheInspector: Reverse engineering cache resources in public clouds. ACM Transactions on Architecture and Code Optimization, 18(3): 35:1–35:25, June 2021. CODEN ???? ISSN 1544-3566 (print),

1544-3973 (electronic). URL https://dl.acm.org/doi/10.
1145/3457373.

Sartor:2012:EMT

[SE12] Jennfer B. Sartor and Lieven Eeckhout. Exploring multithreaded Java application performance on multicore hardware. ACM SIGPLAN Notices, 47(10):281–296, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sedighi:2007:EV

[Sed07] Art Sedighi. Editorial: Virtualization. Scalable Computing: Practice and Experience, 8(2):i–ii, June 2007. CODEN ????

ISSN 1895-1767. URL http://www.scpe.org/vols/vol08/
no2/vol08no2editorial.html.

Seecker:2008:EGS

[See08a] Robert Seecker. Erstellung einer gehärteten Systemarchitecktur mit der Virtualisierungslösung Xen zur Konsolidierung von Kundenserversystemen in einem Rechenzentrum. Bachelor thesis, Fachhochschule, Stralsund, Germany, 2008. vii + 86 pp.

Seeling:2008:L

[See08b] Patrick Seeling. Labs@home. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4): 75–77, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Seely:2010:BVD

[See10] Andrew Seely. Building a virtual DNS appliance using Solaris 10, BIND, and VMware. ;login: the USENIX Association newsletter, 35(3):27–34, June 2010. CODEN LOGNEM. ISSN 1044-6397. URL https://www.usenix.

org/publications/login/june-2010-volume-35-number-
3/building-virtual-dns-appliance-using-solaris-10-
bind.

Smith:2006:SID

[SEF^{+}06] Matthew Smith, Michael Engel, Thomas Friese, Bernd Freisleben, Gregory A. Koenig, and William Yurcik. Security issues in on-demand grid and cluster computing. In Turner et al. [TLC06], pages 24–?? ISBN 0-7695-2585-7. LCCN QA76.9.C58. IEEE Computer Society Order Number P2585.

Staples:2019:SAB

[SEK^{+}19] J. Staples, C. Endicott, L. Krause, P. Pal, P. Samouelian, R. Schantz, and A. Wellman. A semi-autonomic bytecode repair framework. IEEE Software, 36(2):97–102, March/April 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Sha:2020:MVM

[SEM^{+}20] Jing Sha, Abdol Ghaffar Ebadi, Dinesh Mavaluru, Mohmmed Alshehri, Osama Alfarraj, and Lila Rajabion. A method for virtual machine migration in cloud computing using a collective behavior-based metaheuristics algorithm. Concurrency and Computation: Practice and Experience, 32(2):e5441:1– e5441:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Salimian:2016:AFT

[SENS16] Leili Salimian, Faramarz Safi Esfahani, and Mohammad-Hossein Nadimi-Shahraki. An adaptive fuzzy threshold-based approach for energy and performance efficient consolidation of virtual machines. Computing: Archiv für Informatik und Numerik, 98(6):641–660, June 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Simao:2019:GWS

[SEPV19] J. Simão, S. Esteves, André Pires, and L. Veiga. GC-Wise: a self-adaptive approach for memory-performance efficiency in Java VMs. Future Generation Computer Systems, 100(??):674–688, November 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X18304898.

Seth:2013:UJV

[Set13] Sachin Seth. Understanding Java Virtual Machine. Alpha Science International, Oxford, UK, 2013. ISBN 1-84265-815-8. 318 pp. LCCN QA76.73.J38 S437 2013.

Sturley:2024:VVC

[SFSN^{+}24] Hamish Sturley, Augustin Fournier, Andoni Salcedo-Navarro, Miguel Garcia-Pineda, and Jaume Segura-Garcia. Virtualization vs. containerization, a comparative approach for application deployment in the computing continuum focused on the edge. Future Internet, 16(11):427, November 19, 2024. CODEN ???? ISSN 1999-5903. URL https://www.mdpi.

com/1999-5903/16/11/427.

Spinellis:2009:BA

[SG09] Diomidis Spinellis and Georgios Gousios, editors. Beautiful architecture. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-15578-6. xix + 404 pp. LCCN QA76.754 .B43 2009. URL http://

proquest.safaribooksonline.com/9780596155780.

Schmidt:2010:VSB

[SG10a] René W. Schmidt and Steffen Grarup. vApp: a standardsbased container for cloud providers. Operating Systems Review, 44(4):115–123, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Soundararajan:2010:CBS

[SG10b] Vijayaraghavan Soundararajan and Kinshuk Govil. Challenges in building scalable virtualized datacenter management. Operating Systems Review, 44(4):95–102, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shuja:2016:SMD

[SGB^{+}16] Junaid Shuja, Abdullah Gani, Kashif Bilal, Atta Ur Rehman Khan, Sajjad A. Madani, Samee U. Khan, and Albert Y. Zomaya. A survey of mobile device virtualization: Taxonomy and state of the art. ACM Computing Surveys, 49(1):1:1– 1:??, July 2016. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Sirer:1999:DID

[SGGB99] Emin Gün Sirer, Robert Grimm, Arthur J. Gregory, and Brian N. Bershad. Design and implementation of a distributed virtual machine for networked computers. Operating Systems Review, 33(5):202–216, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sirer:2000:DID

[SGGB00] Emin Gün Sirer, Robert Grimm, Arthur J. Gregory, and Brian N. Bershad. Design and implementation of a distributed virtual machine for networked computers. Operating Systems Review, 34(2):23, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saravanakumar:2023:ETV

[SGK^{+}23] C. Saravanakumar, M. Geetha, S. Manoj Kumar, S. Manikandan, C. Arun, and K. Srivatsan. An efficient technique for virtual machine clustering and communications using task-based scheduling in cloud computing. Scientific Programming, 2023(1):5586521:1–5586521:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1155/2021/5586521.

Saeed:1992:ICM

[SGS92] Faisel Saeed, K. M. George, and M. H. Samadzadeh. Implementation of classical mutual exclusion algorithms in Ada. ACM SIGADA Ada Letters, 12(1):73–84, January/February 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Simao:2012:CER

[SGV12] José Simão, Tiago Garrochinho, and Luís Veiga. A checkpointing-enabled and resource-aware Java Virtual Machine for efficient and robust e-Science applications in grid environments. Concurrency and Computation: Practice and Experience, 24(13):1421–1442, September 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shanmuganathan:2013:DCU

[SGV13] Ganesha Shanmuganathan, Ajay Gulati, and Peter Varman. Defragmenting the cloud using demand-based resource allocation. ACM SIGMETRICS Performance Evaluation Review, 41(1):67–80, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schmalenbach:2004:JVM

[SH04] C. Schmalenbach and C. Hofig. The Java Virtual Machine profiler interface. Dr. Dobb’s Journal of Software Tools, 29 (7):28–33, 2004. CODEN DDJOEB. ISSN 1044-789X.

Stefanovic:2003:OFG

[SHB^{+}03] Darko Stefanović, Matthew Hertz, Stephen M. Blackburn, Kathryn S. McKinley, and J. Eliot B. Moss. Older-first garbage collection in practice: Evaluation in a Java virtual machine. ACM SIGPLAN Notices, 38(2s):25–36, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Son:2019:CNM

[SHB19] Jungmin Son, TianZhang He, and Rajkumar Buyya. CloudSimSDN-NFV: Modeling and simulation of network function virtualization and service function chaining in edge computing environments. Software—Practice and Experience, 49(12):1748–1764, December 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shen:1991:VTD

[She91] Shioupyn Shen. The virtual-time data-parallel machine. Thesis (Ph.D.), Department of Computer Science, University of California, Los Angeles, Los Angeles, CA, USA, 1991. xi + 115 pp.

Shelburne:2002:PEP

[She02] Brian J. Shelburne. A PDP-8 emulator program. ACM Journal on Educational Resources in Computing (JERIC), 2(1): 17–47, March 2002. CODEN ???? ISSN 1531-4278.

Shippy:2003:PGT

[Shi03] P. J. Shippy. Porting the Gnat tasking runtime system to the Java Virtual Machine. Literary Review, 47(2):119–120, 2003. CODEN ???? ISSN 0024-4589.

Shao:2013:VOS

[SHLJ13] Zhiyuan Shao, Ligang He, Zhiqiang Lu, and Hai Jin. VSA: an offline scheduling analyzer for Xen virtual machine monitor. Future Generation Computer Systems, 29(8):2067–2076, October 2013. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167739X12002245.

Shriver:1989:PTA

[Shr89] B. D. Shriver, editor. Proceedings of the Twenty-Second Annual Hawaii International Conference on System Sciences. Vol.II: Software Track, Kailua-Kona, HI, USA, January 3– 6, 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-1912-0. LCCN ???? IEEE catalog number 89TH0243-6.

Sayadnavard:2019:CRE

[SHR19a] Monireh H. Sayadnavard, Abolfazl Toroghi Haghighat, and Amir Masoud Rahmani. Correction to: A reliable energyaware approach for dynamic virtual machine consolidation in cloud data centers. The Journal of Supercomputing, 75(4):2148, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11227-018-02733-

1.pdf. See [SHR19b].

Sayadnavard:2019:REA

[SHR19b] Monireh H. Sayadnavard, Abolfazl Toroghi Haghighat, and Amir Masoud Rahmani. A reliable energy-aware approach for dynamic virtual machine consolidation in cloud data centers. The Journal of Supercomputing, 75(4):2126–2147, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See [SHR19a].

Svard:2011:EDC

[SHTE11] Petter Svärd, Benoit Hudzia, Johan Tordsson, and Erik Elmroth. Evaluation of delta compression techniques for efficient live migration of large virtual machines. ACM SIGPLAN Notices, 46(7):111–120, July 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sard:2015:PPC

[SHW^{+}15] Petter Särd, Benoit Hudzia, Steve Walsh, Johan Tordsson, and Erik Elmroth. Principles and performance characteristics of algorithms for live VM migration. Operating Systems Review, 49(1):142–155, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Song:2014:OBS

[SHZ^{+}14] Fei Song, Daochao Huang, Huachun Zhou, Hongke Zhang, and Ilsun You. An optimization-based scheme for efficient virtual machine placement. International Journal of Parallel Programming, 42(5):853–872, October 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10766-013-0274-5.

Sarda:1981:CAD

[SI81] N. L. Sarda and J. R. Isaac. Computer aided design of database internal schema. International Journal of Computer and Information Sciences, 10(4):219–234, August 1981. CO-DEN IJCIAH. ISSN 0091-7036.

Suneja:2015:EVI

[SIdLB15] Sahil Suneja, Canturk Isci, Eyal de Lara, and Vasanth Bala. Exploring VM introspection: Techniques and tradeoffs. ACM SIGPLAN Notices, 50(7):133–146, July 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Signorini:1989:HSM

[Sig89] J. Signorini. How a SIMD machine can implement a complex cellular automaton? A case study: von Neumann’s 29-state cellular automaton. In ACM [ACM89], pages 175–186. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

So-In:2011:VAU

[SIJPP11] Chakchai So-In, Raj Jain, Subharthi Paul, and Jianli Pan. Virtualization architecture using the ID/Locator split concept for Future Wireless Networks (FWNs). Computer Networks (Amsterdam, Netherlands: 1999), 55(2):415–430, February 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Solaimani:2016:OAD

[SIK^{+}16] Mohiuddin Solaimani, Mohammed Iftekhar, Latifur Khan, Bhavani Thuraisingham, Joe Ingram, and Sadi Evren Seker. Online anomaly detection for multi-source VMware using a distributed streaming framework. Software—Practice and Experience, 46(11):1479–1497, November 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Simpkins:1992:AVP

[Sim92] N. K. Simpkins. ALEP-0 Version 2.2: Prototype Virtual Machine. Technical report, CEC, Luxembourg, Luxembourg, 1992. (User guide for the Advanced Language Engineering Platform). BIM.

Santanna:2017:DIS

[SIR^{+}17] Francisco Sant’anna, Roberto Ierusalimschy, Noemi Rodriguez, Silvana Rossetto, and Adriano Branco. The design and implementation of the synchronous language CéU. ACM Transactions on Embedded Computing Systems, 16(4):98:1– 98:26, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Silla:2017:BRG

[SIRP17] Federico Silla, Sergio Iserte, Carlos Reaño, and Javier Prades. On the benefits of the remote GPU virtualization mechanism: The rCUDA case. Concurrency and Computation: Practice and Experience, 29(13), July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Siveroni:2004:OSJ

[Siv04] I. A. Siveroni. Operational semantics of the Java Card Virtual Machine. Journal of Logic and Algebraic Programming, 58(1– 2):3–25, 2004. CODEN ???? ISSN 1567-8326.

Sivakumar:2007:CCA

[Siv07] Nishant Sivakumar. C++/CLI in action. Manning Publications, Greenwich, CT, USA, 2007. ISBN 1-932394-81-8 (paperback). xxiii + 391 pp. LCCN QA76.73.C153 S52 2007.

Sallam:2021:JPA

[SJ21] Gamal Sallam and Bo Ji. Joint placement and allocation of VNF nodes with budget and capacity constraints. IEEE/ ACM Transactions on Networking, 29(3):1238–1251, June 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-

2566 (electronic). URL https://dl.acm.org/doi/10.1109/
TNET.2021.3058378.

Song:2017:EPU

[SJA^{+}17] Wonjun Song, Hyung-Joon Jung, Jung Ho Ahn, Jae W. Lee, and John Kim. Evaluation of performance unfairness in NUMA system architecture. IEEE Computer Architecture Letters, 16(1):26–29, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Salehi:2014:RPB

[SJB14] Mohsen Amini Salehi, Bahman Javadi, and Rajkumar Buyya. Resource provisioning based on preempting virtual machines in distributed systems. Concurrency and Computation: Practice and Experience, 26(2):412–433, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shi:2012:TSW

[SJJ^{+}12] Xuanhua Shi, Hai Jin, Hongbo Jiang, Xiaodong Pan, Dachuan Huang, and Bo Yu. Toward scalable Web systems on multicore clusters: making use of virtual machines.

The Journal of Supercomputing, 61(1):27–45, July 2012. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&volume=61&issue=
1&spage=27.

Son:2020:DII

[SJL20] Yunsik Son, Junho Jeong, and YangSun Lee. Design and implementation of an IoT–cloud converged virtual machine system. The Journal of Supercomputing, 76(7):5259–5275, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sharma:2024:NCM

[SJMG24] Sidharth Sharma, Admela Jukan, Aashi Malik, and Ashwin Gumaste. A network calculus model for SFC realization and traffic bounds estimation in data centers. ACM Transactions on Internet Technology (TOIT), 24(4):21:1–21:??, November 2024. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3700440.

Sem-Jacobsen:2013:ELC

[SJRS^{+}13] Frank Olaf Sem-Jacobsen, Samuel Rodrigo, Tor Skeie, Alessandro Strano, and Davide Bertozzi. An efficient, lowcost routing framework for convex mesh partitions to support virtualization. ACM Transactions on Embedded Computing Systems, 12(4):107:1–107:??, June 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shen:2017:SLC

[SJS^{+}17] Zhiming Shen, Qin Jia, Gur-Eyal Sela, Weijia Song, Hakim Weatherspoon, and Robbert Van Renesse. Supercloud: a library cloud for exploiting cloud diversity. ACM Transactions on Computer Systems, 35(2):6:1–6:??, October 2017. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Sailer:2005:BMB

[SJV^{+}05] Reiner Sailer, Trent Jaeger, Enriquillo Valdez, Ramon Caceres, Ronald Perez, Stefan Berger, John Linwood Griffin, and Leendert van Doorn. Building a MAC-based security architecture for the Xen open-source hypervisor. In IEEE [IEE05], pages 276–285. ISBN 0-7695-2461-3. ISSN 1063-9527. LCCN L787.5. IEEE Computer Society Order Number P2461.

Shi:2013:AGC

[SJW^{+}13] Xuanhua Shi, Hai Jin, Song Wu, Wei Zhu, and Li Qi. Adapting grid computing environments dependable with virtual machines: design, implementation, and evaluations. The Journal of Supercomputing, 66(3):1152–1166, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL http://link.springer.com/article/
10.1007/s11227-011-0664-7.

Salkeld:2013:IDO

[SK13a] Robin Salkeld and Gregor Kiczales. Interacting with dead objects. ACM SIGPLAN Notices, 48(10):203–216, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Sanchez:2013:ZFA

[SK13b] Daniel Sanchez and Christos Kozyrakis. ZSim: fast and accurate microarchitectural simulation of thousand-core systems. ACM SIGARCH Computer Architecture News, 41(3): 475–486, June 2013. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ICSA ’13 conference proceedings.

Sudevalayam:2013:AAM

[SK13c] Sujesha Sudevalayam and Purushottam Kulkarni. Affinityaware modeling of CPU usage with communicating virtual machines. The Journal of Systems and Software, 86(10):2627–2638, October 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0164121213001246.

Skapinetz:2007:VBT

[Ska07] Kevin Skapinetz. Virtualisation as a blackhat tool. Network Security, 2007(10):4–7, October 2007. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1353485807700922.

Sitton:1973:PEL

[SKC73] Gary A. Sitton, Thomas A. Kendrick, and A. Gil Carrick. The PL/EXUS language and virtual machine. ACM SIG-PLAN Notices, 8(11):124–130, November 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Suneja:2017:SIL

[SKI^{+}17] Sahil Suneja, Ricardo Koller, Canturk Isci, Eyal de Lara, Ali Hashemi, Arnamoy Bhattacharyya, and Cristiana Amza. Safe inspection of live virtual machines. ACM SIGPLAN Notices, 52(7):97–111, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Song:2017:HBA

[SKJ^{+}17] Wonjun Song, Gwangsun Kim, Hyungjoon Jung, Jongwook Chung, Jung Ho Ahn, Jae W. Lee, and John Kim. Historybased arbitration for fairness in processor-interconnect of NUMA servers. ACM SIGARCH Computer Architecture News, 45(1):765–777, March 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Skrien:2001:CST

[Skr01] Dale Skrien. CPU Sim 3.1: a tool for simulating computer architectures for computer organization classes. ACM Journal on Educational Resources in Computing (JERIC), 1(4): 46–59, December 2001. CODEN ???? ISSN 1531-4278. URL

http://www.cs.colby.edu/djskrien/CPUSim/.

Stamou:2019:ANM

[SKT^{+}19] Adamantia Stamou, Grigorios Kakkavas, Konstantinos Tsitseklis, Vasileios Karyotis, and Symeon Papavassiliou. Autonomic network management and cross-layer optimization in software defined radio environments. Future Internet, 11(2): 37, February 03, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/1999-5903/11/2/37.

Suzuki:2016:GGV

[SKYK16] Yusuke Suzuki, Shinpei Kato, Hiroshi Yamada, and Kenji Kono. GPUvm: GPU virtualization at the hypervisor. IEEE Transactions on Computers, 65(9):2752–2766, ???? 2016.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shyu:2000:APV

[SL00] Shyong-Jian Shyu and B. M. T. Lin. An application of parallel virtual machine framework to film production problem. Computers and Mathematics with Applications, 39(12):53– 62, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0898122100001292.

Szefer:2012:ASH

[SL12] Jakub Szefer and Ruby B. Lee. Architectural support for hypervisor-secure virtualization. ACM SIGARCH Computer Architecture News, 40(1):437–450, March 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ASPLOS ’12 conference proceedings.

Sallam:2014:MOV

[SL14] Ahmed Sallam and Kenli Li. A multi-objective virtual machine migration policy in cloud systems. The Computer Journal, 57(2):195–204, February 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/57/2/195.full.
pdf+html.

Sgandurra:2016:EAT

[SL16] Daniele Sgandurra and Emil Lupu. Evolution of attacks, threat models, and solutions for virtualized systems. ACM Computing Surveys, 48(3):46:1–46:??, February 2016. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Sun:2016:NTE

[SLA^{+}16] Gang Sun, Dan Liao, Vishal Anand, Dongcheng Zhao, and Hongfang Yu. A new technique for efficient live migration of multiple virtual machines. Future Generation Computer Systems, 55(??):74–86, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X15002848.

Shirinbab:2020:PEC

[SLC20] Sogand Shirinbab, Lars Lundberg, and Emiliano Casalicchio. Performance evaluation of containers and virtual machines when running Cassandra workload concurrently. Concurrency and Computation: Practice and Experience, 32(17): e5693:1–e5693:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Scott:1989:EOS

[SLM89] M. L. Scott, T. J. LeBlanc, and B. D. Marsh. Evolution of an operating system for large-scale shared-memory multiprocessors. TR TR309, University of Rochester, Computer Science Department, March 1989. URL ftp:/

/ftp.cs.rochester.edu/pub/papers/systems/89.TR309.

Psyche_Evolution.ps.Z. Thu, 17 Jul 97 09:00:00 GMT.

Sha:2024:HSC

[SLW^{+}24] Sai Sha, Chuandong Li, Xiaolin Wang, Zhenlin Wang, and Yingwei Luo. Hardware–software collaborative tieredmemory management framework for virtualization. ACM Transactions on Computer Systems, 42(1–2):4:1–4:??, May 2024. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-

7333 (electronic). URL https://dl.acm.org/doi/10.1145/
3639564.

Singh:2025:EGC

[SLZ^{+}25] Radheshyam Singh, Line M. P. Larsen, Eder Ollora Zaballa, Michael Stübert Berger, Christian Kloch, and Lars Dittmann. Enabling green cellular networks: a review and proposal leveraging software-defined networking, network function virtualization, and cloud-radio access network. Future Internet, 17(4):161, April 5, 2025. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.com/1999-5903/17/4/161.

Seawright:1979:VSM

[SM79] L. H. Seawright and R. A. MacKinnon. VM/370 — a study of multiplicity and usefulness. IBM Systems Journal, 18(1):4–17, 1979. CODEN IBMSA7. ISSN 0018-

8670. URL https://pages.cs.wisc.edu/~stjones/proj/
vm_reading/ibmsj1801C.pdf.

Seiden:1990:AFV

[SM90] K. F. Seiden and J. P. Melanson. The auditing facility for a VMM security kernel. In IEEE [IEE90a], pages 262–277. ISBN 0-8186-2060-9, 0-8186-6060-0, 0-8186-9060-7. LCCN QA76.9.A25; QA76.9.A25 S95 1990eb; QA76.9.A25 I34 1990. IEEE Computer Society Order Number 2060. IEEE Catalog Number 90CH2884-5.

Sterrett:1992:PMA

[SM92] Anthony Sterrett and Marvin Minei. Performance measures of the Ada Rendezvous. ACM SIGADA Ada Letters, 12(2): 97–101, March/April 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Shudo:2001:AME

[SM01] Kazuyuki Shudo and Yoichi Muraoka. Asynchronous migration of execution context in Java Virtual Machines. Future Generation Computer Systems, 18(2):225–233, October 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115

(electronic). URL http://www.elsevier.com/gej-ng/10/
19/19/60/31/30/abstract.html.

Surdeanu:2002:DPA

[SM02] Mihai Surdeanu and Dan Moldovan. Design and performance analysis of a distributed Java Virtual Machine. IEEE Transactions on Parallel and Distributed Systems, 13(6):611–627, June 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-

2183 (electronic). URL http://dlib.computer.org/td/
books/td2002/pdf/l0611.pdf;   http://www.computer.
org/tpds/td2002/l0611abs.htm.

Seetharaman:2006:TOU

[SM06] Swaminathan Seetharaman and Krishna Murthy. Test optimization using software virtualization. IEEE Software, 23(5): 66–69, September/October 2006. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Sheeba:2023:EFT

[SM23a] Adlin Sheeba and B. Uma Maheswari. An efficient fault tolerance scheme based enhanced firefly optimization for virtual machine placement in cloud computing. Concurrency and Computation: Practice and Experience, 35(7):e7610:1– e7610:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Siavashi:2023:GMO

[SM23b] Ahmad Siavashi and Mahmoud Momtazpour. gVMP: a multi-objective joint VM and vGPU placement heuristic for API remoting-based GPU virtualization and disaggregation in cloud data centers. Journal of Parallel and Distributed Computing, 172(??):97–113, February 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002234.

Soror:2010:AVM

[SMA^{+}10] Ahmed A. Soror, Umar Farooq Minhas, Ashraf Aboulnaga, Kenneth Salem, Peter Kokosielis, and Sunil Kamath. Automatic virtual machine configuration for database workloads. ACM Transactions on Database Systems, 35(1):7:1– 7:??, February 2010. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Shi:2018:HAV

[SMA18] Hao Shi, Jelena Mirkovic, and Abdulla Alwabel. Handling anti-virtual machine techniques in malicious software. ACM Transactions on Privacy and Security (TOPS), 21(1): 2:1–2:??, January 2018. ISSN 2471-2566 (print), 2471-2574

(electronic). URL https://dl.acm.org/citation.cfm?id=
3139292.

Schneider:2001:APM

[SMES01] Daniel Schneider, Bernd Mathiske, Matthias Ernst, and Matthew Seidl. Automatic persistent memory management for the Spotless Java^{T M} Virtual Machine on the Palm Connected Organizer. In USENIX [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL http:

//www.usenix.org/publications/library/proceedings/
jvm01/schneider.html.

Smith:1997:JNV

[Smi97] Gregory S. Smith. Java’s new virtual machine. Java-World: IDG’s magazine for the Java community, 2(5):??, May 1997. CODEN ???? ISSN 1091-8906. URL http:

//www.javaworld.com/javaworld/jw-05-1997/jw-05-jo-
vm.htm.

S:2002:SPI

[SMK02] Venugopal K. S., Geetha Manjunath, and Venkatesh Krishnan. sEc: a portable interpreter optimizing technique for embedded Java Virtual Machine. In USENIX [USE02], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38

J42 2002. URL http://www.usenix.org/publications/
library/proceedings/javavm02/venugopal.html.

Silva:2018:FPD

[SML18] Nuno Silva, Eduardo R. B. Marques, and Luís M. B. Lopes. Flux: a platform for dynamically reconfigurable mobile crowd-sensing. ACM Transactions on Sensor Networks, 14 (3–4):20:1–20:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Steensgaard-Madsen:1984:DPL

[SMO84] J. Steensgaard-Madsen and L. M. Olsen. Definition of the programming language MODEF. ACM SIGPLAN Notices, 19(2):92–110, February 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sa:2022:FLR

[SMP22] Bruno Sá, José Martins, and Sandro Pinto. A first look at RISC-V virtualization from an embedded systems perspective. IEEE Transactions on Computers, 71(9):2177–2190, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Silva:2024:EDE

[SMR24] Lucas Silva, José Metrôlho, and Fernando Ribeiro. Efficient data exchange between WebAssembly modules. Future Internet, 16(9):341, September 20, 2024. CODEN ???? ISSN

1999-5903. URL https://www.mdpi.com/1999-5903/16/9/
341.

Sewe:2011:CCS

[SMSB11] Andreas Sewe, Mira Mezini, Aibek Sarimbekov, and Walter Binder. Da capo con Scala: design and analysis of a Scala benchmark suite for the Java Virtual Machine. ACM SIG-PLAN Notices, 46(10):657–676, October 2011. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’11 conference proceedings.

Santos:2018:HDD

[SMSH18] Eddie Antonio Santos, Carson McLean, Christopher Solinas, and Abram Hindle. How does Docker affect energy consumption? Evaluating workloads in and out of Docker containers. The Journal of Systems and Software, 146(??):14–25, December 2018. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0164121218301456.

Smith:2005:AVM

[SN05a] James E. Smith and Ravi Nair. The architecture of virtual machines. Computer, 38(5):32–??, May 2005. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.org/comp/mags/co/
2005/05/r5032abs.htm;   http://csdl.computer.org/
dl/mags/co/2005/05/r5032.pdf.

Smith:2005:VMV

[SN05b] James E. (James Edward) Smith and Ravi Nair. Virtual Machines: Versatile Platforms for Systems and Processes. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2005. ISBN 1-55860-910-5 (hardcover), 0-08-052540-7 (e-book), 1-4933-0376-7. xxii + 638 pp. LCCN

QA76.9.V5 S54 2005. URL http://books.elsevier.
com/us/bookscat/search/details.asp?country=United+
States&community=mk&imprint=Morgan+Kaufmann&isbn=
1558609105.

Sharma:2023:CSD

[SN23] Sachin Sharma and Avishek Nag. Cognitive software defined networking and network function virtualization and applications. Future Internet, 15(2):78, February 17, 2023. CODEN

???? ISSN 1999-5903. URL https://www.mdpi.com/1999-
5903/15/2/78.

Stone:1991:VCS

[SNC91] R. L. Stone, T. S. Nettleship, and J. Curtiss. VM/ESA CMS Shared File System. IBM Systems Journal, 30(1):52–71, 1991. CODEN IBMSA7. ISSN 0018-8670.

Stutz:2003:SSC

[SNS03] David Stutz, Ted Neward, and Geoff Shilling. Shared Source CLI Essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00351-X. xviii + 357 pp. LCCN QA76.625 .S76 2003. US$34.95.

URL http://www.oreilly.com/catalog/9780596003517;
 http://www.oreilly.com/catalog/sscliess.

Scales:2010:DPS

[SNV10] Daniel J. Scales, Mike Nelson, and Ganesh Venkitachalam. The design of a practical system for fault-tolerant virtual machines. Operating Systems Review, 44(4):30–39, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sahni:2023:AAS

[SOAK23] Abdul Rasheed Sahni, Hamza Omar, Usman Ali, and Omer Khan. ASM: an adaptive secure multicore for co-located mutually distrusting processes. ACM Transactions on Architecture and Code Optimization, 20(3):32:1–32:??, September 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3587480.

STUG:1983:PUA

[Sof83] Software Tools Users Group, editor. Proceedings: USENIX Association [and] Software Tools Users Group Summer Conference, Toronto 1983, July 1983, Toronto, Ontario, Canada. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1983. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

St-Onge:2023:NMR

[SOKE23] Cédric St-Onge, Nadjia Kara, and Claes Edstrom. NFVLearn: a multi-resource, long short-term memory-based virtual network function resource usage prediction architecture. Software—Practice and Experience, 53(3):555–578, March 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schaefer:1983:IPU

[SP83] Mark T. Schaefer and Yale N. Patt. Improving the performance of UCSD Pascal via microprogramming on the PDP-11/60. ACM SIGMICRO Newsletter, 14(4):140–148, December 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/10.1145/1096419.
1096440.

Surantha:2022:ISN

[SP22] Nico Surantha and Noffal A. Putra. Integrated SDN-NFV 5G network performance and management-complexity evaluation. Future Internet, 14(12):378, December 14, 2022. CO-DEN ???? ISSN 1999-5903. URL https://www.mdpi.com/

1999-5903/14/12/378.

Sunil:2023:EEV

[SP23] Shilpa Sunil and Sanjeev Patel. Energy-efficient virtual machine placement algorithm based on power usage. Computing: Archiv für Informatik und Numerik, 105(7):1597–1621, July 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057

(electronic). URL https://link.springer.com/article/
10.1007/s00607-023-01152-2.

Sparks:2019:EDH

[Spa19] Jonathan Sparks. Enabling Docker for HPC. Concurrency and Computation: Practice and Experience, 31(16):e5018:1– e5018:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Song:2018:GAH

[SPAK18] Tae-Geon Song, Mehdi Pirahandeh, Cheong-Jin Ahn, and Deok-Hwan Kim. GPU-accelerated high-performance encoding and decoding of hierarchical RAID in virtual machines. The Journal of Supercomputing, 74(11):5865–5888, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Soltesz:2007:CBO

[SPF^{+}07] Stephen Soltesz, Herbert Pötzl, Marc E. Fiuczynski, Andy Bavier, and Larry Peterson. Container-based operating system virtualization: a scalable, high-performance alternative to hypervisors. Operating Systems Review, 41(3):275–287, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spivey:2006:VHH

[Spi06] Mark D. Spivey. Virtually hacking: hacking the virtual computer. Taylor and Francis, Boca Raton, FL, USA, 2006. ISBN 0-8493-7057-4. ???? pp. LCCN QA76.76.O63 S6755

2006. URL http://www.loc.gov/catdir/toc/ecip0612/
2006013484.html.

Sprang:2006:XVL

[Spr06] Henning Sprang. Xen: Virtualisierung unter Linux. Open Source Press, München, Germany, 2006. ISBN 3-937514-29-5. 350 pp. LCCN ???? ca. EUR 39.90, EUR 41.35 (AT).

Sprang:2007:XVL

[Spr07] Henning Sprang, editor. Xen: Virtualisierung unter Linux. (German) [Xen: Virtualization under Linux]. Open Source Press, München, Germany, 2007. ISBN 3-937514-29-5. 350

pp. LCCN ???? URL http://deposit.ddb.de/cgi-bin/
dokserv?id=2809360.

Swain:2025:IVM

[SPSL25] Smruti Rekha Swain, Anshu Parashar, Ashutosh Kumar Singh, and Chung Nan Lee. An intelligent virtual machine allocation optimization model for energy-efficient and reliable cloud environment. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL https://link.springer.com/
article/10.1007/s11227-024-06734-1.

Stagner:2009:PHV

[SRS09] Harley Stagner, Jon Rolfe, and Greg Shields. Pro Hyper-V. Expert’s voice in virtualization. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-1908-4, 1-4302-1909-2 (electronic). xxii + 425 pp. LCCN QA76.9.V5 S83 2009.

Stoy:1972:OEOa

[SS72] J. E. Stoy and C. Strachey. OS6 — an experimental operating system for a small computer. Part 1: general principles and structure. The Computer Journal, 15(2): 117–124, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_15/Issue_02/150117.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/117.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/
tiff/118.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/119.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/
tiff/120.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/121.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/
tiff/122.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/123.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/
tiff/124.tif.

Saltzer:1975:PIC

[SS75] Jerome H. Saltzer and Michael D. Schroeder. The protection of information in computer systems. Proceedings of the IEEE, 63(9):1278–1308, September 1975. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic). URL http://www.

mediacity.com/~norm/CapTheory/ProtInf/.

Shih:2005:ICA

[SS05] Timothy K. Shih and Yoshitaka Shibata, editors. 19th International Conference on Advanced Information Networking and Applications: proceedings, AINA 2005, 28–30 March, 2005, Taipei, Taiwan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2249-1 (paperback). ISSN 1550-445X. LCCN TK5105.5 .I5616 2005. URL http://ieeexplore.

ieee.org/servlet/opac?punumber=9746. IEEE Computer

Society Order Number P2249.

Salimi:2013:BSC

[SS13] Hadi Salimi and Mohsen Sharifi. Batch scheduling of consolidated virtual machines based on their workload interference model. Future Generation Computer Systems, 29(8):2057–2066, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X13000332.

Soundararajan:2017:SFC

[SS17] Vijayaraghavan Soundararajan and Joshua Schnee. Sustainability as a first-class metric for developers and end-users. Operating Systems Review, 51(1):60–66, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shooshtarian:2019:MRE

[SS19] L. Shooshtarian and F. Safaei. A maximally robustness embedding algorithm in virtual data centers with multiattribute node ranking based on TOPSIS. The Journal of Supercomputing, 75(12):8059–8093, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Singh:2022:OSA

[SS22] Garima Singh and Anil Kumar Singh. Optimization of SLA aware live migration of multiple virtual machines using Lagrange multiplier. Future Generation Computer Systems, 130(??):279–291, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X22000073.

Stark:2001:JJV

[SSB01] Robert F. Stärk, Joachim Schmid, and Egon Börger. Java and the Java Virtual Machine: definition, verification, validation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42088-6. x + 381 pp. LCCN QA76.73.J38 S785 2001. US$49.95. Includes CD-ROM with the entire text of the book and numerous examples and exercises.

Shaylor:2003:JVM

[SSB03] Nik Shaylor, Douglas N. Simon, and William R. Bush. A Java virtual machine architecture for very small devices. ACM SIGPLAN Notices, 38(7):34–41, July 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sarimbekov:2014:JCS

[SSB^{+}14a] Aibek Sarimbekov, Andreas Sewe, Walter Binder, Philippe Moret, and Mira Mezini. JP2: Call-site aware calling context profiling for the Java Virtual Machine. Science of Computer Programming, 79(??):146–157, January 1, 2014. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167642311002036.

Stark:2014:JJV

[SSB14b] Robert F. Stärk, Joachim Schmid, and Egon Börger. Java and the Java Virtual Machine: Definition, Verification, Validation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-642-63997-6. x + 381 pp. LCCN QA76.76.C65. URL http://deposit.d-

nb.de/cgi-bin/dokserv?id=4745749\%26prov=M\%26dok\

_var=1\%26dok\_ext=htm. Softcover reprint of [SSB01].

Sarimbekov:2016:WCJ

[SSB^{+}16] Aibek Sarimbekov, Lukas Stadler, Lubomír Bulej, Andreas Sewe, Andrej Podzimek, Yudi Zheng, and Walter Binder. Workload characterization of JVM languages. Software— Practice and Experience, 46(8):1053–1089, August 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Shojaei:2018:VVM

[SSEA18] Kiamars Shojaei, Faramarz Safi-Esfahani, and Saeed Ayat. VMDFS: virtual machine dynamic frequency scaling framework in cloud computing. The Journal of Supercomputing, 74(11):5944–5979, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Smith:1990:PTL

[SSG90] Robert Smith, Aaron Sloman, and John Gibson. POPLOG’s two-level virtual machine support for interactive languages. Cognitive Science Research Report 153, University-of-Sussex, Brighton, January 1990.

Saharan:2020:QEV

[SSG^{+}20] Shweta Saharan, Gaurav Somani, Gaurav Gupta, Robin Verma, Manoj Singh Gaur, and Rajkumar Buyya. QuickDedup: Efficient VM deduplication in cloud computing environments. Journal of Parallel and Distributed Computing, 139(??):18–31, May 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0743731519303442.

Srikanth:2017:CVU

[SSH17] Akhilesh Srikanth, Burak Sahin, and William R. Harris. Complexity verification using guided theorem enumeration. ACM SIGPLAN Notices, 52(1):639–652, January 2017. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Song:2013:PLM

[SSL^{+}13] Xiang Song, Jicheng Shi, Ran Liu, Jian Yang, and Haibo Chen. Parallelizing live migration of virtual machines. ACM SIGPLAN Notices, 48(7):85–96, July 2013. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Sciampacone:2010:EMS

[SSMGD10] R. A. Sciampacone, V. Sundaresan, D. Maier, and T. Gray-Donald. Exploitation of multicore systems in a Java virtual machine. IBM Journal of Research and Development, 54(5): 1:1–1:11, ???? 2010. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Stone:1994:PSO

[SSN94] L. C. Stone, S. B. Shukla, and B. Neta. Parallel satellite orbit prediction using a workstation cluster. Computers and Mathematics with Applications, 28(8):1–8, October 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-

7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0898122194001650.

Sharifi:2012:PED

[SSN12] Mohsen Sharifi, Hadi Salimi, and Mahsa Najafzadeh. Powerefficient distributed scheduling of virtual machines using workload-aware consolidation techniques. The Journal of Supercomputing, 61(1):46–66, July 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&
issn=0920-8542&volume=61&issue=1&spage=46.

Stefanovici:2017:TSS

[SSOT17] Ioan Stefanovici, Bianca Schroeder, Greg O’Shea, and Eno Thereska. Treating the storage stack like a network. ACM Transactions on Storage, 13(1):2:1–2:??, March 2017. CO-DEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Shabestari:2025:IVP

[SSS25] Ehsan Rasoulpour Shabestari and Alireza Shameli-Sendi. An intelligent VM placement method for minimizing energy cost and carbon emission in distributed cloud data centers. Journal of Grid Computing, 23(1):??, March 2025. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10723-025-09798-2.

Stoess:2012:LVM

[SSU^{+}12] Jan Stoess, Udo Steinberg, Volkmar Uhlig, Jens Kehne, Jonathan Appavoo, and Amos Waterland. A lightweight virtual machine monitor for Blue Gene/P. The International Journal of High Performance Computing Applications, 26 (2):95–109, May 2012. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://hpc.sagepub.

com/content/26/2/95.full.pdf+html.

Sadolewski:2024:VCS

[ST24] Jan Sadolewski and Bartosz Trybus. Verification of control system runtime using an executable semantic model. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/7/273.

Stankovic:1997:VRR

[Sta97] John A. Stankovic. Virtual roundtable: Real-time global virtual machines. IEEE Concurrency, 5(3):26–27, July/ September 1997. CODEN IECMFX. ISSN 1092-3063 (print),

1558-0849 (electronic). URL http://dlib.computer.org/
pd/books/pd1997/pdf/p3026.pdf.

Stanik:2007:NVR

[Sta07] John Stanik. News 2.0: Virtualization reconsidered; the power of PS3; Anti-P2P software targets universities. ACM Queue: Tomorrow’s Computing Today, 5(3):10, April 2007.

CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Steil:2005:MMM

[Ste05] Michael Steil. 17 mistakes Microsoft made in the Xbox security system. Report, Xbox Linux Project, December

2005. 13 pp. URL http://events.ccc.de/congress/2005/
fahrplan/attachments/591-paper_xbox.pdf.

Stecklina:2014:SHO

[Ste14] Julian Stecklina. Shrinking the hypervisor one subsystem at a time: a userspace packet switch for virtual machines. ACM SIGPLAN Notices, 49(7):189–200, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steinert:2015:OVS

[STFH15] Bastian Steinert, Lauritz Thamsen, Tim Felgentreff, and Robert Hirschfeld. Object versioning to support recovery needs: using proxies to preserve previous development states in lively. ACM SIGPLAN Notices, 50(2):113–124, February 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saber:2018:VRH

[STMV18] Takfarinas Saber, James Thorburn, Liam Murphy, and Anthony Ventresque. VM reassignment in hybrid clouds for large decentralised companies: A multi-objective challenge. Future Generation Computer Systems, 79 (part 2)(??):751– 764, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167739X17301164.

Stoess:2007:TEU

[Sto07] Jan Stoess. Towards effective user-controlled scheduling for microkernel-based systems. Operating Systems Review, 41(4): 59–68, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Strongin:2005:TCU

[Str05] Geoffrey Strongin. Trusted computing using AMD “Pacifica” and “Presidio” secure virtual machine technology. Information Security Technical Report, 10(2):120–132, ???? 2005.

CODEN ISTRFR. ISSN 1363-4127 (print), 1873-605X (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1363412705000257.

Strauss:2013:FCC

[Str13] David Strauss. The future cloud is container, not virtual machines. Linux Journal, 2013(228):5:1–5:??, April 2013. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Sun:2013:BJW

[STS^{+}13] Mengtao Sun, Gang Tan, Joseph Siefers, Bin Zeng, and Greg Morrisett. Bringing Java’s wild native world under control. ACM Transactions on Information and System Security, 16 (3):9:1–9:??, November 2013. CODEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Su:2014:RVP

[STY^{+}14] Tzu-Hsiang Su, Hsiang-Jen Tsai, Keng-Hao Yang, Po-Chun Chang, Tien-Fu Chen, and Yi-Ting Zhao. Reconfigurable vertical profiling framework for the Android runtime system. ACM Transactions on Embedded Computing Systems, 13(2s): 59:1–59:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Subramaniam:2008:PST

[Sub08] Venkat Subramaniam. Programming Scala: tackle multicore complexity on the JVM. Pragmatic Bookshelf, Raleigh, NC, USA, 2008. ISBN 1-934356-31-X (paperback). x + 221 pp. LCCN QA76.73.J38 S83 2008.

Subramaniam:2011:PCJ

[Sub11] Venkat Subramaniam. Programming concurrency on the JVM: mastering synchronization, STM, and actors. The pragmatic programmers. Pragmatic Bookshelf, Dallas, TX, 2011. ISBN 1-934356-76-X. xvii + 270 pp. LCCN QA76.73.J38 S8467 2011.

Samples:1986:SSB

[SUH86] A. Dain Samples, David Ungar, and Paul Hilfinger. SOAR: Smalltalk without bytecodes. ACM SIGPLAN Notices, 21 (11):107–118, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’86 Conference Proceedings, Norman Meyrowitz (editor), September 1986, Portland, Oregon.

Sun:1995:JVMb

[Sun95a] Sun Microsystems. Java Virtual Machine Profiling Interface (JVMPI). Sun Microsystems, 1995. URL http://java.sun. com/j2se/1.3/docs/guide/jvmpi/. http://java.sun.com/j2se/1.3/docs/guide/jvmpi/.

Sun:1995:JVMa

[Sun95b] Sun Microsystems. The Java Virtual Machine Specification, 1.0 beta edition, August 1995. URL http://java.sun.com/

doc/vmspec/VMSpec.ps.

Sun:1997:JCL

[SUN97] SUN Microsystems, Inc., Palo Alto/CA. Java Card 2.0 Language Subset and Virtual Machine Specification, revision 1.0 final edition, October 13, 1997. URL ftp://ftp.javasoft.

com/docs/javacard/JC20-Language.pdf.

Sun:1999:JPD

[Sun99] Sun Microsystems. Java Platform Debugger Architecture (JPDA). Sun Microsystems, Mountain View, CA, USA, 1999.

URL http://java.sun.com/j2se/1.3/docs/guide/jpda/.

Supnik:2004:SVM

[Sup04] Bob Supnik. Simulators: Virtual machines of the past (and future). ACM Queue: Tomorrow’s Computing Today, 2(5): 52–58, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Suri:2001:SCR

[Sur01] Niranjan Suri. State capture and resource control for Java: The design and implementation of the Aroma Virtual Machine. In USENIX [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL http://www.

usenix.org/publications/library/proceedings/jvm01/
JVM_wips/S15.pdf.

Suski:1976:AGC

[Sus76] Gregory J. Suski. Automatic generation of computer graphics languages. ACM SIGPLAN Notices, 11(6):113–122, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Simao:2013:ADQ

[SV13] José Simão and Luís Veiga. Adaptability driven by quality of execution in high level virtual machines for shared cloud environments. International Journal of Computer Systems Science and Engineering, 28(6):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Steindorfer:2015:OHA

[SV15] Michael J. Steindorfer and Jurgen J. Vinju. Optimizing hasharray mapped tries for fast and lean immutable JVM collections. ACM SIGPLAN Notices, 50(10):783–800, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steindorfer:2017:TSP

[SV17] Michael J. Steindorfer and Jurgen J. Vinju. Towards a software product line of trie-based collections. ACM SIG-PLAN Notices, 52(3):168–172, March 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shingarov:2024:TDV

[SV24] Boris Shingarov and Jan Vraný. Towards a Dynabook for verified VM construction. Journal of Computer Languages, 80(??):1–14, August 2024. CODEN ???? ISSN 2665-9182 (print), 2590-1184 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S2590118424000182.

Sebes:1993:MAL

[SVB93] E. J. Sebes and T. C. Vickers-Benzel. Modularity of assemblylanguage implementations of trusted systems. In Anonymous [Ano93], pages 173–184.

Sugerman:2001:VDV

[SVL01] Jeremy Sugerman, Ganesh Venkitachalam, and Beng-Hong Lim. Virtualizing I/O devices on VMware Workstation’s hosted virtual machine monitor. In USENIX [USE01a], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65

U84 2001. URL http://www.usenix.org/publications/
library/proceedings/usenix01/sugerman.html.

Sa:2023:CRV

[SVM^{+}23] Bruno Sá, Luca Valente, José Martins, Davide Rossi, Luca Benini, and Sandro Pinto. CVA6 RISC-V virtualization: Architecture, microarchitecture, and design space exploration. IEEE Transactions on Very Large Scale Integration (VLSI) Systems, 31(11):1713–1726, 2023. CODEN IEVSE9. ISSN 1063-8210 (print), 1557-9999 (electronic).

Scott:2010:SLV

[SVN^{+}10] Stephen L. Scott, Geoffroy Vallée, Thomas Naughton, Anand Tikotekar, Christian Engelmann, and Hong Ong. Systemlevel virtualization research at Oak Ridge National Laboratory. Future Generation Computer Systems, 26(3):304–307, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Swaine:2006:VR

[Swa06] Michael Swaine. Is virtualization real? Dr. Dobb’s Journal of Software Tools, 31(12):18–19, 21–22, December 2006. CODEN DDJOEB. ISSN 1044-789X.

Steinder:2008:SVA

[SWC08] Małgorzata Steinder, Ian Whalley, and David Chess. Server virtualization in autonomic management of heterogeneous workloads. Operating Systems Review, 42(1):94–95, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shan:2012:FIA

[SWcCM12] Zhiyong Shan, Xin Wang, Tzi cker Chiueh, and Xiaofeng Meng. Facilitating inter-application interactions for OS-level virtualization. ACM SIGPLAN Notices, 47(7):75–86, July 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Spink:2016:HAC

[SWF16] Tom Spink, Harry Wagstaff, and Björn Franke. Hardwareaccelerated cross-architecture full-system virtualization. ACM Transactions on Architecture and Code Optimization, 13(4):36:1–36:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Shih:2013:FSV

[SWH^{+}13] Chi-Sheng Shih, Jie-Wen Wei, Shih-Hao Hung, Joen Chen, and Norman Chang. Fairness scheduler for virtual machines on heterogeneous multi-core platforms. ACM SIGAPP Applied Computing Review, 13(1):28–40, March 2013. CO-DEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/
2460136.2460139.

Sun:2023:SID

[SWL^{+}23] Jie Sun, Tianyu Wo, Xudong Liu, Tianjiao Ma, Xudong Mou, Jinghong Lan, Nan Zhang, and Jianwei Niu. Scalable inter-domain network virtualization. Journal of Network and Computer Applications, 218(??):??, September 2023. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001200.

Song:2018:FRD

[SWW^{+}18] Tao Song, Jiajun Wang, Jiewei Wu, Ruhui Ma, Alei Liang, Tao Gu, and Zhengwei Qi. FastDesk: a remote desktop virtualization system for multi-tenant. Future Generation Computer Systems, 81(??):478–491, April 2018. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X17304776.

Song:2014:ARP

[SXCL14] Weijia Song, Zhen Xiao, Qi Chen, and Haipeng Luo. Adaptive resource provisioning for the cloud using online bin packing. IEEE Transactions on Computers, 63(11):2647–2660, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sha:2019:CED

[SXH^{+}19] Le-Tian Sha, Fu Xiao, Hai-Ping Huang, Yu Chen, and Ru-Chuan Wang. Catching escapers: a detection method for advanced persistent escapers in industry Internet of Things based on identity-based broadcast encryption (IBBE). ACM Transactions on Embedded Computing Systems, 18(3):

29:1–29:??, June 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3319615.

Sotiriou-Xanthopoulos:2018:OBV

[SXMX^{+}18] Efstathios Sotiriou-Xanthopoulos, Leonard Masing, Sotirios Xydis, Kostas Siozios, Jürgen Becker, and Dimitrios Soudris. OpenCL-based virtual prototyping and simulation of manyaccelerator architectures. ACM Transactions on Embedded Computing Systems, 17(5):86:1–86:??, November 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3242179.

Shuo:2012:PKR

[SYB12] Tian Shuo, He Yeping, and Ding Baozeng. Prevent kernel return-oriented programming attacks using hardware virtualization. Lecture Notes in Computer Science, 7232:289–300, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-29101-2_20/.

Song:2014:AFB

[SYC14] Xiang Song, Jian Yang, and Haibo Chen. Architecting flash-based solid-state drive for high-performance I/O virtualization. IEEE Computer Architecture Letters, 13(2):61–64, July/December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Sohrabi:2017:EEA

[SYMA17] Sahar Sohrabi, Yun Yang, Irene Moser, and Aldeida Aleti. Energy-efficient adaptive virtual machine migration mechanism for private clouds. Concurrency and Computation: Practice and Experience, 29(18), September 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Syropoulos:2007:PMV

[Syr07] Apostolos Syropoulos. Π machines: virtual machines realizing graph structured transition P systems. ACM SIGPLAN Notices, 42(12):15–22, December 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Savrun-Yeniceri:2014:EHI

[SYZZ^{+}14] Gülfem Savrun-Yeniçeri, Wei Zhang, Huahan Zhang, Eric Seckler, Chen Li, Stefan Brunthaler, Per Larsen, and Michael Franz. Efficient hosted interpreters on the JVM. ACM Transactions on Architecture and Code Optimization, 11(1): 9:1–9:24, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/2532642.

So:1988:PLV

[SZ88] K. So and V. Zecca. Program locality of vectorized applications running on the IBM 3090 with Vector Facility. IBM Systems Journal, 27(4):436–452, November 1988. CODEN IBMSA7. ISSN 0018-8670.

Stolyar:2013:LSS

[SZ13] Alexander L. Stolyar and Yuan Zhong. A large-scale service system with packing constraints: minimizing the number of occupied servers. ACM SIGMETRICS Performance Evaluation Review, 41(1):41–52, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sadegh:2021:TPV

[SZKY21] Samaneh Sadegh, Kamran Zamanifar, Piotr Kasprzak, and Ramin Yahyapour. A two-phase virtual machine placement policy for data-intensive applications in cloud. Journal of Network and Computer Applications, 180(??):??, April 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1084804521000515.

Su:2014:EAV

[SZL^{+}14] Sen Su, Zhongbao Zhang, Alex X. Liu, Xiang Cheng, Yiwen Wang, and Xinchao Zhao. Energy-aware virtual network embedding. IEEE/ACM Transactions on Networking, 22(5): 1607–1620, October 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Shi:2016:OAF

[SZW^{+}16] Weijie Shi, Linquan Zhang, Chuan Wu, Zongpeng Li, and Francis C. M. Lau. An online auction framework for dynamic resource provisioning in cloud computing. IEEE/ACM Transactions on Networking, 24(4):2060–2073, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Taft:2011:EPP

[Taf11] S. Tucker Taft. Experimenting with ParaSail: parallel specification and implementation language. ACM SIGADA Ada Letters, 31(3):11–12, December 2011. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Taivalsaari:1998:IJV

[Tai98] Antero Taivalsaari. Implementing a Java Virtual Machine in the Java programming language. Technical report, Sun Microsystems, Palo Alto, CA, USA, March 1998. 23 pp.

Taylor:1976:RRH

[Tay76] John McMay Taylor. Redundancy and recovery in the HIVE virtual machine. Report 76010, Procurement executive, Ministry of Defence, Royal Signals and Radar Establishment, London, UK, 1976.

Torlak:2014:LSV

[TB14] Emina Torlak and Rastislav Bodik. A lightweight symbolic virtual machine for solver-aided host languages. ACM SIG-PLAN Notices, 49(6):530–541, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tighe:2017:TAA

[TB17] Michael Tighe and Michael Bauer. Topology and application aware dynamic VM management in the cloud. Journal of Grid Computing, 15(2):273–294, June 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

https://link.springer.com/article/10.1007/s10723-
017-9397-z;   https://link.springer.com/content/pdf/
10.1007/s10723-017-9397-z.pdf.

Tsai:2017:JSD

[TBS17] Po-An Tsai, Nathan Beckmann, and Daniel Sanchez. Jenga: Software-defined cache hierarchies. ACM SIGARCH Computer Architecture News, 45(2):652–665, May 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Takemura:2010:BXP

[TC10] Chris Takemura and Luke S. (Luke Seidel) Crawford. The book of Xen: a practical guide for the system administrator. No Starch Press, San Francisco, CA, USA, 2010. ISBN 1-59327-186-7. xxiv + 281 pp. LCCN QA76.9.V5 C83

2009. URL http://proquest.safaribooksonline.com/?
fpi=9781593271862.

Trinder:2017:SRI

[TCP^{+}17] Phil Trinder, Natalia Chechina, Nikolaos Papaspyrou, Konstantinos Sagonas, Simon Thompson, Stephen Adams, Stavros Aronis, Robert Baker, Eva Bihari, Olivier Boudeville, Francesco Cesarini, Maurizio Di Stefano, Sverker Eriksson, Viktória Fördős, Amir Ghaffari, Aggelos Giantsios, Rickard Green, Csaba Hoch, David Klaftenegger, Huiqing Li, Kenneth Lundin, Kenneth Mackenzie, Katerina Roukounaki, Yiannis Tsiouris, and Kjell Winblad. Scaling reliably: Improving the scalability of the Erlang distributed actor platform. ACM Transactions on Programming Languages and Systems, 39(4):17:1–17:??, September 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Tournaire:2023:ECO

[TCTH23] Thomas Tournaire, Hind Castel-Taleb, and Emmanuel Hyon. Efficient computation of optimal thresholds in cloud autoscaling systems. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 8 (4):9:1–9:??, December 2023. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://dl.acm.org/

doi/10.1145/3603532.

Tarafdar:2020:EQS

[TDD20] Anurina Tarafdar, Mukta Debnath, and Rajib K. Das. Energy and quality of service-aware virtual machine consolidation in a cloud data center. The Journal of Supercomputing, 76(11):9095–9126, November 2020. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-020-03203-3.

Travostino:2006:SLM

[TDG^{+}06] Franco Travostino, Paul Daspit, Leon Gommans, Chetan Jog, Cees de Laat, Joe Mambretti, Inder Monga, Bas van Oudenaarde, Satish Raghunath, and Phil Yonghui Wang. Seamless live migration of virtual machines over the MAN/WAN. Future Generation Computer Systems, 22(8):901–907, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tan:2018:UVQ

[TDG^{+}18] Li Tan, Nathan DeBardeleben, Qiang Guan, Sean Blanchard, and Michael Lang. Using virtualization to quantify power conservation via near-threshold voltage reduction for inherently resilient applications. Parallel Computing, 74(??):3– 15, ???? 2018. CODEN PACOEJ. ISSN 0167-8191 (print),

1872-7336 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167819117300996.

Tapwal:2023:SBV

[TDMP23] Riya Tapwal, Pallav Kumar Deb, Sudip Misra, and Surjya Kanta Pal. Shadows: Blockchain virtualization for interoperable computations in IIoT environments. IEEE Transactions on Computers, 72(3):868–879, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tennenhouse:2017:RV

[Ten17] David Tennenhouse. Research at VMware. Operating Systems Review, 51(1):1–4, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Trajano:2016:TPL

[TF16] Alex F. R. Trajano and Marcial P. Fernandez. Two-phase load balancing of In-Memory Key-Value Storages using Network Functions Virtualization (NFV). Journal of Network and Computer Applications, 69(??):1–13, July 2016. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300789.

Tu:2015:CIE

[TFtLcC15] Cheng-Chun Tu, Michael Ferdman, Chao tung Lee, and Tzi cker Chiueh. A comprehensive implementation and evaluation of direct interrupt delivery. ACM SIGPLAN Notices, 50(7):1–15, July 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thomas:2008:DHF

[TGCF08] Gaël Thomas, Nicolas Geoffray, Charles Clément, and Bertil Folliot. Designing highly flexible virtual machines: the Jn-JVM experience. Software—Practice and Experience, 38(15): 1643–1675, December ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Troy:2010:VC

[TH10] Ryan Troy and Matthew Helmke. VMware cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2010. ISBN 0-596-15725-8 (paperback). xv + 280 pp. LCCN QA76.9.V5 T76 2010.

Tanenbaum:2006:CWM

[THB06] Andrew S. Tanenbaum, Jorrit N. Herder, and Herbert Bos. Can we make operating systems reliable and secure? Computer, 39(5):44–51, May 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://www.computer.org/csdl/mags/co/2006/05/
r5044-abs.html. Cover feature.

Thabet:2022:SBO

[THB22] Marwa Thabet, Brahim Hnich, and Mouhebeddine Berrima. A sampling-based online Co-Location-Resistant Virtual Machine placement strategy. The Journal of Systems and Software, 187(??):??, May 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0164121222000012.

Tu:2014:PPP

[THC^{+}14] Chia-Heng Tu, Hui-Hsin Hsu, Jen-Hao Chen, Chun-Han Chen, and Shih-Hao Hung. Performance and power profiling for emulated Android systems. ACM Transactions on Design Automation of Electronic Systems., 19(2):10:1–10:??, March 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tian:2018:MTE

[THG^{+}18] Wenhong Tian, Majun He, Wenxia Guo, Wenqiang Huang, Xiaoyu Shi, Mingsheng Shang, Adel Nadjaran Toosi, and Rajkumar Buyya. On minimizing total energy consumption in the scheduling of virtual machine reservations. Journal of Network and Computer Applications, 113(??):64–74, July 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1084804518301267.

Tan:2014:DBD

[THH^{+}14] Huailiang Tan, Lianjun Huang, Zaihong He, Youyou Lu, and Xubin He. DMVL: an I/O bandwidth dynamic allocation method for virtual networks. Journal of Network and Computer Applications, 39(??):104–116, March 2014. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001380.

Tikir:2003:RDS

[THL03] Mustafa M. Tikir, Jeffrey K. Hollingsworth, and Guei-Yuan Lueh. Recompilation for debugging support in a JITcompiler. ACM SIGSOFT Software Engineering Notes, 28 (1):10–17, January 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Thiruvathukal:2010:VCS

[THLK10] George K. Thiruvathukal, Konrad Hinsen, Konstantin Laufer, and Joe Kaylor. Virtualization for computational scientists. Computing in Science and Engineering, 12(4):52– 61, July/August 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thompson:1968:PTR

[Tho68] Ken Thompson. Programming techniques: Regular expression search algorithm. Communications of the ACM, 11(6): 419–422, June 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://patft.uspto. gov/. See also [KP99, Cox07, Cox09, Cox10, Cox12].

Thomas:1973:COA

[Tho73] Richard T. Thomas. Computer organization for allowing dynamic user microprogramming. ACM SIGMICRO Newsletter, 4(2):28–42, July 1973. CODEN SIGMDJ. ISSN 0163-

5751, 1050-916X. URL https://dl.acm.org/doi/10.1145/
1217124.1217129.

Thomas:1993:PIS

[Tho93] I. Thomas, editor. Proceedings of the 7th International Software Process Workshop: communication and coordination in the software process: October 15–18, 1991, Yountville, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4050-2, 0-8186-4051-0. LCCN QA76.755 .I58 1991.

Thorns:2008:VBK

[Tho08] Fabian Thorns, editor. Das Virtualisierungs-Buch: [Konzepte, Techniken und Lösungen: VMware, MS, Parallels, Xen u.v.a.]. Computer- und Literatur-Verlag, Böblingen, Germany, second edition, 2008. ISBN 3-936546-56-8. 799 pp. LCCN ????

Tickoo:2009:MVM

[TIIN09] Omesh Tickoo, Ravi Iyer, Ramesh Illikkal, and Don Newell. Modeling virtual machine performance: challenges and approaches. ACM SIGMETRICS Performance Evaluation Review, 37(3):55–60, December 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thakur:2020:MDV

[TK20] Dipanwita Thakur and Manas Khatua. Multi-domain virtual network embedding with dynamic flow migration in software-defined networks. Journal of Network and Computer Applications, 162(??):??, July 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804520301132.

Tetzlaff:1989:ABS

[TKG89] William H. Tetzlaff, Martin G. Kienzle, and Juan A. Garay. Analysis of block-paging strategies. IBM Journal of Research and Development, 33(1):51–59, January 1989. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Tuch:2012:BSV

[TLBW12] Harvey Tuch, Cyprien Laplace, Kenneth C. Barr, and Bi Wu. Block storage virtualization with commodity secure digital cards. ACM SIGPLAN Notices, 47(7):191–202, July 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Turner:2006:SIS

[TLC06] Stephen John Turner, Bu Sung Lee, and Wientong Cai, editors. Sixth International Symposium on Cluster Computing and the Grid CCGrid 06: 16–19 May, 2006, Singapore. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 0-7695-2585-7. LCCN QA76.9.C58. IEEE Computer Society Order Number P2585.

Thomas:1989:AMM

[TLD^{+}89] David A. Thomas, Wilf R. LaLonde, John Duimovich, Michael Wilson, Jeff McAffer, and Brian Barry. Actra — a multitasking/multiprocessing Smalltalk. ACM SIG-PLAN Notices, 24(4):87–90, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org:80/pubs/citations/
proceedings/plan/67386/p87-thomas/.

Tan:2017:EPP

[TLX17] Tian Tan, Yue Li, and Jingling Xue. Efficient and precise points-to analysis: modeling the heap by merging equivalent automata. ACM SIGPLAN Notices, 52(6):278–291, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tiburski:2021:LVM

[TMJ^{+}21] Ramão T. Tiburski, Carlos R. Moratelli, Sérgio F. Johann, Everton de Matos, and Fabiano Hessel. A lightweight virtualization model to enable edge computing in deeply embedded systems. Software—Practice and Experience, 51(9): 1964–1981, September 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tang:2014:DFS

[TMLL14] Zhuo Tang, Yanqing Mo, Kenli Li, and Keqin Li. Dynamic forecast scheduling algorithm for virtual machine placement in cloud computing environment. The Journal of Supercomputing, 70(3):1279–1296, December 2014. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11227-014-1227-5.

Tordsson:2012:CBM

[TMMVL12] Johan Tordsson, Rubén S. Montero, Rafael Moreno-Vozmediano, and Ignacio M. Llorente. Cloud brokering mechanisms for optimized placement of virtual machines across multiple providers. Future Generation Computer Systems, 28(2):358–367, February 2012. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X11001373.

Tavakoli:2012:FSC

[TMV12] Zahra Tavakoli, Sebastian Meier, and Alexander Vensmer. A framework for security context migration in a firewall secured virtual machine environment. Lecture Notes in Computer Science, 7479:41–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32808-4_5/.

Torquato:2025:ETB

[TMV25] Matheus Torquato, Paulo Maciel, and Marco Vieira. Evaluation of time-based virtual machine migration as moving target defense against host-based attacks. The Journal of Systems and Software, 219(??):??, January 2025. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121224002668.

Tollenaere:1991:SMN

[TO91] T. Tollenaere and G. A. Orban. Simulating modular neural networks on message-passing multiprocessors. Parallel Computing, 17(4–5):361–379, July 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tremblay:1996:PHI

[TO96] Marc Tremblay and Michael O’Connor. PicoJava: a hardware implementation of the Java Virtual Machine. In IEEE [IEE96a], pages 131–144. ISBN ???? LCCN ????

Tolksdorf:1998:PLJ

[Tol98] Robert Tolksdorf. Programming languages for the Java Virtual Machine. Technical report, Technische Universität Berlin, Fachbereich 13, Informatik, Formale Methoden, Logik und Programmierung (FLP), Sekr. FR 6-10, Franklinstraße 28/29, D-10587 Berlin, Germany, 1998. URL http://

grunge.cs.tu-berlin.de/~tolk/vmlanguages.html. World-

Wide Web document with pointers to more than 60 compilers and translators between various programming languages and Java.

Tikhe:2025:FAA

[TP25] Gajanan Nanaji Tikhe and Pushpinder Singh Patheja. Feature attention assisted convolutional stacked sparse autoencoder model for intrusion detection in network function virtualization environment. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002846.

Tucker:1988:AAC

[TR88] Lewis W. Tucker and George G. Robertson. Architecture and applications of the Connection Machine. Computer, 21 (8):26–38, August 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Treese:2005:VVE

[Tre05] Win Treese. Virtualization virtually everywhere. netWorker, 9(2):13–15, June 2005.

Thorat:2013:OMV

[TRG13] Nishant Thorat, Arvind Raghavendran, and Nigel Groves. Offline management in virtualized environments. Communications of the ACM, 56(4):75–81, April 2013. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tsafrir:2014:ELV

[Tsa14] Dan Tsafrir. Experiences in the land of virtual abstractions. ACM SIGPLAN Notices, 49(7):1–2, July 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Toosi:2019:EAS

[TSCB19] Adel Nadjaran Toosi, Jungmin Son, Qinghua Chi, and Rajkumar Buyya. ElasticSFC: Auto-scaling techniques for elastic service function chaining in network functions virtualization-based clouds. The Journal of Systems and Software, 152(??):108–119, June 2019. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121219300421.

Ta-Shma:2008:VMT

[TSLBYF08] Paula Ta-Shma, Guy Laden, Muli Ben-Yehuda, and Michael Factor. Virtual machine time travel using continuous data protection and checkpointing. Operating Systems Review, 42 (1):127–134, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Troia:2023:PCP

[TSN^{+}23] Sebastian Troia, Marco Savi, Giulia Nava, Ligia Maria Moreira Zorello, Thomas Schneider, and Guido Maier. Performance characterization and profiling of chained CPUbound Virtual Network Functions. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128623002608.

Tu:2017:BEO

[TSP17] Cheng-Chun Tu, Joe Stringer, and Justin Pettit. Building an extensible Open vSwitch datapath. Operating Systems Review, 51(1):72–77, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tavakoli-Someh:2019:MOV

[TSR19] Sanaz Tavakoli-Someh and Mohammad Hossein Rezvani. Multi-objective virtual network function placement using NSGA–II meta-heuristic approach. The Journal of Supercomputing, 75(10):6451–6487, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tsai:1993:LMM

[TT93] Shang Rong Tsai and Lian-Jou Tsai. A logical machine monitor supporting an environment for development and execution of operating systems. The Journal of Systems and Software, 21(1):27–39, April 1993. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tamm:1996:LBV

[TT96] Boris Tamm and Kuldar Taveter. A list-based virtual machine for COBOL. Software—Practice and Experience, 26 (12):1347–1371, December 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

//www3.interscience.wiley.com/cgi-bin/abstract?ID=
16778.

Tan:2019:VMC

[TTH^{+}19] Huailiang Tan, Yanjie Tan, Xiaofei He, Kenli Li, and Keqin Li. A virtual multi-channel GPU fair scheduling method for virtual machines. IEEE Transactions on Parallel and Distributed Systems, 30(2):257–270, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2019/02/08434359-abs.html.

Tu:2013:SDS

[TtLcC13] Cheng-Chun Tu, Chao tang Lee, and Tzi cker Chiueh. Secure I/O device sharing among virtual machines on multiple hosts. ACM SIGARCH Computer Architecture News, 41(3): 108–119, June 2013. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ICSA ’13 conference proceedings.

Thanh:1982:ITC

[tTR82] Nguyen the Thanh and E. Walter Raschner. Indirect threaded code used to emulate a virtual machine. ACM SIG-PLAN Notices, 17(5):80–89, May 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Torquato:2018:MAP

[TUM18] Matheus Torquato, I M Umesh, and Paulo Maciel. Models for availability and power consumption evaluation of a private cloud with VMM rejuvenation enabled by VM Live Migration. The Journal of Supercomputing, 74(9):4817–4841, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Turek:1984:IDV

[Tur84] John Joseph E. Turek. Issues in the design of a virtual network for the connection machine. Thesis (B.S.), Department of Electrical Engineering and Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, 1984. 67 pp. Supervised by Thomas Knight.

Turega:1992:CAS

[Tur92] M. Turega. A computer architecture to support neural net simulation. The Computer Journal, 35(4):353–360, August 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/Volume_35/Issue_04/Vol35_04.body.
html#AbstractTurega.

Tupakula:2012:DSB

[TV12] Udaya Kiran Tupakula and Vijay Varadharajan. Dynamic state-based security architecture for detecting security attacks in virtual machines. The Computer Journal, 55(4): 397–409, April 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/55/4/397.full.pdf+html.

Tsiftes:2018:VVS

[TV18] Nicolas Tsiftes and Thiemo Voigt. Velox VM: a safe execution environment for resource-constrained IoT applications. Journal of Network and Computer Applications, 118(??):61–73, September 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804518302017.

Toosi:2016:AMC

[TVKB16] Adel Nadjaran Toosi, Kurt Vanmechelen, Farzad Khodadadi, and Rajkumar Buyya. An auction mechanism for cloud spot markets. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(1):2:1–2:??, April 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Tollenaere:1992:PIC

[TVO92] Tom Tollenaere, Marc M. Van Hulle, and Guy A. Orban. Parallel implementation and capabilities of entropy-driven artificial neural networks. Journal of Parallel and Distributed Computing, 14(3):286–305, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tian:2024:OPM

[TXD^{+}24] Shu-Juan Tian, Ke-Ke Xu, Wen-Jian Ding, Yan-Chun Li, and De-Ze Zeng. An offloading and pricing mechanism based on virtualization in edge-cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128624003001.

Tien:2014:EOS

[TY14] Tsan-Rong Tien and Yi-Ping You. Enabling OpenCL support for GPGPU in kernel-based virtual machine. Software— Practice and Experience, 44(5):483–510, May 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tekinerdogan:2019:SIA

[TZB19] Bedir Tekinerdogan, Uwe Zdun, and M. Ali Babar. Special issue on architecting for hyper connectivity and hyper virtualization. The Journal of Systems and Software, 149(??):531– 532, March 2019. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0164121218302802.

Taheri:2017:VBB

[TZK17] Javid Taheri, Albert Y. Zomaya, and Andreas Kassler. vmBBProfiler: a black-box profiling approach to quantify sensitivity of virtual machines to shared cloud resources. Computing: Archiv für Informatik und Numerik, 99(12): 1149–1177, December 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s00607-017-
0552-y.pdf.

Tan:2025:HPR

[TZY^{+}25] Kai Tan, Dongyang Zhan, Lin Ye, Hongli Zhang, Binxing Fang, and Zhihong Tian. A high-performance real-time container file monitoring approach based on virtual machine introspection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL https://link.springer.com/
article/10.1007/s11227-024-06779-2.

Ungar:1998:PNC

[UBF^{+}98] David Ungar, Lars Bak, Jesse Fang, John Duimovich, and Scott Meyer. Panel 2: The new crop of Java Virtual Machines. ACM SIGPLAN Notices, 33(10):179–182, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Unger:1982:OSZ

[UBL^{+}82] B. Unger, D. Bidulock, G. Lomow, P. Belanger, C. Hankins, and N. Jain. An OASIS simulation of the ZNET microcomputer network. IEEE Micro, 2(3):70–84, July/September 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Uhlig:2006:F

[Uhl06] Rich Uhlig. Forward: Intel Virtualization Technology: Taking virtualization mainstream on Intel architecture platforms. Intel Technology Journal, 10(3):v–vi, August 10,

2006. ISSN 1535-766X. URL http://developer.intel.
com/technology/itj/2006/v10i3/foreword.htm.

Uhlig:2007:MKS

[Uhl07] Volkmar Uhlig. The mechanics of in-kernel synchronization for a scalable microkernel. Operating Systems Review, 41(4): 49–58, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ugawa:2019:EBJ

[UIK19] Tomoharu Ugawa, Hideya Iwasaki, and Takafumi Kataoka. eJSTK: Building JavaScript virtual machines with customized datatypes for embedded systems. Journal of Computer Languages, 51(??):261–279, April 2019. CODEN ???? ISSN 2665-9182 (print), 2590-1184 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1045926X18302416.

Uhlig:2005:IVT

[UNR^{+}05] Rich Uhlig, Gil Neiger, Dion Rodgers, Amy L. Santoni, Fernando C. M. Martins, Andrew V. Anderson, Steven M. Bennett, Alain Kägi, Felix H. Leung, and Larry Smith. Intel virtualization technology. Computer, 38(5):48–??, May 2005.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.org/comp/mags/co/
2005/05/r5048abs.htm;   http://csdl.computer.org/
dl/mags/co/2005/05/r5048.pdf.

Uehara:1984:BPB

[UOKT84] K. Uehara, R. Ochitani, O. Kakusho, and J. Toyoda. A bottom-up parser base on predicate logic: a survey of the formalism and its implementation techniques. In 1984 International Symposium on Logic Programming. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. ISBN 0-8186-0522-7. CH2007-3/84/0000-0220$01.00.

Upadhyaya:2015:EML

[UR15] Ganesha Upadhyaya and Hridesh Rajan. Effectively mapping linguistic abstractions for message-passing concurrency to threads on the Java Virtual Machine. ACM SIGPLAN Notices, 50(10):840–859, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ugawa:2018:TSL

[URJ18] Tomoharu Ugawa, Carl G. Ritson, and Richard E. Jones. Transactional Sapphire: Lessons in high-performance, onthe-fly garbage collection. ACM Transactions on Programming Languages and Systems, 40(4):15:1–15:??, December 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Qaiser:2020:NEB

[uRQS20] Hammad ur Rehman Qaiser and Gao Shu. Nash equilibrium based replacement of virtual machines for efficient utilization of cloud data centers. Computing: Archiv für Informatik und Numerik, 102(12):2521–2540, December 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

USENIX:1985:SCP

[USE85] USENIX Association, editor. Summer conference proceedings, Portland 1985: June 11–14, 1985, Portland, Oregon, USA. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1985. LCCN QA76.8.U65 U8 1985.

USENIX:1986:SCP

[USE86] USENIX Association, editor. Summer conference proceedings, Atlanta 1986: June 9–13, 1986, Atlanta, Georgia, USA. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1986.

USENIX:1991:PUM

[USE91] USENIX, editor. Proceedings of the USENIX Mach Symposium: November 20–22, 1991, Monterey, California, USA. USENIX, San Francisco, CA, USA, 1991. LCCN QAX 27.

USENIX:1993:PUM

[USE93] USENIX, editor. Proceedings of the USENIX Mobile and Location-Independent Computing Symposium: August 2– 3, 1993, Cambridge, Massachusetts, USA. USENIX, San Francisco, CA, USA, 1993. ISBN 1-880446-51-0. LCCN

QA 76.76 O63 U86 1993. URL http://www.usenix.org/
publications/library/proceedings/mobile93/.

USENIX:1999:PFU

[USE99] USENIX, editor. Proceedings of the fifth USENIX Conference on Object-Oriented Technologies and Systems (COOTS ’99): May 3–7, 1999, San Diego, California, USA. USENIX, San Francisco, CA, USA, 1999. ISBN ???? LCCN QA76.64

.U84 1999. URL http://www.usenix.org/publications/
library/proceedings/coots99/.

USENIX:2000:PAL

[USE00a] USENIX, editor. Proceedings of the 4th Annual Linux Showcase and Conference, Atlanta, October 10–14, 2000, Atlanta, Georgia, USA. USENIX, San Francisco, CA, USA, 2000. ISBN 1-880446-17-0. LCCN ???? URL http://www.usenix.

org/publications/library/proceedings/als2000/.

USENIX:2000:PNU

[USE00b] USENIX, editor. Proceedings of the Ninth USENIX Security Symposium, August 14–17, 2000, Denver, Colorado. USENIX, San Francisco, CA, USA, 2000. ISBN 1-880446-

18-9. LCCN ????  URL http://www.usenix.org/
publications/library/proceedings/sec2000.

USENIX:2001:PUA

[USE01a] USENIX, editor. Proceedings of the 2001 USENIX Annual Technical Conference: June 25–30, 2001, Marriott Copley Place Hotel, Boston, Massachusetts, USA. USENIX, San Francisco, CA, USA, 2001. ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL http:

//www.usenix.org/publications/library/proceedings/
usenix01/technical.html.

USENIX:2001:PUC

[USE01b] USENIX, editor. Proceedings of the 6th USENIX Conference on Object-Oriented Technologies and Systems, January 29–February 2, 2001, San Antonio, Texas, USA. USENIX, San Francisco, CA, USA, 2001. ISBN 1-880446-12-X.

LCCN ???? URL http://www.usenix.org/publications/
library/proceedings/coots01/.

USENIX:2001:PJV

[USE01c] USENIX, editor. Proceedings of the Java Virtual Machine Research and Technology Sy[m]posium (JVM ’01): April 23–24, 2001, Monterey, California, USA. Berkeley, CA. USENIX, San Francisco, CA, USA, 2001. ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL http://www.usenix.

org/publications/library/proceedings/jvm01/.

USENIX:2001:UJV

[USE01d] USENIX, editor. Usenix Java Virtual Machine Research and Technology Symposium (JVM ’01). USENIX, San Francisco, CA, USA, April 2001.

USENIX:2002:PJV

[USE02] USENIX, editor. Proceedings of the Java Virtual Machine Research and Technology Symposium (JVM ’02): August 1–2, 2002, San Francisco, California, US. USENIX, San Francisco, CA, USA, August 1–2, 2002. ISBN 1-931971-01-3. LCCN QA76.73 .J38 J42 2002. URL http://www.usenix.

org/publications/library/proceedings/javavm02/.

USENIX:2006:PUA

[USE06] USENIX, editor. Proceedings of the 2006 USENIX Annual Technical Conference: May 30–June 3, 2006, Boston, MA, USA. USENIX, San Francisco, CA, USA, 2006. ISBN 1-931971-44-7. LCCN ????

Umeno:1987:NMR

[UT87] Hidenori Umeno and Shunji Tanaka. New methods for realizing plural near-native performance virtual machines. IEEE Transactions on Computers, C-36(9):1076–1087, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956

(electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009538.

Ureche:2013:MIS

[UTO13] Vlad Ureche, Cristian Talau, and Martin Odersky. Miniboxing: improving the speed to code size tradeoff in parametric polymorphism translations. ACM SIGPLAN Notices, 48(10): 73–92, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Unnikrishnan:2013:RDP

[UVL^{+}13] Deepak Unnikrishnan, Ramakrishna Vadlamani, Yong Liao, Jeremie Crenne, Lixin Gao, and Russell Tessier. Reconfigurable data planes for scalable network virtualization. IEEE Transactions on Computers, 62(12):2476–2488, 2013. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vachon:2006:DBV

[Vac06] Travis Vachon. Dynamic balancing of virtual operating systems. Journal of Computing Sciences in Colleges, 21(6):309– 310, June 2006. CODEN ???? ISSN ????

Vaghani:2010:VMF

[Vag10] Satyam B. Vaghani. Virtual machine file system. Operating Systems Review, 44(4):57–70, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vanhelsuwe:1998:BRJb

[Van98] Laurence Vanhelsuwé. Book review: Java virtual machine books — a comparative review. JavaWorld: IDG’s magazine for the Java community, 3(3):??, March 1998. CODEN

???? ISSN 1091-8906. URL http://www.javaworld.com/
javaworld/jw-03-1998/jw-03-bookreview.htm.

VanHensbergen:2006:PRP

[Van06] Eric Van Hensbergen. P.R.O.S.E.: partitioned reliable operating system environment. Operating Systems Review, 40 (2):12–15, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Varian:1991:VVC

[Var91] Melinda Varian. VM and the VM community: Past, present, and future. Technical report, Office of Computing and Information Technology, Princeton University, Princeton, NJ 08544, USA, April 1991. 168 pp. URL http://www.

leeandmelindavarian.com/Melinda/neuvm.pdf. Original

presented at Australasian SHARE/GUIDE in Melbourne, Victoria, Australia in 1989. This is a detailed history of the development of virtual machine technology on IBM System/ 360 and later mainframes, and of the opposition by much of IBM to that technology until it was demonstrated that their performance could equal, or even exceed, that of an operating system running on bare hardware, and also allow a single physical host to support multiple operating systems, and software development, simultaneously. There are also several comments about the development of the REXX language, and about the influence of Unix on IBM’s software development.

vanCleeff:2011:IAM

[vCPWvT11] André van Cleeff, Wolter Pieters, Roel Wieringa, and Frits van Tiel. Integrated assessment and mitigation of physical and digital security threats: Case studies on virtualization. Information Security Technical Report, 16(3–4):142–149, August 2011. CODEN ISTRFR. ISSN 1363-4127 (print), 1873-

605X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1363412711000483.

vanDoorn:2000:SJV

[vD00] Leendert van Doorn. A secure Java^{T M} Virtual Machine. In USENIX [USE00b], page ?? ISBN 1-880446-18-9.

LCCN ???? URL http://www.usenix.org/publications/
library/proceedings/sec2000/vandoorn.html.

vanDoorn:2006:HVT

[vD06] Leendert van Doorn. Hardware virtualization trends. In ACM [ACM06f], page 45. ISBN 1-59593-332-6 (??invalid ISBN??). LCCN QA76.9.V4.

vanderKouwe:2009:PQV

[vdK09] Erik van der Kouwe. Porting the QEMU virtualization software to MINIX 3. Master’s thesis, ????, ????, ???? 2009.

???? pp. URL http://www.few.vu.nl/%7Evdkouwe/doc/
msc-thesis-cs-presentation-erik-van-der-kouwe.pdf;
http://www.minix3.org/theses/kouwe-qemu.pdf.

Villadeamigo:1997:EES

[VdlFCC97] José M. Pérez Villadeamigo, Santiago Rodríguez de la Fuente, Rafael Méndez Cavanillas, and M. Isabel García Clemente. The em88110: emulating a superscalar processor. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):45–50, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Visegrady:2014:SCV

[VDO14] T. Visegrady, S. Dragone, and M. Osborne. Stateless cryptography for virtual environments. IBM Journal of Research and Development, 58(1):5:1–5:10, January–February 2014. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Venstermans:2006:BVB

[VED06] Kris Venstermans, Lieven Eeckhout, and Koen De Bosschere. 64-bit versus 32-bit Virtual Machines for Java. Software— Practice and Experience, 36(1):1–26, January 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Venstermans:2007:JOH

[VED07] Kris Venstermans, Lieven Eeckhout, and Koen De Bosschere. Java object header elimination for reduced memory consumption in 64-bit virtual machines. ACM Transactions on Architecture and Code Optimization, 4(3):17:1–17:??, September 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Venners:1996:UHL

[Ven96] Bill Venners. Under the hood: The lean, mean, virtual machine. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CODEN ???? ISSN 1091-

8906. URL http://www.javaworld.com/javaworld/jw-06-
1996/jw-06-vm.htm.

Venners:1997:IJV

[Ven97a] Bill Venners. Inside the Java Virtual Machine. Mc-Graw-Hill, New York, NY, USA, November 1997. ISBN 0-07-913248-0. 384 pp. LCCN QA76.73.J38 V46 1998.

US$39.95. URL http://mcgraw-hill.inforonics.com/
cgi/getarec?mgh31406%comp.

Venners:1997:UHHa

[Ven97b] Bill Venners. Under the hood: How the Java virtual machine handles exceptions. JavaWorld: IDG’s magazine for the Java community, 2(1):??, January 1997. CODEN ???? ISSN 1091-

8906. URL http://www.javaworld.com/javaworld/jw-01-
1997/jw-01-hood.htm.

Venners:1997:UHHb

[Ven97c] Bill Venners. Under the hood: How the Java virtual machine handles method invocation and return. JavaWorld: IDG’s magazine for the Java community, 2(6):??, June 1997. CO-DEN ???? ISSN 1091-8906. URL http://www.javaworld.

com/javaworld/jw-06-1997/jw-06-hood.htm.

Venners:1997:UHHc

[Ven97d] Bill Venners. Under the hood: How the Java virtual machine performs thread synchronization. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CODEN

???? ISSN 1091-8906. URL http://www.javaworld.com/
javaworld/jw-07-1997/jw-07-hood.htm.

Venners:1999:IJV

[Ven99a] Bill Venners. Inside the Java Virtual Machine. McGraw-Hill, New York, NY, USA, second edition, 1999. ISBN 0-07-135093-4. xxxi + 703 pp. LCCN QA76.73.J38 V46

1999. URL http://www.loc.gov/catdir/description/
mh024/00269375.html.

Venners:1999:SJV

[Ven99b] Bill Venners. The state of the Java virtual machine (JVM). JavaWorld: IDG’s magazine for the Java community, 4(1): ??, January 1999. CODEN ???? ISSN 1091-8906. URL http:

//www.javaworld.com/javaworld/jbe/jw-jbe-jvm.htm.

Veglis:2020:SEO

[VG20] Andreas Veglis and Dimitrios Giomelakis. Search engine optimization. Future Internet, 12(1):6, December 31, 2020. CO-DEN ???? ISSN 1999-5903. URL https://www.mdpi.com/

1999-5903/12/1/6.

Vinco:2016:CMV

[VGF16] Sara Vinco, Valerio Guarnieri, and Franco Fummi. Code manipulation for virtual platform integration. IEEE Transactions on Computers, 65(9):2694–2708, September 2016. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vila:2023:CCV

[VGL23] Sergi Vila, Fernando Guirado, and Josep L. Lérida. Cloud computing virtual machine consolidation based on stock trading forecast techniques. Future Generation Computer Systems, 145(??):321–336, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X23000961.

vonHagen:2008:PXV

[vH08] William von Hagen. Professional Xen virtualization. Wiley, New York, NY, USA, 2008. ISBN 0-470-13811-4 (paperback), 0-470-28918-X (electronic). xxiii + 405 pp. LCCN QA76.9.V5 V665 2008. URL http://

www.loc.gov/catdir/enhancements/fy0806/2007049359-
d.html;   http://www.loc.gov/catdir/enhancements/
fy0806/2007049359-t.html;   http://www.loc.gov/catdir/
enhancements/fy0809/2007049359-b.html.

Vitek:2014:CTR

[Vit14] Jan Vitek. The case for the three R’s of systems research: repeatability, reproducibility and rigor. ACM SIGPLAN Notices, 49(7):115–116, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vonKoch:2013:LRB

[vKF13] Tobias J. K. Edler von Koch and Björn Franke. Limits of region-based dynamic binary parallelization. ACM SIGPLAN Notices, 48(7):13–22, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Viswanathan:2000:JVM

[VL00] D. Viswanathan and S. Liang. Java Virtual Machine Profiler Interface. IBM Systems Journal, 39(1):82–95, ???? 2000.

CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/sj/391/viswanathan.html.

vonLaszewski:2001:GBA

[vLSM01] Gregor von Laszewski, Kazuyuki Shudo, and Yoichi Muraoka. Grid-based asynchronous migration of execution context in Java virtual machines. Lecture Notes in Computer Science, 1900:22–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-ny.

com/link/service/series/0558/bibs/1900/19000022.htm;
 http://link.springer-ny.com/link/service/series/
0558/papers/1900/19000022.pdf.

Varvello:2016:MPC

[VLZL16] Matteo Varvello, Rafael Laufer, Feixiong Zhang, and T. V. Lakshman. Multilayer packet classification with graphics processing units. IEEE/ACM Transactions on Networking, 24 (5):2728–2741, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

vanMoolenbroek:2014:TFL

[vMAT14] David C. van Moolenbroek, Raja Appuswamy, and Andrew S. Tanenbaum. Towards a flexible, lightweight virtualization alternative. In ????, editor, SYSTOR ’14: proceedings of the 7th ACM International Systems and Storage Conference: June 10–12, 2014, Haifa, Israel, pages 1–7. ACM Press, New York, NY 10036, USA, 2014. ISBN 1-4503-2920-9.

LCCN ???? URL http://dl.acm.org/citation.cfm?id=
2611354;   http://www.minix3.org/docs/conf/systor-
2014.pdf.

Vicente:2012:ECS

[VMBM12] Elder Vicente, Rivalino Matias, Lúcio Borges, and Autran Macêdo. Evaluation of compound system calls in the Linux kernel. Operating Systems Review, 46(1):53–63, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bulck:2019:BVM

[VMW^{+}19] J. Van Bulck, M. Minkin, O. Weisse, D. Genkin, B. Kasikci, F. Piessens, M. Silberstein, T. F. Wenisch, Y. Yarom, and R. Strackx. Breaking virtual memory protection and the SGX ecosystem with foreshadow. IEEE Micro, 39(3):66–74, May/ June 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Vaughan-Nichols:2006:NAV

[VN06] Stephen J. Vaughan-Nichols. New approach to virtualization is a lightweight. Computer, 39(11):12–14, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.org/comp/mags/co/
2006/11/ry012.pdf.

Vaughan-Nichols:2008:VSS

[VN08] Steven J. Vaughan-Nichols. Virtualization sparks security concerns. Computer, 41(8):13–15, August 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voelcker:1986:MYP

[Voe86] J. Voelcker. Making your PC behave like another. IEEE Spectrum, 23(10):61–67, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vogels:2003:HNC

[Vog03] Werner Vogels. HPC.NET — are CLI-based virtual machines suitable for high performance computing? In ACM [ACM03a], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.sc-conference.org/sc2003/
inter_cal/inter_cal_detail.php?eventid=10710#2; http:
//www.sc-conference.org/sc2003/paperpdfs/pap251.pdf.

Volz:1990:VNU

[Vol90] Richard A. Volz. Virtual nodes and units of distribution for distributed Ada. ACM SIGADA Ada Letters, 10(4):85–96, Spring 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Voith:2012:QSP

[VOS12] Thomas Voith, Karsten Oberle, and Manuel Stein. Quality of service provisioning for distributed data center interconnectivity enabled by network virtualization. Future Generation Computer Systems, 28(3):554–562, March 2012. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X11000392.

Verdu:2016:PSA

[VP16] Javier Verdu and Alex Pajuelo. Performance scalability analysis of JavaScript applications with Web Workers. IEEE Computer Architecture Letters, 15(2):105–108, July/ December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Vrable:2005:SPA

[Vra05] Michael Daniel Vrable, editor. SOSP ’05: proceedings of the 20th ACM Symposium on Operating Systems Principles: October 23–26, 2005, Brighton, United Kingdom, volume 39, no. 5 (Dec. 2005) of Operating Systems Review. ACM Press, New York, NY 10036, USA, 2005. ISBN 1-59593-079-5. LCCN QA76.6 .S9196 2005; QA76.6; QA76.6 .S9196 2005eb; Internet.

Vallee:2006:OTX

[VS06] Geoffroy Vallee and Stephen L. Scott. OSCAR testing with Xen. In IEEE [IEE06a], pages 43–? ISBN 0-7695-2582-2. ISSN 1550-5243. LCCN QA76.88. ACM product number E2582.

Varshney:2019:ARC

[VS19] P. Varshney and Y. Simmhan. AutoBoT: Resilient and costeffective scheduling of a bag of tasks on spot VMs. IEEE Transactions on Parallel and Distributed Systems, 30(7): 1512–1527, July 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Victor:2010:OSS

[VSC^{+}10] Jeff Victor, Jeff Savit, Gary Combs, Simon Hayler, and Bob Netherton, editors. Oracle Solaris 10 system virtualization essentials. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2010. ISBN 0-13-708188-X (paperback). 384 (est.) pp. LCCN QA76.9.V5 O73 2010.

Valsamas:2023:VTB

[VSMC23] Polychronis Valsamas, Sotiris Skaperas, Lefteris Mamatas, and Luis M. Contreras. Virtualization Technology Blending for resource-efficient edge clouds. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128623000919.

vanSchaik:2020:CLD

[vSMK^{+}20] Stephan van Schaik, Marina Minkin, Andrew Kwong, Daniel Genkin, and Yuval Yarom. CacheOut: Leaking data on Intel CPUs via cache evictions. Report, University of Michigan and University of Adelaide and Data61, Ann Arbor, MI, USA and Adelaide, Australia, January 27, 2020. 16 pp. URL https:/

/cacheoutattack.com/CacheOut.pdf.

Varadharajan:2014:CSA

[VT14] Vijay Varadharajan and Udaya Tupakula. Counteracting security attacks in virtual machines in the cloud using property based attestation. Journal of Network and Computer Applications, 40(??):31–45, April 2014. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001768.

Venkatesan:2016:SCA

[VTW16] Vimalraj Venkatesan, Y. C. Tay, and Qingsong Wei. Sizing cleancache allocation for virtual machines’ transcendent memory. IEEE Transactions on Computers, 65(6):1949–1963, ???? 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

VanDijkhuizen:2018:SNT

[VV18] Niels Van Dijkhuizen and Jeroen Van Der Ham. A survey of network traffic anonymisation techniques and implementations. ACM Computing Surveys, 51(3):52:1–52:??, July 2018. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Verboven:2013:BBS

[VVB13] Sam Verboven, Kurt Vanmechelen, and Jan Broeckhove. Black box scheduling for resource intensive virtual machine workloads with interference models. Future Generation Computer Systems, 29(8):1871–1884, October 2013. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X1300099X.

Vissicchio:2017:SUH

[VVC^{+}17] Stefano Vissicchio, Laurent Vanbever, Luca Cittadini, Geoffrey G. Xie, and Olivier Bonaventure. Safe update of hybrid SDN networks. IEEE/ACM Transactions on Networking, 25 (3):1649–1662, June 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Varman:2008:SVP

[VW08] Peter Varman and Jun Wang. Storage and I/O virtualization, performance, energy, evaluation and dependability (SPEED08). Operating Systems Review, 42(6):1–2, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Versick:2013:PCE

[VWT13] Daniel Versick, Ingolf Waßmann, and Djamshid Tavangarian. Power consumption estimation of CPU and peripheral components in virtual machines. ACM SIGAPP Applied Computing Review, 13(3):17–25, September 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:

//dl.acm.org/doi/abs/10.1145/2537728.2537730.

Wakeling:1999:CLF

[Wak99] David Wakeling. Compiling lazy functional programs for the Java virtual machine. Journal of Functional Programming, 9 (6):579–603, November 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/AC59FAFC68063D9EC71E4FCF283E7EFC.

Wallace:1976:SGI

[Wal76] Victor L. Wallace. The semantics of graphic input devices. ACM SIGPLAN Notices, 11(6):61–65, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walters:1999:VVP

[Wal99] Brian Walters. VMware virtual platform. Linux Journal, 63:??, July 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://noframes.

linuxjournal.com/lj-issues/3458.html.

Waldspurger:2002:MRM

[Wal02] Carl A. Waldspurger. Memory resource management in VMware ESX server. Operating Systems Review, 36(5S):181– 194, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Waldspurger:2010:VEM

[Wal10] Carl Waldspurger. A VM ’engine’ that makes a difference: technical perspective. Communications of the ACM, 53(10): 84, October 2010. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Warren:1980:IPI

[War80] D. H. D. Warren. An improved Prolog implementation which optimises tail recursion. Research Paper 156, Department of AI, Univeristy of Edinburgh, 1980.

Ward:2002:BVC

[War02] Brian Ward. The book of VMware: the complete guide to VMware workstation. No Starch Press, San Francisco, CA, USA, 2002. ISBN 1-886411-72-7. xv + 249 pp. LCCN QA76.76.O63 W3653 2002. US$39.95.

Warren:2005:VWH

[War05] Steven S. Warren. The VMWare Workstation 5 Handbook. Charles River Media, Hingham, MA, USA, 2005. ISBN 1-58450-393-9. xvii + 334 pp. LCCN QA76.76.O63 W3665

2005. URL http://www.loc.gov/catdir/toc/ecip0510/
2005010053.html.

Ward:2011:KRC

[War11] Bob Ward. Kuck receives Computer Pioneer Award; VMware’s Greene and Rosenblum win Computer Entrepreneur Award; dozens of new CSDP and CSDA holders named. Computer, 44(6):89–91, June 2011. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Watson:1986:PRL

[Wat86] Paul Watson. The Parallel Reduction of Lambda Calculus Expression. PhD thesis, University of Manchester, July 1986. 133 pp. Also available as UMCS-87-2-1.

Watson:1987:PRL

[Wat87] Paul Watson. The Parallel Reduction of Lambda Calculus Expression. Ph.D. thesis, University of Manchester, Computer Science Department, February 1987. URL http:/

/www.cs.man.ac.uk/csonly/cstechrep/Abstracts/UMCS-
87-2-1.html;   mailto:techreports@cs.man.ac.uk.

Wang:1981:VMB

[WB81] Richard T. Wang and James C. Browne. Virtual machinebased simulation of distributed computing and network computing. In ACM [ACM81], pages 154–156. ISBN ???? LCCN ????

Wei:2016:PVR

[WB16] Yi Wei and M. Brian Blake. Proactive virtualized resource management for service workflows in the cloud. Computing: Archiv für Informatik und Numerik, 98(5):523–538, May 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Waldspurger:2016:SSL

[WBB^{+}16] Carl Waldspurger, Emery Berger, Abhishek Bhattacharjee, Kevin Pedretti, Simon Peter, and Chris Rossbach. Sweet spots and limits for virtualization. ACM SIGPLAN Notices, 51(7):177, July 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2018:HSA

[WBHN18] Kunshan Wang, Stephen M. Blackburn, Antony L. Hosking, and Michael Norrish. Hop, skip, & jump: Practical on-stack replacement for a cross-platform language-neutral VM. ACM SIGPLAN Notices, 53(3):1–16, March 2018. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2019:VTV

[WBW^{+}19] Sen Wang, Jun Bi, Jianping Wu, Athanasios V. Vasilakos, and Qilin Fan. VNE-TD: a virtual network embedding algorithm based on temporal-difference learning. Computer Networks (Amsterdam, Netherlands: 1999), 161(??):251–263, October 9, 2019. CODEN ???? ISSN 1389-1286 (print),

1872-7069 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S138912861830584X.

Wu:1991:NNS

[WC91] J.-. L. Wu and S.-T. Cheng. A neural network simulating system. The Computer Journal, 34(1):92–93, February 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_01/tiff/92.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_34/Issue_01/
tiff/93.tif.

Welsh:2001:VCH

[WC01] M. Welsh and D. Culler. Virtualization considered harmful: OS design directions for well-conditioned services. In IEEE [IEE01], pages 139–146. ISBN 0-7695-1040-X.

US$135.00. URL http://computer.org/CSPRESS/CATALOG/

pr01040.htm. IEEE catalog number PR01040.

Wang:2016:ECA

[WCC^{+}16a] Bei Wang, Yuxia Cheng, Wenzhi Chen, Qinming He, Yang Xiang, Mohammad Mehedi Hassan, and Abdulhameed Alelaiwi. Efficient consolidation-aware VCPU scheduling on multicore virtualization platform. Future Generation Computer Systems, 56(??):229–237, March 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X15002605.

Wang:2016:CIK

[WCC16b] Huang Wang, Xianglan Chen, and Huaping Chen. A cross-ISA kernelized high-performance parallel emulator. International Journal of Parallel Programming, 44(6):1118–1141, December 2016. CODEN IJPPE5. ISSN 0885-7458 (print),

1573-7640 (electronic). URL http://link.springer.com/
article/10.1007/s10766-015-0379-0.

Wang:2016:TSN

[WCC16c] Yang Wang, Phanvu Chau, and Fuyu Chen. Towards a secured network virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 104(??):55–65, July 20, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128616301219.

Wang:2020:ESF

[WCC20] Meng Wang, Bo Cheng, and Junliang Chen. An efficient service function chaining placement algorithm in mobile edge computing. ACM Transactions on Internet Technology (TOIT), 20(4):32:1–32:21, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.1145/3388241.

Wood:2014:LLD

[WCG14] Benjamin P. Wood, Luis Ceze, and Dan Grossman. Lowlevel detection of language-level data races with LARD. ACM SIGARCH Computer Architecture News, 42(1):671– 686, March 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Wu:2021:SLS

[WCG21] Hao Wu, Xin Chen, and He Guo. Scheduling large-scale scientific workflow on virtual machines with different numbers of vCPUs. The Journal of Supercomputing, 77(1):679–710, January 2021. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL https://link.springer.com/
article/10.1007/s11227-020-03273-3.

Wells:2006:HSS

[WCS06] Philip M. Wells, Koushik Chakraborty, and Gurindar S. Sohi. Hardware support for spin management in overcommitted virtual machines. In ACM [ACM06b], pages 124–133. ISBN 1-59593-264-X. LCCN ????

Wells:2009:DHN

[WCS09] Philip M. Wells, Koushik Chakraborty, and Gurindar S. Sohi. Dynamic heterogeneity and the need for multicore virtualization. Operating Systems Review, 43(2):5–14, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Whitaker:2005:RDV

[WCSG05] Andrew Whitaker, Richard S. Cox, Marianne Shaw, and Steven D. Gribble. Rethinking the design of virtual machine monitors. Computer, 38(5):57–??, May 2005. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.org/comp/mags/co/

2005/05/r5057abs.htm;   http://csdl.computer.org/
dl/mags/co/2005/05/r5057.pdf.

Wang:2017:DCT

[WCY^{+}17] Xiumin Wang, Xiaoming Chen, Chau Yuen, Weiwei Wu, Meng Zhang, and Cheng Zhan. Delay-cost tradeoff for virtual machine migration in cloud data centers. Journal of Network and Computer Applications, 78(??):62–72, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1084804516302739.

Wu:2023:LPC

[WCZ^{+}23] Hao Wu, Yuqi Chen, Chi Zhang, Jiangchao Dong, and Yuxin Wang. Loads prediction and consolidation of virtual machines in cloud. Concurrency and Computation: Practice and Experience, 35(23):e7760:1–e7760:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2008:VBA

[WDCL08] Xiaoying Wang, Zhihui Du, Yinong Chen, and Sanli Li. Virtualization-based autonomic resource management for multi-tier Web applications in shared data center. The Journal of Systems and Software, 81(9):1591–1608, September 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wan:2020:MAC

[WDL^{+}20] B. Wan, J. Dang, Z. Li, H. Gong, F. Zhang, and S. Oh. Modeling analysis and cost-performance ratio optimization of virtual machine scheduling in cloud computing. IEEE Transactions on Parallel and Distributed Systems, 31(7):1518–1532, July 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wainer:2001:UAS

[WDSW01] Gabriel A. Wainer, Sergio Daicz, Luis F. De Simoni, and Demian Wassermann. Using the Alfa-1 simulated processor for educational purposes. ACM Journal on Educational Resources in Computing (JERIC), 1(4):111–151, December 2001. CODEN ???? ISSN 1531-4278.

Watanabe:2018:SEE

[WDT18] Ryo Watanabe, Dilawaer Duolikun, and Makoto Takizawa. Simple estimation and energy-aware migration models of virtual machines in a server cluster. Concurrency and Computation: Practice and Experience, 30(21):e4771:1–e4771:??, November 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Weber:2010:EVM

[Web10] Michael Weber. An embeddable virtual machine for state space generation. International Journal on Software Tools for Technology Transfer: STTT, 12(2):97–111, May 2010. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1433-2779&volume=12&issue=
2&spage=97.

Welch:1994:PVM

[Wel94] L. R. Welch. A parallel virtual machine for programs composed of abstract data types. IEEE Transactions on Computers, 43(11):1249–1261, November 1994. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=324558.

Wells:2002:HMA

[Wel02] Paul Damian Wells. The HSSM macro-architecture, Virtual Machine and H languages. ACM SIGPLAN Notices, 37(4):74–82, April 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Westley:1998:WJA

[Wes98] Terry J. Westley. Writing Java applets in Ada: a beginner’s guide. Ada User Journal, 19(1):57–82, April 1998. CODEN AUJOET. ISSN 0268-652X.

Ward:2003:VWH

[WF03] Brian Ward and Gerhard Franken. VMware Workstation: [das Handbuch; Installation, Konfiguration, Anwendung und Troubleshooting; Gast-Systeme: Windows, Linux, BSD, Novell NetWare, Solaris, FreeDOS und Oberon; virtuelle Netzwerke, Netzwerkkonfiguration und -dienste]. mitp-Verl., Landsberg, Germany, 2003. ISBN 3-8266-0964-6. 336 pp. LCCN ???? EUR 32.00 (DE). URL http://www.gbv.de/

du/services/agi/FCC0A57071BE8695C125704A0029797F/
FLMA122525.

Wires:2007:SFS

[WF07] Jake Wires and Michael J. Feeley. Secure file system versioning at the block level. Operating Systems Review, 41(3): 203–215, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wen:2025:TER

[WFL^{+}25] Xiaoqing Wen, Quanbi Feng, Hanzheng Lyu, Jianyu Niu, Yinqian Zhang, and Chen Feng. TeeRollup: Efficient rollup design using heterogeneous TEE. IEEE Transactions on Computers, 74(10):3546–3558, October 2025. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Williams:2007:VXI

[WG07] David E. Williams and Juan R. Garcia. Virtualization with Xen: including XenEnterprise, XenServer, and XenExpress. Syngress Publishing, Inc., Rockland, MA, USA, 2007. ISBN 1-59749-167-5. xx + 364 pp. LCCN QA76.9.V5 W55

2007. URL http://www.loc.gov/catdir/enhancements/
fy0745/2007278278-d.html.

Wagner:2011:SJV

[WGF11] Gregor Wagner, Andreas Gal, and Michael Franz. “slimming” a Java virtual machine by way of cold code removal and optimistic partial program loading. Science of Computer Programming, 76(11):1037–1053, November 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Weng:2013:HCM

[WGLL13] Chuliang Weng, Minyi Guo, Yuan Luo, and Minglu Li. Hybrid CPU management for adapting to the diversity of virtual machines. IEEE Transactions on Computers, 62(7):1332– 1344, July 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wan:2018:ADU

[WGW^{+}18] Xili Wan, Xinjie Guan, Tianjing Wang, Guangwei Bai, and Baek-Yong Choi. Application deployment using Microservice and Docker containers: Framework and optimization. Journal of Network and Computer Applications, 119(??):97–109, October 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print),

1095-8592 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1084804518302273.

Wei:2025:EEM

[WGXC25] Wenting Wei, Huaxi Gu, Zhe Xiao, and Yi Chen. Energy efficient and multi-resource optimization for virtual machine placement by improving MOEA/D. IEEE Transactions on Parallel and Distributed Systems, 36(6):1087–1099, June 2025. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2020:OVR

[WGY20] Xiaohui Wang, Haoran Gu, and YuXian Yue. The optimization of virtual resource allocation in cloud computing based on RBPSO. Concurrency and Computation: Practice and Experience, 32(16):e5113:1–e5113:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Waldron:1999:AVM

[WH99] John Waldron and Owen Harrison. Analysis of virtual machine stack frame usage by Java methods. In Anonymous [Ano99b], pages 271–274. ISBN ???? LCCN ????

Wolf:2005:VDE

[WH05] Chris Wolf and Erick M. Halter. Virtualization: from the desktop to the enterprise. The expert’s voice in networking. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-495-9. xxxiv + 559 pp. LCCN QA76.76.O63 W614

2005. URL http://www.loc.gov/catdir/enhancements/
fy0663/2006296449-d.html.

Weinhold:2008:VBV

[WH08] Carsten Weinhold and Hermann Härtig. VPFS: building a virtual private file system with a small trusted computing base. Operating Systems Review, 42(4):81–93, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2016:BPF

[WHC16] Yang Wang, Qian Hu, and Xiaojun Cao. A branch-and-price framework for optimal virtual network embedding. Computer Networks (Amsterdam, Netherlands: 1999), 94(??):318–326, January 15, 2016. CODEN ???? ISSN 1389-1286 (print),

1872-7069 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1389128615004089.

Williams:2009:STD

[WHD^{+}09] Daniel Williams, Wei Hu, Jack W. Davidson, Jason D. Hiser, John C. Knight, and Anh Nguyen-Tuong. Security through diversity: Leveraging virtual machine technology. IEEE Security & Privacy, 7(1):26–33, January/February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Williams:2016:EEH

[WHD^{+}16] Dan Williams, Yaohui Hu, Umesh Deshpande, Piush K. Sinha, Nilton Bila, Kartik Gopalan, and Hani Jamjoom. Enabling efficient hypervisor-as-a-service clouds with ephemeral virtualization. ACM SIGPLAN Notices, 51(7):79–92, July 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wessel:2015:IMD

[WHSE15] Sascha Wessel, Manuel Huber, Frederic Stumpf, and Claudia Eckert. Improving mobile device security with operating system-level virtualization. Computers & Security, 52(??):207–220, July 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404815000206.

Wei:2020:EAE

[WHW20] Chen Wei, Zhi-Hua Hu, and You-Gan Wang. Exact algorithms for energy-efficient virtual machine placement in data centers. Future Generation Computer Systems, 106(??):77– 91, May 2020. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167739X19319594.

Widdowson:2001:JIS

[Wid01] Liam Widdowson. Jailed Internet services. Sys Admin: The Journal for UNIX Systems Administrators, 10(8):39, 41–42, 44–45, August 2001. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Wagelaar:2012:TSC

[WIDP12] Dennis Wagelaar, Ludovico Iovino, Davide Di Ruscio, and Alfonso Pierantonio. Translational semantics of a co-evolution specific language with the EMF transformation virtual machine. Lecture Notes in Computer Science, 7307:192–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-30476-7_13/.

Wilson:2001:UVD

[Wil01] Brian Wilson. Using VMWare as a development tool. Sys Admin: The Journal for UNIX Systems Administrators, 10 (3):57–59, March 2001. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/.

Wills:2006:PVC

[Wil06] Dean C. Wills. Pro Visual C++ 2005 for developers: featuring C++/CLI. Books for professionals by professionals. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-608-0. xxxii + 379 pp. LCCN QA76.73.C153 W5526 2006. URL http:/

/www.loc.gov/catdir/toc/fy0803/2008295944.html.

Winett:1971:VMD

[Win71] J. M. Winett. Virtual machines for developing systems software. In ????, editor, Proceedings IEEE Computer Society Conference, Boston, MA, September 1971, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1971.

Wang:2015:DAA

[WIS^{+}15] Hui Wang, Canturk Isci, Lavanya Subramanian, Jongmoo Choi, Depei Qian, and Onur Mutlu. A-DRM: Architectureaware distributed resource management of virtualized clusters. ACM SIGPLAN Notices, 50(7):93–106, July 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2010:HLA

[WJ10] Zhi Wang and Xuxian Jiang. HyperSafe: a lightweight approach to provide lifetime hypervisor control-flow integrity. Report, Department of Computer Science, North Carolina State University, Raleigh, NC, USA, 2010. 16

pp.  URL http://www.csc.ncsu.edu/faculty/jiang/
pubs/OAKLAND10.pdf.

Wentzlaff:2012:CFG

[WJGA12] David Wentzlaff, Christopher J. Jackson, Patrick Griffin, and Anant Agarwal. Configurable fine-grain protection for multicore processor virtualization. ACM SIGARCH Computer Architecture News, 40(3):464–475, June 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ISCA ’12 conference proceedings.

Whang:1990:QOM

[WK90] Kyu-Young Y. Whang and Ravi Krishnamurthy. Query optimization in a memory-resident domain relational calculus database system. ACM Transactions on Database Systems, 15(1):67–95, March 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/tods/1990-15-1/p67-
whang/p67-whang.pdf;   http://www.acm.org/pubs/citations/
journals/tods/1990-15-1/p67-whang/;   http://www.
acm.org/pubs/toc/Abstracts/tods/77646.html.

Wegiel:2008:MCV

[WK08] Michal Wegiel and Chandra Krintz. The Mapping Collector: virtual memory support for generational, parallel, and concurrent compaction. Operating Systems Review, 42(2):91– 102, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wein:2009:VGT

[WKC^{+}09] Joel Wein, Kirill Kourtchikov, Yan Cheng, Ron Gutierez, Roman Khmelichek, Matthew Topol, and Chris Sherman. Virtualized games for teaching about distributed systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):246–250, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Wang:2017:JRJ

[WKG17] Kaiyuan Wang, Sarfraz Khurshid, and Milos Gligoric. JPR: Replaying JPF traces using standard JVM. ACM SIGSOFT Software Engineering Notes, 42(4):1–5, October 2017. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wang:2015:IJV

[WKJ15] Yang Wang, Kenneth B. Kent, and Graeme Johnson. Improving J9 virtual machine with LTTng for efficient and effective tracing. Software—Practice and Experience, 45(7):973–987, July 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wade:2017:AVJ

[WKJ17] April W. Wade, Prasad A. Kulkarni, and Michael R. Jantz. AOT vs. JIT: impact of profile data on code quality. ACM SIGPLAN Notices, 52(4):1–10, May 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wade:2020:EIP

[WKJ20] April W. Wade, Prasad A. Kulkarni, and Michael R. Jantz. Exploring impact of profile data on code quality in the HotSpot JVM. ACM Transactions on Embedded Computing Systems, 19(6):48:1–48:26, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:

//dl.acm.org/doi/10.1145/3391894.

Wang:2008:PEV

[WKT08] Lizhe Wang, Marcel Kunze, and Jie Tao. Performance evaluation of virtual machine-based Grid workflow system. Concurrency and Computation: Practice and Experience, 20(15): 1759–1771, October 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Waddington:1996:JVM

[WL96] Simon Waddington and Stephen Li. Java: Virtual machine for virtually any platform. Embedded Systems Programming, 9(6):26–28, 30, 32, 34, 36, 38, 40, 42, June 1, 1996. CODEN EYPRE4. ISSN 1040-3272.

Wu:2017:ACM

[WLCS17] Jiang Wu, Zhou Lei, Shengbo Chen, and Wenfeng Shen. An access control model for preventing virtual machine escape attack. Future Internet, 9(2):20, June 02, 2017. CODEN

???? ISSN 1999-5903. URL https://www.mdpi.com/1999-
5903/9/2/20.

Wei:2011:LPV

[WLG^{+}11] Xiaohui Wei, Hongliang Li, Qingnan Guo, Na Jiang, and Liang Hu. LimeVI: A platform for virtual cluster live migration over WAN. International Journal of Computer Systems Science and Engineering, 26(5):??, September 2011. CODEN CSSEEI. ISSN 0267-6192.

Wen:2013:MPA

[WLL^{+}13] Yuanfeng Wen, JongHyuk Lee, Ziyi Liu, Qingji Zheng, Weidong Shi, Shouhuai Xu, and Taeweon Suh. Multi-processor architectural support for protecting virtual machine privacy in untrusted cloud environment. In Hubertus Franke, editor, Proceedings of the ACM International Conference on Computing Frontiers, CF’ 13, Ischia, Italy. 14–16 May 2013, pages 25:1–25:10. ACM Press, New York, NY 10036, USA, 2013. ISBN 1-4503-2053-8. LCCN QA75.5. URL https://

dl.acm.org/citation.cfm?id=2482799.

Weng:2016:CMV

[WLLZ16] Chuliang Weng, Qian Liu, Kenli Li, and Deqing Zou. Cloud-Mon: Monitoring virtual machines in clouds. IEEE Transactions on Computers, 65(12):3787–3793, ???? 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

West:2016:VSK

[WLMD16] Richard West, Ye Li, Eric Missimer, and Matthew Danish. A virtualized separation kernel for mixed-criticality systems. ACM Transactions on Computer Systems, 34(3):8:1– 8:??, September 2016. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Wang:2018:TCB

[WLS^{+}18] Haiyang Wang, Tong Li, Ryan Shea, Xiaoqiang Ma, Feng Wang, Jiangchuan Liu, and Ke Xu. Toward cloud-based distributed interactive applications: Measurement, modeling, and analysis. IEEE/ACM Transactions on Networking, 26 (1):3–16, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2015:HPI

[WLW^{+}15] Zhe Wang, Jianjun Li, Chenggang Wu, Dongyan Yang, Zhenjiang Wang, Wei-Chung Hsu, Bin Li, and Yong Guan. HSPT: Practical implementation and efficient management of embedded shadow page tables for cross-ISA system virtual machines. ACM SIGPLAN Notices, 50(7):53–64, July 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wu:2017:VPP

[WLW^{+}17] Song Wu, Yongchang Li, Xinhou Wang, Hai Jin, and Hanhua Chen. Vshadow: Promoting physical servers into virtualization world. International Journal of Parallel Programming, 45(1):45–66, February 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://link.

springer.com/article/10.1007/s10766-015-0385-2.

Whaley:2002:AEO

[WML02] John Whaley, Michael C. Martin, and Monica S. Lam. Automatic extraction of object-oriented component interfaces. ACM SIGSOFT Software Engineering Notes, 27(4):218–228, July 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wenzl:2019:HET

[WMUW19] Matthias Wenzl, Georg Merzdovnik, Johanna Ullrich, and Edgar Weippl. From hack to elaborate technique — a survey on binary rewriting. ACM Computing Surveys, 52(3): 49:1–49:37, July 2019. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3316415.

Wei:2017:DCS

[WN17] Xiaohan Wei and Michael J. Neely. Data center server provision: Distributed asynchronous control for coupled renewal systems. IEEE/ACM Transactions on Networking, 25(4): 2180–2194, August 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wulf:1983:SFR

[WNL^{+}83] William A. Wulf, Joe Newcomer, Bruce Leverett, Rick Cattell, and Paul Knueven. Surveyor’s forum: Retargetable code generators. ACM Computing Surveys, 15(3):279–280, September 1983. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). See [GFH82, GHF83a, Fra83, GHF83b].

Williams:1975:CMI

[WO75] M. H. Williams and H. L. Ossher. A computer model for instructional purposes. The Computer Journal, 18(4):333–341, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_18/Issue_04/tiff/333.
tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/334.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_18/Issue_04/tiff/
335.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/336.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/337.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/338.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/339.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/340.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/341.tif.

Wolczko:1999:UTJ

[Wol99] M. Wolczko. Using a Tracing Java Virtual Machine to gather data on the behavior of Java programs. Technical report, Sun Microsystems, Inc., Menlo Park, CA, USA, 1999.

???? pp. URL http://research.sun.com/people/mario/
tracing-jvm/.

Wong:1997:MHJ

[Won97] Wylie Wong. Microsoft hones Java strategy: Virtual machine, class libraries are on top. ComputerWorld, 31(14):113, April 7, 1997. CODEN CMPWAB. ISSN 0010-4841.

Winterbottom:1997:DIV

[WP97] Phil Winterbottom and Rob Pike. The design of the Inferno virtual machine. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Wang:2015:HRR

[WQG15] Jihe Wang, Meikang Qiu, and Bing Guo. High reliable realtime bandwidth scheduling for virtual machines with hidden Markov predicting in telehealth platform. Future Generation Computer Systems, 49(??):68–76, August 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X1400154X.

Warnke:2007:QVC

[WR07] Robert Warnke and Thomas Ritzau. QEMU virtuelle Computer für viele Betriebssysteme; QEMU Version 0.9.0. (German) [QEMU virtual computer for many operating systems]. Books on Demand GmbH, Norderstedt, 2007. ISBN 3-8370-0876-2. 292 pp. LCCN ???? URL http://d-nb.info/

986260371/04.

Warnke:2008:QVC

[WR08] Robert Warnke and Thomas Ritzau. QEMU: virtuelle Computer für viele Betriebssysteme; QEMU Version 0.9.1. Books on Demand GmbH, Norderstedt, Germany, 2008. ISBN 3-8370-0876-2. 300 pp. LCCN ???? URL http://d-nb.info/

986260371/04.

Waldspurger:2012:V

[WR12] Carl Waldspurger and Mendel Rosenblum. I/O virtualization. Communications of the ACM, 55(1):66–73, January 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wang:2013:VPD

[WRS13] Di Wang, Chuangang Ren, and Anand Sivasubramaniam. Virtualizing power distribution in datacenters. ACM SIGARCH Computer Architecture News, 41(3):595–606, June 2013. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ICSA ’13 conference proceedings.

Wood:2015:CDP

[WRS^{+}15] Timothy Wood, K. K. Ramakrishnan, Prashant Shenoy, Jacobus Van Der Merwe, Jinho Hwang, Guyue Liu, and Lucas Chaufournier. CloudNet: dynamic pooling of cloud resources by live WAN migration of virtual machines. IEEE/ ACM Transactions on Networking, 23(5):1568–1583, October 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wood:2011:CDP

[WRSvdM11] Timothy Wood, K. K. Ramakrishnan, Prashant Shenoy, and Jacobus van der Merwe. CloudNet: dynamic pooling of cloud resources by live WAN migration of virtual machines. ACM SIGPLAN Notices, 46(7):121–132, July 2011. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2011:RVM

[WRX11] Kun Wang, Jia Rao, and Cheng-Zhong Xu. Rethink the virtual machine template. ACM SIGPLAN Notices, 46(7):39– 50, July 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

White:2013:CTP

[WSAJ13] David R. White, Jeremy Singer, Jonathan M. Aitken, and Richard E. Jones. Control theory for principled heap sizing. ACM SIGPLAN Notices, 48(11):27–38, November 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ISMM ’13 conference proceedings.

Wood:2009:SBB

[WSVY09] Timothy Wood, Prashant Shenoy, Arun Venkataramani, and Mazin Yousif. Sandpiper: Black-box and gray-box resource management for virtual machines. Computer Networks (Amsterdam, Netherlands: 1999), 53(17):2923–2938, December 3, 2009. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2019:ATA

[WSX^{+}19] Zhong Wang, Daniel Sun, Guangtao Xue, Shiyou Qian, Guoqiang Li, and Minglu Li. Ada-Things: an adaptive virtual machine monitoring and migration strategy for Internet of Things applications. Journal of Parallel and Distributed Computing, 132(??):164–176, October 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304404.

Wejchert:1991:VPN

[WT91] J. Wejchert and G. Tesauro. Visualizing processes in neural networks. IBM Journal of Research and Development, 35(1/2):244–253, January/March 1991. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Woldeyohannes:2022:CEA

[WTJR22] Yordanos Tibebu Woldeyohannes, Besmir Tola, Yuming Jiang, and K. K. Ramakrishnan. CoShare: an efficient approach for redundancy allocation in NFV. IEEE/ACM Transactions on Networking, 30(3):1014–1028, June 2022. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.acm.org/doi/10.1109/TNET.
2021.3132279.

Wu:2016:IBP

[WTL^{+}16] Song Wu, Songqiao Tao, Xiao Ling, Hao Fan, Hai Jin, and Shadi Ibrahim. iShare: Balancing I/O performance isolation and disk I/O efficiency in virtualized environments. Concurrency and Computation: Practice and Experience, 28(2): 386–399, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wood:2009:MBE

[WTLS^{+}09] Timothy Wood, Gabriel Tarasuk-Levin, Prashant Shenoy, Peter Desnoyers, Emmanuel Cecchet, and Mark D. Corner. Memory buddies: exploiting page sharing for smart colocation in virtualized data centers. Operating Systems Review, 43(3):27–36, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Win:2018:BDB

[WTM18] T. Y. Win, H. Tianfield, and Q. Mair. Big data based security analytics for protecting virtualized infrastructures in cloud computing. IEEE Transactions on Big Data, 4(1):11– 25, March 2018. ISSN 2332-7790.

Wu:2013:HSC

[Wu13] Youfeng Wu. HW/SW co-designed acceleration of dynamic languages. ACM SIGPLAN Notices, 48(5):1–2, May 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2018:ECM

[WUK^{+}18] Cheng Wang, Bhuvan Urgaonkar, George Kesidis, Aayush Gupta, Lydia Y. Chen, and Robert Birke. Effective capacity modulation as an explicit control knob for public cloud profitability. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(1):2:1–2:??, May 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wunschiers:2013:CBP

[Wün13] Röbbe Wünschiers. Computational biology: a practical introduction to biodata processing and analysis with Linux, MySQL, and R. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2013. ISBN 3-642-34748-7, 3-642-34749-5 (e-book). xxix + 449 pp. LCCN QH324.2 .W86 2013.

Wang:2017:UBI

[WUNK17] Cheng Wang, Bhuvan Urgaonkar, Neda Nasiriani, and George Kesidis. Using burstable instances in the public cloud: Why, when and how? Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1): 11:1–11:??, June 2017. CODEN ???? ISSN 2476-1249. URL

http://dl.acm.org/citation.cfm?id=3084448.

Wang:2017:SMC

[WVT^{+}17] Yang Wang, Bharadwaj Veeravalli, Chen-Khong Tham, Shuibing He, and Chengzhong Xu. On service migrations in the cloud for mobile accesses: a distributed approach. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(2):6:1–6:??, May 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wilding-White:1977:MBI

[WW77] Arthur Llewellyn Wilding-White. A microprocessor BCPL implementation based on a virtual stack machine. Thesis (B.S.), Department of Electrical Engineering and Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, 1977. 143 pp. Supervised by Stephen A. Ward.

Wang:2016:DMB

[WWH^{+}16] Zhigang Wang, Xiaolin Wang, Fang Hou, Yingwei Luo, and Zhenlin Wang. Dynamic memory balancing for virtualization.

ACM Transactions on Architecture and Code Optimization, 13(1):2:1–2:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wurthinger:2017:PPE

[WWH^{+}17] Thomas Würthinger, Christian Wimmer, Christian Humer, Andreas Wöß, Lukas Stadler, Chris Seaton, Gilles Duboscq, Doug Simon, and Matthias Grimmer. Practical partial evaluation for high-performance dynamic language runtimes. ACM SIGPLAN Notices, 52(6):662–676, June 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2017:RLW

[WWL^{+}17a] Zhe Wang, Chenggang Wu, Jianjun Li, Yuanming Lai, Xiangyu Zhang, Wei-Chung Hsu, and Yueqiang Cheng. ReRanz: a light-weight virtual machine to mitigate memory disclosure attacks. ACM SIGPLAN Notices, 52(7):143–156, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wu:2017:AOA

[WWL^{+}17b] Song Wu, Yihong Wang, Wei Luo, Sheng Di, Haibao Chen, Xiaolin Xu, Ran Zheng, and Hai Jin. ACStor: Optimizing access performance of virtual disk images in clouds. IEEE Transactions on Parallel and Distributed Systems, 28 (9):2414–2427, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:

//www.computer.org/csdl/trans/td/2017/09/07866840-
abs.html.

Wang:2025:TDV

[WWL^{+}25] Dan Wang, Jinjiang Wang, Xize Liu, Junyang Yu, Hangyu Gu, Congyang Wang, Jinghan Liu, and Yanhao Zhang. Towards dynamic virtual machine placement based on safety parameters and resource utilization fluctuation for energy savings and QoS improvement in cloud computing. Future Generation Computer Systems, 171(??):??, October 2025. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X25001487.

Wright:2006:IJV

[WWMG06] Greg Wright, Mario Wolczko, Phil McGachey, and Erika Gunadi. Introspection of a Java Virtual Machine under simulation. Technical Report TR-2006-159, Sun Microsystems, Inc., Menlo Park, CA, USA, September 2006.

URL http://research.sun.com/techrep/2006/smli_tr-
2006-159.pdf.

Wang:1989:NNS

[WWS89] C.-J. Wang, C.-H. Wu, and S. Sivasundaram. Neural network simulation on shared-memory vector multiprocessors. In ACM [ACM89], pages 197–204. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Wendorf:1989:SOS

[WWT89] J. W. Wendorf, R. G. Wendorf, and H. Tokuda. Scheduling operating system processing on small-scale multiprocessors. In Shriver [Shr89], pages 904–913 (vol. 2). ISBN 0-8186-1912-0. LCCN ???? IEEE catalog number 89TH0243-6.

Wang:2013:RMM

[WWWL13] Xiaolin Wang, Lingmei Weng, Zhenlin Wang, and Yingwei Luo. Revisiting memory management on virtualized environments. ACM Transactions on Architecture and Code Optimization, 10(4):48:1–48:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wang:2025:BEC

[WXG^{+}25] Yicheng Wang, Lijie Xu, Tian Guo, Wensheng Dou, Hongbin Zeng, Wei Wang, Jun Wei, and Tao Huang. BridgeGC: an efficient cross-level garbage collector for big data frameworks. ACM Transactions on Architecture and Code Optimization, 22(2):68:1–68:??, June 2025. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wu:2015:WHS

[WXW15] Zhenyu Wu, Zhang Xu, and Haining Wang. Whispers in the hyper-space: high-bandwidth and reliable covert channel attacks inside the cloud. IEEE/ACM Transactions on Networking, 23(2):603–614, April 2015. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wang:2017:RES

[WXZ^{+}17] Tao Wang, Jiwei Xu, Wenbo Zhang, Jianhua Zhang, Jun Wei, and Hua Zhong. ReSeer: Efficient search-based replay for multiprocessor virtual machines. The Journal of Systems and Software, 126(??):101–112, April 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121216301248.

Wu:2020:ESI

[WYZAD20] Yulei Wu, Zheng Yan, Zhiwei Zhao, and Ahmed Al-Dubai. Editorial: Special issue on SDN-based wireless network virtualization. Concurrency and Computation: Practice and Experience, 32(16):e5444:1–e5444:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2024:OPT

[WYZY24] Liang Wang, Jinzhe Yang, Jidong Zhai, and Guangwen Yang. Optimizing I/O performance through effective vCPU scheduling interference management. IEEE Transactions on Parallel and Distributed Systems, 35(12):2315–2330, December 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2016:PDI

[WZH^{+}16] Shangguang Wang, Ao Zhou, Ching-Hsien Hsu, Xuanyu Xiao, and Fangchun Yang. Provision of data-intensive services through energy- and QoS-aware virtual machine placement in national cloud data centers. IEEE Transactions on Emerging Topics in Computing, 4(2):290–300, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2019:MTR

[WZKP19] Qingyang Wang, Shungeng Zhang, Yasuhiko Kanemasa, and Calton Pu. Mitigating tail response time of n-tier applications: The impact of asynchronous invocations. ACM Transactions on Internet Technology (TOIT), 19(3):36:1–36:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-

6051 (electronic). URL https://dl.acm.org/ft_gateway.
cfm?id=3340462.

Weng:2015:TEI

[WZL15] Chuliang Weng, Jianfeng Zhan, and Yuan Luo. TSAC: Enforcing isolation of virtual machines in clouds. IEEE Transactions on Computers, 64(5):1470–1482, ???? 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2018:EBJ

[WZL^{+}18] Mingyu Wu, Ziming Zhao, Haoyu Li, Heting Li, Haibo Chen, Binyu Zang, and Haibing Guan. Espresso: Brewing Java for more non-volatility with non-volatile memory. ACM SIG-PLAN Notices, 53(2):70–83, February 2018. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wu:2023:PTL

[WZL^{+}23] Qiang Wu, Xiangping Bryce Zhai, Xi Liu, Chun-Ming Wu, Fangliang Lou, and Hongke Zhang. Performance tuning via lean measurements for acceleration of network functions virtualization. IEEE/ACM Transactions on Networking, 31(1): 366–379, 2023. CODEN IEANEP. ISSN 1063-6692 (print),

1558-2566 (electronic). URL https://dl.acm.org/doi/10.
1109/TNET.2022.3193686.

Wei:2019:HBS

[WZT19] Song Wei, Kun Zhang, and Bibo Tu. HyperBench: a benchmark suite for virtualization capabilities. ACM SIGMET-RICS Performance Evaluation Review, 47(1):73–74, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3376930.3376977.

Wang:2013:JVM

[WZV^{+}13] Lin Wang, Fa Zhang, Athanasios V. Vasilakos, Chenying Hou, and Zhiyong Liu. Joint virtual machine assignment and traffic engineering for green data center networks. ACM SIGMETRICS Performance Evaluation Review, 41(3):107– 112, December 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wang:2011:SHS

[WZW^{+}11] Xiaolin Wang, Jiarui Zang, Zhenlin Wang, Yingwei Luo, and Xiaoming Li. Selective hardware/software memory virtualization. ACM SIGPLAN Notices, 46(7):217–226, July 2011. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2020:MVN

[WZZ^{+}20] Cong Wang, Fanghui Zheng, Guangcong Zheng, Sancheng Peng, Zejie Tian, Yujia Guo, Guorui Li, and Ying Yuan. Modeling on virtual network embedding using reinforcement learning. Concurrency and Computation: Practice and Experience, 32(23):e6020:1–e6020:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2022:NMB

[XA22] Yanfei Xu and Karlo Abnoosian. A new metaheuristic-based method for solving the virtual machines migration problem in the green cloud computing. Concurrency and Computation: Practice and Experience, 34(3):e6579:1–e6579:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2014:DIP

[XCJ^{+}14] Xia Xie, Wenzhi Cao, Hai Jin, Xijiang Ke, and Shuwen Luo. Design and implementation of process-aware predictive scheduling scheme for virtual machine. The Journal of Supercomputing, 70(3):1577–1587, December 2014. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11227-014-1254-2.

Xie:2018:NSA

[XCSM18] Lei Xie, Shengbo Chen, Wenfeng Shen, and Huaikou Miao. A novel self-adaptive VM consolidation strategy using dynamic multi-thresholds in IaaS clouds. Future Internet, 10(6):52, June 13, 2018. CODEN ???? ISSN 1999-5903. URL https:

//www.mdpi.com/1999-5903/10/6/52.

Xu:2016:SHS

[XD16] Xin Xu and Bhavesh Davda. SRVM: Hypervisor support for live migration with passthrough SR-IOV network devices. ACM SIGPLAN Notices, 51(7):65–77, July 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2017:HAE

[XD17] Xin Xu and Bhavesh Davda. A hypervisor approach to enable live migration with passthrough SR-IOV network devices. Operating Systems Review, 51(1):15–23, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xie:2015:PDC

[XDLS15] Qiaomin Xie, Xiaobo Dong, Yi Lu, and Rayadurgam Srikant. Power of d choices for large-scale bin packing: a loss model. ACM SIGMETRICS Performance Evaluation Review, 43(1): 321–334, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Xu:1990:HMD

[XH90] J. Xu and K. Hwang. Heuristic methods for dynamic load balancing in a message-passing supercomputer. In IEEE [IEE90b], pages 888–897. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Xu:2016:SER

[XH16] Xin Xu and H. Howie Huang. On soft error reliability of virtualization infrastructure. IEEE Transactions on Computers, 65(12):3727–3739, ???? 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xing:2015:OIB

[XHCL15] Bin Xing, Zhen Han, Xiaolin Chang, and Jiqiang Liu. OB–IMA: out-of-the-box integrity measurement approach for guest virtual machines. Concurrency and Computation: Practice and Experience, 27(5):1092–1109, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiao:2013:VMP

[XHL^{+}13] Peng Xiao, Zhigang Hu, Dongbo Liu, Guofeng Yan, and Xilong Qu. Virtual machine power measuring technique with bounded error in cloud environments. Journal of Network and Computer Applications, 36(2):818–828, March 2013. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512002494.

Xie:2019:DDS

[XHW^{+}19] An Xie, Huawei Huang, Xiaoliang Wang, Song Guo, Zhuzhong Qian, and Sanglu Lu. Dual: Deploy stateful virtual network function chains by jointly allocating datacontrol traffic. Computer Networks (Amsterdam, Netherlands: 1999), 162(??):Article 106868, ???? 2019. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1389128619303962.

Xu:2017:BBD

[XJR^{+}17] Quanqing Xu, Chao Jin, Mohamed Faruq Bin Mohamed Rasid, Bharadwaj Veeravalli, and Khin Mi Mi Aung. Blockchain-based decentralized content trust for Docker images. Multimedia Tools and Applications, ??(??), October 2017. CODEN MTAPFB. ISSN 1380-7501 (print), 1573-7721

(electronic). URL http://link.springer.com/article/
10.1007/s11042-017-5224-6.

Xu:2018:MVM

[XJW^{+}18] Chuan Xu, Wenqiang Jin, Xinheng Wang, Guofeng Zhao, and Shui Yu. MC-VAP: a multi-connection virtual access point for high performance software-defined wireless networks. Journal of Network and Computer Applications, 122(??):88–98, November 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804518302662.

Xu:2015:RSV

[XJWW15] Xiaolin Xu, Hai Jin, Song Wu, and Yihong Wang. Rethink the storage of virtual machine images in clouds. Future Generation Computer Systems, 50(??):75–86, September 2015. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X14001885.

Xevgenis:2023:AZS

[XKKL23] Michael Xevgenis, Dimitrios G. Kogias, Panagiotis A. Karkazis, and Helen C. Leligou. Addressing ZSM security issues with blockchain technology. Future Internet, 15(4): 129, March 28, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/1999-5903/15/4/129.

Xia:2011:VWB

[XKY^{+}11] Lei Xia, Sanjay Kumar, Xue Yang, Praveen Gopalakrishnan, York Liu, Sebastian Schoenberg, and Xingang Guo. Virtual WiFi: bring virtualization from wired to wireless. ACM SIG-PLAN Notices, 46(7):181–192, July 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2016:HIA

[XLJ16] Fei Xu, Fangming Liu, and Hai Jin. Heterogeneity and interference-aware virtual machine provisioning for predictable performance in the cloud. IEEE Transactions on Computers, 65(8):2470–2483, ???? 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2014:IML

[XLL^{+}14] Fei Xu, Fangming Liu, Linghui Liu, Hai Jin, Bo Li, and Baochun Li. iAware: Making live migration of virtual machines interference-aware in the cloud. IEEE Transactions on Computers, 63(12):3012–3025, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2020:AAR

[XLL^{+}20] Yajing Xu, Junnan Li, Zhihui Lu, Jie Wu, Patrick C. K. Hung, and Abdulhameed Alelaiwi. ARVMEC: Adaptive recommendation of virtual machines for IoT in edgecloud environment. Journal of Parallel and Distributed Computing, 141(??):23–34, July 2020. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731520302045.

Xu:2018:EVC

[XLQL18] Xinping Xu, Wenxin Li, Heng Qi, and Keqiu Li. On efficient virtual cluster scaling across geo-distributed data centers. Concurrency and Computation: Practice and Experience, 30(10):??, May 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cpe.4383.

Xu:2019:MCE

[XLWX19] Heyang Xu, Yang Liu, Wei Wei, and Ying Xue. Migration cost and energy-aware virtual machine consolidation under cloud environments considering remaining runtime. International Journal of Parallel Programming, 47(3):481–501, June 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Xu:2018:IAV

[XLWZ18] Heyang Xu, Yang Liu, Wei Wei, and Wenqiang Zhang. Incentive-aware virtual machine scheduling in cloud computing. The Journal of Supercomputing, 74(7):3016–3038, July 2018. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11227-018-2349-y.pdf.

Xue:2018:SGV

[XML^{+}18] Mochi Xue, Jiacheng Ma, Wentai Li, Kun Tian, Yaozu Dong, Jinyu Wu, Zhengwei Qi, Bingsheng He, and Haibing Guan. Scalable GPU virtualization with dynamic sharing of graphics memory space. IEEE Transactions on Parallel and Distributed Systems, 29(8):1823–1836, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2018/08/08247267-abs.html.

Xiao:2021:IEE

[XNH21] Peng Xiao, Zhenyu Ni, and Zhigang Hu. Improving the energy-efficiency of virtual machines by I/O compensation. The Journal of Supercomputing, 77(10):11135–11159, October 2021. CODEN JOSUED. ISSN 0920-8542 (print),

1573-0484 (electronic). URL https://link.springer.com/
article/10.1007/s11227-021-03678-8.

Xu:2022:NOL

[XRL^{+}22] Zichuan Xu, Haozhe Ren, Weifa Liang, Qiufen Xia, Wanlei Zhou, Pan Zhou, Wenzheng Xu, Guowei Wu, and Mingchu Li. Near optimal learning-driven mechanisms for stable NFV markets in multitier cloud networks. IEEE/ACM Transactions on Networking, 30(6):2601–2615, 2022. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.acm.org/doi/10.1109/TNET.
2022.3179295.

Xiao:2013:DRA

[XSC13] Zhen Xiao, Weijia Song, and Qi Chen. Dynamic resource allocation using virtual machines for cloud computing environment. IEEE Transactions on Parallel and Distributed Systems, 24(6):1107–1117, June 2013. CODEN ITDSEO. ISSN 1045-9219.

Xu:2017:SLB

[XTB17] Minxian Xu, Wenhong Tian, and Rajkumar Buyya. A survey on load balancing algorithms for virtual machines placement in cloud computing. Concurrency and Computation: Practice and Experience, 29(12), June 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2016:GCF

[XWH^{+}16] Yi Xie, Yu Wang, Haitao He, Yang Xiang, Shunzheng Yu, and Xincheng Liu. A general collaborative framework for modeling and perceiving distributed network behavior. IEEE/ACM Transactions on Networking, 24(5):3162–3176, October 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Xie:2015:SSV

[XWJX15] Ruitao Xie, Yonggang Wen, Xiaohua Jia, and Haiyong Xie. Supporting seamless virtual machine migration via named data networking in cloud data center. IEEE Transactions on Parallel and Distributed Systems, 26(12):3485–3497, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print),

1558-2183 (electronic). URL http://csdl.computer.org/
csdl/trans/td/2015/12/06975186-abs.html.

Xie:2021:OAM

[XWW^{+}21] Yanghao Xie, Sheng Wang, Binbin Wang, Shizhong Xu, Xiong Wang, and Jing Ren. Online algorithm for migration aware virtualized network function placing and routing in dynamic 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128621001924.

Xu:2017:EIR

[XWX^{+}17] Jiwei Xu, Tao Wang, Xiaozhao Xing, Wenbo Zhang, and Hua Zhong. Efficient image restoration of virtual machines with reference count based rewriting and caching. Future Generation Computer Systems, 77(??):87–96, December 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X17301218.

Xu:2023:FTQ

[XXWG23] Heyang Xu, Sen Xu, Wei Wei, and Naixuan Guo. Fault tolerance and quality of service aware virtual machine scheduling algorithm in cloud data centers. The Journal of Supercomputing, 79(3):2603–2625, February 2023. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11227-022-04760-5.

Xie:2013:AAE

[XXZ13] Xinwei Xie, Jingling Xue, and Jie Zhang. Acculock: accurate and efficient detection of data races. Software—Practice and Experience, 43(5):543–576, May 2013. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xu:2018:DES

[XYD^{+}18] Yu Xu, Jianguo Yao, Yaozu Dong, Kun Tian, Xiao Zheng, and Haibing Guan. Demon: an efficient solution for on-device MMU virtualization in mediated pass-through. ACM SIG-PLAN Notices, 53(3):57–70, March 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2017:OCV

[XYYY17] Cong Xu, Jiahai Yang, Kevin Yin, and Hui Yu. Optimal construction of virtual networks for Cloud-based MapReduce workflows. Computer Networks (Amsterdam, Netherlands: 1999), 112(??):194–207, January 15, 2017. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S138912861630367X.

Xiao:2011:HLM

[XZ11] Jing Xiao and Min Zhu. HyperMonitor: A lightweight multiplatform monitor based on hardware virtualization. International Journal of Computer Systems Science and Engineering, 26(5):??, September 2011. CODEN CSSEEI. ISSN 0267-6192.

Xu:2020:BVM

[XZK^{+}20] Xiaolong Xu, Xuyun Zhang, Maqbool Khan, Wanchun Dou, Shengjun Xue, and Shui Yu. A balanced virtual machine scheduling method for energy-performance trade-offs in cyber-physical cloud systems. Future Generation Computer Systems, 105(??):789–799, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X17318927.

Xu:2020:QAV

[XZL^{+}20] Zichuan Xu, Zhiheng Zhang, Weifa Liang, Qiufen Xia, Omer Rana, and Guowei Wu. QoS-aware VNF placement and service chaining for IoT applications in multi-tier mobile edge networks. ACM Transactions on Sensor Networks, 16(3): 23:1–23:27, August 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3387705.

Xu:2016:CBA

[XZZ^{+}16] Jiwei Xu, Wenbo Zhang, Zhenyu Zhang, Tao Wang, and Tao Huang. Clustering-based acceleration for virtual machine image deduplication in the cloud environment. The Journal of Systems and Software, 121(??):144–156, November 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-

1228 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0164121216000534.

Yousefi:2024:HEA

[YB24] Malek Yousefi and Seyed Morteza Babamir. A hybrid energyaware algorithm for virtual machine placement in cloud computing. Computing: Archiv für Informatik und Numerik, 106 (5):1297–1320, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https://link.

springer.com/article/10.1007/s00607-024-01280-3.

Yao:2015:MEV

[YBZ^{+}15] Hong Yao, Changmin Bai, Deze Zeng, Qingzhong Liang, and Yuanyuan Fan. Migrate or not? Exploring virtual machine migration in roadside cloudlet-based vehicular cloud. Concurrency and Computation: Practice and Experience, 27(18): 5780–5792, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yalamanchilli:1998:CPJa

[YC98a] Narendar Yalamanchilli and William Cohen. Communication performance of Java based Parallel Virtual Machines. In ACM [ACM98], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ???? URL http://www.

cs.ucsb.edu/conferences/java98/papers/passing.pdf;
 http://www.cs.ucsb.edu/conferences/java98/papers/

passing.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Yalamanchilli:1998:CPJb

[YC98b] Narendar Yalamanchilli and William Cohen. Communication performance of Java-based parallel virtual machines. Concurrency: Practice and Experience, 10(11–13):1189– 1196, September 1998. CODEN CPEXEI. ISSN 1040-

3108. URL http://www3.interscience.wiley.com/cgi-
bin/abstract?ID=10050407;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=10050407&PLACEBO=IE.

pdf. Special Issue: Java for High-performance Network Computing.

You:2016:SRB

[YC16] Yi-Ping You and Jian-Ru Chen. A static region-based compiler for the Dalvik virtual machine. Software—Practice and Experience, 46(8):1109–1130, August 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yang:2018:CVG

[YCL^{+}18] Chao-Tung Yang, Shuo-Tsung Chen, Yu-Sheng Lo, Endah Kristiani, and Yu-Wei Chan. On construction of a virtual GPU cluster with InfiniBand and 10 Gb Ethernet virtualization. The Journal of Supercomputing, 74(12):6876–6897, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2019:IRT

[YCL^{+}19] Chao-Tung Yang, Shuo-Tsung Chen, Jung-Chun Liu, Yao-Yu Yang, Karan Mitra, and Rajiv Ranjan. Implementation of a real-time network traffic monitoring service with network functions virtualization. Future Generation Computer Systems, 93(??):687–701, April 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167739X1830311X.

Yuan:2018:ASP

[YDW18] Xingliang Yuan, Huayi Duan, and Cong Wang. Assuring string pattern matching in outsourced middleboxes. IEEE/ ACM Transactions on Networking, 26(3):1362–1375, June 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yelland:1999:CAJ

[Yel99] Phillip M. Yelland. A compositional account of the Java virtual machine. In ACM [ACM99], pages 57–69. ISBN 1-58113-

095-3. LCCN ???? URL http://www.acm.org:80/pubs/
citations/proceedings/plan/292540/p57-yelland/.

Yang:2021:FGR

[YGLY21] Yang Yang, Songtao Guo, Guiyan Liu, and Lin Yi. Fine granularity resource allocation of virtual data center with consideration of virtual switches. Journal of Network and Computer Applications, 175(??):??, February 1, 2021. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520303751.

Yu:2006:FWV

[YGN^{+}06] Yang Yu, Fanglu Guo, Susanta Nanda, Lap chung Lam, and Tzi cker Chiueh. A feather-weight virtual machine for Windows applications. In ACM [ACM06f], pages 24–34. ISBN 1-59593-332-6 (??invalid ISBN??). LCCN QA76.9.V4.

Younes:2024:FDA

[YI24] Sana Younes and Maroua Idi. Formal dependability analysis of fault tolerant Virtual Machine allocation strategies in Cloud Radio Access Network. Journal of Network and Computer Applications, 229(??):??, September 2024. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000948.

Younes:2024:DTM

[YIR24] Sana Younes, Maroua Idi, and Riadh Robbana. Discretetime Markov decision process for performance analysis of virtual machine allocation schemes in C-RAN. Journal of Network and Computer Applications, 225(??):??, May 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000365.

Yang:2021:SHS

[YJZ^{+}21] Ye Yang, Haiyang Jiang, Guangxing Zhang, Xin Wang, Yilong Lv, Xing Li, Serge Fdida, and Gaogang Xie. S2H: Hypervisor as a setter within Virtualized Network I/O for VM isolation on cloud platform. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128621004862.

Yan:2012:VCH

[YJZY12] Lok-Kwong Yan, Manjukumar Jayachandra, Mu Zhang, and Heng Yin. V2E: combining hardware virtualization and software emulation for transparent and extensible malware analysis. ACM SIGPLAN Notices, 47(7):227–238, July 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Yamada:2013:TFT

[YK13] Hiroshi Yamada and Kenji Kono. Traveling forward in time to newer operating systems using ShadowReboot. ACM SIG-PLAN Notices, 48(7):121–130, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Yang:2017:EJV

[YKM17] Byung-Sun Yang, Jae-Yun Kim, and Soo-Mook Moon. Exceptionization: a Java VM optimization for non-Java languages. ACM Transactions on Architecture and Code Optimization, 14(1):5:1–5:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yoo:2024:EAO

[YKP^{+}24] Seehwan Yoo, Youngpil Kim, Hyunchan Park, Jaehyun Hwang, and Kitak Kim. enCloud: Aspect-oriented trusted service migration on SGX-enabled cloud VM. Software— Practice and Experience, 54(12):2454–2480, December 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yamanaka:2016:TFF

[YKS16] Hiroaki Yamanaka, Eiji Kawai, and Shinji Shimojo. A technique for full flow virtualization of multi-tenant Open-Flow networks. Computer Networks (Amsterdam, Netherlands: 1999), 102(??):1–19, June 19, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1389128616300585.

Yang:2017:VMM

[YLCH17] Chao-Tung Yang, Jung-Chun Liu, Shuo-Tsung Chen, and Kuan-Lung Huang. Virtual machine management system based on the power saving algorithm in cloud. Journal of Network and Computer Applications, 80(??):165–180, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print),

1095-8592 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1084804516302971.

Yang:2014:ICV

[YLH14] Chao-Tung Yang, Jung-Chun Liu, and Ching-Hsien Hsu. On improvement of cloud virtual machine availability with virtualization fault tolerance mechanism. The Journal of Supercomputing, 69(3):1103–1122, September 2014. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11227-013-1045-1.

Yan:2017:CAE

[YLH17] Fangfang Yan, Tony T. Lee, and Weisheng Hu. Congestionaware embedding of heterogeneous bandwidth virtual data centers with hose model abstraction. IEEE/ACM Transactions on Networking, 25(2):806–819, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yang:2014:MMG

[YLHJ14] Chao-Tung Yang, Jung-Chun Liu, Kuan-Lung Huang, and Fuu-Cheng Jiang. A method for managing green power of a virtual machine cluster in cloud. Future Generation Computer Systems, 37(??):26–36, July 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X14000466.

Yin:2022:VDC

[YLJ22] Chunxia Yin, Jian Liu, and Shunfu Jin. A virtualized data center energy-saving mechanism based on switching operating mode of physical servers and reserving virtual machines. Concurrency and Computation: Practice and Experience, 34 (9):e5785:1–e5785:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ye:2010:EES

[YLK^{+}10] Lei Ye, Gen Lu, Sushanth Kumar, Chris Gniady, and John H. Hartman. Energy-efficient storage in virtual machine environments. ACM SIGPLAN Notices, 45(7):75–84, July 2010. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yi:2017:CDC

[YLN^{+}17] Xiaomeng Yi, Fangming Liu, Di Niu, Hai Jin, and John C. S. Lui. Cocoa: Dynamic container-based group buying strategies for cloud computing. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 2(2):8:1–8:31, May 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://dl.acm.

org/citation.cfm?id=3022876.

Yang:2025:TDV

[YLQL25] Ran Yang, Zhaonan Li, Junhao Qian, and Zhihua Li. Taskdriven virtual machine optimization placement model and algorithm. Future Internet, 17(2):73, February 7, 2025. CO-DEN ???? ISSN 1999-5903. URL https://www.mdpi.com/

1999-5903/17/2/73.

Yao:2023:EOG

[YLT^{+}23] Jianguo Yao, Qiumin Lu, Run Tian, Keqin Li, and Haibing Guan. An economy-oriented GPU virtualization with dynamic and adaptive oversubscription. IEEE Transactions on Computers, 72(5):1371–1383, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2020:TRS

[YLTF20] Song Yang, Fan Li, Stojan Trajanovski, and Xiaoming Fu. Traffic routing in stochastic network function virtualization networks. Journal of Network and Computer Applications, 169(??):??, November 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804520302393.

Yang:2014:IGV

[YLWH14] Chao-Tung Yang, Jung-Chun Liu, Hsien-Yi Wang, and Ching-Hsien Hsu. Implementation of GPU virtualization using PCI pass-through mechanism. The Journal of Supercomputing, 68(1):183–213, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

link.springer.com/article/10.1007/s11227-013-1034-
4.

Yang:2005:LMJ

[YME05] Byung-Sun Yang, Soo-Mook Moon, and Kemal Ebcioğlu. Lightweight monitors for the Java virtual machine. Software —Practice and Experience, 35(3):281–299, March 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yao:2018:SSG

[YML^{+}18] Zhihao Yao, Zongheng Ma, Yingtong Liu, Ardalan Amiri Sani, and Aparna Chandramowlishwaran. Sugar: Secure GPU acceleration in Web browsers. ACM SIGPLAN Notices, 53(2):519–534, February 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Youssef:2017:WGE

[YMY17] Ahmed Youssef, Muthucumaru Maheswaran, and Lamis Youssef. Wireless GINI: an educational platform for hosting virtual wireless networks. Software—Practice and Experience, 47(1):21–59, January 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Young:1973:EAH

[You73] C. J. Young. Extended architecture and hypervisor performance. In ????, editor, Proceedings ACM SIGARCH-SIGOPS Workshop on Virtual Computer Systems, Cambridge, MA, 1973, page ?? ACM Press, New York, NY 10036, USA, 1973.

Yoginath:2015:EPD

[YP15] Srikanth B. Yoginath and Kalyan S. Perumalla. Efficient parallel discrete event simulation on cloud/virtual machine platforms. ACM Transactions on Modeling and Computer Simulation, 26(1):5:1–5:??, December 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Yang:2017:EEV

[YPLZ17] Ting Yang, Haibo Pen, Wei Li, and Albert Y. Zomaya. An energy-efficient virtual machine placement and route scheduling scheme in data center networks. Future Generation Computer Systems, 77(??):1–11, December 2017. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X17310579.

Yu:2014:MPP

[YQZ14] Chao Yu, Leihua Qin, and Jingli Zhou. A multicore periodical preemption virtual machine scheduling scheme to improve the performance of computational tasks. The Journal of Supercomputing, 67(1):254–276, January 2014. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11227-013-0998-4.

Yu:2019:LAV

[YQZ19] Chao Yu, Leihua Qin, and Jingli Zhou. A lock-aware virtual machine scheduling scheme for synchronization performance. The Journal of Supercomputing, 75(1):20–32, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yousefipour:2018:ECA

[YRJ18] Amin Yousefipour, Amir Masoud Rahmani, and Mohsen Jahanshahi. Energy and cost-aware virtual machine consolidation in cloud computing. Software—Practice and Experience, 48(10):1758–1774, October 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yilmaz:2021:FGC

[YSM^{+}21] Fadi Yilmaz, Meera Sridhar, Abhinav Mohanty, Vasant Tendulkar, and Kevin W. Hamlen. A fine-grained classification and security analysis of web-based virtual machine vulnerabilities. Computers & Security, 105(??):Article 102246, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000705.

Yeh:2017:PFG

[YSS^{+}17] Tsung Tai Yeh, Amit Sabne, Putt Sakdhnagool, Rudolf Eigenmann, and Timothy G. Rogers. Pagoda: Fine-grained GPU resource virtualization for narrow tasks. ACM SIG-PLAN Notices, 52(8):221–234, August 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yan:2014:EFG

[YTS14] Jie Yan, Guangming Tan, and Ninghui Sun. Exploiting fine-grained parallelism in graph traversal algorithms via lock virtualization on multi-core architecture. The Journal of Supercomputing, 69(3):1462–1490, September 2014. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11227-014-1239-1.

Yutaka:2000:EJV

[YTY00] Oiwa Yutaka, Kenjiro Taura, and Akinori Yonezawa. Extending Java virtual machine with integer-reference conversion. Concurrency: Practice and Experience, 12(6): 407–422, May 2000. CODEN CPEXEI. ISSN 1040-

3108. URL http://www3.interscience.wiley.com/cgi-
bin/abstract/72515729/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=72515729&PLACEBO=IE.
pdf.

Yu:2020:CCC

[Yu20] Linchen Yu. CCHybrid: CPU co-scheduling in virtualization environment. Concurrency and Computation: Practice and Experience, 32(3):e4213:1–e4213:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yurcik:2002:SIS

[Yur02] William (Bill) Yurcik. Special issue on specialized computer architecture simulators that see the present and may hold the future. ACM Journal on Educational Resources in Computing (JERIC), 2(1):1–3, March 2002. CODEN ???? ISSN 1531-4278.

Yan:2017:HTC

[YVCB17] Zi Yan, Ján Veselý, Guilherme Cox, and Abhishek Bhattacharjee. Hardware translation coherence for virtualized systems. ACM SIGARCH Computer Architecture News, 45(2): 430–443, May 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Yan:2018:HTC

[YVCB18] Zi Yan, Ján Veselý, Guilherme Cox, and Abhishek Bhattacharjee. Hardware translation coherence for virtualized systems. Operating Systems Review, 52(1):57–70, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yang:2020:IES

[YW20] Chao-Tung Yang and Tsung-Yueh Wan. Implementation of an energy saving cloud infrastructure with virtual machine power usage monitoring and live migration on Open-Stack. Computing: Archiv für Informatik und Numerik, 102 (6):1547–1566, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Younge:2015:SHP

[YWCF15] Andrew J. Younge, John Paul Walters, Stephen P. Crago, and Geoffrey C. Fox. Supporting high performance molecular dynamics in virtualized clusters using IOMMU, SR-IOV, and GPUDirect. ACM SIGPLAN Notices, 50(7):31–38, July 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yermolovich:2009:ODL

[YWF09] Alexander Yermolovich, Christian Wimmer, and Michael Franz. Optimization of dynamic languages using hierarchical layering of virtual machines. ACM SIGPLAN Notices, 44 (12):79–88, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yu:2013:OSI

[YWGH13] Yulong Yu, Yuxin Wang, He Guo, and Xubin He. Optimisation schemes to improve hybrid co-scheduling for concurrent virtual machines. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(1):46–66, 2013. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Ye:2021:SSD

[YWH^{+}21] Z. Ye, Y. Wang, S. He, C. Xu, and X. Sun. Sova: a software-defined autonomic framework for virtual network allocations. IEEE Transactions on Parallel and Distributed Systems, 32(1):116–130, 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yao:2023:EEL

[YWH^{+}23] Wenbin Yao, Zhuqing Wang, Yingying Hou, Xikang Zhu, Xiaoyong Li, and Yamei Xia. An energy-efficient load balance strategy based on virtual machine consolidation in cloud environment. Future Generation Computer Systems, 146(??):222–233, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X23001498.

Yi:2018:CSN

[YWL^{+}18] Bo Yi, Xingwei Wang, Keqin Li, Sajal K. Das, and Min Huang. A comprehensive survey of Network Function Virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 133(??):212–262, March 14, 2018. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128618300306.

Yao:2014:GFT

[YWR^{+}14] Lin Yao, Guowei Wu, Jiankang Ren, Yanwei Zhu, and Ying Li. Guaranteeing fault-tolerant requirement load balancing scheme based on VM migration. The Computer Journal, 57 (2):225–232, February 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/57/2/225.full.pdf+html.

You:2015:VFO

[YWTC15] Yi-Ping You, Hen-Jung Wu, Yeh-Ning Tsai, and Yen-Ting Chao. VirtCL: a framework for OpenCL device abstraction and management. ACM SIGPLAN Notices, 50(8):161–172, August 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ye:2015:PBW

[YWW^{+}15] Kejiang Ye, Zhaohui Wu, Chen Wang, Bing Bing Zhou, Weisheng Si, Xiaohong Jiang, and Albert Y. Zomaya. Profiling-based workload consolidation and migration in virtualized data centers. IEEE Transactions on Parallel and Distributed Systems, 26(3):878–890, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://www.computer.org/csdl/trans/td/
2015/03/06777568-abs.html.

Yang:2017:RVM

[YWY^{+}17] Song Yang, Philipp Wieder, Ramin Yahyapour, Stojan Trajanovski, and Xiaoming Fu. Reliable virtual machine placement and routing in clouds. IEEE Transactions on Parallel and Distributed Systems, 28(10):2965–2978, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2017/10/07896612-abs.html.

Yao:2020:JOF

[YXL^{+}20] Hong Yao, Muzhou Xiong, Hui Li, Lin Gu, and Deze Zeng. Joint optimization of function mapping and preemptive scheduling for service chains in network function virtualization. Future Generation Computer Systems, 108(??):1112– 1118, July 2020. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167739X17324305.

Yao:2019:EVM

[YYC^{+}19] Feng Yao, Yiping Yao, Huangke Chen, Tianlin Li, Menglong Lin, and Xiaoxiong Zhang. An efficient virtual machine allocation algorithm for parallel and distributed simulation applications. Concurrency and Computation: Practice and Experience, 31(17):e5237:1–e5237:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yi:2015:ESF

[YYL^{+}15] Qiuping Yi, Zijiang Yang, Jian Liu, Chen Zhao, and Chao Wang. Explaining software failures by cascade fault localization. ACM Transactions on Design Automation of Electronic Systems., 20(3):41:1–41:??, June 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yehezkel:2001:TST

[YYPA01] Cecile Yehezkel, William Yurcik, Murray Pearson, and Dean Armstrong. Three simulator tools for teaching computer architecture: Little Man computer, and RTLSim. ACM Journal on Educational Resources in Computing (JERIC), 1(4): 60–80, December 2001. CODEN ???? ISSN 1531-4278.

Yang:2023:HIH

[YYY^{+}23] Xingyuan Yang, Jie Yuan, Hao Yang, Ya Kong, Hao Zhang, and Jinyu Zhao. A highly interactive honeypot-based approach to network threat management. Future Internet, 15 (4):127, March 28, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.com/1999-5903/15/4/127.

Yang:2014:IIV

[YZLQ14] Hailong Yang, Qi Zhao, Zhongzhi Luan, and Depei Qian. iMeter: an integrated VM power model based on performance profiling. Future Generation Computer Systems, 36(??):267– 286, July 2014. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167739X13001532.

Yut:2017:LRL

[YZSC17] Lele Yut, Ce Zhang, Yingxia Shao, and Bin Cui. LDA*: a robust and large-scale topic modeling system. Proceedings of the VLDB Endowment, 10(11):1406–1417, August 2017. CODEN ???? ISSN 2150-8097.

Yang:2013:QSE

[YZW^{+}13] Yan Yang, Yulong Zhang, Alex Hai Wang, Meng Yu, Wanyu Zang, Peng Liu, and Sushil Jajodia. Quantitative survivability evaluation of three virtual machine-based server architectures. Journal of Network and Computer Applications, 36(2):781–790, March 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1084804512002536.

Zhao:2016:SHC

[ZAI^{+}16] Qian Zhao, Motoki Amagasaki, Masahiro Iida, Morihiro Kuga, and Toshinori Sueyoshi. A study of heterogeneous computing design method based on virtualization technology. ACM SIGARCH Computer Architecture News, 44(4): 86–91, September 2016. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Ziafat:2018:OSV

[ZB18] Hassan Ziafat and Seyed Morteza Babamir. Optimal selection of VMs for resource task scheduling in geographically distributed clouds using fuzzy c-mean and MOLP. Software —Practice and Experience, 48(10):1820–1846, October 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhong:2020:CEC

[ZB20] Zhiheng Zhong and Rajkumar Buyya. A cost-efficient container orchestration strategy in Kubernetes-based cloud computing infrastructures with heterogeneous resources. ACM Transactions on Internet Technology (TOIT), 20(2):15:1– 15:24, May 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3378447.

Zhang:2005:FVM

[ZBG^{+}05] Yuting Zhang, Azer Bestavros, Mina Guirguis, Ibrahim Matta, and Richard West. Friendly virtual machines: leveraging a feedback-control model for application adaptation. In ACM [ACM05d], pages 2–12. ISBN 1-59593-047-7. LCCN

QA76.9.V5 I575 2005. URL http://www.loc.gov/catdir/
toc/fy0611/2006530661.html. ACM order number 548059.

Zhao:2005:SSV

[ZBP05] Xin Zhao, Kevin Borders, and Atul Prakash. SVGrid: a secure virtual environment for untrusted grid applications. In ACM [ACM05b], pages 1–6. ISBN 1-59593-269-0. LCCN ????

Zhao:2007:UVM

[ZBP07] Xin Zhao, Kevin Borders, and Atul Prakash. Using a virtual machine to protect sensitive Grid resources. Concurrency and Computation: Practice and Experience, 19(14): 1917–1935, September 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zou:2015:CDA

[ZBS^{+}15] Tao Zou, Ronan Bras, Marcos Vaz Salles, Alan Demers, and Johannes Gehrke. ClouDiA: a deployment advisor for public clouds. VLDB Journal: Very Large Data Bases, 24(5): 633–653, October 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhao:2022:SEE

[ZBS^{+}22] Zihan Zhao, Sidi Mohamed Beillahi, Ryan Song, Yuxi Cai, Andreas Veneris, and Fan Long. SigVM: enabling eventdriven execution for truly decentralized smart contracts. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):149:1–149:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL https://dl.acm.org/

doi/10.1145/3563312.

Zhang:2017:CAV

[ZCG^{+}17] Weizhan Zhang, Yuxuan Chen, Xiang Gao, Zhichao Mo, Qinghua Zheng, and Zongqing Lu. Cluster-aware virtual machine collaborative migration in media cloud. IEEE Transactions on Parallel and Distributed Systems, 28(10):2808–2822, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print),

1558-2183 (electronic). URL https://www.computer.org/
csdl/trans/td/2017/10/07909001-abs.html.

Zhan:2021:CAW

[ZCJ^{+}21] Yang Zhan, Alex Conway, Yizheng Jiao, Nirjhar Mukherjee, Ian Groombridge, Michael A. Bender, Martin Farach-Colton, William Jannen, Rob Johnson, Donald E. Porter, and Jun Yuan. Copy-on-abundant-write for nimble file system clones. ACM Transactions on Storage, 17(1):5:1–5:27, February 2021. CODEN ???? ISSN 1553-3077 (print), 1553-

3093 (electronic). URL https://dl.acm.org/doi/10.1145/
3423495.

Zhang:2021:TVM

[ZCL^{+}21] Jiachen Zhang, Lixiao Cui, Peng Li, Xiaoguang Liu, and Gang Wang. Toward virtual machine image management for persistent memory. ACM Transactions on Storage, 17 (3):20:1–20:24, August 2021. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic). URL https://dl.acm.org/

doi/10.1145/3450976.

Zhao:2018:FFI

[ZD18] Siqi Zhao and Xuhua Ding. FIMCE: A fully isolated microcomputing environment for multicore systems. ACM Transactions on Privacy and Security (TOPS), 21(3):15:1–15:??, June 2018. ISSN 2471-2566 (print), 2471-2574 (electronic).

URL https://dl.acm.org/citation.cfm?id=3195181.

Zeuch:2019:AES

[ZDK^{+}19] Steffen Zeuch, Bonaventura Del Monte, Jeyhun Karimov, Clemens Lutz, Manuel Renz, Jonas Traub, Sebastian Breß, Tilmann Rabl, and Volker Markl. Analyzing efficient stream processing on modern hardware. Proceedings of the VLDB Endowment, 12(5):516–530, January 2019. CODEN ???? ISSN 2150-8097.

Zeng:2022:ADB

[ZDK^{+}22] Jing Zeng, Ding Ding, Kaixuan Kang, HuaMao Xie, and Qian Yin. Adaptive DRL-based virtual machine consolidation in energy-efficient cloud data center. IEEE Transactions on Parallel and Distributed Systems, 33(11):2991–3002, November 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2017:MAP

[ZDLG17] Jinshi Zhang, Eddie Dong, Jian Li, and Haibing Guan. MigVisor: Accurate prediction of VM live migration behavior using a working-set pattern model. ACM SIGPLAN Notices, 52(7):30–43, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zeng:2022:UFV

[ZDS^{+}22] Shulin Zeng, Guohao Dai, Hanbo Sun, Jun Liu, Shiyao Li, Guangjun Ge, Kai Zhong, Kaiyuan Guo, Yu Wang, and Huazhong Yang. A unified FPGA virtualization framework for general-purpose deep neural networks in the cloud. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(3):24:1–24:31, September 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3480170.

Zamorano:2013:ART

[ZEdlP13] Juan Zamorano, ’Angel Esquinas, and Juan A. de la Puente. Ada real-time services and virtualization. ACM SIGADA Ada Letters, 33(1):128–133, April 2013. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Zhang:2022:CST

[ZFH^{+}22] Yao Zhang, Wenping Fan, Qichen Hao, Xinya Wu, and Min-Ling Zhang. CAFE and SOUP: Toward adaptive VDI workload prediction. ACM Transactions on Intelligent Systems and Technology (TIST), 13(6):94:1–94:??, December 2022. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3529536.

Zeng:2015:PPH

[ZFL15] Junyuan Zeng, Yangchun Fu, and Zhiqiang Lin. PEMU: a pin highly compatible out-of-VM dynamic binary instrumentation framework. ACM SIGPLAN Notices, 50(7):147–160, July 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhao:2023:VPA

[ZFL^{+}23] Hui Zhao, Nanzhi Feng, Jianhua Li, Guobin Zhang, Jing Wang, Quan Wang, and Bo Wan. VM performance-aware virtual machine migration method based on ant colony optimization in cloud environment. Journal of Parallel and Distributed Computing, 176(??):17–27, June 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000187.

Zhang:2018:LFV

[ZFY18] Fei Zhang, Xiaoming Fu, and Ramin Yahyapour. Layer-Mover: Fast virtual machine migration over WAN with three-layer image structure. Future Generation Computer Systems, 83(??):37–49, June 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167739X1630797X.

Zaman:2013:CAB

[ZG13] Sharrukh Zaman and Daniel Grosu. Combinatorial auctionbased allocation of virtual machine instances in clouds. Journal of Parallel and Distributed Computing, 73(4):495–508, April 2013. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731512002870.

Zinner:2017:DTM

[ZGL^{+}17] Thomas Zinner, Stefan Geissler, Stanislav Lange, Steffen Gebert, Michael Seufert, and Phuoc Tran-Gia. A discretetime model for optimizing the processing time of virtualized network functions. Computer Networks (Amsterdam, Netherlands: 1999), 125(??):4–14, October 9, 2017. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1389128617301809.

Zimmermann:2006:AHM

[ZGW^{+}06] Alexander Zimmermann, Mesut Günes, Martin Wenig, Jan Ritzerfeld, and Ulrich Meis. Architecture of the hybrid MCGmesh testbed. In ACM [ACM06c], pages 88–89. ISBN 1-59593-540-0. LCCN ????

Zhang:2015:LOS

[ZHCB15] Minjia Zhang, Jipeng Huang, Man Cao, and Michael D. Bond. Low-overhead software transactional memory with progress guarantees and strong semantics. ACM SIGPLAN Notices, 50(8):97–108, August 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2024:AAD

[ZHH^{+}24] Kai Zhang, Jiahui Hong, Zhengying He, Yinan Jing, and X. Sean Wang. AdaptChain: Adaptive data sharing and synchronization for NFV systems on heterogeneous architectures. IEEE Transactions on Parallel and Distributed Systems, 35(7):1281–1292, July 2024. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2017:NAV

[ZHHC17] Weizhe Zhang, Shuo Han, Hui He, and Huixiang Chen. Network-aware virtual machine migration in an overcommitted cloud. Future Generation Computer Systems, 76(??):428–442, November 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X1630053X.

Zhou:2016:VMP

[ZHL16] Zhou Zhou, Zhigang Hu, and Keqin Li. Virtual machine placement algorithm for both energy-awareness and SLA violation reduction in cloud data centers. Scientific Programming, 2016(??):5612039:1–5612039:11, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.hindawi.com/journals/sp/
2016/5612039/.

Zhou:2010:VN

[Zho10] Shudong Zhou. Virtual networking. Operating Systems Review, 44(4):80–85, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2017:OAI

[ZHW^{+}17] Xiaoxi Zhang, Zhiyi Huang, Chuan Wu, Zongpeng Li, and Francis C. M. Lau. Online auctions in IaaS clouds: Welfare and profit maximization with server costs. IEEE/ACM Transactions on Networking, 25(2):1034–1047, April 2017. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zimmer:2005:VMV

[Zim05] Dennis Zimmer. VMware and Microsoft Virtual Server: virtuelle Server im professionellen Einsatz; [VMware GSX, ESX und Microsoft Virtual Server; Virtualisierungssoftware im Vergleich; Planung, Installation und Verwaltung]. Galileo Computing. Galileo Press, Bonn, Germany, 2005. ISBN 3-89842-701-3. 612 pp. LCCN ???? EUR 49.90. URL http:/

/www.galileocomputing.de.

Zimmer:2006:VSV

[Zim06] Dennis Zimmer. VMware Server and VMware Player: [Installation, Anwendung und Konfiguration; Konzeption und Einsatzmöglichkeiten; virtuelle Maschinen erstellen und nutzen]. Galileo Computing. Galileo Press, Bonn, Germany, 2006. ISBN 3-89842-822-2. 358 pp. LCCN ????

Zhang:2019:TVN

[ZJRW19] Chuanji Zhang, Harshvardhan P. Joshi, George F. Riley, and Steven A. Wright. Towards a virtual network function research agenda: a systematic literature review of VNF design considerations. Journal of Network and Computer Applications, 146(??):Article 102417, ???? 2019. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519302516.

Zhu:2011:OPV

[ZJXL11] Jun Zhu, Zhefu Jiang, Zhen Xiao, and Xiaoming Li. Optimizing the performance of virtual machine synchronization for fault tolerance. IEEE Transactions on Computers, 60(12):1718–1729, December 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5629326.

Zhu:2017:NFV

[ZKWH17] Xiaoqing Zhu, Harilaos Koumaras, Mea Wang, and David Hausheer. Network function virtualization and softwaredefined networking: Advancing multimedia distribution. IEEE MultiMedia, 24(3):16–18, July/September 2017. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (elec-

tronic). URL https://www.computer.org/csdl/mags/mu/
2017/03/mmu2017030016.html.

Zhou:2013:LPC

[ZL13] Ruijin Zhou and Tao Li. Leveraging phase change memory to achieve efficient virtual machine execution. ACM SIGPLAN Notices, 48(7):179–190, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Zhang:2016:MAV

[ZL16] Tianwei Zhang and Ruby B. Lee. Monitoring and attestation of virtual machine security health in cloud computing. IEEE Micro, 36(5):28–37, September/October 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.computer.org/csdl/mags/mi/
2016/05/mmi2016050028-abs.html.

Zha:2018:LSM

[ZL18a] Yue Zha and Jing Li. Liquid Silicon-Monona: a reconfigurable memory-oriented computing fabric with scalable multi-context support. ACM SIGPLAN Notices, 53(2):214– 228, February 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2018:DIV

[ZL18b] Tianwei Zhang and Ruby B. Lee. Design, implementation and verification of cloud architecture for monitoring a virtual machine’s security health. IEEE Transactions on Computers, 67(6):799–815, June 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://ieeexplore.

ieee.org/document/8169039/.

Zhang:2014:AIO

[ZLBF14] Wei Zhang, Per Larsen, Stefan Brunthaler, and Michael Franz. Accelerating iterators in optimizing AST interpreters. ACM SIGPLAN Notices, 49(10):727–743, October 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhou:2018:DNA

[ZLCZ18] Hang Zhou, Qing Li, Kim-Kwang Raymond Choo, and Hai Zhu. DADTA: a novel adaptive strategy for energy and performance efficient virtual machine consolidation. Journal of Parallel and Distributed Computing, 121(??):53–70, November 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731518304520.

Zhang:2020:PER

[ZLG^{+}20] Sheng Zhang, Yu Liang, Jidong Ge, Mingjun Xiao, and Jie Wu. Provably efficient resource allocation for edge service entities using Hermes. IEEE/ACM Transactions on Networking, 28(4):1684–1697, August 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://dl.

acm.org/doi/10.1109/TNET.2020.2989307.

Zhang:2015:SSP

[ZLH^{+}15] Yonglong Zhang, Bin Li, Zhiqiu Huang, Jin Wang, and Junwu Zhu. SGAM: strategy-proof group buying-based auction mechanism for virtual machine allocation in clouds. Concurrency and Computation: Practice and Experience, 27(18): 5577–5589, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zabolotnyi:2015:JCG

[ZLHD15] Rostyslav Zabolotnyi, Philipp Leitner, Waldemar Hummer, and Schahram Dustdar. JCloudScale: Closing the gap between IaaS and PaaS. ACM Transactions on Internet Technology (TOIT), 15(3):10:1–10:??, September 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zheng:2016:VMC

[ZLL^{+}16] Qinghua Zheng, Rui Li, Xiuqi Li, Nazaraf Shah, Jianke Zhang, Feng Tian, Kuo-Ming Chao, and Jia Li. Virtual machine consolidated placement based on multi-objective biogeography-based optimization. Future Generation Computer Systems, 54(??):95–122, January 2016. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X15000564.

Zhang:2020:PEE

[ZLL^{+}20] Yiming Zhang, Huiba Li, Shengyun Liu, Jiawei Xu, and Guangtao Xue. PBS: an efficient erasure-coded block storage system based on speculative partial writes. ACM Transactions on Storage, 16(1):6:1–6:25, April 2020. CODEN ????

ISSN 1553-3077 (print), 1553-3093 (electronic). URL https:

//dl.acm.org/doi/abs/10.1145/3365839.

Zhou:2013:OVM

[ZLLL13] Ruijin Zhou, Fang Liu, Chao Li, and Tao Li. Optimizing virtual machine live storage migration in heterogeneous storage environment. ACM SIGPLAN Notices, 48(7):73–84, July 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’13 Conference proceedings.

Zhang:2017:MSM

[ZLSI17] Qi Zhang, Ling Liu, Gong Su, and Arun Iyengar. Mem-Flex: A shared memory swapper for high performance VM execution. IEEE Transactions on Computers, 66(9):1645– 1652, September 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://ieeexplore.

ieee.org/document/7885535/.

Zou:2012:CDA

[ZLV^{+}12] Tao Zou, Ronan Le Bras, Marcos Vaz Salles, Alan Demers, and Johannes Gehrke. ClouDiA: a deployment advisor for public clouds. Proceedings of the VLDB Endowment, 6(2): 121–132, December 2012. CODEN ???? ISSN 2150-8097.

Zhang:2014:VFP

[ZLW^{+}14] Zhaoning Zhang, Ziyang Li, Kui Wu, Dongsheng Li, Huiba Li, Yuxing Peng, and Xicheng Lu. VMThunder: Fast provisioning of large-scale virtual machine clusters. IEEE Transactions on Parallel and Distributed Systems, 25(12):3328– 3338, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.computer.

org/csdl/trans/td/2014/12/06719385-abs.html.

Zhou:2018:SFC

[ZLW18] Ruiting Zhou, Zongpeng Li, and Chuan Wu. Scheduling frameworks for cloud container services. IEEE/ACM Transactions on Networking, 26(1):436–450, February 2018. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Zhao:2019:RUC

[ZLW^{+}19a] Yan Zhao, Hongwei Liu, Yan Wang, Zhan Zhang, and Decheng Zuo. Reducing the upfront cost of private clouds with clairvoyant virtual machine placement. The Journal of Supercomputing, 75(1):340–369, January 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhong:2019:TFL

[ZLW^{+}19b] K. Zhong, D. Liu, Y. Wu, L. Long, W. Liu, J. Ren, R. Liu, L. Liang, Z. Shao, and T. Li. Towards fast and lightweight checkpointing for mobile virtualization using NVRAM. IEEE Transactions on Parallel and Distributed Systems, 30(6): 1421–1433, June 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2013:ASD

[ZLZ13] Youhui Zhang, Yanhua Li, and Weimin Zheng. Automatic software deployment using user-level virtualization for cloud-computing. Future Generation Computer Systems, 29(1):323–329, January 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X11001518.

Zhang:2015:MCV

[ZLZ15] Xinyan Zhang, Keqiu Li, and Yong Zhang. Minimum-cost virtual machine migration strategy in datacenter. Concurrency and Computation: Practice and Experience, 27(17): 5177–5187, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2019:RNO

[ZLZ^{+}19a] Fei Zhang, Guangming Liu, Bo Zhao, Xiaoming Fu, and Ramin Yahyapour. Reducing the network overhead of user mobility-induced virtual machine migration in mobile edge computing. Software—Practice and Experience, 49(4):673– 693, April 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2019:CFV

[ZLZ^{+}19b] Fei Zhang, Guangming Liu, Bo Zhao, Piotr Kasprzak, Xiaoming Fu, and Ramin Yahyapour. CBase: Fast virtual machine storage data migration with a new data center structure. Journal of Parallel and Distributed Computing, 124(??):14–26, February 2019. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0743731518307342.

Zhang:2021:OAI

[ZLZ21a] Qixia Zhang, Fangming Liu, and Chaobing Zeng. Online adaptive interference-aware VNF deployment and migration for 5G network slice. IEEE/ACM Transactions on Networking, 29(5):2115–2128, October 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://dl.

acm.org/doi/10.1109/TNET.2021.3080197.

Zhang:2021:VGA

[ZLZ^{+}21b] Xiaoli Zhang, Qi Li, Zeyu Zhang, Jianping Wu, and Jiahai Yang. vSFC: Generic and agile verification of service function chains in the cloud. IEEE/ACM Transactions on Networking, 29(1):78–91, February 2021. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https://dl.acm.

org/doi/10.1109/TNET.2020.3028846.

Zerouali:2021:MDA

[ZMD^{+}21] Ahmed Zerouali, Tom Mens, Alexandre Decan, Jesus Gonzalez-Barahona, and Gregorio Robles. A multidimensional analysis of technical lag in Debian-based Docker images. Empirical Software Engineering, 26(2):??, March 2021. CODEN ESENFW. ISSN 1382-3256 (print),

1573-7616 (electronic). URL http://link.springer.com/
article/10.1007/s10664-020-09908-6;   http://link.
springer.com/content/pdf/10.1007/s10664-020-09908-
6.pdf.

Zheng:2014:CCM

[ZNSL14] Jie Zheng, Tze Sing Eugene Ng, Kunwadee Sripanidkulchai, and Zhaolei Liu. COMMA: coordinating the migration of multi-tier applications. ACM SIGPLAN Notices, 49(7):153– 164, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zakkak:2014:JJM

[ZP14] Foivos S. Zakkak and Polyvios Pratikakis. JDMM: a Java memory model for non-cache-coherent memory architectures. ACM SIGPLAN Notices, 49(11):83–92, November 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2016:CGS

[ZQCZ16] Youhui Zhang, Peng Qu, Jiang Cihang, and Weimin Zheng. A cloud gaming system based on user-level virtualization and its resource scheduling. IEEE Transactions on Parallel and Distributed Systems, 27(5):1239–1252, May 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://www.computer.org/csdl/trans/td/
2016/05/07109163.pdf.

Zoppke:2006:VLE

[ZR06] Till Zoppke and Raúl Rojas. The virtual life of ENIAC: Simulating the operation of the first electronic computer. IEEE Annals of the History of Computing, 28(2):18–25, April/June 2006. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Zhang:2015:MIM

[ZRD^{+}15] Wei Zhang, Sundaresan Rajasekaran, Shaohua Duan, Timothy Wood, and Mingfa Zhuy. Minimizing interference and maximizing progress for Hadoop virtual machines. ACM SIGMETRICS Performance Evaluation Review, 42(4):62–71, March 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2016:GDL

[ZRS^{+}16] Jiao Zhang, Fengyuan Ren, Ran Shu, Tao Huang, and Yunjie Liu. Guaranteeing delay of live virtual machine migration by determining and provisioning appropriate bandwidth. IEEE Transactions on Computers, 65(9):2910–2917, ???? 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2015:UPP

[ZRZY15] Yong Zhao, Jia Rao, Xiaobo Zhou, and Qing Yi. Understanding parallel performance under interferences in multitenant clouds. ACM SIGMETRICS Performance Evaluation Review, 43(1):447–448, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zhang:2001:HJAb

[ZS01] Xiaolan Zhang and Margo Seltzer. HBench:Java: an application-specific benchmarking framework for Java Virtual Machines. Concurrency and Computation: Practice and Experience, 13(8–9):775–792, July/August 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/84503219/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=84503219&PLACEBO=IE.
pdf.

Zahedi:2025:BEI

[ZSE^{+}25] Mahdi Zahedi, Taha Shahroodi, Carlos Escuin, Georgi Gaydadjiev, Stephan Wong, and Said Hamdioui. BCIM: Efficient implementation of binary neural network based on computation in memory. IEEE Transactions on Emerging Topics in Computing, 13(2):395–408, 2025. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2021:CHP

[ZSP^{+}21] Xuzhi Zhang, Xiaozhe Shao, George Provelengios, Naveen Kumar Dumpala, Lixin Gao, and Russell Tessier. CoNFV: a heterogeneous platform for scalable network function virtualization. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 14(1):1:1–1:29, January 2021. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

URL https://dl.acm.org/doi/10.1145/3409113.

Zhang:2005:ILS

[ZSR^{+}05] Jianyong Zhang, Anand Sivasubramaniam, Alma Riska, Qian Wang, and Erik Riedel. An interposed 2-level I/O scheduling framework for performance virtualization. ACM SIGMET-RICS Performance Evaluation Review, 33(1):406–407, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Zolfaghari:2022:EAV

[ZSRR22] Rahmat Zolfaghari, Amir Sahafi, Amir Masoud Rahmani, and Reza Rezaei. An energy-aware virtual machines consolidation method for cloud computing: Simulation and verification. Software—Practice and Experience, 52(1):194–235, January 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2006:SPV

[ZSW^{+}06] Jianyong Zhang, Anand Sivasubramaniam, Qian Wang, Alma Riska, and Erik Riedel. Storage performance virtualization via throughput and latency control. ACM Transactions on Storage, 2(3):283–308, August 2006. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Zhang:2007:DIB

[ZSXZ07] Guangyan Zhang, Jiwu Shu, Wei Xue, and Weimin Zheng. Design and implementation of an out-of-band virtualization system for large SANs. IEEE Transactions on Computers, 56(12):1654–1665, December 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
4358222.

Zhao:2021:LSA

[ZTA^{+}21] N. Zhao, V. Tarasov, H. Albahar, A. Anwar, L. Rupprecht, D. Skourtis, A. K. Paul, K. Chen, and A. R. Butt. Largescale analysis of Docker images and performance implications for container storage systems. IEEE Transactions on Parallel and Distributed Systems, 32(4):918–930, 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2017:VLV

[ZTWM17] Min Zhu, Bibo Tu, Wei Wei, and Dan Meng. HA-VMSI: a lightweight virtual machine isolation approach with commodity hardware for ARM. ACM SIGPLAN Notices, 52(7):242– 256, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zou:2014:VOV

[ZWC^{+}14] Shihong Zou, Xitao Wen, Kai Chen, Shan Huang, Yan Chen, Yongqiang Liu, Yong Xia, and Chengchen Hu. VirtualKnotter: Online virtual machine shuffling for congestion resolving in virtualized datacenter. Computer Networks (Amsterdam, Netherlands: 1999), 67(??):141–153, July 4, 2014.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S138912861400139X.

Zhang:2019:EAV

[ZWC^{+}19] Xinqian Zhang, Tingming Wu, Mingsong Chen, Tongquan Wei, Junlong Zhou, Shiyan Hu, and Rajkumar Buyya. Energy-aware virtual machine allocation for cloud with resource reservation. The Journal of Systems and Software, 147(??):147–161, January 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0164121218302152.

Zhao:2023:FTM

[ZWC^{+}23] Ziming Zhao, Mingyu Wu, Xujie Cao, Haibo Chen, and Binyu Zang. Flock: Towards multitasking virtual machines for function-as-a-service. IEEE Transactions on Computers, 72 (11):3153–3166, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zeng:2017:RNN

[ZWFX17] Lingfang Zeng, Yang Wang, Xiaopeng Fan, and Chengzhong Xu. Raccoon: A novel network I/O allocation framework for workload-aware VM scheduling in virtual environments. IEEE Transactions on Parallel and Distributed Systems, 28(9):2651–2662, September 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2017/09/07883963-abs.html.

Zhou:2017:NFA

[ZWH^{+}17] Ao Zhou, Shangguang Wang, Ching-Hsien Hsu, Myung Ho Kim, and Kok seng Wong. Network failure-aware redundant virtual machine placement in a cloud data center. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2017:CBV

[ZWHC17] Jiangtao Zhang, Xuan Wang, Hejiao Huang, and Shi Chen. Clustering based virtual machines placement in distributed cloud computing. Future Generation Computer Systems, 66(??):1–10, January 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X16302035.

Zeng:2017:NSD

[ZWKX17] Lingfang Zeng, Yang Wang, Kenneth B. Kent, and Ziliang Xiao. Naplus: a software distributed shared memory for virtual clusters in the cloud. Software—Practice and Experience, 47(9):1201–1220, September 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhao:2009:DMB

[ZWL09] Weiming Zhao, Zhenlin Wang, and Yingwei Luo. Dynamic memory balancing for virtual machines. Operating Systems Review, 43(3):37–47, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhao:2018:PAP

[ZWL^{+}18] Hui Zhao, Jing Wang, Feng Liu, Quan Wang, Weizhan Zhang, and Qinghua Zheng. Power-aware and performanceguaranteed virtual machine placement in the cloud. IEEE Transactions on Parallel and Distributed Systems, 29 (6):1385–1400, June 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:

//www.computer.org/csdl/trans/td/2018/06/08259446-
abs.html.

Zhao:2020:PHV

[ZWZ20] Sicheng Zhao, Xing Wu, and Zuqing Zhu. On parallel and hitless vSDN reconfiguration. IEEE/ACM Transactions on Networking, 28(6):2657–2670, December 2020. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL https://dl.acm.org/doi/10.1109/TNET.
2020.3014655.

Zhang:2024:CEV

[ZXB^{+}24] Zheng Zhang, Jingfeng Xue, Thar Baker, Tian Chen, Yu an Tan, and Yuanzhang Li. COVER: Enhancing virtualization obfuscation through dynamic scheduling using flash controller-based secure module. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0167404824003432.

Zhao:2024:EML

[ZXG^{+}24] Xiaotian Zhao, Ruge Xu, Yimin Gao, Vaibhav Verma, Mircea R. Stan, and Xinfei Guo. Edge-MPQ: Layer-wise mixed-precision quantization with tightly integrated versatile inference units for edge computing. IEEE Transactions on Computers, 73(11):2504–2519, November 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:IAB

[ZXL^{+}24] Zongpu Zhang, Chenbo Xia, Cunming Liang, Jian Li, Chen Yu, Tiwei Bie, Roberts Martin, Daly Dan, Xiao Wang, Yong Liu, and Haibing Guan. Un-IOV: Achieving bare-metal level I/O virtualization performance for cloud usage with migratability, scalability and transparency. IEEE Transactions on Computers, 73(7):1655–1668, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhong:2022:MLB

[ZXR^{+}22] Zhiheng Zhong, Minxian Xu, Maria Alejandra Rodriguez, Chengzhong Xu, and Rajkumar Buyya. Machine learningbased orchestration of containers: a taxonomy and future directions. ACM Computing Surveys, 54(10s):217:1–217:??,

January 2022. URL https://dl.acm.org/doi/10.1145/
3510415.

Zeng:2016:VEF

[ZXW16] Lingfang Zeng, Shijie Xu, and Yang Wang. VMBackup: an efficient framework for online virtual machine image backup and recovery. Concurrency and Computation: Practice and Experience, 28(9):2630–2643, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhong:2015:VBM

[ZXY^{+}15] Xianming Zhong, Chengcheng Xiang, Miao Yu, Zhengwei Qi, and Haibing Guan. A virtualization based monitoring system for mini-intrusive live forensics. International Journal of Parallel Programming, 43(3):455–471, June 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10766-013-0285-2.

Zhang:2019:AMD

[ZYH^{+}19] Lei Zhang, Zhemin Yang, Yuyu He, Mingqi Li, Sen Yang, Min Yang, Yuan Zhang, and Zhiyun Qian. App in the middle: Demystify application virtualization in Android and its security threats. ACM SIGMETRICS Performance Evaluation Review, 47(1):75–76, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:

//dl.acm.org/doi/abs/10.1145/3376930.3376978.

Zhou:2018:VMM

[ZYLY18] Zhou Zhou, Junyang Yu, Fangmin Li, and Fei Yang. Virtual machine migration algorithm for energy efficiency optimization in cloud computing. Concurrency and Computation: Practice and Experience, 31(7):e4942:1–e4942:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zytaruk:1994:WVMa

[Zyt94a] Kelly Zytaruk. The Windows 3.1 virtual machine control block. Part 1. Dr. Dobb’s Journal of Software Tools, 19(1): 115–??, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Zytaruk:1994:WVMb

[Zyt94b] Kelly Zytaruk. The Windows 3.1 virtual machine control block. Part 2. Dr. Dobb’s Journal of Software Tools, 19(2): 107–??, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Zhan:2018:HPV

[ZYZ^{+}18] Dongyang Zhan, Lin Ye, Hongli Zhang, Binxing Fang, Huhua Li, Yang Liu, Xiaojiang Du, and Mohsen Guizani. A high-performance virtual machine filesystem monitor in cloud-assisted cognitive IoT. Future Generation Computer Systems, 88(??):209–219, November 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X18304394.

Zhao:2006:DFS

[ZZF06] Ming Zhao, Jian Zhang, and Renato J. Figueiredo. Distributed file system virtualization techniques supporting ondemand virtual machine environments for grid computing.

Cluster Computing, 9(1):45–56, January 2006. ISSN 1386-7857.

Zeng:2023:EES

[ZZG^{+}23] Deze Zeng, Andong Zhu, Lin Gu, Peng Li, Quan Chen, and Minyi Guo. Enabling efficient spatio-temporal GPU sharing for network function virtualization. IEEE Transactions on Computers, 72(10):2963–2977, October 2023. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2021:KSV

[ZZW^{+}21] Yiming Zhang, Chengfei Zhang, Yaozheng Wang, Kai Yu, Guangtao Xue, and Jon Crowcroft. KylinX: Simplified virtualization architecture for specialized virtual appliances with strong isolation. ACM Transactions on Computer Systems, 37(1–4):2:1–2:27, March 2021. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic). URL https://dl.

acm.org/doi/10.1145/3436512.

Zhou:2022:SRE

[ZZW22] Hang Zhou, Xinying Zhu, and Jian Wang. A specific risk evaluation system for live virtual machine migration based on the uncertain theory. Scientific Programming, 2022(1):6784419:1–6784419:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.wiley.com/doi/epdf/10.
1155/2021/6784419.

Zheng:2024:MBA

[ZZW^{+}24] Luxin Zheng, Jian Zhang, Xiangyi Wang, Faxin Lin, and Zheng Meng. Multimodal-based abnormal behavior detection method in virtualization environment. Computers & Security, 143(??):??, August 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002104.